A Complete Bibliography of Publications in the Journal Family Statistics

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
            WWW URL: https://www.math.utah.edu/~beebe/

17 October 2023 Version 1.31

Title word cross-reference

(0, λ) [Goe75]. 0 [Bär76, Dac11, HDA99, PW76]. 1 [Bär76, PW76]. 2 [Tou75, TW78, UR81]. 2^{k} [Bag90, MK83]. 2 × 1 [Ano72b]. 2 × 2 [PGT08]. ^{1} [Bel83b, Mam95]. _{n} [Lie93]. _{s} [AM90, NR81]. A [PMA17, Von87, Wie88]. A_{2} [CGK17]. α [CD09, JK92, JSW22, Wan15, WYLW22, WS22, WWLH19, XWY^{+}20]. AR(1) [KW15]. AR(2) [And90, AG91]. Arma(m, ∞) [TVL96]. X̄ [AvC89]. BNBU(t_{0}, t_{0}) [TJZ89b]. c [KB00, Puk81b, Sin91]. χ^{2} [BKN70, Jag80, Mat81b, Mat81a, BKN70, GLVV98]. D [Bis95b, Det93, Det95a, DM06, HM89, KHN15, MOS86, Son82, VYD83, MK83, YW12]. E [Bag90, OM91]. E(Y) [GAB85]. E_{k} [KK80, Kre80]. ϵ [Bed81]. F [ASSH13, Ger86, Güv06, Hav76, Jag80, Nad08a, PGT02, Sin05, GR17]. F^{α} [BS13, Enn98]. G [CL12, Ger86, KC17, WYLW22]. Γ [JRM04, Laz16]. ∞ [Dac11, Fre74, Kle77b, Kre80]. K [ASSH13, Ber98, Col85a, Enn98, Hei18, ZW10, AMB11, AF14, BZ15, Gor19, HT16, Jas18, JS22b, KR17, Oko17, RM17, Rie87, SS21, VP04, Wie86, YW12].

L [BCN15, Cig13, Col03, GR06, Oko16, Ryc93b, Ryc93a, Ryc04]. L^{2} [Pet73]. L_{1} [BDG95, TW05, Yat19, Cad01]. L_{2} [Abd99, BL96a, QCO13, Yat19, ZX14, Wil08]. L_{γ} [Ars16]. L_{p} [Moj07, Pad96]. L_{q} [AC23]. L_{r} [YFZ11]. λ [DLO22, Hus88]. λ^{s} [KS86]. M [CD09, DZX18, FRLLT18, JS82, Jur80, RZ11, SL95, TC13, WYLW22, Zha97, ZZW23, Aly87, BM01, BS21, CMLBSM14, CHP03, DS91, Flo93, Kum89, Sar82, Sin81]. R [GFCO04]. MINQE(U, I) [Wen95]. MM [TC13]. n [Hab89, Han83, MK83, Ped86, Sty75, WY04, WBL06]. n − k [WBL06]. n[diff between]1or2 mod 4 [MK83]. N ≡ 0mod3 [PCKMK17]. P [Son20, AX16, Bic23, FYY^{+}23, FL10, WW22a]. P (X > Y) [SM08, Sen11]. P (Y > X) [CS10, PAW05]. p ≫ n [LJ22]. ϕ [GNP08, Tak09, WYW21]. ϕ_{p} [RDLF03]. P r(a′x > b′y) [GG90, RF94]. P {X > Y } [SZ97]. R [AK03, COQ21, HW88, LMV17, Röd87b, Röd04, DST21, EBTB21, Ped86, RY86]. R = P [X < Y] [BKT14]. R^{1} [Sie74]. R^{3} [Kos80]. R^{d} [Yat19]. R^{n} [KZ86]. ρ∗ [LLWW20]. S [Ciu12, GVV17, TC13]. s(x) [Ano73b]. σ [Páz91a]. T [Gen13, GN98, JS97, Zie95, AG09, CY12, Gup03, Kim10, KM20, MCML15, MCCL18, NK04, ND05, PN13, Ras85, RLAV16, Sch75a, YA12]. t(GT) [Nad08b]. T^{2} [GK99, KS16]. τ [Amm20]. θ [Zin81]. × [Led84]. U [Blo07, BW14, Cup21, DRW17, KAM17, MSW03, Ver92, YYVH16]. VARMA(p, q) [VN19]. φ [Man22a, Pel11]. X [KNS72]. X_{1} − X_{3} [SW85]. X_{2} − X_{3} [SW85]. XY [NK05a, Nad06]. Z [Bod75].

-Anpassungstest [BKN70]. -associate [Sin81]. -ball [Ped86]. -category [SS21]. -charts [AvC89]. -contents [Ped86]. -Criterion [Bis95b]. -Dependent [CHP03, Aly87, Sar82]. -designs [Sin91, KAM17]. -dimensional [DS91, WY04]. -discrepancy [QCO13]. -distribution [Gup03, Kim10, YA12]. -distributions [ND05]. -Divergence [GLVV98, Man22a, Tak09]. -divergences [GNP08]. -efficiency [CL12]. -empirical [Cup21]. -Error [BDG95, Yat19]. -estimates [Ryc93b, Ryc93a, SL95, HW88, Oko16]. -estimation [Kum89, AK03, DZX18, Röd87b, Röd04, RZ11, TW05, ZZW23]. -estimator [CD09, Ciu12, JS82]. -estimators [BM01, Flo93, RY86, Col03, Jur80, LMV17, TC13]. -families [Pel11]. -family [Hus88]. -function [ZW10]. -Functions [Ver92, Bod75, Hei18]. -generalized [CMLBSM14]. -median [Cad01]. -Metric [Pad96]. -minimax [JRM04, Laz16]. -MINQUE [Kle77b]. -mixing [CD09, LLWW20, Wan15, WYLW22, WS22, WWLH19, WYW21, XWY^{+}20]. -Model [JS97, ASSH13]. -models [RLAV16]. -moments [BCN15]. -nearest [HT16]. -nonnegative [GVV17]. -norm [AC23, YFZ11]. -Optimal [Det95a, MK83, Puk81b, Det93, Bag90, CGK17, COQ21, DM06, HM89, KHN15, PMA17, Von87, VYD83]. -Optimality [OM91, MOS86, NR81, Son82, Wie88]. -Optimierungsproblem [PW76]. -Optimierungsprobleme [Bär76]. -parameter [BZ15]. -periodogram [FRLLT18]. -phase [Tou75]. -projections [GR17]. -proportional [Wie86].

-record [AMB11, Ber98]. -records [AF14, JS22b, Enn98]. -Sample [KB00, Rie87]. -Scheme [Enn98, BS13]. -silent-vs.-noisy [Sty75]. -size [EBTB21]. -squares [ASSH13]. -Stable [JK92]. -stage-least-squares [TW78]. -statistic [DRW17, Güv06, MSW03, YYVH16]. -statistics [Ryc04, Blo07, BW14, Cig13, GR06, Moj07]. -Steuerungen [Gus71]. -test [KNS72, Ras85, Zha97, Zie95, KNS72]. -Tests [Sch75a]. -th [DLO22, DST21, Gor19, Han83, Jas18, KR17, RM17]. -tuple [Oko17]. -type [Dac11]. -unit [UR81]. -upper [Yat19]. -validity [Bed81]. -value [AX16, FYY^{+}23]. -values [Bic23]. -Variate [GN98]. -Wiener [JSW22].

/GI/1/ [Kre80]. /GX/1 [KK80].

1 [Bel91, Ber74, Dew75, HGLS82a, KK80, KS76b, Nig90, RL98, RS74, Ros71, SD86]. 100 [FG75]. 10th [HW74]. 1982 [War85b].

2 [DF15, Fle76, HGLS82b, TW78]. 2-inverses [DF15]. 20th [Gir74].

3 [HGLS83a].

4 [HGLS83d]. 40 [GSB11]. 48 [SGMFB18].

5 [HGLS83b]. 54 [AE21]. 56 [Nku23]. 5th [HW74].

6 [HGLS83c]. 6th [BM71].

a.s [Rüs86]. a.s. [KT83]. Aalen [SC18, sSW19, ZR10]. Abbruch [Fäh78, Fäh81]. Abhängige [Wei71]. abhängigen [Mec72]. Abhängigkeit [Tam74b]. Abhängigkeiten [JK76, KS76a]. Abhängigkeitsmaße [Röd70]. Abhängigkeitsmaßen [Hös76]. Ablaufpläne [Näg75]. Ablaufplanung [Hof72]. Abschätzung [Düc75b, Wag83, Düc75a]. abscissa [Ban70a]. absolute [LKM14]. absolutely [HE12]. abstract [Bit75]. Abszissenwerte [Ban70a]. Abweichungen [Wol84]. Accelerated [Ach95, BN97, BN00a, LNBR02, AFSS22, CS06, CT03, RALN93, TDY08, WHXC16, WG21]. accelerating [HAMAZ96]. Acceptance [BN00b, CF21, CF22, Cha15a, Eva75, FE85]. access [Hös76]. accessible [SH17]. accidental [Gla76]. according [Har75a, Har75b, Nät75a]. accounting [Dew73]. accuracy [Sty75]. Accurate [WN14a]. ACE [Röd04]. activation [LCY15]. Active [KM17]. activity [Bak76, FN77]. actual [Möh87b]. Adaptation [PVW74, Lia12]. Adaptive [AK03, BJ14, BC02, BC99, CCD15, Ciu16, Ciu20, Gun88b, Gun88a, Hus88, LY20, LGL^{+}17, PS10, BB07, BBM12, BD13, Bun85, Dro93, GMP16, GLZW12, Gün77, Gün83, KFPG21, Sch87f, SKPK21, YC15, ZH16, Gün80]. adaptively [Art19, Gri80]. adaptiven [Gri80, Gün77]. Addendum [Det95a, MB92]. additional [Bun75b, Hen79]. Additive [Sch96, SLS99, BS07a, BS10, EAA16, HCW16, ML21, NB16, PMA17, TT10a, ZWL14, ZL16]. additive-multiplicative [EAA16]. adequacy [EM87]. Adjacent [FR95]. adjoint [Ros91]. adjusted [BD13, CKM09, CM13, PDG13, SH09]. Adjusting [WN12a]. adjustment [Sti78, WT20]. Admissibility [GM99, Hof77a, Hof85, Rod88, DK84, Lät74, Led84, Rei82, Sha83, Tan11, Yu04]. Admissible [AN98, GH89, GM94, Hof80, Led78a, Led78b, Als93, CXH14, Hol75, Hol76, KZ87, LY93, PH16a]. Advances [Del09]. aeronautic [FGK19]. Affine [AS98, LTV03]. Affine-invariant [LTV03]. after [ST86]. Against [BM02, GD96, BMK22, Ger87, HE12, JE90, Kat89, Kla03, KB00, Led84, RG84, Ryc93a]. age [Bei81]. ageing [AN13, FPLS12]. aggregate [TC22]. agreement [GVGP13]. agricultural [TN08]. Ahmed [Tik09]. AIC [TC13]. aircraft [BNS73]. Akaiee [Tom90]. Akaike [BM94]. Aktivitätskoeffizienten [FN77]. Algebraic [Cen78, And78b]. algorithm [ABD93, Bor73, GPS14, Hav88, HT71, HW71, JBC22, Kér72, LPS90, PW76, Pol08, Ric76a, Sch70c]. Algorithmen [Ban70b, Ter72]. algorithms [AP07, Ban70b, Dau82, GBS20, GB23, JLM23, Pet86, Win17, Ter72]. Algorithmus [HT71, HW71, PW76, Ric76a, Bor73]. aliasing [Els17]. aller [Pie72b]. allgemeinen [Bel79a, Ehl75, Ehl76, Sch75a, Tou70b]. allgemeiner [Sch75d, Sch75e]. allgemeines [Hol75, Näg75]. allied [MOS86]. allocation [BB07, HDC20, Jen73, KCNB11, Lei70, Rui85]. allocations [Ruk07]. allow [NCJ23]. allowable [Pré74a]. Almost [AAB18, JC07, Páz90a, Páz90b, SK20, Col85a, DM22, Sah87, Sch89b, WY13, Zan14, Zan15]. alpha [MFBG14a, MFBG14b, SGMFB18, SDJ16, Zha23]. alpha-power [MFBG14a, MFBG14b, SGMFB18]. alternative [BMK16, FG75, Haf75, MKM21, Mur15, Tro91]. Alternativen [Sch75a]. Alternatives [GD96, KB00, TJZ89b, HE12, Hof77b, JY20, JE90, Kat89, Koz78, MM18, Pin75, RG84, Sch75a]. Alternativsatz [FG75]. ambiguity [Páz84]. amended [AAA13]. among [KR78, RPZB18, TL15]. amount [Deu72]. Amplitude [NKI04, NK06]. analog [Deu72]. analogous [CV78]. Analogrechners [Deu72]. Analyse [Atk75]. Analysis [Ant84a, Atk75, BFF00, BN97, BH18, BPK14, Che97, DY02, Fou97, KKG17, Neu02, PJ17, PEB15, RL98, sSLMJ17, Tur90, Aka82, AAA13, AFSS22, And78a, AG91, AINJ14, AS04, BHM04, BCFF88, Bun73b, BF85, Col85a, Dau82, DDT88, Del09, Den88, DMP07, Döh81, DB13, Dry72, DH77, Dry88, FRLLT18, FW13, FY03, FG06, FR84, FP92, FT79, FK90, GVMS09, HS79, HOCC13, Hav88, Hen89, ID03, JR06, Juh89, Kim07, KRC19, KK14, KJKR76, Lau85b, Lau85a, Lee08, LC17, LL18, MB18, Mat75, May11, MP22, MV10, PGT08, PKM78, Puk77, RAM^{+}17, Ras88, RALN93, Rod93, RM16, Sch89a, SKCE19, Sti78, SH86, Sto88b, Sto93, Tit83, Tit90, TB01, TT22, UR81, Van88, XW10, ZT15, Hof84]. analytic [KHN15, Koo86]. Analytical [Cse95, HKMM98, Pad89]. Analyzing [KM21, NKI04]. Andrew [Tik09]. angular [Mic09]. anisotropic [CCM18]. Anisotropy [Sto91]. Ankündigung [KJM^{+}72]. Anmerkung [Sch76a]. Announcement [KJM^{+}72]. ANOVA [Ano19, Wie88, Are12, LaM17, MNvRS10, WS13]. anova-settings [Wie88]. Anpassungstest [Aui90, BKN70]. Anscombe [Sie87]. Antagonistic [Ano73d]. antagonisticheskimi [Ano73d]. antoregressive [Gun88a]. Anwendung [Dew73, Ehl75, Ehl76, Gla76, Gru71, Lie72, Ric84, Sch75d, Sch75e, Tam76, vW72a, vW72b, Sch73, Wit76a]. Anwendungen [Gir74]. Anwendungsbeispielen [Bel79a]. applicability [Gai23, SV10]. Application [Bed81, KN98, KFL97, NOM13, ABP18, Ahm93, Bel79a, Ber79, CLDG16, CB04, CRCN14, DLMR13, Dew73, DL16, Ehl75, Ehl76, FGK19, FN13, Fur15, Ger79, Gla76, GN17, Gru71, KK11, KMSS10, LPS90, Lie72, LH22, NSG07, Nad08a, Nät75b, Nät79, Ouy92, PYHT01, Pon93, Pré73, RD21, Ric84, SV89, Sch82a, Sch74b, Sch75d, Sch75e, SY22, Sha86, Tho78, TZ91, WYW21, WW22b, vW72a, vW72b, BT04, FP92, GCB94, HDA99, Sch73, Wit76a]. Applications [BFF00, EMG98, FZF97, GK95, LPC97, NLP95, PGT02, AL18, AG09, AE21, BLSC09, BGTV06, BBK14, Ber75, Böh83, Bon82, Bos88, Cap73, CHA15b, CZZ19, Cup21, DWHH19, DR15, DF15, FGL17a, Ger78a, GR17, Gir74, GJLT15, Gup79b, Gup82, Hal91, HCW16, JS22b, KT12, KD16, Kim12, KK15, Kim16, KK14, KT83, LLWW20, MA20, Nad03, NK07, NS10a, Nog20, Oui22, PP83, PS14, PMO14, Pes13, PGT05, PG09, QLT17, STS85, SDJ16, Sol19, TH04, VFS^{+}20, Wan07, WDZH14, WWH18, WS22, Win17, WHVW23, ZZWW23, dQLS16, Cap73]. applied [Ano75b, LD21, Mai91]. apply [Rév77]. applying [Tam76]. Approach [JO95, Mar97, Tro00, AS04, AM90, Bel83c, Bic18, Böh81a, Böt20, BG79, Cap19, DH77, Dry84, FK12, Hös74, KD16, Kob87, KY22, Läu75a, LH15, LM23, MM22, MD22, MBF^{+}20, MV10, NW12, Páz90c, PSB23, PP94, SlP17, SL17, Sär95, ST83, SAEAEA13, TT22, Yoo18, YYVH16, Zha12, Zie88b]. approaches [Ant84b, FM89, GVGP13, Páz80a, Tit80]. Approximate [DP17, FZ89, HW07, Kha12, Von87, CM15a, Gir71, GM78, GKM18, HM88, Huz18, MO73]. approximately [SY81]. Approximating [GG15, MG96, BG74b, PHS09]. Approximation [Bit72, Bun73a, ESSB93, Kha04, Khu95, RS74, Sch90b, Tuy76, AL82, Ber17, Blo07, BZ13, BS80, Dup77, FN13, HK09, Has77, Her81, Jai74, KvR00, Lip75b, Lip76, LH91, PH16b, Rei84, Rév77, Roy87, Sch76b, She14, Sto76a, Tau84, TW83, TH08, Tuy74, Wis77, Sch76b]. Approximations [CG95, Enn98, Sto76b, vR97a, Bos88, Cig13, Moj07, Oui22, Pol86, Sch79b, UI13]. Approximationsmethoden [Lip75b, Lip76]. approximative [Bro93, Zwa80]. approximierenden [BG74b]. April [Atk75, NS10b]. Apriori [JN75]. Apriori-Verteilung [JN75]. Arbeit [Gus71, Sch76a]. Arbeitskräfteentwicklung [Tit70]. Arbeitstagung [BK71, BM71, Els75, HW70b, HW70a, HW72, HW74]. arbitrary [AHPS84, BB04, CM15a, Phe89]. ARCH [DMR07, GML23]. ARCH-type [DMR07]. Archimedean [BKHB22, FZLB22]. arcsin [CL12]. arcsine [CL14]. area [BM09, HEHM13, NW12]. argument [Jag80]. ARIMA [FL03, NDDDG16]. arising [BN14]. ARMA [AK03, And81, Cip81, Cip84, DdMT20, Wei09]. Armed [Cha96, Jon76b, Rei82, Rei86]. Arp [Sto93]. arrangement [LF22]. Arrangements [Pla94, EP17]. arrays [EP17, WDZH14]. arrival [NK88]. arrivals [Cha74, FS81, Goe76, Mad76]. Art [Kin75, Sto93]. Aspects [IRB14, FN77, Kum85, Kum89, Lau85b, PN10, Puk81a]. Aspekte [FN77]. assess [Pol08]. assessment [BKS23]. assignment [Ter72, Ull76b]. associate [Sin81]. Associated [Fer99, BB22a, BNS15, ÇO17, CBP23, LMS06, LMV17, RA06]. association [JGAFBCMG14, Juh89, RD21, VT08]. assumption [BGV06, Böh81b]. assumptions [HS15, KLF22, Net90, Rud76]. Asymmetric [Dör15, LMS92, AC23, CAK13, Col91, HB09, KP09, Kie17, Nie06, QCO13, She88, ZZWG09]. asymmetry [Bow21]. asympotically [Arn91]. Asymptotic [BFF00, BCMR90, BN14, BPvZ04, BP16, BBKP19, Bed85a, BB91, BDG95, BGAL01, BS21, Bil97, Bla14, BS80, BDF^{+}18, Car19, Che97, Che12, Cie98, Cup21, Dem14, DJ17, DGP91, Dro06, Fah87, Faz14, FP92, Fin94, Flo93, GLZW12, GM92, GJML00, GJLT15, GT13, HK00, HSW85, HW88, Hei86, Hei88, Hei18, Hos88a, HB09, IDB12, JSW22, Jur80, KB85, KB00, KZ13a, KZ13b, Kun97, Lai00, Lin09, Mie79, MO03, Neu79, Neu92, Nus77a, Nus77b, Nus79, Pho87, Pon84, Rao83, Rei82, Sai86, Sch75b, Sch75c, Sch79a, Sen82, SX20, TVL96, Ver92, XWY^{+}20, XWJ14, XG07, YFZ11, ZFA20, ZW10, BG19b, Bed85b, Bel80, Bel83b, BNS15, BS23, BW81, BRZ93, BCP99, Col91, Del09, DB10, DdMT20, EJ93, Fin87, Fra82, GW04, Hab89, HW14, HR93, IZ83, JC07]. asymptotic [Ker84a, Ker84b, KHN16, Kle80, Lem13, Lie93, LMV17, Mal89, MMdP09, Mas19, Mie91, Müh84, MB21, NFF^{+}23, PMO14, PW15, Ras85, Röd71, Sch75a, Sch76c, ST81b, Sch77b, She14, Tak09, TDM13, VS81, WY04, Wes14, Wit78b, Wit81, YLY18, Zwa83, Zwa85, Sak71, Zäh74]. Asymptotical [BN97, Bel91]. Asymptotically [BL96a, DM09, GKP96, Mil86, Ryc87, Wit78a]. Asymptotics [BC02, Höp98, KS96, Sch77a, SL11, HBCL14, Nie17, OJM15]. asymptotisch [Wit78a]. asymptotische [Röd71, Sch75a, Sch77b, Wit78b, Sak71, Zäh74]. asymptotischen [Wit81]. asynchronicity [AG23]. attainment [Tar88]. attraction [SZ17]. attributable [GW04]. attribute [FE85, LSS12a, LUK13, LUK14, SKV18]. attributes [Krz05]. Aufgabe [vW72a, vW72b]. Aufgaben [Hol75, Sch76b, vW74]. Aufteilungsarten [Deu72]. Aufteilungsprozeß [Bor73]. augmentation [Bak84]. augmenting [KJ14]. Ausbeutung [Ros71]. Ausgangsverteilung [SW85]. ausgewählte [KJKR76]. Auswertung [Hen89, Sch87a, Sch89a]. auto [HC11, MB18]. auto-models [HC11]. auto-odds [MB18]. autocorrelation [Sai86]. Autocorrelations [And76b, And79b, And79c, HB09]. autocovariances [Bos88, TSO07]. Automatic [AC23, Wer78]. automatische [Wer78]. autoregression [Col85a, PJ17, RZ11, TMN00]. autoregressions [TY12]. Autoregressive [AT08, And76a, FR98, Mou02, Sch89c, Tur90, Vie95, And79c, BH18, BR90, Bla14, CLS14, GPS14, GMCT18, Gun88b, KK14, LKK07, LLC16, LLZ19, Mal89, Mie79, MSW13, SS15a, Sch91b, WLW16, YFZ11]. autoregressive-moving [And79c]. autorsgressive [Car84]. auxiliary [CM90, LSS12b, MMRA07]. Availability [Cse95, BM12, GW76, Ger78a]. Available [PP92, Cha74]. Average [AMSV12, And76b, Ger89, And78b, And78c, And79c, EJ93, GPS14, GMCT18, MCNO19, Mie79, Pie75a, Pie76, TJZ89a, WQO22, ZQC23]. averaged [GBS20, GB23]. averages [Ara93]. averaging [Agr82, CLLY22, DZX18, ZZZ17]. axiomatic [Hös76]. axiomatischer [Hös76].

B [Hec86, TG87]. backgrounds [Juh89]. Bahadur [CD09, KS96, WYW21]. balance [HDC20]. Balanced [Das96, HT16, Khu95, MU99, SJK97, AHPS84, CXH14, CGK17, Joz14, Kha89, Kle75b, OM11, PH16a, SK92, Sei78, Sei79]. ball [Ped86]. ballot [SS81]. ban [PS84]. Banach [AK75, BA89, Cad01, Sie87, Trö75]. Banach-Räumen [Trö75]. Bandemer [NS10b, Böh81b]. bandit [Jon76b, Rei82, Rei86]. Bandits [Cha96]. bands [BB72, Ker16, SDC19]. Bandwidth [HK87, LMP10, MP96]. Bandwidths [Jon95, Cao01]. bang [Gus71]. bang-bang [Gus71]. bank [KS83a, Kra76, Rod89]. barrier [EG74]. Barriere [EG74]. Barriere- [EG74]. Based [AJ97, BGP95, Bel91, Ber98, BT97, BC99, CB02b, GM94, GJML00, GC95, GM97, MJA02, Neu02, RO98, Ste95, TP97, VV00, AAFM07, AA12, ASSH13, AJM12, AMB11, AF14, AAR18, AHJN01, Ara93, AA11, AA13, BC06, BZ14, BC23, BDP16, BP16, BBKP19, BG19b, BMMP16, Bei81, BV11, BS21, BP93, Blo07, BW14, BCDD11, BRF20, BE12, CC01, CY12, CKS22, CM05, CK06, Dab81, DHN20, DLO22, DB13, ELV22, EBTB21, Els22, FYY^{+}23, Fer07b, FG06, Fis09, GB17, Gla90, GR06, Gor20, GF09, GT13, GNP08, Hei18, HM02, HY10, HS15, HHM15, IM14, Jas18, JGMO20, Juh89, KYR89a, Kat89, KK14, LLBG10, LSS12b, LF22, LQ12, LBSM00, LM23, MR73b, Man22a, MCCL18, Möh87b, Mou01, Mou03, MAS06, MB21, NRD08, NB16]. based [OJM15, PT83, RA01, RAD07, Ras06, RN18, Ryc04, STY04, SK14a, SL17, SY22, SS04, SA23, SdSSL15, SB17, Sto79, SSN17, Tab13, TE15, TZ13, TF21, TC13, VS81, VP04, VK20, WZXH15, WS22, WZW21, WD22, Wu16, XW10, Yoo18, ZK16, ZH16]. basic [HBN86, Näg75]. Basis [SL17, Hof79b]. Basu [JGMO20]. batch [YS23]. batches [Mad76]. Bates [KP74a, KP74b, Rei86, Tar88]. bathtub [BN16]. bathtub-shaped [BN16]. Battese [Güv14]. Bayes [HK78, Pfu78, Bic23, Bun77b, BG79, Bun05, BJ05, CM90, GH89, GP82a, Hof97, HL13, IR91, JR12, Jon76a, Kle75a, Las91, Mag95, Mar97, MA86, Nät84b, Pho87, Pil81, SP08, SW89b, WS13, ZRAN17]. Bayes-Lösungen [HK78]. Bayes-Verfahren [Pfu78]. Bayesian [AAFM07, AA12, AALA20, Ahm11, AJM12, AMB11, AO07, AHJN01, AL18, AJ97, AAA13, AG91, Ant84b, AINJ14, AN98, ASM83, AEW94, BSS18, Bel74, Bel83c, BT97, Bic18, BG74a, CKM09, CM15a, Det93, Det95a, DT16, Eva75, EJ81, FB22, FE85, GZ23, GVMS09, GVGP13, GM97, HK78, HZ21, IF03, KM13, KP09, KCL16, KRC19, Lee08, MJA02, Mou03, MB21, Nie06, NAHJ03, PMO14, Pfu78, PGT08, Pho87, RAM^{+}17, RLAV16, RALN93, Rod93, RL98, RV08, Rud76, SGG04, Sch77b, SK05, TH04, TB01, TT22, Tur90, Tzi87, YLWL22]. Bayessche [Bel74]. Bayesschen [Sch77b]. BDS [Lai00]. be [PJ05]. Beale [Gud70]. Becker [ASSH13]. Bedeutung [RS74, Sie74]. Bedienungs [KJKR76, Sto71, Sto72b]. Bedienungs- [KJKR76, Sto71, Sto72b]. Bedienungsbeginn [Sie74]. Bedienungsmodelle [KS76a]. Bedienungsprozesses [Fle71]. Bedienungssystem [Fre74]. Bedienungssysteme [Sie74, Zäh74]. Bedienungstheorie [Sto72a, Sto73]. Bedienungszeiten [Mec72]. Bedienungszeitverteilung [Fre74]. before [sS14]. Beginn [RS74]. Behandlung [Hof72, Ric76b, Gud70]. Behavior [Kre85, Gus71, Sch75a, Sch77b]. Behaviour [MP96, BNS15, DJ17, DB92, LLBM^{+}11, Lie93, Mat89, Ras85, RM17, SK20, Sin84, Zuo01]. behaviours [JSW22]. beidseitig [Sch89a]. beim [BKN70, EW75b, Fäh78, Fäh81, Foc71, Foc84, Lie72, RS74, Ros71]. Beitrag [Sti78]. bekannten [Ros71]. bekannter [Zin81]. beliebiger [Bär76]. Bellman [Bor73]. Bellman-Algorithmus [Bor73]. bemerkenswerte [Pie73]. Bemerkung [Hoy79]. Bemerkungen [Gus71, Pie75a, Bel79a, Hol73a, Neh75, Ron76]. beobachtbarer [Fäh81]. Berechnung [Ric84]. Bericht [BK71, BM71, ES74, Els75, HW70b, HW70a, HW72, Kle72, Man72, Mön72, PKN72]. Berichte [ES71, Gir74, HW74, Lom74]. Bernstein [BPR00, Bos93, KRC19, Pho87]. Bernstein-type [Bos93]. Berry [Alb84, FGL17a, LZW23, MS80a, PV98, Ros89, WYLW22, WWLH19, WHVW23]. Berücksichtigung [Düc75a]. Beschäftigungsperiode [RS74]. beschränkten [Hol75]. beschränkter [Zäh74]. beskoalitsionnikh [Ano75a]. Bessel [NSG07]. Best [AHPS84, BK79a, Dry77, Fal95, Kle75b, Liu11, MK95, VV96, WHG87, ZZ05, BB22a, DH81, Gla76, KP74a, KP74b, Mat83, Wie88]. beste [Gla76]. Bestimmung [Deu72, FN77, Gir71, Gri80, HK78, Pie72b, Lom70]. Beta [CRCN14, AFL08, BG19a, BN14, Bar87, CCMdC13, CSO13, Gup05, Iga20, JG85, Koc74, NK05b, NK07, NSN17, Ped86, PG09, RPZB18, SDJ16, SYM16, TSG20]. beta-generated [BN14, RPZB18]. beta-Pareto [AFL08]. beta-Weibull [CSO13]. Betaverteilung [Koc74]. Betrachtungen [Pie75b]. Better [HHJ95, CFR89, Mej78]. Between [AN98, Dre98, Bak84, BT16, BP93, BVP06, DK12, Ger73, GVGP13, Hös74, HTHF16, KSS76, KR07, LST18, Man22b, MB91, MB92, Nät75b, Nät79, PJ05, SP98a, Tor17, Ull76b]. Beweis [Lie72, Nož71]. Bewertung [Hof79b]. bezüglich [Foc75b]. Biadditive [HDA99]. Bias [Col03, GCF04, KS16, MH97, Rie89b, Ryc93a, Sch90a, SL23, And79b, Bie15, Blo03b, DB92, EPT20, GS14, JSV07, NTJ09, Rie89a, Ruk07, Ryc87, Zha13, Zie85]. Bias-robust [Col03, Zie85]. Bias-Robustness [Rie89b, Ryc93a, Rie89a]. bias-stable [Ryc87]. Biased [Kut12, RO98, AFSS22, EBTB21, GS88, ÖA16, SS20a]. biasedness [BL78]. biases [Col91]. bibliography [Col85b, Jiá81, JA89]. bidimensional [RDLF03]. BIFRA [TJZ89b]. bifurcating [Bla14]. big [PJ05]. biigrisha [Ano76]. Bildung [Wer78]. Bilinear [GJML00, BG16, DMR07, Wit89b]. bimodal [AE21, KYR89a, KYR89b, VFS^{+}20]. Binary [AL00, BL06, KHN15, RS09, Wei09]. Binomial [FC89a, VW96, WK13, AWJF22, CB02a, CZZ19, DDD19, Dup17, GML23, Jai74, KW15, LLZ19, LCY15, PPD07, Sha86]. binormal [Szc79]. bins [BC06]. bioequivalence [CWW09]. biology [BGTV06]. biplots [KFH16]. Birnbaum [HOCC14, BZ14, FZB17, Gen13, Kun15, MFBG14a, RVB18, RVBZ20]. Birnbaum-Saunders [RVBZ20]. birth [Fra82, Goe75, WD22, Küc82b]. Birth- [Küc82b]. birth-and-death [Fra82]. Bivariate [Ach95, BSL13, BvET16, CTL14, Han96, Kun15, NK17, Röd77, Tia91, TJZ89b, AAB12, BM12, BNH20, BCN15, CB04, CF78, DSG91, Dom10, EBQY17, FZB17, FV06, GJV16, Gup05, HY10, JK13, JLO12, Kow77, KM21, LTV03, LF22, LBSW19, LGL^{+}17, Mic09, NK04, NK05b, Nad06, Nad08a, Nad10, NSN17, NS03, NS10a, NN23b, NFP18, RPZB18, Röd87b, SK14b, SKCE19, TDM13, WL05, WG21, WLT^{+}22, Xek86]. bivarite [NN91]. blind [BNS73]. Blinded [Gov11]. blinden [BNS73]. Block [GM91, MU99, ASSH13, Düc75a, Gno76, Guo20, HSW85, HPN22, HK15, Jac91, Jas81, KM82, Mej78, Nie17, QXY23, RFK18, SX20, TT10a]. blocked [Jac13]. blocks [ASSH13, DR92, EP17]. Blockstrukturen [Düc75a]. blyzhdanija [Ano81e]. Board [Ano77a, Ano77b, Ano77c, Ano78a, Ano78b, Ano78c, Ano78d, Ano79a, Ano79b, Ano79c, Ano79d, Ano80a, Ano80b, Ano80c, Ano81a, Ano81b, Ano81c, Ano81d, Ano82a, Ano82b, Ano82c, Ano82d, Ano83g, Ano84r, Ano84s, Ano84t, Ano84u, Ano85v, Ano85w, Ano85x, Ano85y, Ano86q, Ano86r, Ano86s, Ano86t, Ano87f, Ano87g, Ano87h, Ano87i, Ano88d, Ano92b, Ano92c, Ano92d, Ano93d, Ano93e, Ano93f, Ano93g, Ano93h, Ano94a, Ano94b, Ano94c, Ano95a, Ano95b, Ano95c, Ano95d, Ano95e, Ano96a, Ano96c, Ano96d, Ano96e, Ano97a, Ano97b, Ano97c, Ano97d, Ano97e, Ano97g, Ano98b, Ano98c, Ano98d, Ano98f, Ano98g, Ano99a, Ano99b, Ano00b, Ano00c, Ano10, Ano11, Ano12a, Ano15b, Ano16b, Ano17, Ano83e, Ano83f, Ano92a, Ano96b, Ano97f, Ano98a, Ano98e]. board [Ano99c, Ano99d, Ano99e, Ano00a, Ano00d, Ano00e]. Bodies [Mol96]. Book [AEH^{+}86, And87, AA98, Ano83a, Ano83b, Ano83c, Ano83d, Ano84b, Ano84c, Ano84d, Ano84e, Ano84a, Ano84f, Ano84g, Ano84h, Ano84i, Ano84j, Ano84k, Ano84l, Ano84m, Ano84n, Ano84o, Ano84p, Ano84q, Ano85o, Ano85p, Ano85a, Ano85q, Ano85b, Ano85r, Ano85s, Ano85t, Ano85u, Ano85c, Ano85d, Ano85e, Ano85f, Ano85g, Ano85h, Ano85i, Ano85j, Ano85k, Ano85l, Ano85m, Ano85n, Ano86l, Ano86m, Ano86n, Ano86o, Ano86p, Ano86a, Ano86b, Ano86c, Ano86d, Ano86e, Ano86f, Ano86g, Ano86h, Ano86i, Ano86j, Ano86k, Ano87a, Ano87b, Ano87c, Ano87d, Ano87e, Ano88a, Ano88b, Ano88c, Ano89a, Ano89b, Ano89c, Ano89h, Ano89d, Ano89e, Ano89f, Ano89g, Ano90a, Ano90b, Ano91a, Ano91b, Ano91c, Ano93a, Ano93b, Ano93c, BvWT^{+}77, BRB^{+}77, BSF78]. Book [Bel83a, Bel84, Bel87, BS87, Bel88, Bel90, Bra88, Bra91, BSW^{+}81, Bun82a, BB82, Bun82b, Bun84, Bun87a, Bun87b, Bun88, Bun90, BKS^{+}91, CPGB98, Die90, DH90, Doh89, DHS82, Dro86, Dro88a, Dro88b, Dro88c, Dro88d, Dro88e, Dro89, Dro90a, Dro91a, DM91, Dro91b, DWP^{+}92, DRR^{+}94, DRR^{+}95, Dro98, Ege87, EHR^{+}80, Eng82, Eng83, Eng88, FRB77, Fis87, Fis88, Fis89, Fis90, Fle91, FSS^{+}94, Fre88, Gär83, Ger91, Gre88, Gui91, HPP^{+}88, Hec83, Hec86, Hei89, HGH^{+}79, HSM^{+}79, Hen82, Hen86a, Hen86b, Hen87a, Hen87b, Hen87c, HHN^{+}87, Hoe83, Hof84, HF88, Hof88, Jah87, JS89, JRR^{+}96, JSAG78, Kir83, KRF^{+}77, Kon89, Lie88, Lie90b, Lin87, Lis83, Lis88, Lom82, Mai91, MGLG78, MS82, Möh87a, Moh89, Mol97, MRR^{+}98, Nag97, NSH^{+}80, Nat89, Nät91, Neu99, Nol87, Nol88, Ohs91, Pil87, PNK78]. Book [Pin89, PHG^{+}79, PB82, Pla83, Pol87a, Pol87b, PSS^{+}94, RND^{+}92, Ric89, RRL^{+}93, RMGR88, Rie90, SKJ92, SFB^{+}82, SB82, Sch83a, SH83, Sch83b, Sch87c, Sch87d, Sch89d, Sch89e, Sch89f, Sch87e, Sch98, SS99, Sch99, Sch82d, SW89a, SSSS96, SCC^{+}00, SSS^{+}00, SSSS99, Sto83a, Sto88a, Sto89, Sto98, SEG^{+}85, Töp83, TBJ^{+}83, TDB^{+}88, VHW^{+}81, WW83, War85a, War85b, War86c, War86a, War86b, WW91, Wil87, WPH^{+}79, WHH^{+}80, WBG^{+}81, Win87, Win89, Win90, WHD^{+}90, Wit83a, Wit89a, Wit90, Wit91, WWW^{+}90, ZS90, BDP^{+}74, BWR^{+}71, BGB^{+}72, Bun74, BHW75, Bur70, DGB^{+}73, DGG^{+}75, DGH^{+}73, EW75a, EGV^{+}73, EvWE^{+}75, FGR^{+}76, GFFS74, HWK^{+}72, HL70, Hol70a, KS71, KPAF75, LSP^{+}73, MSW^{+}72, MSK^{+}76, MSS^{+}71, RWH^{+}72, SHP^{+}71, SND^{+}72, THN^{+}76, TSL^{+}76, VBP^{+}76, WHS^{+}76, WRME72]. Boolean [GCB94, Mol96]. Bootstrap [HHJ95, HDA99, Neu02, Pau12, Que86, Stu90, BR83, Cap19, DDT88, DDT89, DJ86, Guo20, HMP01, HJM90, KPH^{+}16, Nie17, PD91, YL17]. Bootstrapping [SZ17, HZ21, HS17b, Pes13, Shi15]. Border [BS07a]. Borel [Jai74]. both [Hor84, Sch89a, TL15, ZYZ21]. both-sides-censored [Sch89a]. Bound [Alb84, Ban70b, MS98, Pad96, PV98, CAK13, QCO13, Tar88, WYLW22, Zin81]. Bound-Algorithmen [Ban70b]. boundary [Ano81e]. Bounded [Koh12, MK95, BM01, GB23, KP09, Wan14]. Bounds [AHJ99, BJ85, Bis95b, Gor14, IR91, JB83, Kam89, Kre80, Miz96, Oko16, RF94, Ryc93b, VW96, ABP18, BB93, BR12, Bie06, Bie08, Bie15, Bie16, BS18, Dan03, FGL17a, GR06, Gor19, Gor20, KK80, KR17, LJ22, MS80a, MSW13, Nig90, Oko17, RR04, Ryc04, Ryc09, Sch85, Sto74, Sto83b, Wal88, Win17, WWLH19, WHVW23, Yat19]. Box [Tro91, Tro93, TP97]. Branch [Ban70b]. Branching [Mel94, EW75b, MMdP09]. Brasov [PKN72]. Breakdown [VN98, Zha97, Bor73, JBC22]. breaks [HS17a]. Bregman [BB22b]. bridge [Ars16, JSW22]. Brownian [KK11, KMM15, XZZ11]. Bucy [KP94]. bulk [Cha74]. Bunke [GNS15, Sch76a, Sch76a]. Burr [AJM12, AJ97, AEW94, Dom10, Mou01, NPS12, SP01, WWC^{+}10]. Burr-XII [WWC^{+}10]. busy [CT73].

C [Dro90a, FD72, Gus71, Led84, Ano72a, CT73]. C. [FD72, Gus71]. calculating [ANR07, Fer07a]. Calculation [Ric84]. calculations [JY20]. Calibrating [Tas21]. Calibration [KM95, GGL15, Hen80, Pes13, Sin13, ZX17]. called [Pré74a]. candidates [Gov92]. Canonical [SS98]. Capability [Che97, CHP03, Vän98b]. capacities [Bed85b, Bed86, Lei70]. capacity [Wen76b, Wen76a]. capture [SLL15, Zha12]. capture-recapture [SLL15, Zha12]. Carlo [Wag83, BH89, GJK^{+}16, LMS92, Wag83]. Case [Mel94, Ano74, BAO12, BK86, CGK17, DM22, Ege80, Fäh78, Fäh81, Fin94, GGS14a, GG90, Hab89, HAMAZ96, Han83, Jaw13, KZ13a, Mie91, PCKMK17, PR89, Pra86, Pré74a, PV85, Sha21, Tan11, TDY08, Web06, YT01, ZRAN17]. case-cohort [Sha21]. cases [Fer07b, Nad80, Pho87]. catalogue [GGS14b]. CATANOVA [LC10]. catch [Cey20]. Categorical [Pru93, CM90, GN17, HDC20, LC10, MB18]. categories [KI03, TIT08, TMY05]. category [SS21]. Cauchy [BN03]. causal [NB16, PW15]. cause [SKD21]. causes [LCY15]. Caution [Sto94]. cedlovuyu [Ano75c]. Cell [AAJ97]. cells [GC04]. Censored [AJ97, BGP95, CB02b, Ger89, LP97, ZM95, AJM12, BCKS01, BCK05, BCD11, BZ14, BPK14, BRF20, CBP23, Che21, Cup21, DK12, Geb84, Ger80, Ger87, GB17, HOCC13, Hen89, Hen92, JC07, JIP10, JJS06, KYR89b, KKG17, KLF22, Kum92, LdCIP18, LFGR11, MCML15, MCCL18, Mie91, Mou03, MAS06, PS14, RLAV16, SP01, ST23, Sch87a, Sch89a, Sch91a, SPTB18, SP08, sS14, SW16, SC18, sSW19, SAX23, TJZ89a, TPSR17, VK20, VDB13a, WG21, WLT^{+}22, WHT03, ZBF^{+}20, ZL17, dUÁV17]. Censoring [Bel91, AAFM07, Ahm11, AHJN01, BD13, BSS18, Bre05, Bur08, Çet21, CCD15, CBC12, DR03, GJV16, GS14, HW07, KCNB11, KFPG21, LC19, MK19, Nig90, PBK11, Pon04, SKPK21, TDY08, Ver06, VDB13b, WW21, WWC^{+}10, ZL17]. Censorship [CF98, Kal97, PRS95, Stu96, CCH87, Csö88, Dik01, GG91, Hen93, HLsS01, Hor84]. centers [EG74, GR72]. Central [DM22, EW99, KAM17, AAB18, Chr21, De 22, DJ17, DST21, Góm18, Gui86, HGLS82a, HGLS82b, HGLS83a, HGLS83d, HGLS83b, HGLS83c, KK14, Oui22, Zan14, Zan15, BR90, JS81, Pad81, Pad84, Pad87, Sch89b, Wol84]. centre [MO19, YS23]. Centred [Als93, QCO13]. centring [JH09]. Certain [Fak96, AB05, AK75, Fei72, Gno79, GS04, HAMAZ96, Mag72, Mic87, Nag78, Tay83]. Chain [GH02, Mel94, SS94, Ano81e, GH01, KS79]. Chain-Type [SS94]. chains [Gil78, Küc78, Lie92, Win17]. Champernowne [BLNGB05]. Change [HHS97, Mi96, NS14, OK02, Pon02, RV97, Bat10, Ber00, CG04, CT14, Ciu12, DI11, Fre15, HTX23, KK14, LLC16, PN90, QLT17, QLY21, Som71, SK05, VWL^{+}09, ZXT11]. Change-Point [OK02, Pon02, RV97, NS14, Ciu12, DI11, Fre15, HTX23, KK14, PN90, VWL^{+}09]. Changepoint [Koz96, Stu96]. Changes [Hor95, CHPS13, GS21, Sch70c, Wan07]. channel [Cha74]. channels [Mag72]. Chapman [AO07]. character [Sei78]. Characteristic [Fak96, NP12, Zie00, BP93, Hei86, HHM14, JRT21, NPS12, NDDDG16, RDLF03, WLD18]. Characteristics [HKMM98, GOS10, JLO12, Kre80]. Characterization [Bet87, FR95, Hof79a, KP91, LM23, McC03, MJA02, NN91, Ros72, Ros91, AF14, BM18, CB02a, ELV22, Fra77, Ger78a, Gla90, HKT22, JGMO20, KS92, Luc05, Mat87a, McC06, OJM15, RA01, Sha86, Waw81, YY07, ZB11].

Characterization-based [LM23]. Characterizations [BCD11, HS82, HS15, Jas18, NRD08, SU93, SSM07, YA12, Ahm01, ACS08, Gup84, Gup85, NRS07, Deb77]. Characterizing [WNBH17, Gup77, JS22b]. Charakterisierung [Fra77]. Charakterisierungen [Deb77]. chart [RS09]. charts [AK04, AKN05, AvC89]. Chebyshev [Bit72]. checking [BBK14, PN10]. Chernoff [PS77]. Chi [IKR00, Bed81, Bla88, BH89, CV78, DR15, HGLS82a, HGLS82b, HGLS83a, HGLS83b, HGLS83c, JLO12, LT16, Nad10, Oui22, Sch74b, Sta20, WN13a]. chi-processes [LT16]. Chi-Square [IKR00, Bed81, BH89, CV78, DR15, HGLS82a, HGLS82b, HGLS83a, HGLS83b, HGLS83c, JLO12, Nad10, Oui22, Sch74b]. chi-squared [Bla88, Sta20]. chi-squares [WN13a]. chirp [GKM18, LKM14]. chisla [Ano79e]. Chislenniye [Ano73a]. chlena [Ano87j]. Choice [Atk75, BM94, HK87, Vie95, Vie02, Ban70a, BKN70, Bon88, Bro93, Bun73b, BG78, BS06, FP92, GGHS21, Hav88, MB21, Nät75a, Ryc21, Sch74b, Zwa80]. chosen [Bun82c, KL82]. Chvátal [Ros75]. CIR [BBKP19]. circles [Pie72b]. circular [BR15, Gat09, JBC22, Kut12]. circular-circular [JBC22]. Class [BGTV06, Che97, IR91, MU99, OM91, TJZ89b, AA12, Amm20, BMK22, BN14, Bei74, BKN70, CFR89, CM15b, CS89, FC89b, GAJB13, Gno77, GVV13, GS82, GK09, HE12, HS91, Huz18, JK76, Jür84, Kim05, Kim07, Kim10, Kim12, KK15, Kim16, Läu74, MC75, MF17, MOS86, NB21, Nož72, PM21, PKP21, PB10, Pré72, RDLF03, Ros91, Sah87, Sch73, Sch74b, She14, WLT^{+}22, Wei09, WFK82, WN14b, YW12, ZZWW23, ZRAN17, vW74]. Classes [CG95, CR99, Mic86, Ara93, BSL13, Ber79, GKZ77, JGAFBCMG14, MFTFEO22]. classic [BV86]. Classical [HTHF16, Joh80, Ter72, Yu04]. classification [BB13, Cap73, DHN20, Das87, Kle75b, LT85, MPPZ05, Möh87b, MMR11, Rod88, SK92, Sei78, Sei79]. classified [LC10]. classifiers [Art19]. clinical [Gov11]. Cloning [JK07]. close [Ger73]. Closed [NN23a, NN23b, SLL15, SSG76]. closed-form [SSG76]. closeness [AR13, Joz14, JBD14, VDB13a, VDB13b]. closer [Els17]. clucaiiniye [Ano75b]. Cluster [Luc99, BL11, Góm18, Hei88, Juh89]. cluster-limit [BL11]. clustered [Lee08]. clustering [Cap73, HS83]. cm [HGLS83d]. cm-square [HGLS83d]. co [LYB19, LH22]. co-moment [LH22]. co-moments [LYB19]. Coates [PR89]. Coefficient [AL99, Bla14, CXH14, CT03, CT07, FX16, GP82b, HLM21, Kha15, PJ17, QJ08, SS15a, SL23, TW05, Tan14, TY12, Tua81, VT08, XY16, YLY17, ZZL^{+}14, ZZW^{+}19, ZL14, Som71]. Coefficients [Bla93, GK99, Düc75b, FN77, Hub77, Koz85, Spj77, TR78, Web06, Wei70b]. Cohen [Bha88]. coherent [AM90, Jas22]. cohort [Sha21]. cointegration [SL17, Shi15]. cointegration-based [SL17]. collage [Sto93]. collapsing [KI03]. Colloquium [ES74]. columns [Yin20]. combination [BOHJ12, BHS76, Nad07]. Combinations [PV98, KR17, LT85, Mat87a]. combinatorial [Sch70b]. combined [QCO13, VPR21, WQO23]. combining [KM20]. come [Sta17]. comments [Gus71, Bel79a]. Common [GKP96, Hol91, KS95, Bha88, GPK23, JC22, JT22, JS82, Kam22, KZ20, MM21, PK18, ST23, SKCE19, ZRAN17]. communication [ZZW23]. communication-efficient [ZZW23]. commutative [HK00]. Compact [Boe72]. Comparability [Lis78, Kir76]. comparaisons [VO75]. comparative [Moh86]. Comparing [AA03, AA08, KP07, Lis90, LNBR02, Ste98, TE15, Dav94, Gov11]. Comparison [AR13, BMRPZ16, FM89, Jör99, MO19, Miz17, Ohy13, THN13, AAA13, BMS20, GGHS03, GGS14b, GK93, LST18, MMP05, MB91, MB92, RT13, SK92, ST81a, Sch82a, Tak09, TC13, VP04, VFGM04, VDB13b, ZK16, Zwa85]. Comparisons [BT16, CZZ19, FE85, Tay83, ASKS16, BHN15, BS75, FLL16, FW20, HKK23, KR78, LAO21, MB21, TL15, Tor17, VO75, YN23, YYVH16, ZLLX08]. competing [BPK14, CDD17, DP17, FW13, KKG17, KM21, LST18, Mal21, SKD21, WG21, WLT^{+}22]. competitive [SDKM18]. complementarity [Bod75, GR76]. Complete [LLWW20, MU99, WDZH14, WZXH15, WWH18, DWHH19, Gno77, HSW85, RW84, Sch76b, Son82, WW22b, pZ15]. Completeness [Pla02, Fra79, Rüs83]. complex [EW73, MC75]. Complexly [Bel91]. Component [GL94, HA02, Are05, Çet21, CKS22, DDT88, Dry77, FR84, FP92, ID03, ILRMZ15, Jas22, KFPG21, KHN20, May11, QJ08, RN18, Rud76]. Components [Das96, GL94, GM94, GSTUM02, Lav95, MZ96, Tan91, ASSH13, AHJN01, AHPS84, Are05, Azp73, BKHB22, BCFF88, HK00, HKK23, HH13, Jas81, JR06, KK19b, Kle73, KP74b, Kle75a, Kle75b, Kle77a, Kle77b, KZ87, KZ15, LG05, LH22, Mat87a, MNvRS10, MF17, MA86, Mül85, NRS07, NL05, Ouy92, Özk09, Pin74, Pin77, Pol08, RD80, SK92, Tho80, WB05, YDZ13, Zez93, ZB11, ZB14]. Componentwise [PKP21]. Composite [JHYQ18, BMP15, JLM23, MPPV04, Sch91a]. composition [Väd77]. Compositional [KFH16, HTHF16]. Compound [NKT21, ASM83, Bök85, GMLS16, GML18, GN17, SX20, TE15]. Computable [Ped86]. Computation [HKMM98, Ege80, FGK19, Hav76, JBC22]. Computational [Cse95, Lau85b]. computations [Fer07b]. Computer [JK92, Bar87, Deu72, KYR89a, Väd77, WT20]. computer-based [KYR89a]. computing [Dry84]. concavity [FV06]. concentrated [FN77]. Concentration [Cor17, Bel83c]. Concentrations [Hol91]. concept [And77, BG19b, Moe84, Sch70b]. concepts [HBN86, Hof79b]. concerning [Jen90]. conclusions [Lät74]. Concomitants [Ver92, ASKS16, AA08, BCJ14, NS10a]. concurrence [RG84]. condensation [TK22]. condition [CLCP17, MV18, Röd87a, Sie87, Yu04]. Conditional [BN00b, Bun76b, CB02b, KM13, Laz16, LLV16, Lou99, Nog20, SR02a, AJV21, BGAL01, Chr23, DLMR13, FH01, GS11, GOV17, GGS14a, GS17, Jen90, Kim12, LdCIP18, LWL18, TIT08, TA11, TC22, Tom90, WD22, ZY07, Zha23]. conditionally [BZ10, FPPRG16, GK17]. conditionals [ACS05, Gup05]. conditioning [GS14]. Conditions [Lie98, And78c, Ano73c, EG74, KH21, LY93, Lit14, Ric76b, Sha83, Son20, Tan11]. cone [Foc75b]. Conference [Man72, PKN72, BK71, HW70a, HW74, Lom74]. Confidence [BB72, CR99, FC89b, Fis09, HHJ95, Hen79, HDA99, IV02, Neu97, RF94, RL98, Wan14, Are05, BG19b, BAM18, CKM09, CM05, CTW09, Fer07a, GC04, HAMAZ96, Hen80, HJM90, IR05, Ker16, Mau89, Moe84, OB19, Páz91a, Pol85, RVM22, Sha18, SDC19, WN12a, YL17, Zha18, ZZ05, Zvá79]. conflict [Bic18]. conflicting [CY12]. conformance [FE85]. Conjugate [ACS93, AINJ14, HS91]. Conjugated [EN75]. connection [Sär95, Ull76b]. consequences [Gla90, Yat17]. Conservation [Dew75]. conservative [Bic23]. consideration [Düc75a, Dup77]. considerations [Pie75b]. Considering [Ach95, RA02]. Consistency [AK94, Bel79b, Bel80, Car84, FR98, Fra79, GVV17, KM96, LW12, MWZ15, Pau12, Sch76d, Sha91, WS22, Xie09, Blo03b, Cap19, Che21, DR03, DdMT20, Hab89, JRZ03, Kle80, Läu89, LT13, Lit14, dDNN19, PN90, Pon06, Sch76c, WS13, WZXH15, WW22b, dUÁ00, dUÁRC04]. Consistent [DK21, Fur97, GLVV98, Mi96, Wit93, Bun76a, CLWK20, Fur15, Ger79, LQ12, Nov10]. constancy [Sen82]. Constant [Tij76, Eva75, GG15, Pad89]. constrained [Kim07, LST18, CF95a]. constraint [AGS09, AGOP14, KCL16, YYVH16]. Constraints [Hof97, KN98, DM09, Ehl75, Ehl76, Hof72, MP11, MSW13, NL05, Pré74b, STS85, Wei72, Wei74, Wen76b]. constraints-local [STS85]. Construct [CR99, Böt20, HAMAZ96]. Constructing [ETF19, FK12, OB19, DFRLÚF12, YL17]. Construction [GP82b, Haf82a, Ker16, Lis79, SKS93, BMS20, BN78, Böh81a, CCMW09, Haf82b, Hol70b, KHN15, LM23, Sin81, Ste76, SW89b, Tzi87, ZN14, Got74, Mau89]. Constructions [SJK97, Rüs86]. Contaminants [Zie95]. contaminated [MP96]. Contamination [Rie87, ACH20, Col91, Flo93, Haf82a, Vís86]. contents [Ped86]. context [BM01]. Contingency [AAJ97, BFF00, Bun85, GS11, KI03, KP15, Led78a, Mat87b, Moh86, PGT08, Röd71, TIT08, TMY05]. contingeney [Led84]. Continuability [Bar10]. Continuation [SS98]. Continuity [GR17, Gus71, KS76b, RL83, Sto73]. Continuous [BPR00, DY02, FR95, VYD83, Ahm01, BP16, BBKP19, BCP99, CM15b, CFC10, IV90b, IV89, IS84, KA88, Kum85, PM21, Ros71, SS21, Try89, VI90]. continuous-time [BCP99, PM21]. Contoured [GV95, AST13, Bat10, BG11, GN22]. contours [GN16]. contra [BB78]. contractions [CMLBSM12]. Contrast [BC99, BR15, Fav14, GC90, MW80]. contrasts [Mej78, PCKMK17]. contribution [Sti78]. Contributions [ST81b]. control [AK04, AKN05, Bit75, BS17, Dam76, Dav94, Gla76, JY20, RS09, Rui87, SS74, SSG76, ST87, Try86, Try89, Wer78, Wol75, vW72a, vW72b, vW74, Hec86]. Controlled [CFR89, BG78, Gre87, MMdP09]. Controlling [Tij76]. controls [GM78, Gus71]. Convenient [ACS96]. Convergence [BLO03a, CFLDÁ04, Dod83, DST21, GB23, JK92, Lie98, Ruy98, Sie86, dSL22, AL18, BR90, BL82, Bed89, CC13, Col85a, Cou21, DWHH19, Fäh78, Fäh81, Flo93, GBS20, Gün83, HP08, Her81, Hos88b, HWL17, Jen90, Kin75, KT83, LLWW20, Mal21, MB21, Neu73, Pad87, Pon86, Pra86, QLT17, Ros89, Sch89b, STS85, WDZH14, WWH18, Wit83b, YLWL22, pZ15, ZZLL20, ZZWW23, FGR73, Hol76, Müh80]. Convergent [Cad01, SL95, BGAL01]. Convex [Fed80, Tuy76, Gaf85, Ger73, GR76, Gud76, Mel76, Puk81a, Ric76a, Sch75f, Tam74a, Tuy74]. convexity [GB23]. convolution [PN13, RD80]. Conway [GG15]. cooperative [Ano75a]. coordinate [DH77, LNO14]. coordinate-free [DH77]. Copula [Böt20, BCN15, dUÁV17, ANN11, BKHB22, BRF20, BE12, BZ13, FZLB22, GJV16, Oko17, PM16, SlP17, SKCE19]. copula-based [BRF20]. Copula-graphic [dUÁV17]. copulas [DFRLÚF12, FK12, GOV17, Röd04, SKSZ20]. core [HS83]. Cornish [WN12a]. correct [YL17]. corrected [Fik88a]. Correcting [GS14]. Correction [SP98a, KS16, NTJ09, Rat93]. corrections [DB92]. Correlated [AAK98, Bis95b, MU99, Arn91, BB13, BBL20, BKS23, BN82, Cip84, DK14, HK15, JRZ03, LMP10, MO73, McC05, Nät85, PW15, SO09, TSG20, Wu16]. Correlation [Bla93, Fou97, GK99, Luc99, CT07, FK90, Geb84, Ger79, Ger80, Hös74, Hös76, Hös77, JS14, Stu90]. correlation-theoretical [FK90]. correlation-type [Geb84, Ger79, Ger80]. correlations [GF09, GT13, Kim12, Yin20]. correspondence [BHM04]. Corrigenda [Ano19]. Corrigendum [Ano15a, Ano16a, AE21, Nku23, SN14, SGMFB18]. Cost [Mag95, Wäc86, Ban70a, Bei81, Eva75, Gov92]. Cost-optimal [Wäc86]. Costs [Tij76]. cotingeney [BH89]. count [AWJF22, Fok11, LF22, PG19, SKSZ20, TK22, XW10]. count-based [LF22]. countable [RW84]. counter [LO86]. counterexample [Hel93]. counterparts [Kow77, KL82]. counting [FM01]. counts [HHM15]. cousins [RD21]. covariables [AdUÁS15, FSV13, dUÁRC04]. Covariance [CKS22, Fur97, GON95b, IRV98, MD03, Wen95, Zez93, AdUÁS15, BA89, BW81, BS17, BW80, CFS17, CG04, Dry77, Dry84, GK93, GS04, Hös74, Hös78, HB09, Jas81, Ker84a, Ker84b, KP74a, Kle77b, KvR00, Koz86, Lis90, LH15, LGL^{+}17, Mül85, NL05, Pin74, Rod88, RFK18, Rud76, SGG04, ST81b, ŚS05, SX20, XB17, YT01, dUÁ00]. covariance-operators [Dry84]. Covariances [FZF97, Fei72]. covarianee [Kle78]. Covariant [Das96]. Covariate [RA02, AFV15, BD13, DM06, GOV17, GGS14a, MSW03, Zha16]. covariate-adjusted [BD13]. covariate-matched [MSW03]. covariate-missing [Zha16]. covariates [CLLY22, DDD19, GB17, GNYX15, HDC20, Lem13, LQZF17, LY16, MMR11, PW15, Pon04, Pru93, sS15c, YS23, ZWYG14]. covariations [HS17b]. Covariogram [GC95]. covarites1 [Den88]. coverage [STY04]. Cox [GJ19, MA20, Pon02, sS15c, sSW19, ZWL14]. Cramér [Neu73, Neu73, PV85, Son20, GK95]. Criteria [BK76, CFLDÁ04, Gaf85, HS91, MR73b, Nag78, RDLF03, ST81a]. Criterion [Bis95b, BM94, GM99, MG96, AN13, HG05, Kam22, LLL17a, Nät75a, Nie06, PR22, SK14a, TE15, Tom90, ZYZ21]. critical [Sto76a, KNS72]. Cross [ASFKV08, Sto78, Cor17, GW04, KZ20, LC10, PA14, SO09, Shi15, WC21, Zwa85]. cross-ratio [PA14]. cross-sectional [GW04, KZ20, Shi15]. cross-sectionally [SO09]. Cross-validated [ASFKV08]. Cross-validation [Sto78, Cor17, WC21, Zwa85]. crossed [SK92]. crossing [Fra77]. Crossings [KS83a, HY10]. crossover [BBM12]. CSSR [Lom74]. Cucconi [NM18]. cum [AK85, BB07, BCDD11]. cumulants [WK13]. Cumulative [CBP23, ABP18, BD17, CGS23, Dik01, DP17, HKT22, HGLS82a, HGLS82b, HGLS83a, HGLS83d, HGLS83b, HGLS83c, JD20, MH22, SK21, SL12, SSN17, ZFA20]. cure [WH20]. current [Mal21, WH20]. Curvature [Páz93, SS86]. Curve [HHJ95, Wen95, vR97a, Hu10, JRT21, KvR00, Szc79, Von91, Zez93]. Curves [Cro96, BP93, CCH87, LW18, SDC19, VFGM04]. cusp [Dör15, SG91]. customers [Sin84]. CUSUM [BS17]. cut [Har75a, Har75b, Pie73]. cutpoints [SS15b]. cuts [Pie70, Pie74a, Pie74b, Ste76]. cutting [BTS71, Ros75]. CVaR [DM22]. cyclic [Ano73f]. cyclically [ILRMZ15]. cym [Ano87j]. cyshchestvovaniya [Ano73c]. Czanner [MA20].

D [NR81]. Dalenius [Deu72, Deu72]. Dalenius-Gleichungen [Deu72]. dam [Sto71, Sto72b]. damage [JR86]. Dammtheorie [Sto71, Sto72b]. damping [CLCP17]. Dantzig [Jen73]. Data [AJ97, Bel91, CM13, Ger89, JW04a, JW04b, LP97, LTG97, LNBR02, MO03, Nät97, Neu94, OM10, RO98, RF03, Vie02, Yao00, AJM12, AAFG23, AFSS22, AHPS84, BC23, BO01, BS07a, BS10, BPK14, Bic18, BCP99, BRF20, BAO12, BR15, Bun76a, CM15b, CLWK20, CL14, Cup21, DHN20, DRW17, DLMR13, DK12, DT16, Eco13, FSV13, Geb84, GZ23, Ger87, GB17, GJLT15, Guo20, HH16, HOCC13, HVRI20, HTHF16, HS17a, HS17b, IJZ08, JC07, JIP10, JS22a, Jia07, JLM23, KS16, KRC19, KK19b, KKG17, Koo86, Kre18, Kum92, KM21, Lee08, LGZL14, LdCIP18, LQZF17, LW12, LLV16, LV18, LY20, LC10, LGL^{+}17, MB18, Mal21, MP96, MFTFEO22, Mie91, Mou03, MAS06, NSG07, Nad08a, NJKK16, NTJ09, PP83, PG19, QXY23, RAD07, SKSZ20, SPTB18, SM08, SP08, SMW22]. data [sS14, sS15c, SW16, sSLMJ17, SC18, sSW19, sS20b, SA23, SKD21, SAX23, TA21, Tan14, TX17, TJZ89a, Try88, TT22, WY04, WLF13, WH20, WLT^{+}22, WD22, Wit93, XW10, XB17, XCJH19, YLL16, Yin20, Zha12, Zha16, ZZW23, ZT15, ZTJ17, ZL13, ZL14, ZBF^{+}20, ZL17, dUÁV17]. Data-dependent [CM13, OM10]. Data-Driven [Neu94, Vie02, JW04a, JW04b]. dataset [MM22]. datasets [JHYQ18]. date [Col85b]. Davidson [AAA13]. DDR [Atk75, Gir74, HW74]. de-aliasing [Els17]. death [Fra82, Goe75, Küc82b, PDG13]. Decay [Phe96, Mat89]. Decision [AN98, CPS98, GJML00, Kal97, Man72, PKM78, BG74a, Exn76, HBN86, HR80, IV90b, IV89, Jür84, Lus87, Pré72, RH80]. decisions [Bor70]. Decomposition [Azp73, Got76, Ano73e, Cap73, DGCL14, Eic88, Ger75, Jen73, Lud73, Pfl80, ST75, Tom90, Cap73]. decomposition-like [Jen73]. decompositions [TT10a]. deconvolution [BvET16, CDD17, Duv17, MP22, Yat19]. Deconvolving [SC90]. decreasing [BSB13, BN16, Bie16, FPLS12, LMR14, Mat89]. decrement [Fei72]. deficiency [Faz14]. defined [FG06, Kim12, LT13]. Definite [ETD99, Gne98, RR02]. Definiteness [Jür84]. Definitheit [Jür84]. Definition [Bel79a, Kow77]. definitions [FMM85]. degenerate [Ber76, Nad80]. degradation [BD17]. Dekomposition [Lud73]. dekompositsii [Ano73e]. Dekrementtafeln [Fei72]. Delay [CHPS13, BP17, RS74]. Delayed [JRM99, JRM04, Sie74]. Delbrück [Zhe10]. Delta [NOM13]. Densities [CH91, ETD99, Gne98, BCK05, Dev89, Gup84, Hei88, Kha12, Kob87, Lie90a, LBSW19, Pet86, Ros91, RR02, Tho84]. Density [BDG95, Fer99, GW84, Hen92, KT96, KFL97, Koz02, Lie92, Lie98, Mam95, Ruy98, TP97, vR97a, AG23, BMMP16, BLO03a, BLNGB05, Cao01, Car84, CCD15, CHL09, CFC10, DLMR13, Dev89, DW93, tFYK01, Fik88a, Gün77, Hal91, HY10, Iga20, JC07, JIP10, JD20, KPH^{+}16, KvR00, MR73b, MP96, Mic09, MR12, Moj07, MB21, Nad80, NTJ09, PGGF15, PYHT01, Ped86, PS10, Röd87b, Ros91, Sta17, SC90, Tro91, Tro93, WS22, Wer85, YFZ11, Yat19, ZL17, vdVvdL03]. departure [Pol86, TIT08, TMY05]. departures [Mad76]. Dependence [Cra06, Dor85, Led84, Lie98, SR02b, Tjø96, And79a, BKK03, BL82, Böt20, CD09, Dab81, DMR07, DL16, DJ12, FLL18, Hav84, Hös74, Hös76, JK08, Jaw13, Kou11, Kow77, Kow82, LLBM^{+}11, MV18, NKT21, NL79, Röd70, Röd87a, Ryc93a, Shi15, Szc79, Tam74b, Vís86]. dependencies [JK76, KS76a]. Dependency [WL05, MP22]. Dependent [CHP03, Góm18, Aly87, Amm20, AOV19, BM12, BKHB22, Bed86, BMRPZ16, BTV21, Car19, CM13, CGM19, Dam76, DRW17, FRLLT18, FLL16, FW20, FZLB22, FPPRG16, Fre74, GS14, GGHS21, HKK23, ILRMZ15, Kim16, KPH^{+}16, Kla72, KLF22, Kre18, Kum85, Kum89, LdCIP18, LT16, Mec72, NB21, OM10, Pap16, PGT08, Ryc93b, Ryc15, Sar82, Seb76, Sep01, SW16, Sta84, ST87, Ver06, VFGM04, WLF13, WDZH14, WZXH15, WG21, WLT^{+}22, WWLH19, WW22b, YN23, Wei71]. Depending [BC99, Goe75, Goe76, NK88]. depth [GN16, HVRI20, TF21]. depth-based [TF21]. deren [Ric84, Sie74]. derivation [Bor73]. Derivatives [CH91, JO97, SLS99, BO01, DGL16, Gaf85, HS91, Yat19]. Derived [RL98, Ros89]. described [Pen85]. Describing [Sto91]. Descriptive [Dor85]. deserted [Sto88b]. Design [CF95a, DY02, GS12, KE96, LSS12b, SY22, Ban80, BNP87, BB07, BBM12, BD13, BDP16, BM01, Böh81b, BN80, BCDD11, Bun76a, CL12, Fed80, FM89, Gaf85, GGHS21, HR80, HS91, JY20, Joh80, KM82, KD74, Lät74, Läu74, NN82, Nät79, NW12, Páz80b, Pil81, Puk81b, RH80, Sch87f, Sha21, Sin13, Tit80, Wie88, Wit76b, Wyn76, Ull76a]. Design-based [LSS12b, SY22, BDP16]. Designs [Bis95b, Det93, Det95a, Imh00, MU99, MG96, Sch96, SD83, SKS93, SJK97, Afs85, ASSH13, Bag90, BN78, BBL20, Böh81a, CGK17, COQ21, Dav94, DMP07, DMP10, DK14, DM06, ETF19, GK77, GLZW12, GSA14, GP82a, GP82b, Gno76, GGHS03, GGS14b, HSW85, HM88, HM89, Jac91, Jas81, KHN15, Kha89, KN90, KAM17, Läu76, LFGR11, Mel78, MK83, Nag78, Nät75a, Nät85, PMA17, PEB15, Páz80a, PCKMK17, PM16, QCO13, ROM15, Sch77a, SDKM18, Sha18, Sin81, Sin91, TE15, VYD83, WQO23, Wie86, XWQN23, YC15, ZW20, ZYZ21, ZYW21]. desparsified [PR22]. Detect [Khu95, CCM18, RS09, ZXT11]. Detecting [Hor95, MBF^{+}20]. Detection [Cie98, OK02, ADR14, Bat10, Bha07, Fre15, NS14, Pon93, SK05, WWZ22]. determinant [SN81]. Determination [BFR84, Kla72, Bak73, FN77, GM78, Pie72b, SS86, VT08]. determine [Deu72]. determined [Dem14, DJ17, DAA20]. determining [Gir71, Gri80, ST81a, Lom70]. Deterministic [Ric75]. Deterministische [Ric75]. development [GPS14, Sch73, Tit70]. developments [Joh80]. deviance [SK14a]. deviance-based [SK14a]. Deviation [BGP95, Cro96, FL10, Nad80, Bic18, DJ09, LKM14, Lie82, MP10, MW14, SX20, Wol84]. Deviations [Lou99, Bos93, JC22]. device [BC23]. DFR [Bie06, Bie08, Gor20, Ryc09]. DFRA [Bie06, Bie08, Gor20]. dHSIC [Böt20]. diagnosis [Nog20]. Diagnostic [LKK07, VN19]. Diagnostics [CLS14, DT16, HOCC13, MCML15, TWZ06]. Diagonalization [CLDG16]. diagonals [TMY05]. diagonals-parameter [TMY05]. Dickey [Wes14]. did [Sto93]. difference [AAR18, BQSR20, Bel74, Bel79b, BG79, Che12, DST21, KLF22, LW18, LZW23, MWZ15, Tab13, Try91a, Wu16, ZL13]. difference-based [AAR18, Tab13, Wu16]. differences [Ber74, Gre87, Pad87, PG19, Ros72]. different [BG74a, Deu72, DB92, HT15, Rud76, Try88, Yin20]. differentiability [PD91]. differentiable [HR93, Web06]. Differentiablity [HR93]. differential [Bel80, Bel83b, BP17, Bit75, FL10, GCL16, Mus77, NS22, Pen85]. Differentialgleiehungssystemen [Bel78]. Differenzen [Bel78]. Differenzen-und [Bel78]. Differenzengleichungssystemen [Bel74]. Differenzensystem [Gre87]. differing [KR78, WBL06]. Diffusion [AL99, IR91, AG23, ALS11, EMF19, Fav14, FZ89, GC90, KHNT17, Kut84, Lie86, Mic86, Rao83, Rao85, Sch13, SL23, Tua81]. diffusion-type [Mic86]. diffusions [Huz18]. Diggle [PR89]. digraph [Cey20]. dilations [CMLBSM12]. Dimension [Chr23, Lom70, SYW22, AX16, Ano75c, Art19, KMM21, LLL17a, LCY17, TSC20, Yoo16, Yoo18]. Dimension-reduced [SYW22]. Dimensional [Páz93, AC23, And80, BJfYZ13, BQSR20, BS07a, BCN15, DS91, DFRLÚF12, FYY^{+}23, FR20, GKM18, HH16, Jaw13, KH21, LKM14, Mat87b, PR22, PEB15, Páz90a, Pet86, Röd71, Sch90b, SA23, SX20, TA21, WY04, Wit81, Yin20, ZZW^{+}19, Zha23]. dimensionality [FYY^{+}23, PvR01]. Diophantine [Pie74a]. Diophantische [Pie74a]. direct [Ege80, Iga20]. Directed [MPV00, Pie72b]. Direction [Hol91, Foc71]. direction-of-search [Foc71]. Directional [Gaf85, HS91, LC15]. directions [Hol75, Hol76, Ohs81, Pré74a]. Dirichlet [Cha96, HK09, NB21]. Discontinuities [CH91, Lie90a]. discontinuity [Fäh78, Fäh81, Hoy79, Sto72a]. Discontinuous [CH91, Ano72a]. discounted [Pad87, Rei86]. discounting [JS81, Pad84]. Discrepancies [EMG98]. discrepancy [LPS90, QCO13, WQO22, ZQC23]. Discrete [BN00b, EMG98, Fer99, Nag92, NAB13, Saw81, Ahm93, And79a, BJN14, BN78, BM18, BMM13, BF85, CF21, DI11, Ege80, FZ89, GC90, Glo01, Hol75, Huz18, Jür84, KT12, Kie17, Kir76, KN06, Luc05, NK17, PA14, PVW74, Ric76c, RW79, Rus91b, Sch76b, Sch75f, sS15c, SKD21, Zhe10]. discrete-domain [KN06]. discrete-random [Ege80]. discrete-time [FZ89, Rus91b]. discretely [AG23]. discretization [GM78, HT15]. Discretizing [GN17]. Discriminant [FG06, Lau85b, BF85, FT79, Lau85a, LL18]. Discriminating [DK12, KR07]. Discrimination [PJ05, RPZB18, PvR01]. discs [Kos80]. Discussion [Arn92b, Atk95, Bin96, Bis95a, CF95a, Det95b, Elt96, Kra95, Lee95, Mül95, Páz95, PW95, RS96, Sch95, Sha96a, Ski96, Val96, Won95, BNP87, Páz90c]. discussions [Hös77]. disease [BCDD11, Nog20]. Disguised [GN98]. diskrete [Hol75, PVW74, Ric76c, RW79]. diskreten [Sch76b]. diskreter [Jür84, Sch75f]. Diskretisierung [GM78]. Dispersion [Dor85, Phe96, CTW09, Lem13, Lin09, Ryc15, TWZ06, ZT15]. dispersions [Miz17]. displays [BCFF88]. dissipation [CLCP17]. Distance [BP93, Gat94, BG16, Böt20, CWW09, Haj84, JRT21, Lie93, TIQ10, WLD18, WK15]. distances [BCMR90, Bed89, BH89, Liu92, Rat93, SM19]. distinct [Agr82, HH13]. distinguish [AN13]. distinguishable [HC20]. distortion [HLM21]. Distributed [HR96, Ron76, Aui90, BNAA13, Gai23, KZ13a, KZ13b, Mau89, Ped86, Ric84, RH20, Sch76d, Tho13, Wie92, ZZW23, Zin81]. Distributed-Lag-Modellen [Ron76]. Distribution [Ach95, Ahm74, AHJ99, AEW94, BGP95, BN00b, Cau99, Cie98, ESSB93, FC89a, Gup95, GLVV98, JRM99, KS83b, KvR00, Koz02, Lai00, Lik85, MS94, MS98, Oko17, Páz87, PR89, PH16b, RO98, Tho13, TVL96, VW96, Zie88a, Zie00, vR97a, vR97b, AJM12, AMB11, AFL08, ANR07, AGS15, AG09, AA11, AA13, AE21, BKT14, BA08, BJN14, BCD11, BOHJ12, BZ14, BN03, Bar94, BSL13, Bei75, BS07a, BS10, BW81, BS07b, BG11, Bow12, Bow21, BF85, CLDG16, Car19, CHA15b, CB04, CT73, CM05, Che21, CBC12, CK06, CRCN14, CCMdC13, CSO13, CLO13, CL14, CV78, De 22, Deb77, Dom10, ELV22, EJ81, Exn76, FdHP03, FG06, FT79, FE85, Fre74, Geb84, Ger79, GG15, GN17, GF09, GW04, Gui86, GJ73, Gup79a, Gup03, Güv06]. distribution [Haj84, HGLS82a, HGLS82b, HGLS83a, HGLS83d, HGLS83b, HGLS83c, HP08, Hen89, HSG13, HWL17, HM07, IM14, IZ83, JC07, Jai74, JN75, JK13, JW04b, JSV07, JD20, JLO12, JRM04, Ker84a, Ker84b, KHN16, Kha15, Kim10, KM20, KZ86, KK11, Koc74, KA13, KR07, KG13, Kun15, Kun17, KM21, LLBG10, LSS12a, LUK13, LUK14, LC13, LCMA16, LD21, LBSM00, LH91, LMR14, LM23, Mad76, MR73a, MR73b, MR75, MFBG14a, MFTFEO22, MCCL18, Müh80, MKM21, Nad03, NK05a, Nad08a, Nad08b, Nad10, NH11, NP12, Nad12, NSN17, NN91, NS10a, NL79, Nät75b, Nät79, NN23a, NN23b, NJKK16, NAB13, NK17, Neu73, Nig90, NAHJ03, Nov10, OJM15, PB10, PG09, Phe89, Pog15, Pon93, PDG13, PHS09, QXY23, RAM^{+}17, Raq04, RGB13, RNT13, Ric84, Röd76]. distribution [RVB18, RS74, SN14, SS81, SP01, SV89, SW85, SY22, SDJ16, SYM16, Sha86, SKV18, SK21, Sol19, Sta88, Sta90, Sto72b, SD17, TG87, Tay83, THN13, TH04, TJZ89a, TPSR17, Try91b, TDY08, Tzi87, VS81, VFS^{+}20, Vís11, Wal88, WG21, Wes93, Wes14, WBL06, WN13a, Wit78b, Wit83b, Wit86, WHT03, WWC^{+}10, Xek86, YA12, YH15, YY07, Yin20, YT01, Zäh14, Zha13, Zhe10]. Distribution-Free [Zie88a, Ahm74, BF85, FT79, MKM21]. Distributional [Yat17, Böt20, Páz90c]. Distributions [ACS05, BT08, Bed85a, BT97, CB02b, Fak96, Fal95, FR95, GM97, GV95, GK99, IKR00, KT96, Nag92, PGT05, SR91, SR94, SS98, SR02a, AA12, AT13, Ahm01, Ahm93, AT05, AT08, AMA14, Ano73f, BCKS01, BLSC09, BB04, BN14, BKHB22, BGTV06, BL15, Ber79, Bet87, Bie16, BM18, BS18, BS88, BW14, BLNGB05, CB02a, CS10, CG05, CTL14, CF78, CF89, CCMW09, DLO22, DK12, DR15, DSG91, FC89b, tFYK01, FH01, FV06, Gat09, GN22, GAJB13, GVV13, GMLS16, Gor19, Gor20, Gup77, Gup79b, GS82, Gup85, Gup05, GK09, Haf82a, Haf82b, HG08, HZ21, IRB14, Jas18, JG85, JK07, JS22b, Kal76, KC01, KS86, Kim05, Kim07, Kim10, Kim12, KK15, Kim16, KCL16, KSS76, Kow77, KT89, KP91, Led78a, Lem13, Lis78, Lis79, LBSM02]. distributions [Luc05, MO73, MA20, Man22b, Mat85, Mat89, Mat87b, Miz17, NK04, Nad05, ND05, NK05b, Nad06, NK07, NSG07, Nad08c, NPS12, NNM14, NKT21, NRD06, NRS07, NM18, Nus76b, Ohy13, Oui22, PAW05, Pan84, PK18, PKP21, Páz90a, Páz90b, Pel11, Pet86, PJ05, RN18, RPZB18, RLAV13, Röd71, Röd77, Ros72, RW79, Ruk07, RVM22, Rüs81, Ryc21, SU93, SD86, SZ86, She88, SH09, SSM07, Szc79, Tit90, VPR21, VK20, WLD18, WLT^{+}22, Wit81, WW21, XG07, ZHG12, dQLS16]. disturbances [ST87, Try86]. Divergence [GLVV98, BMP15, BB22b, BMMP16, DGL16, JGAFBCMG14, MA20, Man22a, Man22b, SSN17, Tak09, TMY05, WNBH17, Yat17]. divergences [BV11, GNP08, SV10]. divergent [LLL17a]. diverging [Ciu20]. diversity [SlP17]. Divisible [SS98, BL11, LH91]. division [Deu72]. dla [Ano74]. dls [Ano87j]. dlya [Ano79e]. DNA [LDF17, MP11]. DNID [Jas22]. Does [KJ14, PJ05]. domain [BVP06, JK08, KN06, SZ17, Sti78]. domains [NOM13]. Dominance [BM02, Kla03]. Dominance- [BM02]. Dominatability [Döh81, Ric86]. dominate [KZ87]. Domination [Cey20]. Dominierbarkeit [Döh81, Ric86]. doppelt [HW71, Wei70a, Wei71]. Doppelt-einparametrische [Wei70a, Wei71]. doppelt-einparametrischen [HW71]. doppelte [Lud73]. Double [Abd99, Cau99, DY02, CH08, CBF23, HW71, Lud73, PN10, SK21, SD17, WWH18, Wei70a, Wei71]. double-indexed [WWH18]. double-sampling [CH08]. Double-single-parameter [Wei70a]. Doubly [AJ97, Las91, LQZF17, Zha16, Ahm01, Ahm11, AGS15, Kim10, Mou03, SC18, sS20b, TPSR17]. Doukhan [BFN17]. down [BN16]. drawn [GR06, MR73a, MR73b, MR75, Ryc04]. Drift [AM15, Phe96, JSW22, KMM15, Rao85, Sch13, SL23, Tua81]. Driven [MS98, Neu94, Vie02, JW04a, JW04b, NS22]. dropout [ZL14]. dropouts [LQZF17, SMW22]. drought [Nad08a]. dual [Bit72, EBQY17, Mel76, Sin13, SD83, Sto79]. duale [Mel76]. Dualität [Foc71]. Dualitätsaussagen [Trö75]. Dualitätsprinzips [Nož71]. dualitätssatz [EN74a]. Dualitätssätze [Lip75a]. Dualitätstheorie [ES73, EW73, Tam74a]. Duality [Enn98, PM21, Böh83, ES73, EW73, EN74a, Nož71, Trö75, Foc71, Lip75a, Tam74a]. due [Bod75, GS14]. duel [Ano72b, Sty75]. d’un [Pon86]. d’une [Cap73]. Dünne [Nag83]. durch [Düc75a, GM78, Lud73, SW85, Tau84]. Dynamic [Kov74, SL12, Bor73, GK17, Got76, KTW87, LM79, Man22b, May11, Mül74, Ric75, SGG04, SSG76, TW83, TT22, Wer78, Seb76]. dynamics [GH01, GH02]. dynamische [Ric75, Seb76]. dynamischen [Bor73]. dynamischer [Mül74, SSG76, Wer78]. Dzhaparidze [FGL17b].

each [Fin94, Jas81, Oko17]. Eckpunkt [Sch74a]. Eckpunkt-Optimierung [Sch74a]. econometric [ST83]. economic [Mat75]. Economics [Lom74]. Edge [Fik88a, Kla72, Rat93]. Edge-corrected [Fik88a]. edge-lengths [Kla72]. Edgeworth [Hal91, Son20, WG92]. Edgeworth-expansion [WG92]. edinichnom [Ano72a]. Editorial [Ano77a, Ano77b, Ano77c, Ano78a, Ano78b, Ano78c, Ano78d, Ano79a, Ano79b, Ano79c, Ano79d, Ano80a, Ano80b, Ano80c, Ano81a, Ano81b, Ano81c, Ano81d, Ano82a, Ano82b, Ano82c, Ano82d, Ano83e, Ano83f, Ano83g, Ano84r, Ano84s, Ano84t, Ano84u, Ano85v, Ano85w, Ano85x, Ano85y, Ano86q, Ano86r, Ano86s, Ano86t, Ano87f, Ano87g, Ano87h, Ano87i, Ano88d, Ano92a, Ano92b, Ano92c, Ano92d, Ano93d, Ano93e, Ano93f, Ano93g, Ano93h, Ano94a, Ano94b, Ano94c, Ano95a, Ano95b, Ano95c, Ano95d, Ano95e, Ano96a, Ano96b, Ano96c, Ano96d, Ano96e, Ano97a, Ano97b, Ano97c, Ano97d, Ano97e, Ano97f, Ano97g, Ano98a, Ano98b, Ano98c, Ano98d, Ano98e, Ano98f, Ano98g, Ano99a, Ano99b, Ano99c, Ano99d, Ano99e, Ano00a, Ano00b, Ano00c, Ano00d, Ano00e]. Editorial [Ano10, Ano11, Ano12a, Ano15b, Ano16b, Ano17, BFN17, Bun92, Dou11]. Eds [Hec86, War85b]. Effect [OM91, PvR01, CAK13, DW93, GS12, Güv14, Mak06, MOS86, RM16, WN12a, YS23]. effective [BN82]. Effects [BS82, HA02, Mon96, PW15, Are12, CCM18, DT16, EMF19, Els17, FMTZ07, Gai23, GCL16, GOV17, HG05, Jac13, JS14, Kha89, KN90, LLW17, LY93, MNvRS10, NW12, OM11, SK92, Sei79, ZYW21]. Efficiencies [MU99, HSW85, MO19]. Efficiency [Bis95b, KB00, SS94, WY13, BG19b, Bel80, Cao01, CL12, Gno76, HR93, KK19a, Nus77a, Nus79, PW15, Röd71, VS81]. Efficient [EP17, GKP96, GSA14, GP19, Tru18, WH20, ZTJ17, BZ10, Els17, GB17, Lau85a, TMN00, WC21, ZZW23]. Effizienz [Röd71]. eigenelements [DGP91]. Eigenschaft [Pie73]. Eigenschaften [Sak71, Zäh74]. eigenvalue [Juh89]. Einbeziehung [Seb76]. Eindeutigkeit [Nož70]. einem [FR86, Hof79b, Hol76, Neh75, Röd70, Tam74b, Tou70a, Ull76b]. Einflusses [Düc75a]. Einflußparameter [Seb76]. Eingangsinformationen [Düc75a]. Eingangsintensität [Fre74, Sak71, Zäh74]. einige [Lip75b, Lip76, Pie75a]. einigen [Gru71, Kin75]. einiger [Neu73]. einlinige [Sto72a]. einparametrische [Wei70a, Wei71]. einparametrischen [HW71, Nož72, Wei70b]. einparametrischer [GT70]. Einseitige [Pad81]. einzelner [FN77]. eksperimenta [Ano74]. EL-MM [ÖA21]. elastic [Ciu20, LJ22]. elastic-net [Ciu20]. electrolyte [FN77]. Elektrolytlösungen [FN77]. elementary [Hei85]. Elimination [Lus87]. ellipsoidal [ST91]. ellipsoids [Als93, Wil08]. elliptical [BOHJ12, BS07b, BZ10, DGCL14, Fri74]. Elliptically [GV95, AST13, Bat10, BG11, GN22, NK05a, RH20]. embedded [Str82]. emphasis [Gup84]. Empirical [AFSS22, BMP15, BG74a, BG79, CC12, DZ22, GNYX15, HM02, Hof97, LLL17a, Mar97, MA86, ÖA21, Ras06, SV89, TZ13, Vís11, WLD18, YC15, ZL14, AÖ23, BQSR20, BLO03a, BE12, BZ13, BS75, Bun77b, BR83, CC01, CBC17, Cor17, Cup21, DJ09, DM22, Góm18, HE12, HP08, Hei18, HHM14, KDM22, LY16, LY20, NDDDG16, Pau12, Sar82, SYW22, TX17, WW22a, YYVH16, YLWL22, Zäh14, Zan15, Zha16, ZW10, ZMY20]. Empirically [Gla76]. Empirisch [Gla76]. employment [RS74]. Endlichdimensionale [Sch90b]. endlichem [Küc78, KS79]. endlicher [Röd71]. endogenen [Tou70a]. endogenous [Tou70a]. Endpoint [Fal95, Dup17]. endpoint-inflated [Dup17]. Ensuring [HDC20]. entarteter [Ber76]. entities [Ano75a]. entropic [Bow21]. entropy [CBP23, CK06, GGL15, JRCB22, JS22b, RA01, SK21, SL12, ZFA20]. Entscheidungsprobleme [Jür84]. Entscheidungsverfahren [Exn76]. Entwicklung [Sch73]. environment [Nor90]. enzyme [SDKM18]. epochs [Los23]. Equal [SR91, Rod88, Sty75]. equality [DP17, STS85, WLD18, ZY07]. equation [Bel74, BP17, Ber76, FL10, Mus77, NS22, Sch90b, TG87, TKTM15, Tua81]. equations [BK79b, Bel79b, Bel80, Bel83b, Bit75, Bun76b, BG79, CKS22, Deu72, GCL16, Pen85, Pie74a, Pon06, Rus91b, Ter75, Tou70a, TW78, Wag83, Wol75]. equi [JRZ03, MO73]. equi-correlated [JRZ03, MO73]. equiblock [Sin91]. equiblock-sized [Sin91]. equicorrelated [Rud76]. Equiprecisional [Rui85]. Equivalence [IR91, MP82, Bak84, Bed89, NR81, RM17]. Equivariant [MM21, TB78, Bur08, VK20]. Ereignissen [Sch87a]. ergebnis [Neh75]. Ergebnisse [KJKR76]. ergodic [HBCL14, KHNT17, LLV16, Win17]. Ergodicity [FR98, AP07]. Erhaltungssätze [Dew75]. Erkenntnisse [Tam76]. Erlang [Fle71, Mec72]. erlangsche [Mec72]. Erlangschen [Fle71]. Ermittlung [GM78]. Erratum [Ano14, Ano15c, GSB11]. erroneous [Düc75a]. Error [BDG95, GM99, Neu97, Pad96, RF03, AS04, BS07a, BS10, BCP99, BG78, CY12, Das87, EJ93, HEHM13, Hös78, HM07, JRCB22, JS14, Kin75, KY22, LLBG10, Lau85a, LLL17a, LQZF17, LJ22, ML21, MCCL18, Möh87b, MSW03, Nus77b, Pad81, Pad84, PB10, Ras06, SS20a, She88, Vís86, WLW16, Web06, Win17, WA22, XY16, Yat19, ZWYG14, ZL13, ZZWG09, Zwa85]. Error-Bound [Pad96]. error-covariance [Hös78]. error-estimates [Pad84]. error-in-variables [XY16]. error-prone [ZWYG14]. Errors [LMS92, MU99, Van98a, AK03, AST13, BM09, BQSR20, BBL20, BKS23, BL06, BN82, CLS14, Car19, Che12, CC12, Dup77, HLM21, HK15, ILRMZ15, JRZ03, KMSS10, KZ13b, LMP10, LMS06, LT13, LMV17, Mal89, McC05, MWZ15, Nät85, Pen85, RH20, Sch91b, Sep01, Vid23, VFGM04, WZXH15, Wan15, WYLW22, Wu16, WWLH19, XWY^{+}20, ZWFZ22]. errors-in-function [BQSR20]. errors-in-variables [CC12]. errors-some [Mal89]. ersten [Atk75]. erster [Gün80]. Erwärmung [Dew75]. Erwartungswert [Exn76, Mau89]. erweitertes [Wen76b]. Erweiterung [Pré74a]. Erweiterungen [Ehl75, Ehl76]. Erweiterungsmethode [Sch75f]. Erweiterungsprinzip [Sch70b]. Erweiterungsprinzips [Sch73, Sch75d, Sch75e]. Erzeugen [Lys76]. escalation [Pie70]. Esseen [FGL17a, LZW23, Pad96, PV98, WYLW22, WWLH19, Alb84, MS80a, Pad89, Ros89, WHVW23]. Esseen-type [WYLW22]. Essential [Rüs83]. essentially [Gno77, KK81]. establishing [Jac91]. Esti [Cau99]. Estimability [AS79, BB74, Pin74, BK76, DS78, Sch76a, Sch76a]. estimable [Jas81, Mat87a]. estimate [Ano87j, AdUÁS15, BCMR90, DW93, Foc84, Hos88b, KZ13b, Lie82, Liu92, Yat19, Zie85, dUÁ00]. Estimated [Lai00, AK04, NN82, Rod89]. Estimates [Bun93, BC99, Fal95, GLVV98, Mam95, Páz87, SL95, BZ14, Bel74, Bel79a, BGAL01, BBR05, BLO03a, BW81, BCP99, Bun05, BJ05, BDF^{+}18, Che12, Col85a, CS89, Dev89, Dro93, EW75b, Fei72, FR86, GLR23, Gil78, HW88, Hof79b, Ker84a, Ker84b, Koc74, LT13, Lit14, MP96, MW80, Nad80, Nät72, Oko16, Pad81, Pad84, Ryc93b, Ryc93a, Sch77b, Tau84, TMN00, WK85, Wit93]. estimateurs [Pon86]. Estimating [AS83, BS10, BAO12, GG91, GPK23, Hol91, Höp98, Huz18, JT22, KS95, LH22, MSW03, NK23, PK18, Ryd95, ST23, SZ97, TPSR17, Ber00, BN80, DMP10, Dro87, Fra77, HBCL14, Kha15, Kle77a, NJKK16, PCKMK17, Rus91b, Tar88, Waf88, Zha12, Düc75a]. Estimation [AAJ97, AG23, AL99, Arn95, AS98, Atk75, BFF00, BN00a, BGP95, BRKO99, BDG95, BD17, Bil97, BG11, BS75, BDP99, Cav97, CFS17, CG95, CPS98, CH91, CCH87, Csö88, Dac11, Das87, Das96, DDD19, DSG91, Dry75, FZF97, Fer99, Fik88b, GL94, GCL16, Gno79, GM88, GM91, GCB94, GN00, GJ19, GS82, GG90, GON95a, GS04, HK87, HHJ95, HS17a, IRV98, ILRMZ15, JO95, JC22, JD20, JS97, JRM99, Jon76a, KS10, KN98, KE96, Koz86, KM82, LSS12a, Lav95, LUK13, LUK14, LP97, Lik85, LPC97, LMS92, LTG97, LH15, LZW18, LBSM00, Mag95, MV18, Mar97, MD03, May11, Mi96, Mic09, MS94, MKN99, Mou02, MJA02, NK13, NDDDG16, NFP18, Nus76a, OKW86, PSCC95, PAW05, Phe96, PP92, Pon02, Pon04, Pon06, Qin15, Rao85, RT85, RG09, RV97]. Estimation [Rüs90, Sch82c, SLS99, SKCE19, SKV18, Sta88, ST95, Ste95, Tan91, TM94, Try88, Try91a, Web06, WHG87, WW21, WHT03, Yao00, ZWL14, ZWYG14, ZWFZ22, AMP21, AMSV12, AK03, AHPS84, Ano76, Arn92a, AG09, Ars16, AÖ23, AB12, ASM83, ALS11, ADF21, ABD93, AM15, BA08, BK79a, BM12, BC23, BN81, BS85, BC14, BPV11, BB22b, BR90, Bat10, BA89, Bed04, BP18, BBL20, BO01, BS07a, BB91, BCCF14, BG16, BDS22, BH86, BCDD11, BRF20, BCN15, BR15, BLNGB05, Bun73b, Bun76a, BHSW77, Bun81b, Bun81a, BHS76, Bur08, CXW20, Cao01, Çet21, CS06, CS10, CCD15, CKS22, CM90, CF89, CFC10, DI11, DM09, DLMR13, Dew73, Dör15, Dry72, Dry77, Dry88, DZX18, DT16, DS91, EL79, EMF19, FL03, Fah90, FC89b, FGL17a, Fav14]. estimation [FSV13, FdHP03, FS02, FD72, FGK19, Fra79, Fra82, FM01, FAP10, GGL15, GZ23, GC90, GB17, GV19, GW84, GP19, Glo01, GKZ77, GKTZ92, GSTUM02, GGS14a, GCF04, Gov11, Gro82, Gün83, Gun88b, Gun88a, GP07, HKT22, HOCC13, Hen92, HEHM13, HLsS01, HPN22, HY10, HLM21, Hub77, Hus88, HL13, IV90a, IV02, IF03, Iga20, JIP10, JRCB22, JRM04, JR12, JRT21, Jur75, KYR89a, KYR89b, KP09, KS86, KZ78, Kle78, Kle80, Kob87, KFPG21, Koh12, Kon87, KT89, KMM15, Koz85, KZ15, KG81a, Kum89, Kut12, Kut84, KP94, Läu70, Läu71a, Läu75a, LdCIP18, Lie92, LQZF17, LW12, LLV16, LBR23, LY20, LJ22, Log87, LGL^{+}17, Mag88, MZ18, MMRA07, MM21, Mat83, MNvRS10, MCNO19, Mau89, Mei06, MP23, Mic87, MR12, Möh87b]. estimation [Müh80, Mus77, NS10a, NS22, Nät84b, NJKK16, NL05, NTJ09, NW12, Nov10, Ohs81, ÖA21, PKP21, Pet73, Pet86, PS10, Phe89, Pil81, PDG13, Puk81a, QJ08, Raf88, RH80, Rév77, Röd87b, Röd04, RZ11, RVB18, Ron76, RH20, Rud76, SY81, SS15a, SV89, Sch91b, ST83, Sch76c, Sch79a, Sch13, SM08, Sen11, SP08, Sha21, She88, sS14, SAEAEA13, SL23, Sta90, SW89b, SYW22, TE15, TPL15, TPL19, TW05, Tan14, TSO07, THN13, Tho80, Thr79, Tou75, TS87, TY12, Tru18, Try86, Try91b, TT10b, Tua81, UI13, VI90, Wag83, WY04, WHXC16, WH20, WK13, Wil08, Win81, WWC^{+}10, WK15, WW22b, WB05, YLL16, YLY17, YLWL22, Zez93, ZY07, ZR10, Zha13, ZZW23, ZT15, ZL13, Zie88b, ZL03, ZZWG09, ZL17, Zwa85, dUÁV17, vdVvdL03, Lyt77]. Estimation-Optimal [HK87]. Estimation-Theory [LMS92]. estimations [ASFKV08, BFR84, Düc75b, Nku23, YAA22]. Estimator [AK94, BC02, Ber98, FR98, Fur97, Gol91, Hof95, Hof97, JS97, Tak86, VW96, Zie88a, Zie00, vR97a, AK85, AAR18, Als93, ARN12, AK20, BBKP19, BL82, Bel80, BNS15, Bha88, Blo03b, Bun85, Car84, CFR89, CD09, Che21, CLWK20, CD08, Ciu12, Cor17, DM22, DH81, GVV17, GCF04, GS14, Hof77b, Hof80, JSV07, JBC22, JSW22, JS82, KZ20, KK19b, KB85, Kor84, Kra86, Lät74, Läu75b, Lau85a, Läu89, LY11, Mal21, MY11, MW14, dDNN19, Ouy92, Özk09, ÖA16, PH16a, Pil86, QLT17, Rao83, Rie89c, Ryc87, STY04, Sha83, She03, SW16, TR78, Waf88, WZXH15, WC21, WYLW22, WS22, WY13, Wu16, WA22, XWJ14, YW12, ZFA20, ZZW^{+}19, Zie01, Zin81, ZL03, Zwa83]. Estimators [Bel91, BT97, BB00, Bun99, Cro96, GH89, GKP96, GD96, GM94, GC95, GM99, GM97, GON95b, HHS97, IR91, JO97, Jör99, Kun97, Lie98, LPC97, Miz96, Mol96, Neu92, Neu94, Ruy98, SS94, TVL96, VN98, Als88, BC06, BM09, BP16, BMMP16, BM01, Bel79b, Bel83b, BTT11, BTV21, BKS23, Bun75c, Bun75d, Bun77b, BW80, BM84, Cad01, CXH14, Car19, CFR89, CY12, CHL09, Col03, DSK85, DdMT20, DGL16, Fik88a, Fis09, Flo93, FL10, GLZW12, GJV16, GS88, GKZ77, GB78, GKM18, GK93, Hab89, Haj84, Hal91, Hei88, Hof77a, Hof79a, Hof85, Hös78, Hos88a, Hu10, Hue86, IM14, IZ83, JC07, Jur80, KK19a, Kie17, KP74a, KP74b, Kle75a, Kle75b, KZ87, KP07, KJ14, KG81a, KG81b, LKM14, LLBM^{+}11, Lie90a, LY93, Lis90, LGW14, LBSM02, LMV17]. estimators [MP10, Mej78, MWZ15, Mie91, Moh86, Moj07, MA86, Müh84, MB21, Mül85, MSW13, MSCC16, Nät75a, Nät75b, NN23a, NN23b, Nku12, Nus77a, Nus77b, Nus79, Ohs83, Ohy13, ÖCA15, Páz90a, Páz90b, Páz90c, PN90, Pho87, PW15, Pon86, RY86, RH13, Rüs83, SK92, Sah87, SLL15, SS20a, Sch90a, Sch75b, Sch75c, SY22, Sha91, SS04, SC90, SYW22, Tab13, Tak09, TB78, TSC20, TC13, VPR21, Von91, Wan15, Waw81, Wer85, WWLH19, XWY^{+}20, XZZ11, Xie09, YFZ11, YLY18, Zäh14, ZK16, ZRAN17, Zuo01]. ethical [BB07]. EV [MY11, MWZ15, Wan15, XWY^{+}20]. evaluation [Sch87a]. event [Kin75]. events [BGL15, KK81, Sch87a]. Evidence [Sta20, BR19]. Exact [Cau99, CM05, CBC12, DS78, IM14, Imh00, IR05, KS96, MK19, MK83, Nad06, Nät85, Saw81, Sha18, pZ15, ZBF^{+}20, BMK22, BN78, Hof72, JBC22, Kat89, Nag78, Páz90a, Páz90b, PGT08, WBL06, ZZ05, Zvá79]. exakte [Hof72]. example [DB92, Jen90, Kou11]. examples [Bel79a]. Exceedance [AKN05, FB22, Kat89, SB17]. exceedances [BT08]. excess [Sch87b]. exchangeability [CGM19]. exchangeable [RFK18]. exciting [GPS14]. Existence [BB00, Kal76, Mül85, Ano72a, Ano73c, BZ14, DB74, DSK85, Foc84, GAJB13, Hei92, Jas81, RW84, Sch81, ZN14, Zie85, Trö75]. Existenz [DB74, Foc84, Trö75]. Existenz- [Trö75]. expansion [IZ83, Sch75e, Son20, WG92, Zwa83]. Expansions [Bed85a, STY04, Hal91, WN12a, XG07]. Expectation [SN81, BN80, Kha04, Kle78, KB85, Nog20, Ryc93b, Ryc21]. expectations [BB93, Bie06, Bie08, GR06, Gor20, Jen90, Oko16, RR04, Ryc09, TA11]. expected [BG19b, Exn76, Gor14, Gor19, Los23, Mau89, Moe84, Möh87b, Pap16, SP08]. expectiles [Chr23]. Experiment [Ano74, Tit80]. Experimental [BK71, CF95a, HR80, Läu74, RH80, BN78, Ban80, BNP87, BM01, Böh81a, Böh81b, BN80, GH83, HS91, HDC20, LFGR11, Nag78, Nät75a, Nät79, Páz80a, Sch77a, Wer78, Wit78a, NN82, Wit76b]. experimentellen [Wer78]. Experiments [BN00a, MPV00, Mun98, Ste98, Bun76a, BS06, CAK13, Els22, GGS14b, MSCC16, Sch82a, TN08, VYD83, Yat17]. expert [Bic18]. Explanatory [BC99, Dav94]. Explicit [VPR21, Hu10, Sch76b, Rei86]. explicite [Kle80]. explizite [Sch76b]. Exploitation [Ros71]. Exponent [Ber98, Zie00]. Exponential [Ach95, AHJ99, ACS93, BB00, GM92, Han96, Koz02, Küc82a, Küc82b, KM96, Mag95, MS94, MK95, MPV00, SZ97, Tur90, Win17, AMB11, AG09, BCKS01, BHN15, BBK12, BZ15, BS21, CS10, CBC12, Deb77, DLO22, FB16, Fer07a, Fer07b, FV06, GF09, Gün80, Gup79b, GK09, GPK23, HAMAZ96, Hof81, HS82, Hof85, IF03, JK79, JK13, JR12, Kin75, KP07, Kum13, Led78b, LCMA16, Lin09, LBSM02, LMR14, Mag88, MR73b, Mel78, MK19, MPPV04, Nad06, NH11, NNM14, NN91, NAB13, NK17, PAW05, PK18, PKP21, PGT05, Pog15, Raq04, Ras85, Röd77, Ros72, ST23, Saw81, SD86, Sen11, SYM16, Sta90, SD17, SKPK21, TPSR17, VK20, Wan14, WFK82, XCJH19, YDZ13, YH15, ZB11, ZQB16, ZBF^{+}20, dSL22]. exponential-geometric [LMR14]. Exponential-Type [Mag95, Röd77]. exponential-Weibull [FB16]. Exponentiality [BM02, HM02, JGMO20, TJZ89b, BMK22, Ger78b, Ger87, JE90, Kla03]. exponentially [Aui90, Mau89, SP08]. exponentially-distributed [Aui90]. Exponentialverteilte [Aui90]. exponentialverteilter [Mau89]. Exponentiated [Nad05, BKHB22, GAJB13, LMR14, Pog15, WLT^{+}22]. exponentielle [Gün80]. exponentiellen [Kin75]. Expressions [WN13a, Nad10, TSC20]. Extended [Bun93, BB22b, BL15, CLO13, Hu10, LC13, OB19, TB01, WZXH15, WZW21, Wen76b]. Extending [NCJ23]. extension [Bed86, CHA15b, Gun88b, Gun88a, LCMA16, MA20, MFBG14a, NH11, Pré74a, Sch70b, Sch73, Sch75d, Sch75f, VWL^{+}09]. Extensions [GK72, Ehl75, Ehl76, NRD06]. external [GCF04, KFH16]. extra [Luc05]. extra-Poisson [Luc05]. extrapolation [And80]. extremal [AK75, Ehl75, Ehl76, MESS70]. Extremalaufgaben [Ehl75, Ehl76]. Extremalproblemen [MESS70]. extreme [GJ19, JW04b, JB83, MR75, MGM20, RD21, Tor17, ZZQ16]. Extremes [Tia91, FLL16, FLL18, FW20, FZLB22, FV06, TDM13, WW21]. Extremum [Tho84]. extropies [MH22]. extropy [CGS23, HKT22]. Exzeß [Sch87b].

Fabian [Kob87]. Factor [VP04, Ano19, GCT07, LaM17, LYB19, MV10, ROM15, TT22, Van88, WS13, Zha23]. Factor-based [VP04]. factorial [CAK13, Hei88]. factorials [QCO13, WQO22]. factorisation [BS82]. factorization [Ros71]. factors [Bic23, Foc84, GGS14b, May11, PCKMK17, ZYZ21]. fail [BKHB22, HKK23]. fail-safe [BKHB22, HKK23]. Failure [Ger89, RT85, AFSS22, AS83, BSB13, Bie16, BS18, CF22, CF78, CS73, Ger87, KYR89a, KYR89b, Kum92, LC17, LMR14, LCY15, SKD21, TJZ89a, TDY08, WHXC16, WHT03]. failure-censored [WHT03]. failures [AA12, CGM19, Jas22]. Faktorenmodell [Foc84]. Faktorisierung [Ros71]. Fall [Pré74a]. Falle [Fäh78, Fäh81]. Familie [NN82]. Families [BB00, MPV00, Vän98b, AMA14, ACS05, BB22a, Bie06, Bie08, Dan03, GVV13, KP07, Küc82a, Küc82b, Led78b, LBSM02, Pel11, RLAV13, Saw81, Wan14, WN14a]. Family [ACS93, Fak96, TA90, ABP18, BL15, BBK12, BZ15, Cey20, Gen13, Hof81, HS82, Hof85, Hus88, JK79, JW04a, Lin09, Luc05, Man22a, NL03, Nag78, NN82, Ohy13, PGGF15, PGT05, RN18, Sta88, WN13c]. Faserprozessen [Nag83]. fast [Bar87, FGK19]. Faster [ZZLL20]. fatigue [CLO13]. faulty [Düc75b]. Favourable [Bed85a, Haf82a, Haf82b]. Fay [MMP16]. Feasible [And76b]. features [Páz80b]. Fehlerabschätzungen [Pad81]. fehlerbehafteten [Düc75b]. Fehlerhafter [Düc75a]. Fehlerwahrscheinlichkeiten [Kin75]. Felder [Hoy79]. Feller [RR02].

Fenchel [EN74a]. Fertigungsbäume [Boe72]. fertility [TN08]. FGM [SKCE19]. fiber [Nag83]. fibre [MS80b, Mec81, Ohs81, SMP80, Sto83b, Sto84]. fibres [Zäh84]. fidis [Góm18]. fiducial [Are12, Bun75a, Hof79b]. Fiducialize [Bic23]. Fiduzialkonzepts [Hof79b]. Field [GH02, GH01, Hei86, Sär95, TN08]. Fields [BPR00, IRV98, AOV19, Dam76, Hoy79, LBR23, Nät84b, Zan14]. Fiksel [Bil97]. filter [KP94]. filtered [NTJ09]. filtering [Saw81]. final [Ano73c, Ano81e]. finalnikh [Ano73c]. financial [HS17b]. findings [Ano73f, Tam76]. finish [HGLS83c]. Finite [Arn95, Fal95, LP97, Mel94, Pla94, Ras85, Zuo01, AA12, AHJN01, Blo07, BRZ93, Cig13, CM90, Dro93, HTX23, Kie17, Küc78, KS79, LM79, Led78b, Liu11, Los23, OB19, Pan84, Pol08, PV85, Rie89c, Röd71, Rya14, Ryc04, SV10, Thr77, WN12b, Sch90b]. Finite-dimensional [Sch90b]. finite-sample [Dro93, Rie89c]. finite-valued [Rya14]. First [And76b, Atk75, CLS14, FL10, GGHS03, Gün80, WHT03]. first-order [CLS14, FL10, GGHS03]. fish [NJKK16]. Fisher [AA03, AA08, AAB12, ALS11, ABD93, BC12, Hol91, KM78, PBK11, SPTB18, WN12a]. Fit [HM02, LLK98, MO03, Neu79, AWJF22, Ara93, Aui90, BL78, Bed81, BKN70, CC01, Cap19, CK06, DLO22, ELV22, Geb84, Ger79, Ger80, GF09, GT13, Hei91, Hei18, HK15, IJL90, KW15, KPH^{+}16, LM23, OJM15, PSB23, Sch91a, Sha86, Sol19, Sta20]. Fitting [Aka82, Gat94, GPS14, MP11, GR17, MFTFEO22, PA14]. Fixed [AEW94, CR99, HDA99, Tij76, AAB12, Duv17, LC19, LST18, RM17, RM16, Sei79, Wei72, Wei74, YC15]. fixed-point [Duv17]. Fixed-Width [CR99]. flat [Páz90b, Páz91a]. flexible [MFTFEO22, PGGF15, RLAV13]. fluctuating [Goe75]. fluctuation [CLCP17]. fluctuations [SS79]. Flugzeugeinsatzes [BNS73]. focused [PR22]. folded [Kim16]. foldover [Els17]. Folgen [Gün77, Müh80, Wit78b, Wit81]. follow [Els22]. follow-up [Els22]. following [Cap73, KHN20, LC17, MM21]. forecast [Bel74]. forecasts [GPS14, Tou70a]. Foreword [Bun70]. forgotten [Bod75]. Form [AC09, NS03, Bla88, Kra86, Läu75b, LMS06, NN23a, NN23b, Net90, SSG76]. Form-invariance [AC09]. Form-invariant [NS03]. formation [Wer78]. formen [Ric84]. Forms [EW99, GN98, BS88, KT83, LR17, PHS09, Ric84, Wie92, YY07]. Formula [Las91, Gui86, GJ73, WN13b]. formulae [Kle80, Rei86, Sto84, Sto90, VCO09]. Formulas [HS79, MS80b, Mec81, Páz87, SMP80, Bor73, PMS81]. formulation [PP94]. foundation [Ban80, Cen78]. Four [SS94, ASSH13, EP17, KAM17]. four-mixture [ASSH13]. Fourier [TKTM15, Win81]. Fourth [Dre98, PKN72, War85b, HM89]. fourth-order [HM89]. fractal [Sto94]. fractional [AM15, EMF19, KMM15, NS22, NDDDG16, XZZ11]. frailties [Eco13]. frailty [FPPRG16, HX20]. framework [BDP16, Krz05]. Frechet [PV85, Sch22, Wal88]. Free [Zie88a, Ahm74, BF85, DH77, FT79, GF09, MKM21, SMLL18]. freeze [GJK^{+}16]. French [Cap73, Pon86, VO75]. frequencies [Ara93, Wit93]. frequency [BVP06, HS17b, Mas19, NK13, NK23, NJKK16, Sti78, WW22b]. frequency-domain [Sti78]. frequentist [CKM09, CM15a, GVGP13, PMO14]. Frequenzbereichsanalyse [Sti78]. fringe [Lud73]. frontier [BZ10, WZW21]. Frontiers [Hec86]. Frühjahrstagung [Kle72]. Fubini [Tho78]. Fubini-type [Tho78]. fulfillment [Ruk07]. Full [sS15c, BK77]. Fuller [Wes14, Güv14]. Fully [Neu94]. Function [CPS98, CB02b, GN00, Jör99, Mag95, MS98, Phe96, VV96, Ano76, Ant84a, BA08, Ban70a, BQSR20, Bar94, BSS18, BL82, BBL20, BT87, Ber00, BN16, BCP99, BRF20, CXW20, CXH14, Cap73, CS06, Che21, De 22, DM09, DZ22, Dik01, Fäh78, Fäh81, Fer07a, GP19, GJLT15, GAB85, Gup79b, Haj84, Hav84, Hei91, HHM14, Hof81, HHM15, HL13, JIP10, JSV07, Jas81, JD20, JR06, Joz14, KP09, Kaw15, Kha04, Kor84, Kow82, Laz16, LdCIP18, Lie82, LY93, LQ12, MZ18, MB18, Man22a, MYK18, NSG07, NPS12, NP12, NSK12, NJKK16, NDDDG16, Nov10, ÖCA15, PKP21, PH16a, Phe89, Pol85, Raf88, Rei84, Rév77, Ric76b, Ric84, Rüs81, SV89, SY22, SP08, Sol19, Sta88, ST87, Szc79, Tam74b, Tan11, Tar88, TB01, TPSR17, Vís11]. function [WS22, Wei70b, WN13b, Wyn76, XWJ14, YL17, ZW10, ZZWG09]. Functional [DLMR13, FY03, GGL15, Mou02, Nus76a, RF03, RM16, ASFKV08, AFV15, BDS22, Bit75, BR15, CXW20, DHN20, Del09, EL79, FX16, FSV13, FP92, Fin94, Guo20, HVRI20, Hös78, JK76, KHN16, KCL16, KK19b, KK11, KY16, LW12, LLV16, LV18, Nus77b, Nus79, PS10, Qin15, QK17, Ras06, Rod93, SA23, Tan14, VS81, ZZLL20, ZWFZ22]. functionals [Ano81e, BLO03a, CC01, CF78, EN74b, Góm18, HR93]. Functions [EZ92, Gat94, GCB94, Gup95, Man72, RT85, RF03, Tan91, Ver92, Ano72a, Azp73, BB93, Bod75, BB72, Bun73a, BS80, Bun82c, BM84, DJ09, DS78, DP17, Ger73, GJHN18, Gro82, Gup82, Gup05, HP08, Hei18, HS91, HS82, HBCL14, KPH^{+}16, Kow77, LH91, MR73b, MSCC16, NN82, NSK12, Nät75b, Nät79, NRD08, PGGF15, Pré74a, Röd87b, SIM21, Sto72b, Sty75, Tay83, Van85, Wal88, WLD18, Win17, Wis77, WN12b, Zäh14, ZN14, Zvá79]. Fundamental [AN98, KS79, NK23, RT13]. Fundamentalidentität [KS79]. fundamentals [BF85]. funktiayami [Ano72a]. funktionellen [JK76]. funktsii [Ano76]. Further [LCK21, PYHT01, Sto84, TP97, YDZ13, Tro93]. fusion [BR19]. Future [AHJ99, AA12, Ahm11, BB22a, LV18, NAHJ03]. Fuzzy [AS98, EA90, Nät97, BN81, WY04].

G [Ber74, Dew75, Fre74, Hec86, KSS76, RS74, Ros71, SD86, Sto74, Sto76b, Ber74, Dew75, Fle76, KSS76, RL83, Ros71]. G/G/1 [Dew75, Ros71, Ber74, Dew75, Ros71]. G/G/s [KSS76]. G/GI/m [RL83]. G/M/2 [Fle76]. gains [Bak73]. Galton [EW75b, EW75b]. game [Ano72a, Ano73d, Ano76, Ber17, Tho78]. games [Ano72a, Ano75c, Bor70, Rie87]. Gamma [Koc74, VV96, AMA14, AE21, Bar87, CLZ15, CM05, CTL14, CCMdC13, FW20, FH01, HOCC13, KT89, Nad08c, NS10a, NN23a, NN23b, Pro89, RAM^{+}17, Röd76, SN14, TPL15, TPL19, Väd77, VFS^{+}20, WN13b, Kaw15].

Gamma- [Koc74]. Gammaverteilung [Röd76]. ganzzahlige [Har75a, Har75b]. ganzzahligen [Pie70, Pie73, Pie75a, Pie75b, Pie76]. Ganzzahliger [HT71, Bär72, HP73, Ric76b]. Ganzzahligkeitsbedingung [Ric76b]. GARCH [GMLS16, GML18, HB09, Xie09]. garrote [GVV17]. Gauss [Bak84, DdMT20, OKW86, PH16a, STS85]. Gaussian [BLSC09, BA89, BW14, BZ13, CG04, DLO22, DS91, FZLB22, GK99, HT15, HK15, IV90a, IRV98, IV02, KCL16, LH22, MP10, MYK18, Müh80, Müh84, RO98, Ric86, SA23, TSO07, Wit78b, Wit81, Wit83b]. Gaussianity [Hei88]. Gaußsche [Müh80]. Gauss’scher [Ric86, Wit81, Wit78b]. Gebrauch [Ron76]. GEE [Zha12]. gegebener [Ban70a]. Geleitwort [Bun70]. Gemischt [HT71, Pie76, Har75a]. gemischt-ganzzahligen [Pie76]. Gemischt-Ganzzahliger [HT71]. gemischtganzzahlige [Ste76]. gemischtganzzahliger [Ric76a]. Gene [Cie98]. General [CH08, GM99, JO95, JO97, Kam22, Kow77, LH13, Mar97, NS10a, PRS95, UR81, WHG87, AA12, BK79a, BR90, Bei76, Bel79a, Böt20, CFR89, CAK13, CGS23, DWHH19, DK21, DR03, Dry75, Ehl75, Ehl76, GAJB13, Gui86, GK72, HK00, Hav88, HR80, Hol75, Hös74, HCW16, Kal76, KM82, Kum89, LR17, LC17, LGZL14, LLL17a, Log87, LST18, Mad76, Mat83, MS80b, MPPV04, Näg75, OKW86, PA14, PMO14, PKP21, PB10, Ped86, Pet86, Pol86, PD91, Sch75a, Sch75d, Sch75e, Sch76d, Sei79, Tou70b, WLT^{+}22, Waw81, Zan14, Zan15, Zie85, dSL22, Küc82a]. generalisation [LPS90]. generalization [FD72, Hös77, Pie74b, PS77, SH09, Thä75, Sie74]. Generalizations [Pré73, Sch22]. Generalized [Are05, AV12, BFF00, BMMP16, BBK12, BDS22, BM94, DM95, ESSB93, EW99, FC89a, GSB06, GON95a, Hös78, IKR00, Kra79, PGT02, Raq04, RA06, SR02a, TX17, Tro00, WHXC16, ZZL^{+}14, AAFM07, Ahm11, Are12, AG09, AA11, AA13, BKT14, BT08, BBK14, BMK16, BK17, Bed04, Bie06, Bie08, Bie16, BS18, Bök85, Bon88, Bre05, BJ05, BDF^{+}18, BC12, CMLBSM14, Çet21, CTL14, CLLY22, CRCN14, CCMdC13, Cra06, DT16, DF15, Dup17, EBQY17, Fah87, Fah90, Fer07a, FS02, GLZW12, Gen13, GK17, GMCT18, Gor14, GS17, Gup79a, HOCC13, Hof81, HSG13, HWL17, Hus88, JR86, JRCB22, JG85, JS81, KC01, KD16, KP07, Krz05, KA13, KR07, LLC16, Lem13, Lis90, Mag88, Mat81b, Mic09, Nad08b, NAB13, NK17, Pad84, PR22, RAM^{+}17, RN18, Rod93]. generalized [RH13, SPTB18, SYM16, Sha21, Sto79, TA21, TH04, TK22, VK20, WG21, XW10, XG07, YC15, YYVH16, GSB11, Pfu78]. Generalized- [PGT02]. generalized-order [Ahm11]. Generalizing [HA02, FK12]. generate [BSL13]. generated [AMA14, BN14, RPZB18, WN14b]. generating [Gup82, HS82, HHM15, Lys76]. Generation [tFYK01, KFL97, Bar87, PYHT01, Väd77]. generic [Bär76]. genes [Pon93]. genetic [DB92, GPS14, Gno79, PVW74]. genetics [GM88]. Genetische [PVW74]. genomic [RD21]. genotyping [Rab15, RD21]. Geometric [McC03, McC06, Tit80, AP07, BCD11, CSO13, Deb77, Jas18, Kun17, KM21, LBSM00, LMR14, Nož71, Páz90c, VCO09]. geometrically [BCCF14, Win17]. geometrischen [Deb77]. geometrischer [Nož71]. geometry [McC05, Puk81a, Sin05]. Georg [Bod75]. gerichteten [Pie72b]. germ [Hei92, HP08]. germ-grain [Hei92, HP08]. German [Atk75, Aui90, Ban70a, BK71, BM71, BDP^{+}74, Ban70b, BWR^{+}71, Bär72, Bär76, BNS73, BTS71, Bei74, BKN70, Bel74, Bel79a, Ber74, Ber76, BGB^{+}72, Boe72, Bor73, Bun70, Bun74, BHW75, BG74b, Bur70, Dam76, Deb77, Deu72, DB74, Dew73, Dew75, Dew76, Döh81, DGB^{+}73, Düc75a, DGG^{+}75, Düc75b, DGH^{+}73, Ehl75, Ehl76, ES71, ES73, EW73, ES74, EN74a, EN74b, EG74, Els75, EN75, EW75a, EGV^{+}73, EvWE^{+}75, EW75b, Exn76, Fäh78, Fäh81, FM72, Fei72, FN77, Fle71, Foc71, FD72, FGR73, Foc73, FG75, Foc75a, Foc75b, Foc84, FGR^{+}76, Fra77, FR86, Fre74, GH83, Gil78, GV79, Gir71, Gir74, Gla76, GRV75, GR76, Got74, GM78, Gre87, Gri80, GR72, Gru71, GFFS74, Gud70, GT70, Gün77, Gün80, Gus71, HK78, Har75a, Har75b, HWK^{+}72, Hen89]. German [Hof79b, HT71, Hof72, HP73, HW70b, HW70a, Hol70b, HL70, Hol70a, HW71, HW72, Hol73a, Hol73b, HW74, Hol75, Hol76, Hös76, Hoy79, JN75, JK76, Jür84, KS71, KNS72, Kin75, Kin76, KPAF75, KJM^{+}72, Kle72, Koc74, Köc75, Köc76, KS76a, KJKR76, Küc78, KS79, Läu70, Läu71a, Läu71b, Läu71c, LSP^{+}73, Lie72, Lip75a, Lip75b, Lip76, Lom74, Lom70, Lud73, Lys76, Lyt77, Man72, MSW^{+}72, Mat75, Mau89, Mec72, MSK^{+}76, Mel76, MESS70, Mön72, Müh80, MW80, MSS^{+}71, Mül74, NN82, Nag83, Näg75, Nät72, Neh75, Neu73, Nož70, Nož71, Nož72, Pad81, PVW74, PW76, Pfl80, Pfu78, Pie70, Pie71, Pie72b, Pie72a, Pie73, Pie74a, Pie74b, Pie75a, Pie75b, Pie76, PT83, PKN72, Pré74a, Ric75, Ric76a, Ric76b, Ric76c, Ric84, Ric86, Röd70, Röd71]. German [Röd76, Ron76, RWH^{+}72, RS74, Ros71, RW79, Sak71, SW85, Sch87a, Sch87b, Sch89a, Sch76a, Sch76b, Sch75a, Sch90b, Sch70a, Sch70b, SHP^{+}71, SND^{+}72, Sch73, Sch74a, Sch70c, Sch75d, Sch75e, Sch77b, ST75, Sch75f, Seb76, SSS70, SSS74, SS74, SSG76, Sie74, Som71, Ste76, Sti78, Sto71, Sto72a, Sto72b, Sto73, Tam74b, Tam74a, Tam76, THN^{+}76, Tau84, Ter72, Thä75, Thä76, TSL^{+}76, Tit70, Tou70a, Tou70b, Trö75, Ull76a, Ull76b, VBP^{+}76, Wag83, Web79, Wei70a, Wei70b, Wei71, WHS^{+}76, Wen76b, Wen76a, Wer78, WRME72, Wit76a, Wit76b, Wit78a, Wit78b, Wit81, Wol84, Wol85, Wol75, Zäh74, Zin81, vW72a, vW72b, vW74]. gesamten [Deu72]. geschlossene [SSG76]. Gestaltsparameter [Röd76]. gestauchten [NN82]. gesteuertes [Gre87]. gewichtete [Wol85]. Gewinnfunktion [Fäh78, Fäh81]. gewisser [Fei72]. GI [KS76b, Kre80, RL83, RS74, KS76b, RS74, SD86, Sto74]. GI/G/1 [RS74, RS74, SD86]. GI/G/s [Sto74]. GI/GI/1 [KS76b]. Gibbs [Bil97, Sär95]. Gibbsian [Fik88b, Tak86]. GINAR [WW22a]. given [Ban70a, Bic18, BN00b, BB72, Dev89, FK12, Mau89, Nog20, RD80, WN13c, YYVH16, Zin81]. GLARMA [GLLO^{+}22]. Glättung [Gün80]. Gleichung [Sch90b]. Gleichungen [Deu72, Pie74a, Tou70a]. GLM [SK14a]. Global [GK95, BS07b, Hös86, PA14, Pie75b]. Globale [Pie75b]. GLS [Wes14]. GMM [KZ20]. Gomory [Pie70, Har75a, Har75b, Pie70, Pie73, Pie74a, Pie74b]. Gomory-Schnitte [Pie74a]. Gomoryschen [Pie73, Pie74b]. Gompertz [MGM20, WHT03].

Goodness [Ara93, CC01, DLO22, Hei91, HK15, KW15, LLK98, MO03, Nag78, Neu79, OJM15, Sch91a, AWJF22, Aui90, Cap19, CK06, Geb84, Ger79, Ger80, GV79, GF09, GT13, Hei18, KPH^{+}16, LM23, PSB23, Sha86, Sol19, Sta20]. goodness- [Geb84]. Goodness-of-Fit [MO03, CC01, DLO22, Hei91, HK15, KW15, OJM15, AWJF22, Cap19, CK06, Ger79, Ger80, GF09, GT13, Hei18, KPH^{+}16, LM23, PSB23, Sha86]. Goodness-of-fit-tests [Sch91a]. Gordan [FG75]. gordanscher [FG75]. gOSs [Gor20]. governed [Bit75]. grade [KL82]. gradient [FGR73, Foc75a, GBS20, GB23, Lem13, NRS07]. Gradientenverfahren [FGR73]. Gradientenverfahrens [Foc75a]. Gradual [Wan07]. graduation [Eic88]. grain [Hei92, HP08]. Grains [Str95]. granichnikh [Ano81e]. graph [Dew76, Pie72b, Pie72a]. Graphen [Pie72b]. Graphentheorie [Pie72a]. graphic [dUÁV17]. graphical [BCFF88, SA23]. Green [Bre05, Jör99, ZR10, dUÁ00]. Grenzwertsatz [Wol84, Pad81]. Grenzwertsätze [Wol85]. Grenzwertsätzen [Lie72]. grid [ZH16]. grids [HT15, LT16]. großdimensionierter [ST75]. Große [MESS70]. großer [Wol84]. Group [ZX14, Ano73f, Gno78, HH16, JY20, PEB15, Pie73, YT01, YYVH16]. grouped [AC23]. groups [Ciu20, Gov92, GGS14b, TG87]. growing [Lys76, Mik88]. Growth [Wen95, vR97a, BBKP19, Hu10, KvR00, Ras88, SDC19, Von91, Zez93]. Grundlage [Hof79b, PT83]. Grundmodell [Näg75]. gruppe [Ano73f]. GTARCH [GM96]. Guaranteed [GV06]. guided [AX16]. Gumbel [FV06, MJA02, NN91]. Gupta [She03]. Güteeigenschaften [GV79]. GX [KK80].

H [Sch76a]. H. [Dro91a, Hec86, Hof84, Sch76a]. H.-J. [Hec86]. Hadamard [NLP95, PD91]. halbgeordneten [Thä76]. Half [Las91, Foc75b, RD80]. Half-Line [Las91]. half-orderings [Foc75b]. Hamiltonian [CLCP17]. Hammersley [AO07]. Hammerstein [Wol75]. Hammersteinschen [Wol75]. Hannan [Kam22]. Hans [NS10b, Sto93]. Hans-Walter [NS10b]. hard [AX16, HS83, Zha12]. hard-core [HS83]. hard-thresholding [AX16]. hard-to-reach [Zha12]. Hardy [GM88]. harmonic [BS82, ILRMZ15, SL11]. harmonics [NK23]. Harris [BL15]. Hastings [AP07]. Hat-function [Ber00]. having [Ano75c, Bak73, BKHB22, Jac91, MM21]. Hazard [CZZ17, Mi96, BMRPZ16, BN16, Dik01, JMK21, KAR14, NRS07, PK18, PN90, SV89, SW16, YN23]. Hazards [CF98, Csö88, EAA16, FLL18, FW13, Hen93, HLsS01, Sha21]. heaviness [SdSSL15]. heavy [BLNGB05, GJLT15, HPN22, QXY23, SS15a, WW21]. heavy-tailed [BLNGB05, GJLT15, QXY23, SS15a, WW21]. height [KS83a]. Helga [GNS15]. Hellinger [Bed89, Lie86, WK15]. Helmholtz [TKTM15]. help [Lom70]. Henderson [Eic88]. Herleitung [Bor73]. Hermite [Lie98]. Herriot [MMP16]. Heston [BP16]. Heterogeneity [HA02, YS23].

Heterogeneous [Goe76, CZZ19, FB16, FLL16, FZB17, HSBB23, HKK23, Hue86, KHN20, MGM20, YDZ13, YN23, ZB11, ZB14, ZZQ16]. Heteroscedastic [Neu97, Are12, BZ10, Das87, GV19, KZ78, KJ14, MNvRS10, Van91, XWY^{+}20, ZW20]. heteroscedasticity [GS17, LQ12, TC22, Vís11]. heteroscedastie [Bun81b, KP74b]. Heteroskedasticity [Fur97]. Hidden [Ryd95, Wit93]. Hiddenseem [HW74]. Hierarchical [Das96, Cap73, Gno79, KCL16]. hiérarchique [Cap73]. High [Eks99, FR20, KH21, SA23, TA21, VN98, AC23, BJfYZ13, BQSR20, BT04, FYY^{+}23, FdHP03, HH16, HS17b, Jaw13, KMM21, Mas19, PR22, PEB15, QXY23, SX20, ZZW^{+}19, Zha23]. High-dimensional [FR20, KH21, SA23, TA21, AC23, BJfYZ13, BQSR20, HH16, Jaw13, PR22, PEB15, SX20, ZZW^{+}19, Zha23]. high-frequency [Mas19]. Higher [Bla93, Bos88, LO86, Páz93, UI13, DFRLÚF12, LYB19, LH22]. higher-dimensional [DFRLÚF12]. Higher-order [LO86, LH22]. Highly [TMN00]. Hilbert [Wit76a, GRV75, Kle73, Ric86, Röd70, RM16, Ruy98, Wit76a, ZZLL20, Páz80a]. Hilbert-Räume [Wit76a]. Hilbert-Räumen [GRV75]. Hilbert-space [Páz80a]. Hilbert-valued [RM16]. Hilbertraum [Röd70]. Hilberträumen [Ric86]. Hilfe [BNS73, Lom70, Lyt77, Wag83]. Hilfsmittelbeschränkungen [Hof72]. Hill [KK19b]. Histogram [BO01, JO95, BBR05]. histograms [BV01]. hit [Dod83]. hit-and-run [Dod83]. Hochschule [ES74]. Hodges [GD96, Zuo01]. Hoeffding [BGV06, Rao98]. Hoeffding-like [BGV06]. Hold [GLR23]. Hold-out [GLR23]. Homogeneity [GD96, KB00, Pon93, BH89, PG19, TIT08]. homogenen [Gil78, GV79, Küc78, KS79]. homogeneous [ES73, Gil78, GV79, Kir76, Küc78, KS79, RA02, Thä76, WL05]. homogener [ES73, Thä76]. homoscedastic [Bun81b]. honour [BFN17]. Hopf [Ros71, Ros71]. Hopf-Faktorisierung [Ros71]. horizon [Rei86]. Hossain [Tik09]. Hotelling [GK99, Sch76d]. Hotelling- [Sch76d]. HPCs [GZ23]. Hsu [DH77, Puk77, Zie89]. Huber [CXW20]. Hubert [Van85]. humpedness [BT87]. Hunt [Lus87]. Hurst [LZW18]. hybrid [AAB18, BZ14, BSS18, BPK14, Çet21, CBC12, KFPG21, KKG17, PBK11, SPTB18, WG21, ZWFZ22]. hybrid-censored [SPTB18]. hyperbolic [Bit75, Mag88, RA06, Tam74a, TH04]. hyperbolische [Tam74a]. hypergeometric [KP91, SM19, Try91b, WN13b]. hyperplane [Tho84]. hypersphere [FG06]. Hypotheses [BMK16, CH08, Faz08, HA02, MZ96, Ahm74, BJfYZ13, BMP15, BH87, Bre05, BR83, DR15, DJ86, GV81, HK00, Jur75, KM13, Mat81a, MPPV04, Pin75, Que86, Sch91a, Sch79b, Sch82b, SZ86]. Hypothesis [EBTB21, BG19a, BK77, BR19, BH89, DF15, GK72, MYK18, Rya14, Sch76c, Sch76d, Sol19].

I. [Läu75a]. i.i.d [LT13]. i.i.d. [Hab89]. identical [Çet21, KFPG21, RN18, YY07]. identically [BNAA13, Gai23, KZ13b]. Identifiability [BB74, JR86, Sch76a, TSG20, Tia91, Hös86, Sch76a]. Identification [Mül74, Sch89c, BG79, FMM85, Gno77, Gno78, PKM78].

Identifikation [Mül74]. Identities [BB04, SP98b, Bal03]. Identity [Rao98, Hei85, KT12, KD16, KS79, LD21]. IDMRL [NL03]. ierarkhicheskikh [Ano73g]. if [BS07a, Nus77b, WN13c]. IFRA [TZ91]. IFRA-ness [TZ91]. IGFR [Gor19]. ignorable [TSG20, ZL14]. ignorance [BZ15]. igr [Ano75c]. igrami [Ano73d]. igri [Ano72a, Bor70]. igrpax [Ano72a]. ihm [Ros71]. ihre [FD72, Gla76]. II [Wei71, RH80, AAFM07, BGP95, BCKS01, BCD11, BZ14, CBP23, CB02b, DK12, Ehl76, ETD99, GM97, HW07, JA89, KCNB11, KKG17, Küc82b, Läu71c, Led78a, Lip76, MK19, MAS06, Páz90b, PRS95, Rie89b, SP01, Sch75e, SP08, Sin81, SMP80, SKPK21, VK20, VDB13a, VDB13b, Wei71, ZBF^{+}20, vW72b]. iid [AA03]. III [Dom10, Mec81, MK83]. ikh [Ano75b]. illness [PDG13]. illness-death [PDG13]. illuminated [Ano87j]. Ilmenau [ES74]. Im [Fäh78, Dew75, Dew76, Fäh81, Kin75, NN82, Nät72, Pad81, RS74, Sie74, Tou70b, Wit76b, Wit78a]. Image [BC06, CRCN14]. imbedded [KSS76]. imeyushchikh [Ano75c]. Immunes [ZM95]. Impact [BKS23, AG23]. impatient [Sin84]. Implementation [Cse95]. Implications [GK98]. Importance [GVGP13, Lia12]. Impossibility [Nov10]. imprecise [Ryc21]. Improved [BB93, Bun75b, Das96, GFCO04, GON95a, GON95b, LPC97, TSO07, CKS22, Dev89, TPL15, TPL19, WHVW23, ZZWG09]. improvement [Cip81, ZK16]. Improvements [MH97, PSCC95, Hof80, Tau84]. Improving [JS22a]. imputation [HT16, Sin09]. inactivity [Man22b, ZC10]. Inadequacy [Khu95]. inadequate [BB91, BS75, Bun81a, Hos88a, Hos88b, Läu75a, Pet73]. Inadmissibility [JMK21, Hof85]. incidence [DP17, SDC19]. including [KFH16]. inclusion [KKA84]. Incomplete [Yao00, Fäh78, KP15, KM82, MMR11, Sto71, Sto72b]. incompletely [Bun76b]. Increasing [Ger89, BMK22, BN16, BS18, Ger87, GJHN18, Kum92, LF22, LMR14, TJZ89a, ZC10]. increments [DS91, WK85, WFK82]. indefinite [PB21, PHS09, Tam76]. indefiniter [Tam76]. indépendance [VO75]. Independence [Tjø96, YT01, Bla88, BH89, DB10, JH09, Led83, Led84, LCY17, LWL18, Pan84, Röd71, Rod89, SKD21, TG87, TDM13]. Independent [BN00b, GKP96, AB10, Ano87j, BMS20, BNAA13, CM05, DS91, FB16, FW20, Jia07, Kha12, LFGR11, MR73b, Mil86, MV10, Pro89, RW84, WK85, WFK82, Wei70a]. independently [Tho13]. indeterminate [Sch87a]. index [ASFKV08, BGL15, EBQY17, GS11, GCF04, GJLT15, GNYX15, GPK23, HMP01, HPN22, HLM21, LWS21, YLY18, ZWYG14, ZZL^{+}14]. indexed [WWH18, pZ15]. indicator [ZL17]. Indices [Che97, Vän98b, BDP16, CFLDÁ04]. Individual [MKN99, FN77]. individueller [FN77]. induced [Zhe10]. inducing [TK22]. inefficient [Fis09]. Inequalities [Dre98, KT83, Sto83b, Tuy76, AK75, BGV06, Bos93, CF78, Cor17, FGL17b, FG75, Foc75b, Kra86, Läu75b, RS74, Ros89, Thä76, Tuy74, Win17]. Inequality [BPR00, GK95, Pad96, AO07, BDP16, BM84, CFLDÁ04, GK77, Kim07, LCK21, Mec88, Pad89, PV85, dSL22]. inestimability [KS93].

Inference [BKT14, BK17, BT97, CB02b, Dup17, GM97, GV95, Mou01, Nät97, Pel11, PRS95, QXY23, RO98, RLAV13, WLT^{+}22, ZH16, BPvZ04, BSS18, BGL15, BK86, Bun75b, Bun75a, Bun76b, CLCP17, CG04, CT07, CC12, DDLN17, DZ22, Esc84, GJK^{+}16, GP19, GSB06, GSB11, GJLT15, HS83, KMM21, Koo86, LG05, LDF17, LW18, MK19, MAS06, Nog20, Pho87, PW15, RAD07, SGG04, SMW22, sS15c, SYW22, TZ13, TX17, WG21, WW22a, Wie86, WN14a, YLY18, Zha16, ZL14]. Inferences [Ach95, AA11, AA13, AMP21, BZ15, EJ81, LLL17b, RA06, XB17]. infinite [DMR07, DST21, IJZ08, Rei86, RZ11]. Infinitely [SS98, BL11, LH91]. infinitesimal [DDT89]. Infinity [Kre85]. inflated [DDD19, Dup17, GMLS16, GML23, KMSS10, LLC16, TSG20]. inflation [SS21]. Influence [BHM04, JR06, Kal97, MCML15, RF03, Szc79, TWZ06, XW10, BM01, DT16, Düc75a, Koh12, MSCC16, ÖCA15, Seb76, ZT15]. Information [AS98, Bow12, Gat09, GM91, Man72, AA03, Ahm93, AA08, AAB12, Bei75, Bun75b, BHS76, BC12, Düc75a, Els22, Fäh78, GK17, Hen79, KM78, LSS12b, Lie72, Mar19, Mej78, PR22, Pap16, PBK11, PS14, SK89, SPTB18, Sto71, Sto72b, SV10, Tom90, ZMY20, Fäh78, SSS70, SSS74]. Informationen [Sto71, Sto72b, SSS70, SSS74]. Informationstheorie [Lie72]. informative [AAA13, BD13, Bre05, BJ05, RAM^{+}17, Yoo16]. inhibition [SDKM18]. inhomogeneous [Faz08, Faz14, ZW10]. initial [Els22, SW85]. initial-stage [Els22]. Injection [Phe96]. innovations [BH18, De 22, LMS06, NDDDG16, WWLH19]. Input [Fle76, Düc75a, Fre74, Goe75, Sak71, Zäh74]. inputs [Sch79c, Sch81]. insight [Bic18]. inspection [Arn92c, Bei74, Bei75, Men71]. inspections [Krz05]. Inspektionsmodellen [Bei74]. instrumental [Lyt77, SS04]. Instrumentalvariablen [Lyt77]. insufficiency [Döh81]. Insurance [KN98, Nie06]. Integer [Kov74, Bär72, DK21, GMLS16, GML18, GML23, Har75a, Har75b, HT71, HP73, Pie70, Pie73, Pie75a, Pie75b, Pie76, Ric76a, Ric76b, Ros75, Ste76]. integer-valued [DK21, GMLS16, GML18, GML23]. integral [Bit75, LD21, RS74, TG87, Wag83, WN14b, Wol75]. Integralgleichungen [Wag83, Wol75]. integrals [AdUÁS15, Lie86, WN12b, dUÁ00, dUÁRC04]. Integralungleichungen [RS74]. Integrated [Imh00, YS23, And79c, BE12, Glo01, SLL15]. integrating [LWL18]. integrating-up [LWL18]. intensité [Pon86]. Intensities [BRKO99]. Intensity [GCB94, Fre74, Kon87, Pon86, Sak71, Zäh74]. Inter [GM91, Mej78]. Inter- [GM91]. inter-block [Mej78]. interaction [Fik88b, KHN15]. interactions [GGHS03, KD74]. interconnected [Köc75, Köc76]. interdependence [Düc75b]. interest [CKM09]. intermittently [Cha74]. International [ES74]. Internationale [ES74]. Interpenetration [Str95]. interpoint [SM19]. Interpolation [Van85, Van88, Web79]. interpretation [Bun77a, FK90, Tro91]. interquantile [Col91]. Interrupted [Mad76]. intersecting [Sto90]. intersection [Tho84]. intersections [Mec81]. Interval [WWC^{+}10, Ahm11, BM12, CFLDÁ04, Cey20, CLWK20, FC89b, HOCC13, Los23, Mau89, MH22, sSW19, SV18, TDY08]. interval-censored [HOCC13, sSW19]. interval-valued [CFLDÁ04, CLWK20, SV18]. Intervallänge [Mau89]. Intervals [HHJ95, HDA99, Neu97, RL98, Are05, CKM09, Fer07a, FMTZ07, GC04, HJM90, IM14, Moe84, OM11, OB19, PMO14, Sch74b, SV18, Wan14, WN12a, ZZ05]. interview [Gov92]. Intra [GM91, Luc99]. Intra-Block [GM91]. Intra-Cluster [Luc99]. introduction [Ras91]. Invariance [IV90b, Jag80, Ste95, AC09, Bun77a, KJKR76]. Invariant [DJ12, Jaw13, Kle77a, Ouy92, AG23, BB22a, Böt20, HPN22, Kle75a, KZ87, LTV03, LNO14, Lus87, NS03, VI90, WB05, ZN14, JK76]. Invariante [JK76]. Invarianzuntersuchungen [KJKR76]. inventory [Gir74, Gri80, KA88, Kle72, Köc75, Köc76, Pré73, Ric75]. Inverse [GHB02, RO98, BLSC09, BNS15, Cup21, Düc75b, FY03, PW15, She14, WLT^{+}22, ZTJ17, ZHG12]. Inversen [Düc75b]. inverses [DF15, Sto79]. Inverted [vR97b]. invertibility [And77, And78c]. investigation [JK92]. investigations [Ger73]. Invited [CF95a, Páz90c, Sha96a]. involving [ACS08, RDLF03, Seb76]. Ionensorten [FN77]. ions [FN77]. irp [Ano75a]. irreducible [Küc78, KS79]. irreduziblen [Küc78, KS79]. irregular [AAFG23]. Irregularly [Nie17, Wit93]. irrespective [Kum13]. Irrfahrt [KS79, Sie74, Ros71]. isoperimetrlc [Mec88]. Isotonic [KYR89a]. isotropic [Han83]. ispolzovanii [Ano76]. issledovaniya [Ano73a]. issue [BFN17]. issues [CLCP17]. Itô [War85b]. items [Cha15a, CF22, LC17]. Iterated [EMG98, Mik88, Van91]. Iterative [Dau82, Deu72, GR17]. iterativen [Deu72]. Ivanov [Zwa83].

J [Neh75, War85b]. J. [Dro90a, Hec86, Neh75]. jack [Zwa85]. jack-knifing [Zwa85]. Jackknife [Bun93, KDM22, Lav95, Blo03b, DDT89, Pol85, Sha91, dUÁ00, AdUÁS15]. jackknifed [AK20, Wu16]. jackknifing [Rub21]. Jahre [FG75]. Japan [War85b]. Jensen [LCK21]. job [Men73]. job-machine [Men73]. Joint [BA89, SR91, SR94, BCK05, BM23, GZ23, LGL^{+}17, MK19, Nät75a, SS21, SKPK21, YT01, ZBF^{+}20]. July [GNS15]. jump [CHPS13]. jumps [KHNT17, LZW18, RS09, SL23]. just [PJ05].

K. [War85b]. Kalman [KP94]. Kapazitätsbeschränkungen [Wen76b, Wen76a]. Kaplan [AdUÁS15, SW16, dUÁRC04]. Karl [Sta20]. Kegelhalbordnungen [Foc75b]. Kendall [Amm20, NCJ23]. Kernel [Abd99, BLNGB05, CH91, HK87, JO97, KE96, Kob87, KY22, Lie90a, Mam95, BLO03a, Cao01, CHL09, Che21, DW93, DGL16, Iga20, JSV07, Kie17, LMP10, Lie82, LQ12, LW12, LBR23, MP96, Moj07, Nad80, dDNN19, Nog20, Pon86, RLAV13, Sch90a, SC90, Van85, Wit76a, ZZLL20, dQLS16]. kernel-reproducing [Wit76a]. kernel-type [LBR23]. Kernels [PB21, Sch82a, Wol75]. Kernen [Wol75]. kernreproduzierender [Wit76a]. Kette [KS79]. Ketten [Küc78, Gil78]. Ketzel [Böh81b]. Kibble [CTL14].

Kiefer [Her81]. kind [AM15, Gup05]. Kinds [Kin75]. kinetic [SDKM18]. Kingman [RS74]. Kingmans [RS74]. Klasse [Bei74, JK76, Nož72, Sch73, vW74, Jür84]. Klassenwahl [BKN70]. klassische [Ter72]. Kleffe [Dry84]. Kleffe-approach [Dry84]. kleinen [Deu72]. Kleinste [Tau84]. Kleinste-Quadrate-Schätzung [Tau84]. Klelinen [Dro90a]. knifing [Zwa85]. knowledge [BN81, BS85, Wäc86]. Known [GL94, Hol91, CFS17, CT03, FE85, GK93, HDC20, Hös78, Kha15, LLBM^{+}11, Nus77b, Ros71, Sch82c, sS14, Zin81]. Koeffizienten [Wei70b]. Koeffizientenänderung [Som71]. Kolloquium [ES74]. Kolmogorov [Haf75]. kombinatorischer [Sch70b]. Kompakte [Boe72]. Komplementaritätsprobleme [GR76]. komplexen [EW73]. konechnoi [Ano81e]. Konferenz [Lom74]. Konfidenzschätzverfahrens [Mau89]. Konjugierte [EN75]. Konstruktion [Hol70b, Ste76, Got74, Mau89, Ull76a]. Kontingenz [Röd71]. Kontrast [MW80]. Konvergenz [Fäh78, Fäh81]. Konvergenzgeschwindigkeit [Kin75, Müh80]. Konvergenzuntersuchung [Hol76]. Konvergenzuntersuchungen [FGR73]. konvexe [Sch75f, Tam74a]. konvexer [Mel76, Ric76a]. konzentrierten [FN77]. Konzeption [Sch70b]. Kostenfunktion [Ban70a]. Kotz [HK09, Nad03, NK05a, Nad12]. Kotz-type [Nad03]. Kovarianzen [Fei72]. Koziol [Bre05, Jör99, ZR10, dUÁ00]. KPSS [KY16]. Krafft [Pil86]. Kreise [Pie72b]. Kres [Hof84]. Kriging [TA90]. Krishnamoorthy [She03]. Kritische [KNS72]. Kuhn [Hol73a, Hol73a]. Kuks [Hof95]. Kullback [PS14, SSN17, WNBH17]. Kumaraswamy [BSL13, KC17, RN18]. Kumaraswamy- [KC17]. kurtosis [AHPS84, DLO22]. kvadrate [Ano72a].

l [AM90]. L.S. [Páz87]. Lack [DN97]. Lag [Ron76]. Lager [Köc75]. Lagerhaltung [Ric75]. Lagerhaltungsmodell [Köc75, Köc76]. Lagerhaltungsmodelle [Gir74, Kle72]. Lagerhaltungsmodellen [Gri80]. Lagerhaltungsmodells [Gir71]. Lagern [Köc76]. Lagrange [CF89, GJ73]. Laguerre [Hen92]. Lampard [LO86]. LAN [KHNT17]. Lang [Mön72]. Lang-und [Mön72]. Laplace [BM02, CK06, DLO22, HM02, IM14, Kla03, LD21, Nad07]. Large [BG19a, Car88, DJ09, GJV16, IKR00, LLBM^{+}11, Lou99, MP10, TJZ89b, Amm20, Art19, BJ85, Bic18, Bos93, HE12, HCW16, JB83, LCY17, Luc05, PS77, PP94, Sch22, ST75, Wol84, XCJH19, Yin20, MESS70]. large-deviation [Bic18]. large-dimensional [Yin20]. Large-scale [BG19a, XCJH19]. large-sized [ST75]. larger [KS93]. largest [BJ85, BT16, FB16, HSBB23, WBL06]. lasso [PR22, SH17, Ciu16, TSC20]. latent [LCY15, MV10, NB21]. latin [GK77, Son82]. Law [EMG98, GM88, HCW16, MY11, SZ17, Try89, Car88]. laws [ÇO17, CV78, Dew75, PP94, Sch22, Zäh14]. layer [Deu72]. layout [LTV03]. layouts [KD74]. LBI [HG05]. learning [NB16, PB21]. lease [Pra86]. Least [BFF00, Bed85a, BTV21, Bun75c, DI11, Kun97, Nus76a, TVL96, BCMR90, Bel79b, BS80, Car19, CH89, CFR89, Dro93, FS02, GKM18, Haf82a, Haf82b, Hof77b, Hof80, Hös78, Hue86, IZ83, JR06, KJ14, LKM14, Läu89, LT13, Lit14, Liu92, LH15, Páz84, Pes13, Raf88, Rao83, Rod93, Sch75b, STS85, SL11, Tau84, TW78, Van91, Zwa83, Zwa85]. Least-square [BTV21, Sch75b]. Least-squares [DI11, KJ14, Lit14, LH15, Rod93, Tau84]. least-squares-estimators [Bel79b]. Lectures [Atk75]. Lee [ZQC23]. left [KLF22, sS14, sSW19, SAX23, WLT^{+}22, dUÁV17]. left-truncated [KLF22, sS14, sSW19, WLT^{+}22, dUÁV17]. Lehmann [GD96, JRT21, Zuo01]. Leibler [PS14, SSN17, WNBH17]. lemma [Bed86, Thä75]. Lemmas [Thä75]. Length [RO98, AFSS22, Goe75, Goe76, Kut12, Mau89, Moe84, NJKK16, Zäh74]. Length-Biased [RO98, AFSS22]. lengths [Kla72, KSS76]. Lenz [Hec86]. Leontief [Düc75a, Düc75b]. Leontiefschen [Düc75b]. Leontiefscher [Düc75a]. Lepage [JK92]. Level [Cav97, BM09, COQ21, EP17, Fra77, GSA14, HY10, KCNB11, Mat89, NJKK16, PCKMK17, Vís86, WQO22]. level-crossing [Fra77]. levels [BT04]. Lévy [Kaw15, LM23, Mas19]. Life [Ach95, BN97, GK98, MS98, BMK22, CBP23, CLO13, FPLS12, GK17, IRB14, KDM22, NRS07, NFP18, RALN93, SMK22, SU93, TDY08]. Lifetesting [Phe89]. lifetime [AJV21, AB12, BA08, BSB13, Bei75, BN16, CF21, DZ22, EPT20, MA20, Man22b, MSCC16, PN10, PDG13, RNT13]. lifetimes [BC23, HKK23, Jas22, KHN20, NSK12]. light [WW21]. light-tailed [WW21]. like [BGV06, Jen73, STS85]. Likelihood [AK94, ADR14, Bed85a, BC02, BB00, Cie98, EW75b, Fei72, Foc84, FR98, Gil78, IR91, Koc74, LLW17, Nku08, Nus76a, Rüs91a, RR91, Sch77b, TL15, ZQB16, ZM95, vR97a, AFSS22, AÖ23, ABD93, BZ14, BQSR20, BP16, BBKP19, Bat10, Bel80, BT87, BTT11, BR19, Bun75a, Bun82c, BJ05, BDF^{+}18, CM13, CC12, CD08, DZ22, EL79, Fah90, GLZW12, GSTUM02, GNYX15, Guo20, GV06, Hof81, HLsS01, Hue86, Huz18, IM14, JSW22, KYR89a, KYR89b, KDM22, Kaw15, KLF22, LLL17a, LY16, LY20, ML21, Mal21, MW14, Mic87, MSCC16, ÖA21, PN90, SLL15, Sär95, Sch90a, Sch91b, sS15c, SX20, SYW22, TZ13, Tho80, TX17, TY12, Von91, WLD18, WLF13, WW22a, WD22, WK85, WL01, XWJ14, XZZ11, Xie09, YC15, YYVH16, ZK16, Zha13, Zha16, ZT15, ZMY20, ZL14, ZBF^{+}20]. likelihood-based [TZ13]. likelihood-MM [ÖA21]. likelihood-ratio [ZBF^{+}20]. Likelihoods [ACS93, CKM09]. Limit [BNAA13, BNH20, BR90, Eks99, EBQY17, EW99, FZF97, JS81, LBSM02, Lou99, Pad81, Pad84, Pad87, Sch89b, Sch79c, Wol84, AAB18, Ano87j, BL11, BW14, DM22, DST21, Góm18, KK14, KAM17, KK81, Lie72, MV18, SK20, Wol85, Zan14, Zan15]. limited [Goe75, Hol75, Zäh74]. limiting [BM12, CV78, Güv06, KM20, NM18, Sol19, Wyn76]. Limits [AAK98, Bök85, Vän98b, Bar10, Bei81, BJ05, Sha18, Tzi87]. Lindeberg [BV86, Góm18, RR02]. Line [Las91, Arn92c, RD80, Sto72a]. Linear [AS98, BM94, Bun93, Bun99, CH08, CM90, DN97, Det93, Det95a, DM95, EA90, EW99, FZF97, GH89, GM91, GM99, Hof97, Hor95, HHS97, Jon95, KM95, Kov74, Läu75a, LMS92, LTG97, MD03, McC03, MZ96, Nät97, Neu92, NW12, Nus76a, PV98, Pet73, Puk81a, Rod89, Sch96, SL95, ST91, Ste98, Thr79, Tro00, WHG87, AAR18, AS79, AK03, Als88, Als93, ANR07, AFV15, Ano73e, Ano75c, ARN12, Are05, AS04, AÖ23, ABD93, BCMR90, BJfYZ13, BKK03, BK76, BK79a, BB93, BOHJ12, BB22a, Bal03, Ban70a, BN81, BQSR20, Bär72, BTS71, Bel74, Bel79b, Bel80, Bel83b, BP17, Ber75, BV01, BTV21, BLO03a, BDS22, Bon88, Bos88, BR15, BB72, BG74a, Bun75b, Bun75d, BG79, BW80, BR83, Bun05, BDF^{+}18, Bur08, CXW20, Car19, CHPS13, CC12]. linear [Che21, CLLY22, Cou21, DS78, DLMR13, DWHH19, Dro06, Dry75, DH81, DT16, DJ86, Dup77, EL79, EW73, EM87, Eva75, Fah87, Fah90, FX16, FG75, Foc75b, Fur15, FR20, GLZW12, Ger75, GV19, GP82a, GKTZ92, GK72, GV06, Har75a, Har75b, HK00, HW88, Hen79, Hof77a, Hof79a, Hof79b, HP73, HW71, Hol73b, Hös74, Hös78, Hue86, HM07, Jag80, JS22a, JRCB22, JRZ03, Jur75, KG72, Ker84a, KHN16, KP15, KS10, KP74a, KP74b, Kle75a, Kle77a, Kle78, KJ14, KR17, Kra86, KZ13a, KZ13b, KY22, Läu71a, Läu71b, Läu71c, Läu75b, Lem13, LY11, LL18, LMS06, LWS21, LY93, LQZF17, Lip75b, Lip76, Lis90, Liu92, LH13, LGW14, LH15, Log87, LMV17, LST18, Lud73, MP11, MCML15, Mat83, Mat87a, Mat81b, Mat87b, McC05, Men73, MY11, MA86, Nad07, NN82]. linear [Nät72, Nät75a, NL05, NTJ09, Nku23, Nož70, Nož71, Nus76b, Nus77a, Nus77b, Nus79, ÖA21, Pad84, PR22, PCKMK17, PW76, Pie70, Pie75a, Pie76, Pil81, Pil86, Pin75, QLT17, Qin15, QK17, Que86, Rao83, RH80, Röd77, RH20, Ros75, Ros82, Rud76, Rus91b, SY81, SGG04, Sch76b, Sch91b, Sch75b, Sch75c, Sch76c, Sch77a, Sch79b, ST81b, Sch70c, Sch76d, ST75, SYM16, She88, SC18, Sin05, Som71, SW89b, Stu90, Tab13, Tam74a, Thä76, TX17, Tou70a, Tou70b, Tou75, TR78, TS87, TPSR17, Trö75, Try86, Try89, Van91, VS81, Wag83, Waw81, Wei70a, Wei71, Wei72, Wei74, WY13, Wu16, WWLH19, WA22, XG07, YC15, YH15, YLL16, YLY17, YLY18, YS23, YY07, YAA22, ZY07, ZWYG14, Zha18, ZZLL20, ZXT11, ZX17, ZYW21, ZWFZ22, Zvá79]. Linear [GR76, Wei70b]. Lineare [PW76, Ban70a, FG75, Foc75b, Har75a, Har75b, Lip75b, Lip76, Tam74a, Trö75, Wei70a, Wei71, GR76]. linearem [Läu71a, Läu71b, Läu71c]. linearen [Bel74, EW73, Hof79b, HW71, Hol73b, Lip75a, Lud73, NN82, Nät72, Nož70, Nož71, Nož72, Pie70, Pie75a, Pie76, Sch76b, Som71, Tou70a, Tou70b, Wei70b]. linearer [Bär72, BTS71, HP73, Sch75d, Sch75e, ST75, Thä76, Wag83]. lineares [Sch70c]. linearisierung [Bär76]. Linearity [KW17, WLB04, KZ13a]. Linearization [BDP16, Bär76]. Linearly [CXH14, Azp73]. lineiinom [Ano73e]. lineinikh [Ano75c]. lines [Aby72]. LINEX [CS06, HL13, Ohy13, Tan11, ZZWG09]. Linhart [Dro91a]. link [SIM21]. linkage [Sti78]. Liouville [SR02a, Wes93]. Lipschitz [KS96]. lits [Ano75a]. Liu [Nku23, AAR18, KY22, SS20a, YAA22]. Liu-type [Nku23, SS20a, YAA22]. LM [Sin05]. LO-Probleme [Ste76]. LO-Problems [Ste76]. loads [FGK19]. Local [AAJ97, Jon95, KLF22, Mas19, NTJ09, PSCC95, WLF13, Ano87j, BKK03, Bar10, Bed89, Ber00, Cao01, DLMR13, STS85, Wol85]. locality [BCDD11]. Locally [ROM15, GR76, Kra76]. located [FN77]. Location [AM90, BC02, BT97, Cro96, DT96, Dor85, GM97, HPN22, KS95, Lai00, MS94, Rie89b, Als88, BB22a, BKHB22, Dac11, GF09, GPK23, HH13, HW07, JW04a, JS82, Jur80, KK19a, Kat89, PK18, Pon84, Rie89a, Ryc87, Ryc93a, ST23, Sta88, VO78, WN13c, WN14a, Zie85, Zie88b]. location- [GF09]. Location-Scale [BT97, Lai00, GM97, BB22a, BKHB22, HW07, JW04a, Sta88, WN13c]. locus [ADR14, Rab15]. Log [AMA14, BL96b, FV06, KR07, BC23, Bow21, DK12, HOCC13, Kun15, LCY15, MP11, Mat87b, NRD08, SSM07, dQLS16]. Log-concavity [FV06]. Log-gamma-generated [AMA14]. log-generalized [HOCC13]. log-linear [MP11, Mat87b]. Log-normal [KR07, BC23, DK12]. log-odds [NRD08]. log-skewed [dQLS16]. Log-Spline [BL96b]. log-Weibull-negative-binomial [LCY15]. Logarithm [EMG98, Mik88]. logarithmic [BKS23]. logarithmic-type [BKS23]. Logarithms [Bed85a, Eks99]. Logistic [BGP95, ESSB93, GN00, VV00, AAFG23, AA11, AA13, BKT14, Geb84, GNP08, KHN15, Ker16, KM17, MR73a, ÖA16]. loglog [MY11]. lognormal [MO73, NL79]. Lokale [Wol85]. lokalkonvexen [GR76]. lokalnoi [Ano87j]. Lomax [AHJN01, LC13]. Long [GM92, Mön72, AOV19, BTT11, Ciu12, DRW17, FRLLT18, FN13, GS17, HOCC14, HS17a, JK08, LLBM^{+}11, Wan07]. Long- [Mön72]. long-memory [HS17a]. long-range [AOV19, DRW17, JK08, LLBM^{+}11]. long-range-dependent [FRLLT18]. Long-Term [GM92, HOCC14]. longitudinal [DT16, GZ23, LQZF17, LGL^{+}17, PA14, SMW22, SS21, TA21, TSG20, TX17, XB17, YLL16, ZT15, ZL13, ZL14]. look [Els17]. Looking [MMRA07]. lor [Bun85]. Lorenz [CCH87, LW18]. lösbarer [Hol70b]. Lösbarkeitseigenschaften [Thä76]. Loss [Fle76, VV96, BM84, CXH14, CFR89, CS06, HL13, Joz14, KP09, Laz16, LY93, Lis79, Ohy13, PKP21, PH16a, SP08, ST87, Tak09, Tan11, TB01, YH15, ZN14, ZZWG09, Sto72b]. Lösung [Bär72, Ehl75, Ehl76, GT70, HT71, HP73, HW71, Mel76, MESS70, Nož70, Ric76a, Sch76b, Sch90b, Sch70b, Sch75d, Sch75e, ST75, Sch75f, Tam76, vW72a, vW72b, Lud73]. Lösungen [HK78, SSG76, Wag83]. Lösungsalgorithmus [Sch70c]. Lösungsformeln [Bor73]. Lösungsmenge [Deu72]. Lösungsverfahrens [Sch73]. low [LPS90, Páz90a, Ryc09, TSC20]. low-dimension [TSC20]. Lower [Bis95b, GR06, Miz96, CAK13, KAR14, KR17, QCO13, Tar88, TA11, Zin81]. lowest [BS21]. LR [Lem13, Sin05]. LS [MY11, MWZ15]. LSE [AOV19, Hos88b]. LTRC [sS15c]. Luria [Zhe10].

M [CT73, Fle76, Fre74, RL98, SYW22, WHXC16, CT73, RL98, Sto76b, RL83]. M-estimation [WHXC16]. M-estimators [SYW22]. M/G/ [Fre74]. M/G/s [Sto76b]. M/M/1 [RL98]. M/M/C [CT73]. MA [FL03]. machine [Men73, Ter72, Ull76b]. Mahalanobis [TIQ10]. Mai [HW74]. Main [OM91, CAK13, Jac13, Lei70, MOS86]. Main-Effect [OM91, MOS86]. maintenance [Bei76, Bei81]. major [Pon93]. make [Sto93]. Makeham [MGM20]. maksimalnikh [Ano75c]. management [vW74]. manifold [BCCF14]. Mann [NCJ23]. MANOVA [KH21, Mic87]. many [KM13, Sto74]. many-server [Sto74]. mapping [Rab15]. mappings [FK14]. Marcinkiewicz [Zäh14]. margin [Art19]. Marginal [Tro00, Wal88, De 22, Jia07, Kha12, Mat87b, May11, TIT08]. Marginal-Fréchet-bounds [Wal88]. Marked [Sto91, BC06, HS79, HE12, Laz16, Str82]. market [Gru71]. markings [WS97]. markirovanniye [Ano75b]. Markoff [Bak84, OKW86]. Markov [ANN11, Ano73c, Ano73b, Ano75b, Ano81e, BG16, Cse95, DGL16, FR98, FR86, GLR23, Gil78, GV79, GV81, GW84, GH01, GH02, HBN86, JR86, Kir76, Küc78, KS79, Küc82a, Küc82b, Kum85, Lie92, Mat81a, Mel94, Nog20, PM21, PH16a, Pon04, Pon06, Ryd95, Win17]. Markov-dependent [Kum85]. Markov-Switching [FR98, BG16]. Markova [Ano81e]. Markovian [Mil86]. Markovprozeß [FR86]. Markovschen [Gil78, GV79, Küc78, KS79]. Markovskiye [Ano75b]. markovskoi [Ano73b]. Markovskovo [Ano73c]. Markow [GH01]. Marktmodellen [Gru71]. Marshall [JK13, WG21]. Martingale [FGL17b, BQSR20, Che12, DST21, Hel93, HBCL14, LZW23, MWZ15, Pad87, ZL13]. martingales [FGL17a]. Maschinenbelegungsplanung [Ull76b]. Maschinenbelegungsproblem [Ter72]. mass [GVGP13]. massive [JHYQ18, JLM23]. massovovo [Ano75b]. matched [MSW03]. matching [CM13, OM10, OM11, PW15, Vid23]. Math.operationsforsch.u.statist [Sch84]. Mathematical [Gir74, HW74, Lom74, Ras91, War85b, Cen78, SS74, Tit70, Els75, FN77, HW70a, HW72, Mül74, HW70b]. Mathematische [Els75, Gir74, HW70b, HW70a, HW72, HW74, Lom74, FN77, Mül74]. mathematischen [Tit70]. mathematisches [SS74]. Mator [Cau99]. matric [JG85]. Matrices [Das96, GON95b, HR96, BK77, BG74a, BW80, CC13, Düc75a, Gaf85, GSA14, Koz86, NFF^{+}23, Rod88, SGG04, Yin20]. Matrix [BN03, CG05, Düc75b, Fou97, Fur97, GM99, GV95, GN98, JS97, Wen95, BG11, BS17, CLDG16, CFS17, GK77, GK93, GS04, HG08, Hös74, HB09, Juh89, KP74a, Kle78, KvR00, LY93, Lis90, Net90, Pin74, Pré74b, RFK18, Rub21, SGG04, She03, ŚS05, TH04, Wei72, Wei74, Wit83b, ZHG12, ZRAN17, vR88]. matrix-variate [TH04, ZHG12]. Matrizen [Düc75a]. Max [Sch73, HT15, LLL17b]. max-discretization [HT15]. maxima [FZB17, HWL17, LT16]. Maximal [Col91, Ryc15, Ano75c, BB72, Lie82, Nad80]. maximally [Oko16]. maximin [Kra86, Pil86, ROM15]. Maximizing [Pap16]. Maximum [AK94, BTT11, BB00, Cie98, CD08, Cro96, EL79, Fah90, Fei72, Foc84, FR98, GC90, GSTUM02, HLsS01, Hue86, IR91, JR13, KYR89b, Nus76a, Sch91b, Tho80, XZZ11, ZT15, vR97a, ABD93, BS13, BZ14, BNAA13, BP16, BBKP19, BLT20, Bel80, Bun82c, BJ05, BDF^{+}18, CK06, GGL15, GLZW12, GF09, GT13, GV06, Huz18, IM14, JRCB22, JSW22, KYR89a, Kam89, Mal21, MW14, Mic87, Nov10, PN90, Rei84, SLL15, Sch90a, Sch77b, Von91, WD22, WK85, XWJ14, Xie09, Zha13, EW75b, Foc73, Gil78, Koc74, Sch77b].

Maximum-Likelihood [FR98, BTT11, XZZ11, ZT15, BP16, IM14, MW14, Zha13, Gil78]. maximum-likelihood-estimator [Bel80]. Maximum-Likelihood-Schätzung [Foc84]. Maximum-Likelihood-Schätzungen [Fei72, EW75b, Gil78, Koc74, Sch77b]. Maxwell [GG15, HWL17]. May [GNS15, HW74]. McDonald [CL14, CHO14]. MCMC [SAEAEA13]. MDP [Wan15]. Mean [BM09, BGP95, FSV13, Fer99, Gat94, GN00, GM99, GK98, Hol91, IRV98, IV02, LPC97, Mol96, PSCC95, VW96, AALA20, AB12, ADF21, BJ85, BMK22, Bat10, BR12, BA89, BP18, Bha88, Bie16, BS18, BAO12, CS06, CT03, Dan03, DZ22, DST21, EPT20, EBTB21, Fin94, GGL15, GJHN18, Gno76, GB23, GS21, Gup77, GK93, HS17a, IV90a, IRB14, JC22, JT22, Kam89, KK80, KDM22, Kha04, Kha15, KS10, KS76b, Kre80, KMSS10, Kut12, LC15, Lät74, LdCIP18, LLL17b, LT13, LBSM00, LGL^{+}17, Mec84, NRS07, Pap16, PN10, RA10, Sch22, SS86, Sei78, SKCE19, SZ17, TH08, Wan14, WBL06, ZC10, ZRAN17, Zin81]. mean-covariance [LGL^{+}17]. Mean-squared [BM09]. mean-variance [Dan03, KMSS10, Pap16]. meaning [RS74]. means [Ahm93, BGL15, CFS17, Dan03, FYY^{+}23, Ros72, Sta84, TT10b, VO75, Wag83]. Measure [Fer99, TIT08, BKK03, Bow12, Bow21, CAK13, CWW09, Han83, Joz14, MA20, Man22b, NL05, Ohs83, TZ91, TMY05, ZQC23]. measured [BS07a, BS10]. Measurement [PB10, RF03, BKS23, BL06, HK78, HLM21, JRCB22, KMSS10, KY22, LLBG10, LQZF17, ML21, MCCL18, NCJ23, Ras06, SS20a, Sep01, WA22, ZWFZ22]. Measurements [AAK98, Afs85]. Measures [And79a, BRKO99, CHP03, NL79, SS98, Tjø96, AL18, BHM04, BP93, Böt20, Dab81, GK17, Han83, HH16, Hei88, HM89, LUK14, Man22a, PEB15, Puk81b, Rüs87, SK89, SV10, WL05, WNBH17, Wie88, Röd70]. mechanism [TSG20, TK22]. Median [AHJ99, AEW94, Zie88a, Zie00, Cad01, MM18, SV18]. Median-Unbiased [Zie88a, Zie00]. Medians [Alb84]. medium [Mön72]. medium-term [Mön72]. Meeting [Gir74, Kle72, Mön72]. mehrdimensionalen [Wit81]. mehrere [Köc75]. mehreren [Köc76]. Mehrfache [BTS71]. mehrliniger [Sto73]. Mehrproduktmodelle [Ric75]. mehrstufiges [Sch70c]. Meier [AdUÁS15, SW16, dUÁRC04]. meiner [Gus71]. Memory [EZ92, BTT11, Ciu12, FN13, GS17, HS17a, Wan07]. Mengenänderungen [Sch70c]. Menten [DK14]. mercury [NJKK16]. messender [HK78]. meta [SKCE19]. meta-analysis [SKCE19]. Method [Abd99, Bil97, Lav95, TP97, Ano74, AÖ23, Bar87, BSL13, Bel83c, BNS15, Ber00, BCN15, BG74b, DDT88, DDT89, Dew73, DFRLÚF12, Duv17, Ege80, Ehl75, Ehl76, FN77, Fin87, FT79, FS02, FD72, Foc75a, GJK^{+}16, Ger75, Gir71, GR72, Gud70, GT70, HTX23, Hav76, Hof72, HP73, Hol75, Hol76, HPN22, HY10, Huz18, Jen73, KTW87, Koz86, LQ12, LL85, Lom70, Lys76, Mau89, Mel76, NDDDG16, NOM13, PVW74, Pfu78, Pie70, PD91, Pré74a, Rév77, ST75, Sch75f, Sin09, SH09, SMLL18, SG91, Tro91, Tro93, Wag83, XCJH19, ZTJ17, Bär72].

Methode [BG74b, FN77, Gud70, GT70, Mel76, Pie70, Pré74a, Sch75f, Wag83]. Methoden [Lom74]. methodology [PA14]. Methods [ANR07, HH16, Lom74, Sha96a, Ano73a, AL82, BN82, Dup77, EG74, FGR73, Has77, HJM90, Hub77, Jac91, JD20, Kle77a, Lip75b, Lip76, Pie71, STS85, SH86, THN13, Win81, Mai91]. Metod [Ano74]. metodi [Ano73a]. Metric [Pad96, Pet73, FP92]. metrics [Bow12, SDC19]. metrizing [AL18]. Metropolis [AP07, Win17]. Michaelis [DK14]. midrange [MR73b]. Midzuno [MB92, MB91]. min [Sch73, Sch73]. min-max [Sch73]. Min-Max-Problemen [Sch73]. Mini [LLL17b]. Mini-max-risk [LLL17b]. mini-mean-risk [LLL17b]. minima [FZB17]. minimal [FR86, Moe84]. Minimality [Bod75]. minimally [Oko16]. Minimalvarianz [FR86]. Minimalvarianz-Schätzungen [FR86]. Minimax [Als88, Bei75, BL96a, Bun75d, BM84, Bun99, GH89, GON95b, Hol75, Kal97, Köc76, KS96, Miz96, MG96, Rei86, ST83, Sch70a, ST87, Tou75, TR78, Try89, Wil08, vW72a, vW72b, Als93, Bun75c, DM09, Hof79a, JRM04, Kin76, Koz85, Läu75a, Läu75b, Laz16, PH16a, Pet73, Pil81, Rie89c, TS87, ZRAN17, vW74, Nät72, Zin81, vW74]. Minimax-Aufgabe [vW72a, vW72b]. Minimax-Aufgaben [Hol75, vW74]. minimax-Bayes [Pil81]. Minimax-Lagerhaltungsmodell [Köc76]. Minimax-linear [Tou75, TR78]. minimax-tasks [Hol75]. Minimax-Tests [Sch70a]. minimaxity [She03]. Minimaxschätzer [Zin81]. Minimaxschätzungen [Nät72]. minimaxtheorem [Kin76]. Minimising [Imh00]. minimization [CBC17, Fav14]. minimizer [Cor17]. minimizing [BSP76, BH89]. Minimum [BG16, CF89, DGL16, JRT21, KN98, WB05, Zha18, vW74, BMMP16, BS07b, Dal77, GS88, GKZ77, KC17, Ouy92, SP08, Waw81, MW80]. Minimum-contrast [MW80]. Minimum-Kontrast-Schätzungen [MW80]. Minimum-Problemen [vW74]. Minimum-volume [Zha18]. MINQE [ABD93, Dry88]. MINQE-theory [Dry88]. MINQUE [Dew73, FD72, Dew73, FD72, Kle76, Kle77b, Kle80, Koz86, WG92]. misclassification [MPPZ05]. MISE [MB21]. Mises [Hol91, Neu73, PJ05, Pho87, PD91]. Mises-Statistik [Neu73]. mismeasurement [RLAV16]. missing [BAM18, BAO12, CLLY22, Cip91, DDD19, FSV13, GNYX15, KS16, LWS21, LLV16, LY16, OKW86, TSG20, TZ13, TT22, Zha16, ZZW23, ZTJ17, ZL17, ZWFZ22]. Mission [Cse95]. misspecification [LYB19]. Misspecified [AS98, Fah90, KM17, XWJ14]. miteinander [Köc75, Köc76]. mittelfristige [Mön72]. mittels [SSG76]. Mittelwert [Zin81]. Mixed [Bun77b, GM99, HC11, Khu95, MZ96, Tan91, Tro00, AAR18, Are05, DT16, FMTZ07, GCL16, GSTUM02, Gui86, HT71, JRCB22, Kle77a, KY22, LY93, Lis90, LH13, Pie76, Ric76a, Sch79c, Sch81, Sei79, SS21, sSW19, Ste76, TWZ06, TT22, WQO22, XZZ11, ZT15, ZYW21, Ano72b, Har75a]. mixed-effects [DT16]. mixed-integer [HT71, Pie76, Ric76a]. Mixed-state [HC11]. Mixing [AL99, BO01, Bos93, CD09, FL03, Hei92, Ker84b, LLWW20, SV89, Wan15, WYLW22, WS22, WWLH19, WYW21, WHVW23, XWY^{+}20, Yat19]. Mixture [BLSC09, GM92, GK99, LP97, AA12, ASSH13, AJM12, AG09, BP18, JY20, PMA17, Pol08, Ryc21, Tit90, VYD83, WH20, WQO22, WN13c, Yat19, ZL14]. Mixtures [SD86, AHJN01, Bet87, CLS14, CCD15, Eco13, FH01, HH13, KCL16, LLBG10, Miz17, Pon93]. ML [KP07]. MLE [DB92, VS81]. MLEs [CBC12, GAJB13, HW07]. MM [ÖA21, Ars16, ÖA21]. MML [BT97, GM97]. MMLE [Tou75]. mnovostupenchatoii [Ano73e]. mode [LLV16]. Model [AJ97, Atk75, BPH88, BR19, Bon88, BM94, BCDD11, Bun73b, BDP99, CLLY22, CF98, Det93, Det95a, DZX18, FC89a, GL94, GM99, GK98, HA02, JS97, Jör99, KM96, LTG97, Luc99, Mec72, MJA02, Neu02, PRS95, Pon02, Rie89b, SS94, Sch89c, SL95, SZ97, Tan91, Vie02, WHG87, Wen95, ZZZ17, vR97a, ASSH13, ASFKV08, AT05, AHJN01, AC23, AK03, Als88, AAA13, AFSS22, AOV19, AST13, ACH20, AK20, ASM83, AE21, BM09, BK77, BK79a, BK79b, Bak84, BQSR20, BPvZ04, BSB13, BR90, BBK12, BBK14, BMK16, BK17, Bed85b, Bed04, BP18, BP17, BCCF14, BN16, BAM18, BDS22, BGL15, Bre05, BG78, CLS14, Çet21, CY12, CG04, Che12, CC12, CD08, Ciu12, Ciu16, Ciu20, CL14, CHO14, Csö88, DK14, Dew75, DK21, Dro06, DH81, Dup17, EAA16, FX16]. model [FP92, Fin94, Foc84, FMTZ07, FR20, GZ23, GPS14, Gir71, GJ19, GS17, Gun88b, Gun88a, GNYX15, GK72, GCT07, GV06, Güv14, GP07, HS83, HG05, HOCC13, HOCC14, Hav88, HW88, Hen93, HEHM13, Hof79b, HLsS01, Hu10, Hue86, HM07, IZ83, Jag80, JR86, JR12, JJS06, Jur75, KHN15, KA88, KSS11, KAR14, KCL16, KP74a, Kle75b, Köc75, Köc76, KvR00, Kou11, KMSS10, KZ15, KHN20, LKM14, Laz16, LDF17, LMS06, LWS21, LY93, Lis90, LH13, LGW14, LLZ19, LCY15, LYB19, LGL^{+}17, ML21, MMP16, MFBG14b, Mat83, Mat81b, MWZ15, Mou01, MAS06, NN82, Näg75, NKI04, NK13, NK23, Nät72, Nät75a, Nie06, NW12, dDNN19, NS14, Nus76b, Nus77a, OKW86, ÖA21, ÖA16, PS12, PMA17, Pen85, PH16a, PS10, PN90, Pol08, Pon04, PDG13, Puk77, QK17, RAD07, Rie89a]. model [Rie87, RVBZ20, RS74, Ros71, RV08, SK92, SK14a, SLL15, SGMFB18, SGG04, SS15a, SK14b, Sep01, Sha21, SS21, sS15c, SC18, sSW19, SS74, SKV18, SMLL18, SG91, Stu90, TA21, TSG20, TT10a, TX17, Tom90, Tou70b, Tou75, TS87, VFS^{+}20, Von91, WHXC16, WH20, WG21, WYLW22, WW22a, Waw81, WZW21, WL01, WY13, Wu16, WWLH19, WA22, Xie09, YS23, Yoo18, YLWL22, Zez93, ZY07, ZR10, ZZL^{+}14, Zhe10, ZWFZ22, Dro91a, Pie72a]. Model-Based [Neu02, Yoo18]. model-checking [BBK14]. Model-cum-design-based [BCDD11]. model-design [NW12]. model-free [SMLL18]. Modeling [BL96b, Mül74, Tit70]. Modell [Hof79b, RS74, SS74]. Modelle [JK76, PT83, Wer78]. Modellen [Ron76]. Modellierung [Mül74, Tit70]. Modelling [CGM19, Eco13, NSK12, SKSZ20, Arn91, GS11, GN17, LV18, MP11, MV10, NL05, PH16b, SL17, XB17]. Modellsysteme [Pie72a]. Modellwahl [Atk75]. Models [AVB98, AS98, BN97, BM94, Bun99, Das96, Det93, DM95, GH89, GM92, GM94, GH02, Han96, HHS97, HDA99, Khu95, LP97, LMS92, LNBR02, MD03, Mi96, MZ96, Mol96, Mon96, NK06, Neu92, PP92, RF03, Ryd95, Sch96, SLS99, Tia91, TM94, Tro00, Vie95, WS97, Yao00, AMP21, ASSH13, Aka82, AAR18, AT08, AL18, AS79, And89, ANR07, ARN12, AINJ14, Are05, AGS09, AGOP14, ABD93, BCMR90, BK76, BH18, BT16, Ban80, BN81, BS85, BNP87, BC14, BP16, BHN15, Bed81, Bei74, BBL20, BB91, Ber74, BTT11, BTV21, BS07b, BZ10, BN82, Bon88, BCN15, Bro93, BR15, BG74a, BS75, Bun75b, Bun77b, BHSW77, BW80, Bun81a, BD85, Bun05, BDF^{+}18, CXW20, Car19, CHPS13, CF21, CTW09, CLLY22, CGM19, CCMW09, Cox77, DM06]. models [DGCL14, Dik01, DR03, DB92, Dör15, DMR07, Dry75, Dry77, DT16, DdMT20, Dup77, EM87, Esc84, Fah87, Fah90, FPPRG16, Flo93, FS02, Gai23, GLZW12, GLR23, GK17, Gir74, GP82b, Gno79, GKTZ92, GSTUM02, GMCT18, GSB06, GSB11, GLLO^{+}22, GMLS16, GML18, GML23, GGHS21, Gri80, Gru71, GNP08, GH01, HBN86, HMP01, HC11, HK00, Has77, HX20, Hei92, Hen79, HR93, Hös86, HW07, Hos88b, HLM21, JRCB22, JS14, JK76, JRZ03, Ker16, KP15, KRC19, KP74b, Kle75a, Kle76, Kle77a, Kle77b, Kle78, Kle72, KS76a, KMM15, KN06, KY22, LLBG10, LTV03, Läu70, Läu74, Läu75a, Läu76, LM79, LKK07, LLC16, Lem13, LSS12b, LGZL14, Lin09, LQZF17, Liu92, LH15, Log87, LMV17, LST18, MCML15, MNvRS10, MCCL18, MS80a, May11, MP23, MY11, Mic87, Mic09, Mie79]. models [Mön72, MF17, MPPV04, MA86, MSW13, NS03, Nät85, Nku23, NB16, OM11, PM21, PR22, PPD07, PA14, PJ17, PB10, Páz90b, Páz91a, PM16, PSB23, Pet73, PT83, Pin77, PS84, PD80, Pol85, PD91, Pon06, Puk81a, QJ08, Ras91, Ras06, RT13, Ric75, RLAV16, ROM15, Ron76, RH20, Ros82, RH13, Rud76, RM16, Rüs90, SY81, SS20a, Sch91b, ST83, Sch75b, Sch75c, Sch76c, Sch77a, Sch79a, ST81b, SDKM18, Sch76d, Sei79, SH17, She88, sSLMJ17, Sie14, Sin05, SA23, SAEAEA13, SL23, Spj77, Sto72a, Tab13, TB78, TW05, TWZ06, TZ13, Tan14, Tit83, Tor17, TK22, Tru18, TT22, Van91, VN19, WL05, WS13, WZXH15, Wan15, WT20, Web89, Wei09, Wer78, Wit89b, XWY^{+}20, XWJ14, XY16, XW10, XG07, YC15, YLL16, YLY17, YLY18, YL17, YAA22, ZW20]. models [ZWL14, ZWYG14, Zha18, ZT15, ZQB16, ZZW^{+}19, Zha23, ZL13, ZL14, ZYW21, ZZZ17, Zil78, Zwa80, Det95a]. Moderate [MW14, SX20]. Modern [And78a, Ham78]. modes [CF22]. Modification [GT70, Foc75a]. Modified [BM94, Koz85, RGB13, BBR05, CXW20, CL12, FC89b, Fle71, GS82, HEHM13, Kér72, LW12, Mur15, NM18, TKTM15, YN23]. Modifikation [Foc75a]. modifizierten [Fle71]. Modifying [JSV07]. modulated [NKI04]. Möglichkeiten [Tam76]. Moment [ESSB93, KT12, MR12, CKS22, Han83, Hei88, Hei91, KD16, LLWW20, LH22, Mei06, Ohs83, SS86, VCO09, WWH18, pZ15]. Moments [Ahm01, Dre98, HR96, Raq04, SP98a, SP98b, Von91, vR88, vR97b, BLSC09, BB04, BCN15, DST21, Ege80, Gla90, Gui86, Gup79a, Gup85, HS82, JB83, KT83, KTW87, LYB19, SP01, Sch85, Sch81, She14, SH09, WK13, Wie92, dSL22]. Monitoring [GS21, BT04, BS17, CHPS13, HTX23, WWZ22]. monotone [CF78, Hav84, KS16, Kow82, NFP18, Rüs83, vdVvdL03]. Monotonic [MS98, BL82, Kow77, Szc79]. Monotonicity [Kir76, Rüs86, Rüs81, Sto73]. Monotonie [Sto73]. Monotonie- [Sto73]. Monotony [Nät79]. Monro [LPS90]. Monte [BH89, GJK^{+}16, LMS92, Wag83]. Monte-Carlo-Methode [Wag83]. Moran [WN12b]. Morgenstern [NS10a]. mosaics [Mec84]. most [Mil86, Ryc87, ZX17, Zie85]. motion [KMM15, XZZ11]. motivated [BCCF14, FR20]. Moving [And76b, And78b, And78c, And79c, GPS14, GMCT18, Mie79]. moyennes [VO75]. MPSD [Gup82]. MRL [HSBB23]. msep [Dro87]. Muller [Tro91, Tro93, TP97]. Multi [BD13, KFPG21, Ano73e, Arn92a, CD08, Cou21, HH16, KCNB11, MBF^{+}20, Ric75, ROM15, Sch70c, Wit81, YS23]. multi-centre [YS23]. Multi-component [KFPG21]. multi-dimensional [Wit81]. multi-factor [ROM15]. multi-group [HH16]. multi-level [KCNB11]. multi-phase [CD08]. multi-product [Ric75]. multi-scale [MBF^{+}20]. multi-stage [Ano73e, Arn92a, Sch70c]. multi-target [Cou21]. Multi-treatment [BD13]. multichannel [NK23]. multicollinearity [LGW14]. Multicorrelation [Rao98]. multidimension [Ker84a]. multidimensional [Faz08, Hei91, IV90a, IV02, Ker84b, KK81, KN06, Mat85, Sch13, Try89, Wal88]. multidimensionally [pZ15]. multimodal [CHA15b, MFTFEO22, Zil78]. Multinomial [JRM99, Jon76a, Gno77, GC04, JRM04, Möh87b, NB21, Try91b]. multiparameter [BCN15]. multiphase [BS85, Ciu16]. Multiple [AGOP14, BH87, GK99, And81, AGS09, BG19a, BT16, BJ14, BHN15, BPH88, BR19, CZZ17, DR15, Fei72, FS81, FR20, GP82a, Gun88a, Ker16, Köc75, KR78, Kum13, MK19, TT10a, ZQB16, ZZZ17, BTS71]. multiple-outlier [BT16, BHN15, CZZ17, ZQB16]. multiplen [Fei72]. Multiplicative [MP22, CCD15, EAA16, HLM21]. Multiply [BGP95, AAFM07, SP08, VK20]. multipurpose [Läu76]. multiresponse [RV08]. multisample [NM18]. Multitude [NK04, ND05, NK07]. Multivariance [Rao98, Böt20]. Multivariate [BBK14, Bun99, Das96, DF15, GK99, Gup03, Hof84, JS97, KFL97, KT89, Koz78, Koz02, Kun17, LLBG10, LPC97, MCCL18, NRD06, PSCC95, SR02b, Vie02, dQLS16, AT13, ASKS16, And78a, AINJ14, BCMR90, BM09, BK76, Bha07, BZ10, BG11, BS17, Bun05, BDF^{+}18, CB02a, CM15b, DR15, EL79, tFYK01, FP92, Fin94, FK12, GS21, GVMS09, Gui86, GK72, GG90, GS04, GCT07, HG05, HH16, Hav88, Hen80, Kim10, KK15, KMSS10, KZ15, Lät74, LH15, MN08, MCNO19, ND05, NL79, NRS07, NN23a, Nus76b, Nus77a, Nus79, PYHT01, PG09, RVB18, RVBZ20, Roy87, RM16, Sch75b, Sch75c, Sch76c, Sch77a, ŚS05, SZ17, SK05, SM19, TF21, TB01, Try91b, Wes93, YT01, ZT15, ZHG12, ZZZ17]. multiyariate [Sch76d]. mutation [Zhe10]. my [Gus71].

n [Ano75a]. nach [Har75a, Har75b, Lys76]. nad [Ano73d]. Nadarajah [SN14]. Nadaraya [BNS15]. Näherungsverfahren [Gir71]. näherungsweisen [GM78]. natural [KP07, LBSM02]. Nature [CF98]. Navier [Sch90b, Sch90b]. NBU [Ger86]. near [AF14, BS13, BZ15, DB10, IDB12, Rei82, RZ11]. nearest [Haj84, HT16, Lie93, WS22]. nearly [CGK17, YH15]. nearly-balanced [CGK17]. Nebenbedingungen [Ehl75, Ehl76]. necessarily [LT13]. Necessary [LY93, Röd87a]. Negative [FC89a, VW96, AG91, CGS23, CZZ19, Gup79a, Jai74, LLZ19, LCY15, MNvRS10, Puk81a]. negatively [BTV21, LMS06, SW16, WDZH14, WZXH15]. negativeness [DH81]. Neighbor [MU99]. Neighbour [Mon96, Haj84, HT16, Lie93, WS22]. neighbourhoods [Haf82a, Haf82b]. nekotorikh [Ano75c]. nelineinoi [Ano74]. Nelson [ZR10]. ness [TZ91]. nested [BH87, DR92, FMTZ07, HEHM13, Kha89, KN90, OM11, PPD07, Sei78]. nested-error [HEHM13]. net [Ciu20, LJ22]. Network [Aby72, Bak73, Bak76, Dew76, FM89, KK14, Kla72, Lei70]. network-based [KK14]. networks [Pol86]. Netzplans [Dew76]. Neuendorf [HW74]. Neuendorf/Hiddenseem [HW74]. neural [KK14]. Newton [DdMT20, STS85]. Newton-like [STS85]. Neyman [Thä75, Bed86, Kon87, Thä75]. Neytman [IJL90]. Neytman-type [IJL90]. nezavicimikh [Ano87j]. nichtlinear [Wei70b]. nichtlinear-einparametrischen [Wei70b]. nichtlinearen [Ehl75, Ehl76, GR72, Hol76, Lip75b, Lip76, Pré74a]. nichtlinearer [ES73]. Nichtparametrische [Sch87a, Sch89a]. Nikodym [BO01, DGL16, JO97]. Niveaudurchagangszeiten [Fra77]. NN [Col85a]. no [GOV17]. Noise [Tur90, IF03, KP94, LZW18, Mei06, NS22, NB16, Sch85]. noisy [Fav14, Sty75]. Non [BW14, Col85a, DSK85, DH81, EMF19, FGL17a, GCB94, HA02, HBCL14, Kre18, Lee08, MNvRS10, MS98, OK02, RA02, Ros82, Ryc09, TA90, AG91, Ano75a, ABD93, BK77, BNAA13, Bel83c, BLO03a, BV86, Çet21, Dau82, DM09, Del09, EL79, Ger75, GW84, HG05, HH16, HGLS82a, HGLS82b, HGLS83a, HGLS83d, HGLS83b, HGLS83c, HCW16, Hue86, JH09, KP15, KFPG21, KZ13b, KFH16, LO86, LDF17, Log87, LH22, Nad80, NKI04, NM18, Puk81a, RAM^{+}17, Rao83, RN18, Rév77, Ruk07, SDKM18, SA23, SKD21, Sto90, TSG20, Tan11, TIQ10, VO75, WH20, WWZ22, Wie92, Wie88, ZL14, Sch89a]. non-additive [HCW16]. non-central [HGLS82a, HGLS82b, HGLS83a, HGLS83d, HGLS83b, HGLS83c]. non-classic [BV86]. non-competitive [SDKM18]. non-compositional [KFH16]. non-cooperative [Ano75a]. Non-ergodic [HBCL14]. Non-existence [DSK85]. non-full [BK77]. Non-Gaussian [BW14, LH22, SA23]. Non-homogeneous [RA02]. non-identical [Çet21]. non-identical-component [KFPG21]. non-identical-component-strengths [RN18]. non-identically [BNAA13, KZ13b]. non-ignorable [TSG20, ZL14]. non-indépendance [VO75]. non-independence [JH09]. non-informative [RAM^{+}17]. non-intersecting [Sto90]. Non-linear [Ros82, BLO03a, EL79, Ger75, Hue86, Log87, Rao83]. non-mixture [WH20]. Non-Monotonic [MS98]. Non-negative [MNvRS10, AG91, Puk81a]. Non-negativeness [DH81]. non-normality [HG05, HH16, TIQ10]. non-null [NM18]. non-optimality [Wie88]. non-orthogonal [Dau82]. Non-Parametric [GCB94, OK02, TA90, Kre18, Lee08, Bel83c, DM09, Del09, Nad80, Rév77, SKD21, WWZ22, Sch89a]. Non-positive [Ryc09]. non-regular [Tan11]. non-responses [KP15]. non-spherically [Wie92]. non-stationary [GW84, LO86, LDF17, NKI04]. Non-uniform [FGL17a, Ruk07]. Non-Zero [HA02]. nonadmissibility [Lät74]. nonasymptotic [Zie88b]. Noncentral [IKR00, NFF^{+}23, Oui22]. Nonconcave [YLL16, ZL16]. nondifferentiable [MC75]. nonegatively [Mat87a]. Nonexponential [GK95]. nonidentically [KZ13a]. nonignorable [KP15, LY20, SMW22, SYW22, TT22, ZTJ17]. nonindependence [VO75]. noninformative [AAA13]. Noniterative [GKP96]. Nonlinear [And90, BOHJ12, Bun93, BDP99, Cox77, GKTZ92, HDA99, Páz84, Páz87, Páz93, RV97, Tuy76, Vie95, And89, Ano74, Bro93, Bun76a, BHSW77, BS80, Bun81a, BD85, CLS14, CTW09, Dro87, Ehl75, Ehl76, ES73, GPS14, GAB85, GR72, Gro82, GS17, HE12, Has77, Hol76, Hös86, Hos88a, Hos88b, HJM90, IR05, IZ83, Läu89, Lin09, Lip75b, Lip76, MS80a, Nož72, Páz90a, Páz90b, Páz90c, Páz91a, Pol85, Pré74a, ROM15, RV08, Sch82b, SZ86, Sch75d, Sch75e, STS85, ŚS05, SL11, TWZ06, TZ13, TT22, Tuy74, Wei70b, Wis77, XWJ14, Zwa80, Zwa85]. nonlinearly [Kim12]. Nonnegative [BV01, GVV17, Kle80, Ouy92, She14]. nonnormal [Güv06, Jag80, SZ86]. nonnormally [Sch76d]. Nonnull [Lem13]. Nonparametric [AL00, BN00a, BL96a, Bel91, Cav97, Col85b, Dev89, GOV17, GGS14a, Gol91, HK87, Iga20, JO95, Jur75, LLK98, LV18, Lou99, Mai91, Mar97, Neu92, Neu94, Neu97, Neu02, PDG13, Raf88, sS14, ST95, TKTM15, Tua81, Vie02, VFGM04, WLW16, AL18, AAFG23, Ant84b, BBL20, BCP99, Che12, CT14, CD09, GJV16, GP19, Hal91, Hel81, Kie17, KPH^{+}16, LF22, LQ12, Mal21, Mie91, MSW03, Nie06, PP83, PD91, SV89, Sch13, WZXH15, WYLW22, WNBH17, Win81, XB17, YLWL22, Zie01, ZZ05, Sch87a]. nonresponse [LY20, SYW22]. nonstandard [Pho87]. nonstationary [HE12, LR17, Mar19]. Nonuniform [Pad96, Pad89]. nor [GB23]. Norm [KN98, AC23, BJ14, Ouy92, SDC19, YFZ11]. Normal [ESSB93, GON95b, LPC97, LH91, Luc99, PSCC95, Sch79b, SS04, Ste98, AGS15, BG19a, BC23, BB13, BP18, BS10, CLS14, CFR89, CHA15b, CFS17, CS06, CT03, CG05, CRCN14, DK12, Exn76, FE85, FH01, GSTUM02, GSB06, GSB11, Gui86, GG90, GS04, HG08, HSG13, JC22, JT22, Kam89, Kha15, Kim05, Kim07, KK15, KP74a, Kle76, KZ15, KA13, KR07, KG13, Kun17, LLBG10, Lät74, MFBG14b, MFTFEO22, MM21, MA86, Oui22, PJ05, PN13, RLAV13, Rod88, Ros91, Roy87, SGG04, SDJ16, She03, SK05, Sta84, TB01, TT10b, WZW21, WBL06, WN12b, Zha18, ZT15, ZHG12, ZRAN17, dQLS16, vR88, SGMFB18]. normal-beta [BG19a]. Normality [BDG95, MO03, Bel80, BGAL01, BR83, Del09, Dem14, DdMT20, Ger80, HG05, HH16, HW88, Hei86, Kle80, LMV17, MMdP09, Mas19, Net90, NFF^{+}23, Que86, Sch75b, SX20, TIQ10, WB05]. normalization [GG15, Jia07]. normalized [BM84, pZ15]. normalizing [JH09]. Normally [HR96, Ric84, Zin81]. normals [Kim16]. normalverteilen [Ric84]. normalverteilten [Zin81]. Normalverteilung [Exn76]. normed [Röd87b]. Note [AN98, Bed85a, Cha96, GH91, Hof95, Koz96, KM96, Läu89, Log87, RL98, Ruy98, SS98, Sei78, SL95, ARN12, Are12, BL78, BL82, Blo03b, Bun81a, BR83, CS10, CT73, DK84, Dro87, DJ86, Duv17, FN13, GHB02, Hav84, Hos88b, Jen73, Joz14, KZ20, Kle76, Kle77b, Kor84, Led84, LR17, Lie93, Mat83, Moe84, Nku23, Pil86, PS84, Pol08, RA01, Rus91b, SK89, Sch76a, Sta84, Ter75, Tik09, WK85, WG92, ZK16, ZN14, Zie89, Zwa83]. notes [KM78]. notion [FPLS12]. Novel [AWJF22, Els22, YLWL22]. November [NS10b]. noyau [Pon86]. Nuisance [Alb84, Bak84, Hab89, KM78, SD17]. null [BMP15, BR19, GVMS09, NM18]. Nullstellensuche [JN75]. Number [BS13, HDA99, KFL97, AA12, AF14, Cey20, Ciu20, Ege80, Fin94, GJHN18, GC04, Kam89, LM79, Los23, MF17, NK88, Pol08, Sch82c, SMLL18, Thr77]. numbers [Ano19, DB10, Dem14, DAA20, Deu72, HCW16, IDB12, LaM17, PP94, Sch22, Car88]. Numerical [Fou97, HKMM98, BH86, FN77, HBN86, ST81a, Ano73a]. numerischen [FN77].

O [Ano72a, Ano73b, Ano75c, Ano81e, Ano73f, Ano75a]. O. [Sch76a]. O.Bunke [Sch76a]. Ob [Ano73c, Ano76, Ano87j, Ano73b]. Oberer [Zin81]. Obituary [GNS15, NS10b]. objective [Ric76b, Tam74b, Wei70b]. observability [PD80]. observable [Fäh81, KZ20]. observation [BN82, FM89, Kut84]. Observations [Bis95b, EA90, Fer99, HDA99, JRM99, Ste98, AA03, AF14, AA08, BS13, BP16, BBKP19, Bat10, BTV21, CFR89, Cip91, Dac11, DB10, Dem14, DJ17, DAA20, DK14, FLL16, Fav14, Fin94, GC90, Huz18, IV90b, IV89, IDB12, JRM04, KYR89b, LWS21, Mil86, MK83, NAHJ03, Pap16, Pon06, PR89, Sär95, SS15b, SZ17, SK05, Thr77, VO75, YN23]. observed [AG23, DI11, HT15, Nie17, Wit93]. observing [Nät84a]. obslyzhivaniya [Ano75b]. obtain [Mej78]. obtaining [Rei84]. oc [Rüs81, Ege80]. oc-function [Rüs81]. occasions [KG81a]. October [Gir74]. odds [Bow21, KHN20, MB18, NRD08, SSM07]. Odna [Ano79e]. odnoi [Ano87j]. of-fit [Geb84]. off [PP83]. ohne [Ric76b, RS74]. ökonomie [Lom74]. ökonomischer [Mat75]. Oktober [Gir74]. Olech [Gus71]. Olkin [JK13, WG21]. Olman [Hof95]. Omnibus [HM07, ELV22, Hen93]. Once [BNP87, And79c]. One [BP17, Cha96, HA02, Vän98b, Zie95, AALA20, Ano87j, Are12, BC23, BB13, Cou21, Das87, DdMT20, Fer07b, FE85, Flo93, GL94, GKM18, GT70, HG05, Hof85, HW71, LKM14, LTV03, LCY17, MNvRS10, Men73, Mur15, NAHJ03, Nož72, OM11, PEB15, Pet86, Rei86, TL15, Wei70b, Wei71, ZLLX08, Pad81]. one- [Fer07b]. One-Armed [Cha96]. one-armed-bandit [Rei86]. one-dimensional [GKM18, LKM14, Pet86]. One-parameter [BP17, GT70, Hof85, HW71, Nož72, Wei70b, Wei71]. One-Sample [Zie95, AALA20, NAHJ03]. one-shot [BC23]. One-Sided [Vän98b, Zie95, FE85, Mur15, Pad81]. one-step [DdMT20, Flo93]. One-Way [HA02, Are12, Das87, HG05, LTV03, MNvRS10, OM11]. ones [BN78]. only [LCY17]. operating [BP93, Fre74, JRT21, KJKR76, Sto71, Sto72b, Zäh74]. operating-systems [Sto72b]. operation [Fle71, Sie74, Sto72a, Sto73, KS76a]. operational [SW89b]. Operations [PGT02, Ano73a, Ano73d, Pie71, Pie72a]. Operationsforschung [Pie72a, Pie71]. operator [BV86]. operatoren [EN75]. operators [DGP91, Dry84, EN75, WN14b]. Operatsii [Ano73d, Ano73a]. opinion [Bic18]. Opportunities [Tam76]. Optimal [Afs85, ASSH13, Arn95, BK71, BM01, BGV06, Bie15, Bie16, BCP99, Dav94, Det93, Det95a, DMP07, DMP10, DK14, Fle76, Gor19, GGHS03, GGHS21, HK87, HM88, Jac13, JLM23, KCNB11, KD74, LT85, Läu76, LFGR11, MD03, Mel78, MG96, Nät84a, PCKMK17, PM16, Ryc04, SMK22, SS15b, SDKM18, Sch96, Sei79, Thr77, TDY08, VI90, VK20, YH15, YLWL22, ZW20, Zha97, ZYW21, Ano73b, Arn91, Bag90, BM23, Ban80, BNP87, BB07, BB13, Bit75, Böh81a, Böh81b, Bun77c, BS06, Bur08, CBP23, CGK17, COQ21, DB74, DM06, ETF19, Fäh78, Fäh81, FM89, GP82b, Gov92, GM78, Gri80, Hab89, HR80, HT71, HP73, HDC20, HM89, KHN15, KS10, Kra76, Lät74, Lei70, MMRA07, Mat89, MK83, PMA17, Páz80b, Puk81b, RH80, Röd04]. optimal [ROM15, Sch77a, SS74, SSG76, Ull76a, Von87, VYD83, Wäc86, Wit78a, Zie01, vW72a, vW72b, JN75, Läu70, Läu71a, Pfl80, Tou70a, Wol75]. Optimale [BK71, Wit78a, Ban70a, JN75, Läu70, Läu71a, Pfl80, Tou70a, Wol75]. optimalen [Fäh78, Fäh81, Gri80, vW72a, vW72b]. optimaler [DB74, Gir71, GM78, Ull76a]. Optimality [KMSS10, OM91, Pla02, Wie86, BB22a, Bed81, CAK13, Gaf85, HS91, Hos88a, IS84, Jac91, MOS86, Nät75a, NR81, Nus77b, Rei82, RDLF03, Sch75c, Son82, Wie88]. Optimally [RH13]. optimalnoi [Ano73b]. Optimalprobleme [HT71, HP73]. Optimalregelung [SS74, SSG76]. Optimierung [Ber76, Ehl75, Ehl76, Els75, HW70b, HW70a, Hol73b, HW74, Lip75b, Lip76, Pré74a, Seb76, Tam76, Wei70a, Wei70b, Wei71, BNS73, BG74b, EW73, Foc73, HW72, Nož71, Pie70, Pie73, Pie75a, Pie75b, Pie76, Sch74a, SSG76, Som71, GR72]. Optimierungsaufgabe [Lud73]. Optimierungsaufgaben [Ric76b, Ric76c, Sch75d, Sch75e, BTS71, Hol76, ST75, Sch75f]. Optimierungsproblem [PW76]. Optimierungsprobleme [Gud70, Lip75a, Trö75, Bär72, Bär76, ES73, GT70, Hol70b, Mel76, PVW74, Ric76a, Sch70b, Tam74a, Tam76]. Optimierungsproblemen [Nož70, Nož72]. Optimierungsproblems [Tam74b, HW71]. optimising [FdHP03]. Optimization [CF95a, Bär76, BTS71, Ber76, BCFF88, BG74b, Ehl75, Ehl76, ES73, EW73, Foc73, GR72, HTX23, Har75a, Har75b, Hol70b, HW71, Hol73b, Hol76, Lip75a, Lip75b, Lip76, Lud73, Mel76, Nož71, Nož72, PVW74, PW76, Pie70, Pie75a, Pie75b, Pie76, Pré74a, Ric76a, Ric76b, Ric76c, Sch70b, Sch74a, Sch75d, Sch75e, Sch75f, Seb76, SSG76, Som71, Tam74b, Tam74a, Tam76, Trö75, Wei70a, Wei70b, Wei71, Zil78, HW70b, BNS73, Els75, HW70a, HW72]. Optimized [Web89]. Optimum [DR92, GK77, KN90, Gir71, PMA17, Rui85, Ban70a]. Option [IJZ08, Yat17]. optional [AMP21, MP23]. Ord [ABP18]. Order [ASKS16, BM02, Bla93, CG95, Eks99, FR95, HKMM98, LBSW19, Nag92, Neu92, PV98, Ryd95, SR91, SR94, SP98a, SP98b, VV00, Ver92, Vie95, AAFM07, Ahm01, Ahm11, AB10, AMB11, AR13, AV12, AA11, AA13, BT08, BJ85, BB93, BCKS01, BCK05, BCD11, BOHJ12, BCJ14, BT16, BB22a, BM23, Bal03, BB04, BC14, BN14, BNH20, BBK12, BBK14, BMK16, BK17, Bie06, Bie08, BMM13, Blo07, Bos88, BC12, CZZ17, CLS14, CMLBSM12, CMLBSM14, CBP23, CZZ19, Cig13, CCM18, Cra06, DB10, Dem14, DJ17, DL16, EBQY17, FB16, FZLB22, FK14, FL10, GKFMM10, GCF04, Gor14, GGHS03, Gün80, Gup79b, HSBB23, HS79, HS15, HM89, IM14, IRB14, JC22, JR13, JT22, JB83, KC01, KS83a, KS83b, Kla03, LAO21, LO86]. order [LYB19, LH22, MR73a, MR73b, MR75, MGM20, Oko17, OM10, Pol08, Pon84, Raq04, RA06, RM17, RT13, Ros72, Ryc09, Ryc15, ST81a, SP01, Sen11, SdSSL15, Tak09, TPL19, Tor17, UI13, VK20, VDB13a, YN23, YA12, ZZQ16, ZQB16, Zwa85]. Ordered [BLT20, GD96, MS94, CFS17, DSK85, Miz17, NJKK16, PK18, PKP21, RG84, Rüs81, Ryc21, SMK22, ST23, TIT08, Thä76, TMY05, TT22]. Ordering [BKHB22, EZ92, FB16, FLL18, KT96, KC17, MGM20, Rüs91a, RR91, CZZ17, FZB17, HSBB23, KDM22, NRD08, ZB14]. Orderings [FZLB22, Dab81, Dew76, Foc75b, Nät75b]. orders [HX20]. ordinal [SS21]. ordinary [KJ14, WLB04, Zäh14]. ordinate [Tro91]. Ordnung [Gün80]. oriented [SMP80]. oriented-fibre [SMP80]. origin [Kum13]. Ornstein [AM15, Glo01]. Orthogonal [ABP18, ASSH13, Dau82, EP17, HTHF16, PH16b, RALN93, Cap73]. orthogonale [Cap73]. orthonormal [RW84]. orthosymmetrie [SN81]. oscillation [DDLN17]. ostanovkye [Ano73b]. ostatochnovo [Ano87j]. ot [Ano81e]. other [BR83, MR73b]. otsenivaniya [Ano76]. otsenke [Ano87j]. out-of-sample [GPS14]. outbreak [FAP10]. Outcomes [DY02]. outer [LH15]. Outlier [Mat85, Sch89c, WWZ22, BT16, BHN15, CZZ17, GHB02, Mat89, ZQB16]. Outlier-prone [Mat85]. outlier-resistant [Mat85]. outliers [Bha07, Kum13, RG09]. outline [Wer78]. overall [Mak06]. overdispersion [JJS06]. overlapping [LBSW19]. overlook [CLCP17]. overview [FM72]. ozone [BT04].

P [Dro90a, Hec86, War85b]. P. [Mai91]. package [Pet86]. Page [Fre15]. Pair [And76b, Tak86]. pair-potential [Tak86]. paired [AAA13, GGHS03, GGS14b]. pairs [BK86, Haf82a, Haf82b]. Pairwise [SJK97, RW84]. Pairwise-and [SJK97]. panel [JS14, Shi15]. panels [LLW17, SO09]. Panne [GT70, GT70]. Paper [CF95a, Det95a, Sha96a, GH91, MB92, Páz90c, Tik09, Zwa83]. paradigm [BR19]. Paradoxical [CF98]. parallel [KHN20, MF17, PN10, YDZ13, ZB11, ZB14]. parallelism [RG84]. parallelized [GBS20]. Parameter [Atk75, ALS11, BN81, BPV11, Bel83b, BBR05, BPR00, BHSW77, DM95, Fal95, Fav14, Fei72, FM01, GH89, GKP96, Gil78, Gir71, Gla76, Glo01, Höp98, IRV98, KN98, Kut84, KP94, LLC16, MP23, NS22, PP92, Tam74b, ANN11, Als88, AM15, BR90, BK17, BZ15, BP17, Bun73b, BG79, Bun81a, BHS76, CKM09, CY12, CS10, CBC12, Col03, EJ81, Faz08, Fer07a, Fer07b, GG91, Ger79, GB17, GCF04, GT70, Hof77a, Hof85, HW71, Hus88, IS84, JSW22, KFPG21, KMM15, Lau85b, LCMA16, LZW18, Log87, LST18, MMP16, Müh80, Nät75b, Nät84b, Nož72, PAW05, Pet86, PKM78, Pol85, PV85, RNT13, Röd76, Ryc87, Sen11, Sha83, Sta90, SD17, ST87, TPL15, TPL19, TMY05, VO78, VI90, Waw81, Wei70a, Wei70b, Wei71, WWC^{+}10, WY13, WA22, Zie85]. parameter [Zie88b, Wei70a, Wei71]. parameterization [Ano74]. Parametern [Koc74, Wen76b, Wen76a, Hol75]. Parameters [Dor85, GD96, Hof97, Jon76a, MS94, MK95, Neu79, VV96, And76a, Bak84, BZ14, BBK12, BMK16, Bun75d, CLZ15, CM05, CT07, CC12, CBC12, DSK85, Fei72, GAJB13, Gil78, Gir71, Gla76, Gno79, GM88, Gri80, GKM18, Hab89, Hen79, Hol75, HW07, JT22, Kha12, KM78, KP74a, Koc74, Kon87, KL82, KR78, Kra86, KS93, Läu75b, LM79, LMV17, MB21, Mus77, NK13, NK23, ÖA21, PK18, PKP21, Pil86, Pin74, RALN93, RG09, Rus91b, ST23, Seb76, She88, Sie86, SD17, TSG20, Try86, Try88, Try91a, Try91b, VPR21, Wei70a, Wei71, Wen76b, Wen76a, WHT03, ZW20]. Parameterschätzung [Atk75, Müh80]. Parametric [AHJ99, CT07, GCB94, LLK98, Mec84, OK02, Tan91, TA90, AKN05, BB22b, Bed81, Bel83c, BB91, BCCF14, BAM18, BF85, Cap19, Col85a, DM09, Del09, EMF19, Flo93, Foc73, Gre87, Gud70, Gud76, Jas81, Kre18, Lee08, LQ12, Nad80, Pon84, Rév77, Sch89a, SKD21, Tam76, Tar88, Tit83, TPSR17, WT20, WWZ22, Wei72, Wei74, Hen89, Hol73b]. parametrische [Foc73, Hen89, Hol73b]. parametrischen [Tam76]. parametrischer [Gud70]. parametrizatsii [Ano74]. Pareto [AFL08, CB02b, DB13, Gup79b, HVRI20, JD20, Lik85, Mic09, Mou03, Nad05, NAHJ03, OJM15, RAD07, Ros72, RH13, SK14b, Zha13]. Paris [BT04]. part [FL03, Bel91, BT97, Ehl75, Ehl76, GM97, HGLS82a, HGLS82b, HGLS83a, HGLS83d, HGLS83b, HGLS83c, Küc82a, Küc82b, Lip75b, Ric76c, Sch75d, Sch75e]. Partial [GK99, Pan84, Rüs91a, RR91, Sto96, And79c, AFV15, ARN12, Bei75, Bos93, GW04, GGS14b, JR06, Lys76, PD80, Pon06, SS79, TX17, Wu16, ZR10, ZWYG14]. Partially [FX16, AAR18, BQSR20, Bre05, CM15b, CC12, DT16, Fäh81, HC20, KY22, LWS21, LQZF17, LLL17b, NK13, NJKK16, RH20, SMP80, Tab13, Thä76, XY16, YLL16, YLY17, YLY18, YS23]. Partialsummen [Lys76]. particle [GP07]. particular [Jen73]. partition [GJLT15, WC21]. partitioned [TT10a]. Partly [SL95]. parts [HTHF16]. Parzen [NSK12]. Passing [Han96]. past [AB12, DZ22, GK17, KDM22, KAR14, MH22, PN10, WLB04]. path [Bak73, Kla72]. paths [Bak76]. pathway [Tho13]. pattern [Sto88b, WQO22]. patterned [LH15]. patterns [AN13, CCM18, Hei84, Sto94]. Paul [BFN17]. PBIB [Sin81, SD83]. PCA [BVP06, DDT89]. peaks [MBF^{+}20]. Pearson [Bed86, CV78, Sta20, Thä75]. Penalized [AAFG23, Ars16, BQSR20, BR15, Sch13, WW22a, AC23, AÖ23, ÖCA15, YLL16]. penalty [Ars16, Böh81a, ZL16]. pension [SSG76]. Percentile [VW96, FPLS12]. percentiles [WN13a]. Performance [CY12, Tab13, Dev89, DW93, Wäc86]. performances [TE15]. performing [Els22]. period [CT73, RS74]. periodic [Sak71, Zäh74]. periodicity [Cip91]. periodischer [Sak71, Zäh74]. periodogram [FRLLT18, Wit78b, Wit81, Wit83b, Wit86]. Periodogramms [Wit81, Wit78b]. perishable [KA88]. permutation [LSS12b, PSB23]. persistence [QLY21]. persistent [GS21]. pertinent [JBD14]. perturbation [Fin87]. Petersburg [Ber17]. Petersex [Läu75a]. Pfeifer [AV12, BG14, SK20]. Phase [SS94, CD08, KG81a, KG81b, Mag72, MN08, Tou75, ZL03]. phase-randomized [MN08]. phases [BB91]. phi [BMP15]. phi-divergence [BMP15]. PHR [CZZ17]. physical [Kou11]. pick [GJK^{+}16]. pick-freeze [GJK^{+}16]. pictorial [Sto79]. piece [Kim05, KA13, SD17]. piecewise [BV01, Eva75, LLL17b, Oko17, Tam74a]. Pillai [Sch76d]. Pillai-tests [Sch76d]. Piterbarg [HT15]. Pitman [AR13, HSW85, Joz14, JBD14, VDB13a, VDB13b]. Pivot [Bun77a, Sol19]. Pivotal [Páz91a]. pivots [Are12]. plan [Cha15a, CF21, CF22, LC17, SMK22, WHT03]. planar [Han83, McC06, MS80b, Mec81, SMP80, SH86]. plane [FGK19]. planes [Ros75]. planirovaniya [Ano74]. Planner [Sto91]. Planning [BK71, Ano74, GH83, Lei70, Mön72, Ric75, Tit70, Ull76b, Wit78a]. Plans [OM91, CBP23, CAK13, FE85, Jac13, Kum85, MOS86, Von87]. Planung [Tit70]. Planungsmodelle [Mön72]. Plug [Yat19, Zäh14]. Plug-in [Yat19, Zäh14]. PMC [Zie01]. PMC-optimal [Zie01]. Point [Bel91, Bil97, Gol91, Las91, Mi96, OK02, Pon02, RV97, Sto91, VN98, Ano75c, BL11, Bat10, BFR84, BR19, Bök85, CS06, Ciu12, CCM18, DI11, DL16, Duv17, Fik88a, Fik88b, Fre15, GVMS09, Haj84, HTX23, HS79, Han83, Hei84, Hei86, Hei18, HR93, JN75, KK14, Laz16, LL85, NL03, NS14, Ohs83, PN90, Pon86, Sär95, SK05, Sto83b, Sto88b, SG91, Sto94, Str82, Tak86, UI13, VWL^{+}09, WY04, ZW10]. points [Arn92c, Ber00, CG04, Ped86]. Pointwise [Neu97]. Poisson [AMSV12, Bed04, Bet87, CB04, Dac11, Faz08, Faz14, FS81, GG15, GMLS16, GML18, Gup79a, HOCC14, Hei88, Hei91, HKMM98, HL13, JJS06, KS86, KMSS10, LSS12a, Led78a, LUK13, LUK14, LLC16, LL85, LH91, Luc05, MO03, PJ17, Pog15, Pol86, RA02, Sch85, Sch79c, Sch81, SKV18, TSG20, Tho84, WS97, WD22, XW10]. polar [EN74b]. Polarfunktionale [EN74b]. policies [Bei81]. Policy [Tij76, Bei76]. Poljak [Ehl75, Ehl76, Ehl75, Ehl76]. Polya [MB21, JR86]. Polyeders [Lom70]. polygon [WW22b]. polyhedron [Lom70]. Polynomial [AAJ97, Mag95, Wis77, Zar78, EJ93, KRC19, PH16b, Ron76, WLF13, ZWL14]. polynomialen [Ron76]. polynomials [ABP18, DMP07]. polytomous [GNP08]. ponctuel [Pon86]. Pooling [Ahm93, Hor88]. Population [Arn95, MK95, Arn91, Blo07, BAO12, CT03, Cig13, EBTB21, GM88, GC04, GK93, KSS11, Lät74, Sah87, SLL15, VDB13a, Zha12]. Populations [GK99, Hol91, VV96, VV00, AMSV12, BB13, BRZ93, CM90, GPK23, HC20, JC22, JR13, JT22, JBD14, KR78, Liu11, MM21, MK19, OB19, RR04, Rod88, Ryc04, Ryc09, ST23, Sen11, SKPK21, VP04]. portfolio [BS07b, GOS10, KMM21, KS10, LH22]. Positional [Bor70]. Positive [ETD99, Gne98, Led84, SR02b, And90, Düc75b, GR06, JT22, MR12, Röd87a, RR02, Ryc09, Sha86]. Positive-Definite [ETD99]. positiven [Düc75b]. positivity [DM09]. posledovatyelnosti [Ano73b]. Possible [Str95]. Posterior [Fer07b, TH08, Bic23, MMdP09]. Posteriors [ACS96]. postroennii [Ano73b]. potential [Tak86]. potentially [LFGR11]. potentials [Fik88b]. potoki [Ano75b]. Potthoff [BK77]. pour [Pon86]. Power [CH08, CB02b, KP15, KB00, Koz02, KG13, TMY05, TT10b, AG09, BMMP16, Bet87, CF89, DLO22, FC89b, Gup79b, GS82, Kat89, MFBG14a, MFBG14b, Mat89, Mur15, PGT05, SGMFB18, SS74, SH09]. Power-divergence-type [TMY05]. Power-linear [KP15]. Power-normal [KG13]. power-quadratic [PGT05]. powerful [CLZ15, Mil86, ZX17]. powers [ACS08, JT22]. Pozitsionniye [Bor70]. PPS [AK85]. PPSWR [MB92, MB91]. Practioners [Dro90a]. Prague [Man72]. Pratt [Moe84]. precautionary [Laz16]. precedence [BGLS15, CBF23, SB17]. Precision [GON95a, GON95b, MH97, BG11, She03]. predator [Nku08]. predator-prey [Nku08]. predelnoi [Ano87j]. predicting [Ter75]. Prediction [AB10, AHJ99, AJ97, AEW94, GB78, Mou02, MJA02, Nig90, AAFM07, AA12, Ahm11, AJM12, AMB11, AHJN01, BK79b, BB22a, BM23, Bon88, Bun77c, Cip81, Col85a, Dry75, GFCO04, GLR23, Gün77, Gun88a, HS83, IM14, KHN16, KY22, Läu70, Laz16, Mou01, Mou03, Nie06, NW12, NAHJ03, STY04, Thr77, Thr79, Tzi87, VK20, ZYW21, Läu71b, Läu71c, Tou70b]. predictions [Liu11]. Predictive [CM15a]. predictor [Yoo16]. predictors [KSS11, KM17, LH13]. predlnaya [Ano79e]. preliminary [CY12, MPPZ05]. Presence [RA02, ZM95, BS82, Dav94, JK08, KZ20, KM78, KB85, LZW18, LCY15, Mei06, RG09, SD17, VO78, WW21]. Present [Neu79, dUÁRC04]. Preservation [AB05]. preserving [FK14, YT01]. Presmoothed [JIP10, JC07, dUÁRC04]. pretest [Nku12]. prevalence [BCDD11]. preventive [Bei76]. Previous [vW74]. prey [Nku08]. pri [Ano73b]. pricing [Ano73b, IJZ08, Yat17]. prilozhenniye [Ano75b]. primal [Bit72]. primal-dual [Bit72]. prime [BG19a]. Principal [BCFF88, Kle73, Özk09, ZZW^{+}19, DDT88, FR84, FP92, ID03, KK19b, LG05, May11]. principle [Ano73e, Bun77a, GGL15, MW14, Nož71, RW79, Sch70b, Sch73, Sch75d, Sch75e, SX20]. Principles [Ste95, Lei70]. Printsip [Ano73e]. prinyatie [Bor70]. Prior [AS98, BZ15, Ahm93, BG19a, BN81, BS85, Bic18, BHS76, CKM09, Els22, GVGP13, MB21, Ryc21, Wäc86, YLWL22]. prior-data [Bic18]. priori [JN75]. Priors [ACS93, ACS96, BT97, GM97, AAA13, AINJ14, CM13, CM15a, OM10, OM11, RAM^{+}17]. pro [BB78]. probabilistic [Pré74b]. Probabilities [AAJ97, GKFMM10, IKR00, AKN05, Ano73c, Bic23, BH86, Bun85, GS11, Got74, GJ19, GS82, HCW16, JK76, Kin75, Kum87, MPPZ05, Moh86, Roy87, STY04, SLL15, Vís86, Wol84]. Probability [ETD99, Fak96, Gne98, Gup82, OM11, PKN72, AK85, BP93, Bun76b, CM13, Cup21, DK84, FdHP03, HHM15, JD20, KKA84, LUK14, Lie90a, Lie72, Luc05, MR73b, OM10, PW15, Ros91, RW84, Rüs81, Rüs87, Sha18, SKV18, Sta17, WNBH17, Wer85, ZH16, ZTJ17, War85b]. probability-generating [HHM15]. probes [VCO09]. probit [AK20, GGHS21]. Problem [Koz96, RV97, vW72a, vW72b, AR13, AK75, Ano73b, BK79b, BSP76, BM18, CLZ15, CDD17, Foc71, Foc73, GH83, Gno78, Goe75, GVMS09, Gov92, Hel81, Hel93, HW71, Jen73, Jon76b, Juh89, JS82, Kér72, KS92, MM22, Men71, PW76, Pré73, Rei82, Sch70c, Sin13, SB17, Tam74b, Ter72, Ull76b, Van88, VWL^{+}09, Vid23, WZW21, Wei70b, Wen76b, Wen76a, Yu04, ZHG12, Zie89]. Probleme [Atk75, BM71, Dew73, Ste76, Sto71, Wei70b]. Problemen [Hof72, vW74, Sch73]. Problems [Atk75, BM71, Ban80, GV95, Kov74, Ste76, Stu96, Ano73a, Ano75b, Ban70b, Bär72, Bär76, Dam76, Dew73, Ehl75, Ehl76, ES73, Fra79, GR76, Gud70, GT70, GJ73, Hen80, HT71, Hof72, HP73, Hol70b, Jür84, Koo86, Kos80, Lip75a, Lus87, Mat87b, Mel76, Men73, MESS70, MC75, Nož70, Nož72, PVW74, Pfu78, Pré72, Ric76a, Ric76b, SS81, Sch70b, Sch73, Sch75d, Sch75e, Sch75f, Tam74a, Tam76, TH08, Trö75, Ull76a, Wei72, Wei74, Zha16, vW74, Sto71]. procedural [Fra77]. Procedure [GJML00, BS82, BS75, BG78, BD85, DK21, DJ86, DS91, Fre15, GR17, HAMAZ96, Har75a, Har75b, Her81, NL05, Que86, Rei84, Rod93, Sch73, Sta90, UI13, YL17]. Procedures [CG95, CR99, CB02b, Kal97, Mag95, BPH88, Bed89, BS75, BR83, BF85, CM15b, CT03, Dro06, Exn76, Fra77, Fur15, GMP16, Gün83, Hen80, LT85, Mak06, Sti78, Väd77, WFK82, YH15]. Proceedings [War85b]. Process [Che97, IV02, Las91, MG96, MS98, Sto91, Tur90, AAB18, Ano73c, ADR14, AM15, BBKP19, BNS73, BN80, Bor73, BG14, BCP99, BGL15, Cup21, DI11, Dew73, EW75b, FL03, Fav14, Fle71, FR86, Glo01, HE12, Hei91, HR93, HL13, IV90a, IF03, JSW22, KK11, KK81, Koz86, Laz16, LC17, LMS06, LL85, Mag88, Mat81a, NK88, Pau12, Pon86, Pon04, Pra86, Rao85, Ras06, Sch13, Str82, Tak86, TSO07, Thr77, VI90, VYD83, WD22, WWLH19, Zan15, ZXT11]. Processes [And76b, Bil97, CHP03, FZF97, FR98, GM96, IR91, Lai00, Mag95, Mou02, RA02, Ahm74, Aly87, AG23, And79a, And80, And81, And90, AG91, And78b, And78c, And79c, AV12, BL11, BLO03a, BG16, Bit75, Bla14, Bök85, Bos88, Bos93, BE12, BZ13, CC01, CB04, Cip81, Cip84, CFC10, DDLN17, De 22, DRW17, DJ17, DWHH19, DL16, DS91, EMF19, Fäh78, Fäh81, Faz08, Faz14, Fik88a, Fik88b, FZ89, Fra82, FM01, FN13, GLR23, GC90, GV79, GV81, GW84, Góm18, Haj84, HS79, Han83, HT15, Hei88, HP08, Hei18, HB09, IV90b, IV89, IS84, Kaw15, Ker84a, KW15, KK14, Kir76, Kon87, KK81, Küc82a, Küc82b, Kum85, Kut84, Läu71a, Läu71b, Läu71c, LR17, Lie86, LT16, MP10, MMdP09, Mas19, MS80b, Mec81, Mic86, MW80]. processes [Nag83, Nät84a, NB21, Ohs81, Ohs83, PMS81, Pol86, QLT17, Rab15, Rao83, Rus91b, Rya14, ST81a, SS15a, Sär95, Sch85, Sch70a, Sie86, Sie87, Sin84, SMP80, Sto83b, Sto84, Sto88b, Str82, Tho84, Thr79, WS97, Wan07, Web06, WK13, WK85, WFK82, Wit76a, Wit76b, Wit83b, Wit93, ZW10, Man72, PT83]. processing [CRCN14]. processus [Pon86]. Product [NLP95, BSS18, Hei88, LH15, NK05a, Nad07, Ric75]. production [Arn92c, Boe72, DJ09, Kum85, Ric75]. Products [Mos85, BK77, Dry84, NRS07, Tho13, WN13a]. Produktionsplanung [Ric75]. Profile [WK15, WWZ22]. profiles [GGS14b]. profit [Fäh78, Fäh81]. Programming [HW74, Pré74b, Ano73e, Bär72, Ger75, Gud70, GT70, Gud76, Kov74, Men73, MC75, Nož70, Pie73, Pré72, Ros75, Tam76, Wei72, Wei74]. programmirovanii [Ano73e]. programs [Ber75, ST75, The72]. progress [Fok11, Kle80]. Progressive [BCKS01, BSS18, LC19, Bur08, Çet21, KFPG21, MK19, PBK11, SP01, SPTB18, SKPK21, VDB13b, WG21, WWC^{+}10]. progressively [BCK05, BCD11, CBP23, MAS06, VDB13a]. Projection [BRKO99, Bie06, Bie08, BW80, DHN20, ETF19, WQO22, WQO23]. Projection-based [DHN20]. Projections [Rüs87, GJHN18, GR17, KK19b]. Prokhorov [War85b, Haf82b]. Prokhorov-neighbourhoods [Haf82b]. promoters [LDF17]. prone [Mat85, ZWYG14]. proof [BV86, DH77, Hei85, Lie72, Nož71]. proper [RVM22]. Properties [AB12, BFF00, Bel91, Bil97, ESSB93, Fak96, Hu10, Wei09, AB05, Als88, Ano81e, AS04, AE21, BB22a, BP16, BBKP19, BSL13, Bel83b, BS21, BS23, Bod75, CHA15b, CB02a, Che12, Cup21, DJ09, Dod83, Dom10, Dro93, Dro06, FPPRG16, Fra82, GK77, GLZW12, GJV16, GVV17, GV79, GSB06, GSB11, GJLT15, GT13, Hei92, IDB12, JBC22, JBD14, Kal76, Kie17, KB85, KL82, KT89, KZ13b, KA13, KC17, LO86, Lin09, LBSW19, MN08, Mie91, Mie79, Müh84, NS10a, NRD08, Páz90c, Rei86, Ros72, Ruk07, Sai86, Sak71, SU93, SDJ16, Sto73, Thä76, Tho84, VFS^{+}20, WY04, WS22, Xek86, XWY^{+}20, XWJ14, YLY18, Zäh74, ZFA20, ZW10]. Property [CB04, Pla94, KHNT17, LF22, Pie73, PS84]. Proportional [Ano19, CF98, GK98, LaM17, CL14, Csö88, FLL18, GR17, Hen93, HLsS01, KAR14, KKA84, KHN20, LUK14, Sch70c, Sha21, Wie86, Wie88, YN23]. proportionalen [Sch70c]. Proportionality [FW13]. proportions [DJ17, PGT08]. prostranstv [Ano75c]. protsessa [Ano73c]. proving [Cap19]. proximity [BVP06]. Prozeßbilanzierung [Dew73]. Prozesse [MW80]. Prozessen [Läu71a, Läu71b, Läu71c, GV79, Sch70a, PT83]. Prüfung [HK78]. Pruning [AX16]. Pseudo [Sär95]. Pseudo-likelihood [Sär95]. Pseudorandom [EMG98]. pure [Pie70, Pie73, Pie75a, Har75b]. pure-integer [Pie70, Pie75a]. Purely [TVL96]. purported [Hel93]. purpose [HK78, Pet86].

QR [DGCL14]. quadrangle [Mec88]. Quadrate [Tau84]. Quadratic [Bun99, EW99, GM91, GN98, Imh00, MD03, BS88, BCP99, BM84, CC01, CFR89, Dry77, DH81, Fäh81, GS88, GB23, Gud70, GT70, Gud76, HK00, KP74a, KP74b, Kle75a, Kle75b, KZ78, KT83, LR17, Liu11, MA86, Net90, Ouy92, PMA17, PCKMK17, PS10, PGT05, PHS09, Ric84, Tam74b, Tam76, Wie92, Ber76]. Quadratische [Ber76]. quadratischen [Fäh81, Tam74b]. quadratischer [Gud70, GT70, Ric84, Tam76]. Qualitative [BC99, Mam95, LDF17, ZYZ21]. qualitatively [CS89]. Qualitätskontrolle [Gla76]. Qualitätsunterschiede [Ber74]. Quality [Hec86, FE85, Gla76, PDG13, Ber74]. Quantile [LWS21, QK17, sS20b, SSN17, XCJH19, YLY17, Zie88a, Aly87, Chr21, Ciu16, Ciu20, DJ09, FB22, JS22a, JHYQ18, JLM23, KLF22, LY16, Man22a, MMRA07, MD22, NSK12, NFP18, OB19, Sar82, TT22, ZL16, ZTJ17, Zie01]. Quantile-based [SSN17]. quantile-type [DJ09]. Quantiles [AJV21, IKR00, Lik85, BJ14, BGAL01, CBC12, Fer07b, FdHP03, IDB12, LZW23, QXY23, Ryc21, Sta17, Ver06, VDB13a, WYW21, ZZ05]. quantitative [ADR14, KS76b, Rab15, Win17, ZYZ21]. quantity [Sch70c]. Quartiles [MH97]. Quasi [TS87, Tro00, WL01, GLZW12, Ger73, Guo20, KMSS10, TY12, WLF13, XWJ14]. quasi-convex [Ger73]. quasi-likelihood [GLZW12, Guo20, TY12, WLF13, WL01, XWJ14]. Quasi-Score [Tro00, WL01, KMSS10]. quasiconcave [Pré74a]. Quasiconvexity [Dew76]. quasiellipsoids [KZ86]. quasikonkaver [Pré74a]. Quasikonvexität [Dew76]. Quasilinearization [Bit72]. quasimidranges [Bie15]. quaternion [LBAR13]. question [PS12]. questions [Ano75a]. queue [Cha74, FS81, Goe75, Goe76, KSS76, SD86, Zäh74]. Queueing [Tij76, Ber74, Ber79, Esc84, Goe75, Goe76, KSS76, KK81, Mad76, NK88, Pol86, Ros71, Sin84, Sto76a]. Queues [KK80, RL98, Kal76, Kre80, RL83, Sch79c, Sch81, Sto76b]. queuing [Ano75b, CT73, GJ73]. quick [Hen93]. Quinn [Kam22]. quotient [Har75a, Har75b, KS79]. Quotientenoptimierung [Har75a, Har75b]. Quotiententest [Küc78, KS79]. quotients [Ros72].

R [FD72, Led84, Tik09]. Radon [BO01, DGL16, JO97]. rainfall [NSG07]. Random [AEW94, BPR00, BMM13, BN00b, CMLBSM12, CMLBSM14, CF98, EA90, Fer99, GCB94, HA02, IRV98, JK92, KE96, KFL97, Man72, Nät97, Rat93, Spj77, Str95, Stu96, The72, AL18, AA08, And76a, AOV19, Ano75b, Ano81e, Ano87j, Are12, ACS08, BNAA13, Bar87, BL82, Bed86, BMRPZ16, BTT11, BTV21, BAM18, Bla14, BW81, BN80, BAO12, Cad01, ÇO17, CC13, CLLY22, CD08, Ciu12, CCMW09, CCH87, Csö88, Dal77, DR92, Dem14, DJ17, DAA20, Dik01, DGP91, DST21, Ege80, EMF19, EBQY17, EM87, Fäh78, Fäh81, FB16, FZB17, FLL18, FSV13, Fin94, FMTZ07, Gai23, GP82b, GGS14a, Gün77, GNYX15, Güv14, HE12, HG05, HK09, HKK23, Hei86, Hen93, HLsS01, Hor84, Hös74, Hös76, Hös77, Hoy79, JS14, Jas22, JRZ03, Kam89, KT12, Kha89]. random [KN90, Kin75, Kle73, Kos80, KS79, Kum89, KC17, LLW17, Led83, LSS12b, LCY17, LLV16, LBR23, LBAR13, LLWW20, MGM20, MNvRS10, MYK18, Mec84, Mec88, Mik88, MR12, MF17, Müh84, Nad07, Nät84b, OM11, OB19, PYHT01, Ped86, PGT05, PN13, Pré74b, Pro89, RM17, Ric84, Röd70, Ros71, RW84, Rüs90, SK92, SV89, SS15b, She14, Sie74, Sie86, SV18, SH86, Sto90, Sto93, TZ13, Tho13, TY12, TDY08, Väd77, WDZH14, WNBH17, WBL06, WN13b, WN14b, Wit76a, Wit78b, Wit86, WW21, WYW21, WHVW23, Zäh84, Zan14, pZ15, ZZWW23, Zin81, ZL17]. random-effects [MNvRS10]. Randomization [BB78, GMP16, Mon96, Ruk07, Lus87]. Randomized [Arn95, TM94, Arn92a, GS12, LCY15, MN08, PS12, Sei78, SKV18]. Randomly [Ger89, Ger87, Kum92, LLWW20, Mie91, Sch87a, WWH18]. randomness [Hei84, Kra79]. range [AOV19, DRW17, FRLLT18, JK08, LLBM^{+}11, Pap16]. ranges [Col91, MR73b]. Rank [Alb84, KB00, Koz96, Kre85, SR91, SR94, BDS22, BDF^{+}18, Cou21, HSW85, Hav84, KS83b, KM20, KL82, Koz78, KZ13a, KZ13b, LTV03, MM18, Mur15, NJKK16, Pie75a, Pon84, QLY21, Ryc09, SA23]. rank-based [SA23]. rank-one [Cou21]. rank-ordered [NJKK16]. rank-sum [Mur15]. ranked [BLT20, MZ18, SM08, SY22]. ranking [LGZL14]. Rankings [Sto96]. ranks [BK77, BNH20, Kra79]. Rao [FD72, FD72, GK95, PV85, SW89b]. rapidly [Mik88]. rare [KK81, LSS12a, LUK13, LUK14, SKV18]. raspredeleniyaks [Ano73f]. rasreshimosti [Ano75a]. rasrivnimi [Ano72a]. Rate [Ger89, JK92, Mi96, Phe96, RT85, BBKP19, BSB13, Bed89, BN16, Bie16, BS18, BCDD11, CZZ17, Flo93, Ger87, Hos88b, HWL17, KYR89a, KYR89b, KAR14, Kin75, Kum92, Lau85a, LMR14, Mal21, Möh87b, Müh80, Pad87, PK18, PN90, QLT17, Sto72a, SSM07, TJZ89a, Wit83b, YN23, ZZLL20]. Rates [SL95, Tij76, BR90, BMRPZ16, CF78, GBS20, Goe75, JMK21, MB21, NK88, YLWL22, pZ15]. Ratio [CT14, Cie98, ZXT11, ZM95, AK85, ADR14, Bat10, CM13, CHL09, DK84, Iga20, JS82, KDM22, Kim16, Küc78, KG81a, KG81b, LLW17, MR73b, MB18, Nad07, Nad12, Nku08, Nus77b, PA14, Rüs81, Sah87, SX20, TL15, Waf88, ZQB16, ZBF^{+}20, ZL03]. ration [Gup77]. rational [Gün77]. rationaler [Gün77]. Ratios [Bed85a, Rüs91a, RR91, PGT05]. Räume [Wit76a]. Räumen [GRV75, GR76, Hol73b, Lip75a, Thä76, Trö75]. räumlichen [Nag83]. Rayleigh [JD20, KR07, MAS06, SP08]. razmernosti [Ano75c]. re [Gov11]. re-estimation [Gov11]. reach [Zha12]. Real [Las91, LBAR13]. realized [HS17b]. reasonable [Hös76]. recapture [SLL15, Zha12]. receiver [JRT21]. recentred [YL17]. reciprocal [AJM12, Ros91]. reciprocals [ST23]. Recognition [Fri74]. Reconstruction [KAR14]. Record [Ber98, AA03, AB05, AB10, AMB11, AA08, AAB12, Bar10, BN14, Gor19, Jas18, KAR14, RA01, RR04, RAD07, RN18, SK20, TA11, ZLLX08]. Records [DT96, Enn98, MJA02, AF14, AV12, BR12, BS21, BM18, BG14, CMLBSM12, DAA20, DB13, JS22b, KW17, KR17, Mou01, RA10, TF21, WLB04].

Recruitment [EPT20]. rectangle [KK15]. rectangle-screened [KK15]. Rectangular [SKS93, Roy87]. Recurrence [Los23, SP98a, SP01]. recurrent [BGL15]. recursion [HS82]. Recursive [Luc05, ST95, Yao00, Bel79a, BNS15, GMP16, Hav76, Pru93, Ras06, Wer85]. reduce [KJ14]. Reduced [FR84, BDF^{+}18, Foc75a, Hei88, JSV07, Ohs83, SYW22]. reduced-rank [BDF^{+}18]. Reducing [QJ08, Zha13, SMLL18]. Reduction [Han83, AX16, Art19, Chr23, FYY^{+}23, GCF04, Mic86, Sch90a, SL23, WN13b, Yoo16, Yoo18]. redundant [CS73, SMLL18]. reduzierten [Foc75a]. reference [GJ73]. Refined [Oui22]. reflected [TB01]. reflexive [Trö75]. reflexiven [Trö75]. regenerative [Win17]. regime [Xie09]. regime-switching [Xie09]. regimes [Sch82c]. region [CM05, HAMAZ96, STY04]. Regions [And76b, CR99, BG19b, BAM18, CTW09, Dem14, DJ17, DAA20, Fis09, Hen79, IR05, Jiá81, JA89, Ker16, Páz91a, Pol85, Roy87, VK20, Zvá79]. regress [GAB85]. Regression [AL00, AST13, AVB98, AS98, Atk75, BCJ14, Ban70a, BL96a, Bun93, Bun99, BDP99, BC99, Cau99, Cav97, Dab81, Det93, Det95a, EA90, GHB02, HK87, HHJ95, Hof97, HDA99, Imh00, JS97, Jon95, KS96, LTG97, Mal89, McC03, MKN99, Neu92, Neu97, OK02, Páz93, Pon02, RV97, TA90, VN98, Ver06, Vie02, AMP21, AK03, AAFG23, AFV15, AOV19, AINJ14, AS04, Ars16, AÖ23, AK20, AE21, BK79b, BOHJ12, Ban80, BN81, BS85, BNP87, BJ14, BM01, BBL20, BB91, BNS15, BTV21, BM18, BAM18, BL06, BN82, BCP99, BRF20, Bro93, BB72, Bun73a, Bun73b, Bun76a, BHSW77, BS80, Bun81a, BM84, BD85, CXH14, CY12, CT07, Che12, CT14, Che21, CD08, Ciu16, Col85b, Cou21, DMP10, DDD19, Dör15, Dro87, Dro93, Dro06, Dry72, DH77]. regression [DH81, Dup77, Dup17, EJ93, FB22, FY03, Fre15, FAP10, Fur15, GH83, GV19, GP82a, GP82b, GGS14a, Gro82, GNP08, GV06, HOCC13, Has77, HEHM13, Hof79a, Hös86, Hos88a, Hos88b, HTHF16, HJM90, HS15, Hub77, HK15, HM07, IR05, IZ83, ILRMZ15, JS22a, JS14, JBC22, JHYQ18, JLM23, JJS06, Jur75, KCNB11, KHN15, KW17, Ker16, KHN16, KCL16, KRC19, Koh12, Kor84, Koz85, KJ14, Kre18, KM17, KZ13b, KP91, LTV03, Läu70, Läu71a, Läu71b, Läu71c, Läu76, Läu89, LMP10, LY11, LMS06, LWS21, Lie82, LQ12, LLL17b, LW12, Lis90, LGW14, LY16, LBSW19, LMV17, LCY15, LGL^{+}17, MM22, Mar19, MCML15, McC05, McC06, Mel78, MP23, MY11, MWZ15, Mie91, MF17, MSW03, NN82, Nät72, Nät75a, Nät85, dDNN19, NS14, ÖCA15, ÖA21, Özk09, ÖA16]. regression [PPD07, Páz90a, Páz90b, Páz90c, Páz91a, PSB23, Pil81, Pil86, PS84, Puk77, Qin15, QK17, Raf88, RH80, Ras85, Rei84, Rév77, Rie89c, Röd77, RVBZ20, RH20, RV08, RG09, Sch91b, Sch79a, ST81b, SZ86, Sch82c, Sch87f, Sen82, She88, sS20b, Sin05, SL11, Spj77, ST91, TSG20, TZ13, Tan14, Ter75, TT10a, TX17, Tou70b, Tou75, TR78, TS87, TH08, TKTM15, Tru18, Van85, VT08, VFS^{+}20, WLF13, WZXH15, Wan15, WLW16, WYLW22, Web89, Web06, WLB04, Wis77, WY13, XW10, XCJH19, YLY17, YLWL22, ZW20, Zha18, ZZLL20, ZL16, ZTJ17, ZMY20, ZL13, ZZZ17, ZWFZ22, Zvá79, Zwa80, Zwa85].

Regression-based [Dab81]. Regressions [Atk75, Hor95, BJfYZ13, Dry88, RG84, Sch82b, YA12]. Regressions-Analyse [Atk75]. Regressionsanteil [Läu71b, Läu71c, Läu71a]. Regressionsmodell [Tou70b, NN82, Nät72]. Regressionsmodellen [Läu70]. Regressionsproblem [GH83]. Regressograms [BC99]. regressor [BG74a]. Regressoren [Tou70b]. regressors [BS85, Dup77, LST18, Tou70b]. Regular [Ber98, HE12, Hös74, Läu70, Pho87, Tan11]. regulären [Läu70]. Regularitätsbedingungen [EG74]. regularity [EG74]. regularization [TKTM15]. regularized [BJ14]. regularly [PN10]. Reihenfolgen [Ull76a]. Reihenfolgeproblem [Ull76b]. Reihenfolgeprobleme [Ban70b]. Reihenfolgeproblemen [Ull76a]. rein [Pie70, Pie73, Pie75a, Har75b]. rein-ganzzahligen [Pie70, Pie75a]. Rejection [BN00b, Väd77]. Rejoinder [Ano92e, CF95b]. rekursiver [Bel79a]. Related [Mol96, VW96, Ger73, GVV13, HBCL14, KS83a, KS83b, Kim10, TL15, ZZWW23]. Relation [BCK05, ETD99, Gne98, SP98a, Ger73, Nus77b, Nus79]. relations [Bit72, HS82, Hös78, KG72, KSS76, MP82, Nät79, SP01]. Relationship [AN98, DN97, VS81]. Relationships [Gup95, Nus76a, TA90, EL79, Rod93, Sen82]. Relative [Cao01, HX20, KB00, AS04, BG19b, BP93, HSW85, LLL17a, YH15]. relevance [Sie74]. Reliability [Bel91, Çet21, CPS98, Gup05, KHN20, LC17, RA02, SAEAEA13, AB05, Bei74, BD17, Cha15a, CF21, CF22, DSG91, Fer07a, Gup84, KJKR76, MZ18, MD22, PT83, Pré73, RN18, Wäc86]. reliabilty [BA08]. remainder [WN12b]. remark [BH86, Hoy79, Kha15, Ohs81, Sto76a, Tro93]. remarkable [Pie73]. Remarks [GV81, HS83, BN80, GW76, Haj84, HR80, Pie75a, Wes93, Wit89b, Hol73a, Neh75, Ron76]. Remedy [KE96]. removals [TDY08]. Renewal [KK81, MS98, BMK22, CGM19, Gup84, Los23, Sch79c, Sch81]. Rényi [MPPV04, SK21, JS22b, MPV00, SL12]. repair [CS73, Mad76]. Repairable [Cse95, Cha15a, LC17]. Repeated [WN12b, Afs85, HH16, PEB15, Sep01]. repeated-measurement [Sep01]. replaceable [Bei75]. Replacement [Bei81, GR06, Ryc04]. Replicated [RF03, Agr82, Jac91, JS22a, PR89]. Reply [Sha96b]. Report [BK71, BM71, ES74, Els75, HW70b, HW70a, HW72, Kle72, Man72, Mön72, PKN72]. Reports [ES71, Gir74, HW74, Lom74]. Representation [KT96, KFL97, TP97, AT13, BCN15, CD09, Dik01, tFYK01, Flo93, HC20, JC07, Jur80, LBAR13, Mec84, PYHT01, Sch82a, Tro93, WYW21]. representations [Ped86]. reproducing [Van85, Wit76a, ZZLL20]. reproductive [CTW09, TWZ06, ZT15]. requirement [FE85]. resampling [Blo07, Sha96a]. research [Ano73a, Pie71, Pie72a, Tit90]. reshennije [Bor70]. residence [Zäh74]. Residual [GK98, AJV21, Ano87j, BMK22, BR12, CGS23, Dry72, Dry88, EPT20, FPLS12, HKT22, IRB14, KDM22, MA20, MH22, NRS07, NFP18, PS84, RA10, SU93, SK21, SL12, ZFA20]. Residuals [GC95, GNP08, Lai00, Cox77, Ras06]. resistant [Mat85]. resolution [MK83]. resolvability [Ano75a]. resource [Hof72, Jen73, Ric76c]. resource-allocation [Jen73]. Respect [Bis95b, GM99, MG96, CFR89, Hof77a, Lis90, Nag78, VO75]. Response [AL00, Arn95, GN00, TM94, Yoo18, Arn92a, BBM12, Bun82c, GMP16, KM82, PS12, RG09, Sha18, SKV18, Sto73, ZWFZ22]. response-adaptive [BBM12, GMP16]. responses [BAM18, KP15, LLV16, MCCL18, PA14, RLAV16, SS21, TZ13]. Ressourcenaufspaltung [Ric76c]. restoration [SW85]. restrained [TR78]. Restricted [GH89, WHG87, AS79, BK79a, BC14, Dan03, Hof77a, JC22, Ker16, LY11, Mat83, NS14, Pin75, Ros89, Tou75, TW78, YW12]. restriction [JT22, Pré74a]. restrictions [BMK16, BM84, FZLB22, Kra86, Läu75b, Lau85b, Özk09, Ron76, ST91, Waw81, Wen76a, ZY07]. Restriktionen [Ron76]. Restriktionsfunktionen [Pré74a]. result [DWHH19, DR03, Neh75, PS77, Zan14, Zan15]. resultant [Kut12]. Results [CPS98, Cse95, GM92, Pho87, Ver92, AF14, AL82, BMS20, BB91, Bla14, BRZ93, BS80, CC13, DM22, FB16, FLL18, Gat09, Hor88, HJM90, JMK21, Jon76b, Küc82a, MP10, MGM20, Mal89, Mik88, MOS86, PYHT01, PMO14, RA10, Ras88, SS81, Sch79a, TA11, WHVW23, YDZ13, ZB11, KJKR76]. Rethinking [Van88]. Retraction [NN11, Ano12b]. Retrospective [WD22]. reversed [KAR14]. reversible [LO86, PM21]. révész [Kor84]. Review [Ano84a, Ano85a, Ano85b, Ano85h, Ano85k, Ano85l, Ano85m, Ano85n, Ano86a, Bel88, Bra88, BHW75, Bun87a, Bun87b, Dro86, Dro88a, Dro88b, Dro90a, Dro91a, DWP^{+}92, Fis87, Fis88, Fre88, Gre88, Hec86, Hen86a, Hen86b, Hen87a, Hen87b, Hen87c, HHN^{+}87, Hof84, HF88, Hof88, Hol70a, Jah87, LSP^{+}73, Lis88, MS82, Nag97, Pol87a, RND^{+}92, SB82, Sto88a, TDB^{+}88, War85b, War86a, War86b, Win87, And87, Ano85c, Ano85d, Ano85e, Ano85f, Ano85g, Ano85i, Ano85j, Ano86b, Ano86c, Ano86d, Ano86e, Ano86f, Ano86g, Ano86h, Ano86i, Ano86j, Ano86k, Ano87a, Ano87b, Ano87c, Ano87d, Ano87e, Ano88a, Ano88b, Ano88c, Ano89a, Ano89b, Ano89c, Ano89d, Ano89e, Ano89f, Ano89g, Ano90a, Ano90b, Ano91a, Ano91b, BRB^{+}77, BSF78, Bel87, BSW^{+}81, Bun82a, BB82, Bun82b]. review [Bun88, CPGB98, DHS82, Dro88c, Dro88d, Dro88e, Dro89, DRR^{+}94, DRR^{+}95, Dro98, Ege87, Eng82, Eng88, FRB77, Fis89, FSS^{+}94, Hen82, JS89, KA88, Lie88, Lin87, Lom82, Mai91, MGLG78, Möh87a, MRR^{+}98, NRD06, Neu99, Nol87, Pil87, PNK78, Pin89, PB82, Pol87b, PSS^{+}94, Ric89, RRL^{+}93, SKJ92, Sch87c, Sch87d, Sch89d, Sch87e, SS99, Sch99, Sch82d, Spj77, Sto78, War85a, Wil87, WHH^{+}80, Hof79b]. Reviews [BDP^{+}74, BWR^{+}71, BGB^{+}72, Bun74, Bur70, DGB^{+}73, DGG^{+}75, DGH^{+}73, EW75a, EGV^{+}73, EvWE^{+}75, FGR^{+}76, GFFS74, HWK^{+}72, HL70, KS71, KPAF75, MSW^{+}72, MSK^{+}76, Mol97, MSS^{+}71, Nol88, RWH^{+}72, SFB^{+}82, SHP^{+}71, SND^{+}72, THN^{+}76, TSL^{+}76, VBP^{+}76, War86c, WHS^{+}76, WRME72, AEH^{+}86, AA98, Ano83a, Ano83b, Ano83c, Ano83d, Ano84b, Ano84c, Ano84d, Ano84e, Ano84f, Ano84g, Ano84h, Ano84i, Ano84j, Ano84k, Ano84l, Ano84m, Ano84n, Ano84o, Ano84p, Ano84q, Ano85o, Ano85p, Ano85q, Ano85r, Ano85s, Ano85t, Ano85u, Ano86l, Ano86m, Ano86n, Ano86o, Ano86p, Ano89h, Ano91c, Ano93a, Ano93b, Ano93c, BvWT^{+}77, Bel83a, Bel84, BS87, Bel90, Bra91, Bun84, BKS^{+}91, Die90, DH90, Doh89, DM91, Dro91b, EHR^{+}80, Eng83, Fis90, Fle91, Gär83, Ger91, Gui91, HPP^{+}88, Hec83, Hei89]. reviews [HGH^{+}79, HSM^{+}79, Hoe83, JRR^{+}96, JSAG78, Kir83, KRF^{+}77, Kon89, Lie90b, Lis83, Moh89, NSH^{+}80, Nat89, Nät91, Ohs91, PHG^{+}79, Pla83, RMGR88, Rie90, Sch83a, SH83, Sch83b, Sch89e, Sch89f, Sch98, SW89a, SSSS96, SCC^{+}00, SSS^{+}00, SSSS99, Sto83a, Sto89, Sto98, SEG^{+}85, Töp83, TBJ^{+}83, VHW^{+}81, WW83, WW91, WPH^{+}79, WBG^{+}81, Win89, Win90, WHD^{+}90, Wit83a, Wit89a, Wit90, Wit91, WWW^{+}90, ZS90]. revision [Bic18]. revisited [BT87, JR13, KW17, Pog15, VT08]. Rezension [BHW75, Hol70a, LSP^{+}73]. Rezensionen [DGB^{+}73, MSW^{+}72, BDP^{+}74, BWR^{+}71, BGB^{+}72, Bun74, Bur70, DGG^{+}75, DGH^{+}73, EW75a, EGV^{+}73, EvWE^{+}75, FGR^{+}76, GFFS74, HWK^{+}72, HL70, KS71, KPAF75, MSK^{+}76, MSS^{+}71, RWH^{+}72, SHP^{+}71, SND^{+}72, THN^{+}76, TSL^{+}76, VBP^{+}76, WHS^{+}76, WRME72]. Richtungen [Hol75, Hol76, Pré74a]. Richtungssuchproblem [Foc71]. Ridge [JRCB22, Tau84, AK20, Bun75d, LY11, LH13, Rod88, RH20, Wu16, ZY07, Tau84]. Ridge-typ [Tau84]. ridge-type [LY11, ZY07, Tau84]. Riemannian [Páz93]. Right [Bel91, CB02b, Ger89, BCD11, BRF20, Cup21, KLF22, Pon04, SP01, sS14, sSLMJ17, SAX23, VDB13a, WLT^{+}22, ZB14, dUÁV17]. right-censored [BCD11, BRF20, Cup21, KLF22, sS14, VDB13a, WLT^{+}22, dUÁV17]. Right-Censoring [Bel91, Pon04]. right-truncated [sSLMJ17]. Risk [Kal97, KS96, Miz96, Nor90, CBC17, Cor17, DP17, FW13, GN22, GW04, LLL17b, Nie06, Nku12, PG19, SL17, Tar88]. Risks [Neu92, BPK14, CDD17, HW14, KKG17, KM21, Mal21, Ohy13, SKD21, WG21, WLT^{+}22]. Rivew [Bun90]. road [NL05]. roadmaps [Näg75]. Robbins [AO07, LPS90]. Robbins-type [AO07]. Robert [Bed89]. Robust [AÖ23, ADF21, BC23, Bed04, BT97, BAM18, CXW20, Det93, Det95a, Fur97, GB17, GV19, GM97, Hof77b, Hub77, KM95, KZ15, Led83, Pil81, RVM22, SIM21, Sha21, SMW22, She88, Tan14, Zha23, AG09, Ars16, Art19, Bed85b, Bun75c, CT07, Col03, CS89, Haf82a, Haf82b, HTHF16, ID03, KP09, LLBM^{+}11, LQZF17, LGL^{+}17, PR22, Rie87, RVB18, RVBZ20, RH13, TC13, TDM13, YH15, YLY17, Zha16, Zie85]. Robustesse [VO75]. Robustness [Bat10, LC15, Liu92, Neu02, Pla02, Rie89b, AvC89, Bha07, GVV17, Ham78, KK19a, Kha89, Nät75b, Rie89a, Ryc93a, VO75]. role [Mar19]. roles [Bun82c]. root [Hab89, LLW17, SN81]. root- [Hab89]. roots [RZ11, SS04, SO09]. rose [Ohs81]. rotatable [HM89]. Roughness [GCB94]. round [Ull76b]. round-trip [Ull76b]. rounded [PP83]. rounded-off [PP83]. roundoff [WN12a]. rowwise [WDZH14]. Roy [BK77, Sch76d]. Roy- [Sch76d]. rule [BBL20, KS79]. Rules [AN98, BV11, BG74a, GN22, Gno77, Rod88, RVM22]. Rumania [PKN72]. Rumanian [PKN72]. run [Dod83]. Rundreiseproblem [Ull76b]. Russian [Ano72b, Ano72a, Ano73a, Ano73d, Ano73c, Ano73e, Ano73b, Ano73f, Ano74, Ano75a, Ano75b, Ano75c, Ano76, Ano79e, Ano81e, Ano87j, Bor70]. Rv [Aly87].

s [KSS76, Sto74, Sto76b]. s.c [Bod75]. Sackrowitz [Bha88]. Sacks [Wit76b, Wit78a, Wit76b, Wit78a]. saddle [Ano75c]. safe [BKHB22, HKK23]. safety [NL05]. Saison [Ron76]. Saison-Variablen [Ron76]. Saisonbereinigungsverfahren [Sti78]. salagaemiks [Ano79e]. same [Oko17, Try88]. Sample [AEW94, BZ10, FZF97, JY20, KB00, Kre85, MH97, SR91, SR94, Sha96a, TJZ89b, ZM95, Zie95, AALA20, AR13, AAB12, AS83, BHN15, BLT20, BFR84, BS07b, BE12, BHS76, CM15b, Dal77, DRW17, Dro93, Ege80, FLL16, FW20, Fer07b, GJV16, GJHN18, GPS14, Gov11, Hel81, HB09, IDB12, JS82, KCNB11, Kat89, Kha04, Kie17, KRC19, Koo86, Kow77, Kra76, Kum13, KC17, LUK13, LLBM^{+}11, LZW23, MMRA07, Mau89, Möh87b, MM18, MKM21, NAHJ03, Nov10, OB19, PJ05, PGT08, PS77, RG84, Ras85, RM17, Rie87, Rie89c, SM08, sS14, Sta84, SB17, TZ91, VDB13b, WYW21, Yu04, ZLLX08, Zuo01, Páz90c]. sample-based [Möh87b]. sampled [And79b, BCP99]. sampled-data [BCP99]. sampler [Lia12]. samplers [Dod83]. Samples [BGP95, CB02b, GKP96, Agr82, BJ85, BZ14, Blo03b, Blo07, BW14, CZZ17, CM05, EBTB21, FW20, GR06, HTX23, Hen89, Hen92, JB83, KFPG21, LBSW19, MR73a, MR73b, MR75, NJKK16, Oko16, PGT10, Ryc93b, Ryc04, Ryc15, ST23, Sch89a, Sch91a, TL15, WS22, WW22b]. samplig [Arn91]. Sampling [BB13, BN00b, Cau99, CBC17, DY02, SS94, AK85, AC09, AG23, Arn92a, BLT20, BBL20, BAO12, CH08, Cha15a, CF21, CF22, CBF23, EPT20, Eva75, FE85, GS12, GFCO04, Ger80, Glo01, KSS11, Krz05, KG81a, KG81b, KKA84, Kum85, Kum87, LC17, MZ18, Mas19, RD21, Rui87, Sch87a, SY22, Sin09, ST86, Von87, VCO09, WC21, WD22, WHT03, ZL03]. SAR [CRCN14]. Saralees [SN14]. Satterthwaite [Khu95]. Saturated [OM91, Bag90, MOS86]. Satz [GRV75]. Saunders [BZ14, FZB17, Gen13, HOCC14, Kun15, MFBG14a, RVB18, RVBZ20]. SCAD [ZX14]. SCAD- [ZX14]. Scale [Alb84, BT97, DM95, GD96, JS97, Lai00, MK95, VV96, AG09, ADF21, BG19a, BB22a, BKHB22, BM01, CLS14, CLZ15, Col03, GG91, GF09, GM97, HSBB23, HW07, JW04a, JT22, JS82, KCL16, Koh12, KR78, Kra76, LLBG10, MBF^{+}20, PKP21, ST23, Sha83, She88, Sta88, TPL19, Tor17, VO78, VI90, WN13c, WN14a, XCJH19]. scale-free [GF09]. Scale-Rank [Alb84]. scale-ratio [JS82]. scaled [FLL18, Hei18]. scales [ST23]. scalings [CMLBSM14]. Schätzfunktionen [NN82]. Schätzmethode [Dew73, FD72]. Schätzung [Foc84, Läu70, Läu71a, Ron76, Tau84, Lyt77]. Schätzungen [Bel74, Bel78, Bel79a, Fei72, FR86, Hof79b, Tau84, EW75b, Gil78, Koc74, MW80, Sch77b]. Schätzverfahren [Fra77]. scheduling [Hof72, Men73]. Scheme [Enn98, BS13, BSS18, Çet21, Eva75, KKA84, MK19, Pru93, SSG76]. Schemes [Neu02, BS17, FZ89, LCY15, PBK11, PKM78]. Schichtanzahlen [Deu72]. Schichtungen [DB74]. Schiefe [Sch87b]. Schnitte [Nag83, Pie74a, Pie74b]. Schnitten [Pie70, Ste76]. Schnittgruppe [Pie73]. Schnittrang [Pie75a]. Schnittverfahren [Har75a, Har75b]. Schnittverschärfungen [Pie76]. School [Atk75]. Schranke [Zin81]. Schur [FK90]. schwach [Web79, Wol75]. Scientific [ES74]. Score [AVB98, NB16, Tro00, Zha97, ANN11, Faz14, KMSS10, Lem13, Lin09, NSK12, SlP17, WL01]. Score-based [NB16]. scores [Rod89]. scoring [ABD93, RVM22]. Scott [Kon87]. screened [KK15, KCL16]. screening [GSA14]. SDE [FL10]. search [BNS73, Foc71, JN75, PVW74, TE15]. seasonal [Eic88, NDDDG16, Ron76, SO09, Sti78]. secant [Mag88, RA06]. Second [Blo07, CG95, Cig13, Dre98, HKMM98, LAO21, Neu92, RT13, Tak09, VCO09, AM15, BJ85, CZZ17, CCM18, DL16, GCF04, Gup05, HS79, Hei91, NAB13, Ohs83, OM10, RR04, SS86, TPL19, UI13, YN23, ZQB16]. Second-Order [CG95, HKMM98, Blo07, Cig13, LAO21, RT13, Tak09, CZZ17, CCM18, DL16, GCF04, HS79, TPL19, YN23]. secretary [Gov92]. section [Zäh84]. sectional [GW04, KZ20, Shi15]. sectionally [SO09]. sections [Nag83]. seemingly [BK79b, Ter75]. segmented [Sch82c]. seine [Ehl75, Ehl76, Gru71]. Selberg [PG09]. selected [AMSV12, JMK21, KJKR76, LV18]. Selecting [MK95, VV96]. Selection [BV11, BDP99, GAB85, VV00, AC23, AFV15, BPH88, BBR05, BDS22, BG78, BD85, Ciu16, Ciu20, DK21, Dro06, Fur15, FR20, GV19, GLLO^{+}22, Gov92, KRC19, LMP10, LGZL14, LNO14, LH22, ML21, MM21, PS10, RD21, Ras91, RV08, SK14a, SS15a, SMLL18, XY16, YLY17, ZX14, ZWL14, ZWYG14, ZH16, ZL16, Dro91a]. selective [Rab15, RD21]. selectors [MP96]. Self [BJ05, CLWK20, GPS14, Jia07, JH09, Pon06, Ros91, SD83, pZ15]. self-adjoint [Ros91]. self-centring [JH09]. self-consistency [Pon06]. self-consistent [CLWK20]. self-exciting [GPS14]. Self-informative [BJ05]. self-normalization [Jia07]. self-normalized [pZ15]. self-normalizing [JH09]. self-reciprocal [Ros91]. selfadjoint [DGP91]. semes [Pru93]. Semi [Cse95, Nät75b, WT20, BAM18, DGL16, LQ12, Pon04, Pon06, Qin15]. semi-functional [Qin15]. Semi-Markov [Cse95, DGL16, Pon04, Pon06]. Semi-orderings [Nät75b]. Semi-parametric [WT20, BAM18, LQ12]. semimartingale [MP23]. Semiparametric [ANN11, BN97, BP18, BGL15, Bun99, CH89, FGK19, FAP10, HH13, XB17, ZMY20, ZL13, BRF20, Bun05, CTW09, Dik01, KRC19, LMS06, dDNN19, QK17, TA21, Tru18, TT22, WWLH19, XWY^{+}20, ZL14]. sense [AR13, AM90, Wit76b, Wit78a]. sensitive [LSS12a, LUK13, LUK14, SKV18]. sensitivity [BHM04, GHB02]. separability [Mat81a]. Separable [SLS99, Azp73, Kle73, Röd70, Sie87]. separablen [Röd70]. separating [GKFMM10]. Separation [GJML00]. Sequence [MKN99, Ano73b, BM12, Ban70b, Bat10, BG74a, Car88, DGP91, HTX23, PPD07, SK05, Ull76a, Ull76b]. Sequences [EMG98, Fer99, LPS90, AB10, Amm20, BW81, Gün77, Jia07, Ker84b, LDF17, LZW23, MP11, Müh80, Müh84, PH16b, Sar82, Ull76a, Wit78b, Wit81, Wit86, Wyn76, ZLLX08]. Sequential [BM12, BGLS15, CG95, FZF97, Fra82, Gol91, GN00, GC04, Jon76a, Mag88, Mag95, Miz96, Mus77, Wer85, WFK82, CS06, CM15b, CT03, Döh81, DK84, DS91, Fra77, Fra79, Fre15, GH83, HAMAZ96, HL13, IV90a, JR12, KB85, KJ14, Kra79, Küc78, KS79, Mau89, Pfl80, Pfu78, Rüs81, Sta90, UI13, WK85, Exn76, FR86, Sch70a]. Sequentialanalyse [Döh81]. Sequentially [XWQN23]. Sequentielle [Küc78, Pfl80, Exn76, FR86, Sch70a]. sequentiellen [GH83, Mau89, Pfu78]. sequentieller [Fra77, KS79]. ser.statist [Sch84]. serial [JH09, Sin84]. Seriation [Bon82]. Series [CPS98, GJML00, JK92, Lie98, Neu02, ANN11, Aka82, AT05, AWJF22, And76a, And78a, And89, ANR07, And77, BS17, Col85a, CF89, Del09, DK21, DdMT20, Eic88, FRLLT18, FC89b, Fok11, Fre15, GPS14, GS21, GML18, GML23, Gun88b, Gun88a, GS82, HE12, Hen92, HHM15, Kam22, KY16, KHN20, Log87, Mag72, MB18, MCNO19, MBF^{+}20, MF17, Nie17, PKM78, SO09, SD83, ŚS05, Sti78, TC22, Tit83, TH08, Web79, Web89, Win81, WN12b, Wit89b, YFZ11, Zar78, dSL22]. served [NK88]. server [Cha74, FS81, Kal76, Sch79c, Sch81, Sto74]. Service [Tij76, Fre74, Mad76, Mag72, Mec72, NK88, Sie74]. Set [Mol96, Pla94, Str95, BLT20, BH87, Bet87, MZ18, NJKK16, SM08, SY22]. Set-Valued [Mol96]. Sets [Cav97, BS06, CM15a, Dev89, GKFMM10, Hof77a, HS17a, KM17, Lip75a, MP96, Sch22, YL17, Zha18]. settings [Wie88]. settlement [Sto88b]. Several [Hol91, KM78, BS85, CWW09, Dav94, Gai23, Gov11, GPK23, KG72, Köc76, RG84, TG87]. severe [LGW14]. SgenoLasso [RD21]. SGT [NP12]. Shannon [RA01]. Shape [Kow82, DW93, DGCL14, Fri74, KT89, NS14, Röd76, TPL15, VO78, ZMY20]. shape-restricted [NS14]. shaped [BN16]. shapes [Sta17]. Shapley [Gru71]. Shapleyysche [Gru71]. share [Läu71a, Läu71b, Läu71c]. Sharp [BR12, BS18, Dan03, RR04]. sharpening [LCK21]. sheet [BPV11, KK11]. shift [Kat89, Koz78]. shifted [And81, BPV11]. shifts [ADF21, Bow12, CMLBSM14]. shock [KZ20, WL05]. shocks [HKK23]. short [AT05, AT08]. short-tailed [AT05, AT08]. shortest [Bak73, Kla72]. shot [BC23, Sch85]. showing [Luc05]. shrinkage [ARN12, BR19, CFR89, GB78, LGW14, Nku12, Nku23, STY04, YAA22]. shrinkage-type [LGW14]. shrinking [Flo93, KS10]. Shrunken [KSS11, Bun75d]. Sidák [SB17]. Sidák-type [SB17]. Sided [Vän98b, Zie95, FE85, GVGP13, Hel81, Mur15, OM11, Pad81, PMO14, SH09]. sides [Hor84, Sch89a]. sign [CLZ15]. Signal [Kun97, GKM18]. Signals [KM96, NKI04]. Signed [Koz96, BDS22, Koz78]. signed-rank [BDS22]. significance [Bic23]. silent [Sty75]. similar [Bic18]. similarities [FGK19]. Simple [Bun76a, Cip84, Haf82b, Nad10, TSC20, BMP15, BMK16, BAO12, FZLB22, KJ14, LTV03, LMV17, MY11, OB19, PDG13]. Simplex [SR02a, SR02b, Lom70]. Simplexmethode [Lom70]. Simulation [Dro90a, Mel94, HP73, PPD07, ST83, Pie71]. Simulations [Höp98, HJM90]. Simulationsmethoden [Pie71]. Simulationsverfahren [HP73]. Simultaneous [BB22a, Hen80, JS82, KS86, Kle78, SS15a, SDC19, Try86, VDB13a, BJ14, HAMAZ96, Ker16, LZW18, Wie86]. Single [GS11, Men73, SP98b, ASFKV08, BHN15, BT87, BM18, BGL15, Cha74, FS81, GNYX15, HMP01, HLM21, Kal76, LWS21, Sch79c, Sch81, SK05, Sto72a, Wei70a, YLY18, ZWYG14, ZZL^{+}14]. single- [BHN15]. single-functional [ASFKV08]. single-humpedness [BT87]. Single-index [GS11, BGL15, HLM21, LWS21, YLY18, ZZL^{+}14]. single-line [Sto72a]. Single-stage [Men73]. Singular [Páz80a, Gaf85, Hös74, Läu70, Sch75b, Wol75, Yin20, ZY07]. singulären [Läu70, Wol75]. singularity [Dac11]. Sinne [Wit76b, Wit78a]. Sinusoidal [Kun97, KN06, NK13]. sistem [Ano73g]. sites [BT04]. situations [Jag80]. Size [KZ86, AAB12, AS83, BB72, BS06, EBTB21, Gov11, JY20, KCNB11, KG81a, KG81b, KKA84, LUK14, Mad76, PJ05, PGT08, sS14, Zha12]. sized [ST75, Sin91]. Sizes [AEW94, SR91, SR94, BFR84, RM17]. Skew [BL06, GM97, MFBG14b, AGOP14, AG09, BGTV06, CLS14, CG05, GSB06, GSB11, GVV13, Gup03, GCT07, HG08, HSG13, Kim05, KA13, Kun17, LLBG10, NNM14, SDJ16, WZW21, ZT15, ZHG12, SGMFB18]. skew-distributions [BGTV06]. Skew-normal [MFBG14b, CLS14, GSB06, GSB11, HG08, HSG13, Kim05, Kun17, LLBG10, ZT15, SGMFB18]. skew-normal-based [WZW21]. skew-symmetric [GVV13, NNM14]. Skewed [NSG07, AT13, AGS09, AGS15, CHA15b, dQLS16]. skewnes [WZW21]. skewness [AM90, DLO22, SD17, Sch87b]. slash [Atk75, RGB13]. Sliced [GHB02, MCNO19, BNS15, FY03]. slope [DMP10, LMV17, RH80, TSC20]. SLSE [TW78]. sluchaya [Ano74]. slycainovo [Ano79e]. slychainikh [Ano79e, Ano87j]. slychainovo [Ano81e]. Small [DM95, BM09, Deu72, HEHM13, Kaw15, KP94, NS22, NW12, PGT10, RS09, Páz90c]. small-area [BM09]. Small-sample [Páz90c]. smallest [FB16]. Smeshannaya [Ano72b]. Smirnov [Haf75, Pré73]. Smirnov-statistics [Haf75]. Smooth [MZ18, vdVvdL03, Che21, IJL90, JW04a, JW04b, Kob87, NFF^{+}23, Pra86, YL17]. Smoothed [MM18, Neu92, BLO03a]. Smoothing [TN08, Van98a, BCDD11, Bun85, Gün80]. Smoothness [Mam95, GN16]. snacheniya [Ano72a]. Snedecor [ALS11]. so-called [Pré74a]. Sobol [GJK^{+}16]. Software [RA02]. sogenannten [Pré74a]. Soil [GCB94]. solid [BSP76]. Solution [Kov74, Bor73, Cha74, Deu72, FS81, Hel93, MESS70, Nož70, Sch76b, Sch90b, Sch73, Sch70c, Sch75d, Sch75e, Tam76, WZW21, vW72a, Lud73]. solutions [FN77, HK78, Men73, SSG76, Wag83]. solvability [Thä76]. solvable [Hol70b]. solving [Bär72, Ehl75, Ehl76, GT70, HT71, HP73, HW71, Jen73, Mel76, Ric76a, Sch70b, ST75, Sch75f, vW72b]. Some [AS04, ACS08, AL82, Ber79, BRZ93, BN80, BF85, CB02a, CPS98, Dom10, ESSB93, Fak96, Fal95, FPPRG16, Fok11, GW76, Gla90, GJ73, Gup84, Gup85, GV95, Gup95, HBN86, Haj84, Han96, HKMM98, Hor88, HJM90, Jac91, Jon76b, JBD14, KK80, Koo86, Kos80, KB00, KL82, Kum85, Kum89, LM79, Lou99, MN08, MFTFEO22, Mie91, Moj07, NK05b, Nad08c, NRS07, NLP95, Páz80b, Pol86, PN10, RA10, SS81, SP98b, Sin91, Sto74, SH86, TA11, Tia91, Tit90, WY04, Wes93, Wit89b, WHVW23, ZB11, AB05, Ano75c, Bon82, Cap19, Fur15, Ger78b, GK17, Gru71, Gün83, Gup79b, GK93, Hei88, JLO12, JE90, Kal76, KP09, KSS76, Kra76, KA13, Led84, LLBM^{+}11, Lip76, MP10, Mal89, Mat87b, Men73, MO19, Nad06, Nad80, NRD06]. some [Neu73, Nie17, Nog20, Oui22, Pho87, Ros72, Sai86, Sen82, Xek86, XG07, Lip75b, Pie75a, SN14]. Sommerschule [Atk75]. source [Goe75]. Sources [GM91]. Space [GH89, GP07, AK75, Goe75, Kir76, Kle73, Küc78, KS79, Páz80a, Röd70, RW84, Sie87, Waw81]. Space-time [GP07]. spaced [AAFG23]. Spaces [Ruy98, Ano75c, Ant84a, Dry75, GRV75, GR76, Hol73b, Lip75a, MP82, Mos85, Ric86, Sto79, Thä76, Trö75, Van85, Wit76a, ZZLL20]. spacing [ZQB16]. Spacings [Eks99, Bar94, BHN15, BSS18, BS21, BS23, Bie08, TL15, ZLLX08]. Sparse [LL18, MO03, ACH20, GLLO^{+}22, Guo20, PG19, SS15a, XCJH19, ZZW23]. Spatial [HKMM98, BPvZ04, BCDD11, CFC10, Haj84, Hei84, JS14, Nag83, RZ11, SV18, Sto84, Sto93, Tan14, WLW16, ZW10]. spatial-type [SV18]. Spatially [MU99, Kre18]. spatio [CCM18, HC11]. spatio-temporal [CCM18, HC11]. Special [HR96, Bal03, BFN17, Bed85b, FR86, Han83, Lis78, Nku08, SSG76, Ull76a, Wäc86, Gir74]. specific [Sch76b]. Specification [Vän98b, BG78, CKM09]. specified [Bun76b, GK17]. spectral [BS82, CFC10, Gün77, TKTM15]. speech [NKI04]. speed [Her81, Pon86]. Spektraldichte [Gün77]. Spezialtagung [Gir74]. speziellen [FR86, Sch76b, SSG76, Ull76a]. Sphere [Gat94, GP82a]. spheres [AOV19, Sto90]. spherical [HHM14, KK19a]. spherically [Wie92]. sphericity [PGT10]. Spitzer [Hei85]. Spline [BL96b]. Splines [TA90, ZWL14]. splitting [GV06, Ric76c]. spread [ZB14]. Sprent [Mai91]. Spring [Kle72]. SPRT [IS84]. SPRTs [Ege80]. Spurkorrelation [Hös76]. Square [GM99, IKR00, Ano72a, Bed81, BTV21, BH89, Car19, CV78, DR15, GK77, HGLS82a, HGLS82b, HGLS83a, HGLS83d, HGLS83b, HGLS83c, JLO12, Nad10, Oui22, Sch75b, Sch74b, SN81, TIT08, TMY05]. squared [AS04, BM09, Bla88, EJ93, Sta20]. Squares [BFF00, Kun97, Nus76a, TVL96, ASSH13, BK77, Bel79b, Bun75c, BS80, CH89, CFR89, DI11, Dro93, FS02, GKM18, Hof77b, Hof80, Hös78, Hue86, IZ83, JR06, KJ14, Läu89, LT13, Lit14, LH15, Páz84, Pes13, PS84, Pra86, Raf88, Rao83, Rod93, STS85, SL11, Son82, Tau84, TW78, Van91, WN13a, Zwa83, Zwa85]. SR [PKN72]. St. [Ber17]. Stabilitätsanalyse [Mat75]. Stability [DDT88, Gud76, Höp98, YY07, DDT89, Mat75]. Stable [Ber75, JK92, MKN99, Zie88b, Zie00, BH18, BBKP19, BG11, NDDDG16, Oko16, RY86, Ryc87, SZ17, VPR21]. Stage [CR99, Ano73e, Arn92a, CS06, CT03, Els22, Men73, SMK22, Sch70c, Sha18, SKV18, TW78, UI13, ZZW23]. Staistics [Ano88e]. Standard [BGP95, JC22, Páz80a, Ros91, ZZWG09]. standby [CS73]. start [Sie74]. starting [RS74]. State [Mel94, Ber76, Cha74, FS81, HC11, JK76, Kir76, Küc78, KS79, Lis79, Sin84]. Statement [Ano12b, NN11]. statements [Trö75]. stationäre [JK76, Müh80]. stationären [Gri80, Gün77, Nag83, RS74]. stationärer [Web79, Wit78b, Wit81]. Stationarität [Fle71]. Stationarity [GM96, Fle71]. Stationary [CHP03, FZF97, HS17b, Shi15, And79a, And80, ANR07, BM12, BW81, BN80, Car88, CCMW09, DM22, DJ17, GW84, Gri80, Gün77, Haj84, Han83, HT15, Hei88, Hei91, JK76, Kal76, Ker84a, Ker84b, KSS76, KS76b, LO86, LDF17, LLV16, Lis79, MP10, Mat81a, MS80b, Mec81, Mec84, Müh80, Müh84, Nag83, NKI04, PMS81, RS74, Rya14, SMP80, SG91, TSO07, Tho84, Web79, Wit78b, Wit81, Wit86, Wit93]. Statistic [Lai00, AV12, BJ85, Car88, DRW17, GF09, Güv06, KS16, KZ13a, MSW03, Neu73, PR89, SX20, TG87, Wes14, WL01, YYVH16]. Statistical [BK86, Bun75a, CG04, DDLN17, DGCL14, Esc84, GJK^{+}16, Gro82, Hec86, Hof84, HLM21, KMM21, LDF17, LW18, Mai91, Man72, MAS06, Nät97, Páz93, Pla94, RO98, RAD07, WG21, YLY18, AMP21, Ahm74, Ano76, BP17, Bic23, Bor70, Bun76b, CLCP17, DWHH19, DMP07, FGL17a, Ger78a, Gla76, HS83, Jiá81, JA89, Jür84, Koo86, Lau85b, LC17, Lus87, MYK18, MP82, Mos85, Nog20, Pet86, PD80, Ric86, SMW22, SH86, Sto88b, Sto93, TKTM15, WFK82, XB17, Yat17, ZZWW23, Zil78, Ano73f, Dro90a]. statisticheskich [Bor70]. Statisticheskiye [Ano73f]. statisticheskovo [Ano76]. Statistics [Ano86u, Ano89i, Ano90c, Ano91d, AE21, Dro90b, FR95, GSB11, JS97, Lou99, MPV00, MO03, Nag92, NLP95, Nku23, PV98, Páz91b, SGMFB18, SR91, SR94, SP98a, SP98b, SG91, SAX23, VV00, Ver92, War85b, AAFM07, Ahm01, Ahm11, AB05, AB10, AMB11, AR13, ASKS16, AA11, AA13, BT08, BB93, BCKS01, BCK05, BCD11, BOHJ12, BCJ14, BMP15, BT16, BB22a, BM23, Bal03, BB04, BN14, BNH20, BBK12, BBK14, BMK16, BK17, Bie06, Bie08, BMM13, Blo07, Böh83, BW14, BC12, CZZ17, CMLBSM12, CMLBSM14, Cen78, CBP23, CY12, CM13, CZZ19, Cig13, Cra06, CV78, DB10, Dem14, DJ17, DL16, EBQY17, FB16, FK14, FZ89, GKFMM10, GR06, Gor14, Gup79b, GNP08, GH01, GH02, Haf75, HSBB23, Hav76, HS15]. statistics [IM14, JR13, JGAFBCMG14, JB83, KC01, Kat89, KS83a, KS83b, KS92, Koz78, LAO21, Lem13, LBSW19, MR73a, MR73b, MR75, MGM20, Moj07, MPPV04, NK06, Nor90, NFF^{+}23, Oko17, PP83, RA01, Raq04, RA06, RM17, Ric84, Rod89, Ros72, Ryc04, Ryc09, Ryc15, SP01, SW85, Sch87b, Sen11, Sta20, SB17, Sto79, SV10, Tor17, VK20, VDB13a, Wit76a, XG07, YN23, YA12, ZZQ16]. statisties [Ant84b]. Statistik [SW85, Wit76a, Neu73, Ric84]. statistischen [Gla76]. statistischer [Jür84, Ric86]. status [Jas22, Mal21, WH20]. Steady [FS81, Sin84]. Steady-state [FS81, Sin84, Cha74]. Stein [ARN12, Bun77b, GN22, Hof97, Lus87, Tau84, ZZWG09]. Stein-rules [GN22]. Stein-typ [Tau84]. Stein-Type [Hof97, ARN12, Tau84, ZZWG09]. Step [BN00a, BS75, DdMT20, Flo93, KS16, SMK22, WG21]. step-stress [SMK22, WG21]. Step-Stresses [BN00a]. stepdown [Fur15]. stepping [Kér72]. stepping-stone [Kér72]. stepwise [BD85]. Stereological [PMS81, SS86, Sto90, Kos80, Sto84]. stereology [KZ86, Zäh84]. Stetigkeit [Hoy79]. Stetigkeitseigenschaften [Sto73]. Stetigkeitssatz [Sto72a]. Stetigkeitsverhalten [Gus71]. Steuerbereichen [Dam76]. Steuertheorie [vW74]. Steuerung [Wer78, Wol75, vW72a, vW72b]. Steuerungen [GM78, Gus71]. Steuerungsprobleme [Dam76]. Stichproben [Hen89, Sch87a, Sch89a]. Stichprobenwerte [Mau89]. Stieltjes [LD21]. Stochastic [BM71, BHN15, Dor85, Dup77, FC89a, FLL16, FZB17, FW20, Has77, HKK23, Las91, Lia12, MG96, Rüs91a, RR91, Tor17, YN23, ZLLX08, Ahm74, ASKS16, BMS20, Bel74, Bel79b, Bel80, Bel83b, BP17, Ber75, Bun76b, BG79, CXH14, CLCP17, CB04, Dab81, FPPRG16, FL10, GCL16, GB17, GBS20, GB23, Got76, GM78, Gre87, HX20, Her81, JK76, Köc75, Kon87, KS76a, LAO21, Läu71a, Läu71b, Läu71c, LO86, LR17, LY11, LF22, MW80, Mus77, NS22, Nät84a, Pen85, PVW74, PT83, Pol08, Pré72, Rab15, Rei84, Rév77, RH20, Rus91b, Sch89a, Sch82a, Sch90b, Sie86, Sie87, Str82, ST87, TW83, Thr77, Thr79, Tou70b, TK22, Try86, Try89, VI90, WZW21, YW12]. Stochastically [Rüs81]. Stochastik [BM71]. stochastisch [Sch89a]. stochastische [MW80]. Stochastischen [Bel78, Bel74, KS76a, Läu71a, Läu71b, Läu71c, Sch90b, Tou70b]. stochastischer [GM78, JK76, PT83]. stochastisches [Gre87, Köc75, PVW74]. stock [Köc75]. stocks [Köc76]. Stokes [Sch90b]. Stokes-Gleichung [Sch90b]. stone [Kér72]. stopping [Ano73b]. storage [Gir71]. straightforward [Sto73]. Strategieparameter [Gri80]. strategies [Arn91, Bei75, Rei86]. strategy [BB13, Gri80, MB91, MB92, MB92]. stratification [DB74, KG81a, KG81b, Rui85]. Stratified [DY02, Blo03b, Blo07, KSS11, LUK13, Rui87, SKV18]. Strauss [SG91]. streams [Ano75b]. Strenge [Foc75b]. strength [BD17, Çet21, GPK23, KFPG21, MD22, RN18, SAEAEA13]. strengths [Çet21, KFPG21, RN18]. Stress [Han96, MD22, RN18, BD17, Çet21, GPK23, KCNB11, KFPG21, SMK22, SAEAEA13, WG21]. Stress-Passing [Han96]. Stress-strength [MD22, RN18, Çet21, GPK23, KFPG21, SAEAEA13]. Stresses [BN00a]. strict [LF22, Foc75b]. Strong [AL99, Ber17, BZ13, ÇO17, Eks99, Enn98, LT13, dDNN19, PN90, Sch76c, Sch22, dUÁ00, dUÁRC04, BO01, Bos93, Che21, DdMT20, FL03, FN13, GB23, HK09, HP08, HCW16, Läu89, Lit14, QLT17, WHVW23, Zäh14]. Strongly [Mi96]. Structural [AVB98, Nus76b, Bun75a, HS17a, Pol85, Rei86, TB78, Tou70a, TW78]. Structure [FK14, MD03, Neu97, And78b, Cra06, Dew76, DJ12, GS04, Jaw13, Lud73, Pad89, Rud76, ST81b, SX20, WLW16, ZL13]. Structure-preserving [FK14]. Structured [Bel91]. structures [Düc75a, Ric86, YT01]. strukturellen [Tou70a]. Strukturen [Ric86]. Strukturgraphen [Dew76]. stückweise [Tam74a]. Student [Zie95, CY12, Mas19, MCML15, MCCL18, YA12]. Student- [MCML15, MCCL18]. Studentization [Pon84]. studied [DDT88, DDT89]. studies [FGR73, GW04, Koo86, RD21, YS23, EG74]. study [Bak84, BH89, DGP91, FP92, Fin94, Hol76, HSG13, KYR89a, KHN20, Moh86, PPD07, ZQC23]. sub [UR81]. sub-system [UR81]. subclass [Ano19, LaM17]. subdifferentiale [EN75]. subdifferentials [EN75]. subgradients [HS91]. subject [Ros72]. subjective [Got74]. Subjektiver [Got74]. Subsample [Pau12]. Subsampling [BPR00, JLM23]. subsequences [Sch89b]. subseries [Car88]. Subset [VV00, AMSV12]. subspace [Chr21, KS10, LL18, Yoo16]. successful [Els22]. successive [KG81a]. Suche [BNS73]. Suchverfahren [PVW74]. Sufficiency [Hof81, IV89, Pla02, Rüs91a, RR91, Bun77a, Sie14]. Sufficient [Sha83, Tan11, Yoo16, Yu04, AX16, Art19, LY93, Lit14, Mic86]. sufficiently [Tas21]. Suffizienz [Döh81]. suivant [Cap73]. Sum [BN00b, EW99, KM20, Lys76, MO73, MM18, Mur15, NKT21, PS84, WBL06]. Summen [Wol84, Wol85]. Summenwerten [Lys76]. Summer [Atk75]. Sums [BK77, Eks99, Bos93, DWHH19, DST21, LLWW20, Lys76, MV18, Pad87, Pro89, SS79, WWH18, WN13a, Wol84, Wol85, pZ15, ZZWW23]. sup [BJ14, SDC19]. sup-norm [BJ14, SDC19]. super [Arn91, KSS11]. super-population [Arn91, KSS11]. superadditive [BTV21, SW16, WDZH14]. superadditive-dependent [WDZH14]. Superposition [RA02]. superquantiles [DM22]. superquantiles/CVaR [DM22]. supersaturated [COQ21]. Supervised [LNO14, PB21]. Supplementary [Mak06]. Support [Mei06, Led78b, PV85]. supporting [BR19]. sure [AAB18, Col85a, DM22, JC07, Sch89b, SK20, Zan14, Zan15]. Surface [GCB94, KM82, PMS81]. surfaces [Hös86]. surrogates [MN08]. Surveillance [ŚS05, GOS10]. Survey [Ant84b, Neu79, DR15, Koo86, LV18, MMRA07, Páz80b, Sin09, Tit83]. Surveys [Arn95, Sha96a, Arn92a]. Survival [BN97, EZ92, Jör99, LP97, RT85, Eco13, EPT20, GS14, GJ19, Lee08, MP22, NTJ09, WD22]. Survivors [GM92, HOCC14]. sushchestvovanii [Ano72a]. svoictvakh [Ano81e]. sweep [SW89b]. sweep-operational [SW89b]. Switch [Tij76]. Switch-Over [Tij76]. switched [FM01]. Switching [FR98, Azp73, BG16, Xie09]. symm [Ano79e]. Symmetric [ESSB93, Zie00, AT08, BH18, De 22, GMCT18, GVV13, HZ21, JR13, JBD14, KZ13a, KZ13b, NK05a, NNM14, RR04, BT97]. symmetrized [Col91, Rub21]. Symmetry [Han96, BGV06, HHM14, MO19, SX20, Tom90, TMY05]. Symposium [War85b]. System [CPS98, Fle76, Köc76, Tij76, Tou70a, Bei81, CT73, CS73, Fre74, GW76, Gre87, GP07, Jas22, KFPG21, LO86, Nku08, PN10, RN18, Sto76a, Sto90, ST87, TW78, Try89, UR81]. Systematic [Try91b, GS12, GFCO04]. Systeme [MESS70, Mül74, Mat75]. Systemen [Wer78]. Systems [Cse95, BKHB22, Bei75, Bel74, CLCP17, DS78, Got76, HKK23, KJKR76, KSS76, KHN20, Lis79, Mat75, Mec81, MESS70, MF17, Mül74, Pie72a, RW84, Sak71, SS74, Sie74, Sto72b, Sto73, Sto74, SMP80, Try86, UR81, Wer78, YDZ13, Zäh74, ZB11, ZB14].

Table [AAJ97, BH89, Led84, PGT08]. Tables [BFF00, BS88, HGLS82a, HGLS82b, HGLS83a, HGLS83d, HGLS83b, HGLS83c, Hof84, Bun85, Fei72, GS11, KI03, KP15, Led78a, Mat87b, Moh86, TIT08, TMY05]. tagged [Ano75b]. Tagung [Mön72]. Tail [HW14, GCF04, GJLT15, HPN22, RM17, SdSSL15, Zuo01]. tail-heaviness [SdSSL15]. tailed [AT05, AT08, BLNGB05, GJLT15, QXY23, SS15a, WW21]. tails [GGS14a]. tailweight [AM90]. Takacs [Bil97]. tangency [KMM21]. Tanner [Jai74]. target [Cou21]. task [KM13, Lud73, vW72a, vW72b]. tasks [BTS71, Hol75, Hol76, Ric76c, Sch76b, vW74]. Tchebycheff [Mun98]. Tchebycheff-Experiments [Mun98]. Technical [ES74]. Technique [Mel94, Els17]. techniques [Chr23, Els22]. Technischen [ES74]. technology [Pré74b]. Teil [Ehl75, Lip75b, Lip76, Ric76c, Sch75d, Sch75e]. teilweise [Fäh81]. telegraph [DI11]. temporal [CCM18, HC11]. Tends [HDA99]. Tensor [Páz93, Dry84, GN22]. tensor-products [Dry84]. teorema [Ano79e]. teoremye [Ano87j]. teorii [Ano75b, Ano76]. teoriyu [Ano73g]. Term [GM92, Ano87j, BS82, HOCC14, Mön72]. Terms [VV96, Foc75b, NL05, WN12b]. tessellation [RT13]. Tessellations [HKMM98, Mec88, SH86]. Test [AAK98, Bro93, Cie98, HA02, Khu95, Kum13, Zha97, ZM95, Zie95, ANN11, Aui90, ADR14, BMP15, BMK22, BC23, Bat10, BPH88, Bed81, BKN70, Bha07, Bie16, BS18, Bla88, CY12, CM15b, CCM18, CCH87, DK84, ELV22, FYY^{+}23, Faz14, Geb84, Ger79, Ger80, Ger86, GJHN18, GVGP13, GT13, GNP08, HMP01, HE12, Hei84, Hen93, HK15, JK08, JW04a, KDM22, KNS72, KPH^{+}16, Kla03, KY16, Küc78, KS79, Led83, LKK07, LLC16, LF22, Lin09, LQ12, LYB19, LM23, Mak06, MYK18, MPPZ05, Mur15, NCJ23, NM18, Nku08, Özk09, Pfu78, Pol08, Pon93, PR89, QLY21, Ras85, Sch87b, Sch74b, Sei78, SKD21, TZ91, TDM13, VS81, VO75, Vís86, VCO09, WLD18, Wes14, XG07, Yu04, ZXT11, ZX17, Zha23]. testability [AS79]. Testing [Amm20, BN97, BJfYZ13, BM02, BL96b, Bre05, CK06, DN97, EM87, Ger87, GD96, Güv14, Han96, HHM14, Hol91, IV02, Jia07, JH09, JJS06, LCY17, LWL18, MZ96, PG19, PGT10, Pin75, RG84, RFK18, Sch82b, SZ86, Sto96, Str95, TJZ89a, WN13c, ZC10, Ahm74, BG19a, BMP15, BH87, BBK14, BMK16, Bed81, Bed85b, BR19, BR83, Cey20, CBP23, CLZ15, CWW09, DR15, DJ86, DF15, EBTB21, Fah87, Faz08, FR20, GOV17, GV81, GVMS09, HK78, HKT22, JGAFBCMG14, Jur75, KCNB11, KG72, KM13, KDM22, LL85, MBF^{+}20, MPPV04, PAW05, PSB23, Que86, Rod89, Rya14, SMK22, Sch79a, Sei79, Sol19, TG87, TC22, TT10b, VO78]. testing-plans [CBP23]. Testproblemen [Pfu78]. Tests [Ach95, Alb84, AVB98, Cip91, DT96, Ger89, GV79, Gla76, GM94, Han96, HM02, HHM15, JS14, Kin75, KB00, Koz96, Kre85, Kum92, LP97, MMP16, Mat81b, Mat81a, Neu79, Sch75a, Sch70a, SO09, SD17, Tjø96, TJZ89b, AALA20, AWJF22, Ara93, BGLS15, BMMP16, BL78, BE12, CC01, Cap19, CH08, CBF23, CLZ15, CT14, DR92, DLO22, DP17, Gai23, GMP16, Ger78b, GV06, Hab89, Haf82a, Haf82b, HK00, HSW85, Hei91, Hei18, Hel81, Hos88a, HM07, IJL90, IR05, JW04b, JGMO20, JE90, Kat89, Kha89, KN90, KW15, KM20, Kra76, Kra79, LC15, LTV03, LLW17, LM79, Led78a, Led78b, Led84, MMP05, Mat89, Mie79, Mil86, MO19, MM18, MKM21, NL03, OJM15, Pin77, Pon84, Röd71, RALN93, Rüs81, Rüs83, Rüs86]. tests [Sai86, Sch91a, Sch76c, Sch79b, Sch76d, Sen82, SS04, Shi15, Sin05, SB17, TDY08, VO75, Wäc86, ZBF^{+}20, Röd76]. Teststatistiken [Sch87b]. Th [Hec86, BS21, DLO22, DST21, Gor19, Han83, Jas18, KR17, Lie93, RM17]. Theil [GT70, GT70, TR78]. their [AA08, BCJ14, Bel83b, Ber79, BN00b, DR15, Gla76, Gup95, Kim12, KL82, MU99, MSCC16, NS10a, Nät75b, Nät79, OJM15, Ric84, Sto88b, SV10, TA90, VV96, dQLS16]. them [Hös77]. Theorem [EW99, AAB18, Ano87j, BL11, BV86, Bod75, Böh83, DM22, DH77, EN74a, Gla90, Góm18, GRV75, HT15, Kin76, KK14, KS76b, LG05, LZW23, NR81, RR02, Sch82a, Tho78, Tro91, Zan15, Zie85, BR90, Hol73a, JS81, Pad81, Pad84, Pad87, Sch89b, Wol84]. Theorems [Eks99, FZF97, Lou99, BNAA13, BNH20, DST21, FG75, KAM17, KK81, Lie72, Lus87, Pré73, RD80, Sch79c, Wol85]. Theoretic [CPS98, Gat09, SK89]. theoretical [AL82, FK90, Juh89, MN08]. Theorie [EN74b, Wit76a]. Theory [Kun97, LMS92, Man72, Neu79, PP83, PKN72, War85b, Ano75b, Ano76, Ano87j, BCMR90, BGTV06, Bei74, Böh81b, BDF^{+}18, CLDG16, DGCL14, Dry88, EBQY17, ES73, EW73, EN74b, Fah87, Fed80, Fin87, Gaf85, GJ73, Ham78, Hös77, HM88, Kle80, Kum89, Lie72, MS80b, Nor90, Páz80b, Pie72a, PS77, Rao83, Sch76c, Sch79b, ST81b, Sen82, SW89b, Sto71, Sto72a, Sto72b, Sto73, Sto76a, SV10, Tam74a, Tho78, Wit76a, Zan14, vW74]. thermodynamic [FN77]. thermodynamischer [FN77]. Thick [Zäh84]. Thin [Nag83]. things [Sta17]. Thinnings [Sto88b, WS97]. third [Han83, Zwa85]. Three [CS06, CT03, GSA14, GGS14b, Kum87, KD74, LCMA16, Mat87b, PCKMK17, RT13, RNT13, WQO22]. three-dimensional [Mat87b]. three-level [GSA14, WQO22]. three-parameter [LCMA16, RNT13]. Three-stage [CS06, CT03]. Threshold [LLZ19, De 22, GPS14, HB09, LT16]. threshold-asymmetric [HB09]. threshold-type [De 22]. thresholding [AX16]. tie [Sta88]. tied [KYR89b]. tightening [Pie76, Düc75b]. tillages [Sto88b]. Time [AT05, BSP76, GJML00, HHS97, MB18, Neu02, Pon02, ANN11, Aka82, AWJF22, AFSS22, And78a, And89, ANR07, And77, And78b, BP16, BBKP19, BR12, BP17, Bie16, BS18, BS17, BCP99, CHPS13, Col85a, CCH87, CS73, Dam76, Del09, DK21, DdMT20, EPT20, Eic88, FRLLT18, FB22, FZ89, Fok11, Fre15, Fre74, GPS14, GS21, GML18, GML23, Gun88b, Gun88a, GP07, HE12, HHM15, IS84, JS14, JR12, Kam22, KK80, Kaw15, Kir76, Kla72, KY16, KS76b, Log87, Mad76, MCNO19, MBF^{+}20, Nie17, PM21, PKM78, Pru93, RA10, RS74, Ruk07, Rus91b, Saw81, Sch81, Seb76, Sen82, SO09, ŚS05, SKD21, Sti78, TC22, Tit83, TH08, Tru18, Try89, VI90, WHXC16, Web79, Web89, Wit89b, XY16, YFZ11, Zäh74, Zar78, ZC10, Zha23]. time-dependent [Dam76, Fre74, Kla72, Seb76]. time-domain [Sti78]. time-homogeneous [Kir76]. time-series [And78a, GPS14]. time-transformed [JR12]. time-varying [FB22, Tru18, XY16, Zha23]. times [CGM19, DI11, Fra77, LC19, Los23, Mec72]. tipa [Ano72b]. to-asymptotic [ST81b]. tochku [Ano75c]. Toeplitz [MD03]. Tolerance [FMTZ07, Jiá81, JA89, OM11, PMO14]. tool [VN19]. Tools [Dro90a]. topological [Lip75a, PB21]. topologischen [Lip75a]. Total [Arn95, Pes13, Bed89, Bie16, BS18, CCH87, Deu72]. Trace [JS97]. Tracers [Phe96]. track [Hös76]. tracks [LV18]. trait [ADR14, Rab15]. transaction [FS81]. transfer [DS78]. Transform [HM02, Böt20, LD21, LWL18, XCJH19]. Transformation [Ste98, BLNGB05, Bun76a, LGZL14, sSLMJ17]. Transformations [BDP99, ST86, BLSC09, Bar87, Bun82c, Cox77, Röd04]. Transformed [Chr21, GMCT18, LTG97, BMS20, CC01, JR12]. transforming [Bic23]. Transforms [Gup95, CCH87, LD21, ZH16]. Transient [Cha74, Phe96]. Transient/steady [Cha74]. Transient/steady-state [Cha74]. translation [BMM13, PV85, Rüs90]. translations [CMLBSM12]. transport [Wen76a]. transportation [BSP76, Kér72, Sch70c, Wen76b]. Transportproblem [Sch70c, Wen76b, Wen76a]. Trapezoidal [BBL20]. treatment [BD13, HBN86, Hof72, JY20, Mak06, Mej78, Ric76b, Sto79, Gud70]. Treatments [Mon96, CWW09, Dav94, DR92, Gov11, HDC20, Jac91]. tree [BC14, FZLB22]. tree-order [BC14]. trees [Boe72]. Trend [BMS20, BT04, DT96, Eic88, Sto96, JK08, KB85, Nät84b]. Trend-seasonal [Eic88]. Trend-transformed [BMS20]. trends [And78a, Ham78, TN08, Zar78]. trial [BBM12, Hös77]. trials [Gov11]. trigger [Kon87]. Trimmed [BC02, JBC22, MSCC16, BW14, Dan03]. trip [Ull76b]. trivariate [BOHJ12, CB02a]. tropospheric [BT04]. truly [SD83, Sto93]. Truncated [NNM14, Ahm01, ACS05, AGOP14, AA11, AA13, Gla90, GML18, Gup85, JIP10, Kim10, Kim12, KLF22, LdCIP18, Ohy13, sS14, sSLMJ17, sSW19, sS20b, SK21, SAX23, WLF13, WLT^{+}22, Zha13, dUÁV17]. Truncated-exponential [NNM14]. truncation [DJ12, Jaw13, PP94, sS14]. tseni [Ano73b]. tsepi [Ano81e]. tsiklicheskoi [Ano73f]. Tucker [Hol73a, GRV75, Hol73a]. Tucker-Theorem [Hol73a]. Tuckersche [GRV75]. Tukey [GN16, Hus88]. tunktsionalov [Ano81e]. tuple [Oko17]. turning [NL03]. Two [CG95, CR99, DRW17, Den88, GL94, GM94, GK99, JGAFBCMG14, Kre85, LNBR02, MKM21, PMO14, SS94, Tij76, VV96, VDB13b, ZZW23, ASSH13, AB10, AR13, And80, Ano19, Are05, AdUÁS15, BB13, Ban70b, BK17, BL82, BS07a, BB91, BH89, BE12, BCN15, BS75, CF22, CY12, CFS17, COQ21, CM15b, CS10, CM05, CBC12, CS73, DP17, EP17, EJ81, FM89, Fer07a, Fer07b, FMM85, FMTZ07, GS11, Ger79, GJHN18, Gno78, GSTUM02, GVGP13, Gov92, Gre87, GCT07, Hel81, HDC20, JT22, Jon76b, JS82, KHN15, KI03, Kat89, KS16, Kim05, Kin75, Kle75b, Kra76, KG81a, KG81b, KS93, KA13, LaM17, Led83, LCY17, LW18, LD21, LC10, LST18, MM22, MR73b, Man22b, MM21, Mat87b, MM18, OM11, PAW05, PK18, PKP21, Rei82, Rod88]. two [SK92, ST81a, ST23, Sch76b, Sen11, Sha18, SKV18, SH09, SB17, SD17, SKPK21, TZ91, TL15, UI13, WLD18, WQO22, WWC^{+}10, WY13, WA22, YDZ13, YLY18, YT01, Yu04, YYVH16, ZW20, ZLLX08, ZB11, ZB14, ZL03, Röd71]. two- [WQO22]. two-and [Mat87b]. two-dataset [MM22]. two-dimensional [And80, BS07a, BCN15, Röd71]. two-factor [Ano19, GCT07, LaM17]. two-group [Gno78, YT01, YYVH16]. two-level [COQ21, EP17]. two-parameter [BK17, CY12, CS10, CBC12, Fer07a, Fer07b, Ger79, PAW05, Sen11, WWC^{+}10, WY13, WA22]. two-parametric [Gre87]. Two-Phase [SS94, ZL03]. two-piece [Kim05, KA13, SD17]. Two-Sample [Kre85, DRW17, VDB13b, AR13, BE12, CM15b, GJHN18, Hel81, JS82, Kat89, Kra76, MM18, MKM21, SB17, Yu04]. Two-sided [PMO14, GVGP13, Hel81, OM11, SH09]. Two-Stage [CR99, ZZW23, Sha18, SKV18, UI13]. two-step [KS16]. two-variable [KHN15]. two-way [BH89, FMTZ07, GS11, KI03, Kle75b, LC10, OM11, SK92]. Tyler [Rub21]. typ [Ber74, Tau84, Tau84]. Type [Hof97, JK92, JS97, Lai00, Mag95, SS94, AJM12, AO07, Ano72b, ARN12, BCKS01, BMMP16, Ber74, BKS23, Bos93, CBP23, CTL14, CGM19, Dac11, DJ09, De 22, Dom10, DMR07, FGL17b, Geb84, Ger79, Ger80, HK00, HW14, HW07, IJL90, KS16, Kut84, KP91, LY11, LBR23, LGW14, Lus87, Mag72, MS80a, Men73, Mic86, MK19, Nad03, NK05a, Nad12, Nad80, NAB13, Nku23, Pré73, Röd77, SS20a, Sie14, SV18, SB17, SKPK21, TF21, Tau84, Tho78, TMY05, TP97, VK20, WYLW22, WWLH19, YAA22, ZY07, ZR10, ZBF^{+}20, ZZWG09, Zuo01, AAFM07, AJM12, AHJN01, AJ97, AA11, AA13, BGP95, BCD11, BZ14, BSS18, CB02b, DK12, KCNB11, KKG17, Mou01, MAS06, Nig90, PRS95, SP01, Sch89c, SP08, TG87, VDB13a, VDB13b]. Type-B [TG87]. Type-I [AJM12, AA13, BZ14, BSS18]. type-II [CBP23, MK19, SKPK21, VK20, ZBF^{+}20, AAFM07, BGP95, BCD11, BZ14, CB02b, DK12, KCNB11, KKG17, MAS06, SP01, SP08, VDB13a, VDB13b]. types [Deu72, KAM17].

U [SAX23]. U-Statistics [SAX23]. Übersicht [Wer78]. Übersichtsarbeiten [FM72]. ucloviyakh [Ano73c]. Uhlenbeck [AM15, Glo01]. ultivariate [Liu92]. Umbrella [KB00]. Umpu [Han96, Sei78]. UMPU-character [Sei78]. Umrandungsstruktur [Lud73]. UMVIQUE [Wen95]. UMVU [KP07, TM94]. UMVU-Estimation [TM94]. UMVUQE [Jas81]. Unabhängige [Wei70a]. Unabhängigkeitstests [Röd71]. unbalanced [KN90, WS13]. unbekannten [Exn76]. unbestimmten [Sch87a]. Unbiased [Arn92a, KS93, MD03, SM08, Sen11, WHG87, Zie88a, Zie00, BK79a, CF89, DSK85, Dry77, DH81, GKZ77, GKTZ92, KP74a, KP74b, Kle75b, KZ87, Kut12, LBSM02, Mat83, Mül85, Sah87, Waw81, WY13, WB05, ZX17]. Unbiasedness [Hel81, Ger78b, Rod88, Rüs86]. unbounded [Win17]. Uncertain [RH20, Ahm93, KCL16]. Uncertainty [ETD99, Gne98, Jen73, NCJ23, RW79]. unconditional [SS81]. uncorrelated [JR06]. Undamped [KM96]. underdispersion [TK22]. Underestimated [Fra77]. undominated [Sie14]. Unequal [SR94, AK85, BLT20, SLL15]. unequally [Jac91]. unequireplicated [Sin91].

Unforced [Vid23]. Ungleichungen [Foc75b, Thä76, FG75]. Unified [SM19, AT13, Cap19, KD16, SW89b]. Uniform [AK94, Mal21, Rya14, ACS08, Bar87, BvET16, BS07a, BS10, Che21, Col85a, ELV22, EJ81, FGL17a, FK14, JGAFBCMG14, KK14, Oko17, PP94, Ruk07, Son20, XWQN23, Zan15]. uniform-in- [Son20]. Uniformity [BL96b, WQO22, WQO23, ZYZ21, Cey20, MMP05]. Uniformly [GKZ77, WW22b, ZX17, Mej78, Mil86, YL17]. unifying [FK12]. unimodal [LMR14]. unique [BM18]. Uniqueness [Hös86, RD80, BZ14, GAJB13, Kal76, LD21, Nož70, Páz84, Wei74]. unit [Ano72a, BM09, CS73, GP82a, LLW17, RZ11, SS04, SO09, UR81]. unit-level [BM09]. units [Agr82, HDC20, Kum87, NK88]. Univariate [ML21, AAFG23, BNH20, DJ12, Jaw13, WN12b, ZX17]. Universal [LBR23]. universes [Güv06]. University [ES74]. ’unknown [Rei84, Pon02, BS10, Bro93, BW80, Exn76, GAB85, MM21, MO19, NL03, Páz91a, SGG04, Try86, ZRAN17]. unordered [BK86]. unrelated [BK79b, PS12, SKV18, Ter75]. Unschärferelation [RW79]. Unstable [TVL96, BPvZ04, LKK07]. unterer [Zin81]. Untersuchungen [EG74]. unvollstandigen [Sto72b, Sto71]. unvollständiger [Fäh78]. up-to-date [Col85b]. upcrossings [KS83b]. Upper [KR17, BR12, Dan03, Gor19, Gor20, Kum13, RN18, RM17, Ryc09, Yat19, Zin81]. uppermost [BS21]. upravleniya [Ano73g]. upside [BN16]. upside-down [BN16]. Urn [FC89a, MB21]. Use [Mej78, Páz93, Ano76, Ban80, Cey20, Deu72, Ron76, Sto88b, TC22]. Usedom [Atk75]. useful [Kim07, LCMA16]. uses [GJ73]. Using [AAK98, AS98, Art19, GD96, JS97, KPH^{+}16, Pol85, VV96, AAFM07, Ahm11, AAA13, AFSS22, BS85, BSS18, BKS23, BCDD11, BAO12, BLNGB05, CS06, CWW09, Den88, DF15, Eco13, Els17, GSA14, Iga20, JRCB22, KCL16, Kob87, Kum87, KM21, LSS12a, LUK13, LUK14, LY16, Lyt77, MB18, NSK12, NRS07, NJKK16, Nig90, OB19, PGT10, Pol08, RAM^{+}17, ST23, Sch90a, SM08, Sen11, SSG76, SKV18, SSM07, TKTM15, Wäc86, WLD18, WNBH17, WN13b, Zha12, ZL16, ZMY20]. USSR [War85b]. usual [BVP06, Lät74]. utilities [Eva75, GGHS21]. utilization [BNS73].

v [Ano72a, Ano73a, Ano73e, Ano73g, Ano76, Ano87j]. vague [AL18]. Valid [Mon96]. validated [ASFKV08]. validation [Cor17, Sto78, WC21, Zwa85]. validity [Bed81, CM15a, Fin87, Pol08]. value [AX16, Ano72a, Exn76, FYY^{+}23, Gru71, JW04b, MR75, Mau89, Yin20]. Valued [Mol96, BA89, Cad01, CFLDÁ04, CLWK20, DK21, GMLS16, GML18, GML23, RM16, Rya14, SV18]. Values [Ber98, AA03, AB10, AMB11, AA08, AAB12, Ban70a, Bar10, BN14, Bic23, Car88, Fäh78, Fäh81, Gor19, Jas18, KNS72, KAR14, Kle73, LT13, Lys76, Mau89, Mec84, OKW86, RR04, RN18, Röd70, SK20, TA11, WBL06, ZLLX08]. variability [SdSSL15]. Variable [AFV15, GLLO^{+}22, Jon95, LGZL14, XY16, ZL16, BNH20, BG78, Cad01, Ciu20, CM90, FE85, Fur15, GV19, KHN15, KT12, LFGR11, ML21, Mad76, MV10, PS14, Sep01, SS04, SMLL18, YLY17, ZX14, ZWL14, ZWYG14, Zin81, Aui90]. Variablen [Ron76, Tou70a]. Variables [BN00b, BC99, Cha15a, JK92, PGT02, AC23, Ano87j, ACS08, Aui90, Bar87, BL82, Bed86, BMRPZ16, ÇO17, CC12, CZZ19, Dav94, Ege80, FB16, FLL18, FZLB22, GAB85, HSBB23, Jas22, Kam89, Kle73, KZ13a, Kum89, KC17, KFH16, Led83, LBAR13, LLWW20, Lyt77, MGM20, MMRA07, MR12, Nad07, Páz91a, Pen85, PGT05, PN13, Pro89, Ric84, Röd70, Ron76, RW84, SV89, She14, SMLL18, Sta84, TG87, Tho13, Tou70a, Väd77, VYD83, WDZH14, Wie92, WBL06, WN13b, WN14b, WYW21, WHVW23, XY16, pZ15, ZZQ16, ZZWW23, vR88]. Variance [BC14, BDP99, Das96, Gat94, GL94, GM94, GON95a, HA02, HDA99, Imh00, Lav95, LPC97, MZ96, MSW13, Neu94, RY86, SY81, SJK97, ST95, Ste95, Tan91, ABP18, AHPS84, Ant84a, Are05, ACH20, Arn92a, ABD93, BP18, BS10, Blo03b, BS07b, Bro93, CKS22, CT14, CF89, Dan03, Dau82, Dry72, Dry77, DH81, Dry88, EBTB21, FE85, FL10, FH01, FR86, GFCO04, GKZ77, GSTUM02, Gup77, GK93, HK00, HB09, IJZ08, JR13, Jas81, Kal76, Kaw15, KP74b, Kle75a, Kle75b, Kle77a, Kle77b, KZ87, KP07, KMSS10, KZ15, Kut12, LQ12, MV18, MMP16, MM21, Mat87a, MNvRS10, MCNO19, MA86, MSW03, Net90, Ouy92, Pap16, Pin74, Pin77, Puk81a, QLT17, RZ11, Rud76, RM16, SK92, Sch76c, Sha91, Tho80, TH08, Waf88, WC21, Waw81]. variance [WB05, Wyn76, XWJ14, ZXT11, Zin81, ZL03, Zwa85, dUÁ00]. Variance-Balanced [SJK97]. variance-covariance [Dry77, HB09]. variance-mean [BP18]. Variances [BRKO99, Bun81b, Col91, Hue86, KP74b, KZ78, KJ14, KR17, Nus77b, Sch79a, Sta84, Sto83b, ZRAN17]. Varianten [Neu73]. variants [Kim07, Neu73]. Varianz [Zin81]. Variate [GV95, GN98, BN03, CG05, HG08, JG85, TH04, ZHG12]. variates [MO73]. Variation [Ber98, Ryc21, Bed89, CT03, Kha15, Luc05, WNBH17]. Various [PP92, GKZ77]. VARMA [FS02]. varying [ANN11, FX16, FB22, HLM21, Kum87, Lin09, PJ17, QJ08, TW05, Tru18, XY16, YLY17, ZZL^{+}14, ZZW^{+}19, Zha23, ZL14]. varying-coefficient [HLM21, PJ17, YLY17, ZZL^{+}14, ZL14]. VDR [TP97]. VDR-type [TP97]. vec [NFF^{+}23]. Vector [EMG98, LPC97, PSCC95, BA89, Foc73, Gru71, HT15, KS10, Koz86, Lip75a, PYHT01, SKCE19, SZ17, Sto79, Wit76a, Wit76b, Wit83b]. vectors [BMS20, BNAA13, DST21, FZB17, HE12, HK09, Hös74, Hös76, Hös77, LCY17]. Vektormaximumproblem [Foc73]. Vektorprozesse [Wit76a, Wit76b]. velichin [Ano87j]. Verallgemeinerte [Pfu78]. Verallgemeinerung [Pie74b, Thä75, FD72, Sie74]. Verbesserungen [Tau84]. verbundene [Köc75]. verbundenen [Köc76]. Verfahren [Hof72, Hol75, Hol76, Lys76, Pfu78, Bär72]. Verfahrens [BNS73, Ehl75, Ehl76]. Verflechtungskoeffizienten [Düc75b]. Verhalten [Sch75a, Sch77b]. Verknüpfung [Sti78]. Verlustsysteme [Sto72b]. veroyatnosteii [Ano73c]. versatile [MKM21]. Verschärfung [Pie70, Düc75b]. verschiedene [Deu72]. Verschleiß [PT83]. Version [Abd99, Pad96, Ars16, Pad89]. versions [Böt20, TC13]. Versorgungssystemen [SS74]. Versorgungssystems [SSG76]. Versuchsplanung [BK71, GH83, Wit78a, NN82, Wit76b]. Versus [LLK98, TJZ89b, Páz84]. Verteilung [Müh80, Wit78b, Deb77, JN75]. Verteilungen [Röd71, Wit81, RW79]. Verteilungsfunktion [Ric84]. verteilungsfunktionen [Sto72b]. Verteilungskonvergenz [Neu73]. vertex [Sch74a]. Vertical [KT96, KFL97, Koz02, Tro93, tFYK01, PYHT01]. Verweilzeit [Zäh74]. Verwendung [Deu72]. verzögertem [Sie74]. Verzögerung [RS74]. Verzweigungsprozeß [EW75b]. VI [Man72]. VI. [Man72]. Via [BL96b, BOHJ12, BCJ14, Böt20, CKM09, CGM19, CCMW09, DGCL14, HHM14, Huz18, Jia07, LD21, LWL18, Mei06, NDDDG16, NB21, Pon06, PH16b, RM17, Röd04, RALN93, Rüs86, SKSZ20, XCJH19, YA12, Yat19, ZX14]. view [AO07]. viirgish [Ano72a]. viirgisha [Ano72a]. vis [VO75]. Vitesse [Pon86]. vivodi [Ano73f]. volatility [GP19]. vollständige [Sch76b]. volume [Zha18]. volumes [BG19b]. Vom [Lom74, Ber74, Gir74, HW74, PKN72, Tau84]. vorgegebener [Mau89, Zin81]. Vorhersage [Gün77, Läu70, Tou70a, Läu71b, Läu71c, Tou70b]. Vorhersagen [Bel74]. vorliegen [JN75]. Voronoi [HKMM98]. Vorträge [Atk75]. vprosy [Ano75a]. vs [Bun75a, JY20, Sty75]. vsveshennikh [Ano87j]. Vvedeniye [Ano73g].

W [Dro91a]. Wachsenden [Lys76]. Waerden [KNS72]. Wahl [Ban70a]. Wahrscheinlichkeiten [Wol84, Got74]. Wahrscheinlichkeitstheorie [Lie72]. waiting [Aby72, BR12, CS73, Dew75, KK80, KS76b, RA10, RS74, Sch81, Sto72a]. Wald [BMMP16, KS79, Lem13, Sin05]. Wald-type [BMMP16]. Waldsche [KS79]. walk [KS79, Mik88, Ros71, Sie74]. walks [Ano81e, BTT11]. Walter [NS10b]. warming [Dew75, SD86]. Wartemodell [Dew75, Ros71]. Wartemodelle [Sto72a]. Wartemodellen [Ber74]. Warteschlangenlänge [Zäh74]. Wartesysteme [Sto73]. Wartesystemen [Sak71]. Wartezeitverteilung [RS74]. Watson [BNS15, EW75b, FG06]. Wavelet [AL00, JK08, ZL17, AAFG23, CCD15, YLWL22]. Wavelet-domain [JK08]. wavelets [Kre18]. Way [HA02, Are12, BH89, Das87, Den88, FMTZ07, GS11, HG05, KI03, Kle75b, KD74, LTV03, LC10, MNvRS10, OM11, SK92]. Weak [Dik01, DMR07, DL16, HP08, Lie98, Pra86, Ruy98, ZZWW23, BN81, Bar10, BL82, BM18, DK84, JRZ03, KW17, KH21, MV18, Son20, WLB04]. weakly [Web79, Wol75]. wear [PT83]. Weibull [ASM83, AS83, BA08, CSO13, CHO14, DK12, FB16, Ger79, HW14, Hen89, KCNB11, KM21, LNBR02, LCY15, Nig90, THN13, TDY08, Tzi87, WG21, ZZQ16]. Weibull-geometric [KM21]. Weibull-type [HW14]. Weibullverteilung [Hen89]. weighing [GSA14, HM88]. Weighted [BHS76, JK13, LdCIP18, LY16, AAR18, AC09, ACS05, Art19, BS80, CH89, CGS23, Cup21, DWHH19, DST21, GK09, HKT22, HX20, IRB14, Kim07, Kim10, LLWW20, MH22, NS03, NRD06, PS10, SP08, SDC19, TMN00, Van91, WWH18, Wol85, XWQN23, ZZWW23]. weighting [PW15, ZTJ17]. weights [KS10, Sin13]. Weimar [Gir74]. Weinberg [GM88]. Welch [HA02]. well [Dew76, LLBM^{+}11]. well-known [LLBM^{+}11]. well-orderings [Dew76]. Werte [Fäh78, Fäh81, KNS72]. Werten [Röd70]. Wertvektor [Gru71]. Wetherill [Hec86]. where [LCY17]. Whether [Str95, Ger86, LL85]. which [Als93, KZ87, Ros72, UR81]. White [Tur90, IF03]. Whittaker [Eic88]. Whittle [NR81]. whose [LMS06, Yin20]. Width [CR99]. Wiederherstellung [SW85]. Wiener [Ros71, BPV11, JSW22, KK11, Koz86, Ros71, Sch70a]. Wiener-Prozessen [Sch70a]. Wilcoxon [KM20, MM18, Mur15, QLY21, TMN00, Yu04]. wild [KPH^{+}16]. Wilks [Sch76d, TG87]. Willan [Tik09]. Wilrich [Hec86]. window [Col85a]. Winsorized [BW14, LBSM00]. Wintgen [Bod75]. Wishart [NFF^{+}23, vR97b]. Wissenschaftliche [ES74]. within [Ros91]. ’without [BR83, GB23, KHN15, KD74, Que86, Ric76b, RS74, SD86]. WLSEs [TT10a]. Wohlordnungen [Dew76]. Wolfe [Jen73]. Wolfowitz [Her81]. work [Gus71, Sch76a, Sto93, FM72]. workforce [Tit70]. working [BK71, HW70a, HW74, Jas22]. Workload [Tij76]. Workshop [BM71, HW70b, HW72, Els75]. worth [AMSV12]. wrapped [PJ05]. wrong [WZW21].

X [Mou01]. XII [AJM12, AJ97, WWC^{+}10]. XVIII [ES74]. XVIII. [ES74].

years [FG75]. Ylvisaker [Wit76b, Wit76b, Wit78a]. Young [Kin76]. Yule [Xek86].

zadachakh [Ano73a]. zadachye [Ano73b]. Zadov [Lom74]. Zanten [FGL17b]. zdacham [Ano75b]. zeitabhängigen [Dam76]. zeitabhängiger [Fre74, Seb76]. Zeitbereichsanalyse [Sti78]. Zeitreihen [Web79]. Zeitreihenanalyse [Sti78]. zensierten [Sch87a, Sch89a]. zensierter [Hen89]. zentrale [Wol84]. zentralen [Pad81]. zentrenmethode [GR72]. Zentrenmethoden [EG74]. Zerlegung [BTS71, Pfl80]. Zerlegungsmethode [ST75]. Zernike [DMP07]. Zero [GMLS16, GML18, GML23, HA02, DDD19, JN75, KMSS10, LLC16, Men73]. Zero-inflated [GMLS16, GML23, DDD19, KMSS10, LLC16]. zero-one [Men73]. zero-point [JN75]. Zero-truncated [GML18]. Zielfunktion [Ric76b, Tam74b, Wei70b]. Zinnowitz [Atk75]. Zinnowitz/Usedom [Atk75]. Zolotarev [BV86]. Zoutendijk [Foc71]. Zowe [Neh75]. Zucchini [Dro91a]. zufällig [Sch87a]. zufällige [KS79]. zufälligem [Gla76]. zufälligen [Gün77]. zufälliger [Hoy79, Wit76a, Wit78b]. Zufallsgröße [Zin81]. Zufallsgroßen [Ric84, Röd70]. Zufallsprozesse [Fäh81]. Zufallsprozessen [Fäh78]. Zufallsvektoren [Hös76]. Zugang [Hös76]. zulässigen [Hol75, Hol76, Pré74a]. zum [Hol73a, Lys76, Pie75a, Ron76, EN74a]. zur [Bär72, Deu72, Ehl75, Ehl76, FN77, Fra77, Gir71, GT70, HT71, HP73, HW71, Hoy79, Mel76, Ric76a, Sch70b, Sch73, ST75, Sch75f, SS74, SSG76, Sti78, Tam76, Bär76, BG74b, DB74, Dew73, ES73, EW73, EN74b, Fäh78, Fäh81, Fle71, Foc84, GH83, GM78, Gri80, GR72, Gün77, HK78, Kin75, Lie72, Neu73, Pie72b, Ric76b, Sch74a, Ste76, Sti78, Tit70, Wag83, Wer78, vW72a, vW72b]. Zurückführung [vW74]. Zusammenhang [Ull76b]. zusammenhängenden [Ros71]. zustandsgleichung [Ber76]. Zustandsraum [Küc78, KS79]. Zustandswahrscheinlichkeiten [JK76]. Zuverlässigkeitsmodelle [PT83]. Zuverlässigkeitssysteme [KJKR76]. Zuverlässigkeitstheorie [Bei74]. Zwanzig [Zwa83]. zwei [Ban70b, Sch76b]. Zweidimensionale [Röd71]. zweiparametrisches [Gre87]. Zweistichprobenfall [Kin75]. zwischen [Ull76b]. Zygmund [Zäh14].

References

Andel:1998:BR

[AA98] Jiri Andel and Jiri Andel. Book reviews. Statistics: a Journal of Theoretical and Applied Statistics, 31(2):187–188, 1998. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Ahmadi:2003:CFI

[AA03] Jafar Ahmadi and N. R. Arghami. Comparing the Fisher information in record values and iid observations. Statistics: a Journal of Theoretical and Applied Statistics, 37(5):435–441, 2003. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Amini:2008:CFI

[AA08] Morteza Amini and J. Ahmadi. Comparing Fisher information in record values and their concomitants with random observations. Statistics: a Journal of Theoretical and Applied Statistics, 42(5):393–405, 2008. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Ateya:2011:IBG

[AA11] Saieed F. Ateya and Abd El-Baset A. Ahmad. Inferences based on generalized order statistics under truncated Type I generalized logistic distribution. Statistics: a Journal of Theoretical and Applied Statistics, 45(4):389–402, 2011. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Abdel-Aty:2012:BPF

[AA12] Yahia Abdel-Aty. Bayesian prediction of future number of failures based on finite mixture of general class of distributions. Statistics: a Journal of Theoretical and Applied Statistics, 46(1):111–122, 2012. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Ateya:2013:IBG

[AA13] Saieed F. Ateya and Abd El-Baset A. Ahmad. Inferences based on generalized order statistics under truncated Type-I generalized logistic distribution. Statistics: a Journal of Theoretical and Applied Statistics, 47(1):141–155, 2013. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Altaf:2013:BAA

[AAA13] Saima Altaf, Muhammad Aslam, and Muhammad Aslam. Bayesian analysis of the amended Davidson model for paired comparison using noninformative and informative priors. Statistics: a Journal of Theoretical and Applied Statistics, 47(5):1090–1103, 2013. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Amini:2012:FIB

[AAB12] Morteza Amini, Jafar Ahmadi, and N. Balakrishnan. Fisher information in bivariate record values from a sample of fixed size. Statistics: a Journal of Theoretical and Applied Statistics, 46(1):23–39, 2012. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Alvarez-Andrade:2018:ASC

[AAB18] Sergio Alvarez-Andrade and Salim Bouzebda. Almost sure central limit theorem for the hybrid process. Statistics: a Journal of Theoretical and Applied Statistics, 52(3):519–532, 2018. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Amato:2023:PWN

[AAFG23] Umberto Amato, Anestis Antoniadis, Italia De Feis, and Irène Gijbels. Penalized wavelet nonparametric univariate logistic regression for irregular spaced data. Statistics: a Journal of Theoretical and Applied Statistics, 57(5):1037–1060, 2023. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/02331888.2023.2248679.

Abdel-Aty:2007:BPB

[AAFM07] Y. Abdel-Aty, J. Franz, and M. A. W. Mahmoud. Bayesian prediction based on generalized order statistics using multiply Type-II censoring. Statistics: a Journal of Theoretical and Applied Statistics, 41(6):495–504, 2007. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Aerts:1997:LPE

[AAJ97] M. Aerts, I. Augustyns, and P. Janssen. Local polynomial estimation of contingency table Cell probabilities. Statistics: a Journal of Theoretical and Applied Statistics, 30(2):127– 148, 1997. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Albers:1998:TLU

[AAK98] Willem Albers, Gerda R. J. Arts, and Wilbert C. M. Kallenberg. Test limits using correlated measurements. Statistics: a Journal of Theoretical and Applied Statistics, 30(4):307–330, 1998. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Abdelrazeq:2020:OSB

[AALA20] Ibrahim Abdelrazeq, Luai Al-Labadi, and Ayman Alzaatreh. On one-sample Bayesian tests for the mean. Statistics: a Journal of Theoretical and Applied Statistics, 54(2):424–440, 2020. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Akdeniz:2018:NDB

[AAR18] Esra Akdeniz, Fikri Akdeniz, and Mahdi Roozbeh. A new difference-based weighted mixed Liu estimator in partially linear models. Statistics: a Journal of Theoretical and Applied Statistics, 52(6):1309–1327, 2018. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Ahmadi:2005:PSR

[AB05] J. Ahmadi and N. Balakrishnan. Preservation of some reliability properties by certain record statistics. Statistics: a Journal of Theoretical and Applied Statistics, 39(4):347–354, 2005. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Ahmadi:2010:POS

[AB10] Jafar Ahmadi and N. Balakrishnan. Prediction of order statistics and record values from two independent sequences. Statistics: a Journal of Theoretical and Applied Statistics, 44(4): 417–430, 2010. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Asadi:2012:PEM

[AB12] Majid Asadi and Alexandre Berred. Properties and estimation of the mean past lifetime. Statistics: a Journal of Theoretical and Applied Statistics, 46(3):405–417, 2012. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Azais:1993:MML

[ABD93] J. M. Azaïs, A. Bardin, and T. Dhorne. MINQE, maximum likelihood estimation and Fisher scoring algorithm for non linear variance models. Statistics: a Journal of Theoretical and Applied Statistics, 24(3):205–213, 1993. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Abdous:1999:VDK

[Abd99] Belkacem Abdous. L_{2} version of the double kernel method. Statistics: a Journal of Theoretical and Applied Statistics, 32(3):249–266, 1999. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Afendras:2018:OPC

[ABP18] Georgios Afendras, Narayanaswamy Balakrishnan, and Nickos Papadatos. Orthogonal polynomials in the cumulative Ord family and its application to variance bounds. Statistics: a Journal of Theoretical and Applied Statistics, 52(2): 364–392, 2018. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Abya:1972:NWL

[Aby72] K. L. Abya. Network of waiting lines. Mathematische Operationsforschung und Statistik, 3(5):363–371, 1972. CODEN MTOSBI. ISSN 0047-6277.

Alavi:2009:FIU

[AC09] S. M. R. Alavi and R. Chinipardaz. Form-invariance under weighted sampling. Statistics: a Journal of Theoretical and Applied Statistics, 43(1):81–90, 2009. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Alcaraz:2023:ASP

[AC23] Angelo Alcaraz and Gabriela Ciuperca. Automatic selection by penalized asymmetric L_{q}-norm in a high-dimensional model with grouped variables. Statistics: a Journal of Theoretical and Applied Statistics, 57(5):1202–1238, 2023. CO-DEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/02331888.2023.2256948.

Achcar:1995:IAL

[Ach95] Jorge Alberto Achcar. Inferences for accelerated life tests considering a bivariate exponential distribution. Statistics: a Journal of Theoretical and Applied Statistics, 26(3):269–283, 1995. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Arias-Castro:2020:SVC

[ACH20] Ery Arias-Castro and Rong Huang. The sparse variance contamination model. Statistics: a Journal of Theoretical and Applied Statistics, 54(5):1081–1093, 2020. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Arnold:1993:CEF

[ACS93] Barry C. Arnold, Enrique Castillo, and Jose María Sarabia. Conjugate exponential family priors for exponential family likelihoods. Statistics: a Journal of Theoretical and Applied Statistics, 25(1):71–77, 1993. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Arnold:1996:PCP

[ACS96] Barry C. Arnold, Enrique Castillo, and José María Sarabia. Priors with convenient posteriors. Statistics: a Journal of Theoretical and Applied Statistics, 28(4):347–354, 1996. CO-DEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Arnold:2005:DCT

[ACS05] Barry C. Arnold, Enrique Castillo, and José María Sarabia. Distributions with conditionals in truncated weighted families. Statistics: a Journal of Theoretical and Applied Statistics, 39(2):133–147, 2005. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Arnold:2008:SCI

[ACS08] Barry C. Arnold, Enrique Castillo, and José María Sarabia. Some characterizations involving uniform and powers of uniform random variables. Statistics: a Journal of Theoretical and Applied Statistics, 42(6):527–534, 2008. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Axt:2021:RSE

[ADF21] I. Axt, Alexander Dürre, and Roland Fried. Robust scale estimation under shifts in the mean. Statistics: a Journal of Theoretical and Applied Statistics, 55(4):787–830, 2021. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Azais:2014:LRT

[ADR14] Jean-Marc Azaïs, Céline Delmas, and Charles-Elie Rabier. Likelihood ratio test process for quantitative trait locus detection. Statistics: a Journal of Theoretical and Applied Statistics, 48(4):787–801, 2014. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Azarang:2015:JEC

[AdUÁS15] Leyla Azarang, Jacobo de Uña-Álvarez, and Winfried Stute. The Jackknife estimate of covariance of two Kaplan–Meier integrals with covariables. Statistics: a Journal of Theoretical and Applied Statistics, 49(5):1005–1025, 2015. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Azimi:2021:CBG

[AE21] Reza Azimi and Mahdy Esmailian. Corrigendum to: A bimodal gamma distribution: properties, regression model and applications (Statistics, 54, 469–493). Statistics: a Journal of Theoretical and Applied Statistics, 55(4):733–735, 2021. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic). See [VFS^{+}20].

Andel:1986:BR

[AEH^{+}86] J. Andel, H. J. Engelbert, K.-H. Hanisch, K. Wendler, G. Richter, F. Liese, W. Winkler, W. Winkler, O. Bunke, O. Bunke, B. Lisek, R. Cünther, W. Schmidt, W. Schmidt, O. Bunke, W. Schmidt, W. Schmidt, P. Neumann, W. Lipfert, and V. Nollau. Book reviews. Statistics: a Journal of Theoretical and Applied Statistics, 17(3):461–480, 1986. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Ashour:1994:BPM

[AEW94] S. K. Ashour and M. A. M. H. El-Wakeel. Bayesian prediction of the median of the Burr distribution with fixed and random sample sizes. Statistics: a Journal of Theoretical and Applied Statistics, 25(2):113–122, 1994. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Akbari:2014:CRB

[AF14] M. Akbari and M. Fashandi. On characterization results based on the number of observations near the k-records. Statistics: a Journal of Theoretical and Applied Statistics, 48(3):633–640, 2014. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Akinsete:2008:BPD

[AFL08] Alfred Akinsete, Felix Famoye, and Carl Lee. The beta-Pareto distribution. Statistics: a Journal of Theoretical and Applied Statistics, 42(6):547–563, 2008. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Afsarinejad:1985:ORM

[Afs85] K. Afsarinejad. Optimal repeated measurements designs. Statistics: a Journal of Theoretical and Applied Statistics, 16(4):563–568, 1985. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Amiri:2022:ELA

[AFSS22] Narjes Amiri, Vahid Fakoor, Majid Sarmad, and Ali Shariati. Empirical likelihood analysis for accelerated failure time model using length-biased data. Statistics: a Journal of Theoretical and Applied Statistics, 56(3):578–597, 2022. CO-DEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/02331888.2022.2077334.

Aneiros:2015:VSP

[AFV15] G. Aneiros, F. Ferraty, and P. Vieu. Variable selection in partial linear regression with functional covariate. Statistics: a Journal of Theoretical and Applied Statistics, 49(6):1322– 1347, 2015. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Andel:1991:BAN

[AG91] Jiri Andel and Manuel Garrido. Bayesian analysis of nonnegative AR(2) processes. Statistics: a Journal of Theoretical and Applied Statistics, 22(4):579–588, 1991. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Arslan:2009:SGD

[AG09] Olcay Arslan and Ali I. Genç. The skew generalized t distribution as the scale mixture of a skew exponential power distribution and its applications in robust estimation. Statistics: a Journal of Theoretical and Applied Statistics, 43(5): 481–498, 2009. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Amorino:2023:EID

[AG23] Chiara Amorino and Arnaud Gloter. Estimation of the invariant density for discretely observed diffusion processes: impact of the sampling and of the asynchronicity. Statistics: a Journal of Theoretical and Applied Statistics, 57(1):213–259, 2023. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/02331888.2023.2166047.

Arnold:2014:MCT

[AGOP14] Barry C. Arnold, Héctor W. Gómez, and Juan F. Olivares-Pacheco. Multiple constraint and truncated skew models. Statistics: a Journal of Theoretical and Applied Statistics, 48(5):971–982, 2014. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Agrawal:1982:ADU

[Agr82] M. C. Agrawal. On averaging over distinct units in replicated samples. Statistics (Berlin, DDR), 13(1):91–97, 1982. CO-DEN MOSSD5. ISSN 0323-3944.

Arnold:2009:MCS

[AGS09] Barry C. Arnold, Héctor W. Gómez, and Hugo S. Salinas. On multiple constraint skewed models. Statistics: a Journal of Theoretical and Applied Statistics, 43(3):279–293, 2009.

CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Arnold:2015:DSN

[AGS15] Barry C. Arnold, Héctor W. Gómez, and Hugo S. Salinas. A doubly skewed normal distribution. Statistics: a Journal of Theoretical and Applied Statistics, 49(4):842–858, 2015. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Al-Hussaini:1999:PPB

[AHJ99] Essam K. Al-Hussaini and Zeinhum F. Jaheen. Parametric prediction bounds for the future median of the exponential distribution. Statistics: a Journal of Theoretical and Applied Statistics, 32(3):267–275, 1999. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Al-Hussaini:2001:BPB

[AHJN01] Essam K. Al-Hussaini, Z. F. Jaheen, and A. M. Nigm. Bayesian prediction based on finite mixtures of Lomax components model and Type I censoring. Statistics: a Journal of Theoretical and Applied Statistics, 35(3):259–268, 2001. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Ahmad:1974:DFS

[Ahm74] R. Ahmad. Distribution-free statistical hypotheses testing for stochastic processes. Mathematische Operationsforschung und Statistik, 5(7–8):643–656, 1974. CODEN MTOSBI. ISSN 0047-6277.

Ahmed:1993:PMU

[Ahm93] S. E. Ahmed. Pooling means under uncertain prior information with application to discrete distributions. Statistics: a Journal of Theoretical and Applied Statistics, 24(3):265–277, 1993. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Ahmad:2001:MOS

[Ahm01] Abd El-Baset A. Ahmad. Moments of order statistics from doubly truncated continuous distributions and characterizations. Statistics: a Journal of Theoretical and Applied Statistics, 35(4):479–494, 2001. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Ahmad:2011:BIP

[Ahm11] Abd El-Baset A. Ahmad. On Bayesian interval prediction of future generalized-order statistics using doubly censoring. Statistics: a Journal of Theoretical and Applied Statistics, 45(5):413–425, 2011. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Anderson:1984:BEV

[AHPS84] R. D. Anderson, H. V. Henderson, F. Pukelsheim, and S. R. Searle. Best estimation of variance components from balanced data, with arbitrary kurtosis. Statistics (Berlin, DDR), 15(2): 163–176, 1984. CODEN MOSSD5. ISSN 0323-3944.

Arashi:2014:BAM

[AINJ14] M. Arashi, Anis Iranmanesh, M. Norouzirad, and Hashem Salarzadeh Jenatabadi. Bayesian analysis in multivariate regression models with conjugate priors. Statistics: a Journal of Theoretical and Applied Statistics, 48(6):1324–1334, 2014. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

AliMousa:1997:BPB

[AJ97] M. A. M. Ali Mousa and Z. F. Jaheen. Bayesian prediction for the Burr Type XII model based on doubly censored data. Statistics: a Journal of Theoretical and Applied Statistics, 29(3):285–294, 1997. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Ahmad:2012:BPB

[AJM12] K. E. Ahmad, Z. F. Jaheen, and Heba S. Mohammed. Bayesian prediction based on Type-I censored data from a mixture of Burr type XII distribution and its reciprocal. Statistics: a Journal of Theoretical and Applied Statistics, 46(6):787–797, 2012. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Abrams:2021:QCR

[AJV21] S. Abrams, P. Janssen, and N. Veraverbeke. Quantiles of the conditional residual lifetime. Statistics: a Journal of Theoretical and Applied Statistics, 55(6):1271–1290, 2021. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Ahmetov:1975:CEP

[AK75] O. A. Ahmetov and I. A. Krass. On a certain extremal problem in inequalities in a Banach space. Mathematische Operationsforschung und Statistik, 6(3):341–354, 1975. CODEN MTOSBI. ISSN 0047-6277.

Agarwal:1985:RPE

[AK85] S. K. Agarwal and Pranesh Kumar. A ratio cum PPS estimator in unequal probability sampling. Statistics: a Journal of Theoretical and Applied Statistics, 16(4):577–580, 1985. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Akahira:1994:CML

[AK94] Masafumi Akahira and Hiroyuki Kashima. On the consistency of the maximum likelihood estimator through its uniform consistency. Statistics: a Journal of Theoretical and Applied Statistics, 25(4):333–341, 1994. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Allal:2003:AEL

[AK03] J. Allal and A. Kaaouachi. Adaptive R-estimation in a linear regression model with ARMA errors. Statistics: a Journal of Theoretical and Applied Statistics, 37(4):271–286, 2003. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Albers:2004:ECC

[AK04] Willem Albers and Wilbert C. M. Kallenberg. Are estimated control charts in control? Statistics: a Journal of Theoretical and Applied Statistics, 38(1):67–79, 2004. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Asar:2020:JRE

[AK20] Yasin Asar and Kadriye Kilinç. A jackknifed ridge estimator in probit regression model. Statistics: a Journal of Theoretical and Applied Statistics, 54(4):667–685, 2020. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Akaeel:1982:FMT

[Aka82] Jiří Akaeel. Fitting models in time series analysis. Statistics (Berlin, DDR), 13(1):121–143, 1982. CODEN MOSSD5. ISSN 0323-3944.

Albers:2005:EPP

[AKN05] Willem Albers, Wilbert C. M. Kallenberg, and Sri Nurdiati. Exceedance probabilities for parametric control charts. Statistics: a Journal of Theoretical and Applied Statistics, 39(5): 429–443, 2005. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Auert:1982:STR

[AL82] F. Auert and H. Läuter. Some theoretical results on approximation methods. Statistics (Berlin, DDR), 13(2):265–293, 1982. CODEN MOSSD5. ISSN 0323-3944.

Arfi:1999:EDC

[AL99] Mounir Arfi and Jean-Pierre Lecoutre. Estimation of the diffusion coefficient under strong mixing. Statistics: a Journal of Theoretical and Applied Statistics, 33(1):73–84, 1999. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Antoniadis:2000:NWR

[AL00] A. Antoniadis and F. Leblanc. Nonparametric wavelet regression for binary response. Statistics: a Journal of Theoretical and Applied Statistics, 34(3):183–213, 2000. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Al-Labadi:2018:MVC

[AL18] Luai Al-Labadi. On metrizing vague convergence of random measures with applications on Bayesian nonparametric models. Statistics: a Journal of Theoretical and Applied Statistics, 52(2):445–457, 2018. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Albers:1984:BEB

[Alb84] W. Albers. A Berry–Esséen bound for scale-rank tests with nuisance medians. Statistics (Berlin, DDR), 15(3):355–367, 1984. CODEN MOSSD5. ISSN 0323-3944.

Alson:1988:MPL

[Als88] P. Alson. Minimax properties for linear estimators of the location parameter of a linear model. Statistics: a Journal of Theoretical and Applied Statistics, 19(2):163–171, 1988. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Alson:1993:CEW

[Als93] Pedro Alson. Centred ellipsoids for which an admissible linear estimator is the minimax linear estimator. Statistics: a Journal of Theoretical and Applied Statistics, 24(2):85–94, 1993. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Avram:2011:PEF

[ALS11] F. Avram, N. N. Leonenko, and N. Suvak. Parameter estimation for Fisher–Snedecor diffusion. Statistics: a Journal of Theoretical and Applied Statistics, 45(1):27–42, 2011. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Aly:1987:QPD

[Aly87] Eman-Eldin A. A. Aly. On quantile processes for m-dependent Rv’s. Statistics: a Journal of Theoretical and Applied Statistics, 18(3):423–435, 1987. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Averous:1990:LST

[AM90] Jean Averous and Michel Meste. Location, skewness and tailweight in l_{s} sense: a coherent approach. Statistics: a Journal of Theoretical and Applied Statistics, 21(1):57–74, 1990. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Azmoodeh:2015:DPE

[AM15] Ehsan Azmoodeh and José Igor Morlanes. Drift parameter estimation for fractional Ornstein–Uhlenbeck process of the second kind. Statistics: a Journal of Theoretical and Applied Statistics, 49(1):1–18, 2015. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Amini:2014:LGG

[AMA14] Morteza Amini, S. M. T. K. MirMostafaee, and J. Ahmadi. Log-gamma-generated families of distributions. Statistics: a Journal of Theoretical and Applied Statistics, 48(4):913–932, 2014. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Ahmadi:2011:BPO

[AMB11] Jafar Ahmadi, S. M. T. K. MirMostafaee, and N. Balakrishnan. Bayesian prediction of order statistics based on k-record values from exponential distribution. Statistics: a Journal of Theoretical and Applied Statistics, 45(4):375–387, 2011. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Ammous:2020:TKL

[Amm20] Sinda Ammous. Testing Kendall’s τ for a large class of dependent sequences. Statistics: a Journal of Theoretical and Applied Statistics, 54(4):686–713, 2020. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Abdelghani:2021:SEI

[AMP21] Mohamed Abdelghani, Alexander Melnikov, and Andrey Pak. On statistical estimation and inferences in optional regression models. Statistics: a Journal of Theoretical and Applied Statistics, 55(2):445–457, 2021. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Al-Mosawi:2012:AWE

[AMSV12] Riyadh R. Al-Mosawi, A. Shanubhogue, and P. Vellaisamy. Average worth estimation of the selected subset of Poisson populations. Statistics: a Journal of Theoretical and Applied Statistics, 46(6):813–831, 2012. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Asgharian:1998:NFR

[AN98] M. Asgharian and S. Noorbaloochi. Note on a fundamental relationship between admissible and Bayesian decision rules. Statistics: a Journal of Theoretical and Applied Statistics, 31 (1):21–34, 1998. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Abraham:2013:CDA

[AN13] Bovas Abraham and N. Unnikrishnan Nair. A criterion to distinguish ageing patterns. Statistics: a Journal of Theoretical and Applied Statistics, 47(1):85–92, 2013. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Andel:1976:ASR

[And76a] Jirí Andel. Autoregressive series with random parameters. Mathematische Operationsforschung und Statistik, 7(5):735– 741, 1976. CODEN MTOSBI. ISSN 0047-6277.

Anderson:1976:FRF

[And76b] O. D. Anderson. Feasible regions for the first pair of autocorrelations of moving average processes. Mathematische Operationsforschung und Statistik, 7(1):85–93, 1976. CODEN MTOSBI. ISSN 0047-6277.

Anderson:1977:TSC

[And77] O. D. Anderson. The time series concept of invertibility. Statistics (Berlin, DDR), 8(3):399–406, 1977. CODEN MOSSD5. ISSN 0323-3944.

Andel:1978:MTM

[And78a] Jiří Andel. Modern trends in multivariate time-series analysis. Statistics (Berlin, DDR), 9(1):141–158, 1978. CODEN MOSSD5. ISSN 0323-3944.

Anderson:1978:ASM

[And78b] O. D. Anderson. The algebraic structure of moving average time processes. Statistics (Berlin, DDR), 9(2):211–216, 1978. CODEN MOSSD5. ISSN 0323-3944.

Anderson:1978:ICM

[And78c] O. D. Anderson. On the invertibility conditions for moving average processes. Statistics (Berlin, DDR), 9(4):525–529, 1978. CODEN MOSSD5. ISSN 0323-3944.

Andel:1979:MDD

[And79a] Jiří Andel. Measures of dependence in discrete stationary processes. Statistics (Berlin, DDR), 10(1):107–126, 1979. CO-DEN MOSSD5. ISSN 0323-3944.

Anderson:1979:BSA

[And79b] O. D. Anderson. On the bias of sampled autocorrelations. Statistics (Berlin, DDR), 10(2):299–305, 1979. CODEN MOSSD5. ISSN 0323-3944.

Anderson:1979:PAO

[And79c] O. D. Anderson. On the partial autocorrelations of once integrated autoregressive-moving average processes. Statistics (Berlin, DDR), 10(3):389–394, 1979. CODEN MOSSD5. ISSN 0323-3944.

Andel:1980:ETD

[And80] Jiří Andel. On extrapolation in two-dimensional stationary processes. Statistics (Berlin, DDR), 11(3):315–326, 1980. CO-DEN MOSSD5. ISSN 0323-3944.

Andel:1981:SMA

[And81] Jiří Andel. On shifted multiple ARMA processes. Statistics (Berlin, DDR), 12(2):183–191, 1981. CODEN MOSSD5. ISSN 0323-3944.

Andel:1987:BR

[And87] J. Andel. Book review. Statistics: a Journal of Theoretical and Applied Statistics, 18(2):292, 1987. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Andel:1989:NMT

[And89] Jiri Andel. On nonlinear models for time series. Statistics: a Journal of Theoretical and Applied Statistics, 20(4):615–632, 1989. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Andel:1990:NPP

[And90] Jiri Andel. Nonlinear positive AR(2) processes. Statistics: a Journal of Theoretical and Applied Statistics, 21(4):591–600, 1990. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Abegaz:2011:SST

[ANN11] Fentaw Abegaz and U. V. Naik-Nimbalkar. Semiparametric score test for varying copula parameter in Markov time series. Statistics: a Journal of Theoretical and Applied Statistics, 45(3):209–222, 2011. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Anonymous:1972:SSI

[Ano72a] Anonymous. O sushchestvovanii snacheniya igri v igrpax na edinichnom kvadrate c rasrivnimi funktiayami viirgisha. (Russian) [On the existence of the value of the game of games on the unit square with discontinuous viirgish (??) functions]. Mathematische Operationsforschung und Statistik, 3 (2):91–96, 1972. CODEN MTOSBI. ISSN 0047-6277.

Anonymous:1972:SDT

[Ano72b] Anonymous. Smeshannaya duel tipa 2 × 1. (Russian) [Mixed duel type 2 × 1]. Mathematische Operationsforschung und Statistik, 3(2):97–102, 1972. CODEN MTOSBI. ISSN 0047-6277.

Anonymous:1973:CMV

[Ano73a] Anonymous. Chislenniye metodi v zadachakh issledovaniya operatsii. (Russian) [Numerical methods in operations research problems]. Mathematische Operationsforschung und Statistik, 5(4–5):399–412, 1973. CODEN MTOSBI. ISSN 0047-6277.

Anonymous:1973:PTP

[Ano73b] Anonymous. O postroennii tseni s(x) pri zadachye ob optimalnoi ostanovkye markovskoi posledovatyelnosti. (Russian) [On pricing s(x) with the problem of optimal stopping of a Markov sequence]. Mathematische Operationsforschung und Statistik, 6(3):493–498, 1973. CODEN MTOSBI. ISSN 0047-6277.

Anonymous:1973:OUC

[Ano73c] Anonymous. Ob ucloviyakh cyshchestvovaniya finalnikh veroyatnosteii y markovskovo protsessa. (Russian) [On the conditions for the existence of final probabilities and the Markov process]. Mathematische Operationsforschung und Statistik, 4(5):379–390, 1973. CODEN MTOSBI. ISSN 0047-6277.

Anonymous:1973:ONA

[Ano73d] Anonymous. Operatsii nad antagonisticheskimi igrami. (Russian) [Antagonistic game operations]. Mathematische Operationsforschung und Statistik, 4(5):371–377, 1973. CODEN MTOSBI. ISSN 0047-6277.

Anonymous:1973:PMD

[Ano73e] Anonymous. Printsip mnovostupenchatoii dekompositsii v lineiinom programmirovanii. (Russian) [The principle of multi-stage decomposition in linear programming]. Mathematische Operationsforschung und Statistik, 4(6):427–443, 1973. CODEN MTOSBI. ISSN 0047-6277.

Anonymous:1973:SVR

[Ano73f] Anonymous. Statisticheskiye vivodi o raspredeleniyaks na tsiklicheskoi gruppe. (Russian) [Statistical findings for distributions on a cyclic group]. Mathematische Operationsforschung und Statistik, 6(2):225–238, 1973. CODEN MTOSBI. ISSN 0047-6277.

Anonymous:1973:VVT

[Ano73g] Anonymous. Vvedeniye v teoriyu ierarkhicheskikh sistem upravleniya. Mathematische Operationsforschung und Statistik, 4(2):133–154, 1973. CODEN MTOSBI. ISSN 0047-6277.

Anonymous:1974:MPE

[Ano74] Anonymous. Metod planirovaniya eksperimenta dla sluchaya nelineinoi parametrizatsii. (Russian) [experiment planning method for the case of nonlinear parameterization]. Mathematische Operationsforschung und Statistik, 5(7–8):625–636, 1974. CODEN MTOSBI. ISSN 0047-6277.

Anonymous:1975:KVR

[Ano75a] Anonymous. K vprosy o rasreshimosti beskoalitsionnikh irp n lits. (Russian) [On questions about the resolvability of noncooperative entities]. Mathematische Operationsforschung und Statistik, 6(4):627–639, 1975. CODEN MTOSBI. ISSN 0047-6277.

Anonymous:1975:MCM

[Ano75b] Anonymous. Markovskiye clucaiiniye markirovanniye potoki i ikh prilozhenniye k zdacham teorii massovovo obslyzhivaniya. (Russian) [Markov random tagged streams applied to queuing theory problems]. Mathematische Operationsforschung und Statistik, 6(3):445–477, 1975. CODEN MTOSBI. ISSN 0047-6277.

Anonymous:1975:RNM

[Ano75c]  Anonymous.  O razmernosti nekotorikh maksimalnikh

lineinikh prostranstv igr, imeyushchikh cedlovuyu tochku. (Russian) [On the dimension of some maximal linear spaces of games having a saddle point]. Mathematische Operationsforschung und Statistik, 6(3):355–365, 1975. CODEN MTOSBI. ISSN 0047-6277.

Anonymous:1976:OIF

[Ano76] Anonymous. Ob ispolzovanii funktsii biigrisha v teorii statisticheskovo otsenivaniya. (Russian) [On the use of the game function in the theory of statistical estimation]. Mathematische Operationsforschung und Statistik, 7(5):729–733, 1976. CODEN MTOSBI. ISSN 0047-6277.

Anonymous:1977:EBa

[Ano77a] Anonymous. Editorial board. Statistics (Berlin, DDR), 8(2): ebi, 1977. CODEN MOSSD5. ISSN 0323-3944.

Anonymous:1977:EBb

[Ano77b] Anonymous. Editorial board. Statistics (Berlin, DDR), 8(3): ebi, 1977. CODEN MOSSD5. ISSN 0323-3944.

Anonymous:1977:EBc

[Ano77c] Anonymous. Editorial board. Statistics (Berlin, DDR), 8(4): ebi, 1977. CODEN MOSSD5. ISSN 0323-3944.

Anonymous:1978:EBa

[Ano78a] Anonymous. Editorial board. Statistics (Berlin, DDR), 9(1): ebi, 1978. CODEN MOSSD5. ISSN 0323-3944.

Anonymous:1978:EBb

[Ano78b] Anonymous. Editorial board. Statistics (Berlin, DDR), 9(2): ebi, 1978. CODEN MOSSD5. ISSN 0323-3944.

Anonymous:1978:EBc

[Ano78c] Anonymous. Editorial board. Statistics (Berlin, DDR), 9(3): ebi, 1978. CODEN MOSSD5. ISSN 0323-3944.

Anonymous:1978:EBd

[Ano78d] Anonymous. Editorial board. Statistics (Berlin, DDR), 9(4): ebi, 1978. CODEN MOSSD5. ISSN 0323-3944.

Anonymous:1979:EBa

[Ano79a] Anonymous. Editorial board. Statistics (Berlin, DDR), 10 (1):ebi, 1979. CODEN MOSSD5. ISSN 0323-3944.

Anonymous:1979:EBb

[Ano79b] Anonymous. Editorial board. Statistics (Berlin, DDR), 10 (2):ebi, 1979. CODEN MOSSD5. ISSN 0323-3944.

Anonymous:1979:EBc

[Ano79c] Anonymous. Editorial board. Statistics (Berlin, DDR), 10 (3):ebi, 1979. CODEN MOSSD5. ISSN 0323-3944.

Anonymous:1979:EBd

[Ano79d] Anonymous. Editorial board. Statistics (Berlin, DDR), 10 (4):ebi, 1979. CODEN MOSSD5. ISSN 0323-3944.

Anonymous:1979:OPT

[Ano79e] Anonymous. Odna predlnaya teorema dlya symm slycainovo chisla slychainikh salagaemiks. (Russian) []. Mathematische Operationsforschung und Statistik, 10(3):461–467, 1979. CO-DEN MTOSBI. ISSN 0047-6277.

Anonymous:1980:EBa

[Ano80a] Anonymous. Editorial board. Statistics (Berlin, DDR), 11 (1):ebi, 1980. CODEN MOSSD5. ISSN 0323-3944.

Anonymous:1980:EBb

[Ano80b] Anonymous. Editorial board. Statistics (Berlin, DDR), 11 (3):ebi, 1980. CODEN MOSSD5. ISSN 0323-3944.

Anonymous:1980:EBc

[Ano80c] Anonymous. Editorial board. Statistics (Berlin, DDR), 11 (4):ebi, 1980. CODEN MOSSD5. ISSN 0323-3944.

Anonymous:1981:EBa

[Ano81a] Anonymous. Editorial board. Statistics (Berlin, DDR), 12 (1):ebi, 1981. CODEN MOSSD5. ISSN 0323-3944.

Anonymous:1981:EBb

[Ano81b] Anonymous. Editorial board. Statistics (Berlin, DDR), 12 (2):ebi, 1981. CODEN MOSSD5. ISSN 0323-3944.

Anonymous:1981:EBc

[Ano81c] Anonymous. Editorial board. Statistics (Berlin, DDR), 12 (3):ebi, 1981. CODEN MOSSD5. ISSN 0323-3944.

Anonymous:1981:EBd

[Ano81d] Anonymous. Editorial board. Statistics (Berlin, DDR), 12 (4):ebi, 1981. CODEN MOSSD5. ISSN 0323-3944.

Anonymous:1981:SGT

[Ano81e] Anonymous. O svoictvakh granichnikh tunktsionalov ot slychainovo blyzhdanija na konechnoi tsepi Markova. (Russian) [On properties of boundary functionals from random walks on the final Markov chain]. Mathematische Operationsforschung und Statistik, 12(1):85–100, 1981. CODEN MTOSBI. ISSN 0047-6277.

Anonymous:1982:EBa

[Ano82a] Anonymous. Editorial board. Statistics (Berlin, DDR), 13

(1):ebi, 1982. CODEN MOSSD5. ISSN 0323-3944.

Anonymous:1982:EBb

[Ano82b] Anonymous. Editorial board. Statistics (Berlin, DDR), 13

(2):ebi, 1982. CODEN MOSSD5. ISSN 0323-3944.

Anonymous:1982:EBc

[Ano82c] Anonymous. Editorial board. Statistics (Berlin, DDR), 13

(3):ebi, 1982. CODEN MOSSD5. ISSN 0323-3944.

Anonymous:1982:EBd

[Ano82d] Anonymous. Editorial board. Statistics (Berlin, DDR), 13

(4):ebi, 1982. CODEN MOSSD5. ISSN 0323-3944.

Anonymous:1983:BRa

[Ano83a] Anonymous. Book reviews. Statistics (Berlin, DDR), 14(4):

534, 1983. CODEN MOSSD5. ISSN 0323-3944.

Anonymous:1983:BRb

[Ano83b] Anonymous. Book reviews. Statistics (Berlin, DDR), 14(4):

558, 1983. CODEN MOSSD5. ISSN 0323-3944.

Anonymous:1983:BRc

[Ano83c] Anonymous. Book reviews. Statistics (Berlin, DDR), 14(4):

568, 1983. CODEN MOSSD5. ISSN 0323-3944.

Anonymous:1983:BRd

[Ano83d] Anonymous. Book reviews. Statistics (Berlin, DDR), 14(4):

588, 1983. CODEN MOSSD5. ISSN 0323-3944.

Anonymous:1983:EBa

[Ano83e] Anonymous. Editorial board. Statistics (Berlin, DDR), 14 (1):ebi, 1983. CODEN MOSSD5. ISSN 0323-3944.

Anonymous:1983:EBb

[Ano83f] Anonymous. Editorial board. Statistics (Berlin, DDR), 14

(2):ebi, 1983. CODEN MOSSD5. ISSN 0323-3944.

Anonymous:1983:EBc

[Ano83g] Anonymous. Editorial board. Statistics (Berlin, DDR), 14

(4):ebi, 1983. CODEN MOSSD5. ISSN 0323-3944.

Anonymous:1984:BRe

[Ano84a] Anonymous. Book review. Statistics (Berlin, DDR), 15(3):

368, 1984. CODEN MOSSD5. ISSN 0323-3944.

Anonymous:1984:BRa

[Ano84b] Anonymous. Book reviews. Statistics (Berlin, DDR), 15(1):

143–156, 1984. CODEN MOSSD5. ISSN 0323-3944.

Anonymous:1984:BRb

[Ano84c] Anonymous. Book reviews. Statistics (Berlin, DDR), 15(2):

248, 1984. CODEN MOSSD5. ISSN 0323-3944.

Anonymous:1984:BRc

[Ano84d] Anonymous. Book reviews. Statistics (Berlin, DDR), 15(2):

276, 1984. CODEN MOSSD5. ISSN 0323-3944.

Anonymous:1984:BRd

[Ano84e] Anonymous. Book reviews. Statistics (Berlin, DDR), 15(2):

288, 1984. CODEN MOSSD5. ISSN 0323-3944.

Anonymous:1984:BRf

[Ano84f] Anonymous. Book reviews. Statistics (Berlin, DDR), 15(3):

372, 1984. CODEN MOSSD5. ISSN 0323-3944.

Anonymous:1984:BRg

[Ano84g] Anonymous. Book reviews. Statistics (Berlin, DDR), 15(3):

382, 1984. CODEN MOSSD5. ISSN 0323-3944.

Anonymous:1984:BRh

[Ano84h] Anonymous. Book reviews. Statistics (Berlin, DDR), 15(3): 388, 1984. CODEN MOSSD5. ISSN 0323-3944.

Anonymous:1984:BRi

[Ano84i] Anonymous. Book reviews. Statistics (Berlin, DDR), 15(3):

396, 1984. CODEN MOSSD5. ISSN 0323-3944.

Anonymous:1984:BRj

[Ano84j] Anonymous. Book reviews. Statistics (Berlin, DDR), 15(3):

404, 1984. CODEN MOSSD5. ISSN 0323-3944.

Anonymous:1984:BRk

[Ano84k] Anonymous. Book reviews. Statistics (Berlin, DDR), 15(3):

408, 1984. CODEN MOSSD5. ISSN 0323-3944.

Anonymous:1984:BRl

[Ano84l] Anonymous. Book reviews. Statistics (Berlin, DDR), 15(3):

436, 1984. CODEN MOSSD5. ISSN 0323-3944.

Anonymous:1984:BRm

[Ano84m] Anonymous. Book reviews. Statistics (Berlin, DDR), 15(3):

450, 1984. CODEN MOSSD5. ISSN 0323-3944.

Anonymous:1984:BRn

[Ano84n] Anonymous. Book reviews. Statistics (Berlin, DDR), 15(3):

463–468, 1984. CODEN MOSSD5. ISSN 0323-3944.

Anonymous:1984:BRo

[Ano84o] Anonymous. Book reviews. Statistics (Berlin, DDR), 15(4):

552, 1984. CODEN MOSSD5. ISSN 0323-3944.

Anonymous:1984:BRp

[Ano84p] Anonymous. Book reviews. Statistics (Berlin, DDR), 15(4):

580, 1984. CODEN MOSSD5. ISSN 0323-3944.

Anonymous:1984:BRq

[Ano84q] Anonymous. Book reviews. Statistics (Berlin, DDR), 15(4):

611–631, 1984. CODEN MOSSD5. ISSN 0323-3944.

Anonymous:1984:EBa

[Ano84r] Anonymous. Editorial board. Statistics (Berlin, DDR), 15 (1):ebi, 1984. CODEN MOSSD5. ISSN 0323-3944.

Anonymous:1984:EBb

[Ano84s] Anonymous. Editorial board. Statistics (Berlin, DDR), 15 (2):ebi, 1984. CODEN MOSSD5. ISSN 0323-3944.

Anonymous:1984:EBc

[Ano84t] Anonymous. Editorial board. Statistics (Berlin, DDR), 15 (3):ebi, 1984. CODEN MOSSD5. ISSN 0323-3944.

Anonymous:1984:EBd

[Ano84u] Anonymous. Editorial board. Statistics (Berlin, DDR), 15 (4):ebi, 1984. CODEN MOSSD5. ISSN 0323-3944.

Anonymous:1985:BRc

[Ano85a] Anonymous. Book review. Statistics: a Journal of Theoretical and Applied Statistics, 16(1):46, 1985. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Anonymous:1985:BRe

[Ano85b] Anonymous. Book review. Statistics: a Journal of Theoretical and Applied Statistics, 16(1):106, 1985. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Anonymous:1985:BRj

[Ano85c] Anonymous. Book review. Statistics: a Journal of Theoretical and Applied Statistics, 16(2):232, 1985. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Anonymous:1985:BRk

[Ano85d] Anonymous. Book review. Statistics: a Journal of Theoretical and Applied Statistics, 16(2):242, 1985. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Anonymous:1985:BRl

[Ano85e] Anonymous. Book review. Statistics: a Journal of Theoretical and Applied Statistics, 16(2):276, 1985. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Anonymous:1985:BRm

[Ano85f] Anonymous. Book review. Statistics: a Journal of Theoretical and Applied Statistics, 16(2):296, 1985. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Anonymous:1985:BRn

[Ano85g] Anonymous. Book review. Statistics: a Journal of Theoretical and Applied Statistics, 16(2):308, 1985. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Anonymous:1985:BRo

[Ano85h] Anonymous. Book review. Statistics: a Journal of Theoretical and Applied Statistics, 16(3):350, 1985. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Anonymous:1985:BRp

[Ano85i] Anonymous. Book review. Statistics: a Journal of Theoretical and Applied Statistics, 16(3):362, 1985. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Anonymous:1985:BRq

[Ano85j] Anonymous. Book review. Statistics: a Journal of Theoretical and Applied Statistics, 16(3):372, 1985. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Anonymous:1985:BRr

[Ano85k] Anonymous. Book review. Statistics: a Journal of Theoretical and Applied Statistics, 16(4):494, 1985. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Anonymous:1985:BRs

[Ano85l] Anonymous. Book review. Statistics: a Journal of Theoretical and Applied Statistics, 16(4):520, 1985. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Anonymous:1985:BRt

[Ano85m] Anonymous. Book review. Statistics: a Journal of Theoretical and Applied Statistics, 16(4):548, 1985. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Anonymous:1985:BRu

[Ano85n] Anonymous. Book review. Statistics: a Journal of Theoretical and Applied Statistics, 16(4):636, 1985. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Anonymous:1985:BRa

[Ano85o] Anonymous. Book reviews. Statistics: a Journal of Theoretical and Applied Statistics, 16(1):14, 1985. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Anonymous:1985:BRb

[Ano85p] Anonymous. Book reviews. Statistics: a Journal of Theoretical and Applied Statistics, 16(1):34, 1985. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Anonymous:1985:BRd

[Ano85q] Anonymous. Book reviews. Statistics: a Journal of Theoretical and Applied Statistics, 16(1):96, 1985. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Anonymous:1985:BRf

[Ano85r] Anonymous. Book reviews. Statistics: a Journal of Theoretical and Applied Statistics, 16(1):120, 1985. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Anonymous:1985:BRg

[Ano85s] Anonymous. Book reviews. Statistics: a Journal of Theoretical and Applied Statistics, 16(1):138, 1985. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Anonymous:1985:BRh

[Ano85t] Anonymous. Book reviews. Statistics: a Journal of Theoretical and Applied Statistics, 16(2):202, 1985. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Anonymous:1985:BRi

[Ano85u] Anonymous. Book reviews. Statistics: a Journal of Theoretical and Applied Statistics, 16(2):212, 1985. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Anonymous:1985:EBa

[Ano85v] Anonymous. Editorial board. Statistics: a Journal of Theoretical and Applied Statistics, 16(1):ebi, 1985. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Anonymous:1985:EBb

[Ano85w] Anonymous. Editorial board. Statistics: a Journal of Theoretical and Applied Statistics, 16(2):ebi, 1985. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Anonymous:1985:EBc

[Ano85x] Anonymous. Editorial board. Statistics: a Journal of Theoretical and Applied Statistics, 16(3):ebi, 1985. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Anonymous:1985:EBd

[Ano85y] Anonymous. Editorial board. Statistics: a Journal of Theoretical and Applied Statistics, 16(4):ebi, 1985. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Anonymous:1986:BRf

[Ano86a] Anonymous. Book review. Statistics: a Journal of Theoretical and Applied Statistics, 17(3):336, 1986. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Anonymous:1986:BRg

[Ano86b] Anonymous. Book review. Statistics: a Journal of Theoretical and Applied Statistics, 17(3):356, 1986. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Anonymous:1986:BRh

[Ano86c] Anonymous. Book review. Statistics: a Journal of Theoretical and Applied Statistics, 17(3):392, 1986. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Anonymous:1986:BRi

[Ano86d] Anonymous. Book review. Statistics: a Journal of Theoretical and Applied Statistics, 17(3):406, 1986. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Anonymous:1986:BRj

[Ano86e] Anonymous. Book review. Statistics: a Journal of Theoretical and Applied Statistics, 17(3):428, 1986. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Anonymous:1986:BRk

[Ano86f] Anonymous. Book review. Statistics: a Journal of Theoretical and Applied Statistics, 17(3):434, 1986. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Anonymous:1986:BRl

[Ano86g] Anonymous. Book review. Statistics: a Journal of Theoretical and Applied Statistics, 17(3):461–480, 1986. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Anonymous:1986:BRm

[Ano86h] Anonymous. Book review. Statistics: a Journal of Theoretical and Applied Statistics, 17(4):532, 1986. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Anonymous:1986:BRn

[Ano86i] Anonymous. Book review. Statistics: a Journal of Theoretical and Applied Statistics, 17(4):548, 1986. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Anonymous:1986:BRo

[Ano86j] Anonymous. Book review. Statistics: a Journal of Theoretical and Applied Statistics, 17(4):576, 1986. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Anonymous:1986:BRp

[Ano86k] Anonymous. Book review. Statistics: a Journal of Theoretical and Applied Statistics, 17(4):616–632, 1986. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Anonymous:1986:BRa

[Ano86l] Anonymous. Book reviews. Statistics: a Journal of Theoretical and Applied Statistics, 17(2):188, 1986. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Anonymous:1986:BRb

[Ano86m] Anonymous. Book reviews. Statistics: a Journal of Theoretical and Applied Statistics, 17(2):200, 1986. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Anonymous:1986:BRc

[Ano86n] Anonymous. Book reviews. Statistics: a Journal of Theoretical and Applied Statistics, 17(2):220, 1986. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Anonymous:1986:BRd

[Ano86o] Anonymous. Book reviews. Statistics: a Journal of Theoretical and Applied Statistics, 17(2):248, 1986. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Anonymous:1986:BRe

[Ano86p] Anonymous. Book reviews. Statistics: a Journal of Theoretical and Applied Statistics, 17(2):290, 1986. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Anonymous:1986:EBa

[Ano86q] Anonymous. Editorial board. Statistics: a Journal of Theoretical and Applied Statistics, 17(1):ebi, 1986. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Anonymous:1986:EBb

[Ano86r] Anonymous. Editorial board. Statistics: a Journal of Theoretical and Applied Statistics, 17(2):ebi, 1986. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Anonymous:1986:EBc

[Ano86s] Anonymous. Editorial board. Statistics: a Journal of Theoretical and Applied Statistics, 17(3):ebi, 1986. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Anonymous:1986:EBd

[Ano86t] Anonymous. Editorial board. Statistics: a Journal of Theoretical and Applied Statistics, 17(4):ebi, 1986. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Anonymous:1986:S

[Ano86u] Anonymous. Statistics. Statistics: a Journal of Theoretical and Applied Statistics, 17(2):179–178, 1986. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Anonymous:1987:BRa

[Ano87a] Anonymous. Book review. Statistics: a Journal of Theoretical and Applied Statistics, 18(4):526, 1987. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Anonymous:1987:BRb

[Ano87b] Anonymous. Book review. Statistics: a Journal of Theoretical and Applied Statistics, 18(4):572, 1987. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Anonymous:1987:BRc

[Ano87c] Anonymous. Book review. Statistics: a Journal of Theoretical and Applied Statistics, 18(4):586, 1987. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Anonymous:1987:BRd

[Ano87d] Anonymous. Book review. Statistics: a Journal of Theoretical and Applied Statistics, 18(4):598, 1987. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Anonymous:1987:BRe

[Ano87e] Anonymous. Book review. Statistics: a Journal of Theoretical and Applied Statistics, 18(4):622, 1987. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Anonymous:1987:EBa

[Ano87f] Anonymous. Editorial board. Statistics: a Journal of Theoretical and Applied Statistics, 18(1):ebi, 1987. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Anonymous:1987:EBb

[Ano87g] Anonymous. Editorial board. Statistics: a Journal of Theoretical and Applied Statistics, 18(2):ebi, 1987. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Anonymous:1987:EBc

[Ano87h] Anonymous. Editorial board. Statistics: a Journal of Theoretical and Applied Statistics, 18(3):ebi, 1987. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Anonymous:1987:EBd

[Ano87i] Anonymous. Editorial board. Statistics: a Journal of Theoretical and Applied Statistics, 18(4):ebi, 1987. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Anonymous:1987:OOO

[Ano87j] Anonymous. Ob odnoi otsenke ostatochnovo chlena v lokalnoi predelnoi teoremye dls vsveshennikh cym nezavicimikh slychainikh velichin. (Russian) [On one estimate of the residual term in the local limit theorem for the theory of illuminated independent random variables]. Mathematische Operationsforschung und Statistik, 18(3):415–422, 1987. CODEN MTOSBI. ISSN 0047-6277.

Anonymous:1988:BRa

[Ano88a] Anonymous. Book review. Statistics: a Journal of Theoretical and Applied Statistics, 19(3):348, 1988. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Anonymous:1988:BRb

[Ano88b] Anonymous. Book review. Statistics: a Journal of Theoretical and Applied Statistics, 19(3):476, 1988. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Anonymous:1988:BRc

[Ano88c] Anonymous. Book review. Statistics: a Journal of Theoretical and Applied Statistics, 19(4):502, 1988. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Anonymous:1988:EB

[Ano88d] Anonymous. Editorial board. Statistics: a Journal of Theoretical and Applied Statistics, 19(1):ebi, 1988. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Anonymous:1988:S

[Ano88e] Anonymous. Staistics. Statistics: a Journal of Theoretical and Applied Statistics, 19(2):172, 1988. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Anonymous:1989:BRa

[Ano89a] Anonymous. Book review. Statistics: a Journal of Theoretical and Applied Statistics, 20(2):222, 1989. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Anonymous:1989:BRb

[Ano89b] Anonymous. Book review. Statistics: a Journal of Theoretical and Applied Statistics, 20(2):234, 1989. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Anonymous:1989:BRc

[Ano89c] Anonymous. Book review. Statistics: a Journal of Theoretical and Applied Statistics, 20(2):296, 1989. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Anonymous:1989:BRe

[Ano89d] Anonymous. Book review. Statistics: a Journal of Theoretical and Applied Statistics, 20(3):352, 1989. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Anonymous:1989:BRf

[Ano89e] Anonymous. Book review. Statistics: a Journal of Theoretical and Applied Statistics, 20(3):382, 1989. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Anonymous:1989:BRg

[Ano89f] Anonymous. Book review. Statistics: a Journal of Theoretical and Applied Statistics, 20(3):396, 1989. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Anonymous:1989:BRh

[Ano89g] Anonymous. Book review. Statistics: a Journal of Theoretical and Applied Statistics, 20(3):406, 1989. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Anonymous:1989:BRd

[Ano89h] Anonymous. Book reviews. Statistics: a Journal of Theoretical and Applied Statistics, 20(2):331–336, 1989. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Anonymous:1989:S

[Ano89i] Anonymous. Statistics. Statistics: a Journal of Theoretical and Applied Statistics, 20(1):40, 1989. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Anonymous:1990:BRa

[Ano90a] Anonymous. Book review. Statistics: a Journal of Theoretical and Applied Statistics, 21(1):20, 1990. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Anonymous:1990:BRb

[Ano90b] Anonymous. Book review. Statistics: a Journal of Theoretical and Applied Statistics, 21(4):520, 1990. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Anonymous:1990:S

[Ano90c] Anonymous. Statistics. Statistics: a Journal of Theoretical and Applied Statistics, 21(2):238, 1990. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Anonymous:1991:BRa

[Ano91a] Anonymous. Book review. Statistics: a Journal of Theoretical and Applied Statistics, 22(2):190, 1991. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Anonymous:1991:BRb

[Ano91b] Anonymous. Book review. Statistics: a Journal of Theoretical and Applied Statistics, 22(4):650, 1991. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Anonymous:1991:BRc

[Ano91c] Anonymous. Book reviews. Statistics: a Journal of Theoretical and Applied Statistics, 22(4):655–656, 1991. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Anonymous:1991:S

[Ano91d] Anonymous. Statistics. Statistics: a Journal of Theoretical and Applied Statistics, 22(1):16, 1991. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Anonymous:1992:EBa

[Ano92a] Anonymous. Editorial board. Statistics: a Journal of Theoretical and Applied Statistics, 23(1):ebi, 1992. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Anonymous:1992:EBb

[Ano92b] Anonymous. Editorial board. Statistics: a Journal of Theoretical and Applied Statistics, 23(2):ebi, 1992. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Anonymous:1992:EBc

[Ano92c] Anonymous. Editorial board. Statistics: a Journal of Theoretical and Applied Statistics, 23(3):ebi, 1992. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Anonymous:1992:EBd

[Ano92d] Anonymous. Editorial board. Statistics: a Journal of Theoretical and Applied Statistics, 23(4):ebi, 1992. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Anonymous:1992:R

[Ano92e] Anonymous. Rejoinder. Statistics: a Journal of Theoretical and Applied Statistics, 23(3):214–216, 1992. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Anonymous:1993:BRa

[Ano93a] Anonymous. Book reviews. Statistics: a Journal of Theoretical and Applied Statistics, 24(2):175–179, 1993. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Anonymous:1993:BRb

[Ano93b] Anonymous. Book reviews. Statistics: a Journal of Theoretical and Applied Statistics, 24(3):279–287, 1993. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Anonymous:1993:BRc

[Ano93c] Anonymous. Book reviews. Statistics: a Journal of Theoretical and Applied Statistics, 24(3):289–302, 1993. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Anonymous:1993:EBa

[Ano93d] Anonymous. Editorial board. Statistics: a Journal of Theoretical and Applied Statistics, 24(1):ebi, 1993. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Anonymous:1993:EBb

[Ano93e] Anonymous. Editorial board. Statistics: a Journal of Theoretical and Applied Statistics, 24(2):ebi, 1993. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Anonymous:1993:EBc

[Ano93f] Anonymous. Editorial board. Statistics: a Journal of Theoretical and Applied Statistics, 24(3):ebi, 1993. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Anonymous:1993:EBd

[Ano93g] Anonymous. Editorial board. Statistics: a Journal of Theoretical and Applied Statistics, 24(4):ebi, 1993. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Anonymous:1993:EBe

[Ano93h] Anonymous. Editorial board. Statistics: a Journal of Theoretical and Applied Statistics, 25(1):ebi, 1993. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Anonymous:1994:EBa

[Ano94a] Anonymous. Editorial board. Statistics: a Journal of Theoretical and Applied Statistics, 25(2):ebi, 1994. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Anonymous:1994:EBb

[Ano94b] Anonymous. Editorial board. Statistics: a Journal of Theoretical and Applied Statistics, 25(3):ebi, 1994. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Anonymous:1994:EBc

[Ano94c] Anonymous. Editorial board. Statistics: a Journal of Theoretical and Applied Statistics, 25(4):ebi, 1994. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Anonymous:1995:EBa

[Ano95a] Anonymous. Editorial board. Statistics: a Journal of Theoretical and Applied Statistics, 26(1):ebi, 1995. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Anonymous:1995:EBb

[Ano95b] Anonymous. Editorial board. Statistics: a Journal of Theoretical and Applied Statistics, 26(2):ebi, 1995. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Anonymous:1995:EBc

[Ano95c] Anonymous. Editorial board. Statistics: a Journal of Theoretical and Applied Statistics, 26(3):ebi, 1995. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Anonymous:1995:EBd

[Ano95d] Anonymous. Editorial board. Statistics: a Journal of Theoretical and Applied Statistics, 26(4):ebi, 1995. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Anonymous:1995:EBe

[Ano95e] Anonymous. Editorial board. Statistics: a Journal of Theoretical and Applied Statistics, 27(1–2):ebi, 1995. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Anonymous:1996:EBa

[Ano96a] Anonymous. Editorial board. Statistics: a Journal of Theoretical and Applied Statistics, 27(3–4):ebi, 1996. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Anonymous:1996:EBb

[Ano96b] Anonymous. Editorial board. Statistics: a Journal of Theoretical and Applied Statistics, 28(1):ebi, 1996. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Anonymous:1996:EBc

[Ano96c] Anonymous. Editorial board. Statistics: a Journal of Theoretical and Applied Statistics, 28(2):ebi, 1996. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Anonymous:1996:EBd

[Ano96d] Anonymous. Editorial board. Statistics: a Journal of Theoretical and Applied Statistics, 28(3):ebi, 1996. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Anonymous:1996:EBe

[Ano96e] Anonymous. Editorial board. Statistics: a Journal of Theoretical and Applied Statistics, 28(4):ebi, 1996. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Anonymous:1997:EBa

[Ano97a] Anonymous. Editorial board. Statistics: a Journal of Theoretical and Applied Statistics, 29(1):ebi, 1997. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Anonymous:1997:EBb

[Ano97b] Anonymous. Editorial board. Statistics: a Journal of Theoretical and Applied Statistics, 29(2):ebi, 1997. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Anonymous:1997:EBc

[Ano97c] Anonymous. Editorial board. Statistics: a Journal of Theoretical and Applied Statistics, 29(3):ebi, 1997. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Anonymous:1997:EBd

[Ano97d] Anonymous. Editorial board. Statistics: a Journal of Theoretical and Applied Statistics, 29(4):ebi, 1997. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Anonymous:1997:EBe

[Ano97e] Anonymous. Editorial board. Statistics: a Journal of Theoretical and Applied Statistics, 30(1):ebi, 1997. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Anonymous:1997:EBf

[Ano97f] Anonymous. Editorial board. Statistics: a Journal of Theoretical and Applied Statistics, 30(2):ebi, 1997. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Anonymous:1997:EBg

[Ano97g] Anonymous. Editorial board. Statistics: a Journal of Theoretical and Applied Statistics, 30(3):ebi, 1997. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Anonymous:1998:EBa

[Ano98a] Anonymous. Editorial board. Statistics: a Journal of Theoretical and Applied Statistics, 30(4):ebi, 1998. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Anonymous:1998:EBb

[Ano98b] Anonymous. Editorial board. Statistics: a Journal of Theoretical and Applied Statistics, 31(1):ebi, 1998. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Anonymous:1998:EBc

[Ano98c] Anonymous. Editorial board. Statistics: a Journal of Theoretical and Applied Statistics, 31(2):ebi, 1998. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Anonymous:1998:EBd

[Ano98d] Anonymous. Editorial board. Statistics: a Journal of Theoretical and Applied Statistics, 31(3):ebi, 1998. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Anonymous:1998:EBe

[Ano98e] Anonymous. Editorial board. Statistics: a Journal of Theoretical and Applied Statistics, 31(4):ebi, 1998. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Anonymous:1998:EBf

[Ano98f] Anonymous. Editorial board. Statistics: a Journal of Theoretical and Applied Statistics, 32(1):ebi, 1998. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Anonymous:1998:EBg

[Ano98g] Anonymous. Editorial board. Statistics: a Journal of Theoretical and Applied Statistics, 32(2):ebi, 1998. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Anonymous:1999:EBa

[Ano99a] Anonymous. Editorial board. Statistics: a Journal of Theoretical and Applied Statistics, 32(3):ebi, 1999. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Anonymous:1999:EBb

[Ano99b] Anonymous. Editorial board. Statistics: a Journal of Theoretical and Applied Statistics, 32(4):ebi, 1999. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Anonymous:1999:EBc

[Ano99c] Anonymous. Editorial board. Statistics: a Journal of Theoretical and Applied Statistics, 33(1):ebi, 1999. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Anonymous:1999:EBd

[Ano99d] Anonymous. Editorial board. Statistics: a Journal of Theoretical and Applied Statistics, 33(2):ebi, 1999. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Anonymous:1999:EBe

[Ano99e] Anonymous. Editorial board. Statistics: a Journal of Theoretical and Applied Statistics, 33(3):ebi, 1999. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Anonymous:2000:EBa

[Ano00a] Anonymous. Editorial board. Statistics: a Journal of Theoretical and Applied Statistics, 33(4):ebi, 2000. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Anonymous:2000:EBb

[Ano00b] Anonymous. Editorial board. Statistics: a Journal of Theoretical and Applied Statistics, 34(1):ebi, 2000. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Anonymous:2000:EBc

[Ano00c] Anonymous. Editorial board. Statistics: a Journal of Theoretical and Applied Statistics, 34(2):ebi, 2000. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Anonymous:2000:EBd

[Ano00d] Anonymous. Editorial board. Statistics: a Journal of Theoretical and Applied Statistics, 34(3):ebi, 2000. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Anonymous:2000:EBe

[Ano00e] Anonymous. Editorial board. Statistics: a Journal of Theoretical and Applied Statistics, 34(4):ebi, 2000. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Anonymous:2010:EB

[Ano10] Anonymous. Editorial board. Statistics: a Journal of Theoretical and Applied Statistics, 44(6):ebi, 2010. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Anonymous:2011:EB

[Ano11] Anonymous. Editorial board. Statistics: a Journal of Theoretical and Applied Statistics, 45(6):ebi, 2011. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Anonymous:2012:EB

[Ano12a] Anonymous. Editorial board. Statistics: a Journal of Theoretical and Applied Statistics, 46(6):ebi, 2012. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Anonymous:2012:SR

[Ano12b] Anonymous. Statement of retraction. Statistics: a Journal of Theoretical and Applied Statistics, 46(4):563, 2012. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Anonymous:2014:E

[Ano14] Anonymous. Erratum. Statistics: a Journal of Theoretical and Applied Statistics, 48(3):710, 2014. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Anonymous:2015:C

[Ano15a] Anonymous. Corrigendum. Statistics: a Journal of Theoretical and Applied Statistics, 49(6):1422, 2015. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Anonymous:2015:EB

[Ano15b] Anonymous. Editorial board. Statistics: a Journal of Theoretical and Applied Statistics, 49(6):ebi, 2015. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Anonymous:2015:E

[Ano15c] Anonymous. Erratum. Statistics: a Journal of Theoretical and Applied Statistics, 49(1):i, 2015. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Anonymous:2016:C

[Ano16a] Anonymous. Corrigendum. Statistics: a Journal of Theoretical and Applied Statistics, 50(3):i, 2016. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Anonymous:2016:EB

[Ano16b] Anonymous. Editorial board. Statistics: a Journal of Theoretical and Applied Statistics, 50(6):ebi, 2016. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Anonymous:2017:EB

[Ano17] Anonymous. Editorial board. Statistics: a Journal of Theoretical and Applied Statistics, 51(6):ebi, 2017. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Anonymous:2019:CPS

[Ano19] Anonymous. Corrigenda to “Proportional subclass numbers in two-factor ANOVA”. Statistics: a Journal of Theoretical and Applied Statistics, 53(3):696–698, 2019. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic). See [LaM17].

Andel:2007:MCS

[ANR07] J. Andel, I. Netuka, and P. Ranocha. Methods for calculating stationary distribution in linear models of time series. Statistics: a Journal of Theoretical and Applied Statistics, 41(4): 279–287, 2007. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Antoniadis:1984:AVF

[Ant84a] A. Antoniadis. Analysis of variance on function spaces. Statistics (Berlin, DDR), 15(1):59–71, 1984. CODEN MOSSD5. ISSN 0323-3944.

Antoniadis:1984:SBA

[Ant84b] A. Antoniadis. A survey of Bayesian approaches to nonparametric statisties. Statistics (Berlin, DDR), 15(1):121–142, 1984. CODEN MOSSD5. ISSN 0323-3944.

Akahira:2007:BVH

[AO07] M. Akahira and N. Ohyauchi. A Bayesian view of the Hammersley–Chapman–Robbins-type inequality. Statistics: a Journal of Theoretical and Applied Statistics, 41(2):137– 144, 2007. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Arslan:2023:RPE

[AÖ23] Olcay Arslan and Senay Özdemir. Robust penalized empirical likelihood estimation method for linear regression. Statistics: a Journal of Theoretical and Applied Statistics, 57(2):423– 443, 2023. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-

4910 (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/02331888.2023.2179054.

Anh:2019:LRM

[AOV19] Vo Anh, Andriy Olenko, and Volodymyr Vaskovych. On LSE in regression model for long-range dependent random fields on spheres. Statistics: a Journal of Theoretical and Applied Statistics, 53(5):1131–1151, 2019. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Atchade:2007:GEM

[AP07] Yves F. Atchadé and François Perron. On the geometric ergodicity of Metropolis–Hastings algorithms. Statistics: a Journal of Theoretical and Applied Statistics, 41(1):77–84, 2007. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Ahmadi:2013:COS

[AR13] Jafar Ahmadi and Mohammad Z. Raqab. Comparison of order statistics in two-sample problem in the sense of Pitman closeness. Statistics: a Journal of Theoretical and Applied Statistics, 47(4):729–743, 2013. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Aragon:1993:GFT

[Ara93] Yves Aragon. Goodness of fit tests based on frequencies and averages of classes. Statistics: a Journal of Theoretical and Applied Statistics, 24(2):117–135, 1993. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Arendacka:2005:GCI

[Are05] Barbora Arendacká. Generalized confidence intervals on the variance component in mixed linear models with two variance components. Statistics: a Journal of Theoretical and Applied Statistics, 39(4):275–286, 2005. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Arendacka:2012:NFG

[Are12] Barbora Arendacká. A note on fiducial generalized pivots for in one-way heteroscedastic ANOVA with random effects. Statistics: a Journal of Theoretical and Applied Statistics, 46(4):489–504, 2012. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Arnab:1991:AOS

[Arn91] Raghunath Arnab. On asympotically optimal samplig strategies under correlated super-population modelling. Statistics: a Journal of Theoretical and Applied Statistics, 22(2):295– 298, 1991. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Arnab:1992:UEV

[Arn92a] Raghunath Arnab. Unbiased estimation of variance in multistage sampling under randomized response surveys. Statistics: a Journal of Theoretical and Applied Statistics, 23(4):357– 364, 1992. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Arnold:1992:D

[Arn92b] Barry C. Arnold. Discussion. Statistics: a Journal of Theoretical and Applied Statistics, 23(3):209–211, 1992. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Arnold:1992:IPP

[Arn92c] Bernhard F. Arnold. The inspection points of a production line. Statistics: a Journal of Theoretical and Applied Statistics, 23(4):347–356, 1992. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Arnab:1995:OEF

[Arn95] Raghunath Arnab. Optimal estimation of a finite population total under randomized response surveys. Statistics: a Journal of Theoretical and Applied Statistics, 27(1–2):175– 180, 1995. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Arashi:2012:NST

[ARN12] M. Arashi, M. Roozbeh, and H. A. Niroumand. A note on Stein-type shrinkage estimator in partial linear models. Statistics: a Journal of Theoretical and Applied Statistics, 46(5): 673–685, 2012. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Arslan:2016:PMR

[Ars16] Olcay Arslan. Penalized MM regression estimation with L_{γ} penalty: a robust version of bridge regression. Statistics: a Journal of Theoretical and Applied Statistics, 50(6):1236– 1260, 2016. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Artemiou:2019:UAW

[Art19] Andreas Artemiou. Using adaptively weighted large margin classifiers for robust sufficient dimension reduction. Statistics: a Journal of Theoretical and Applied Statistics, 53(5):1037– 1051, 2019. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Alalouf:1979:ETR

[AS79] I. S. Alalouf and G. P. H. Styan. Estimability and testability in restricted linear models. Statistics (Berlin, DDR), 10(2): 189–201, 1979. CODEN MOSSD5. ISSN 0323-3944.

Ashour:1983:ESS

[AS83] S. K. Ashour and O. A. Shalaby. Estimating sample size with Weibull failure. Statistics (Berlin, DDR), 14(2):263–268, 1983. CODEN MOSSD5. ISSN 0323-3944.

Arnold:1998:LAE

[AS98] Bernhard F. Arnold and Peter Stahlecker. Linear affine estimation in misspecified linear regression models using fuzzy prior information. Statistics: a Journal of Theoretical and Applied Statistics, 32(1):1–12, 1998. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Arnold:2004:SPR

[AS04] Bernhard F. Arnold and Peter Stahlecker. Some properties of the relative squared error approach to linear regression analysis. Statistics: a Journal of Theoretical and Applied Statistics, 38(3):183–193, 2004. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Ait-Saidi:2008:CVE

[ASFKV08] Ahmed Ait-Saïdi, Frédéric Ferraty, Rabah Kassa, and Philippe Vieu. Cross-validated estimations in the singlefunctional index model. Statistics: a Journal of Theoretical and Applied Statistics, 42(6):475–494, 2008. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Amini-Seresht:2016:OSM

[ASKS16] Ebrahim Amini-Seresht, Baha-Eldin Khaledi, and Moshe Shaked. Order statistics with multivariate concomitants: stochastic comparisons. Statistics: a Journal of Theoretical and Applied Statistics, 50(1):190–205, 2016. CO-DEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/02331888.2015.1022550.

Ashour:1983:BEC

[ASM83] S. K. Ashour, E. Shoukry, and T. Mohamed. Bayesian estimation for compound Weibull model. Statistics (Berlin, DDR), 14(3):381–386, 1983. CODEN MOSSD5. ISSN 0323-3944.

Aggarwal:2013:OOB

[ASSH13] M. L. Aggarwal, Poonam Singh, Vandana Sarin, and Bushra Husain. Optimal orthogonal block designs for four-mixture components in two blocks based on F -squares for Becker’s models and K-model. Statistics: a Journal of Theoretical and Applied Statistics, 47(5):1003–1021, 2013. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Arashi:2013:RME

[AST13] M. Arashi, A. K. Md. E. Saleh, and S. M. M. Tabatabaey. Regression model with elliptically contoured errors. Statistics: a Journal of Theoretical and Applied Statistics, 47(6):1266– 1284, 2013. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Akkaya:2005:TSA

[AT05] Aysen D. Akkaya and Moti L. Tiku. Time series AR(1) model for short-tailed distributions. Statistics: a Journal of Theoretical and Applied Statistics, 39(2):117–132, 2005. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Akkaya:2008:AMS

[AT08] Aysen D. Akkaya and Moti L. Tiku. Autoregressive models with short-tailed symmetric distributions. Statistics: a Journal of Theoretical and Applied Statistics, 42(3):207–221, 2008. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Abtahi:2013:NUR

[AT13] Asieh Abtahi and Mina Towhidi. The new unified representation of multivariate skewed distributions. Statistics: a Journal of Theoretical and Applied Statistics, 47(1):126–140, 2013. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Atkinson:1975:VES

[Atk75] A. C. Atkinson. Vorträge auf der ersten Sommerschule über Probleme der Modellwahl und Parameterschätzung in der Regressions-Analyse, April 1974 in Zinnowitz/Usedom, DDR. (German) [Lectures at the first Summer School on Problems of Model Choice and Parameter Estimation in Regression Analysis, April 1974 in Zinnowitz slash Usedom, DDR]. Mathematische Operationsforschung und Statistik, 6(2):253– 267, 1975. CODEN MTOSBI. ISSN 0047-6277.

Atkinson:1995:D

[Atk95] A. C. Atkinson. Discussion. Statistics: a Journal of Theoretical and Applied Statistics, 26(2):149–150, 1995. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Auinger:1990:AEV

[Aui90] Kabl Auinger. Ein Anpassungstest für Exponentialverteilte Variable. (German) [A goodness of fit test for exponentiallydistributed variables]. Statistics: a Journal of Theoretical and Applied Statistics, 21(1):75–90, 1990. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Arnold:2012:GOS

[AV12] Barry C. Arnold and Jose A. Villasenor. Generalized order statistic processes and Pfeifer records. Statistics: a Journal of Theoretical and Applied Statistics, 46(3):373–385, 2012. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Arellano-Valle:1998:STS

[AVB98] Reinaldo B. Arellano-Valle and Heleno Bolfarine. On score tests in structural regression models. Statistics: a Journal of Theoretical and Applied Statistics, 32(2):131–149, 1998. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Arnold:1989:RC

[AvC89] B. F. Arnold and E. v. Collani. On the robustness of X̄-charts. Statistics: a Journal of Theoretical and Applied Statistics, 20(1):149–159, 1989. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Aleksandrov:2022:NGF

[AWJF22] Boris Aleksandrov, Christian H. Weiß, Carsten Jentsch, and Maxime Faymonville. Novel goodness-of-fit tests for binomial count time series. Statistics: a Journal of Theoretical and Applied Statistics, 56(5):957–990, 2022. CO-DEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/02331888.2022.2134384.

Adragni:2016:PSD

[AX16] Kofi P. Adragni and Mingyu Xi. Pruning a sufficient dimension reduction with a p-value guided hard-thresholding. Statistics: a Journal of Theoretical and Applied Statistics, 50(2):254–270, 2016. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic). URL http://

www.tandfonline.com/doi/abs/10.1080/02331888.2015.
1050019.

Azpeitla:1973:DSF

[Azp73] A. G. Azpeitla. Decomposition of switching functions into linearly separable components. Mathematische Operationsforschung und Statistik, 4(3):179–182, 1973. CODEN MTOSBI. ISSN 0047-6277.

Beder:1989:JEM

[BA89] J. H. Beder and A. Antoniadis. Joint estimation of the mean and the covariance of a Banach valued Gaussian vector. Statistics: a Journal of Theoretical and Applied Statistics, 20(1): 77–93, 1989. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Baklizi:2008:ERF

[BA08] Ayman Baklizi and S. E. Ahmed. On the estimation of reliabilty function in a Weibull lifetime distribution. Statistics: a Journal of Theoretical and Applied Statistics, 42(4):351–362, 2008. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Bagiatis:1990:OSD

[Bag90] C. Bagiatis. E-optimal 2^{k} saturated designs. Statistics: a Journal of Theoretical and Applied Statistics, 21(1):35–44, 1990. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Bako:1973:DSP

[Bak73] A. Bakó. On the determination of the shortest path in a network having gains. Mathematische Operationsforschung und Statistik, 4(1):63–68, 1973. CODEN MTOSBI. ISSN 0047-6277.

Bako:1976:APA

[Bak76] A. Bakó. All paths in an activity network. Mathematische Operationsforschung und Statistik, 7(6):851–858, 1976. CODEN MTOSBI. ISSN 0047-6277.

Baksalary:1984:SEB

[Bak84] Jerzy K. Baksalary. A study of the equivalence between a Gauss–Markoff model and its augmentation by nuisance parameters. Statistics (Berlin, DDR), 15(1):3–35, 1984. CO-DEN MOSSD5. ISSN 0323-3944.

Balasubramanian:2003:SLI

[Bal03] K. Balasubramanian. On special linear identities for order statistics. Statistics: a Journal of Theoretical and Applied Statistics, 37(4):335–339, 2003. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Bindele:2018:RCR

[BAM18] Huybrechts F. Bindele, Asheber Abebe, and Nicole K. Meyer. Robust confidence regions for the semi-parametric regression model with responses missing at random. Statistics: a Journal of Theoretical and Applied Statistics, 52(4):885–900, 2018. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Bandemer:1970:OWA

[Ban70a] Hans Bandemer. Optimale Wahl der Abszissenwerte für die lineare Regression bei gegebener Kostenfunktion. (German) [Optimum choice of the abscissa values for the linear regression in a given cost function]. Mathematische Operationsforschung und Statistik, 1(4):297–308, 1970. CODEN MTOSBI. ISSN 0047-6277.

Bank:1970:BBA

[Ban70b] Bernd Bank. Branch and Bound-Algorithmen für zwei Reihenfolgeprobleme. (German) [Branch and bound algorithms for two sequence problems]. Mathematische Operationsforschung und Statistik, 1(3):217–228, 1970. CODEN MTOSBI. ISSN 0047-6277.

Bandemer:1980:PFU

[Ban80] H. Bandemer. Problems in foundation and use of optimal experimental design in regression models. Statistics (Berlin, DDR), 11(1):89–113, 1980. CODEN MOSSD5. ISSN 0323-3944.

Bouza:2012:EPM

[BAO12] Carlos N. Bouza and Amer Ibrahim Al-Omari. Estimating the population mean in the case of missing data using simple random sampling. Statistics: a Journal of Theoretical and Applied Statistics, 46(2):279–290, 2012. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Bar:1972:VLG

[Bär72] Gunter Bär. Verfahren zur Lösung ganzzahliger linearer Optimierungsprobleme. (German) [Method for solving integer linear programming problems]. Mathematische Operationsforschung und Statistik, 3(1):3–17, 1972. CODEN MTOSBI. ISSN 0047-6277.

Bar:1976:LBO

[Bär76] Gunter Bär. Zur linearisierung beliebiger 0–1-Optimierungsprobleme. (German) [On linearization of generic 0–1 optimization problems]. Mathematische Operationsforschung und Statistik, 7 (2):181–195, 1976. CODEN MTOSBI. ISSN 0047-6277.

Barbu:1987:NFM

[Bar87] Gh. Barbu. A new fast method for computer generation of gamma and beta random variables by transformations of uniform variables. Statistics: a Journal of Theoretical and Applied Statistics, 18(3):453–464, 1987. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Barbe:1994:DFS

[Bar94] Philippe Barbe. On the distribution function for the spacings. Statistics: a Journal of Theoretical and Applied Statistics, 25(4):367–373, 1994. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Barakat:2010:CLW

[Bar10] H. M. Barakat. Continuability of local weak limits for record values. Statistics: a Journal of Theoretical and Applied Statistics, 44(3):269–274, 2010. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Batsidis:2010:RLR

[Bat10] Apostolos Batsidis. Robustness of the likelihood ratio test for detection and estimation of a mean change point in a sequence of elliptically contoured observations. Statistics: a Journal of Theoretical and Applied Statistics, 44(1):17–24, 2010. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Bunke:1972:CBG

[BB72] H. Bunke and O. Bunke. Confidence bands with a given maximal size for linear regression functions. Mathematische Operationsforschung und Statistik, 3(4):273–280, 1972. CODEN MTOSBI. ISSN 0047-6277.

Bunke:1974:IE

[BB74] H. Bunke and O. Bunke. Identifiability and estimability. Mathematische Operationsforschung und Statistik, 5(3):223– 233, 1974. CODEN MTOSBI. ISSN 0047-6277. See note [Sch76a].

Bunke:1978:RPC

[BB78] Helga Bunke and Olaf Bunke. Randomization. pro and contra. Statistics (Berlin, DDR), 9(4):607–623, 1978. CODEN MOSSD5. ISSN 0323-3944.

Bunke:1982:BRb

[BB82] O. Bunke and H. Bandemer. Book review. Statistics (Berlin, DDR), 13(3):444, 1982. CODEN MOSSD5. ISSN 0323-3944.

Benzekri:1991:ARP

[BB91] S. Benzekri and F. Brodeau. Asymptotic results for parametric estimation in inadequate two phases regression models. Statistics: a Journal of Theoretical and Applied Statistics, 22(3):331–348, 1991. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Balakrishnan:1993:IBE

[BB93] N. Balakrishnan and S. M. Bendre. Improved bounds for expectations of linear functions of order statistics. Statistics: a Journal of Theoretical and Applied Statistics, 24(2):161– 165, 1993. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Bogdan:2000:EML

[BB00] Krzysztof Bogdan and Małgorzata Bogdan. On existence of maximum likelihood estimators in exponential families. Statistics: a Journal of Theoretical and Applied Statistics, 34(2):137–149, 2000. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Balasubramanian:2004:IOS

[BB04] K. Balasubramanian and M. I. Beg. Identities in order statistics — from moments to arbitrary distributions. Statistics: a Journal of Theoretical and Applied Statistics, 38(1):43–48, 2004. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Bandyopadhyay:2007:EOA

[BB07] Uttam Bandyopadhyay and Rahul Bhattacharya. On an ethical cum optimal adaptive allocation design. Statistics: a Journal of Theoretical and Applied Statistics, 41(6):471–483, 2007. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Bandyopadhyay:2013:SSO

[BB13] Subhadip Bandyopadhyay and Shibdas Bandyopadhyay. Sampling strategy for optimal classification into one of two correlated normal populations. Statistics: a Journal of Theoretical and Applied Statistics, 47(5):1116–1127, 2013. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Balakrishnan:2022:SBL

[BB22a] Narayanaswamy Balakrishnan and Ritwik Bhattacharya. Simultaneous best linear invariant prediction of future order statistics for location-scale and scale families and associated optimality properties. Statistics: a Journal of Theoretical and Applied Statistics, 56(6):1409–1426, 2022. CO-DEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/02331888.2022.2144858.

Basak:2022:EBD

[BB22b] Sancharee Basak and Ayanendranath Basu. The extended Bregman divergence and parametric estimation. Statistics: a Journal of Theoretical and Applied Statistics, 56(3):699– 718, 2022. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-

4910 (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/02331888.2022.2070622.

Bedbur:2012:GOS

[BBK12] Stefan Bedbur, Eric Beutner, and Udo Kamps. Generalized order statistics: an exponential family in model parameters. Statistics: a Journal of Theoretical and Applied Statistics, 46(2):159–166, 2012. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Bedbur:2014:MTM

[BBK14] Stefan Bedbur, Eric Beutner, and Udo Kamps. Multivariate testing and model-checking for generalized order statistics with applications. Statistics: a Journal of Theoretical and Applied Statistics, 48(6):1297–1310, 2014. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Barczy:2019:APM

[BBKP19] Mátyás Barczy, Mohamed Ben Alaya, Ahmed Kebaier, and Gyula Pap. Asymptotic properties of maximum likelihood estimator for the growth rate of a stable CIR process based on continuous time observations. Statistics: a Journal of Theoretical and Applied Statistics, 53(3):533–568, 2019. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Benelmadani:2020:TRS

[BBL20] D. Benelmadani, K. Benhenni, and S. Louhichi. Trapezoidal rule and sampling designs for the nonparametric estimation of the regression function in models with correlated errors. Statistics: a Journal of Theoretical and Applied Statistics, 54 (1):59–96, 2020. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Bandyopadhyay:2012:RAD

[BBM12] Uttam Bandyopadhyay, Atanu Biswas, and Shirsendu Mukherjee. A response-adaptive design in crossover trial. Statistics: a Journal of Theoretical and Applied Statistics, 46(5):645–661, 2012. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Berlinet:2005:PSM

[BBR05] Alain Berlinet, Gérard Biau, and Laurent Rouvière. Parameter selection in modified histogram estimates. Statistics: a Journal of Theoretical and Applied Statistics, 39(2):91–105, 2005. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Bunke:1999:RCE

[BC99] Olaf Bunke and Ernestina Castell. Regression and contrast estimates based on adaptive regressograms depending on qualitative explanatory variables. Statistics: a Journal of Theoretical and Applied Statistics, 33(1):37–56, 1999. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Bednarski:2002:AAT

[BC02] Tadeusz Bednarski and Brenton R. Clarke. Asymptotics for an adaptive trimmed likelihood location estimator. Statistics: a Journal of Theoretical and Applied Statistics, 36(1):

1–8, 2002. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Baillo:2006:IEB

[BC06] Amparo Baíllo and Antonio Cuevas. Image estimators based on marked bins. Statistics: a Journal of Theoretical and Applied Statistics, 40(4):277–288, 2006. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Burkschat:2012:FIG

[BC12] M. Burkschat and E. Cramer. Fisher information in generalized order statistics. Statistics: a Journal of Theoretical and Applied Statistics, 46(6):719–743, 2012. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Bandyopadhyay:2014:VET

[BC14] Antar Bandyopadhyay and Sanjay Chaudhuri. Variance estimation for tree-order restricted models. Statistics: a Journal of Theoretical and Applied Statistics, 48(5):1122–1137, 2014. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Balakrishnan:2023:REB

[BC23] N. Balakrishnan and E. Castilla. Robust estimation based on one-shot device test data under log-normal lifetimes. Statistics: a Journal of Theoretical and Applied Statistics, 57(5): 1061–1086, 2023. CODEN MOSSD5. ISSN 0233-1888 (print),

1029-4910 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/02331888.2023.2240925.

Berrendero:2014:GMP

[BCCF14] José R. Berrendero, Alejandro Cholaquidis, Antonio Cuevas, and Ricardo Fraiman. A geometrically motivated parametric model in manifold estimation. Statistics: a Journal of Theoretical and Applied Statistics, 48(5):983–1004, 2014. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Balakrishnan:2011:CGD

[BCD11] N. Balakrishnan, E. Cramer, and A. Dembińska. Characterizations of geometric distribution through progressively Type-II right-censored order statistics. Statistics: a Journal of Theoretical and Applied Statistics, 45(6):559–573, 2011. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Bose:2011:MDB

[BCDD11] Mausumi Bose, Arijit Chaudhuri, Kajal Dihidar, and Shyamal Das. Model-cum-design-based estimation of the prevalence rate of a disease in a locality using spatial smoothing. Statistics: a Journal of Theoretical and Applied Statistics, 45(3):293–305, 2011. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Besse:1988:PCA

[BCFF88] P. H. Besse, H. Caussinus, L. Ferre, and J. Fine. Principal components analysis and optimization of graphical displays. Statistics: a Journal of Theoretical and Applied Statistics, 19(2):301–312, 1988. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Balakrishnan:2014:ROS

[BCJ14] N. Balakrishnan, H. R. Chareh, and Ahad Jamalizadeh. Regression via order statistics and their concomitants. Statistics: a Journal of Theoretical and Applied Statistics, 48(2): 436–446, 2014. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Balakrishnan:2005:RJD

[BCK05] N. Balakrishnan, E. Cramer, and U. Kamps. Relation for joint densities of progressively censored order statistics. Statistics: a Journal of Theoretical and Applied Statistics, 39(6):529– 536, 2005. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Balakrishnan:2001:PTI

[BCKS01] N. Balakrishnan, E. Cramer, U. Kamps, and N. Schenk. Progressive type II censored order statistics from exponential distributions. Statistics: a Journal of Theoretical and Applied Statistics, 35(4):537–556, 2001. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Bai:1990:ATL

[BCMR90] Z. D. Bai, X. R. Chen, B. Q. Miao, and C. Radhakrishna Rao. Asymptotic theory of least distances estimate in multivariate linear models. Statistics: a Journal of Theoretical and Applied Statistics, 21(4):503–519, 1990. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Brahimi:2015:CRB

[BCN15] Brahim Brahimi, Fateh Chebana, and Abdelhakim Necir. Copula representation of bivariate L-moments: a new estimation method for multiparameter two-dimensional copula models. Statistics: a Journal of Theoretical and Applied Statistics, 49(3):497–521, 2015. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Bosq:1999:OAQ

[BCP99] Denis Bosq and Nathalie Cheze-Payaud. Optimal asymptotic quadratic error of nonparametric regression function estimates for a continuous-time process from sampled-data. Statistics: a Journal of Theoretical and Applied Statistics, 32(3):229–247, 1999. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Bunke:1985:SPS

[BD85] H. Bunke and B. Droge. A stepwise procedure for the selection of nonlinear regression models. Statistics: a Journal of Theoretical and Applied Statistics, 16(1):35–45, 1985. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Bandyopadhyay:2013:MTC

[BD13] Uttam Bandyopadhyay and Saurav De. Multi-treatment covariate-adjusted adaptive design under informative censoring. Statistics: a Journal of Theoretical and Applied Statistics, 47(6):1285–1304, 2013. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Bhuyan:2017:ERC

[BD17] P. Bhuyan and A. Dewanji. Estimation of reliability with cumulative stress and strength degradation. Statistics: a Journal of Theoretical and Applied Statistics, 51(4):766–781, 2017. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Bura:2018:ATM

[BDF^{+}18] E. Bura, S. Duarte, L. Forzani, E. Smucler, and M. Sued. Asymptotic theory for maximum likelihood estimates in reduced-rank multivariate generalized linear models. Statistics: a Journal of Theoretical and Applied Statistics, 52(5):

1005–1024, 2018. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Berlinet:1995:ANE

[BDG95] Alain Berlinet, Luc Devroye, and László Györfi. Asymptotic normality of L_{1}-error in density estimation. Statistics: a Journal of Theoretical and Applied Statistics, 26(4):329–343, 1995. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Bandemer:1974:RGB

[BDP^{+}74] H. Bandemer, W. Dück, J. Piehler, N. Sieber, O. Bunke, K. Lommatzsch, H. Hollatz, H. J. Girlich, E. Läuter, P. Rudolph, H. Toutenburg, R. Thrum, and A. M. Kagan. Rezensionen. (German) [Book reviews]. Mathematische Operationsforschung und Statistik, 5(1):69–80, 1974. CODEN MTOSBI. ISSN 0047-6277.

Bunke:1999:MST

[BDP99] Olaf Bunke, Bernd Droge, and Jörg Polzehl. Model selection, transformations and variance estimation in nonlinear regression. Statistics: a Journal of Theoretical and Applied Statistics, 33(3):197–240, 1999. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Barabesi:2016:LII

[BDP16] Lucio Barabesi, Giancarlo Diana, and Pier Francesco Perri. Linearization of inequality indices in the design-based framework. Statistics: a Journal of Theoretical and Applied Statistics, 50(5):1161–??, 2016. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Bindele:2022:GSR

[BDS22] Huybrechts F. Bindele, Melody Denhere, and Wei Sun. Generalized signed-rank estimation and selection for the functional linear model. Statistics: a Journal of Theoretical and Applied Statistics, 56(4):719–738, 2022. CO-DEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/02331888.2022.2084094.

Bouzebda:2012:NTS

[BE12] Salim Bouzebda and Nour-Eddin El Faouzi. New two-sample tests based on the integrated empirical copula processes. Statistics: a Journal of Theoretical and Applied Statistics, 46(3):313–324, 2012. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Bednarski:1981:AOC

[Bed81] Tadeusz Bednarski. Application and optimality of the chisquare test of fit for testing ϵ-validity of parametric models. Statistics (Berlin, DDR), 12(1):31–41, 1981. CODEN MOSSD5. ISSN 0323-3944.

Bednarski:1985:NAE

[Bed85a] Tadeusz Bednarski. A note on asymptotic expansions of logarithms of likelihood ratios of least favourable distributions. Statistics: a Journal of Theoretical and Applied Statistics, 16(4):521–526, 1985. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Bednarski:1985:RAT

[Bed85b] Tadeusz Bednarski. A robust asymptotic testing model for special capacities. Statistics: a Journal of Theoretical and Applied Statistics, 16(4):507–519, 1985. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Bednarski:1986:ENP

[Bed86] Tadesuz Bednarski. On extension of the Neyman–Pearson lemma for capacities to dependent random variables. Statistics: a Journal of Theoretical and Applied Statistics, 17(4): 549–556, 1986. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Bednarski:1989:CRR

[Bed89] T. Bednarski. On the convergence rate of Robert procedures and local equivalence of Hellinger and total variation distances. Statistics: a Journal of Theoretical and Applied Statistics, 20(1):71–76, 1989. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Bednarski:2004:REG

[Bed04] Tadeusz Bednarski. Robust estimation in the generalized Poisson model. Statistics: a Journal of Theoretical and Applied Statistics, 38(2):149–159, 2004. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Beichelt:1974:KIZ

[Bei74] F. Beichelt. Über eine Klasse von Inspektionsmodellen der Zuverlässigkeitstheorie. (German) [On a class of inspection models of reliability theory]. Mathematische Operationsforschung und Statistik, 5(4–5):315–332, 1974. CODEN MTOSBI. ISSN 0047-6277.

Beichelt:1975:MIS

[Bei75] F. Beichelt. Minimax inspection strategies for replaceable systems with partial information on lifetime distribution. Mathematische Operationsforschung und Statistik, 6(3):479–492, 1975. CODEN MTOSBI. ISSN 0047-6277.

Beichelt:1976:GPM

[Bei76] F. Beichelt. A general preventive maintenance policy. Mathematische Operationsforschung und Statistik, 7(6):927–932, 1976. CODEN MTOSBI. ISSN 0047-6277.

Beichelt:1981:RPB

[Bei81] Frank Beichelt. Replacement policies based on system age and maintenance cost limits. Statistics (Berlin, DDR), 12(4): 621–627, 1981. CODEN MOSSD5. ISSN 0323-3944.

Bellach:1974:BSV

[Bel74] Joachim Bellach. Bayessche Schätzungen und Vorhersagen bei stochastischen linearen Differenzengleichungssystemen. (German) [Bayesian estimates and forecast for stochastic linear difference equation systems]. Mathematische Operationsforschung und Statistik, 5(7–8):599–623, 1974. CODEN MTOSBI. ISSN 0047-6277.

Bellach:1978:SSD

[Bel78] Joachim Bellach. Schätzungen in stochastischen differenzenund differentialgleiehungssystemen. Statistics (Berlin, DDR), 9(2):277–291, 1978. CODEN MOSSD5. ISSN 0323-3944.

Bellach:1979:BAD

[Bel79a] Joachim Bellach. Bemerkungen zu einer allgemeinen Definition rekursiver Schätzungen mit Anwendungsbeispielen. (German) [Comments on a general definition of recursive estimates with application examples]. Statistics (Berlin, DDR), 10(3): 367–379, 1979. CODEN MOSSD5. ISSN 0323-3944.

Bellach:1979:CLS

[Bel79b] Joachim Bellach. Consistency of the least-squares-estimators for linear stochastic difference equations. Statistics (Berlin, DDR), 10(1):79–106, 1979. CODEN MOSSD5. ISSN 0323-3944.

Bellach:1980:CAN

[Bel80] Bärbel Bellach. Consistency, asymptotic normality and asymptotic efficiency of the maximum-likelihood-estimator in linear stochastic differential equations. Statistics (Berlin, DDR), 11(2):227–266, 1980. CODEN MOSSD5. ISSN 0323-3944.

Bellach:1983:BR

[Bel83a] B. Bellach. Book reviews. Statistics (Berlin, DDR), 14(3): 420, 1983. CODEN MOSSD5. ISSN 0323-3944.

Bellach:1983:PEL

[Bel83b] Bärbel Bellach. Parameter estimators in linear stochastic differential equations and their asymptotic properties^{1}. Statistics (Berlin, DDR), 14(1):141–191, 1983. CODEN MOSSD5. ISSN 0323-3944.

Belyaev:1983:CMN

[Bel83c] Yu. K. Belyaev. The concentration method in a nonparametric Bayesian approach. Statistics (Berlin, DDR), 14 (1):73–74, 1983. CODEN MOSSD5. ISSN 0323-3944.

Bellach:1984:BR

[Bel84] B. Bellach. Book reviews. Statistics (Berlin, DDR), 15(1):90,

1984. CODEN MOSSD5. ISSN 0323-3944.

Bellach:1987:BRa

[Bel87] B. Bellach. Book review. Statistics: a Journal of Theoretical and Applied Statistics, 18(1):112, 1987. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Bellach:1988:BR

[Bel88] B. Bellach. Book review. Statistics: a Journal of Theoretical and Applied Statistics, 19(1):116, 1988. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Bellach:1990:BR

[Bel90] B. Bellach. Book reviews. Statistics: a Journal of Theoretical and Applied Statistics, 21(1):98, 1990. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Belyaev:1991:APN

[Bel91] Yu. K. Belyaev. Asymptotical properties of nonparametric point estimators based on complexly structured reliability data with right-censoring (Part 1). Statistics: a Journal of Theoretical and Applied Statistics, 22(4):589–612, 1991. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Bergmann:1974:QWT

[Ber74] Reinhard Bergmann. Qualitätsunterschiede bei Wartemodellen vom typ G/G/1. (German) [Quality differences of queueing models of type G/G/1]. Mathematische Operationsforschung und Statistik, 5(9):709–724, 1974. CODEN MTOSBI. ISSN 0047-6277.

Bereanu:1975:SSL

[Ber75] Bernard Bereanu. Stable stochastic linear programs and applications. Mathematische Operationsforschung und Statistik, 6(4):593–607, 1975. CODEN MTOSBI. ISSN 0047-6277.

Bergmann:1976:QOE

[Ber76] Joachim Bergmann. Quadratische Optimierung mit entarteter zustandsgleichung. (German) [Quadratic optimization with degenerate state equation]. Mathematische Operationsforschung und Statistik, 7(3):405–413, 1976. CODEN MTOSBI. ISSN 0047-6277.

Bergmann:1979:SCD

[Ber79] R. Bergmann. Some classes of distributions and their application in queueing. Statistics (Berlin, DDR), 10(4):583–600, 1979. CODEN MOSSD5. ISSN 0323-3944.

Berred:1998:EER

[Ber98] Alexandre M. Berred. An estimator of the exponent of regular variation based on K-record values. Statistics: a Journal of Theoretical and Applied Statistics, 32(2):93–109, 1998. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Bertrand:2000:LME

[Ber00] Pierre Bertrand. A local method for estimating change points: the “hat-function”. Statistics: a Journal of Theoretical and Applied Statistics, 34(3):215–235, 2000. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Berkes:2017:SAS

[Ber17] István Berkes. Strong approximation of the St. Petersburg game. Statistics: a Journal of Theoretical and Applied Statistics, 51(1):3–10, 2017. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Bethmann:1987:CPD

[Bet87] J. Bethmann. Characterization of Poisson distributions in the set of power mixtures. Statistics: a Journal of Theoretical and Applied Statistics, 18(2):309–320, 1987. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Bunke:1985:SFP

[BF85] Olaf Bunke and Klaus Fischer. Some fundamentals and procedures of parametric distribution-free and discrete discriminant analysis. Statistics: a Journal of Theoretical and Applied Statistics, 16(2):185–201, 1985. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Baccini:2000:GLS

[BFF00] A. Baccini, M. Fekri, and J. Fine. Generalized least squares estimation in contingency tables analysis: Asymptotic properties and applications. Statistics: a Journal of Theoretical and Applied Statistics, 34(4):267–300, 2000. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Bardet:2017:ESI

[BFN17] Jean-Marc Bardet, Konstantinos Fokianos, and Michael H. Neumann. Editorial for the special issue in honour of Paul Doukhan. Statistics: a Journal of Theoretical and Applied Statistics, 51(1):1–2, 2017. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Bergmann:1984:DSS

[BFR84] R. Bergmann, B. Fritzsche, and M. Riedel. Determination of sample sizes in point estimations. Statistics (Berlin, DDR), 15(1):73–89, 1984. CODEN MOSSD5. ISSN 0323-3944.

Bunke:1974:ELB

[BG74a] H. Bunke and J. Gladitz. Empirical linear Bayesian decision rules for a sequence of linear models with different regressor matrices. Mathematische Operationsforschung und Statistik, 5(3):235–244, 1974. CODEN MTOSBI. ISSN 0047-6277.

Burkard:1974:MAO

[BG74b] Rainer E. Burkard and Bernd Genser. Zur Methode der approximierenden Optimierung. (German) [On the method of approximating optimization]. Mathematische Operationsforschung und Statistik, 5(1):23–33, 1974. CODEN MTOSBI. ISSN 0047-6277.

Bunke:1978:PMC

[BG78] O. Bunke and B. Grabowski. A procedure for model choice or variable selection with controlled model specification error. Statistics (Berlin, DDR), 9(4):483–497, 1978. CODEN MOSSD5. ISSN 0323-3944.

Bunke:1979:EBA

[BG79] H. Bunke and J. Gladitz. Empirical Bayes approach to parameter identification in linear stochastic difference equations. Statistics (Berlin, DDR), 10(1):63–78, 1979. CODEN MOSSD5. ISSN 0323-3944.

Bodnar:2011:EPM

[BG11] Taras Bodnar and Arjun K. Gupta. Estimation of the precision matrix of a multivariate elliptically contoured stable distribution. Statistics: a Journal of Theoretical and Applied Statistics, 45(2):131–142, 2011. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Bose:2014:PRP

[BG14] Arup Bose and Sreela Gangopadhyay. Pfeifer records process. Statistics: a Journal of Theoretical and Applied Statistics, 48(1):129–141, 2014. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Bibi:2016:MDE

[BG16] Abdelouahab Bibi and Ahmed Ghezal. Minimum distance estimation of Markov-switching bilinear processes. Statistics: a Journal of Theoretical and Applied Statistics, 50(6):1290– 1309, 2016. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Bai:2019:LSM

[BG19a] Ray Bai and Malay Ghosh. Large-scale multiple hypothesis testing with the normal-beta prime prior. Statistics: a Journal of Theoretical and Applied Statistics, 53(6):1210–1233, 2019. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Baringhaus:2019:ARE

[BG19b] Ludwig Baringhaus and Daniel Gaigall. On an asymptotic relative efficiency concept based on expected volumes of confidence regions. Statistics: a Journal of Theoretical and Applied Statistics, 53(6):1396–1436, 2019. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Berlinet:2001:ANC

[BGAL01] Alain Berlinet, Ali Gannoun, and Eric Atzner-Løber. Asymptotic normality of convergent estimates of conditional quantiles. Statistics: a Journal of Theoretical and Applied Statistics, 35(2):139–169, 2001. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Bittner:1972:RGB

[BGB^{+}72] L. Bittner, A. Göpfert, H. L. Burmeister, P. H. Müller, H. Läuter, W. Richter, M. Schoch, O. Bunke, G. Wintgen, and J. Kunze. Rezensionen. (German) [Book reviews]. Mathematische Operationsforschung und Statistik, 3(4):309–316, 1972. CODEN MTOSBI. ISSN 0047-6277.

Bouaziz:2015:SIR

[BGL15] Olivier Bouaziz, Ségolen Geffray, and Olivier Lopez. Semiparametric inference for the recurrent events process by means of a single-index model. Statistics: a Journal of Theoretical and Applied Statistics, 49(2):361–385, 2015. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Balakrishnan:2015:SPT

[BGLS15] N. Balakrishnan, R. Gouet, F. J. López, and G. Sanz. Sequential precedence tests. Statistics: a Journal of Theoretical and Applied Statistics, 49(1):224–238, 2015. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Balakrishnan:1995:EMS

[BGP95] N. Balakrishnan, Shanti S. Gupta, and S. Panchapakesan. Estimation of the mean and standard deviation of the logistic distribution based on multiply Type-II censored samples. Statistics: a Journal of Theoretical and Applied Statistics, 27 (1–2):127–142, 1995. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Barrera:2006:CSD

[BGTV06] Nelson P. Barrera, Manuel Galea, Soledad Torres, and Manuel Villalón. Class of skew-distributions: theory and applications in biology. Statistics: a Journal of Theoretical and Applied Statistics, 40(4):365–375, 2006. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Bentkus:2006:OHL

[BGV06] V. Bentkus, G. D. C. Geuze, and M. C. A. Van Zuijlen. Optimal Hoeffding-like inequalities under a symmetry assumption. Statistics: a Journal of Theoretical and Applied Statistics, 40(2):159–164, 2006. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Bohning:1986:RNE

[BH86] Dankmar Böhning and Karl-Heinz Hoffmann. A remark on the numerical estimation of probabilities. Statistics: a Journal of Theoretical and Applied Statistics, 17(2):231–236, 1986. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Bauer:1987:MTS

[BH87] Peter Bauer and Peter Hackl. Multiple testing in a set of nested hypotheses. Statistics: a Journal of Theoretical and Applied Statistics, 18(3):345–349, 1987. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Bohning:1989:MCS

[BH89] D. Böhning and H. Holling. A Monte Carlo study on minimizing chi-square distances under the hypothesis of homogeneity or independence for a two-way cotingeney table. Statistics: a Journal of Theoretical and Applied Statistics, 20(1):55–70, 1989. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Balakrishna:2018:AAM

[BH18] N. Balakrishna and G. Hareesh. Analysis of autoregressive models with symmetric stable innovations. Statistics: a Journal of Theoretical and Applied Statistics, 52(2):288–302, 2018. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Bhattacharya:1988:CSE

[Bha88] C. G. Bhattacharya. On the Cohen–Sackrowitz estimator of a common mean. Statistics: a Journal of Theoretical and Applied Statistics, 19(4):493–501, 1988. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Bhandary:2007:RTD

[Bha07] Madhusudan Bhandary. On robustness of the test for detection of multivariate outliers. Statistics: a Journal of Theoretical and Applied Statistics, 41(5):377–384, 2007. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Bar-Hen:2004:ISM

[BHM04] Avner Bar-Hen and Frédéric Mortier. Influence and sensitivity measures in correspondence analysis. Statistics: a Journal of Theoretical and Applied Statistics, 38(3):207–215, 2004. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Barmalzan:2015:SCS

[BHN15] Ghobad Barmalzan, Abedin Haidari, and Amir T. Payandeh Najafabadi. Stochastic comparisons of sample spacings in single- and multiple-outlier exponential models. Statistics: a Journal of Theoretical and Applied Statistics, 49(4):815–830, 2015. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Burke:1976:WCP

[BHS76] H. Burke, C. Hennig, and W. H. Schmidt. Weighted combination of prior and sample information and parameter estimation. Mathematische Operationsforschung und Statistik, 7 (5):665–672, 1976. CODEN MTOSBI. ISSN 0047-6277.

Bunke:1977:PEN

[BHSW77] H. Bunke, K. Henschke, R. Strüby, and C. Wisotzki. Parameter estimation in nonlinear regression models. Statistics (Berlin, DDR), 8(1):23–40, 1977. CODEN MOSSD5. ISSN 0323-3944.

Bunke:1975:RGB

[BHW75] O. Bunke, H. Hollatz, and H. Weinert. Rezension. (German) [Book review]. Mathematische Operationsforschung und Statistik, 6(1):155–160, 1975. CODEN MTOSBI. ISSN 0047-6277.

Bickel:2018:BRP

[Bic18] David R. Bickel. Bayesian revision of a prior given priordata conflict, expert opinion, or a similar insight: a largedeviation approach. Statistics: a Journal of Theoretical and Applied Statistics, 52(3):552–570, 2018. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Bickel:2023:FSS

[Bic23] David R. Bickel. Fiducialize statistical significance: transforming p-values into conservative posterior probabilities and Bayes factors. Statistics: a Journal of Theoretical and Applied Statistics, 57(4):941–959, 2023. CO-DEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/02331888.2023.2232912.

Bieniek:2006:PBE

[Bie06] Mariusz Bieniek. Projection bounds on expectations of generalized order statistics from DFR and DFRA families. Statistics: a Journal of Theoretical and Applied Statistics, 40(4): 339–351, 2006. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Bieniek:2008:PBE

[Bie08] Mariusz Bieniek. Projection bounds on expectations of spacings of generalized order statistics from DFR and DFRA families. Statistics: a Journal of Theoretical and Applied Statistics, 42(3):231–243, 2008. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Bieniek:2015:OBB

[Bie15] Mariusz Bieniek. Optimal bounds on the bias of quasimidranges. Statistics: a Journal of Theoretical and Applied Statistics, 49 (6):1382–1399, 2015. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Bieniek:2016:OBM

[Bie16] Mariusz Bieniek. Optimal bounds for the mean of the total time on test for distributions with decreasing generalized failure rate. Statistics: a Journal of Theoretical and Applied Statistics, 50(6):1206–1220, 2016. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Billiot:1997:APT

[Bil97] Jean-Michel Billiot. Asymptotic properties of Takacs–Fiksel estimation method for Gibbs point processes. Statistics: a Journal of Theoretical and Applied Statistics, 30(1):69–89, 1997. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Binder:1996:D

[Bin96] David A. Binder. Discussion. Statistics: a Journal of Theoretical and Applied Statistics, 27(3–4):237–241, 1996. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Bischoff:1995:D

[Bis95a] Wolfgang Bischoff. Discussion. Statistics: a Journal of Theoretical and Applied Statistics, 26(2):150–153, 1995. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Bischoff:1995:LBE

[Bis95b] Wolfgang Bischoff. Lower bounds for the efficiency of designs with respect to the D-criterion when the observations are correlated. Statistics: a Journal of Theoretical and Applied Statistics, 27(1–2):27–44, 1995. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Bittner:1972:QPD

[Bit72] L. Bittner. Quasilinearization and primal-dual relations of Chebyshev approximation. I. Mathematische Operationsforschung und Statistik, 3(6):431–451, 1972. CODEN MTOSBI. ISSN 0047-6277.

Bittner:1975:OCP

[Bit75] L. Bittner. On optimal control of processes governed by abstract functional, integral and hyperbolic differential equations. Mathematische Operationsforschung und Statistik, 6 (1):107–134, 1975. CODEN MTOSBI. ISSN 0047-6277.

Balakrishnan:1985:BMS

[BJ85] N. Balakrishnan and P. C. Joshi. Bounds for the mean of second largest order statistic in large samples. Statistics: a Journal of Theoretical and Applied Statistics, 16(3):457–464, 1985. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Bunke:2005:SIL

[BJ05] Olaf Bunke and Jan Johannes. Self-informative limits of Bayes estimates and generalized maximum likelihood. Statistics: a Journal of Theoretical and Applied Statistics, 39(6):483–502, 2005. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Bang:2014:ASN

[BJ14] Sungwan Bang and Myoungshic Jhun. Adaptive sup-norm regularized simultaneous multiple quantiles regression. Statistics: a Journal of Theoretical and Applied Statistics, 48(1): 17–33, 2014. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Bai:2013:TLH

[BJfYZ13] Zhidong Bai, Dandan Jiang, Jian feng Yao, and Shurong Zheng. Testing linear hypotheses in high-dimensional regressions. Statistics: a Journal of Theoretical and Applied Statistics, 47(6):1207–1223, 2013. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Bakouch:2014:NDD

[BJN14] Hassan S. Bakouch, M. Aghababaei Jazi, and Saralees Nadarajah. A new discrete distribution. Statistics: a Journal of Theoretical and Applied Statistics, 48(1):200–240, 2014. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Bandemer:1971:BAO

[BK71] H. Bandemer and H. Klemm. Bericht über die Arbeitstagung “Optimale Versuchsplanung”. (German) [Report on the working conference “Optimal Experimental Planning”]. Mathematische Operationsforschung und Statistik, 2(3):255–256, 1971. CODEN MTOSBI. ISSN 0047-6277.

Baksalary:1976:CEM

[BK76] J. K. Baksalary and R. Kala. Criteria for estimability in multivariate linear models. Mathematische Operationsforschung und Statistik, 7(1):5–9, 1976. CODEN MTOSBI. ISSN 0047-6277.

Baksalary:1977:SSP

[BK77] J. K. Baksalary and R. Kala. Sums of squares and products matrices for a non-full ranks hypothesis in the model of Potthoff and Roy. Statistics (Berlin, DDR), 8(4):459–465, 1977. CODEN MOSSD5. ISSN 0323-3944.

Baksalary:1979:BLU

[BK79a] J. K. Baksalary and R. Kala. Best linear unbiased estimation in the restricted general linear model. Statistics (Berlin, DDR), 10(1):27–35, 1979. CODEN MOSSD5. ISSN 0323-3944.

Baksalary:1979:PPS

[BK79b] J. K. Baksalary and R. Kala. On the prediction problem in the seemingly unrelated regression equations model. Statistics (Berlin, DDR), 10(2):203–208, 1979. CODEN MOSSD5. ISSN 0323-3944.

Bromek:1986:SIC

[BK86] T. Bromek and T. Kowalczyk. Statistical inference in the case of unordered pairs. Statistics: a Journal of Theoretical and Applied Statistics, 17(3):357–364, 1986. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Bedbur:2017:ITP

[BK17] S. Bedbur and U. Kamps. Inference in a two-parameter generalized order statistics model. Statistics: a Journal of Theoretical and Applied Statistics, 51(5):1132–1142, 2017. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Barmalzan:2022:OFS

[BKHB22] Ghobad Barmalzan, Sajad Kosari, Ali Akbar Hosseinzadeh, and Narayanaswamy Balakrishnan. Ordering fail-safe systems having dependent components with Archimedean copula and exponentiated location-scale distributions. Statistics: a Journal of Theoretical and Applied Statistics, 56(3):631–661, 2022. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/02331888.2022.2061488.

Bairamov:2003:NML

[BKK03] Ismihan Bairamov, Samuel Kotz, and Tomasz Kozubowski. A new measure of linear local dependence. Statistics: a Journal of Theoretical and Applied Statistics, 37(3):243–258, 2003. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Beier-Kuchler:1970:KBA

[BKN70] Ingeborg Beier-Küchler and Peter Neumann. Über die Klassenwahl beim χ^{2}-Anpassungstest. (German) [On the class choice in the χ^{2} fit test]. Mathematische Operationsforschung und Statistik, 1(1):55–68, 1970. CODEN MTOSBI. ISSN 0047-6277.

Bunke:1991:BR

[BKS^{+}91] O. Bunke, U. Küchler, D. Stoyan, K. H. Fichtner, W. Kölzow, O. Bunke, B. Droge, H. Bandemer, H. Bandemer, B. Droge, Hans M. Dietz, B. Selfert, R. Döhler, K. Fischer, Regina Storm, B. Droge, and F. Liese. Book reviews. Statistics: a Journal of Theoretical and Applied Statistics, 22(2):313–328, 1991. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Bhushan:2023:IAC

[BKS23] Shashi Bhushan, Anoop Kumar, and Shivam Shukla. Impact assessment of correlated measurement errors using logarithmic-type estimators. Statistics: a Journal of Theoretical and Applied Statistics, 57(5):1010–1036, 2023. CO-DEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/02331888.2023.2260915.

Babayi:2014:IGL

[BKT14] Salman Babayi, Esmaile Khorram, and Farzad Tondro. Inference of R = P [X < Y] for generalized logistic distribution. Statistics: a Journal of Theoretical and Applied Statistics, 48(4):862–871, 2014. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Bednarski:1978:NBT

[BL78] T. Bednarski and T. Ledwtna. A note on biasedness of tests of fit. Statistics (Berlin, DDR), 9(2):191–193, 1978. CODEN MOSSD5. ISSN 0323-3944.

Bednarski:1982:NWC

[BL82] Tadeusz Bednarski and Teresa Ledwina. A note on the weak convergence of an estimator of monotonic dependence function of two random variables. Statistics (Berlin, DDR), 13 (4):555–560, 1982. CODEN MOSSD5. ISSN 0323-3944.

Belitser:1996:AMN

[BL96a] E. N. Belitser and B. Y. Levit. Asymptotically minimax nonparametric regression in L_{2}. Statistics: a Journal of Theoretical and Applied Statistics, 28(2):105–122, 1996. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Bogdan:1996:TUL

[BL96b] M. Bogdan and T. Ledwina. Testing uniformity via log-spline modeling. Statistics: a Journal of Theoretical and Applied Statistics, 28(2):131–157, 1996. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Bolfarine:2006:SBR

[BL06] Heleno Bolfarine and V. H. Lachos. Skew binary regression with measurement errors. Statistics: a Journal of Theoretical and Applied Statistics, 40(6):485–494, 2006. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Balan:2011:CLT

[BL11] Raluca Balan and Sana Louhichi. A cluster-limit theorem for infinitely divisible point processes. Statistics: a Journal of Theoretical and Applied Statistics, 45(1):3–18, 2011. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Batsidis:2015:HEF

[BL15] Apostolos Batsidis and Artur J. Lemonte. On the Harris extended family of distributions. Statistics: a Journal of Theoretical and Applied Statistics, 49(6):1400–1421, 2015. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Blacker:1988:NFC

[Bla88] R. Blacker. A new form for the chi-squared test of independence. Statistics: a Journal of Theoretical and Applied Statistics, 19(4):519–536, 1988. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Blacher:1993:HOC

[Bla93] R. Blacher. Higher order correlation coefficients. Statistics: a Journal of Theoretical and Applied Statistics, 25(1):1–15, 1993. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Blandin:2014:ARR

[Bla14] Vassili Blandin. Asymptotic results for random coefficient bifurcating autoregressive processes. Statistics: a Journal of Theoretical and Applied Statistics, 48(6):1202–1232, 2014. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Buch-Larsen:2005:KDE

[BLNGB05] Tine Buch-Larsen, Jens Perch Nielsen, Montserrat Guillén, and Catalina Bolancé. Kernel density estimation for heavytailed distributions using the Champernowne transformation. Statistics: a Journal of Theoretical and Applied Statistics, 39(6):503–516, 2005. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Berzin:2003:CNL

[BLO03a] Corinne Berzin, José León, and Joaquín Ortega. Convergence of non-linear functionals of smoothed empirical processes and kernel density estimates. Statistics: a Journal of Theoretical and Applied Statistics, 37(3):217–242, 2003. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Bloznelis:2003:NBC

[Blo03b] M. Bloznelis. A note on the bias and consistency of the jackknife variance estimator in stratified samples. Statistics: a Journal of Theoretical and Applied Statistics, 37(6):489–504, 2003. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Bloznelis:2007:SOR

[Blo07] Mindaugas Bloznelis. Second-order and resampling approximation of finite population U-statistics based on stratified samples. Statistics: a Journal of Theoretical and Applied Statistics, 41(4):321–332, 2007. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Balakrishnan:2009:MIG

[BLSC09] N. Balakrishnan, Víctor Leiva, Antonio Sanhueza, and Enrique Cabrera. Mixture inverse Gaussian distributions and its transformations, moments and applications. Statistics: a Journal of Theoretical and Applied Statistics, 43(1):91–104, 2009. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Basikhasteh:2020:OMR

[BLT20] Mehdi Basikhasteh, Fazlollah Lak, and Saeid Tahmasebi. Ordered maximum ranked set sampling with unequal sample. Statistics: a Journal of Theoretical and Applied Statistics, 54(4):856–870, 2020. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Bandemer:1971:BAP

[BM71] H. Bandemer and K. Matthes. Bericht über die “6. Arbeitstagung über Probleme der Stochastik”. (German) [Report on the “6th Workshop on Stochastic Problems”]. Mathematische Operationsforschung und Statistik, 2(2):167–169, 1971. CODEN MTOSBI. ISSN 0047-6277.

Bunke:1984:MER

[BM84] O. Bunke and M. Möhner. Minimax estimators of regression functions under normalized quadratic loss functions and inequality restrictions. Statistics (Berlin, DDR), 15(4):471–482, 1984. CODEN MOSSD5. ISSN 0323-3944.

Bonneu:1994:MAC

[BM94] Michel Bonneu and Xavier Milhaud. A modified Akaike criterion for model choice in generalized linear models. Statistics: a Journal of Theoretical and Applied Statistics, 25(3): 225–238, 1994. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Bednarski:2001:OBI

[BM01] Tadeusz Bednarski and Christine H. Müller. Optimal bounded influence regression and scale m-estimators in the context of experimental design. Statistics: a Journal of Theoretical and Applied Statistics, 35(4):349–369, 2001. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Basu:2002:TEA

[BM02] Sujit Kumar Basu and Murari Mitra. Testing exponentiality against Laplace order dominance-. Statistics: a Journal of Theoretical and Applied Statistics, 36(3):223–229, 2002. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Baillo:2009:MSE

[BM09] Amparo Baíllo and Isabel Molina. Mean-squared errors of small-area estimators under a unit-level multivariate model. Statistics: a Journal of Theoretical and Applied Statistics, 43(6):553–569, 2009. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Balakrishna:2012:SIE

[BM12] N. Balakrishna and Angel Mathew. Sequential interval estimation of the limiting interval availability for a bivariate stationary dependent sequence. Statistics: a Journal of Theoretical and Applied Statistics, 46(2):185–196, 2012. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Bieniek:2018:PUC

[BM18] Mariusz Bieniek and Krystyna Maciag. On the problem of the unique characterization of discrete distributions by single regression of weak records. Statistics: a Journal of Theoretical and Applied Statistics, 52(3):533–551, 2018. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Balakrishnan:2023:OJP

[BM23] N. Balakrishnan and Rahul Mukerjee. On optimal joint prediction of order statistics. Statistics: a Journal of Theoretical and Applied Statistics, 57(5):1239–1250, 2023. CO-DEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/02331888.2023.2249572.

Bedbur:2016:HTG

[BMK16] Stefan Bedbur, Norman Müller, and Udo Kamps. Hypotheses testing for generalized order statistics with simple order restrictions on model parameters under the alternative. Statistics: a Journal of Theoretical and Applied Statistics, 50(4): 775–790, 2016. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Balakrishnan:2022:ETE

[BMK22] N. Balakrishnan, Deemat C. Mathew, and Sudheesh K. Kattumannil. An exact test for exponentiality against renewal increasing mean residual life class. Statistics: a Journal of Theoretical and Applied Statistics, 56(1):164–181, 2022. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Blazquez:2013:RTD

[BMM13] Fernando López Blázquez, Antonia Castaño Martínez, and Begoña Salamanca Miño. Random translation of discrete order statistics. Statistics: a Journal of Theoretical and Applied Statistics, 47(6):1241–1248, 2013. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Basu:2016:GWT

[BMMP16] A. Basu, A. Mandal, N. Martin, and L. Pardo. Generalized Wald-type tests based on minimum density power divergence estimators. Statistics: a Journal of Theoretical and Applied Statistics, 50(1):1–26, 2016. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic). URL http://

www.tandfonline.com/doi/abs/10.1080/02331888.2015.
1016435.

Balakrishnan:2015:EPD

[BMP15] N. Balakrishnan, N. Martín, and L. Pardo. Empirical phidivergence test statistics for testing simple and composite null hypotheses. Statistics: a Journal of Theoretical and Applied Statistics, 49(5):951–977, 2015. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Belzunce:2016:CHR

[BMRPZ16] Félix Belzunce, Carolina Martínez-Riquelme, Franco Pellerey, and Saeed Zalzadeh. Comparison of hazard rates for dependent random variables. Statistics: a Journal of Theoretical and Applied Statistics, 50(3):630–648, 2016. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Badia:2020:TTI

[BMS20] F. G. Badía, Sophie Mercier, and C. Sangüesa. Trendtransformed independent vectors: construction and stochastic comparison results. Statistics: a Journal of Theoretical and Applied Statistics, 54(4):778–804, 2020. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Bandemer:1978:CEE

[BN78] Hans Bandemer and Sabine Nagel. On construction of exact experimental designs from discrete ones. Statistics (Berlin, DDR), 9(4):507–517, 1978. CODEN MOSSD5. ISSN 0323-3944.

Boltze:1980:SRE

[BN80] Lothar Boltze and Wolfgang Näther. Some remarks on experimental design for estimating the expectation of a stationary random process. Statistics (Berlin, DDR), 11(4):475–481, 1980. CODEN MOSSD5. ISSN 0323-3944.

Bandemer:1981:PEL

[BN81] H. Bandemer and W. Nagel. Parameter estimation in linear regression models with weak and fuzzy prior knowledge. Statistics (Berlin, DDR), 12(3):297–305, 1981. CODEN MOSSD5. ISSN 0323-3944.

Boltze:1982:EOM

[BN82] L. Boltze and W. Näther. On effective observation methods in regression models with correlated errors. Statistics (Berlin, DDR), 13(4):507–519, 1982. CODEN MOSSD5. ISSN 0323-3944.

Bagdonavicius:1997:AAS

[BN97] V. Bagdonavicius and M. Nikulin. Asymptotical analysis of semiparametric models in survival analysis and accelerated life testing. Statistics: a Journal of Theoretical and Applied Statistics, 29(3):261–283, 1997. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Bagdonavicius:2000:NEA

[BN00a] V. Bagdonavicius and M. Nikulin. On nonparametric estimation in accelerated experiments with step-stresses. Statistics: a Journal of Theoretical and Applied Statistics, 33(4): 349–365, 2000. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Brostrom:2000:ARS

[BN00b] Göran Broström and Leif Nilsson. Acceptance–rejection sampling from the conditional distribution of independent discrete random variables, given their sum. Statistics: a Journal of Theoretical and Applied Statistics, 34(3):247–257, 2000. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Bandekar:2003:MVC

[BN03] Rajesh R. Bandekar and Daya K. Nagar. Matrix variate Cauchy distribution. Statistics: a Journal of Theoretical and Applied Statistics, 37(6):537–550, 2003. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Barakat:2014:ADO

[BN14] H. M. Barakat and E. M. Nigm. Asymptotic distributions of order statistics and record values arising from the class of beta-generated distributions. Statistics: a Journal of Theoretical and Applied Statistics, 48(5):1005–1012, 2014. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Bidram:2016:NLM

[BN16] Hamid Bidram and Saralees Nadarajah. A new lifetime model with decreasing, increasing, bathtub-shaped, and upsidedown bathtub-shaped hazard rate function. Statistics: a Journal of Theoretical and Applied Statistics, 50(1):139–156, 2016. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910

(electronic). URL http://www.tandfonline.com/doi/abs/
10.1080/02331888.2015.1006224.

Barakat:2013:LTR

[BNAA13] H. M. Barakat, E. M. Nigm, and M. A. Al-Awady. Limit theorems for random maximum of independent and nonidentically distributed random vectors. Statistics: a Journal of Theoretical and Applied Statistics, 47(3):546–557, 2013. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Barakat:2020:LTU

[BNH20] H. M. Barakat, E. M. Nigm, and M. H. Harpy. Limit theorems for univariate and bivariate order statistics with variable ranks. Statistics: a Journal of Theoretical and Applied Statistics, 54(4):737–755, 2020. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Bandemer:1987:OMO

[BNP87] Hans Bandemer, Wolfgang Näther, and Jürgen Pilz. Once more: optimal experimental design for regression models (with discussion). Statistics: a Journal of Theoretical and Applied Statistics, 18(2):171–217, 1987. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Baumgart:1973:OFH

[BNS73] Gunther Baumgart, Peter Neumann, and Dieter Schramme. Optimierung des Flugzeugeinsatzes mit Hilfe eines Verfahrens der blinden Suche. (German) [Optimization of aircraft utilization by a process of blind search]. Mathematische Operationsforschung und Statistik, 4(1):51–62, 1973. CODEN MTOSBI. ISSN 0047-6277.

Bercu:2015:ABR

[BNS15] Bernard Bercu, Thi Mong Ngoc Nguyen, and Jerome Saracco. On the asymptotic behaviour of the recursive Nadaraya– Watson estimator associated with the recursive sliced inverse regression method. Statistics: a Journal of Theoretical and Applied Statistics, 49(3):660–679, 2015. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Bensaid:2001:HER

[BO01] Nadia Bensaïd and Paulo Eduardo Oliveira. Histogram estimation of Radon–Nikodym derivatives for strong mixing data. Statistics: a Journal of Theoretical and Applied Statistics, 35(4):569–592, 2001. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Bod:1975:MCP

[Bod75] Péter Bod. Minimality and complementarity properties of s.c. Z-functions and a forgotten theorem due to Georg Wintgen. Mathematische Operationsforschung und Statistik, 6(6):867– 872, 1975. CODEN MTOSBI. ISSN 0047-6277.

Boeselt:1972:KFG

[Boe72] Martin R. J. Boeselt. Kompakte Fertigungsbäume. (German) [Compact production trees]. Mathematische Operationsforschung und Statistik, 3(5):333–338, 1972. CODEN MTOSBI. ISSN 0047-6277.

Bohing:1981:COE

[Böh81a] Dankmar Böhing. On the construction of optimal experimental designs: a penalty approach. Statistics (Berlin, DDR), 12 (4):487–495, 1981. CODEN MOSSD5. ISSN 0323-3944.

Bohning:1981:ABK

[Böh81b] Dankmar Böhning. On an assumption of Bandemer and Ketzel in optimal experimental design theory. Statistics (Berlin, DDR), 12(4):497–502, 1981. CODEN MOSSD5. ISSN 0323-3944.

Bohning:1983:DTA

[Böh83] Dankmar Böhning. A duality theorem with applications to statistics. Statistics (Berlin, DDR), 14(4):551–557, 1983. CO-DEN MOSSD5. ISSN 0323-3944.

Balakrishnan:2012:NRT

[BOHJ12] N. Balakrishnan, S. H. Ong, M. Habibi, and A. Jamalizadeh. Nonlinear regression in a trivariate elliptical distribution via a linear combination of order statistics. Statistics: a Journal of Theoretical and Applied Statistics, 46(5):605–619, 2012. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Boker:1985:LGC

[Bök85] Fred Böker. Limits of generalized compound point processes. Statistics: a Journal of Theoretical and Applied Statistics, 16 (1):97–105, 1985. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Boneva:1982:SSA

[Bon82] L. Boneva. Seriation and some applications. Statistics (Berlin, DDR), 13(3):473–485, 1982. CODEN MOSSD5. ISSN 0323-3944.

Bonneu:1988:MCP

[Bon88] Michel Bonneu. Model choice for prediction in generalized linear models. Statistics: a Journal of Theoretical and Applied Statistics, 19(3):369–382, 1988. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Borobev:1970:PIP

[Bor70] N. N. Borobev. Pozitsionniye igri i prinyatie statisticheskich reshennije. (Russian) [Positional games and statistical decisions]. Mathematische Operationsforschung und Statistik, 1 (1):45–53, 1970.

Borngraber:1973:HLD

[Bor73] Ernst-Egon Borngräber. Die Herleitung von Lösungsformeln für einen dynamischen Aufteilungsprozeß mit dem Bellman-Algorithmus. (German) [The derivation of solution formulas for a dynamic breakdown process with the Bellman algorithm]. Mathematische Operationsforschung und Statistik, 4(6):445–452, 1973. CODEN MTOSBI. ISSN 0047-6277.

Bose:1988:HOA

[Bos88] Arup Bose. Higher order approximations for autocovariances from linear processes with applications. Statistics: a Journal of Theoretical and Applied Statistics, 19(2):259–269, 1988. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Bosq:1993:BTL

[Bos93] Denis Bosq. Bernstein-type large deviations inequalities for partial sums of strong mixing processes. Statistics: a Journal of Theoretical and Applied Statistics, 24(1):59–70, 1993.

CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Bottcher:2020:CVD

[Böt20] Björn Böttcher. Copula versions of distance multivariance and dHSIC via the distributional transform — a general approach to construct invariant dependence measures. Statistics: a Journal of Theoretical and Applied Statistics, 54(3): 577–594, 2020. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Bowden:2012:IMS

[Bow12] Roger J. Bowden. Information, measure shifts and distribution metrics. Statistics: a Journal of Theoretical and Applied Statistics, 46(2):249–262, 2012. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Bowden:2021:ELO

[Bow21] Roger J. Bowden. The entropic log odds as a measure of distribution asymmetry. Statistics: a Journal of Theoretical and Applied Statistics, 55(2):243–249, 2021. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Biscay:1993:DBP

[BP93] Rolando Biscay and Roberto Pascual. Distance between probability measures based on the relative operating characteristic curves. Statistics: a Journal of Theoretical and Applied Statistics, 24(4):371–376, 1993. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Barczy:2016:APM

[BP16] Mátyás Barczy and Gyula Pap. Asymptotic properties of maximum-likelihood estimators for Heston models based on continuous time observations. Statistics: a Journal of Theoretical and Applied Statistics, 50(2):389–417, 2016. CO-DEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/02331888.2015.1044991.

Benke:2017:OPS

[BP17] János Marcell Benke and Gyula Pap. One-parameter statistical model for linear stochastic differential equation with time delay. Statistics: a Journal of Theoretical and Applied Statistics, 51(3):510–531, 2017. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Belomestny:2018:SEN

[BP18] Denis Belomestny and Vladimir Panov. Semiparametric estimation in the normal variance-mean mixture model. Statistics: a Journal of Theoretical and Applied Statistics, 52(3): 571–589, 2018. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Bauer:1988:MSM

[BPH88] Peter Bauer, Benedikt M. Pötscher, and Peter Hackl. Model selection by multiple test procedures. Statistics: a Journal of Theoretical and Applied Statistics, 19(1):39–44, 1988. CO-DEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Bhattacharya:2014:AHC

[BPK14] Shrijita Bhattacharya, Biswabrata Pradhan, and Debasis Kundu. Analysis of hybrid censored competing risks data. Statistics: a Journal of Theoretical and Applied Statistics, 48 (5):1138–1154, 2014. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Bertail:2000:SCP

[BPR00] Patrice Bertail, Dimitris N. Politis, and Nourheddine Rhomari. Subsampling continuous parameter random fields and a Bernstein inequality. Statistics: a Journal of Theoretical and Applied Statistics, 33(4):367–392, 2000. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Baran:2011:PES

[BPV11] Sándor Baran, Gyula Pap, and Martien C. A. Van Zuijlen. Parameter estimation of a shifted Wiener sheet. Statistics: a Journal of Theoretical and Applied Statistics, 45(4):319–335, 2011. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Baran:2004:AIU

[BPvZ04] Sándor Baran, Gyula Pap, and Martien C. A. van Zuijlen. Asymptotic inference for an unstable spatial AR model.

Statistics: a Journal of Theoretical and Applied Statistics, 38(6):465–482, 2004. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Bang-Qiang:2020:PEL

[BQSR20] He Bang-Qiang and Lv Sheng-Ri. Penalized empirical likelihood for high-dimensional partially linear errors-in-function model with martingale difference errors. Statistics: a Journal of Theoretical and Applied Statistics, 54(6):1214–1231, 2020. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Bunke:1983:NBO

[BR83] O. Bunke and S. Riemer. A note on bootstrap and other empirical procedures for testing linear hypotheses ’without normality. Statistics (Berlin, DDR), 14(4):517–526, 1983. CO-DEN MOSSD5. ISSN 0323-3944.

Basu:1990:RCC

[BR90] A. K. Basu and S. Sen Roy. On rates of convergence in the Central Limit Theorem for parameter estimation in general autoregressive model. Statistics: a Journal of Theoretical and Applied Statistics, 21(3):461–470, 1990. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Bdair:2012:SUB

[BR12] Omar M. Bdair and Mohammad Z. Raqab. Sharp upper bounds for the mean residual waiting time of records. Statistics: a Journal of Theoretical and Applied Statistics, 46(1): 69–84, 2012. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Brunel:2015:PCE

[BR15] E. Brunel and A. Roche. Penalized contrast estimation in functional linear models with circular data. Statistics: a Journal of Theoretical and Applied Statistics, 49(6):1298–1321, 2015. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Bickel:2019:MFM

[BR19] David R. Bickel and Abbas Rahal. Model fusion and multiple testing in the likelihood paradigm: shrinkage and evidence supporting a point null hypothesis. Statistics: a Journal of Theoretical and Applied Statistics, 53(6):1187–1209, 2019. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Brandt:1988:BR

[Bra88] A. Brandt. Book review. Statistics: a Journal of Theoretical and Applied Statistics, 19(2):282, 1988. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Brandt:1991:BR

[Bra91] A. Brandt. Book reviews. Statistics: a Journal of Theoretical and Applied Statistics, 22(1):68, 1991. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Bandemer:1977:BRb

[BRB^{+}77] H. Bandemer, P. Rudolph, O. Bunke, G. Herrendörfer, H. Toutenburg, M. Nussbaum, K. Henschke, B. Seifert, F. Liese, D. König, D. Stoyan, and S. Pfeil. Book review. Statistics (Berlin, DDR), 8(4):577–588, 1977. CODEN MOSSD5. ISSN 0323-3944.

Brendel:2005:THU

[Bre05] Michael Brendel. Testing hypotheses under a generalized Koziol–Green model with partially informative censoring. Statistics: a Journal of Theoretical and Applied Statistics, 39(4):329–345, 2005. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Bouezmarni:2020:SCB

[BRF20] Taoufik Bouezmarni, Yassir Rabhi, and Charles Fontaine. A semiparametric copula-based estimation of the regression function for right-censored data. Statistics: a Journal of Theoretical and Applied Statistics, 54(1):46–58, 2020. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Benesa:1999:PME

[BRKO99] Viktor Benesa, Jan Ratajb, Pavel Krejcír, and Joachim Ohserc. Projection measures and estimation variances of intensities. Statistics: a Journal of Theoretical and Applied Statistics, 32(4):369–393, 1999. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Brodeau:1993:TCA

[Bro93] F. Brodeau. Test for the choice of approximative models in nonlinear regression when the variance is unknown. Statistics: a Journal of Theoretical and Applied Statistics, 24(2):95–106, 1993. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Bolfarine:1993:SAR

[BRZ93] H. Bolfarine, J. Rodrigues, and S. Zacks. Some asymptotic results in finite populations. Statistics: a Journal of Theoretical and Applied Statistics, 24(4):359–370, 1993. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Bunke:1975:EPI

[BS75] H. Bunke and R. Strüby. Estimation procedures in inadequate models: comparisons and empirical two step procedure. Mathematische Operationsforschung und Statistik, 6(2):167– 177, 1975. CODEN MTOSBI. ISSN 0047-6277.

Bunke:1980:ARN

[BS80] H. Bunke and W. H. Schmidt. Asymptotic results on nonlinear approximation of regression functions and weighted least squares. Statistics (Berlin, DDR), 11(1):3–22, 1980. CODEN MOSSD5. ISSN 0323-3944.

Bhansali:1982:EPH

[BS82] R. J. Bhansali and S. H. Sarhan. Effects of the presence of a harmonic term on the spectral factorisation procedure. Statistics (Berlin, DDR), 13(2):317–349, 1982. CODEN MOSSD5. ISSN 0323-3944.

Bandemer:1985:EMR

[BS85] H. Bandemer and U. Schulze. On estimation in multiphase regression models with several regressors using prior knowledge. Statistics: a Journal of Theoretical and Applied Statistics, 16 (1):3–13, 1985. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Bellach:1987:BRb

[BS87] R. Bellach and A. Stereller. Book reviews. Statistics: a Journal of Theoretical and Applied Statistics, 18(1):166–167, 1987. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Bock:1988:TDQ

[BS88] Mary Ellen Bock and Herbert Solomon. Tables for distributions of quadratic forms. Statistics: a Journal of Theoretical and Applied Statistics, 19(3):399–435, 1988. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Burgess:2006:OSC

[BS06] Leonie Burgess and Deborah J. Street. The optimal size of choice sets in choice experiments. Statistics: a Journal of Theoretical and Applied Statistics, 40(6):507–515, 2006. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Bensic:2007:BET

[BS07a] Mirta Bensić and Kristian Sabo. Border estimation of a twodimensional uniform distribution if data are measured with additive error. Statistics: a Journal of Theoretical and Applied Statistics, 41(4):311–319, 2007. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Bodnar:2007:DSV

[BS07b] Taras Bodnar and Wolfgang Schmid. The distribution of the sample variance of the global minimum variance portfolio in elliptical models. Statistics: a Journal of Theoretical and Applied Statistics, 41(1):65–75, 2007. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Bensic:2010:EUD

[BS10] Mirta Bensić and Kristian Sabo. Estimating a uniform distribution when data are measured with a normal additive error with unknown variance. Statistics: a Journal of Theoretical and Applied Statistics, 44(3):235–246, 2010. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Bairamov:2013:NON

[BS13] I. Bairamov and A. Stepanov. Number of observations near the maximum in an F^{α}-scheme. Statistics: a Journal of Theoretical and Applied Statistics, 47(1):191–201, 2013. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Bodnar:2017:CCS

[BS17] Olha Bodnar and Wolfgang Schmid. CUSUM control schemes for monitoring the covariance matrix of multivariate time series. Statistics: a Journal of Theoretical and Applied Statistics, 51(4):722–744, 2017. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Bieniek:2018:SBM

[BS18] Mariusz Bieniek and Maria Szpak. Sharp bounds for the mean of the total time on test for distributions with increasing generalized failure rate. Statistics: a Journal of Theoretical and Applied Statistics, 52(4):818–828, 2018. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Berred:2021:APU

[BS21] Alexandre Berred and Alexei Stepanov. Asymptotic properties of the uppermost and the lowest m th exponential spacings based on records. Statistics: a Journal of Theoretical and Applied Statistics, 55(5):1094–1106, 2021. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Berred:2023:APS

[BS23] Alexandre Berred and Alexei Stepanov. On asymptotic properties of spacings. Statistics: a Journal of Theoretical and Applied Statistics, 57(5):1267–1283, 2023. CO-DEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/02331888.2023.2260037.

Barreto-Souza:2013:NLM

[BSB13] Wagner Barreto-Souza and Hassan S. Bakouch. A new lifetime model with decreasing failure rate. Statistics: a Journal of Theoretical and Applied Statistics, 47(2):465–476, 2013. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Bandemer:1978:BR

[BSF78] H. Bandemer, P. SchÖnfeld, and K. Fleischmann. Book review. Statistics (Berlin, DDR), 9(3):479–480, 1978. CODEN MOSSD5. ISSN 0323-3944.

Barreto-Souza:2013:BKD

[BSL13] Wagner Barreto-Souza and Artur J. Lemonte. Bivariate Kumaraswamy distribution: properties and a new method to generate bivariate classes. Statistics: a Journal of Theoretical and Applied Statistics, 47(6):1321–1342, 2013. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Bhatia:1976:TMS

[BSP76] H. L. Bhatia, Kanti Swarup, and M. C. Puri. Time minimizing solid transportation problem. Mathematische Operationsforschung und Statistik, 7(3):395–403, 1976. CODEN MTOSBI. ISSN 0047-6277.

Basu:2018:BIU

[BSS18] Suparna Basu, Sanjay K. Singh, and Umesh Singh. Bayesian inference using product of spacings function for progressive hybrid Type-I censoring scheme. Statistics: a Journal of Theoretical and Applied Statistics, 52(2):345–363, 2018. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Bunke:1981:BR

[BSW^{+}81] O. Bunke, W. Schmidt, W. Winkler, G. Christoph, and H. P. Höschel. Book review. Statistics (Berlin, DDR), 12(3):457– 460, 1981. CODEN MOSSD5. ISSN 0323-3944.

Bertin:1987:SHL

[BT87] Emile M. J. Bertin and Radu Theodoresch. The singlehumpedness of likelihood function revisited. Statistics: a Journal of Theoretical and Applied Statistics, 18(4):629–636, 1987. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Bian:1997:BIB

[BT97] Guorui Bian and Moti L. Tiku. Bayesian inference based on robust priors and MML estimators: Part I, Symmetric location-scale distributions. Statistics: a Journal of Theoretical and Applied Statistics, 29(4):317–345, 1997. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Bellanger:2004:THT

[BT04] Lise Bellanger and Richard Tomassone. Trend in high tropospheric ozone levels. Application to Paris monitoring sites. Statistics: a Journal of Theoretical and Applied Statistics, 38(3):217–241, 2004. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Bairamov:2008:DEG

[BT08] Ismihan Bairamov and Halil Tanil. Distributions of exceedances of generalized order statistics. Statistics: a Journal of Theoretical and Applied Statistics, 42(1):67–76, 2008. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Balakrishnan:2016:CBL

[BT16] Narayanaswamy Balakrishnan and Nuria Torrado. Comparisons between largest order statistics from multiple-outlier models. Statistics: a Journal of Theoretical and Applied Statistics, 50(1):176–189, 2016. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic). URL http://

www.tandfonline.com/doi/abs/10.1080/02331888.2015.
1038268.

Beer:1971:MZL

[BTS71] K. Beer, R. Tichatschke, and B. Schwartz. Mehrfache Zerlegung linearer Optimierungsaufgaben. (German) [Multiple linear cutting optimization tasks]. Mathematische Operationsforschung und Statistik, 2(2):95–115, 1971. CODEN MTOSBI. ISSN 0047-6277.

Bertin:2011:MLE

[BTT11] Karine Bertin, Soledad Torres, and Ciprian A. Tudor. Maximum-likelihood estimators and random walks in long memory models. Statistics: a Journal of Theoretical and Applied Statistics, 45(4):361–374, 2011. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Bertin:2021:LSE

[BTV21] Karine Bertin, Soledad Torres, and Lauri Viitasaari. Leastsquare estimators in linear regression models under negatively superadditive dependent random observations. Statistics: a Journal of Theoretical and Applied Statistics, 55(5): 1018–1034, 2021. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Bunke:1970:GGF

[Bun70] Olaf Bunke. Geleitwort. (German) [Foreword]. Mathematische Operationsforschung und Statistik, 1(1):1–4, 1970. CO-DEN MTOSBI. ISSN 0047-6277.

Bunke:1973:ARF

[Bun73a] Helga Bunke. Approximation of regression functions. Mathematische Operationsforschung und Statistik, 4(4):315–325, 1973. CODEN MTOSBI. ISSN 0047-6277.

Bunke:1973:MCP

[Bun73b] Olaf Bunke. Model choice and parameter estimation in regression analysis. Mathematische Operationsforschung und Statistik, 4(5):407–423, 1973. CODEN MTOSBI. ISSN 0047-6277.

Bunke:1974:RGB

[Bun74] O. Bunke. Rezensionen. (German) [Book reviews]. Mathematische Operationsforschung und Statistik, 5(3):255–256, 1974. CODEN MTOSBI. ISSN 0047-6277.

Bunke:1975:SIF

[Bun75a] H. Bunke. Statistical inference: fiducial and structural vs. likelihood. Mathematische Operationsforschung und Statistik, 6(5):667–676, 1975. CODEN MTOSBI. ISSN 0047-6277.

Bunke:1975:IIL

[Bun75b] O. Bunke. Improved inference in linear models with additional information. Mathematische Operationsforschung und Statistik, 6(5):817–829, 1975. CODEN MTOSBI. ISSN 0047-6277.

Bunke:1975:LSE

[Bun75c] O. Bunke. Least squares estimators as robust and minimax estimators. Mathematische Operationsforschung und Statistik, 6(5):687–688, 1975. CODEN MTOSBI. ISSN 0047-6277.

Bunke:1975:MLR

[Bun75d] O. Bunke. Minimax linear, ridge and shrunken estimators for linear parameters. Mathematische Operationsforschung und Statistik, 6(5):697–701, 1975. CODEN MTOSBI. ISSN 0047-6277.

Bunke:1976:SCE

[Bun76a] Helga Bunke. Simple consistent estimation in nonlinear regression by data transformation and design of experiments. Mathematische Operationsforschung und Statistik, 7(5):715– 719, 1976. CODEN MTOSBI. ISSN 0047-6277.

Bunke:1976:CPI

[Bun76b] O. Bunke. Conditional probability in incompletely specified stochastic equations and statistical inference. Mathematische Operationsforschung und Statistik, 7(5):673–678, 1976. CO-DEN MTOSBI. ISSN 0047-6277.

Bunke:1977:PSI

[Bun77a] J. Bunke. Pivot sufficiency and an interpretation of the invariance principle. Statistics (Berlin, DDR), 8(3):299–304, 1977. CODEN MOSSD5. ISSN 0323-3944.

Bunke:1977:MME

[Bun77b] O. Bunke. Mixed models, empirical Bayes and Stein estimators. Statistics (Berlin, DDR), 8(1):55–68, 1977. CODEN MOSSD5. ISSN 0323-3944.

Bunke:1977:OP

[Bun77c] O. Bunke. On optimal prediction. Statistics (Berlin, DDR), 8(4):453–455, 1977. CODEN MOSSD5. ISSN 0323-3944.

Bunke:1981:NPE

[Bun81a] H. Bunke. A note on parameter estimation in inadequate nonlinear regression models. Statistics (Berlin, DDR), 12(1): 7–11, 1981. CODEN MOSSD5. ISSN 0323-3944.

Bunke:1981:HEH

[Bun81b] O. Bunke. On the homoscedastic estimation of heteroscedastie variances. Statistics (Berlin, DDR), 12(2):179–181, 1981. CO-DEN MOSSD5. ISSN 0323-3944.

Bunke:1982:BRa

[Bun82a] O. Bunke. Book review. Statistics (Berlin, DDR), 13(3):386, 1982. CODEN MOSSD5. ISSN 0323-3944.

Bunke:1982:BRc

[Bun82b] O. Bunke. Book review. Statistics (Berlin, DDR), 13(4):554, 1982. CODEN MOSSD5. ISSN 0323-3944.

Bunke:1982:RTR

[Bun82c] Olaf Bunke. The roles of transformations and response functions chosen by maximum likelihood. Statistics (Berlin, DDR), 13(3):395–404, 1982. CODEN MOSSD5. ISSN 0323-

Bunke:1984:BR

[Bun84] O. Bunke. Book reviews. Statistics (Berlin, DDR), 15(2): 299–319, 1984. CODEN MOSSD5. ISSN 0323-3944.

Bunke:1985:ASE

[Bun85] O. Bunke. An adaptive smoothing estimator lor probabilities in contingency tables. Statistics: a Journal of Theoretical and Applied Statistics, 16(1):55–62, 1985. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Bunke:1987:BRa

[Bun87a] O. Bunke. Book review. Statistics: a Journal of Theoretical and Applied Statistics, 18(3):332, 1987. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Bunke:1987:BRb

[Bun87b] O. Bunke. Book review. Statistics: a Journal of Theoretical and Applied Statistics, 18(3):350, 1987. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Bunke:1988:BR

[Bun88] O. Bunke. Book review. Statistics: a Journal of Theoretical and Applied Statistics, 19(3):436, 1988. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Bunke:1990:BR

[Bun90] O. Bunke. Book rivew. Statistics: a Journal of Theoretical and Applied Statistics, 21(4):532, 1990. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Bunke:1992:E

[Bun92] Olaf Bunke. Editorial. Statistics: a Journal of Theoretical and Applied Statistics, 23(1):1–3, 1992. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Bunke:1993:EJE

[Bun93] O. Bunke. Extended jackknife estimates in linear or nonlinear regression. Statistics: a Journal of Theoretical and Applied Statistics, 25(1):47–61, 1993. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Bunke:1999:MLQ

[Bun99] Olaf Bunke. Minimax linear and quadratic estimators in semiparametric multivariate regression models. Statistics: a Journal of Theoretical and Applied Statistics, 33(1):1–35, 1999. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Bunke:2005:BEM

[Bun05] Olaf Bunke. Bayes estimates in multivariate semiparametric linear models. Statistics: a Journal of Theoretical and Applied Statistics, 39(6):467–481, 2005. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Burmeister:1970:RGB

[Bur70] H. L. Burmeister. Rezensionen. (German) [Book reviews]. Mathematische Operationsforschung und Statistik, 1(3):245, 1970. CODEN MTOSBI. ISSN 0047-6277.

Burkschat:2008:OLE

[Bur08] M. Burkschat. On optimal linear equivariant estimation under progressive censoring. Statistics: a Journal of Theoretical and Applied Statistics, 42(5):383–392, 2008. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Bischoff:1986:OPL

[BV86] D. Bischoff and E. Voigtlander. An operator proof for Lindeberg theorem in the non-classic from of Zolotarev. Statistics: a Journal of Theoretical and Applied Statistics, 17(4): 609–615, 1986. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Berlinet:2001:NPL

[BV01] Alain Berlinet and Igor Vajda. Nonnegative piecewise linear histograms. Statistics: a Journal of Theoretical and Applied Statistics, 35(4):295–317, 2001. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Berlinet:2011:SRB

[BV11] A. Berlinet and I. Vajda. Selection rules based on divergences. Statistics: a Journal of Theoretical and Applied Statistics, 45(5):479–495, 2011. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Benesova:2016:BUD

[BvET16] Martina Benesová, Bert van Es, and Peter Tegelaar. Bivariate uniform deconvolution. Statistics: a Journal of Theoretical and Applied Statistics, 50(4):812–840, 2016. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Boudou:2006:PBP

[BVP06] Alain Boudou and Sylvie Viguier-Pla. On proximity between PCA in the frequency domain and usual PCA. Statistics: a Journal of Theoretical and Applied Statistics, 40(5):447–464, 2006. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Bandemer:1977:BRa

[BvWT^{+}77] H. Bandemer, L. v. Wolfersdorf, H. Toutenburg, J. Groh, R. Bartoszyński, O. Bunke, H. Láuter, F. Beichelt, G. Enderlein, K. Fischer, W. Randt, W. Schmidt, J. Kleffe, W. Näther, H. J. Rossberg, and D. Stoyan. Book reviews. Statistics (Berlin, DDR), 8(3):407–418, 1977. CODEN MOSSD5. ISSN 0323-3944.

Bunke:1980:PEU

[BW80] O. Bunke and H. Wandl. Projection estimators for unknown covariance matrices in linear models. Statistics (Berlin, DDR), 11(2):175–191, 1980. CODEN MOSSD5. ISSN 0323-3944.

Blume:1981:ADC

[BW81] Wolfgang Blume and Gisela Wittwer. On the asymptotic distribution of covariance estimates of stationary random sequences. Statistics (Berlin, DDR), 12(2):193–199, 1981. CO-DEN MOSSD5. ISSN 0323-3944.

Borovskikh:2014:NGL

[BW14] Yuri V. Borovskikh and N. C. Weber. Non-Gaussian limit distributions for U-statistics based on trimmed and Winsorized samples. Statistics: a Journal of Theoretical and Applied Statistics, 48(3):609–632, 2014. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Bank:1971:RGB

[BWR^{+}71] B. Bank, H. Weinert, R. Reinhardt, L. Bitter, H.-J Roßberg, J. Grimm, H. Toutenburg, and W. Dück. Rezensionen. (German) [Book reviews]. Mathematische Operationsforschung und Statistik, 2(4):311–317, 1971. CODEN MTOSBI. ISSN 0047-6277.

Bodnar:2010:SEF

[BZ10] Taras Bodnar and Taras Zabolotskyy. Sample efficient frontier in multivariate conditionally heteroscedastic elliptical models. Statistics: a Journal of Theoretical and Applied Statistics, 44 (1):1–15, 2010. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Bouzebda:2013:SAE

[BZ13] Salim Bouzebda and Tarek Zari. Strong approximation of empirical copula processes by Gaussian processes. Statistics: a Journal of Theoretical and Applied Statistics, 47(5):1047– 1063, 2013. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Balakrishnan:2014:EUM

[BZ14] N. Balakrishnan and Xiaojun Zhu. On the existence and uniqueness of the maximum likelihood estimates of the parameters of Birnbaum–Saunders distribution based on Type-I, Type-II and hybrid censored samples. Statistics: a Journal of Theoretical and Applied Statistics, 48(5):1013–1032, 2014. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Benavoli:2015:PNI

[BZ15] A. Benavoli and M. Zaffalon. Prior near ignorance for inferences in the k-parameter exponential family. Statistics: a Journal of Theoretical and Applied Statistics, 49(5):1104– 1140, 2015. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Cadre:2001:CEM

[Cad01] Benoît Cadre. Convergent estimators for the l_{1}-median of Banach valued random variable. Statistics: a Journal of Theoretical and Applied Statistics, 35(4):509–521, 2001. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Chatterjee:2013:LBM

[CAK13] K. Chatterjee, P. Angelopoulos, and C. Koukouvinos. A lower bound to the measure of optimality for main effect plans in the general asymmetric factorial experiments. Statistics: a Journal of Theoretical and Applied Statistics, 47(2):405–410, 2013. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Cao:2001:REL

[Cao01] Ricardo Cao. Relative efficiency of local bandwidths in kernel density estimation. Statistics: a Journal of Theoretical and Applied Statistics, 35(2):113–137, 2001. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Caperaa:1973:DOF

[Cap73] Philippe Capéraà. Décomposition orthogonale d’une function suivant un classification hiérarchique et applications. (French) [Orthogonal decomposition of a function following a hierarchical clustering and applications]. Mathematische Operationsforschung und Statistik, 4(3):213–228, 1973. CODEN MTOSBI. ISSN 0047-6277.

Capanu:2019:UAP

[Cap19] Marinela Capanu. A unified approach to proving parametric bootstrap consistency for some goodness-of-fit tests. Statistics: a Journal of Theoretical and Applied Statistics, 53(1): 58–80, 2019. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Carmichael:1984:CAD

[Car84] J. P. Carmichael. Consistency of an autorsgressive density estimator. Statistics (Berlin, DDR), 15(3):383–387, 1984. CO-DEN MOSSD5. ISSN 0323-3944.

Carlstein:1988:LLN

[Car88] E. Carlstein. Law of Large Numbers for the subseries values of a statistic from a stationary sequence. Statistics: a Journal of Theoretical and Applied Statistics, 19(2):295–299, 1988. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Caron:2019:ADL

[Car19] Emmanuel Caron. Asymptotic distribution of least square estimators for linear models with dependent errors. Statistics: a Journal of Theoretical and Applied Statistics, 53(4):885– 902, 2019. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Causeur:1999:EDR

[Cau99] David Causeur. Exact distribution of the regression esti mator in double sampling. Statistics: a Journal of Theoretical and Applied Statistics, 32(4):297–315, 1999. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Cavalier:1997:NER

[Cav97] Laurent Cavalier. Nonparametric estimation of regression level sets. Statistics: a Journal of Theoretical and Applied Statistics, 29(2):131–160, 1997. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Chandrasekar:2002:SPC

[CB02a] B. Chandrasekar and N. Balakrishnan. Some properties and a characterization of trivariate and multivariate binomial distributions. Statistics: a Journal of Theoretical and Applied Statistics, 36(3):211–218, 2002. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Childs:2002:CIP

[CB02b] Aaron Childs and N. Balakrishnan. Conditional inference procedures for the Pareto and power function distributions based on Type-II right censored samples. Statistics: a Journal of Theoretical and Applied Statistics, 36(3):247–257, 2002. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Chandrasekar:2004:PBP

[CB04] B. Chandrasekar and N. Balakrishnan. Property of bivariate Poisson distribution and its application to stochastic processes. Statistics: a Journal of Theoretical and Applied Statistics, 38(2):161–165, 2004. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Childs:2012:EDM

[CBC12] A. Childs, N. Balakrishnan, and B. Chandrasekar. Exact distribution of the MLEs of the parameters and of the quantiles of two-parameter exponential distribution under hybrid censoring. Statistics: a Journal of Theoretical and Applied Statistics, 46(4):441–458, 2012. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Clemencon:2017:SER

[CBC17] Stephan Clémençon, Patrice Bertail, and Emilie Chautru. Sampling and empirical risk minimization. Statistics: a Journal of Theoretical and Applied Statistics, 51(1):30–42, 2017. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Chakraborty:2023:PTD

[CBF23] Niladri Chakraborty, Narayanaswamy Balakrishnan, and Maxim Finkelstein. On precedence tests with double sampling. Statistics: a Journal of Theoretical and Applied Statistics, 57(3):554–576, 2023. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/02331888.2023.2203491.

Chakraborty:2023:CEP

[CBP23] Siddhartha Chakraborty, Ritwik Bhattacharya, and Biswabrata Pradhan. Cumulative entropy of progressively type-II censored order statistics and associated optimal life testingplans. Statistics: a Journal of Theoretical and Applied Statistics, 57(1):161–174, 2023. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/02331888.2023.2168666.

Cabana:2001:GFT

[CC01] A. Cabaña and E. M. Cabaña. Goodness-of-fit tests based on quadratic functionals of transformed empirical processes. Statistics: a Journal of Theoretical and Applied Statistics, 35(2):171–189, 2001. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Chen:2012:ELI

[CC12] Xia Chen and Hengjian Cui. Empirical likelihood inference for parameters in a partially linear errors-in-variables model. Statistics: a Journal of Theoretical and Applied Statistics, 46(6):745–757, 2012. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Cerqueti:2013:NRC

[CC13] Roy Cerqueti and Mauro Costantini. New results on the convergence of random matrices. Statistics: a Journal of Theoretical and Applied Statistics, 47(4):663–671, 2013. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Chaubey:2015:AWE

[CCD15] Yogendra P. Chaubey, Christophe Chesneau, and Hassan Doosti. Adaptive wavelet estimation of a density from mixtures under multiplicative censoring. Statistics: a Journal of Theoretical and Applied Statistics, 49(3):638–659, 2015. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Csorgo:1987:ETT

[CCH87] Miklós Csörgő, Sándor Csörgő, and Lajos Horváth. Estimation of total time on test transforms and Lorenz curves under random censorship. Statistics: a Journal of Theoretical and Applied Statistics, 18(1):77–97, 1987. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Comas:2018:SOT

[CCM18] C. Comas, J. Conde, and J. Mateu. A second-order test to detect spatio-temporal anisotropic effects in point patterns. Statistics: a Journal of Theoretical and Applied Statistics, 52(4):717–733, 2018. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Cordeiro:2013:BGG

[CCMdC13] Gauss M. Cordeiro, Fredy Castellares, Lourdes C. Montenegro, and Mário de Castro. The beta generalized gamma distribution. Statistics: a Journal of Theoretical and Applied Statistics, 47(4):888–900, 2013. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Contreras-Cristan:2009:CSA

[CCMW09] Alberto Contreras-Cristán, Ramsés H. Mena, and Stephen G. Walker. On the construction of stationary AR(1) models via random distributions. Statistics: a Journal of Theoretical and Applied Statistics, 43(3):227–240, 2009. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Ciuperca:2008:MLE

[CD08] Gabriela Ciuperca and Nicolas Dapzol. Maximum likelihood estimator in a multi-phase random regression model. Statistics: a Journal of Theoretical and Applied Statistics, 42(4): 363–381, 2008. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Cheng:2009:BRN

[CD09] Yebin Cheng and Jan G. De Gooijer. Bahadur representation for the nonparametric M-estimator under α-mixing dependence. Statistics: a Journal of Theoretical and Applied Statistics, 43(5):443–462, 2009. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Chesneau:2017:DPU

[CDD17] Christophe Chesneau, Isha Dewan, and Hassan Doosti. On a deconvolution problem under competing risks. Statistics: a Journal of Theoretical and Applied Statistics, 51(2):331–346, 2017. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Cencov:1978:AFM

[Cen78] N. N. Cencov. Algebraic foundation of mathematical statistics. Statistics (Berlin, DDR), 9(2):267–276, 1978. CODEN MOSSD5. ISSN 0323-3944.

Cetinkaya:2021:RES

[Çet21]  Çagatay Çetinkaya. Reliability estimation of a stress-strength

model with non-identical component strengths under generalized progressive hybrid censoring scheme. Statistics: a Journal of Theoretical and Applied Statistics, 55(2):250–275, 2021. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Ceyhan:2020:DNI

[Cey20] Elvan Ceyhan. Domination number of an interval catch digraph family and its use for testing uniformity. Statistics: a Journal of Theoretical and Applied Statistics, 54(2):310–339, 2020. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Chepurin:1978:IFB

[CF78] E. V. Chepurin and P. Fehrmann. On inequalities for functionals of bivariate distributions with monotone failure rates. Statistics (Berlin, DDR), 9(3):395–404, 1978. CODEN MOSSD5. ISSN 0323-3944.

Consul:1989:MVU

[CF89] P. C. Consul and Famoye Felix. Minimum variance unbiased estimation for the Lagrange power series distributions. Statistics: a Journal of Theoretical and Applied Statistics, 20(3): 407–415, 1989. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Cook:1995:IDP

[CF95a] Dennis Cook and Valery Fedorov. Invited discussion paper: Constrained optimization of experimental design. Statistics: a Journal of Theoretical and Applied Statistics, 26(2):129–148, 1995. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Cook:1995:R

[CF95b] Dennis Cook and Valery Fedorov. Rejoinder. Statistics: a Journal of Theoretical and Applied Statistics, 26(2):175–178, 1995. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Csorgo:1998:PNP

[CF98] Sándor Csörgő and Julian J. Faraway. The paradoxical nature of the proportional hazards model of random censorship. Statistics: a Journal of Theoretical and Applied Statistics, 31 (1):67–78, 1998. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Cha:2021:ARS

[CF21] Ji Hwan Cha and Maxim Finkelstein. Acceptance reliability sampling plan for discrete lifetime models. Statistics: a Journal of Theoretical and Applied Statistics, 55(6):1291–1309, 2021. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Cha:2022:ARS

[CF22] Ji Hwan Cha and Maxim Finkelstein. Acceptance reliability sampling plan for items with two failure modes. Statistics: a Journal of Theoretical and Applied Statistics, 56(6):1345– 1363, 2022. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-

4910 (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/02331888.2022.2154354.

Crujeiras:2010:ESD

[CFC10] Rosa M. Crujeiras and Rubén Fernández-Casal. On the estimation of the spectral density for continuous spatial processes. Statistics: a Journal of Theoretical and Applied Statistics, 44(6):587–600, 2010. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Cascos-Fernandez:2004:CCI

[CFLDÁ04] Ignacio Cascos-Fernández, Miguel López-Díaz, and Maria Ángeles Gil. Convergence criteria for interval-valued inequality indices. Statistics: a Journal of Theoretical and Applied Statistics, 38(1):59–66, 2004. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Cellieb:1989:CSE

[CFR89] Dominique Cellieb, Dominique Fouedrinier, and Christian Robert. Controlled shrinkage estimators (a class of estimators better than the least squares estimator, with respect to a general quadratic loss, for normal observations). Statistics: a Journal of Theoretical and Applied Statistics, 20(1):13–22, 1989. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Chang:2017:ETO

[CFS17] Yuan-Tsung Chang, Kazufumi Fukuda, and Nobuo Shinozaki. Estimation of two ordered normal means when a covariance matrix is known. Statistics: a Journal of Theoretical and Applied Statistics, 51(5):1095–1104, 2017. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Chaturvedi:1995:SOA

[CG95] Ajit Chaturvedi and Rahul Gupta. Second-order approximations to two classes of sequential estimation procedures. Statistics: a Journal of Theoretical and Applied Statistics, 27 (1–2):159–173, 1995. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Chen:2004:SIC

[CG04] Jie Chen and A. K. Gupta. Statistical inference of covariance change points in Gaussian model. Statistics: a Journal of Theoretical and Applied Statistics, 38(1):17–28, 2004. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Chen:2005:MVS

[CG05] John T. Chen and Arjun K. Gupta. Matrix variate skew normal distributions. Statistics: a Journal of Theoretical and Applied Statistics, 39(3):247–253, 2005. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Chatterjee:2017:ODN

[CGK17] K. Chatterjee, S. D. Georgiou, and C. Koukouvinos. A_{2}- optimal designs: the nearly-balanced case. Statistics: a Journal of Theoretical and Applied Statistics, 51(2):235–246, 2017. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Coen:2019:MFT

[CGM19] Arrigo Coen, Luis Gutiérrez, and Ramsés H. Mena. Modelling failures times with dependent renewal type models via exchangeability. Statistics: a Journal of Theoretical and Applied Statistics, 53(5):1112–1130, 2019. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Chaudhary:2023:GWC

[CGS23] Santosh Kumar Chaudhary, Nitin Gupta, and Pradeep Kumar Sahu. On general weighted cumulative residual extropy and general weighted negative cumulative extropy. Statistics: a Journal of Theoretical and Applied Statistics, 57(5): 1117–1141, 2023. CODEN MOSSD5. ISSN 0233-1888 (print),

1029-4910 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/02331888.2023.2241595.

Carroll:1989:SWL

[CH89] R. J. Carroll and W. Härdle. Semiparametric weighted least squares. Statistics: a Journal of Theoretical and Applied Statistics, 20(1):173–175, 1989. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Cline:1991:KED

[CH91] Daren B. H. Cline and Jeffrey D. Hart. Kernel estimation of densities with discontinuities or discontinuous derivatives. Statistics: a Journal of Theoretical and Applied Statistics, 22 (1):69–84, 1991. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Causeur:2008:PDS

[CH08] David Causeur and François Husson. Power of doublesampling tests for general linear hypotheses. Statistics: a Journal of Theoretical and Applied Statistics, 42(2):115–125, 2008. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Chaudhry:1974:TSS

[Cha74] M. L. Chaudhry. Transient/steady-state solution of a single channel queue with bulk arrivals and intermittently available server. Mathematische Operationsforschung und Statistik, 5 (4–5):307–314, 1974. CODEN MTOSBI. ISSN 0047-6277.

Chattopadhyay:1996:NDO

[Cha96] Manas K. Chattopadhyay. A note on Dirichlet one-armed bandits. Statistics: a Journal of Theoretical and Applied Statistics, 27(3–4):395–402, 1996. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Cha:2015:VAR

[Cha15a] Ji Hwan Cha. Variables acceptance reliability sampling plan for repairable items. Statistics: a Journal of Theoretical and Applied Statistics, 49(5):1141–1156, 2015. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Chakraborty:2015:MSE

[CHA15b] Subrata Chakraborty, Partha Jyoti Hazarika, and M. Masoom Ali. A multimodal skewed extension of normal distribution: its properties and applications. Statistics: a Journal of Theoretical and Applied Statistics, 49(4):859–877, 2015. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Chen:1997:AAC

[Che97] Hanfeng Chen. Asymptotic analysis of a class of process capability indices. Statistics: a Journal of Theoretical and Applied Statistics, 30(2):149–162, 1997. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Chen:2012:APE

[Che12] Xia Chen. Asymptotic properties for estimates of nonparametric regression model with martingale difference errors. Statistics: a Journal of Theoretical and Applied Statistics, 46(5):687–696, 2012. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Cheng:2021:SUC

[Che21] Fuxia Cheng. The strong uniform consistency of kernel estimator of a smooth distribution function in censored linear regression. Statistics: a Journal of Theoretical and Applied Statistics, 55(1):110–119, 2021. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Chen:2009:KED

[CHL09] Sy-Mien Chen, Yu-Sheng Hsu, and Jian-Tong Liaw. On kernel estimators of density ratio. Statistics: a Journal of Theoretical and Applied Statistics, 43(5):463–479, 2009. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Cordeiro:2014:MWM

[CHO14] Gauss M. Cordeiro, Elizabeth M. Hashimoto, and Edwin M. M. Ortega. The McDonald Weibull model. Statistics: a Journal of Theoretical and Applied Statistics, 48(2):256–278, 2014. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Chen:2003:CMD

[CHP03] Sy-Mien Chen, Yu-Sheng Hsu, and W. L. Pearn. Capability measures for m-dependent stationary processes. Statistics: a Journal of Theoretical and Applied Statistics, 37(2):145–168, 2003. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Cernikova:2013:DTM

[CHPS13] Alena Cerníková, Marie Husková, Zuzana Prásková, and Josef G. Steinebach. Delay time in monitoring jump changes in linear models. Statistics: a Journal of Theoretical and Applied Statistics, 47(1):1–25, 2013. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Christou:2021:TCQ

[Chr21] Eliana Christou. Transformed central quantile subspace. Statistics: a Journal of Theoretical and Applied Statistics, 55(2):350–366, 2021. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Christou:2023:DRT

[Chr23] Eliana Christou. Dimension reduction techniques for conditional expectiles. Statistics: a Journal of Theoretical and Applied Statistics, 57(4):960–985, 2023. CO-DEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/02331888.2023.2236745.

Cierco:1998:ADM

[Cie98] Christine Cierco. Asymptotic distribution of the maximum likelihood ratio test for gene detection. Statistics: a Journal of Theoretical and Applied Statistics, 31(3):261–285, 1998. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Ciginas:2013:SOA

[Cig13] Andrius Ciginas. Second-order approximations of finite population L-statistics. Statistics: a Journal of Theoretical and Applied Statistics, 47(5):954–965, 2013. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Cipra:1981:IPA

[Cip81] Tomás Cipra. On improvement of prediction in ARMA processes. Statistics (Berlin, DDR), 12(4):567–580, 1981. CO-DEN MOSSD5. ISSN 0323-3944.

Cipra:1984:SCA

[Cip84] Tomas Cipra. Simple correlated ARMA processes. Statistics (Berlin, DDR), 15(4):513–524, 1984. CODEN MOSSD5. ISSN 0323-3944.

Cipra:1991:TPM

[Cip91] T. Cipra. Tests of periodicity with missing observations. Statistics: a Journal of Theoretical and Applied Statistics, 22(2):233–243, 1991. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Ciuperca:2012:ECP

[Ciu12] Gabriela Ciuperca. The S-estimator in the change-point random model with long memory. Statistics: a Journal of Theoretical and Applied Statistics, 46(6):697–718, 2012. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Ciuperca:2016:ALM

[Ciu16] Gabriela Ciuperca. Adaptive LASSO model selection in a multiphase quantile regression. Statistics: a Journal of Theoretical and Applied Statistics, 50(5):1100–??, 2016. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Ciuperca:2020:AEN

[Ciu20] Gabriela Ciuperca. Adaptive elastic-net selection in a quantile model with diverging number of variable groups. Statistics: a Journal of Theoretical and Applied Statistics, 54(5):1147– 1170, 2020. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Choi:2006:TGF

[CK06] Byungjin Choi and Keeyoung Kim. Testing goodness-of-fit for Laplace distribution based on maximum entropy. Statistics: a Journal of Theoretical and Applied Statistics, 40(6):517–531, 2006. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Chang:2009:BFC

[CKM09] In Hong Chang, Byung Hwee Kim, and Rahul Mukerjee. Bayesian and frequentist confidence intervals via adjusted likelihoods under prior specification on the interest parameter. Statistics: a Journal of Theoretical and Applied Statistics, 43(2):203–211, 2009. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Chaudhuri:2022:CBM

[CKS22] Sanjay Chaudhuri, Tatsuya Kubokawa, and Shonosuke Sugasawa. Covariance based moment equations for improved variance component estimation. Statistics: a Journal of Theoretical and Applied Statistics, 56(6):1290–1318, 2022. CO-DEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/02331888.2022.2144856.

Cho:2012:EMA

[CL12] Jaerin Cho and Jung-Jo Lee. On the G-efficiency of the modified arcsin design. Statistics: a Journal of Theoretical and Applied Statistics, 46(3):343–350, 2012. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Cordeiro:2014:MAD

[CL14] Gauss M. Cordeiro and Artur J. Lemonte. The McDonald arcsine distribution: a new model to proportional data. Statistics: a Journal of Theoretical and Applied Statistics, 48(1): 182–199, 2014. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Cattiaux:2017:OSI

[CLCP17] P. Cattiaux, J. R. León, A. A. Pineda Centeno, and C. Prieur. An overlook on statistical inference issues for stochastic damping Hamiltonian systems under the fluctuation–dissipation condition. Statistics: a Journal of Theoretical and Applied Statistics, 51(1):11–29, 2017. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Caro-Lopera:2016:DMA

[CLDG16] Francisco J. Caro-Lopera and José A. Díaz-García. Diagonalization matrix and its application in distribution theory. Statistics: a Journal of Theoretical and Applied Statistics, 50(4):870–880, 2016. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Cheng:2022:MAG

[CLLY22] Weili Cheng, Xiaorui Li, Xiaoxia Li, and Xiaodong Yan. Model averaging for generalized linear models with missing at random covariates. Statistics: a Journal of Theoretical and Applied Statistics, 57(1):26–52, 2022. CO-DEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/02331888.2022.2161094.

Cordeiro:2013:EFL

[CLO13] Gauss M. Cordeiro, Artur J. Lemonte, and Edwin M. M. Ortega. An extended fatigue life distribution. Statistics: a Journal of Theoretical and Applied Statistics, 47(3):626–653, 2013. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Cao:2014:DNM

[CLS14] Chun-Zheng Cao, Jin-Guan Lin, and Jian-Qing Shi. Diagnostics on nonlinear model with scale mixtures of skew-normal and first-order autoregressive errors. Statistics: a Journal of Theoretical and Applied Statistics, 48(5):1033–1047, 2014. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Choi:2020:SCE

[CLWK20] Hyejeong Choi, Johan Lim, Xinlei Wang, and Minjung Kwak. A self-consistent estimator for interval-valued data. Statistics: a Journal of Theoretical and Applied Statistics, 54(5):1005– 1029, 2020. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Chan:2015:MPT

[CLZ15] Chia-Hao Chan, Huimei Liu, and Mei-Mei Zen. More powerful tests for the sign testing problem about gamma scale parameters. Statistics: a Journal of Theoretical and Applied Statistics, 49(3):564–577, 2015. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Cocchi:1990:LBE

[CM90] D. Cocchi and M. Motjchart. Linear Bayes estimation in finite populations with a categorical auxiliary variable. Statistics: a Journal of Theoretical and Applied Statistics, 21(3):437–454, 1990. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Chen:2005:ECR

[CM05] Zhenmin Chen and Jie Mi. Exact confidence region for the parameters of the gamma distribution based on two independent samples. Statistics: a Journal of Theoretical and Applied Statistics, 39(6):519–527, 2005. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Chang:2013:DDP

[CM13] In Hong Chang and Rahul Mukerjee. Data-dependent probability matching priors for likelihood ratio and adjusted likelihood ratio statistics. Statistics: a Journal of Theoretical and Applied Statistics, 47(2):294–305, 2013. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Chang:2015:PSA

[CM15a] In Hong Chang and Rahul Mukerjee. Predictive sets with approximate frequentist and Bayesian validity for arbitrary priors. Statistics: a Journal of Theoretical and Applied Statistics, 49(5):1095–1103, 2015. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Chattopadhyay:2015:CPS

[CM15b] Gopaldeb Chattopadhyay and Indranil Mukhopadhyay. On a class of partially sequential two-sample test procedures for multivariate continuous data. Statistics: a Journal of Theoretical and Applied Statistics, 49(2):455–473, 2015. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Castano-Martinez:2012:RTC

[CMLBSM12] A. Castaño-Martínez, F. López-Blázquez, and B. Salamanca-Miño. Random translations, contractions and dilations of order statistics and records. Statistics: a Journal of Theoretical and Applied Statistics, 46(1):57–67, 2012. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Castano-Martinez:2014:RSS

[CMLBSM14] Antonia Castaño-Martínez, Fernando López-Blázquez, and Begoña Salamanca-Miño. Random shifts and scalings of mgeneralized order statistics. Statistics: a Journal of Theoretical and Applied Statistics, 48(3):558–574, 2014. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Cagin:2017:SLA

[ÇO17] Tonguç Çagin and Paulo Eduardo Oliveira. Strong laws for associated random variables. Statistics: a Journal of Theoretical and Applied Statistics, 51(6):1259–1279, 2017. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Collomb:1985:NPT

[Col85a] Gérard Collomb. Non parametric time series analysis and prediction: uniform almost sure convergence of the window and K–NN autoregression estimates. Statistics: a Journal of Theoretical and Applied Statistics, 16(2):297–307, 1985. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Collomb:1985:NRD

[Col85b] Gérard Collomb. Nonparametric regression: An up-to-date bibliography. Statistics: a Journal of Theoretical and Applied Statistics, 16(2):309–324, 1985. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Collins:1991:MAB

[Col91] John R. Collins. Maximal asymptotic biases and variances of symmetrized interquantile ranges under asymmetric contamination. Statistics: a Journal of Theoretical and Applied Statistics, 22(3):379–402, 1991. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Collins:2003:BRE

[Col03] John R. Collins. Bias-robust L-estimators of a scale parameter. Statistics: a Journal of Theoretical and Applied Statistics, 37(4):287–304, 2003. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Chatterjee:2021:OTL

[COQ21] Kashinath Chatterjee, Zujun Ou, and Hong Qin. R-optimal two-level supersaturated designs. Statistics: a Journal of Theoretical and Applied Statistics, 55(2):458–473, 2021. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Cornec:2017:CIC

[Cor17] Matthieu Cornec. Concentration inequalities of the crossvalidation estimator for empirical risk minimizer. Statistics: a Journal of Theoretical and Applied Statistics, 51(1):43–60, 2017. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Courrieu:2021:CRO

[Cou21] Pierre Courrieu. On the convergence of rank-one multi-target linear regression. Statistics: a Journal of Theoretical and Applied Statistics, 55(1):68–89, 2021. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Cox:1977:NMR

[Cox77] D. R. Cox. Nonlinear models, residuals and transformations. Statistics (Berlin, DDR), 8(1):3–22, 1977. CODEN MOSSD5. ISSN 0323-3944.

Cao:1998:BR

[CPGB98] Ricardo Cao, O. Pons, Arjun K. Gupta, and O. Bunke. Book review. Statistics: a Journal of Theoretical and Applied Statistics, 32(1):87–88, 1998. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Chaudhuri:1998:ERF

[CPS98] Gopal Chaudhuri, Nabendu Pal, and Jyotirmoy Sarkar. Estimation of the reliability function of a series system: Some decision theoretic results. Statistics: a Journal of Theoretical and Applied Statistics, 32(1):59–74, 1998. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Chaturvedi:1999:CTS

[CR99] Ajit Chaturvedi and Uma Rani. On the classes of two-stage procedures to construct fixed-width confidence regions. Statistics: a Journal of Theoretical and Applied Statistics, 32(4): 341–352, 1999. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Cramer:2006:DSG

[Cra06] Erhard Cramer. Dependence structure of generalized order statistics. Statistics: a Journal of Theoretical and Applied Statistics, 40(5):409–413, 2006. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Cintra:2014:BGN

[CRCN14] R. J. Cintra, L. C. Rêgo, G. M. Cordeiro, and A. D. C. Nascimento. Beta generalized normal distribution with an application for SAR image processing. Statistics: a Journal of Theoretical and Applied Statistics, 48(2):279–294, 2014. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Croux:1996:MDC

[Cro96] Christophe Croux. Maximum deviation curves for location estimators. Statistics: a Journal of Theoretical and Applied Statistics, 28(4):285–305, 1996. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Cunningham:1973:TUS

[CS73] L. Cunningham and N. Singh. A two unit standby redundant system with standby failure and repair waiting time. Mathematische Operationsforschung und Statistik, 4(6):453– 462, 1973. CODEN MTOSBI. ISSN 0047-6277.

Cuevas:1989:CQR

[CS89] Antonio Cuevas and Paloma Sanz. A class of qualitatively robust estimates. Statistics: a Journal of Theoretical and Applied Statistics, 20(4):509–520, 1989. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Chattopadhyay:2006:TSA

[CS06] Saibal Chattopadhyay and Raghu Nandan Sengupta. Threestage and accelerated sequential point estimation of the normal mean using LINEX loss function. Statistics: a Journal of Theoretical and Applied Statistics, 40(1):39–49, 2006. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Chaturvedi:2010:NET

[CS10] Ajit Chaturvedi and Vandana Sharma. A note on the estimation of P (Y > X) in two-parameter exponential distributions. Statistics: a Journal of Theoretical and Applied Statistics, 44 (1):73–75, 2010. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Csenki:1995:MAR

[Cse95] Attila Csenki. Mission availability for repairable semi-Markov systems: Analytical results and computational implementation. Statistics: a Journal of Theoretical and Applied Statistics, 26(1):75–87, 1995. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Csorgo:1988:EPH

[Csö88] Sándor Csörgő. Estimation in the proportional hazards model of random censorship. Statistics: a Journal of Theoretical and Applied Statistics, 19(3):437–463, 1988. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Cordeiro:2013:BWG

[CSO13] Gauss M. Cordeiro, Giovana O. Silva, and Edwin M. M. Ortega. The beta-Weibull geometric distribution. Statistics: a Journal of Theoretical and Applied Statistics, 47(4):817–834, 2013. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Chaudhry:1973:NDB

[CT73] M. L. Chaudhry and J. G. C. Templeton. A note on the distribution of a busy period for M/M/C queuing system. Mathematische Operationsforschung und Statistik, 4(1):75–79, 1973. CODEN MTOSBI. ISSN 0047-6277.

Chaturvedi:2003:TSA

[CT03] Ajit Chaturvedi and Sanjeev K. Tomer. Three-stage and ‘accelerated’ sequential procedures for the mean of a normal population with known coefficient of variation. Statistics: a Journal of Theoretical and Applied Statistics, 37(1):51–64, 2003. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Chen:2007:PRI

[CT07] Chien-Hung Chen and Tsung-Shan Tsou. Parametric robust inference about regression parameters for the correlation coefficient. Statistics: a Journal of Theoretical and Applied Statistics, 41(1):1–9, 2007. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Chen:2014:RTV

[CT14] Zhanshou Chen and Zheng Tian. Ratio tests for variance change in nonparametric regression. Statistics: a Journal of Theoretical and Applied Statistics, 48(1):1–16, 2014. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Chen:2014:BGG

[CTL14] Li-Shya Chen, I-Shiang Tzeng, and Chien-Tai Lin. Bivariate generalized gamma distributions of Kibble’s type. Statistics: a Journal of Theoretical and Applied Statistics, 48(4):933– 949, 2014. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Chen:2009:CRS

[CTW09] Xue-Dong Chen, Nian-Sheng Tang, and Xue-Ren Wang. On confidence regions of semiparametric nonlinear reproductive dispersion models. Statistics: a Journal of Theoretical and Applied Statistics, 43(6):583–595, 2009. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Cuparic:2021:API

[Cup21] Marija Cuparić. Asymptotic properties of inverse probability of censored weighted U-empirical process for right-censored data with applications. Statistics: a Journal of Theoretical and Applied Statistics, 55(5):1035–1057, 2021. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Csaki:1978:LDL

[CV78] E. Csáki and I. Vincze. On limiting distribution laws of statistics analogous to Pearson’s chi-square. Statistics (Berlin, DDR), 9(4):531–548, 1978. CODEN MOSSD5. ISSN 0323-3944.

Chen:2009:TBS

[CWW09] Hubert J. Chen, Miin-Jye Wen, and Andrew Minglong Wang. On testing the bioequivalence of several treatments using the measure of distance. Statistics: a Journal of Theoretical and Applied Statistics, 43(5):513–530, 2009. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Cao:2014:LAE

[CXH14] Mingxiang Cao, Xingzhong Xu, and Daojiang He. Linearly admissible estimators of stochastic regression coefficient under balanced loss function. Statistics: a Journal of Theoretical and Applied Statistics, 48(2):359–366, 2014. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Cai:2020:REM

[CXW20] Xiong Cai, Liugen Xue, and Zhaoliang Wang. Robust estimation with modified Huber’s function for functional linear models. Statistics: a Journal of Theoretical and Applied Statistics, 54(6):1276–1286, 2020. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Chang:2012:PPT

[CY12] Xinfeng Chang and Hu Yang. Performance of the preliminary test two-parameter estimators based on the conflicting test statistics in a regression model with Student’s t error. Statistics: a Journal of Theoretical and Applied Statistics, 46(3): 291–303, 2012. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Cai:2017:HRO

[CZZ17] Xiong Cai, Yiying Zhang, and Peng Zhao. Hazard rate ordering of the second-order statistics from multiple-outlier PHR samples. Statistics: a Journal of Theoretical and Applied Statistics, 51(3):615–626, 2017. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Chen:2019:COS

[CZZ19] Jianbin Chen, Yiying Zhang, and Peng Zhao. Comparisons of order statistics from heterogeneous negative binomial variables with applications. Statistics: a Journal of Theoretical and Applied Statistics, 53(5):990–1011, 2019. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Dembinska:2020:NOR

[DAA20] Anna Dembińska, Masoumeh Akbari, and Jafar Ahmadi. On numbers of observations in random regions determined by records. Statistics: a Journal of Theoretical and Applied Statistics, 54(4):805–829, 2020. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Dabrowska:1981:RBO

[Dab81] D. Dabrowska. Regression-based orderings and measures of stochastic dependence. Statistics (Berlin, DDR), 12(3):317– 325, 1981. CODEN MOSSD5. ISSN 0323-3944.

Dachian:2011:ELT

[Dac11] Sergueï Dachian. Estimation of the location of a 0-type or ∞- type singularity by Poisson observations. Statistics: a Journal of Theoretical and Applied Statistics, 45(5):509–523, 2011. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Dallas:1977:MRS

[Dal77] Agis Dallas. On the minimum of a random sample. Statistics (Berlin, DDR), 8(4):511–513, 1977. CODEN MOSSD5. ISSN 0323-3944.

Dammi:1976:SZS

[Dam76] Heinz Dammi. Über Steuerungsprobleme mit zeitabhängigen Steuerbereichen. (German) [On control problems with timedependent control fields]. Mathematische Operationsforschung und Statistik, 7(2):307–318, 1976. CODEN MTOSBI. ISSN 0047-6277.

Danielak:2003:SUM

[Dan03] Katarzyna Danielak. Sharp upper mean-variance bounds for trimmed means from restricted families. Statistics: a Journal of Theoretical and Applied Statistics, 37(4):305–324, 2003.

CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Das:1987:EOW

[Das87] Kalyan Das. Estimation in one-way classification with heteroscedastic error. Statistics: a Journal of Theoretical and Applied Statistics, 18(2):247–257, 1987. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Das:1996:IEC

[Das96] Kalyan Das. Improved estimation of covariant matrices in balanced hierarchical multivariate variance components models. Statistics: a Journal of Theoretical and Applied Statistics, 28 (1):73–82, 1996. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Daudin:1982:IAN

[Dau82] J. J. Daudin. Iterative algorithms for non-orthogonal analysis of variance. Statistics (Berlin, DDR), 13(3):355–370, 1982. CODEN MOSSD5. ISSN 0323-3944.

David:1994:ODC

[Dav94] O. David. Optimal designs for comparing several treatments with a control in the presence of explanatory variables. Statistics: a Journal of Theoretical and Applied Statistics, 25(4): 325–331, 1994. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Deutler:1974:EOS

[DB74] T. Deutler and W. Bühler. Zur Existenz optimaler Schichtungen. (German) [The existence of optimal stratification]. Mathematische Operationsforschung und Statistik, 5(7–8):637–642, 1974. CODEN MTOSBI. ISSN 0047-6277.

Dominique:1992:BME

[DB92] Picard Dominique and Prum Bernard. The bias of the MLE, an example of the behaviour of different corrections in genetic models. Statistics: a Journal of Theoretical and Applied Statistics, 23(2):159–169, 1992. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Dembinska:2010:AIN

[DB10] Anna Dembińska and N. Balakrishnan. On the asymptotic independence of numbers of observations near order statistics.

Statistics: a Journal of Theoretical and Applied Statistics, 44(5):517–528, 2010. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Doostparast:2013:PAB

[DB13] Mahdi Doostparast and Narayanaswamy Balakrishnan. Pareto analysis based on records. Statistics: a Journal of Theoretical and Applied Statistics, 47(5):1075–1089, 2013. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Diallo:2019:EZI

[DDD19] Alpha Oumar Diallo, Aliou Diop, and Jean-François Dupuy. Estimation in zero-inflated binomial regression with missing covariates. Statistics: a Journal of Theoretical and Applied Statistics, 53(4):839–865, 2019. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Dahlhaus:2017:SIO

[DDLN17] Rainer Dahlhaus, Thierry Dumont, Sylvain Le Corff, and Jan C. Neddermeyer. Statistical inference for oscillation processes. Statistics: a Journal of Theoretical and Applied Statistics, 51(1):61–83, 2017. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Duchesne:2020:SCA

[DdMT20] Pierre Duchesne, Pierre Lafaye de Micheaux, and Joseph François Tagne Tatsinkou. On strong consistency and asymptotic normality of one-step Gauss–Newton estimators in ARMA time series models. Statistics: a Journal of Theoretical and Applied Statistics, 54(5):1030–1057, 2020. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Nkou:2019:SCK

[dDNN19] Emmanuel de Dieu Nkou and Guy Martial Nkiet. Strong consistency of kernel estimator in a semiparametric regression model. Statistics: a Journal of Theoretical and Applied Statistics, 53(6):1289–1305, 2019. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Daudin:1988:SPC

[DDT88] J. J. Daudin, C. Duby, and P. Trecourt. Stability of principal component analysis studied by the bootstrap method. Statistics: a Journal of Theoretical and Applied Statistics, 19(2): 241–258, 1988. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Daudin:1989:PSS

[DDT89] J. J. Daudin, C. Duby, and P. Trécourt. PCA stability studied by the bootstrap and the infinitesimal jackknife method. Statistics: a Journal of Theoretical and Applied Statistics, 20(2):255–270, 1989. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

DeGooijer:2022:MDF

[De 22] Jan G. De Gooijer. The marginal distribution function of threshold-type processes with central symmetric innovations. Statistics: a Journal of Theoretical and Applied Statistics, 56 (1):1–33, 2022. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Debes:1977:CEG

[Deb77] Hermann Debes. Charakterisierungen der exponential — und der geometrischen Verteilung. (German) [Characterizations of the exponential and the geometric distribution]. Statistics (Berlin, DDR), 8(4):515–522, 1977. CODEN MOSSD5. ISSN 0323-3944.

Delsol:2009:AAN

[Del09] Laurent Delsol. Advances on asymptotic normality in nonparametric functional time series analysis. Statistics: a Journal of Theoretical and Applied Statistics, 43(1):13–33, 2009. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Dembinska:2014:ANN

[Dem14] Anna Dembińska. Asymptotic normality of numbers of observations in random regions determined by order statistics. Statistics: a Journal of Theoretical and Applied Statistics, 48(3):508–523, 2014. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Denis:1988:TWA

[Den88] Jean Baptiste Denis. Two way analysis using covarites1. Statistics: a Journal of Theoretical and Applied Statistics, 19(1):123–132, 1988. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Dette:1993:BOM

[Det93] Holger Dette. Bayesian D-optimal and model robust designs in linear regression models. Statistics: a Journal of Theoretical and Applied Statistics, 25(1):27–46, 1993. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic). See addendum [Det95a].

Dette:1995:APB

[Det95a] Holger Dette. Addendum to the paper, Bayesian D-Optimal and Model Robust Designs in Linear Regression Models. Statistics: a Journal of Theoretical and Applied Statistics, 26(3):285–286, 1995. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic). See [Det93].

Dette:1995:D

[Det95b] Holger Dette. Discussion. Statistics: a Journal of Theoretical and Applied Statistics, 26(2):153–161, 1995. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Deutler:1972:VIA

[Deu72] T. Deutler. Die Verwendung eines iterativen Analogrechners zur Bestimmung der gesamten Lösungsmenge der Dalenius-Gleichungen (für verschiedene Aufteilungsarten) bei kleinen Schichtanzahlen. (German) [The use of an iterative analog computer to determine the total amount of solution of Dalenius equations (for different division types) at small layer numbers]. Mathematische Operationsforschung und Statistik, 3(6):465–474, 1972. CODEN MTOSBI. ISSN 0047-6277.

Devroye:1989:NDE

[Dev89] Luc Devroye. Nonparametric density estimates with improved performance on given sets of densities. Statistics: a Journal of Theoretical and Applied Statistics, 20(3):357–376, 1989. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Dewess:1973:ASM

[Dew73] Günter Dewess. Zur Anwendung der Schätzmethode MINQUE auf Probleme der Prozeßbilanzierung. (German)

[On application of the MINQUE estimation method to problems of process accounting]. Mathematische Operationsforschung und Statistik, 4(4):299–313, 1973. CODEN MTOSBI. ISSN 0047-6277.

Dewess:1975:EIW

[Dew75] M. Dewess. Erhaltungssätze im Wartemodell G/G/1 mit “Erwärmung”. (German) [Conservation laws in the waiting model G/G/1 with “warming”]. Mathematische Operationsforschung und Statistik, 6(3):427–436, 1975. CODEN MTOSBI. ISSN 0047-6277.

Dewess:1976:QIS

[Dew76] Günter Dewess. Quasikonvexität im Strukturgraphen der Wohlordnungen eines Netzplans. (German) [Quasiconvexity in the structure graph of well-orderings of a network]. Mathematische Operationsforschung und Statistik, 7(3):415–425, 1976. CODEN MTOSBI. ISSN 0047-6277.

Duchesne:2015:MHT

[DF15] Pierre Duchesne and Christian Francq. Multivariate hypothesis testing using generalized and 2-inverses — with applications. Statistics: a Journal of Theoretical and Applied Statistics, 49(3):475–496, 2015. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Durante:2012:MCH

[DFRLÚF12] Fabrizio Durante, Enrico Foscolo, José Antonio Rodríguez-Lallena, and Manuel Úbeda-Flores. A method for constructing higher-dimensional copulas. Statistics: a Journal of Theoretical and Applied Statistics, 46(3):387–404, 2012. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Duck:1973:RGB

[DGB^{+}73] W. Dück, H.-J. Girlich, O. Bunke, J. Piehler, and E. Griepentrog. Rezensionen. (German) [Book reviews]. Mathematische Operationsforschung und Statistik, 4(1):89–94, 1973. CODEN MTOSBI. ISSN 0047-6277.

Diaz-Garcia:2014:STS

[DGCL14] José A. Díaz-García and Francisco J. Caro-Lopera. Statistical theory of shape under elliptical models via QR decomposition. Statistics: a Journal of Theoretical and Applied Statistics, 48(2):456–472, 2014. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Duck:1975:RGB

[DGG^{+}75] W. Dück, A. Göpfert, Ch. Großmann, K. Tammer, K. H. Elster, K. Tammer, K. Richter, L. V. Wolfersdorf, L. V. Wolfersdorf, H. Hollatz, J. Focke, N. Sieber, H. W. Meier, N. Sieber, H. J. Sebastian, W. S. Witting, G. Nägler, and Klaus Wendler. Rezensionen. (German) [Book reviews]. Mathematische Operationsforschung und Statistik, 6(6):939–946, 1975. CODEN MTOSBI. ISSN 0047-6277.

Duick:1973:RGB

[DGH^{+}73] W. Düick, A. Göpfert, G. Hebermehl, J. Piehler, H. L. Burmeister, W. D. Müller, and H. J. Girlich. Rezensionen. (German) [Book reviews]. Mathematische Operationsforschung und Statistik, 4(3):243–248, 1973. CODEN MTOSBI. ISSN 0047-6277.

Dumitrescu:2016:MDE

[DGL16] M. Dumitrescu, M. L. Gámiz, and N. Limnios. Minimum divergence estimators for the Radon–Nikodym derivatives of the semi-Markov kernel. Statistics: a Journal of Theoretical and Applied Statistics, 50(3):486–504, 2016. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Dossou-Gbete:1991:ASE

[DGP91] Simplice Dossou-Gbete and Alain Pousse. Asymptotic study of eigenelements of a sequence of random selfadjoint operators. Statistics: a Journal of Theoretical and Applied Statistics, 22(3):479–491, 1991. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Drygas:1977:NPH

[DH77] Hilmar Drygas and Georges Hupet. A new proof of Hsu’s theorem in regression analysis — a coordinate-free approach. Statistics (Berlin, DDR), 8(3):333–335, 1977. CODEN MOSSD5. ISSN 0323-3944.

Drygas:1981:NNB

[DH81] Hilmar Drygas and Georges Hupet. Non-negativeness of the best quadratic unbiased estimator of the variance in the linear regression model. Statistics (Berlin, DDR), 12(1):3–5, 1981. CODEN MOSSD5. ISSN 0323-3944.

Dietz:1990:BRb

[DH90] Hans M. Dietz and K. Hoffmann. Book reviews. Statistics: a Journal of Theoretical and Applied Statistics, 21(3):483–484, 1990. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Darabi:2020:PBC

[DHN20] Nadiyeh Darabi and S. Mohammad E. Hosseini-Nasab. Projection-based classification for functional data. Statistics: a Journal of Theoretical and Applied Statistics, 54(3): 544–558, 2020. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Droge:1982:BR

[DHS82] B. Droge, K. Hoffmann, and W. Schmidt. Book review. Statistics (Berlin, DDR), 13(3):486–488, 1982. CODEN MOSSD5. ISSN 0323-3944.

DeGregorio:2011:LSC

[DI11] Alessandro De Gregorio and Stefano M. Iacus. Least-squares change-point estimation for the telegraph process observed at discrete times. Statistics: a Journal of Theoretical and Applied Statistics, 45(4):349–359, 2011. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Dietz:1990:BRa

[Die90] Hans M. Dietz. Book reviews. Statistics: a Journal of Theoretical and Applied Statistics, 21(1):158–159, 1990. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Dikta:2001:WRC

[Dik01] Gerhard Dikta. Weak representation of the cumulative hazard function under semiparametric random censorship models. Statistics: a Journal of Theoretical and Applied Statistics, 35(4):395–409, 2001. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Ducharme:1986:NBP

[DJ86] G. R. Ducharme and M. Jhun. A note on the bootstrap procedure for testing linear hypotheses. Statistics: a Journal of Theoretical and Applied Statistics, 17(4):527–531, 1986. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Daouia:2009:LDP

[DJ09] Abdelaati Daouia and Cyrille Joutard. Large deviation properties for empirical quantile-type production functions. Statistics: a Journal of Theoretical and Applied Statistics, 43(3): 267–277, 2009. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Durante:2012:IDS

[DJ12] Fabrizio Durante and Piotr Jaworski. Invariant dependence structure under univariate truncation. Statistics: a Journal of Theoretical and Applied Statistics, 46(2):263–277, 2012. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Dembinska:2017:ABP

[DJ17] Anna Dembińska and Krzysztof Jasiński. Asymptotic behaviour of proportions of observations in random regions determined by central order statistics from stationary processes. Statistics: a Journal of Theoretical and Applied Statistics, 51(3):591–608, 2017. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Dohler:1984:NWA

[DK84] R. Döhler and I. Küchler. A note on weak admissibility of the sequential probability ratio test. Statistics (Berlin, DDR), 15 (4):559–564, 1984. CODEN MOSSD5. ISSN 0323-3944.

Dey:2012:DBW

[DK12] Arabin Kumar Dey and Debasis Kundu. Discriminating between the Weibull and log-normal distributions for Type-II censored data. Statistics: a Journal of Theoretical and Applied Statistics, 46(2):197–214, 2012. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Dette:2014:ODM

[DK14] Holger Dette and Joachim Kunert. Optimal designs for the Michaelis–Menten model with correlated observations. Statistics: a Journal of Theoretical and Applied Statistics, 48(6): 1254–1267, 2014. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Diop:2021:CMS

[DK21] Mamadou Lamine Diop and William Kengne. Consistent model selection procedure for general integer-valued time series. Statistics: a Journal of Theoretical and Applied Statistics, 55(6):1207–1230, 2021. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Doukhan:2016:WDP

[DL16] Paul Doukhan and Gabriel Lang. Weak dependence of point processes and application to second-order statistics. Statistics: a Journal of Theoretical and Applied Statistics, 50(6):1221– 1235, 2016. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Demongeot:2013:FDL

[DLMR13] Jacques Demongeot, Ali Laksaci, Fethi Madani, and Mustapha Rachdi. Functional data: local linear estimation of the conditional density and its application. Statistics: a Journal of Theoretical and Applied Statistics, 47(1):26–44, 2013. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Desgagne:2022:GFT

[DLO22] Alain Desgagné, Pierre Lafaye de Micheaux, and Frédéric Ouimet. Goodness-of-fit tests for Laplace, Gaussian and exponential power distributions based on λ-th power skewness and kurtosis. Statistics: a Journal of Theoretical and Applied Statistics, 57(1):94–122, 2022. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
02331888.2022.2144859.

Droge:1991:BRa

[DM91] B. Droge and K. Mai. Book reviews. Statistics: a Journal of Theoretical and Applied Statistics, 22(1):148, 1991. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Dorkenoo:1995:GLM

[DM95] K. M. Dorkenoo and J.-R. Mathieu. Generalized linear models with a small scale parameter. Statistics: a Journal of Theoretical and Applied Statistics, 27(1–2):107–113, 1995. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Dey:2006:ODC

[DM06] Aloke Dey and Rahul Mukerjee. D-optimal designs for covariate models. Statistics: a Journal of Theoretical and Applied Statistics, 40(4):297–305, 2006. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Dechevsky:2009:AMN

[DM09] Lubomir Dechevsky and Brenda MacGibbon. Asymptotically minimax non-parametric function estimation with positivity constraints I. Statistics: a Journal of Theoretical and Applied Statistics, 43(5):423–442, 2009. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Dedecker:2022:CLT

[DM22] Jérôme Dedecker and Florence Merlevède. Central limit theorem and almost sure results for the empirical estimator of superquantiles/CVaR in the stationary case. Statistics: a Journal of Theoretical and Applied Statistics, 56(1):53–72, 2022. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Dette:2007:ODS

[DMP07] Holger Dette, Viatcheslav B. Melas, and Andrey Pepelyshev. Optimal designs for statistical analysis with Zernike polynomials. Statistics: a Journal of Theoretical and Applied Statistics, 41(6):453–470, 2007. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Dette:2010:ODE

[DMP10] Holger Dette, Viatcheslav B. Melas, and Andrey Pepelyshev. Optimal designs for estimating the slope of a regression. Statistics: a Journal of Theoretical and Applied Statistics, 44(6):617–628, 2010. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Doukhan:2007:WDI

[DMR07] Paul Doukhan, Hélène Madre, and Mathieu Rosenbaum. Weak dependence for infinite ARCH-type bilinear models. Statistics: a Journal of Theoretical and Applied Statistics, 41 (1):31–45, 2007. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Dauxois:1997:TLL

[DN97] Jacques Dauxois and Guy Martial Nkiet. Testing for the lack of a linear relationship. Statistics: a Journal of Theoretical and Applied Statistics, 30(1):1–23, 1997. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Dodynekova:1983:CPH

[Dod83] P. Dodynekova. Convergence properties of hit-and-run samplers. Statistics (Berlin, DDR), 14(4):509–515, 1983. CODEN MOSSD5. ISSN 0323-3944.

Dohler:1981:DSS

[Döh81] R. Döhler. Dominierbarkeit und Suffizienz in der Sequentialanalyse. (German) [Dominatability and insufficiency in sequential analysis]. Statistics (Berlin, DDR), 12(1):101–134, 1981. CODEN MOSSD5. ISSN 0323-3944.

Dohler:1989:BR

[Doh89] R. Dohler. Book reviews. Statistics: a Journal of Theoretical and Applied Statistics, 20(4):592, 1989. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Domma:2010:SPB

[Dom10] Filippo Domma. Some properties of the bivariate Burr type III distribution. Statistics: a Journal of Theoretical and Applied Statistics, 44(2):203–215, 2010. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Dorota:1985:DPL

[Dor85] Dabrowska Dorota. Descriptive parameters of location, dispersion and stochastic dependence. Statistics: a Journal of Theoretical and Applied Statistics, 16(1):63–88, 1985. CO-DEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Doring:2015:ACE

[Dör15] Maik Döring. Asymmetric cusp estimation in regression models. Statistics: a Journal of Theoretical and Applied Statistics, 49(6):1279–1297, 2015. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Doukhan:2011:E

[Dou11] Paul Doukhan. Editorial. Statistics: a Journal of Theoretical and Applied Statistics, 45(1):1–2, 2011. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Dobler:2017:ATE

[DP17] Dennis Dobler and Markus Pauly. Approximate tests for the equality of two cumulative incidence functions of a competing risk. Statistics: a Journal of Theoretical and Applied Statistics, 51(6):1238–1258, 2017. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

deQueiroz:2016:MLS

[dQLS16] Marina M. de Queiroz, Rosangela H. Loschi, and Roger W. C. Silva. Multivariate log-skewed distributions with normal kernel and their applications. Statistics: a Journal of Theoretical and Applied Statistics, 50(1):157–175, 2016. CO-DEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/02331888.2015.1032972.

Dayanand:1992:OTT

[DR92] N. Naik Dayanand and Khattree Ravindra. Optimum tests for treatments when blocks are nested and random. Statistics: a Journal of Theoretical and Applied Statistics, 23(2):101–108, 1992. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Dohler:2003:CRG

[DR03] Sebastian Döhler and Ludger Rüschendorf. A consistency result in general censoring models. Statistics: a Journal of Theoretical and Applied Statistics, 37(3):205–216, 2003. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Dickhaus:2015:SMC

[DR15] Thorsten Dickhaus and Thomas Royen. A survey on multivariate chi-square distributions and their applications in testing multiple hypotheses. Statistics: a Journal of Theoretical and Applied Statistics, 49(2):427–454, 2015. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Dreier:1998:IBS

[Dre98] Ilona Dreier. Inequalities between the second and fourth moments. Statistics: a Journal of Theoretical and Applied Statistics, 32(2):189–198, 1998. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Droge:1986:BR

[Dro86] B. Droge. Book review. Statistics: a Journal of Theoretical and Applied Statistics, 17(1):48, 1986. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Droge:1987:NEM

[Dro87] Bernd Droge. A note on estimating the msep in nonlinear regression. Statistics: a Journal of Theoretical and Applied Statistics, 18(4):499–520, 1987. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Droge:1988:BRa

[Dro88a] B. Droge. Book review. Statistics: a Journal of Theoretical and Applied Statistics, 19(2):214, 1988. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Droge:1988:BRb

[Dro88b] B. Droge. Book review. Statistics: a Journal of Theoretical and Applied Statistics, 19(2):228, 1988. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Droge:1988:BRc

[Dro88c] B. Droge. Book review. Statistics: a Journal of Theoretical and Applied Statistics, 19(3):368, 1988. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Droge:1988:BRd

[Dro88d] B. Droge. Book review. Statistics: a Journal of Theoretical and Applied Statistics, 19(4):556, 1988. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Droge:1988:BRe

[Dro88e] B. Droge. Book review. Statistics: a Journal of Theoretical and Applied Statistics, 19(4):584, 1988. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Droge:1989:BR

[Dro89] B. Droge. Book review. Statistics: a Journal of Theoretical and Applied Statistics, 20(1):108, 1989. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Droge:1990:BRJ

[Dro90a] B. Droge. Book review: J. P. C. Klelinen: Statistical Tools for Simulation Practioners. Statistics: a Journal of Theoretical and Applied Statistics, 21(3):368, 1990. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Droge:1990:S

[Dro90b] B. Droge. Statistics. Statistics: a Journal of Theoretical and Applied Statistics, 21(1):1, 1990. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Droge:1991:BRH

[Dro91a] B. Droge. Book review: H. Linhart and W. Zucchini: Model Selection. Statistics: a Journal of Theoretical and Applied Statistics, 22(3):418, 1991. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Droge:1991:BRb

[Dro91b] B. Droge. Book reviews. Statistics: a Journal of Theoretical and Applied Statistics, 22(3):492, 1991. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Droge:1993:FSP

[Dro93] Bernd Droge. On finite-sample properties of adaptive least squares regression estimates. Statistics: a Journal of Theoretical and Applied Statistics, 24(3):181–203, 1993. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Droge:1998:BR

[Dro98] B. Droge. Book review. Statistics: a Journal of Theoretical and Applied Statistics, 31(3):287–288, 1998. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Droge:2006:APM

[Dro06] Bernd Droge. Asymptotic properties of model selection procedures in linear regression. Statistics: a Journal of Theoretical and Applied Statistics, 40(1):1–38, 2006. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Droge:1994:BR

[DRR^{+}94] B. Droge, M. Reidel, M. Reidel, M. Reidel, F. Liese, Thomas Mikosch, I. Küclker, W. Kuhlisch, and D. Stoyan. Book review. Statistics: a Journal of Theoretical and Applied Statistics, 25(4):375–382, 1994. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Droge:1995:BR

[DRR^{+}95] B. Droge, W.-D. Richter, M. Riedel, B. Prum, B. Droge, M. Husková, and U. Lorz. Book review. Statistics: a Journal of Theoretical and Applied Statistics, 26(2):179–183, 1995. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Dehling:2017:TSS

[DRW17] H. Dehling, A. Rooch, and M. Wendler. Two-sample Ustatistic processes for long-range dependent data. Statistics: a Journal of Theoretical and Applied Statistics, 51(1):84–104, 2017. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Drygas:1972:ERV

[Dry72] Hilmar Drygas. The estimation of residual variance in regression analysis. Mathematische Operationsforschung und Statistik, 3(5):373–388, 1972. CODEN MTOSBI. ISSN 0047-6277.

Drygas:1975:EPL

[Dry75] Hilmar Drygas. Estimation and prediction for linear models in general spaces. Mathematische Operationsforschung und Statistik, 6(2):301–324, 1975. CODEN MTOSBI. ISSN 0047-6277.

Drygas:1977:BQU

[Dry77] Hilmar Drygas. Best quadratic unbiased estimation in variance-covariance component models. Statistics (Berlin, DDR), 8(2):211–231, 1977. CODEN MOSSD5. ISSN 0323-3944.

Drygas:1984:KAC

[Dry84] Hilmar Drygas. On the Kleffe-approach. for computing covariance-operators of tensor-products. Statistics (Berlin, DDR), 15(4):487–492, 1984. CODEN MOSSD5. ISSN 0323-3944.

Drygas:1988:MTE

[Dry88] Hilmar Drygas. MINQE-theory and the estimation of residual variance in regressions analysis. Statistics: a Journal of Theoretical and Applied Statistics, 19(3):341–347, 1988. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Deistler:1978:EET

[DS78] Manfred Deistler and Jürgen Schrader. Exact estimability of the transfer functions of linear systems. Statistics (Berlin, DDR), 9(3):411–417, 1978. CODEN MOSSD5. ISSN 0323-3944.

Dumitrescu:1991:SEP

[DS91] Monica E. Bad Dumitrescu and Stefan Corneliu V. Stefanescu. A sequential estimation procedure for m-dimensional Gaussian processes with independent increments. Statistics: a Journal of Theoretical and Applied Statistics, 22(2):269–282, 1991. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Dinh:1991:ERB

[DSG91] Khoan T. Dinh, Jagbir Singh, and Ramesh C. Gupta. Estimation of reliability in bivariate distributions. Statistics: a Journal of Theoretical and Applied Statistics, 22(3):409–417, 1991. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Dhabiyal:1985:NEU

[DSK85] Ishwari D. Dhabiyal, Divakar Shabma, and K. Krishnamoorthy. Non-existence of unbiased estimators of ordered parameters. Statistics: a Journal of Theoretical and Applied Statistics, 16(1):89–95, 1985. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

daSilva:2022:CSM

[dSL22] João Lita da Silva and Vanda Lourenço. Convergence of series of moments on general exponential inequality. Statistics: a Journal of Theoretical and Applied Statistics, 56(1):73–96, 2022. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Dung:2021:CMC

[DST21] L. V. Dung, T. C. Son, and T. T. Tu. Convergence in mean and central limit theorems for weighted sums of martingale difference random vectors with infinite r-th moments. Statistics: a Journal of Theoretical and Applied Statistics, 55(2): 386–408, 2021. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Diersen:1996:RTT

[DT96] J. Diersen and G. Trenkler. Records tests for trend in location. Statistics: a Journal of Theoretical and Applied Statistics, 28 (1):1–12, 1996. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Duan:2016:BEI

[DT16] Xing-De Duan and Nian-Sheng Tang. Bayesian estimation and influence diagnostics of generalized partially linear mixedeffects models for longitudinal data. Statistics: a Journal of Theoretical and Applied Statistics, 50(3):525–539, 2016. CO-DEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

deUna-Alvarez:2000:SCJ

[dUÁ00] Jacobo de Uña-Álvarez. Strong consistency of the jackknife estimate of variance and covariance for Koziol–Green integrals. Statistics: a Journal of Theoretical and Applied Statistics, 34(4):301–352, 2000. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

deUna-Alvarez:2004:SCP

[dUÁRC04] Jacobo de Uña-Álvarez and M. Celia Rodríguez-Campos. Strong consistency of presmoothed Kaplan–Meier integrals when covariables are present. Statistics: a Journal of Theoretical and Applied Statistics, 38(6):483–496, 2004. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

deUna-Alvarez:2017:CGE

[dUÁV17] Jacobo de Uña-Álvarez and Noël Veraverbeke. Copulagraphic estimation with left-truncated and right-censored data. Statistics: a Journal of Theoretical and Applied Statistics, 51(2):387–403, 2017. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Duck:1975:AEF

[Düc75a] Werner Dück. Abschätzung des Einflusses Fehlerhafter Eingangsinformationen Leontiefscher Matrizen durch Berücksichtigung von Blockstrukturen. (German) [Estimating the influence of erroneous input information of Leontief matrices by consideration of block structures]. Mathematische Operationsforschung und Statistik, 6(1):135–139, 1975. CODEN MTOSBI. ISSN 0047-6277.

Duck:1975:VAI

[Düc75b] Werner Dück. Verschärfung der Abschätzung der Inversen einer Leontiefschen Matrix mit fehlerbehafteten positiven Verflechtungskoeffizienten. (German) [Tightening estimations of the inverse of a Leontief matrix with faulty positive interdependence coefficients]. Mathematische Operationsforschung und Statistik, 6(3):327–330, 1975. CODEN MTOSBI. ISSN 0047-6277.

Dupac:1977:SAM

[Dup77] V. Dupac. Stochastic approximation methods in linear regression models (with consideration of errors in the regressors). Statistics (Berlin, DDR), 8(1):107–117, 1977. CODEN MOSSD5. ISSN 0323-3944.

Dupuy:2017:IGE

[Dup17] Jean-François Dupuy. Inference in a generalized endpointinflated binomial regression model. Statistics: a Journal of Theoretical and Applied Statistics, 51(4):888–903, 2017. CO-DEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Duval:2017:NFP

[Duv17] Céline Duval. A note on a fixed-point method for deconvolution. Statistics: a Journal of Theoretical and Applied Statistics, 51(2):347–362, 2017. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Devroye:1993:EDS

[DW93] Luc Devroye and Matthew P. Wand. On the effect of density shape on the performance of its kernel estimate. Statistics: a Journal of Theoretical and Applied Statistics, 24(3):215–233, 1993. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Deng:2019:GRC

[DWHH19] Xin Deng, Xuejun Wang, Shuhe Hu, and Ming Hu. A general result on complete convergence for weighted sums of linear processes and its statistical applications. Statistics: a Journal of Theoretical and Applied Statistics, 53(4):903–920, 2019. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Droge:1992:BR

[DWP^{+}92] B. Droge, W. Winkler, J. Pilz, J. Pilz, J. Pilz, W. Näther, B. Droge, J. Frauz, W. Näther, and O. Beyer. Book review. Statistics: a Journal of Theoretical and Applied Statistics, 23 (4):375, 1992. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Davidov:2002:SDS

[DY02] Ori Davidov and Chang Yu. Stratified double sampling with continuous outcomes: Design and analysis. Statistics: a Journal of Theoretical and Applied Statistics, 36(2):163–173, 2002. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Defor:2022:ELI

[DZ22] Edem Defor and Yichuan Zhao. Empirical likelihood inference for the mean past lifetime function. Statistics: a Journal of Theoretical and Applied Statistics, 56(2):329–346, 2022. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Du:2018:MAE

[DZX18] Jiang Du, Zhongzhan Zhang, and Tianfa Xie. Model averaging for M-estimation. Statistics: a Journal of Theoretical and Applied Statistics, 52(6):1417–1432, 2018. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Eather:1990:LRR

[EA90] Wolfgang Eather and Matthias Albbecht. Linear regression with random fuzzy observations. Statistics: a Journal of Theoretical and Applied Statistics, 21(4):521–531, 1990. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Esna-Ashari:2016:AMH

[EAA16] Maryam Esna-Ashari and Majid Asadi. On additivemultiplicative hazards model. Statistics: a Journal of Theoretical and Applied Statistics, 50(6):1421–1433, 2016. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Elgawad:2017:LTB

[EBQY17] M. A. Abd Elgawad, H. M. Barakat, Hong Qin, and Ting Yan. Limit theory of bivariate dual generalized order statistics with random index. Statistics: a Journal of Theoretical and Applied Statistics, 51(3):572–590, 2017. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Economou:2021:HTP

[EBTB21] P. Economou, A. Batsidis, G. Tzavelas, and D. Bagkavos. Hypothesis testing for the population mean and variance based on r -size biased samples. Statistics: a Journal of Theoretical and Applied Statistics, 55(4):894–924, 2021. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Economou:2013:MSD

[Eco13] P. Economou. Modelling survival data using mixtures of frailties. Statistics: a Journal of Theoretical and Applied Statistics, 47(2):453–464, 2013. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Elster:1974:URB

[EG74] K. H. Elster and C. Grossmann. Untersuchungen zu Regularitätsbedingungen bei den Barriere- und Zentrenmethoden. (German) [Studies on regularity conditions at the barrier and centers methods]. Mathematische Operationsforschung und Statistik, 5(3):191–206, 1974. CODEN MTOSBI. ISSN 0047-6277.

Eger:1980:DMC

[Ege80] Karl-Heinz Eger. A direct method of the computation of the OC and of the moments of the sample number for SPRTs in the case of discrete-random variables. Statistics (Berlin, DDR), 11(4):499–514, 1980. CODEN MOSSD5. ISSN 0323-

Eger:1987:BR

[Ege87] K. H. Eger. Book review. Statistics: a Journal of Theoretical and Applied Statistics, 18(1):122, 1987. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Erfurth:1973:RGB

[EGV^{+}73] H. Erfurth, G. Deweßand G. Grützmacher, H.-J Voss, H. Walther, B. Bank, W. Heinrich, H. Klemm, D. Stoyan, O. Bunke, L. Bellach, J. Kleffe, P. H. Müller, P. Neumann, and R. Sender. Rezensionen. (German) [Book reviews]. Mathematische Operationsforschung und Statistik, 4(4):327–340, 1973. CODEN MTOSBI. ISSN 0047-6277.

Ehlert:1975:EAV

[Ehl75] J. Ehlert. Über Erweiterungen eines allgemeinen Verfahrens von Poljak zur Lösung von Extremalaufgaben mit Nebenbedingungen und seine Anwendung in der nichtlinearen Optimierung. Teil I. (German) [On extensions of a general method of Poljak for solving extremal problems with constraints and its application in nonlinear optimization. Part I]. Mathematische Operationsforschung und Statistik, 6(1):91–105, 1975. CODEN MTOSBI. ISSN 0047-6277.

Ehlert:1976:EAV

[Ehl76] Jürgen Ehlert. Über Erweiterungen eines allgemeinen Verfahrens von Poljak zur Lösung von Extremalaufgaben mit Nebenbedingungen und seine Anwendung in der nichtlinearen Optimierung. II. (German) [On extensions of a general method of Poljak for solving extremal problems with constraints and its application in nonlinear optimization. Part II]. Mathematische Operationsforschung und Statistik, 7(2): 257–268, 1976. CODEN MTOSBI. ISSN 0047-6277.

Enderlein:1980:BR

[EHR^{+}80] G. Enderlein, H. Heckendorff, H. J. Rossbero, E. Platen, W. Watmuth, D. Stoyan, and B. Geabowski. Book reviews. Statistics (Berlin, DDR), 11(3):447–452, 1980. CO-DEN MOSSD5. ISSN 0323-3944.

Eicker:1988:TSD

[Eic88] Friedhelm Eicker. Trend-seasonal decomposition of time series as Whittaker–Henderson graduation. Statistics: a Journal of Theoretical and Applied Statistics, 19(2):313–338, 1988.

CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Evans:1981:BIT

[EJ81] I. G. Evans and P. W. Jones. Bayesian inferences for the two parameter uniform distribution. Statistics (Berlin, DDR), 12 (3):403–408, 1981. CODEN MOSSD5. ISSN 0323-3944.

Eubank:1993:AAS

[EJ93] R. L. Eubank and B. R. Jayasuriya. The asymptotic average squared error for polynomial regression. Statistics: a Journal of Theoretical and Applied Statistics, 24(4):311–319, 1993. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Ekstrom:1999:SLT

[Eks99] Magnus Ekström. Strong limit theorems for sums of logarithms of high order spacings. Statistics: a Journal of Theoretical and Applied Statistics, 33(2):153–169, 1999. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Egerton:1979:MLE

[EL79] M. F. Egerton and P. J. Laycock. Maximum likelihood estimation of multivariate non-linear functional relationships. Statistics (Berlin, DDR), 10(2):273–280, 1979. CODEN MOSSD5. ISSN 0323-3944.

Elster:1975:BAM

[Els75] K. H. Elster. Bericht über die Arbeitstagung “Mathematische Optimierung”. (German) [Report on the “Mathematical Optimization” workshop]. Mathematische Operationsforschung und Statistik, 6(4):663–664, 1975. CODEN MTOSBI. ISSN 0047-6277.

Elsawah:2017:CLA

[Els17] A. M. Elsawah. A closer look at de-aliasing effects using an efficient foldover technique. Statistics: a Journal of Theoretical and Applied Statistics, 51(3):532–557, 2017. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Elsawah:2022:NTP

[Els22] A. M. Elsawah. Novel techniques for performing successful follow-up experiments based on prior information from initial-stage experiments. Statistics: a Journal of Theoretical and Applied Statistics, 56(5):1133–1165, 2022. CO-DEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/02331888.2022.2133122.

Eltinge:1996:D

[Elt96] John L. Eltinge. Discussion. Statistics: a Journal of Theoretical and Applied Statistics, 27(3–4):241–244, 1996. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Ebner:2022:NOT

[ELV22] B. Ebner, S. C. Liebenberg, and I. J. H. Visagie. A new omnibus test of fit based on a characterization of the uniform distribution. Statistics: a Journal of Theoretical and Applied Statistics, 56(6):1364–1384, 2022. CO-DEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/02331888.2022.2133121.

Es:1987:TAL

[EM87] T. N. Es and H. Martens. Testing adequacy of linear random models. Statistics: a Journal of Theoretical and Applied Statistics, 18(3):323–331, 1987. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

ElOmari:2019:NPE

[EMF19] Mohamed El Omari, Hamid El Maroufy, and Christiane Fuchs. Non parametric estimation for fractional diffusion processes with random effects. Statistics: a Journal of Theoretical and Applied Statistics, 53(4):753–769, 2019. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Emmerich:1998:LIL

[EMG98] Frank Emmerich and Thomas Müller-Gronbach. A law of the iterated logarithm for discrete discrepancies and its applications to pseudorandom vector sequences. Statistics: a Journal of Theoretical and Applied Statistics, 32(1):75–85, 1998. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Elster:1974:DFG

[EN74a] Karl-Heinz Elster and Reinhard Nehse. Zum dualitätssatz von Fenchel. (German) [On Fenchel’s duality theorem]. Mathematische Operationsforschung und Statistik, 5(4–5):269–280, 1974. CODEN MTOSBI. ISSN 0047-6277.

Elster:1974:TPG

[EN74b] Karl-Heinz Elster and Reinhard Nehse. Zur Theorie der Polarfunktionale. (German) [On the theory of polar functionals]. Mathematische Operationsforschung und Statistik, 5(1):3–21, 1974. CODEN MTOSBI. ISSN 0047-6277.

Elster:1975:KOS

[EN75] Karl-Heinz Elster and Reinhard Nehse. Konjugierte operatoren und subdifferentiale. (German) [Conjugated operators and subdifferentials]. Mathematische Operationsforschung und Statistik, 6(4):641–657, 1975. CODEN MTOSBI. ISSN 0047-6277.

Engelbert:1982:BR

[Eng82] H. J. Engelbert. Book review. Statistics (Berlin, DDR), 13 (4):520, 1982. CODEN MOSSD5. ISSN 0323-3944.

Engelbert:1983:BR

[Eng83] H. J. Engelbert. Book reviews. Statistics (Berlin, DDR), 14 (2):350–351, 1983. CODEN MOSSD5. ISSN 0323-3944.

Engelbert:1988:BR

[Eng88] H. J. Engelbert. Book review. Statistics: a Journal of Theoretical and Applied Statistics, 19(4):518, 1988. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Ennadifi:1998:SAR

[Enn98] Gratiane Ennadifi. Strong approximations for K-records and records in the F^{α}-scheme — a duality. Statistics: a Journal of Theoretical and Applied Statistics, 31(4):359–382, 1998. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Evangelaras:2017:EAT

[EP17] H. Evangelaras and C. Peveretos. Efficient arrangements of two-level orthogonal arrays in two and four blocks. Statistics: a Journal of Theoretical and Applied Statistics, 51(6):1326– 1341, 2017. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Economou:2020:RST

[EPT20] Polychronis Economou, Georgios Psarrakos, and Abdolsaeed Toomaj. Recruitment and survival time under mean residual lifetime bias sampling. Statistics: a Journal of Theoretical and Applied Statistics, 54(4):885–907, 2020. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Elster:1971:BGR

[ES71] K. H. Elster and M. Schäuble. Berichte. (German) [Reports]. Mathematische Operationsforschung und Statistik, 2(4):319– 322, 1971. CODEN MTOSBI. ISSN 0047-6277.

Elster:1973:DHN

[ES73] Karl-Heinz Elster and Conrad Suppe. Zur Dualitätstheorie homogener nichtlinearer Optimierungsprobleme. (German) [On the duality theory of homogeneous nonlinear optimization problems]. Mathematische Operationsforschung und Statistik, 4(5):351–363, 1973. CODEN MTOSBI. ISSN 0047-6277.

Elster:1974:BXI

[ES74] K. H. Elster and H. Sachs. Bericht über das XVIII. Internationale Wissenschaftliche Kolloquium 1973 der Technischen Hochschule Ilmenau. (German) [Report on the XVIII. International Scientific Colloquium in 1973 at the Technical University of Ilmenau]. Mathematische Operationsforschung und Statistik, 5(4–5):439–446, 1974. CODEN MTOSBI. ISSN 0047-6277.

Eschenbach:1984:SIQ

[Esc84] W. Eschenbach. Statistical inference for queueing models. Statistics (Berlin, DDR), 15(3):451–462, 1984. CODEN MOSSD5. ISSN 0323-3944.

El-Saidi:1993:SGL

[ESSB93] Mohammed A. El-Saidi, Karan P. Singh, and Alfred A. Bartolucci. Symmetric generalized logistic distribution: Some moment properties and approximation to the normal distribution. Statistics: a Journal of Theoretical and Applied Statistics, 25(1):79–87, 1993. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Ehm:1999:URP

[ETD99] Werner Ehm, Gneiting Tilmann, and Richards Donald. On the uncertainty relation for positive-definite probability densities, II. Statistics: a Journal of Theoretical and Applied Statistics, 33(3):267–286, 1999. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Elsawah:2019:COP

[ETF19] A. M. Elsawah, Yu Tang, and Kai-Tai Fang. Constructing optimal projection designs. Statistics: a Journal of Theoretical and Applied Statistics, 53(6):1357–1385, 2019. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Evans:1975:BAS

[Eva75] I. G. Evans. A Bayesian acceptance sampling scheme when utilities are piecewise constant and the sampling cost is linear. Mathematische Operationsforschung und Statistik, 6(5):797– 807, 1975. CODEN MTOSBI. ISSN 0047-6277.

Erfurth:1975:RGB

[EvWE^{+}75] H. Erfurth, L. von Wolfersdorf, H. Erfurth, P. Franken, W. Randt, and E. Griepentrog. Rezensionen. (German) [Book reviews]. Mathematische Operationsforschung und Statistik, 6 (3):503–508, 1975. CODEN MTOSBI. ISSN 0047-6277.

Elster:1973:DKL

[EW73] Karl-Heinz Elster and Ulrich Wendt. Zur dualitätstheorie der komplexen linearen Optimierung. (German) [On the duality theory of complex linear optimization]. Mathematische Operationsforschung und Statistik, 4(3):183–200, 1973. CODEN MTOSBI. ISSN 0047-6277.

Elster:1975:RGB

[EW75a] K. H. Elster and L. V. Wolfersdorf. Rezensionen. (German) [Book reviews]. Mathematische Operationsforschung und Statistik, 6(4):659–661, 1975. CODEN MTOSBI. ISSN 0047-6277.

Eschenbach:1975:MLS

[EW75b] Walter Eschenbach and Wolfgang Winkler. Maximum-Likelihood-Schätzungen beim Verzweigungsprozeß von Galton– Watson. (German) [Maximum likelihood estimates of the branching process of Galton–Watson]. Mathematische Operationsforschung und Statistik, 6(2):213–224, 1975. CODEN MTOSBI. ISSN 0047-6277.

Eubank:1999:CLT

[EW99] R. L. Eubank and Suojin Wang. A central limit theorem for the sum of generalized linear and quadratic forms. Statistics: a Journal of Theoretical and Applied Statistics, 33(1):85–91, 1999. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Exner:1976:SEU

[Exn76] H. Exner. Sequentielle Entscheidungsverfahren für den unbekannten Erwartungswert der Normalverteilung. (German) [Sequential decision procedures for the unknown expected value of the normal distribution]. Mathematische Operationsforschung und Statistik, 7(5):773–787, 1976. CODEN MTOSBI. ISSN 0047-6277.

Ebrahimi:1992:MOS

[EZ92] N. B. Ebrahimi and Hassan Zahedi. Memory ordering of survival functions. Statistics: a Journal of Theoretical and Applied Statistics, 23(4):337–345, 1992. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Fahrmann:1978:KOW

[Fäh78] Hans-Heiner Fährmann. Zur Konvergenz der optimalen Werte der Gewinnfunktion beim Abbruch von Zufallsprozessen Im Falle von unvollständiger Information. (German) [On the convergence of the optimal values of the profit function from the discontinuity of random processes in the case of incomplete information]. Statistics (Berlin, DDR), 9(2):241–253, 1978. CODEN MOSSD5. ISSN 0323-3944.

Fahrmann:1981:KWB

[Fäh81] Hans-Heiner Fährmann. Zur Konvergenz der Werte beim optimalen Abbruch teilweise beobachtbarer Zufallsprozesse im Falle einer quadratischen Gewinnfunktion. (German) [On the convergence of the values at the optimal discontinuity of partially observable random processes in the case of a quadratic profit function]. Statistics (Berlin, DDR), 12(1):75–84, 1981. CODEN MOSSD5. ISSN 0323-3944.

Fahrmeir:1987:ATT

[Fah87] L. Fahrmeir. Asymptotic testing theory for generalized linear models. Statistics: a Journal of Theoretical and Applied Statistics, 18(1):65–76, 1987. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Fahrmeir:1990:MLE

[Fah90] Ludwig Fahrmeir. Maximum likelihood estimation in misspecified generalized linear models. Statistics: a Journal of Theoretical and Applied Statistics, 21(4):487–502, 1990. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Fakhry:1996:SCP

[Fak96] Mohammed E. Fakhry. Some characteristic properties of a certain family of probability distributions. Statistics: a Journal of Theoretical and Applied Statistics, 28(2):179–185, 1996. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Falk:1995:SBP

[Fal95] Michael Falk. Some best parameter estimates for distributions with finite endpoint. Statistics: a Journal of Theoretical and Applied Statistics, 27(1–2):115–125, 1995. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Frisen:2010:SEO

[FAP10] Marianne Frisén, Eva Andersson, and Kjell Pettersson. Semiparametric estimation of outbreak regression. Statistics: a Journal of Theoretical and Applied Statistics, 44(2):107–117, 2010. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Favetto:2014:PEC

[Fav14] Benjamin Favetto. Parameter estimation by contrast minimization for noisy observations of a diffusion process. Statistics: a Journal of Theoretical and Applied Statistics, 48(6): 1344–1370, 2014. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Fazli:2008:HTM

[Faz08] Kh. Fazli. Hypotheses testing for a multidimensional parameter of inhomogeneous Poisson processes. Statistics: a Journal of Theoretical and Applied Statistics, 42(5):407–428, 2008. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Fazli:2014:ADS

[Faz14] Kh. Fazli. Asymptotic deficiency of score test for inhomogeneous Poisson processes. Statistics: a Journal of Theoretical and Applied Statistics, 48(3):473–486, 2014. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Fang:2016:ORS

[FB16] Longxiang Fang and N. Balakrishnan. Ordering results for the smallest and largest order statistics from independent heterogeneous exponential-Weibull random variables. Statistics: a Journal of Theoretical and Applied Statistics, 50(6):1195– 1205, 2016. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

FerrazdoNascimento:2022:BTV

[FB22] Fernando Ferraz do Nascimento and Marcelo Bourguignon. Bayesian time-varying quantile regression on exceedance. Statistics: a Journal of Theoretical and Applied Statistics, 56(6):1207–1224, 2022. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/02331888.2022.2146691.

Famonye:1989:SUM

[FC89a] F. Famonye and P. C. Consul. A stochastic urn model for the generalized negative binomial distribution. Statistics: a Journal of Theoretical and Applied Statistics, 20(4):607–613, 1989. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Famoye:1989:CIE

[FC89b] Felix Famoye and P. C. Consul. Confidence interval estimation in the class of modified power series distributions. Statistics: a Journal of Theoretical and Applied Statistics, 20(1): 141–143, 1989. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Focke:1972:SMC

[FD72] J. Focke and G. Dewess. Über die schätzmethode MINQUE von C. R. Rao und ihre Verallgemeinerung. (German) [On the MINQUE estimation method of C. R. Rao and its generalization]. Mathematische Operationsforschung und Statistik, 3(2): 129–143, 1972. CODEN MTOSBI. ISSN 0047-6277.

Ferreira:2003:OEH

[FdHP03] A. Ferreira, L. de Haan, and L. Peng. On optimising the estimation of high quantiles of a probability distribution. Statistics: a Journal of Theoretical and Applied Statistics, 37(5): 401–434, 2003. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Flobens:1985:CBV

[FE85] P. Flobens and I. G. Evans. Comparisons of Bayesian variable and attribute sampling acceptance plans when the quality distribution is normal with known variance and the conformance requirement is one-sided. Statistics: a Journal of Theoretical and Applied Statistics, 16(2):219–231, 1985. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Fedorov:1980:CDT

[Fed80] V. V. Fedorov. Convex design theory. Statistics (Berlin, DDR), 11(3):403–413, 1980. CODEN MOSSD5. ISSN 0323-3944.

Feichtinger:1972:KML

[Fei72] Gustav Feichtinger. Über die Kovarianzen von Maximum-Likelihood-Schätzungen gewisser Parameter bei multiplen Dekrementtafeln. (German) [On the covariances of the maximum likelihood estimates of certain parameters for multiple decrement tables]. Mathematische Operationsforschung und Statistik, 3(3):217–226, 1972. CODEN MTOSBI. ISSN 0047-6277.

Ferrieux:1999:EMM

[Fer99] Dominique Ferrieux. Estimation of the mean measure density of a discrete random measure through associated sequences of observations. Statistics: a Journal of Theoretical and Applied Statistics, 33(2):129–152, 1999. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Fernandez:2007:CGC

[Fer07a] Arturo J. Fernández. On calculating generalized confidence intervals for the two-parameter exponential reliability function. Statistics: a Journal of Theoretical and Applied Statistics, 41(2):129–135, 2007. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Fernandez:2007:PCB

[Fer07b] Arturo J. Fernández. Posterior computations based on sample quantiles: one- and two-parameter exponential cases. Statistics: a Journal of Theoretical and Applied Statistics, 41(5): 409–420, 2007. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Focke:1975:JGA

[FG75] J. Focke and A. Göpfert. 100 Jahre gordanscher Alternativsatz für lineare Ungleichungen. (German) [100 years of Gordan alternative theorems for linear inequalities]. Mathematische Operationsforschung und Statistik, 6(6):873–880, 1975. CODEN MTOSBI. ISSN 0047-6277.

Figueiredo:2006:DAB

[FG06] Adelaide Figueiredo and Paulo Gomes. Discriminant analysis based on the Watson distribution defined on the hypersphere. Statistics: a Journal of Theoretical and Applied Statistics, 40(5):435–445, 2006. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Fournier:2019:SEP

[FGK19] Edouard Fournier, Stéphane Grihon, and Thierry Klein. Semiparametric estimation of plane similarities: application to fast computation of aeronautic loads. Statistics: a Journal of Theoretical and Applied Statistics, 53(5):1168–1186, 2019. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Fan:2017:NUB

[FGL17a] X. Fan, I. Grama, and Q. Liu. Non-uniform Berry–Esseen bounds for martingales with applications to statistical estimation. Statistics: a Journal of Theoretical and Applied Statistics, 51(1):105–122, 2017. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Fan:2017:MIT

[FGL17b] Xiequan Fan, Ion Grama, and Quansheng Liu. Martingale inequalities of type Dzhaparidze and van Zanten. Statistics: a Journal of Theoretical and Applied Statistics, 51(6):1200– 1213, 2017. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Focke:1973:KGG

[FGR73] J. Focke, A. Göpfert, and H. Rudolph. Konvergenzuntersuchungen bei Gradientenverfahren. (German) [Convergence studies of gradient methods]. Mathematische Operationsforschung und Statistik, 4(1):35–50, 1973. CODEN MTOSBI. ISSN 0047-6277.

Franken:1976:RGB

[FGR^{+}76] P. Franken, J. Gladitz, W. Randt, O. Bunke, P. P. Mager, R. Pincus, O. Bunke, H. Bunke, K. Fischer, and U. Schulze. Rezensionen. (German) [Book reviews]. Mathematische Operationsforschung und Statistik, 7(1):169–175, 1976. CODEN MTOSBI. ISSN 0047-6277.

Fotopoulos:2001:CVG

[FH01] Stergios B. Fotopoulos and Lijian He. The conditional variance for gamma mixtures of normal distributions. Statistics: a Journal of Theoretical and Applied Statistics, 35(4):319–330, 2001. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Fiksel:1988:ECD

[Fik88a] Thomas Fiksel. Edge-corrected density estimators for point processes. Statistics: a Journal of Theoretical and Applied Statistics, 19(1):67–75, 1988. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Fiksel:1988:EIP

[Fik88b] Thomas Fiksel. Estimation of interaction potentials of Gibbsian point processes. Statistics: a Journal of Theoretical and Applied Statistics, 19(1):77–86, 1988. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Fine:1987:VPM

[Fin87] Jeanne Fine. On the validity of the perturbation method in asymptotic theory. Statistics: a Journal of Theoretical and Applied Statistics, 18(3):401–414, 1987. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Fine:1994:ASM

[Fin94] Jeanne Fine. Asymptotic study of the multivariate functional model in the case of a random number of observations for each mean. Statistics: a Journal of Theoretical and Applied Statistics, 25(4):285–306, 1994. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Fischer:1987:BR

[Fis87] K. Fischer. Book review. Statistics: a Journal of Theoretical and Applied Statistics, 18(3):452, 1987. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Fischer:1988:BR

[Fis88] K. Fischer. Book review. Statistics: a Journal of Theoretical and Applied Statistics, 19(1):76, 1988. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Fischer:1989:BR

[Fis89] K. Fischer. Book review. Statistics: a Journal of Theoretical and Applied Statistics, 20(1):94, 1989. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Fischer:1990:BR

[Fis90] K. Fischer. Book reviews. Statistics: a Journal of Theoretical and Applied Statistics, 21(3):460, 1990. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Fiserova:2009:CRB

[Fis09] Eva Fiserová. Confidence regions based on inefficient estimators. Statistics: a Journal of Theoretical and Applied Statistics, 43(3):241–251, 2009. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Fritzsche:1990:CTI

[FK90] Bernd Fritzsche and Bernd Kirstein. A correlation-theoretical interpretation of Schur analysis. Statistics: a Journal of Theoretical and Applied Statistics, 21(1):137–148, 1990. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Fischer:2012:CGG

[FK12] Matthias Fischer and Christian Köck. Constructing and generalizing given multivariate copulas: a unifying approach.

Statistics: a Journal of Theoretical and Applied Statistics, 46 (1):1–12, 2012. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Fischer:2014:SPM

[FK14] Tim Fischer and Udo Kamps. Structure-preserving mappings of uniform order statistics. Statistics: a Journal of Theoretical and Applied Statistics, 48(1):142–158, 2014. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Fabre:2003:AEA

[FL03] Jean-Pierre Fabre and Irène Larramendy. About estimation of ARIMA process with strong mixing MA part. Statistics: a Journal of Theoretical and Applied Statistics, 37(1):17–32, 2003. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Flores:2010:DOE

[FL10] Elvia Flores and José R. León. Deviation of order p for estimators of the variance in first-order stochastic differential equation (SDE). Statistics: a Journal of Theoretical and Applied Statistics, 44(5):431–454, 2010. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Fleischmann:1971:SME

[Fle71] Klaus Fleischmann. Zur Stationarität eines modifizierten Erlangschen Bedienungsprozesses. (German) [On the stationarity of a modified Erlang operation process]. Mathematische Operationsforschung und Statistik, 2(4):283–290, 1971. CO-DEN MTOSBI. ISSN 0047-6277.

Fleischmann:1976:OIL

[Fle76] Klaus Fleischmann. Optimal input for the loss system G/M/2. Mathematische Operationsforschung und Statistik, 7(1):129–137, 1976. CODEN MTOSBI. ISSN 0047-6277.

Fleischmann:1991:BR

[Fle91] K. Fleischmann. Book reviews. Statistics: a Journal of Theoretical and Applied Statistics, 22(1):58, 1991. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Fang:2016:SCS

[FLL16] Rui Fang, Chen Li, and Xiaohu Li. Stochastic comparisons on sample extremes of dependent and heterogeneous observations. Statistics: a Journal of Theoretical and Applied Statistics, 50(4):930–955, 2016. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Fang:2018:ORE

[FLL18] Rui Fang, Chen Li, and Xiaohu Li. Ordering results on extremes of scaled random variables with dependence and proportional hazards. Statistics: a Journal of Theoretical and Applied Statistics, 52(2):458–478, 2018. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Florczak:1993:ARE

[Flo93] Witold J. Florczak. Asymptotic representation of mestimators and rate of convergence of one-step m-estimators for parametric models with shrinking contamination. Statistics: a Journal of Theoretical and Applied Statistics, 24(2): 107–115, 1993. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Fedorov:1972:UGW

[FM72] Y. V. Fedorov and M. B. Malyutov. Übersichtsarbeiten. (German) [Work overview]. Mathematische Operationsforschung und Statistik, 3(4):281–308, 1972. CODEN MTOSBI. ISSN 0047-6277.

Fedorov:1989:CTA

[FM89] V. Fedorov and W. Mueller. Comparison of two approaches in the optimal design of an observation network. Statistics: a Journal of Theoretical and Applied Statistics, 20(3):339–351, 1989. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Franza:2001:PES

[FM01] Jürgen Franza and Ryszard Magiera. Parameter estimation for switched counting processes. Statistics: a Journal of Theoretical and Applied Statistics, 35(4):371–393, 2001. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Flobens:1985:TDI

[FMM85] J. P. Flobens, M. Mouchabt, and Rolin J. M. On two definitions of identification. Statistics: a Journal of Theoretical and Applied Statistics, 16(2):213–218, 1985. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Fonseca:2007:TIT

[FMTZ07] Miguel Fonseca, Thomas Mathew, João Tiago Mexia, and Roman Zmyślony. Tolerance intervals in a two-way nested model with mixed or random effects. Statistics: a Journal of Theoretical and Applied Statistics, 41(4):289–300, 2007. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Ferse:1977:MAM

[FN77] Armin Ferse and Peter Neuman. Mathematische Aspekte zu einer Methode zur numerischen Bestimmung individueller thermodynamischer Aktivitätskoeffizienten einzelner Ionensorten in konzentrierten Elektrolytlösungen. (German) [Mathematical aspects of a method for the numerical determination of individual thermodynamic activity coefficients of individual ions located in concentrated electrolyte solutions]. Statistics (Berlin, DDR), 8(4):529–543, 1977. CODEN MOSSD5. ISSN 0323-3944.

Fu:2013:NSA

[FN13] Ke-Ang Fu and Andrew Cheuk-Yin Ng. A note on the strong approximation for long memory processes and its application. Statistics: a Journal of Theoretical and Applied Statistics, 47(3):511–520, 2013. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Focke:1971:DBR

[Foc71] J. Focke. Dualität beim Richtungssuchproblem von Zoutendijk. (German) [Duality in the direction-of-search problem of Zoutendijk]. Mathematische Operationsforschung und Statistik, 2(3):213–219, 1971. CODEN MTOSBI. ISSN 0047-6277.

Focke:1973:VPO

[Foc73] J. Focke. Vektormaximumproblem und parametrische Optimierung. (German) [Maximum vector problem and parametric optimization]. Mathematische Operationsforschung und Statistik, 4(5):365–369, 1973. CODEN MTOSBI. ISSN 0047-6277.

Focke:1975:MRG

[Foc75a] J. Focke. Eine Modifikation des reduzierten Gradientenverfahrens. (German) [A modification of the reduced gradient method]. Mathematische Operationsforschung und Statistik, 6(4):561–569, 1975. CODEN MTOSBI. ISSN 0047-6277.

Focke:1975:SLU

[Foc75b] J. Focke. Strenge lineare Ungleichungen bezüglich Kegelhalbordnungen. (German) [Strict linear inequalities in terms of cone half-orderings]. Mathematische Operationsforschung und Statistik, 6(6):881–900, 1975. CODEN MTOSBI. ISSN 0047-6277.

Focke:1984:EML

[Foc84] Joachim Focke. Zur Existenz der Maximum-Likelihood-Schätzung beim Faktorenmodell. (German) [On the existence of the maximum likelihood estimate of the model factors]. Statistics (Berlin, DDR), 15(1):51–57, 1984. CODEN MOSSD5. ISSN 0323-3944.

Fokianos:2011:SRP

[Fok11] Konstantinos Fokianos. Some recent progress in count time series. Statistics: a Journal of Theoretical and Applied Statistics, 45(1):49–58, 2011. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Foucart:1997:NAC

[Fou97] Thierry Foucart. Numerical analysis of a correlation matrix. Statistics: a Journal of Theoretical and Applied Statistics, 29(4):347–361, 1997. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Fine:1992:ASM

[FP92] Jeanne Fine and Alain Pousse. Asymptotic study of the multivariate functional model. Application to the metric choice in principal component analysis. Statistics: a Journal of Theoretical and Applied Statistics, 23(1):63–83, 1992. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Franco-Pereira:2012:DPR

[FPLS12] Alba M. Franco-Pereira, Rosa E. Lillo, and Moshe Shaked. The decreasing percentile residual life ageing notion. Statistics: a Journal of Theoretical and Applied Statistics, 46(5): 587–603, 2012. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Fernandez-Ponce:2016:SSP

[FPPRG16] J. M. Fernández-Ponce, F. Pellerey, and M. R. Rodríguez-Griñolo. Some stochastic properties of conditionally dependent frailty models. Statistics: a Journal of Theoretical and Applied Statistics, 50(3):649–666, 2016. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Fine:1984:RPC

[FR84] Jeanne Fine and Yves Romain. Reduced principal component analysis. Statistics (Berlin, DDR), 15(4):493–512, 1984. CODEN MOSSD5. ISSN 0323-3944.

Franz:1986:SMS

[FR86] J. Franz and W. Rüger. Sequentielle Minimalvarianz-Schätzungen in einem speziellen Markovprozeß. (German) [Sequential minimal variance estimates in a special Markov process]. Statistics: a Journal of Theoretical and Applied Statistics, 17(1):105–119, 1986. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Franco:1995:CCD

[FR95] M. Franco and J. M. Ruiz. On characterization of continuous distributions with adjacent order statistics. Statistics: a Journal of Theoretical and Applied Statistics, 26(4):375–385, 1995. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Francq:1998:EAP

[FR98] Christian Francq and Michel Roussignol. Ergodicity of autoregressive processes with Markov-switching and consistency of the maximum-likelihood estimator. Statistics: a Journal of Theoretical and Applied Statistics, 32(2):151–173, 1998. CO-DEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Furmanczyk:2020:HDL

[FR20] Konrad Furmańczyk and Wojciech Rejchel. High-dimensional linear model selection motivated by multiple testing. Statistics: a Journal of Theoretical and Applied Statistics, 54(1): 152–166, 2020. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Franz:1977:NCS

[Fra77] Jürgen Franz. Niveaudurchagangszeiten zur Charakterisierung sequentieller Schätzverfahren. (German) [Underestimated procedural level-crossing times for characterization of sequential estimating procedures]. Statistics (Berlin, DDR), 8(4):499–510, 1977. CODEN MOSSD5. ISSN 0323-3944.

Franz:1979:CCS

[Fra79] Jürgen Franz. Consistency and completeness in sequential estimation problems. Statistics (Berlin, DDR), 10(1):127– 139, 1979. CODEN MOSSD5. ISSN 0323-3944.

Franz:1982:SEA

[Fra82] Jürgen Franz. Sequential estimation and asymptotic properties in birth-and-death processes. Statistics (Berlin, DDR), 13(2):231–244, 1982. CODEN MOSSD5. ISSN 0323-3944.

Fischer:1977:BR

[FRB77] K. Fischer, P. Rudolph, and O. Bunke. Book review. Statistics (Berlin, DDR), 8(1):127–128, 1977. CODEN MOSSD5. ISSN 0323-3944.

Freyer:1974:BMG

[Fre74]  Bernd Freyer.  Ein Bedienungssystem (M/G/∞) mit

zeitabhängiger Eingangsintensität und Bedienungszeitverteilung. (German) [An operating system (M/G/∞) with timedependent input intensity and service time distribution]. Mathematische Operationsforschung und Statistik, 5(9):701– 708, 1974. CODEN MTOSBI. ISSN 0047-6277.

Freudenberg:1988:BR

[Fre88] W. Freudenberg. Book review. Statistics: a Journal of Theoretical and Applied Statistics, 19(2):240, 1988. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Fremdt:2015:PSP

[Fre15] Stefan Fremdt. Page’s sequential procedure for change-point detection in time series regression. Statistics: a Journal of Theoretical and Applied Statistics, 49(1):128–155, 2015. CO-DEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Frisen:1974:RES

[Fri74] Marianne Frisén. Recognition of elliptical shape. Mathematische Operationsforschung und Statistik, 5(3):213–221, 1974. CODEN MTOSBI. ISSN 0047-6277.

Fajardo:2018:PAL

[FRLLT18] F. A. Fajardo, V. A. Reisen, C. Lévy-Leduc, and M. S. Taqqu. M-periodogram for the analysis of long-range-dependent time series. Statistics: a Journal of Theoretical and Applied Statistics, 52(3):665–683, 2018. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Fujisawa:1981:SSS

[FS81] T. Fujisawa and Y. Sugizaki. Steady-state solution of a single server queue with Poisson arrivals and multiple transaction. Statistics (Berlin, DDR), 12(2):233–241, 1981. CODEN MOSSD5. ISSN 0323-3944.

FloresDeFrutos:2002:GLS

[FS02] Rafael Flores De Frutos and Gregorio Serrano. A generalized least squares estimation method for VARMA models. Statistics: a Journal of Theoretical and Applied Statistics, 36(4): 303–316, 2002. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Fleischmann:1994:BR

[FSS^{+}94] Klaus Fleischmann, V. Schmidt, V. Schmidt, H.-J. Girlich, H. J. Engelbert, J. Andel, H.-J. Girlich, Walter Warmuth, Walter Warmuth, B. Kirstein, and M. Weese. Book review. Statistics: a Journal of Theoretical and Applied Statistics, 25(2):177–187, 1994. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Ferraty:2013:MED

[FSV13] Frédéric Ferraty, Mariela Sued, and Philippe Vieu. Mean estimation with data missing at random for functional covariables. Statistics: a Journal of Theoretical and Applied Statistics, 47(4):688–706, 2013. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Fischer:1979:DFM

[FT79] K. Fischer and Chr Thiele. On a distribution-free method in discriminant analysis. Statistics (Berlin, DDR), 10(2):281– 289, 1979. CODEN MOSSD5. ISSN 0323-3944.

Furno:1997:RHC

[Fur97] Marilena Furno. A robust heteroskedasticity consistent covariance matrix estimator. Statistics: a Journal of Theoretical and Applied Statistics, 30(3):201–219, 1997. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Furmanczyk:2015:SSP

[Fur15] Konrad Furmańczyk. On some stepdown procedures with application to consistent variable selection in linear regression. Statistics: a Journal of Theoretical and Applied Statistics, 49(3):614–628, 2015. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Franco:2006:LCE

[FV06] Manuel Franco and Juana María Vivo. Log-concavity of the extremes from Gumbel bivariate exponential distributions. Statistics: a Journal of Theoretical and Applied Statistics, 40(5):415–433, 2006. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Feng:2013:PHC

[FW13] Changyong Feng and Hongyue Wang. Proportionality of hazards in competing risk analysis. Statistics: a Journal of Theoretical and Applied Statistics, 47(3):654–661, 2013. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Fang:2020:SCS

[FW20] Rui Fang and Boyang Wang. Stochastic comparisons on sample extremes from independent or dependent gamma samples. Statistics: a Journal of Theoretical and Applied Statistics, 54(4):841–855, 2020. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Feng:2016:PFL

[FX16] Sanying Feng and Liugen Xue. Partially functional linear varying coefficient model. Statistics: a Journal of Theoretical and Applied Statistics, 50(4):717–732, 2016. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Ferre:2003:FSI

[FY03] L. Ferré and A. F. Yao. Functional sliced inverse regression analysis. Statistics: a Journal of Theoretical and Applied Statistics, 37(6):475–488, 2003. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Fang:2023:VBD

[FYY^{+}23] Hongyan Fang, Chunyu Yao, Wenzhi Yang, Xuejun Wang, and Huang Xu. A p-value based dimensionality reduction test for high dimensional means. Statistics: a Journal of Theoretical and Applied Statistics, 57(2):282–299, 2023. CO-DEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/02331888.2023.2179627.

Florens-Zmirou:1989:ADT

[FZ89] Danielle Florens-Zmirou. Approximate discrete-time schemes for statistics of diffusion processes. Statistics: a Journal of Theoretical and Applied Statistics, 20(4):547–557, 1989. CO-DEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Fang:2017:SOM

[FZB17] Longxiang Fang, Xiaojun Zhu, and N. Balakrishnan. Stochastic ordering of minima and maxima from heterogeneous bivariate Birnbaum–Saunders random vectors. Statistics: a Journal of Theoretical and Applied Statistics, 52(1):147–155, 2017. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Fakhre-Zakeri:1997:LTS

[FZF97] Issa Fakhre-Zakeri and Jamshid Farshidi. Limit theorems for sample covariances of stationary linear processes with applications to sequential estimation. Statistics: a Journal of Theoretical and Applied Statistics, 29(3):251–260, 1997. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Fang:2022:OED

[FZLB22] Longxiang Fang, Shuai Zhang, Jinling Lu, and N. Balakrishnan. Orderings of extremes from dependent Gaussian variables with Archimedean copula under simple tree order restrictions. Statistics: a Journal of Theoretical and Applied Statistics, 56(1):134–146, 2022. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Green:1985:SRV

[GAB85] J. R. Green and M. F. Al-Bayatti. Selection of regress or variables when E(Y) is an unknown nonlinear function. Statistics: a Journal of Theoretical and Applied Statistics, 16(1):15–33, 1985. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Gaffke:1985:DDO

[Gaf85] N. Gaffke. Directional derivatives of optimality criteria at singular matrices in convex design theory. Statistics: a Journal of Theoretical and Applied Statistics, 16(3):373–388, 1985. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Gaigall:2023:AST

[Gai23] Daniel Gaigall. On the applicability of several tests to models with not identically distributed random effects. Statistics: a Journal of Theoretical and Applied Statistics, 57(2):300– 327, 2023. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-

4910 (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/02331888.2023.2193748.

Ghitany:2013:EUM

[GAJB13] M. E. Ghitany, R. A. Al-Jarallah, and N. Balakrishnan. On the existence and uniqueness of the MLEs of the parameters of a general class of exponentiated distributions. Statistics: a Journal of Theoretical and Applied Statistics, 47(3):605–612, 2013. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Gartner:1983:BR

[Gär83] J. Gärtner. Book reviews. Statistics (Berlin, DDR), 14(4): 516, 1983. CODEN MOSSD5. ISSN 0323-3944.

Gates:1994:DMV

[Gat94] J. Gates. Distance mean and variance functions and sphere fitting. Statistics: a Journal of Theoretical and Applied Statistics, 25(3):251–266, 1994. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Gatto:2009:ITR

[Gat09] Riccardo Gatto. Information theoretic results for circular distributions. Statistics: a Journal of Theoretical and Applied Statistics, 43(4):409–421, 2009. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Goldstein:1978:PSE

[GB78] M. Goldstein and P. J. Brown. Prediction with shrinkage estimators. Statistics (Berlin, DDR), 9(1):3–7, 1978. CODEN MOSSD5. ISSN 0323-3944.

Ghosh:2017:REP

[GB17] Abhik Ghosh and Ayanendranath Basu. Robust and efficient parameter estimation based on censored data with stochastic covariates. Statistics: a Journal of Theoretical and Applied Statistics, 51(4):801–823, 2017. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Godichon-Baggioni:2023:CQM

[GB23] Antoine Godichon-Baggioni. Convergence in quadratic mean of averaged stochastic gradient algorithms without strong convexity nor bounded gradient. Statistics: a Journal of Theoretical and Applied Statistics, 57(3):637–668, 2023. CO-DEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/02331888.2023.2213371.

Godichon-Baggioni:2020:RCP

[GBS20] Antoine Godichon-Baggioni and Sofiane Saadane. On the rates of convergence of parallelized averaged stochastic gradient algorithms. Statistics: a Journal of Theoretical and Applied Statistics, 54(3):618–635, 2020. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Genon-Catalot:1990:MCE

[GC90] V. Genon-Catalot. Maximum contrast estimation for diffusion processes from discrete observations. Statistics: a Journal of Theoretical and Applied Statistics, 21(1):99–116, 1990. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Grondona:1995:RBE

[GC95] Martin O. Grondona and Noel Cressie. Residuals based estimators of the covariogram. Statistics: a Journal of Theoretical and Applied Statistics, 26(3):209–218, 1995. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Govindarajulu:2004:SCI

[GC04] Z. Govindarajulu and Hokwon A. Cho. Sequential confidence intervals for the number of cells in a multinomial population. Statistics: a Journal of Theoretical and Applied Statistics, 38(4):357–370, 2004. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Goulard:1994:RBF

[GCB94] Michel Goulard, Joël Chadoeuf, and Patrick Bertuzzi. Random Boolean functions: Non-parametric estimation of the intensity. Application to soil surface roughness. Statistics: a Journal of Theoretical and Applied Statistics, 25(2):123–136, 1994. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Gomes:2004:BRT

[GCF04] M. Ivette Gomes, Frederico Caeiro, and Fernanda Figueiredo. Bias reduction of a tail index estimator through an external estimation of the second-order parameter. Statistics: a Journal of Theoretical and Applied Statistics, 38(6):497–510, 2004. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Genon-Catalot:2016:ESD

[GCL16] V. Genon-Catalot and C. Larédo. Estimation for stochastic differential equations with mixed effects. Statistics: a Journal of Theoretical and Applied Statistics, 50(5):1014–??, 2016. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Gupta:2007:MTF

[GCT07] Arjun K. Gupta, John T. Chen, and J. Tang. A multivariate two-factor skew model. Statistics: a Journal of Theoretical and Applied Statistics, 41(4):301–309, 2007. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Gill:1996:THS

[GD96] A. N. Gill and A. K. Dhawan. Testing homogeneity of scale parameters against ordered alternatives using Hodges– Lehmann estimators. Statistics: a Journal of Theoretical and Applied Statistics, 28(2):159–170, 1996. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Geblach:1984:CTG

[Geb84] B. Geblach. A correlation-type goodness- of-fit test for the logistic distribution with censored data. Statistics (Berlin, DDR), 15(3):373–381, 1984. CODEN MOSSD5. ISSN 0323-3944.

Genc:2013:GBS

[Gen13] Ali I. Genç. The generalized T Birnbaum–Saunders family. Statistics: a Journal of Theoretical and Applied Statistics, 47(3):613–625, 2013. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Gerencser:1973:CRB

[Ger73] L. Gerencsér. On a close relation between quasi-convex and convex functions and related investigations. Mathematische Operationsforschung und Statistik, 4(3):201–211, 1973. CO-DEN MTOSBI. ISSN 0047-6277.

Gerencser:1975:DMN

[Ger75] L. Gerencsér. A decomposition method in non-linear programming. Mathematische Operationsforschung und Statistik, 6(4):549–559, 1975. CODEN MTOSBI. ISSN 0047-6277.

Gerlach:1978:CAS

[Ger78a] B. Gerlach. A characterization of availability with statistical applications. Statistics (Berlin, DDR), 9(4):563–586, 1978. CODEN MOSSD5. ISSN 0323-3944.

Gerlach:1978:UST

[Ger78b] B. Gerlach. The unbiasedness of some tests for exponentiality. Statistics (Berlin, DDR), 9(3):405–410, 1978. CODEN MOSSD5. ISSN 0323-3944.

Gerlach:1979:CCT

[Ger79] B. Gerlach. A consistent correlation-type goodness-of-fit test;with application to the two-parameter Weibull distribution. Statistics (Berlin, DDR), 10(3):427–452, 1979. CODEN MOSSD5. ISSN 0323-3944.

Gerlach:1980:CTG

[Ger80] B. Gerlach. A correlation-type goodness-of-fit test for normality with censored sampling. Statistics (Berlin, DDR), 11 (2):207–218, 1980. CODEN MOSSD5. ISSN 0323-3944.

Gerlach:1986:NTW

[Ger86] B. Gerlach. A new test for whether F is “more NBU” than G. Statistics: a Journal of Theoretical and Applied Statistics, 17 (1):79–86, 1986. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Gerlach:1987:TEA

[Ger87] B. Gerlach. Testing exponentiality against increasing failure rate with randomly censored data. Statistics: a Journal of Theoretical and Applied Statistics, 18(2):275–286, 1987. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Gerlach:1989:TIF

[Ger89] B. Gerlach. Tests for increasing failure rate average with randomly right censored data. Statistics: a Journal of Theoretical and Applied Statistics, 20(2):287–295, 1989. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Gerlach:1991:BR

[Ger91] B. Gerlach. Book reviews. Statistics: a Journal of Theoretical and Applied Statistics, 22(1):94, 1991. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Grane:2009:LSF

[GF09] Aurea Grané and Josep Fortiana. A location- and scale-free goodness-of-fit statistic for the exponential distribution based on maximum correlations. Statistics: a Journal of Theoretical and Applied Statistics, 43(1):1–12, 2009. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Garcia-Finana:2004:IVP

[GFCO04] Marta García-Fiñana and Luis M. Cruz-Orive. Improved variance prediction for systematic sampling on R. Statistics: a Journal of Theoretical and Applied Statistics, 38(3):243–272, 2004. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Grutzmacher:1974:RGB

[GFFS74] G. Grützmacher, P. Franken, P. Franken, and D. Stoyan. Rezensionen. (German) [Book reviews]. Mathematische Operationsforschung und Statistik, 5(4–5):447–448, 1974. CODEN MTOSBI. ISSN 0047-6277.

Gupta:1990:EMN

[GG90] R. D. Gupta and Ramesh C. Gupta. Estimation of P r(a′x > b′y) in the multivariate normal case. Statistics: a Journal of Theoretical and Applied Statistics, 21(1):91–97, 1990. CO-DEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Gajek:1991:ESP

[GG91] L. Gajek and U. Gather. Estimating a scale parameter under censorship. Statistics: a Journal of Theoretical and Applied Statistics, 22(4):529–549, 1991. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Gillispie:2015:ACM

[GG15] Steven B. Gillispie and Christopher G. Green. Approximating the Conway–Maxwell–Poisson distribution normalization constant. Statistics: a Journal of Theoretical and Applied Statistics, 49(5):1062–1073, 2015. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Grasshoff:2003:OPC

[GGHS03] Ulrike Grasshoff, Heiko Grossmann, Heinz Holling, and Rainer Schwabe. Optimal paired comparison designs for firstorder interactions. Statistics: a Journal of Theoretical and Applied Statistics, 37(5):373–386, 2003. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Grasshoff:2021:ODP

[GGHS21] Ulrike Graßhoff, Heiko Großmann, Heinz Holling, and Rainer Schwabe. Optimal design for probit choice models with dependent utilities. Statistics: a Journal of Theoretical and Applied Statistics, 55(1):173–194, 2021. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Gallon:2015:FCE

[GGL15] S. Gallón, F. Gamboa, and J. M. Loubes. Functional calibration estimation by the maximum entropy on the mean principle. Statistics: a Journal of Theoretical and Applied Statistics, 49(5):989–1004, 2015. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Goegebeur:2014:NRE

[GGS14a] Yuri Goegebeur, Armelle Guillou, and Antoine Schorgen. Nonparametric regression estimation of conditional tails: the random covariate case. Statistics: a Journal of Theoretical and Applied Statistics, 48(4):732–755, 2014. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Grossmann:2014:CDP

[GGS14b] Heiko Großmann, Ulrike Graßhoff, and Rainer Schwabe. A catalogue of designs for partial profiles in paired comparison experiments with three groups of factors. Statistics: a Journal of Theoretical and Applied Statistics, 48(6):1268–1281, 2014. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Gebhardt:1983:SVR

[GH83] R. Gebhardt and H. Heckendorff. Zur sequentiellen Versuchsplanung für das Regressionsproblem. (German) [On sequential experimental planning for the regression problem]. Statistics (Berlin, DDR), 14(3):353–366, 1983. CODEN MOSSD5. ISSN 0323-3944.

Gaffke:1989:BAM

[GH89] Norbert Gaffke and Berthold Heiligers. Bayes, admissible, and minimax linear estimators in linear models with restricted parameter space. Statistics: a Journal of Theoretical and Applied Statistics, 20(4):487–508, 1989. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Gaffke:1991:NP

[GH91] Norbert Gaffke and Berthold Heiligers. Note on a paper. Statistics: a Journal of Theoretical and Applied Statistics, 22 (1):3–8, 1991. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Guyo:2001:MCM

[GH01] X. Guyo and C. Hardouin. Markow chain Markov field dynamics: models and statistics. Statistics: a Journal of Theoretical and Applied Statistics, 35(4):593–627, 2001. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Guyon:2002:MCM

[GH02] X. Guyon and C. Hardouin. Markov chain Markov field dynamics: Models and statistics. Statistics: a Journal of Theoretical and Applied Statistics, 36(4):339–363, 2002. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Gather:2002:NOS

[GHB02] Ursula Gather, Torsten Hilker, and Claudia Becker. A note on outlier sensitivity of sliced inverse regression. Statistics: a Journal of Theoretical and Applied Statistics, 36(4):271–281, 2002. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Gillert:1978:MLS

[Gil78] Heinz Gillert. Maximum-Likelihood-Schätzungen für Parameter in homogenen Markovschen Ketten. (German) [Maximumlikelihood estimates for parameters in homogeneous Markov chains]. Statistics (Berlin, DDR), 9(2):217–226, 1978. CO-DEN MOSSD5. ISSN 0323-3944.

Girlich:1971:NBO

[Gir71] Hans-Joachim Girlich. Ein Näherungsverfahren zur Bestimmung optimaler Parameter eines Lagerhaltungsmodells. (German) [An approximate method for determining optimum parameters of a storage model]. Mathematische Operationsforschung und Statistik, 2(1):41–49, 1971. CODEN MTOSBI. ISSN 0047-6277.

Girlich:1974:BSM

[Gir74] H. J. Girlich. Berichte über die Spezialtagung “Mathematische Lagerhaltungsmodelle und Anwendungen” vom 15.–20. Oktober 1973 in Weimar, DDR. (German) [Reports of the Special Meeting “Mathematical inventory models and applications” by 15.–20th October 1973 in Weimar, DDR]. Mathematische Operationsforschung und Statistik, 5(3):252, 1974. CODEN MTOSBI. ISSN 0047-6277.

Gupta:1973:SUL

[GJ73] R. P. Gupta and G. C. Jain. Some uses of Lagrange’s formula in the distribution theory with reference to queuing problems. Mathematische Operationsforschung und Statistik, 4(1): 69–73, 1973. CODEN MTOSBI. ISSN 0047-6277.

Grama:2019:EES

[GJ19] Ion Grama and Kévin Jaunâtre. Estimation of extreme survival probabilities with Cox model. Statistics: a Journal of Theoretical and Applied Statistics, 53(4):807–838, 2019. CO-DEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Ghale-Joogh:2018:TST

[GJHN18] Hassan Sharghi Ghale-Joogh and S. Mohammad E. Hosseini-Nasab. A two-sample test for mean functions with increasing number of projections. Statistics: a Journal of Theoretical and Applied Statistics, 52(4):852–873, 2018. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Gamboa:2016:SIS

[GJK^{+}16] F. Gamboa, A. Janon, T. Klein, A. Lagnoux, and C. Prieur. Statistical inference for Sobol pick-freeze Monte Carlo method. Statistics: a Journal of Theoretical and Applied Statistics, 50(4):881–902, 2016. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Grahovac:2015:APP

[GJLT15] Danijel Grahovac, Mofei Jia, Nikolai Leonenko, and Emanuele Taufer. Asymptotic properties of the partition function and applications in tail index inference of heavy-tailed data. Statistics: a Journal of Theoretical and Applied Statistics, 49 (6):1221–1242, 2015. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Goncalves:2000:DPB

[GJML00] E. Goncalves, P. Jacob, and N. Mendes-Lopes. A decision procedure for bilinear time series based on the asymptotic separation. Statistics: a Journal of Theoretical and Applied Statistics, 33(4):333–348, 2000. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Geerdens:2016:LSP

[GJV16] Candida Geerdens, Paul Janssen, and Noël Veraverbeke. Large sample properties of nonparametric copula estimators under bivariate censoring. Statistics: a Journal of Theoretical and Applied Statistics, 50(5):1036–??, 2016. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Gupta:1972:EGM

[GK72] R. P. Gupta and I. G. Kabe. Extensions of the general multivariate linear hypothesis model. Mathematische Operationsforschung und Statistik, 3(1):61–65, 1972. CODEN MTOSBI. ISSN 0047-6277.

Gaffke:1977:OPL

[GK77] N. Gaffke and O. Krafft. Optimum properties of latin square designs and a matrix inequality. Statistics (Berlin, DDR), 8 (3):345–350, 1977. CODEN MOSSD5. ISSN 0323-3944.

Gupta:1993:CSE

[GK93] Rameshwar D. Gupta and Ravindra Khattree. A comparison of some estimators of the variance covariance matrix when the population mean is known. Statistics: a Journal of Theoretical and Applied Statistics, 24(4):331–345, 1993. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Gajek:1995:NAG

[GK95] Leslaw Gajek and Marek Kaluszka. Nonexponential applications of a global Cramèr–Rao inequality. Statistics: a Journal of Theoretical and Applied Statistics, 26(2):111–122, 1995. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Gupta:1998:PMR

[GK98] Ramesh C. Gupta and S. N. U. A. Kirmani. On the proportional mean residual life model and its implications. Statistics: a Journal of Theoretical and Applied Statistics, 32(2): 175–187, 1998. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Gupta:1999:DHM

[GK99] A. K. Gupta and D. G. Kabe. Distributions of Hotelling’s T^{2} and multiple and partial correlation coefficients for the mixture of two multivariate Gaussian populations. Statistics: a Journal of Theoretical and Applied Statistics, 32(4):331– 339, 1999. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Gupta:2009:NCW

[GK09] Rameshwar D. Gupta and Debasis Kundu. A new class of weighted exponential distributions. Statistics: a Journal of Theoretical and Applied Statistics, 43(6):621–634, 2009. CO-DEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Ghosh:2017:SDG

[GK17] Amit Ghosh and Chanchal Kundu. On some dynamic generalized measures of information for conditionally specified models in past life. Statistics: a Journal of Theoretical and Applied Statistics, 51(6):1398–1418, 2017. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Glueck:2010:PSS

[GKFMM10] D. H. Glueck, A. Karimpour-Fard, J. Mandel, and K. E. Muller. Probabilities for separating sets of order statistics. Statistics: a Journal of Theoretical and Applied Statistics, 44(2):145–153, 2010. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Grover:2018:ALS

[GKM18] Rhythm Grover, Debasis Kundu, and Amit Mitra. On approximate least squares estimators of parameters of onedimensional chirp signal. Statistics: a Journal of Theoretical and Applied Statistics, 52(5):1060–1085, 2018. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Gather:1996:AEN

[GKP96] Ursula Gather, Stephan Köhne, and Iris Pigeot. Asymptotically efficient noniterative estimators of a common parameter from independent samples. Statistics: a Journal of Theoretical and Applied Statistics, 28(3):187–200, 1996. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Gnot:1992:NUE

[GKTZ92] S. Gnot, H. Knautz, G. Trenkler, and R. Zmyslony. Nonlinear unbiased estimation in linear models. Statistics: a Journal of Theoretical and Applied Statistics, 23(1):5–16, 1992. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Gnot:1977:UMV

[GKZ77] Stanisław Gnot, Witold Klonecki, and Roman Zmyślony. Uniformly minimum variance unbiased estimation in various classes of estimators, I. Statistics (Berlin, DDR), 8(2):199– 210, 1977. CODEN MOSSD5. ISSN 0323-3944.

Gaudoin:1994:ETV

[GL94] Olivier Gaudoin and Christian Lavergne. Estimation in a two variance components model when one component is known. Statistics: a Journal of Theoretical and Applied Statistics, 25 (2):97–106, 1994. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Gladitz:1976:EBT

[Gla76] J. Gladitz. Empirisch beste Tests bei zufälligem Parameter und ihre Anwendung in der statistischen Qualitätskontrolle. (German) [Empirically best tests for accidental parameters and their application in statistical quality control]. Mathematische Operationsforschung und Statistik, 7(5):685–696, 1976. CODEN MTOSBI. ISSN 0047-6277.

Glanzel:1990:SCC

[Gla90] Wolfgang Glanzel. Some consequences of a characterization theorem based on truncated moments. Statistics: a Journal of Theoretical and Applied Statistics, 21(4):613–618, 1990. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Gomtsyan:2022:VSS

[GLLO^{+}22] Marina Gomtsyan, Céline Lévy-Leduc, Sarah Ouadah, Laure Sansonnet, and Thomas Blein. Variable selection in sparse GLARMA models. Statistics: a Journal of Theoretical and Applied Statistics, 56(4):755–784, 2022. CO-DEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/02331888.2022.2090943.

Gloter:2001:PED

[Glo01] Arnaud Gloter. Parameter estimation for a discrete sampling of an integrated Ornstein–Uhlenbeck process. Statistics: a Journal of Theoretical and Applied Statistics, 35(3):225–243, 2001. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Garnier:2023:HEP

[GLR23] Remy Garnier, Raphaël Langhendries, and Joseph Rynkiewicz. Hold-out estimates of prediction models for Markov processes. Statistics: a Journal of Theoretical and Applied Statistics, 57(2):458–481, 2023. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/02331888.2023.2183203.

Gyorfi:1998:DEC

[GLVV98] László Györfi, Friedrich Liese, Igor Vajda, and Edward C. Van Der Meulen. Distribution estimates consistent in χ^{2}- divergence. Statistics: a Journal of Theoretical and Applied Statistics, 32(1):31–57, 1998. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Gao:2012:APM

[GLZW12] Qi-Bing Gao, Jin-Guan Lin, Chun-Hua Zhu, and Yao-Hua Wu. Asymptotic properties of maximum quasi-likelihood estimators in generalized linear models with adaptive designs. Statistics: a Journal of Theoretical and Applied Statistics, 46(6):833–846, 2012. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Grecksch:1978:NEO

[GM78] Wilfried Grecksch and Heinz Müller. Zur näherungsweisen Ermittlung optimaler stochastischer Steuerungen durch Diskretisierung. (German) [On the approximate determination of optimal stochastic controls by discretization]. Statistics (Berlin, DDR), 9(3):383–393, 1978. CODEN MOSSD5. ISSN 0323-3944.

Gnot:1988:EPU

[GM88] Stanisław Gnot and Andrzej Michalski. Estimation of parameters under Hardy–Weinberg law in population genetics. Statistics: a Journal of Theoretical and Applied Statistics, 19(3):477–487, 1988. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Gnot:1991:LQE

[GM91] S. Gnot and A. Michalski. Linear and quadratic estimation from inter- and intra-block sources of information. Statistics: a Journal of Theoretical and Applied Statistics, 22(1):17–32, 1991. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Ghitany:1992:ARE

[GM92] M. E. Ghitany and R. A. Maller. Asymptotic results for exponential mixture models with long-term survivors. Statistics: a Journal of Theoretical and Applied Statistics, 23(4):321–336, 1992. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Gnot:1994:TBA

[GM94] S. Gnot and A. Michalski. Tests based on admissible estimators in two variance components models. Statistics: a Journal of Theoretical and Applied Statistics, 25(3):213–223, 1994. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Goncalves:1996:SGP

[GM96] E. Goncalves and N. Mendes Lopes. Stationarity of GTARCH processes. Statistics: a Journal of Theoretical and Applied Statistics, 28(2):171–178, 1996. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Guorui:1997:BIB

[GM97] Bian Guorui and L. Tiku Moti. Bayesian inference based on robust priors and MML estimators: Part II, skew locationscale distributions. Statistics: a Journal of Theoretical and Applied Statistics, 29(1):81–99, 1997. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Gross:1999:ALE

[GM99] Jürgen Groß and Augustyn Markiewicz. On admissibility of linear estimators with respect to the mean square error matrix criterion under the general mixed linear model. Statistics: a Journal of Theoretical and Applied Statistics, 33(1):57–71, 1999. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Gomes:2018:TSG

[GMCT18] Amanda S. Gomes, Pedro A. Morettin, Gauss M. Cordeiro, and Marcelo M. Taddeo. Transformed symmetric generalized autoregressive moving average models. Statistics: a Journal of Theoretical and Applied Statistics, 52(3):643–664, 2018. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Goncalves:2018:ZTC

[GML18] Esmeralda Gonçalves and Nazaré Mendes-Lopes. Zerotruncated compound Poisson integer-valued GARCH models for time series. Statistics: a Journal of Theoretical and Applied Statistics, 52(3):619–642, 2018. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Goncalves:2023:ZIB

[GML23] Esmeralda Gonçalves and Nazaré Mendes-Lopes. Zeroinflated binomial integer-valued ARCH models for time series. Statistics: a Journal of Theoretical and Applied Statistics, 57(4):764–784, 2023. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/02331888.2023.2221363.

Goncalves:2016:ZIC

[GMLS16] Esmeralda Gonçalves, Nazaré Mendes-Lopes, and Filipa Silva. Zero-inflated compound Poisson distributions in integer-valued GARCH models. Statistics: a Journal of Theoretical and Applied Statistics, 50(3):558–578, 2016. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Galbete:2016:RTR

[GMP16] A. Galbete, J. A. Moler, and F. Plo. Randomization tests in recursive response-adaptive randomization procedures. Statistics: a Journal of Theoretical and Applied Statistics, 50(2):418–434, 2016. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic). URL http://

www.tandfonline.com/doi/abs/10.1080/02331888.2015.
1050020.

Gupta:1998:QFD

[GN98] A. K. Gupta and D. K. Nagar. Quadratic forms in disguised matrix T -variate. Statistics: a Journal of Theoretical and Applied Statistics, 30(4):357–374, 1998. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Govindarajulu:2000:SEM

[GN00] Z. Govindarajulu and A. Nanthakumar. Sequential estimation of the mean logistic response function. Statistics: a Journal of Theoretical and Applied Statistics, 33(4):309–332, 2000. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Gijbels:2016:STD

[GN16] Irène Gijbels and Stanislav Nagy. On smoothness of Tukey depth contours. Statistics: a Journal of Theoretical and Applied Statistics, 50(5):1075–??, 2016. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Graf:2017:DCD

[GN17] Monique Graf and Desislava Nedyalkova. Discretizing a compound distribution with application to categorical modelling. Statistics: a Journal of Theoretical and Applied Statistics, 51(3):685–710, 2017. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Ghannam:2022:RTS

[GN22] Mai Ghannam and Sévérien Nkurunziza. The risk of tensor Stein-rules in elliptically contoured distributions. Statistics: a Journal of Theoretical and Applied Statistics, 56(2):421–454, 2022. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Gneiting:1998:URP

[Gne98] Tilmann Gneiting. On the uncertainty relation for positive definite probability densities. Statistics: a Journal of Theoretical and Applied Statistics, 31(1):83–88, 1998. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Gnot:1976:MEB

[Gno76] S. Gnot. The mean efficiency of block designs. Mathematische Operationsforschung und Statistik, 7(1):75–84, 1976. CODEN MTOSBI. ISSN 0047-6277.

Gnot:1977:ECC

[Gno77] Stanisław Gnot. The essentially complete class of rules in multinomial identification. Statistics (Berlin, DDR), 8(3): 381–386, 1977. CODEN MOSSD5. ISSN 0323-3944.

Gnot:1978:PTG

[Gno78] Stanisław Gnot. The problem of two-group identification. Statistics (Berlin, DDR), 9(3):343–349, 1978. CODEN MOSSD5. ISSN 0323-3944.

Gnot:1979:EPC

[Gno79] Stanisaw Gnot. Estimation of parameters in certain hierarchical genetic models. Statistics (Berlin, DDR), 10(4):529–539, 1979. CODEN MOSSD5. ISSN 0323-3944.

Gupta:2008:RPL

[GNP08] A. K. Gupta, T. Nguyen, and L. Pardo. Residuals for polytomous logistic regression models based on ϕ-divergences test statistics. Statistics: a Journal of Theoretical and Applied Statistics, 42(6):495–514, 2008. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Gladitz:2015:OHB

[GNS15] Johannes Gladitz, Michael Nussbaum, and Zwanzig Silvelyn. Obituary for Helga Bunke (13 July 1938–4 May 2014). Statistics: a Journal of Theoretical and Applied Statistics, 49(6): 1187–1188, 2015. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Guo:2015:ELS

[GNYX15] Xu Guo, Cuizhen Niu, Yiping Yang, and Wangli Xu. Empirical likelihood for single index model with missing covariates at random. Statistics: a Journal of Theoretical and Applied Statistics, 49(3):588–601, 2015. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Goel:1975:LSF

[Goe75] L. R. Goel. A limited space, fluctuating (0, λ) input source queueing problem with death and birth rates of the input source depending on queue length. Mathematische Operationsforschung und Statistik, 6(3):437–444, 1975. CODEN MTOSBI. ISSN 0047-6277.

Goel:1976:HQA

[Goe76] L. R. Goel. Heterogeneous queueing with arrivals depending on queue length. Mathematische Operationsforschung und Statistik, 7(6):945–952, 1976. CODEN MTOSBI. ISSN 0047-6277.

Goldmank:1991:NSP

[Gol91] Werner Goldmank. A nonparametric sequential point estimator. Statistics: a Journal of Theoretical and Applied Statistics, 22(1):133–141, 1991. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Gomez:2018:DLC

[Góm18] José G. Gómez. Dependent Lindeberg central limit theorem for the fidis of empirical processes of cluster functionals. Statistics: a Journal of Theoretical and Applied Statistics, 52(5):955–979, 2018. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Gupta:1995:IEG

[GON95a] A. K. Gupta and S. Ofori-Nyarko. Improved estimation of generalized variance and precision. Statistics: a Journal of Theoretical and Applied Statistics, 26(2):99–109, 1995. CO-DEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Gupta:1995:IME

[GON95b] A. K. Gupta and Samuel Ofori-Nyarko. Improved minimax estimators of normal covariance and precision matrices. Statistics: a Journal of Theoretical and Applied Statistics, 26(1): 19–25, 1995. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Goroncy:2014:BEG

[Gor14] Agnieszka Goroncy. Bounds on expected generalized order statistics. Statistics: a Journal of Theoretical and Applied Statistics, 48(3):593–608, 2014. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Goroncy:2019:OUB

[Gor19] Agnieszka Goroncy. Optimal upper bounds on expected k-th record values from IGFR distributions. Statistics: a Journal of Theoretical and Applied Statistics, 53(5):1012–1036, 2019. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Goroncy:2020:UBE

[Gor20] Agnieszka Goroncy. On upper bounds on expectations of gOSs based on DFR and DFRA distributions. Statistics: a Journal of Theoretical and Applied Statistics, 54(2):402–414, 2020. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Golosnoy:2010:NCP

[GOS10] Vasyl Golosnoy, Iryna Okhrin, and Wolfgang Schmid. New characteristics for portfolio surveillance. Statistics: a Journal of Theoretical and Applied Statistics, 44(3):303–321, 2010. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Gottinger:1974:KSW

[Got74] Hens-Werner Gottinger. Konstruktion Subjektiver Wahrscheinlichkeiten. (German) [Construction of subjective probabilities]. Mathematische Operationsforschung und Statistik, 5(6): 509–539, 1974. CODEN MTOSBI. ISSN 0047-6277.

Gottinger:1976:DSD

[Got76] Hans W. Gottinger. Decomposition for stochastic dynamic systems. Mathematische Operationsforschung und Statistik, 7 (6):889–903, 1976. CODEN MTOSBI. ISSN 0047-6277.

Govindarajulu:1992:SPO

[Gov92] Z. Govindarajulu. The secretary problem: optimal selection with interview cost from two groups of candidates. Statistics: a Journal of Theoretical and Applied Statistics, 23(2):171– 184, 1992. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Govindarajulu:2011:BSS

[Gov11] Z. Govindarajulu. Blinded sample size re-estimation in clinical trials comparing several treatments. Statistics: a Journal of Theoretical and Applied Statistics, 45(6):575–591, 2011. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Gijbels:2017:NTN

[GOV17] Irène Gijbels, Marek Omelka, and Noël Veraverbeke. Nonparametric testing for no covariate effects in conditional copulas. Statistics: a Journal of Theoretical and Applied Statistics, 51(3):475–509, 2017. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Gladitz:1982:BDM

[GP82a] J. Gladitz and J. Pilz. Bayes designs for multiple linear regression on the unit sphere. Statistics (Berlin, DDR), 13(4): 491–506, 1982. CODEN MOSSD5. ISSN 0323-3944.

Gladitz:1982:COD

[GP82b] J. Gladitz and J. Pilz. Construction of optimal designs in random coefficient regression models. Statistics (Berlin, DDR), 13(3):371–385, 1982. CODEN MOSSD5. ISSN 0323-3944.

Guyon:2007:STE

[GP07] Xavier Guyon and Besnik Pumo. Space-time estimation of a particle system model. Statistics: a Journal of Theoretical and Applied Statistics, 41(5):395–407, 2007. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Giordano:2019:ENE

[GP19] Francesco Giordano and Maria Lucia Parrella. Efficient nonparametric estimation and inference for the volatility function. Statistics: a Journal of Theoretical and Applied Statistics, 53(4):770–791, 2019. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Gupta:2023:ENS

[GPK23] Tulika Rudra Gupta, Markus Pauly, and Somesh Kumar. Estimating a new stress-strength index for several exponential populations with a common location. Statistics: a Journal of Theoretical and Applied Statistics, 57(3):669–693, 2023.

CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/02331888.2023.2203492.

Ghosh:2014:FSE

[GPS14] Himadri Ghosh, Prajneshu, and Sandipan Samanta. Fitting of self-exciting threshold autoregressive moving average nonlinear time-series model through genetic algorithm and development of out-of-sample forecasts. Statistics: a Journal of Theoretical and Applied Statistics, 48(5):1166–1184, 2014. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Grossmann:1972:ZNO

[GR72] Christian Grossmann and Rüdiger Reinhardt. Zur zentrenmethode in der nichtlinearen optimierung. (German) [On the centers method in nonlinear optimization]. Mathematische Operationsforschung und Statistik, 3(6):419–430, 1972. CO-DEN MTOSBI. ISSN 0047-6277.

Gopfert:1976:LKL

[GR76] A. Göpfert and H. Rudolph. Lineare Komplementaritätsprobleme in lokalkonvexen Räumen. (German) [Linear complementarity problems in locally convex spaces]. Mathematische Operationsforschung und Statistik, 7(2):295–300, 1976. CODEN MTOSBI. ISSN 0047-6277.

Goroncy:2006:LBE

[GR06] Agnieszka Goroncy and Tomasz Rychlik. Lower bounds on expectations of positive L-statistics based on samples drawn with replacement. Statistics: a Journal of Theoretical and Applied Statistics, 40(5):389–408, 2006. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Gietl:2017:CPA

[GR17] Christoph Gietl and Fabian P. Reffel. Continuity of fprojections and applications to the iterative proportional fitting procedure. Statistics: a Journal of Theoretical and Applied Statistics, 51(3):668–684, 2017. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Grecksch:1987:GZS

[Gre87] W. Grecksch. Über ein gesteuertes zweiparametrisches stochastisches Differenzensystem. (German) [On a controlled two-parametric stochastic differences system]. Statistics: a Journal of Theoretical and Applied Statistics, 18(2):293–308, 1987. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Grecksch:1988:BR

[Gre88] W. Grecksch. Book review. Statistics: a Journal of Theoretical and Applied Statistics, 19(1):122, 1988. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Griessbach:1980:ABO

[Gri80] G. Griessbach. Zur adaptiven Bestimmung der optimalen Strategieparameter in stationären Lagerhaltungsmodellen. (German) [On adaptively determining the optimal strategy parameters in stationary inventory models]. Statistics (Berlin, DDR), 11(4):521–544, 1980. CODEN MOSSD5. ISSN 0323-3944.

Grossmann:1982:SEN

[Gro82] Wilfried Grossmann. Statistical estimation of nonlinear regression functions. Statistics (Berlin, DDR), 13(3):455–471, 1982. CODEN MOSSD5. ISSN 0323-3944.

Grund:1971:SWS

[Gru71] J. Grund. Der Shapleyysche Wertvektor und seine Anwendung in einigen Marktmodellen. (German) [The Shapley value vector and its application in some market models]. Mathematische Operationsforschung und Statistik, 2(1):61–67, 1971. CODEN MTOSBI. ISSN 0047-6277.

Gopfert:1975:TSH

[GRV75] A. Göpfert, H. Rudolph, and H. Voigt. Der Tuckersche Satz in Hilbert-Räumen. (German) [The Tucker theorem in Hilbert spaces]. Mathematische Operationsforschung und Statistik, 6 (4):571–580, 1975. CODEN MTOSBI. ISSN 0047-6277.

Gupta:1982:EPC

[GS82] Ramesh C. Gupta and Jagbib Sihgh. Estimation of probabilities in the class of modified power series distributions. Statistics (Berlin, DDR), 13(1):71–77, 1982. CODEN MOSSD5. ISSN 0323-3944.

Ginot:1988:MBQ

[GS88] S. Ginot and J. Sezedinicka. On minimum biased quadratic estimators. Statistics: a Journal of Theoretical and Applied Statistics, 19(2):215–222, 1988. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Gupta:2004:EMN

[GS04] Arjun K. Gupta and Yo Sheena. Estimation of a multivariate normal covariance matrix under a certain structure. Statistics: a Journal of Theoretical and Applied Statistics, 38(5):371– 379, 2004. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Geenens:2011:SIM

[GS11] Gery Geenens and Léopold Simar. Single-index modelling of conditional probabilities in two-way contingency tables. Statistics: a Journal of Theoretical and Applied Statistics, 45(5):451–478, 2011. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Gabler:2012:DER

[GS12] Siegfried Gabler and Horst Stenger. Design effect of randomized systematic sampling. Statistics: a Journal of Theoretical and Applied Statistics, 46(1):131–148, 2012. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Gomez:2014:CBD

[GS14] Guadalupe Gómez and Carles Serrat. Correcting the bias due to dependent censoring of the survival estimator by conditioning. Statistics: a Journal of Theoretical and Applied Statistics, 48(2):295–314, 2014. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Grublyte:2017:GNM

[GS17] Ieva Grublyte and Andrius Skarnulis. A generalized nonlinear model for long memory conditional heteroscedasticity. Statistics: a Journal of Theoretical and Applied Statistics, 51(1): 123–140, 2017. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Golosnoy:2021:MMC

[GS21] Vasyl Golosnoy and Miriam Isabel Seifert. Monitoring mean changes in persistent multivariate time series. Statistics: a Journal of Theoretical and Applied Statistics, 55(3):475–488, 2021. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Georgiou:2014:ETL

[GSA14] Stelios D. Georgiou, Stella Stylianou, and Manohar Aggarwal. Efficient three-level screening designs using weighing matrices. Statistics: a Journal of Theoretical and Applied Statistics, 48(4):815–833, 2014. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Gomez:2006:GSN

[GSB06] Héctor W. Gómez, Hugo S. Salinas, and Heleno Bolfarine. Generalized skew-normal models: properties and inference. Statistics: a Journal of Theoretical and Applied Statistics, 40(6):495–505, 2006. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic). See erratum [GSB11].

Gomez:2011:EGS

[GSB11] Héctor W. Gómez, Hugo S. Salinas, and Heleno Bolfarine. Erratum: Generalized skew-normal models: properties and inference [Statistics 40 (2006) 495–505]. Statistics: a Journal of Theoretical and Applied Statistics, 45(3):307–308, 2011. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic). See [GSB06].

Gnot:2002:MLE

[GSTUM02] S. Gnot, D. Stemann, G. Trenkler, and A. Urbanśka-Motyka. Maximum likelihood estimation in mixed normal models with two variance components. Statistics: a Journal of Theoretical and Applied Statistics, 36(4):283–302, 2002. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Guddat:1970:MMT

[GT70] Jürgen Guddat and Klaus Tammer. Eine Modification der Methode von Theil und van de Panne zur Lösung einparametrischer quadratischer Optimierungsprobleme. (German) [A modification of the method of Theil and van de Panne for solving one-parameter quadratic programming problems].

Mathematische Operationsforschung und Statistik, 1(3):199– 206, 1970. CODEN MTOSBI. ISSN 0047-6277.

Grane:2013:APG

[GT13] Aurea Grané and Anna V. Tchirina. Asymptotic properties of a goodness-of-fit test based on maximum correlations. Statistics: a Journal of Theoretical and Applied Statistics, 47(1): 202–215, 2013. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Guddat:1970:BPQ

[Gud70] Jürgen Guddat. Behandlung parametrischer quadratischer Optimierungsprobleme mit der Methode von Beale. (German) [Treatment of parametric quadratic programming problems with the method of Beale]. Mathematische Operationsforschung und Statistik, 1(2):81–86, 1970. CODEN MTOSBI. ISSN 0047-6277.

Guddat:1976:SCQ

[Gud76] Jürgen Guddat. Stability in convex quadratic parametric programming. Mathematische Operationsforschung und Statistik, 7(2):223–245, 1976. CODEN MTOSBI. ISSN 0047-6277.

Guiard:1986:GFC

[Gui86] V. Guiard. A general formula for the central mixed moments of the multivariate normal distribution. Statistics: a Journal of Theoretical and Applied Statistics, 17(2):279–289, 1986. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Guiard:1991:BR

[Gui91] V. Guiard. Book reviews. Statistics: a Journal of Theoretical and Applied Statistics, 22(4):626, 1991. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Gunther:1977:AVS

[Gün77] Roland Günther. Zur adaptiven Vorhersage von stationären zufälligen Folgen mit rationaler Spektraldichte. (German) [On adaptive prediction of stationary random sequences with rational spectral density]. Statistics (Berlin, DDR), 8(4):467– 482, 1977. CODEN MOSSD5. ISSN 0323-3944.

Gunther:1980:AEG

[Gün80] Roland Günther. Adaptive exponentielle Glättung erster Ordnung. (German) [Adaptive exponential smoothing of first order]. Statistics (Berlin, DDR), 11(3):327–340, 1980. CODEN MOSSD5. ISSN 0323-3944.

Gunther:1983:CSA

[Gün83] Roland Günther. On the convergence of some adaptive estimation procedures. Statistics (Berlin, DDR), 14(4):535–550, 1983. CODEN MOSSD5. ISSN 0323-3944.

Gunther:1988:AEP

[Gun88a] Roland Gunther. Adaptive estimation and prediction for an extension of the antoregressive multiple time series model. Statistics: a Journal of Theoretical and Applied Statistics, 19(4):557–564, 1988. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Gunther:1988:AEE

[Gun88b] Roland Gunther. Adaptive estimation for an extension of the autoregressive time series model. Statistics: a Journal of Theoretical and Applied Statistics, 19(4):545–555, 1988. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Guo:2020:BBQ

[Guo20] Guangbao Guo. A block bootstrap for quasi-likelihood in sparse functional data. Statistics: a Journal of Theoretical and Applied Statistics, 54(5):909–925, 2020. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Gupta:1977:CDR

[Gup77] Ramesh C. Gupta. On characterizing distributions by the ration of variance and mean. Statistics (Berlin, DDR), 8(4): 523–527, 1977. CODEN MOSSD5. ISSN 0323-3944.

Gupta:1979:NMG

[Gup79a] Ramesh C. Gupta. On negative moments of generalized Poisson distribution. Statistics (Berlin, DDR), 10(1):169–172, 1979. CODEN MOSSD5. ISSN 0323-3944.

Gupta:1979:OSE

[Gup79b] Ramesh C. Gupta. The order statistics of exponential, power function and Pareto distributions and some applications. Statistics (Berlin, DDR), 10(4):551–554, 1979. CODEN MOSSD5. ISSN 0323-3944.

Gupta:1982:PGF

[Gup82] Pushpa Lata Gupta. Probability generating functions of a MPSD with applications. Statistics (Berlin, DDR), 13(1): 99–103, 1982. CODEN MOSSD5. ISSN 0323-3944.

Gupta:1984:SCR

[Gup84] Ramesh C. Gupta. Some characterizations of renewal densities with emphasis in reliability. Statistics (Berlin, DDR), 15 (4):571–579, 1984. CODEN MOSSD5. ISSN 0323-3944.

Gupta:1985:SCD

[Gup85] Pushpa Lata Gupta. Some characterizations of distributions by truncated moments. Statistics: a Journal of Theoretical and Applied Statistics, 16(3):465–473, 1985. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Gupta:1995:STD

[Gup95] Ramesh C. Gupta. On some transforms of distribution functions and their relationships. Statistics: a Journal of Theoretical and Applied Statistics, 26(2):123–128, 1995. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Gupta:2003:MSD

[Gup03] A. K. Gupta. Multivariate skew t-distribution. Statistics: a Journal of Theoretical and Applied Statistics, 37(4):359–363, 2003. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Gupta:2005:RFB

[Gup05] R. C. Gupta. Reliability functions of bivariate distributions with second kind of beta conditionals. Statistics: a Journal of Theoretical and Applied Statistics, 39(1):43–52, 2005. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Gussefeldt:1971:SBB

[Gus71] Gisela Gussefeldt. Über das Stetigkeitsverhalten von “bangbang”-Steuerungen: Bemerkungen zu meiner Arbeit von C. Olech. (German) [On the continuity behavior of “bang-bang” controls — comments on my work with C. Olech]. Mathematische Operationsforschung und Statistik, 2(3):221–226, 1971. CODEN MTOSBI. ISSN 0047-6277.

Guven:2006:LDS

[Güv06] Bilgehan Güven. The limiting distribution of the F -statistic from nonnormal universes. Statistics: a Journal of Theoretical and Applied Statistics, 40(6):545–557, 2006. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Guven:2014:TRE

[Güv14] Bilgehan Güven. Testing for random effect in the Fuller– Battese model. Statistics: a Journal of Theoretical and Applied Statistics, 48(4):802–814, 2014. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Gillert:1979:GTH

[GV79] H. Gillert and J. Vogel. Über Güteeigenschaften von Tests bei homogenen Markovschen Prozessen. (German) [On goodness properties of tests for homogeneous Markov processes]. Statistics (Berlin, DDR), 10(1):141–152, 1979. CODEN MOSSD5. ISSN 0323-3944.

Gillert:1981:RTH

[GV81] Heinz Gillert and Jürgen Vogel. Remarks on testing hypotheses for Markov processes. Statistics (Berlin, DDR), 12(1): 53–59, 1981. CODEN MOSSD5. ISSN 0323-3944.

Gupta:1995:SIP

[GV95] A. K. Gupta and T. Varga. Some inference problems for matrix variate elliptically contoured distributions. Statistics: a Journal of Theoretical and Applied Statistics, 26(3):219–229, 1995. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Gurevich:2006:GML

[GV06] Gregory Gurevich and Albert Vexler. Guaranteed maximum likelihood splitting tests of a linear regression model. Statistics: a Journal of Theoretical and Applied Statistics, 40(6): 465–484, 2006. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Gijbels:2019:REV

[GV19] I. Gijbels and I. Vrinssen. Robust estimation and variable selection in heteroscedastic linear regression. Statistics: a Journal of Theoretical and Applied Statistics, 53(3):489–532, 2019. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Gomez-Villegas:2013:IPM

[GVGP13] Miguel A. Gómez-Villegas and Beatriz González-Pérez. Importance of the prior mass for agreement between frequentist and Bayesian approaches in the two-sided test. Statistics: a Journal of Theoretical and Applied Statistics, 47(3):558–565, 2013. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Gomez-Villegas:2009:BAM

[GVMS09] Miguel A. Gómez-Villegas, Paloma Maín, and Luis Sanz. A Bayesian analysis for the multivariate point null testing problem. Statistics: a Journal of Theoretical and Applied Statistics, 43(4):379–391, 2009. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Gomez:2013:NCS

[GVV13] Héctor W. Gómez, Héctor Varela, and Ignacio Vidal. A new class of skew-symmetric distributions and related families. Statistics: a Journal of Theoretical and Applied Statistics, 47(2):411–421, 2013. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Gijbels:2017:CRP

[GVV17] I. Gijbels, A. Verhasselt, and I. Vrinssen. Consistency and robustness properties of the S-nonnegative garrote estimator. Statistics: a Journal of Theoretical and Applied Statistics, 51(4):921–947, 2017. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Gerlach:1976:SRS

[GW76] B. Gerlach and W. Warmuth. Some remarks to system availability. Mathematische Operationsforschung und Statistik, 7 (4):601–606, 1976. CODEN MTOSBI. ISSN 0047-6277.

Gillert:1984:DEN

[GW84] H. Gillert and A. Wartenberg. Density estimation for nonstationary Markov processes. Statistics (Berlin, DDR), 15 (2):263–275, 1984. CODEN MOSSD5. ISSN 0323-3944.

Gromping:2004:ADP

[GW04] Ulrike Grömping and Ulla Weimann. The asymptotic distribution of the partial attributable risk in cross-sectional studies. Statistics: a Journal of Theoretical and Applied Statistics, 38(5):427–438, 2004. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Geng:2023:BEL

[GZ23] Shuli Geng and Lixin Zhang. Bayesian estimation for longitudinal data in a joint model with HPCs. Statistics: a Journal of Theoretical and Applied Statistics, 57(2):375–387, 2023. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/02331888.2023.2185243.

Hartung:2002:GWT

[HA02] Joachim Hartung and Doğan Argaç. Generalizing the Welch test to non-zero hypotheses on the variance component in the one-way random effects model under variance heterogeneity. Statistics: a Journal of Theoretical and Applied Statistics, 36 (2):89–99, 2002. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Habib:1989:RCE

[Hab89] Mohameb G. Habib. On root-n consistency of estimators for nuisance parameters of asymptotic optimal tests: the i.i.d. case. Statistics: a Journal of Theoretical and Applied Statistics, 20(3):377–381, 1989. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Hafner:1975:KSS

[Haf75] Robert Hafner. Kolmogorov–Smirnov-statistics under the alternative. Mathematische Operationsforschung und Statistik, 6(5):787–796, 1975. CODEN MTOSBI. ISSN 0047-6277.

Hafner:1982:CLF

[Haf82a] Robert Hafner. Construction of least favourable pairs of distributions and of robust tests for contamination neighbourhoods. Statistics (Berlin, DDR), 13(1):47–56, 1982. CODEN MOSSD5. ISSN 0323-3944.

Hafner:1982:SCL

[Haf82b] Robert Hafner. Simple construction of least favourable pairs of distributions and of robust tests for Prokhorovneighbourhoods. Statistics (Berlin, DDR), 13(1):33–46, 1982. CODEN MOSSD5. ISSN 0323-3944.

Hajstcsch:1984:SRE

[Haj84] K. H. Hajstcsch. Some remarks on estimators of the distribution function of nearest neighbour distance in stationary spatial point processes. Statistics (Berlin, DDR), 15(3):409– 412, 1984. CODEN MOSSD5. ISSN 0323-3944.

Hall:1991:EEN

[Hal91] Peter Hall. Edgeworth expansions for nonparametric density estimators, with applications. Statistics: a Journal of Theoretical and Applied Statistics, 22(2):215–232, 1991. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Hampel:1978:MTT

[Ham78] Frank Hampel. Modern trends in the theory of robustness. Statistics (Berlin, DDR), 9(3):425–442, 1978. CODEN MOSSD5. ISSN 0323-3944.

Hamdy:1996:CAS

[HAMAZ96] H. I. Hamdy, M. Al-Mahmeed, and Y. Al-Zalzala. A certain accelerating sequential procedure to construct simultaneous confidence region: the exponential case. Statistics: a Journal of Theoretical and Applied Statistics, 28(1):57–71, 1996. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Hanisch:1983:RMM

[Han83] K.-H. Hanisch. Reduction of n-th moment measures and the special case of the third moment measure of stationary and isotropic planar point processes. Statistics (Berlin, DDR), 14 (3):421–435, 1983. CODEN MOSSD5. ISSN 0323-3944.

Hanagal:1996:UTT

[Han96] David D. Hanagal. Umpu tests for testing symmetry and stress-passing in some bivariate exponential models. Statistics: a Journal of Theoretical and Applied Statistics, 28(3):

227–239, 1996. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Hartmann:1975:GGL

[Har75a] Karl Hartmann. Gemischt ganzzahlige lineare Quotientenoptimierung nach dem Schnittverfahren von Gomory. (German) [Mixed integer linear optimization quotient according to the cut procedure of Gomory]. Mathematische Operationsforschung und Statistik, 6(6):845–854, 1975. CODEN MTOSBI. ISSN 0047-6277.

Hartmann:1975:RGL

[Har75b] Karl Hartmann. Rein ganzzahlige lineare Quotientenoptimierung nach dem Schnittverfahren von Gomory. (German) [Pure integer linear optimization quotient according to the cut procedure of Gomory]. Mathematische Operationsforschung und Statistik, 6(1):33–53, 1975. CODEN MTOSBI. ISSN 0047-6277.

Hasminskii:1977:SAM

[Has77] R. Z. Hasminskii. Stochastic approximation methods in nonlinear regression models. Statistics (Berlin, DDR), 8(1):95– 106, 1977. CODEN MOSSD5. ISSN 0323-3944.

Haverland:1976:MRC

[Hav76] U. Haverland. A method for the recursive,computation of F statistics. Mathematische Operationsforschung und Statistik, 7(5):721–727, 1976. CODEN MTOSBI. ISSN 0047-6277.

Havranek:1984:NRM

[Hav84] T. Havranek. A note on the rank monotone dependence function. Statistics (Berlin, DDR), 15(3):369–371, 1984. CODEN MOSSD5. ISSN 0323-3944.

Havranek:1988:GAM

[Hav88] Tomas Havranek. On a general algorithm for model choice in multivariate analysis. Statistics: a Journal of Theoretical and Applied Statistics, 19(3):465–475, 1988. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Hwang:2009:AVC

[HB09] S. Y. Hwang and J. S. Baek. Asymptotic variance-covariance matrix of sample autocorrelations for threshold-asymmetric GARCH processes. Statistics: a Journal of Theoretical and Applied Statistics, 43(1):35–51, 2009. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Hwang:2014:NEM

[HBCL14] S. Y. Hwang, I. V. Basawa, M. S. Choi, and S. D. Lee. Nonergodic martingale estimating functions and related asymptotics. Statistics: a Journal of Theoretical and Applied Statistics, 48(3):487–507, 2014. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Hahnewald-Busch:1986:SBC

[HBN86] A. Hahnewald-Busch and V. Nollau. Some basic concepts of numerical treatment of Markov decision models. Statistics: a Journal of Theoretical and Applied Statistics, 17(1):147–164, 1986. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Hardouin:2011:MSS

[HC11] C. Hardouin and T. Crivelli. Mixed-state spatio-temporal auto-models. Statistics: a Journal of Theoretical and Applied Statistics, 45(1):87–100, 2011. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Houssineau:2020:RPD

[HC20] Jeremie Houssineau and Daniel E. Clark. On a representation of partially distinguishable populations. Statistics: a Journal of Theoretical and Applied Statistics, 54(1):23–45, 2020. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Hu:2016:GSL

[HCW16] Feng Hu, Zengjing Chen, and Panyu Wu. A general strong law of large numbers for non-additive probabilities and its applications. Statistics: a Journal of Theoretical and Applied Statistics, 50(4):733–749, 2016. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Huet:1999:BCI

[HDA99] S. Huet, J.-B. Denis, and K. Adamczyk. Bootstrap confidence intervals in nonlinear regression models when the number of observations is fixed and the variance tends to 0. Application to biadditive models. Statistics: a Journal of Theoretical and Applied Statistics, 32(3):203–227, 1999. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Hore:2020:EBT

[HDC20] Samrat Hore, Anup Dewanji, and Aditya Chatterjee. Ensuring balance through optimal allocation of experimental units with known categorical covariates into two treatments. Statistics: a Journal of Theoretical and Applied Statistics, 54(3): 559–576, 2020. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Harel:2012:MEP

[HE12] Michel Harel and Echarif Elharfaoui. The marked empirical process to test nonlinear time series against a large class of alternatives when the random vectors are nonstationary and absolutely regular. Statistics: a Journal of Theoretical and Applied Statistics, 46(2):231–247, 2012. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Heckendorff:1983:BR

[Hec83] H. Heckendorff. Book reviews. Statistics (Berlin, DDR), 14 (1):52, 1983. CODEN MOSSD5. ISSN 0323-3944.

Heckendorff:1986:BRH

[Hec86] H. Heckendorff. Book review: H.-J. Lenz, G. B. Wetherill, and P. Th. Wilrich (Eds.), Frontiers in Statistical Quality Control. Statistics: a Journal of Theoretical and Applied Statistics, 17 (4):504, 1986. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Herrador:2013:MNE

[HEHM13] M. Herrador, M. D. Esteban, T. Hobza, and D. Morales. A modified nested-error regression model for small area estimation. Statistics: a Journal of Theoretical and Applied Statistics, 47(2):258–273, 2013. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Heinrich:1984:TRS

[Hei84] L. Heinrich. On a test of randomness of spatial point patterns. Statistics (Berlin, DDR), 15(3):413–420, 1984. CO-DEN MOSSD5. ISSN 0323-3944.

Heinbich:1985:EPS

[Hei85] Lothar Heinbich. An elementary proof of Spitzer’s identity. Statistics: a Journal of Theoretical and Applied Statistics, 16(2):249–252, 1985. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Heinrich:1986:ANR

[Hei86] L. Heinrich. Asymptotic normality of a random point field characteristic. Statistics: a Journal of Theoretical and Applied Statistics, 17(3):453–460, 1986. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Heinrich:1988:AGS

[Hei88] L. Heinrich. Asymptotic Gaussianity of some estimators for reduced factorial moment measures and product densities of stationary Poisson cluster processes. Statistics: a Journal of Theoretical and Applied Statistics, 19(1):87–106, 1988. CO-DEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Heinrich:1989:BR

[Hei89] L. Heinrich. Book reviews. Statistics: a Journal of Theoretical and Applied Statistics, 20(4):606, 1989. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Heinrich:1991:GFT

[Hei91] L. Heinrich. Goodness-of-fit tests for the second moment function of a stationary multidimensional Poisson process. Statistics: a Journal of Theoretical and Applied Statistics, 22(2): 245–268, 1991. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Heinrich:1992:EMP

[Hei92] L. Heinrich. On existence and mixing properties of germgrain models. Statistics: a Journal of Theoretical and Applied Statistics, 23(3):271–286, 1992. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Heinrich:2018:AGF

[Hei18] L. Heinrich. Asymptotic goodness-of-fit tests for point processes based on scaled empirical K-functions. Statistics: a Journal of Theoretical and Applied Statistics, 52(4):829–851, 2018. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Hellmann:1981:UTS

[Hel81] Piotr Hellmann. Unbiasedness of two-sided nonparametric tests in the two-sample problem. Statistics (Berlin, DDR), 12(3):357–360, 1981. CODEN MOSSD5. ISSN 0323-3944.

Helms:1993:CPS

[Hel93] L. L. Helms. A counterexample to a purported solution of a martingale problem. Statistics: a Journal of Theoretical and Applied Statistics, 24(2):171–174, 1993. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Henschke:1979:CRP

[Hen79] K. Henschke. Confidence regions for parameters in linear models with additional information. Statistics (Berlin, DDR), 10(2):257–272, 1979. CODEN MOSSD5. ISSN 0323-3944.

Henschke:1980:SCP

[Hen80] K. Henschke. Simultaneous confidence procedures in multivariate calibration problems. Statistics (Berlin, DDR), 11(2): 193–206, 1980. CODEN MOSSD5. ISSN 0323-3944.

Henschke:1982:BR

[Hen82] K. Henschke. Book review. Statistics (Berlin, DDR), 13(4): 542, 1982. CODEN MOSSD5. ISSN 0323-3944.

Henschke:1986:BRa

[Hen86a] K. Henschke. Book review. Statistics: a Journal of Theoretical and Applied Statistics, 17(1):120, 1986. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Henschke:1986:BRb

[Hen86b] K. Henschke. Book review. Statistics: a Journal of Theoretical and Applied Statistics, 17(1):146, 1986. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Henschke:1987:BRa

[Hen87a] K. Henschke. Book review. Statistics: a Journal of Theoretical and Applied Statistics, 18(3):360, 1987. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Henschke:1987:BRb

[Hen87b] K. Henschke. Book review. Statistics: a Journal of Theoretical and Applied Statistics, 18(3):378, 1987. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Henschke:1987:BRc

[Hen87c] K. Henschke. Book review. Statistics: a Journal of Theoretical and Applied Statistics, 18(3):436, 1987. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Hentzschel:1989:PAZ

[Hen89] J. Hentzschel. Parametrische Auswertung zensierter Stichproben: Die Weibullverteilung. (German) [Parametric analysis of censored samples: the Weibull distribution]. Statistics: a Journal of Theoretical and Applied Statistics, 20(3): 383–395, 1989. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Hentzschel:1992:DEL

[Hen92] Jürgen Hentzschel. Density estimation with Laguerre series and censored samples. Statistics: a Journal of Theoretical and Applied Statistics, 23(1):49–61, 1992. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Henze:1993:QOT

[Hen93] Norbert Henze. A quick omnibus test for the proportional hazards model of random censorship. Statistics: a Journal of Theoretical and Applied Statistics, 24(3):253–263, 1993. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Herkenrath:1981:SCK

[Her81] Ulrich Herkenrath. On the speed of convergence of the Kiefer–Wolfowitz stochastic approximation procedure. Statistics (Berlin, DDR), 12(3):377–392, 1981. CODEN MOSSD5. ISSN 0323-3944.

Hoffmann:1988:BRa

[HF88] K. Hoffmann and W. Freudenberg. Book review. Statistics: a Journal of Theoretical and Applied Statistics, 19(2):232, 1988. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Harrar:2005:LCM

[HG05] Solomon W. Harrar and Arjun K. Gupta. On the LBI criterion for the multivariate one-way random effects model under non-normality. Statistics: a Journal of Theoretical and Applied Statistics, 39(5):405–414, 2005. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Harrar:2008:MVS

[HG08] Solomon W. Harrar and Arjun K. Gupta. On matrix variate skew-normal distributions. Statistics: a Journal of Theoretical and Applied Statistics, 42(2):179–194, 2008. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Henschke:1979:BRa

[HGH^{+}79] K. Henschke, B. Grabowski, K. Henschke, D. Stoyan, W. Warmuth, H.-J Girlich, P. Neumann, P. H. Müller, M. Weese, H. Gillert, H.-J Girlich, and H. W. Meier. Book reviews. Statistics (Berlin, DDR), 10(3):495–502, 1979. CODEN MOSSD5. ISSN 0323-3944.

Haynam:1982:TCNa

[HGLS82a] G. E. Haynam, Z. Govindarajulu, F. C. Leone, and P. Siefert. Tables of the cumulative non-central chi-square distribution — Part 1. Statistics (Berlin, DDR), 13(3):413–443, 1982. CODEN MOSSD5. ISSN 0323-3944.

Haynam:1982:TCNb

[HGLS82b] G. E. Haynam, Z. Govindarajulu, F. C. Leone, and P. Siefert. Tables of the cumulative non-central chi-square distribution — Part 2. Statistics (Berlin, DDR), 13(4):577–634, 1982. CODEN MOSSD5. ISSN 0323-3944.

Haynam:1983:TCNa

[HGLS83a] G. E. Haynam, Z. Govindarajulu, F. C. Leone, and P. Siefert. Tables of the cumulative non-central chi-square distribution — Part 3. Statistics (Berlin, DDR), 14(1):75–139, 1983. CO-DEN MOSSD5. ISSN 0323-3944.

Haynam:1983:TCNc

[HGLS83b] G. E. Haynam, Z. Govindarajulu, F. C. Leone, and P. Siefert. Tables of the cumulative non-central chi-square distribution — Part 5. Statistics (Berlin, DDR), 14(3):457–484, 1983. CODEN MOSSD5. ISSN 0323-3944.

Haynam:1983:TCNd

[HGLS83c] G. E. Haynam, Z. Govindarajulu, F. C. Leone, and P. Siefert. Tables of the cumulative non-central chi-square distribution — Part 6 (finish). Statistics (Berlin, DDR), 14(4):589–603, 1983. CODEN MOSSD5. ISSN 0323-3944.

Haynam:1983:TCNb

[HGLS83d] G. E. Haynam, Z. Govindarajulu, F. C. Leone, and P. Siefert. Tables of the cumulative non-central cm-square distribution — Part 4. Statistics (Berlin, DDR), 14(2):269–300, 1983. CODEN MOSSD5. ISSN 0323-3944.

Hohmann:2013:SLM

[HH13] Daniel Hohmann and Hajo Holzmann. Semiparametric location mixtures with distinct components. Statistics: a Journal of Theoretical and Applied Statistics, 47(2):348–362, 2013. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Harrar:2016:MHD

[HH16] Solomon W. Harrar and John Z. Hossler. Methods for highdimensional multivariate and multi-group repeated measures data under non-normality. Statistics: a Journal of Theoretical and Applied Statistics, 50(5):1056–??, 2016. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Hardle:1995:BBC

[HHJ95] W. Härdle, S. Huet, and E. Jolivet. Better bootstrap confidence intervals for regression curve estimation. Statistics: a Journal of Theoretical and Applied Statistics, 26(4):287–306, 1995. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Henze:2014:TSS

[HHM14] N. Henze, Z. Hlávka, and S. G. Meintanis. Testing for spherical symmetry via the empirical characteristic function. Statistics: a Journal of Theoretical and Applied Statistics, 48(6): 1282–1296, 2014. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Hudecova:2015:TTS

[HHM15] Sárka Hudecová, Marie Husková, and Simos G. Meintanis. Tests for time series of counts based on the probabilitygenerating function. Statistics: a Journal of Theoretical and Applied Statistics, 49(2):316–337, 2015. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Henschke:1987:BRd

[HHN^{+}87] K. Henschke, K. Henschke, P. Neumann, L. Heinrich, and Hans M. Dietz. Book review. Statistics: a Journal of Theoretical and Applied Statistics, 18(3):465–480, 1987. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Horvath:1997:ETC

[HHS97] Lajos Horváth, Marie Husková, and Monika Serbinowska. Estimators for the time of change in linear models. Statistics: a Journal of Theoretical and Applied Statistics, 29(2):109–130, 1997. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Htjet:1990:SSR

[HJM90] S. Htjet, E. Jolivet, and A. Messean. Some simulations results about confidence intervals and bootstrap methods in nonlinear regression. Statistics: a Journal of Theoretical and Applied Statistics, 21(3):369–432, 1990. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Hamerle:1978:BBL

[HK78] A. Hamerle and P. Kemény. Zur Bestimmung von Bayes-Lösungen bei messender Prüfung. (German) [On the purpose of Bayesian solutions for measurement testing]. Statistics (Berlin, DDR), 9(4):599–605, 1978. CODEN MOSSD5. ISSN 0323-3944.

Hardle:1987:NKR

[HK87] Wolfgang Härdle and Gabrielle Kelly. Nonparametric kernel regression estimation-optimal choice of bandwidth. Statistics: a Journal of Theoretical and Applied Statistics, 18(1):21–35, 1987. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Hartung:2000:ATG

[HK00] Joachim Hartung and Guido Knapp. Asymptotic tests for general linear hypotheses on variance components in models of commutative quadratic type. Statistics: a Journal of Theoretical and Applied Statistics, 35(1):23–44, 2000. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Hashorva:2009:SKA

[HK09] Enkelejd Hashorva and Samuel Kotz. On the strong Kotz approximation of Dirichlet random vectors. Statistics: a Journal of Theoretical and Applied Statistics, 43(4):393–408, 2009. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Huet:2015:GFT

[HK15] S. Huet and E. Kuhn. Goodness-of-fit test for Gaussian regression with block correlated errors. Statistics: a Journal of Theoretical and Applied Statistics, 49(2):239–266, 2015. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Hawlader:2023:SCL

[HKK23] Biplab Hawlader, Pradip Kundu, and Amarjit Kundu. Stochastic comparisons of lifetimes of fail-safe systems with dependent and heterogeneous components under random shocks. Statistics: a Journal of Theoretical and Applied Statistics, 57(3):694–709, 2023. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/02331888.2023.2203926.

Heinrich:1998:NAC

[HKMM98] L. Heinrich, R. Körner, N. Mehlhorn, and L. Muche. Numerical and analytical computation of some second-order characteristics of spatial Poisson–Voronoi tessellations. Statistics: a Journal of Theoretical and Applied Statistics, 31(3):235–259, 1998. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Hashempour:2022:WCR

[HKT22] M. Hashempour, M. R. Kazemi, and S. Tahmasebi. On weighted cumulative residual extropy: characterization, estimation and testing. Statistics: a Journal of Theoretical and Applied Statistics, 56(3):681–698, 2022. CO-DEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/02331888.2022.2072505.

Hollatz:1970:RGBa

[HL70] H. Hollatz and K. Lommatzsch. Rezensionen. (German) [Book reviews]. Mathematische Operationsforschung und Statistik, 1(1):77–78, 1970. CODEN MTOSBI. ISSN 0047-6277.

Hwang:2013:BSE

[HL13] Leng-Cheng Hwang and Cheng-Hung Lee. Bayes sequential estimation for a Poisson process under a LINEX loss function. Statistics: a Journal of Theoretical and Applied Statistics, 47(4):672–687, 2013. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Huang:2021:SES

[HLM21] Zhensheng Huang, Wen Lou, and Shuyu Meng. Statistical estimation for single-index varying-coefficient models with multiplicative distortion measurement errors. Statistics: a Journal of Theoretical and Applied Statistics, 55(4):851–874, 2021. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Hollander:2001:MLE

[HLsS01] Myles Hollander, Glen Laird, and Kai sheng Song. Maximum likelihood estimation in the proportional hazards model of random censorship. Statistics: a Journal of Theoretical and Applied Statistics, 35(3):245–258, 2001. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Huda:1988:OWD

[HM88] S. Huda and Rahul Mukerjee. Optimal weighing designs: approximate theory. Statistics: a Journal of Theoretical and Applied Statistics, 19(4):513–517, 1988. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Huda:1989:OMF

[HM89] S. Huda and Rahul Mukerjee. D-optimal measures for fourthorder rotatable designs. Statistics: a Journal of Theoretical and Applied Statistics, 20(3):353–356, 1989. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Henze:2002:TFE

[HM02] Norbert Henze and Simos G. Meintanis. Tests of fit for exponentiality based on the empirical Laplace transform. Statistics: a Journal of Theoretical and Applied Statistics, 36(2): 147–161, 2002. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Huskova:2007:OTE

[HM07] Marie Husková and Simos G. Meintanis. Omnibus tests for the error distribution in the linear regression model. Statistics: a Journal of Theoretical and Applied Statistics, 41(5):363– 376, 2007. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Hardle:2001:BTS

[HMP01] Wolfgang Härdle, Enno Mammen, and Isabel Proenca. A bootstrap test for single index models. Statistics: a Journal of Theoretical and Applied Statistics, 35(4):427–451, 2001. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Hashimoto:2013:EDA

[HOCC13] Elizabeth M. Hashimoto, Edwin M. M. Ortega, Vicente G. Cancho, and Gauss M. Cordeiro. On estimation and diagnostics analysis in log-generalized gamma regression model for interval-censored data. Statistics: a Journal of Theoretical and Applied Statistics, 47(2):379–398, 2013. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Hashimoto:2014:PBS

[HOCC14] Elizabeth M. Hashimoto, Edwin M. M. Ortega, Gauss M. Cordeiro, and Vicente G. Cancho. The Poisson Birnbaum– Saunders model with long-term survivors. Statistics: a Journal of Theoretical and Applied Statistics, 48(6):1394–1413, 2014. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Hoeschel:1983:BR

[Hoe83] H.-P. Hoeschel. Book reviews. Statistics (Berlin, DDR), 14 (2):262, 1983. CODEN MOSSD5. ISSN 0323-3944.

Hofstedt:1972:VEB

[Hof72] Klaus Hofstedt. Ein Verfahren für die exakte Behandlung von Problemen der Ablaufplanung mit Hilfsmittelbeschränkungen. (German) [A method for the exact treatment of problems of scheduling with resource constraints].

Mathematische Operationsforschung und Statistik, 3(4):239– 253, 1972. CODEN MTOSBI. ISSN 0047-6277.

Hoffmann:1977:ALE

[Hof77a] Kurt Hoffmann. Admissibility of linear estimators with respect to restricted parameter sets. Statistics (Berlin, DDR), 8(4):425–438, 1977. CODEN MOSSD5. ISSN 0323-3944.

Hoffmann:1977:RAL

[Hof77b] Kurt Hoffmann. Robust alternatives of the least squares estimator. Statistics (Berlin, DDR), 8(3):305–311, 1977. CODEN MOSSD5. ISSN 0323-3944.

Hoffmann:1979:CML

[Hof79a] Kurt Hoffmann. Characterization of minimax linear estimators in linear regression. Statistics (Berlin, DDR), 10(1):19– 26, 1979. CODEN MOSSD5. ISSN 0323-3944.

Hofmann:1979:BSE

[Hof79b] Bernd Hofmann. Bewertung von Schätzungen bei einem linearen Modell auf der Grundlage des Fiduzialkonzepts. (German) [Review of estimates in a linear model on the basis of fiducial concepts]. Statistics (Berlin, DDR), 10(1):37–46, 1979. CODEN MOSSD5. ISSN 0323-3944.

Hoffmann:1980:AIL

[Hof80] Kurt Hoffmann. Admissible improvements of the least squares estimator. Statistics (Berlin, DDR), 11(3):373–388, 1980. CODEN MOSSD5. ISSN 0323-3944.

Hoffmann:1981:SGL

[Hof81] Kurt Hoffmann. Sufficiency, generalized likelihood function and exponential family. Statistics (Berlin, DDR), 12(4):559– 566, 1981. CODEN MOSSD5. ISSN 0323-3944.

Hoffmann:1984:BRH

[Hof84] K. Hoffmann. Book review: H. Kres: Statistical Tables for Multivariate Analysis. Statistics (Berlin, DDR), 15(4):532, 1984. CODEN MOSSD5. ISSN 0323-3944.

Hoffmann:1985:AIE

[Hof85] Kurt Hoffmann. Admissibility and inadmissibility of estimators in the one-parameter exponential family. Statistics: a Journal of Theoretical and Applied Statistics, 16(3):327–350, 1985. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Hoffmann:1988:BRb

[Hof88] K. Hoffmann. Book review. Statistics: a Journal of Theoretical and Applied Statistics, 19(2):270, 1988. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Hoffmann:1995:NKO

[Hof95] Kurt Hoffmann. A note on the Kuks–Olman estimator. Statistics: a Journal of Theoretical and Applied Statistics, 26(3): 185–187, 1995. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Hoffmann:1997:EBS

[Hof97] Kurt Hoffmann. An empirical Bayes Stein-type estimator for regression parameters under linear constraints. Statistics: a Journal of Theoretical and Applied Statistics, 30(2):91–98, 1997. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Hollatz:1970:RGBb

[Hol70a] H. Hollatz. Rezension. (German) [Book review]. Mathematische Operationsforschung und Statistik, 1(4):327–328, 1970. CODEN MTOSBI. ISSN 0047-6277.

Hollatz:1970:KLO

[Hol70b] Horst Hollatz. Die Konstruktion lösbarer Optimierungsprobleme. (German) [The construction of solvable optimization problems]. Mathematische Operationsforschung und Statistik, 1(4):255–263, 1970. CODEN MTOSBI. ISSN 0047-6277.

Hollatz:1973:BKT

[Hol73a] Horst Hollatz. Bemerkungen zum Kuhn–Tucker-Theorem. (German) [Remarks on the Kuhn–Tucker Theorem]. Mathematische Operationsforschung und Statistik, 4(4):275–281, 1973. CODEN MTOSBI. ISSN 0047-6277.

Hollatz:1973:POL

[Hol73b] Horst Hollatz. Parametrische Optimierung in linearen Räumen. (German) [Parametric optimization in linear spaces]. Mathematische Operationsforschung und Statistik, 4(2):107–125, 1973. CODEN MTOSBI. ISSN 0047-6277.

Hollatz:1975:AVZ

[Hol75] Horst Hollatz. Ein allgemeines Verfahren der zulässigen Richtungen für diskrete Minimax-Aufgaben mit beschränkten Parametern. (German) [A general method of admissible directions for discrete minimax-tasks with limited parameters]. Mathematische Operationsforschung und Statistik, 6(4):535– 547, 1975. CODEN MTOSBI. ISSN 0047-6277.

Hollatz:1976:KEV

[Hol76] Hobst Hollatz. Konvergenzuntersuchung zu einem Verfahren der zulässigen Richtungen bei nichtlinearen Optimierungsaufgaben. (German) [Convergence study of a method of admissible directions in nonlinear optimization tasks]. Mathematische Operationsforschung und Statistik, 7(2):247–255, 1976. CODEN MTOSBI. ISSN 0047-6277.

Holmquist:1991:ETC

[Hol91] B. Holmquist. Estimating and testing the common mean direction of several von Mises—Fisher populations with known concentrations. Statistics: a Journal of Theoretical and Applied Statistics, 22(3):369–378, 1991. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Hopfner:1998:ESP

[Höp98] Reinhard Höpfner. Estimating a stability parameter: Asymptotics and simulations. Statistics: a Journal of Theoretical and Applied Statistics, 30(4):291–305, 1998. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Horvath:1984:RCB

[Hor84] Lajos Horváth. On random censorship from both sides. Statistics (Berlin, DDR), 15(4):581–594, 1984. CODEN MOSSD5. ISSN 0323-3944.

Horn:1988:SRP

[Hor88] Paul S. Horn. Some results on pooling. Statistics: a Journal of Theoretical and Applied Statistics, 19(1):9–14, 1988. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Horvath:1995:DCL

[Hor95] Lajos Horváth. Detecting changes in linear regressions. Statistics: a Journal of Theoretical and Applied Statistics, 26(3): 189–208, 1995. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Hoschel:1974:GAC

[Hös74] Hans-Peter Höschel. A general approach to correlation and linear dependence between random vectors with regular or singular covariance matrix. Mathematische Operationsforschung und Statistik, 5(6):487–507, 1974. CODEN MTOSBI. ISSN 0047-6277.

Hoschel:1976:AZA

[Hös76] Hans-Peter Höschel. Ein axiomatischer Zugang zu Abhängigkeitsmaßen für Zufallsvektoren und die Spurkorrelation. (German) [An axiomatic access to reasonable dependence for random vectors and the track correlation]. Mathematische Operationsforschung und Statistik, 7(5):789–802, 1976. CODEN MTOSBI. ISSN 0047-6277.

Hoschel:1977:GRV

[Hös77] Hans-Peter Höschel. A generalization of random vectors and a correlation theory for them — a trial (with discussions). Statistics (Berlin, DDR), 8(3):263–297, 1977. CODEN MOSSD5. ISSN 0323-3944.

Hoschel:1978:GLS

[Hös78] H. P. Höschel. Generalized least squares estimators of linear functional relations with known error-covariance. Statistics (Berlin, DDR), 9(1):9–26, 1978. CODEN MOSSD5. ISSN 0323-3944.

Hoschel:1986:USG

[Hös86] Hans-Peter Höschel. Uniqueness of surfaces and global identifiability in nonlinear regression models. Statistics: a Journal of Theoretical and Applied Statistics, 17(1):15–24, 1986. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Hostmann:1988:AOE

[Hos88a] Uwe Hostmann. Asymptotic optimality of estimators and tests in inadequate nonlinear regression. Statistics: a Journal of Theoretical and Applied Statistics, 19(2):185–206, 1988.

CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Hostmann:1988:NEC

[Hos88b] Uwe Hostmann. A note on an estimate for the convergence rate of the LSE in inadequate nonlinear regression models. Statistics: a Journal of Theoretical and Applied Statistics, 19 (1):3–8, 1988. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Hoy:1979:BSZ

[Hoy79] Ludwig Hoy. Eine Bemerkung zur Stetigkeit zufälliger Felder. (German) [A remark to the discontinuity of random fields]. Statistics (Berlin, DDR), 10(3):453–459, 1979. CODEN MOSSD5. ISSN 0323-3944.

Hofstedt:1973:SLG

[HP73] Klaus Hofstedt and Joachim Piehler. Ein Simulationsverfahren zur Lösung ganzzahliger linearer Optimalprobleme. (German) [A simulation method for solving integer linear optimal problems]. Mathematische Operationsforschung und Statistik, 4(2):99–105, 1973. CODEN MTOSBI. ISSN 0047-6277.

Heinrich:2008:WSC

[HP08] Lothar Heinrich and Zbynek Pawlas. Weak and strong convergence of empirical distribution functions from germ-grain processes. Statistics: a Journal of Theoretical and Applied Statistics, 42(1):49–65, 2008. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Hu:2022:LIH

[HPN22] Shuang Hu, Zuoxiang Peng, and Saralees Nadarajah. Location invariant heavy tail index estimation with block method. Statistics: a Journal of Theoretical and Applied Statistics, 56(3):479–497, 2022. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/02331888.2022.2071898.

Hansel:1988:BR

[HPP^{+}88] R. Hansel, J. Polzehl, J. Pilz, Jürgen Franz, K. Hoffmann, K.-H. Eger, B. Seigert, W. Schmidt, W. Schmidt, R. Trommer, K. Matthes, R. Pincus, D. Stoyan, M. Nagel, B. Droge, W. Näther, G. Herrendörfer, M. Nagel, Hans M. Dietz, H. U. Kuenle (??), D. Stoyan, B. Droge, P. Neumann, B. Droge, Walter Warmuth, T. Mikosch, V. Schmidt, V. Nollau, H.-J. Girlich, Johannes Terno, K. Schwarz, and O. Bunke. Book reviews. Statistics: a Journal of Theoretical and Applied Statistics, 19(4):605–631, 1988. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Herrendorfer:1980:EDO

[HR80] G. Herrendörfer and D. Rasch. Experimental design and optimal decision i general remarks. Statistics (Berlin, DDR), 11 (1):115–123, 1980. CODEN MOSSD5. ISSN 0323-3944.

Holtrode:1993:DPP

[HR93] R. Holtrode and L. Rüschendorf. Differentiablity of point process models and asymptotic efficiency of differentiable functionals. Statistics: a Journal of Theoretical and Applied Statistics, 24(1):17–42, 1993. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Hudak:1996:MSN

[HR96] Dietmar Hudak and Gerhard Richter. Moments of special normally distributed matrices. Statistics: a Journal of Theoretical and Applied Statistics, 27(3–4):363–378, 1996. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Hanisch:1979:FSO

[HS79] K. H. Hanisch and D. Stoyan. Formulas for the second-order analysis of marked point processes. Statistics (Berlin, DDR), 10(4):555–560, 1979. CODEN MOSSD5. ISSN 0323-3944.

Hoffmann:1982:CEF

[HS82] K. Hoffmann and C. Schmidt. Characterizations of the exponential family by generating functions and recursion relations for moments. Statistics (Berlin, DDR), 13(1):79–90, 1982. CODEN MOSSD5. ISSN 0323-3944.

Hakisch:1983:RSI

[HS83] K. H. Hakisch and D. Stoyan. Remarks on statistical inference and prediction for a hard-core clustering model. Statistics (Berlin, DDR), 14(4):559–567, 1983. CODEN MOSSD5. ISSN 0323-3944.

Hoang:1991:CFS

[HS91] Thanh Hoang and Alberto Seeger. On conjugate functions, subgradients, and directional derivatives of a class of optimality criteria in experimental design. Statistics: a Journal of Theoretical and Applied Statistics, 22(3):349–368, 1991. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Huang:2015:CBR

[HS15] Wen-Jang Huang and Nan-Cheng Su. Characterizations based on regression assumptions of order statistics. Statistics: a Journal of Theoretical and Applied Statistics, 49(3): 578–587, 2015. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Hwang:2017:ESM

[HS17a] Eunju Hwang and Dong Wan Shin. Estimation of structural mean breaks for long-memory data sets. Statistics: a Journal of Theoretical and Applied Statistics, 51(4):904–920, 2017. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Hwang:2017:SBR

[HS17b] Eunju Hwang and Dong Wan Shin. Stationary bootstrapping for realized covariations of high frequency financial data. Statistics: a Journal of Theoretical and Applied Statistics, 51(4):844–861, 2017. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Haidari:2023:MOL

[HSBB23] Abedin Haidari, Mostafa Sattari, Ghobad Saadat Kia (Barmalzan), and Narayanaswamy Balakrishnan. MRL ordering of largest order statistics from heterogeneous scale variables. Statistics: a Journal of Theoretical and Applied Statistics, 57(2):354–374, 2023. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
02331888.2023.2193407.

Huang:2013:SGS

[HSG13] Wen-Jang Huang, Nan-Cheng Su, and Arjun K. Gupta. A study of generalized skew-normal distribution. Statistics: a Journal of Theoretical and Applied Statistics, 47(5):942–953, 2013. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Henschke:1979:BRb

[HSM^{+}79] K. Henschke, W. Schmidt, P. H. Müller, H. Bandemer, H.- P. Höschel, V. Dupac, G. Maibaum, H.-P. Höschel, and A. Prékopa. Book reviews. Statistics (Berlin, DDR), 10(4): 613–622, 1979. CODEN MOSSD5. ISSN 0323-3944.

Haux:1985:ARP

[HSW85] R. Haux, M. Schumacher, and G. Weckesser. Asymptotic relative Pitman efficiencies of rank tests for complete block designs. Statistics: a Journal of Theoretical and Applied Statistics, 16(3):427–444, 1985. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Hofstedt:1971:ALG

[HT71] Klaus Hofstedt and Wolfgang Thämelt. Über einen Algorithmus zur Lösung Gemischt-Ganzzahliger Optimalprobleme. (German) [An algorithm for solving mixed-integer optimal problems]. Mathematische Operationsforschung und Statistik, 2(3):199–212, 1971. CODEN MTOSBI. ISSN 0047-6277.

Hashorva:2015:PMD

[HT15] Enkelejd Hashorva and Zhongquan Tan. Piterbarg’s maxdiscretization theorem for stationary vector Gaussian processes observed on different grids. Statistics: a Journal of Theoretical and Applied Statistics, 49(2):338–360, 2015. CO-DEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Hasler:2016:BNN

[HT16] Caren Hasler and Yves Tillé. Balanced k-nearest neighbour imputation. Statistics: a Journal of Theoretical and Applied Statistics, 50(6):1310–1331, 2016. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Hruzova:2016:CRO

[HTHF16] K. Hruzová, V. Todorov, K. Hron, and P. Filzmoser. Classical and robust orthogonal regression between parts of compositional data. Statistics: a Journal of Theoretical and Applied Statistics, 50(6):1261–1275, 2016. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Han:2023:OMC

[HTX23] Dong Han, Fugee Tsung, and Jinguo Xian. An optimization method for change-point monitoring in finite samples sequence. Statistics: a Journal of Theoretical and Applied Statistics, 57(5):1251–1266, 2023. CO-DEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/02331888.2023.2258249.

Hu:2010:PEE

[Hu10] Jianhua Hu. Properties of the explicit estimators in the extended growth curve model. Statistics: a Journal of Theoretical and Applied Statistics, 44(5):477–492, 2010. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Huber:1977:RME

[Hub77] P. J. Huber. Robust methods of estimation of regression coefficients. Statistics (Berlin, DDR), 8(1):41–53, 1977. CODEN MOSSD5. ISSN 0323-3944.

Huet:1986:MLL

[Hue86] Sylvie Huet. Maximum likelihood and least squares estimators for a non-linear model with heterogeneous variances. Statistics: a Journal of Theoretical and Applied Statistics, 17(4): 517–526, 1986. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Huskova:1988:APE

[Hus88] Marie Huskova. Adaptive parameter estimation for generalized Tukey’s λ-family. Statistics: a Journal of Theoretical and Applied Statistics, 19(1):15–26, 1988. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Huzak:2018:ECD

[Huz18] Miljenko Huzak. Estimating a class of diffusions from discrete observations via approximate maximum likelihood method. Statistics: a Journal of Theoretical and Applied Statistics, 52(2):239–272, 2018. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Helander:2020:PDF

[HVRI20] Sami Helander, Germain Van Bever, Sakke Rantala, and Pauliina Ilmonen. Pareto depth for functional data. Statistics: a Journal of Theoretical and Applied Statistics, 54(1): 182–204, 2020. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Hollatz:1970:BAMb

[HW70a] H. Hollatz and H. Weinert. Bericht über die Arbeitstagung “Mathematische Optimierung”. (German) [Report on the working conference “Mathematical Optimization”]. Mathematische Operationsforschung und Statistik, 1(4):329–331, 1970. CODEN MTOSBI. ISSN 0047-6277.

Hollatz:1970:BAMa

[HW70b] H. Hollatz and H. Weinert. Bericht über die Arbeitstagung “Mathematische Optimierung”. (German) [Report on the workshop ‘Mathematical Optimization’]. Mathematische Operationsforschung und Statistik, 1(3):247–248, 1970. CO-DEN MTOSBI. ISSN 0047-6277.

Hollatz:1971:ALD

[HW71] Horst Hollatz and Horst Weinert. Ein Algorithmus zur Lösung des doppelt-einparametrischen linearen Optimierungsproblems. (German) [An algorithm for solving the double oneparameter linear optimization problem]. Mathematische Operationsforschung und Statistik, 2(3):181–197, 1971. CODEN MTOSBI. ISSN 0047-6277.

Hollatz:1972:BAM

[HW72] H. Hollatz and H. Weinert. Bericht über die Arbeitstagung Mathematische Optimierung. (German) [Report on the workshop Mathematical Optimization]. Mathematische Operationsforschung und Statistik, 3(1):76–78, 1972. CODEN MTOSBI. ISSN 0047-6277.

Hollatz:1974:BAM

[HW74] H. Hollatz and H. Weinert. Berichte über die Arbeitstagung “Mathematische Optimierung” vom 5.–10. Mai 1973 in Neuendorf/Hiddenseem DDR. (German) [Reports of the working conference “Mathematical Programming”, 5th–10th May 1973 in Neuendorf/Hiddenseem DDR]. Mathematische Operationsforschung und Statistik, 5(3):249–251, 1974. CO-DEN MTOSBI. ISSN 0047-6277.

Heiler:1988:ANE

[HW88] Siegfried Heiler and Reinhart Willers. Asymptotic normality of R-estimates in the linear model. Statistics: a Journal of Theoretical and Applied Statistics, 19(2):173–184, 1988. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Hossain:2007:AMP

[HW07] Ahmed Hossain and Andrew R. Willan. Approximate MLEs of the parameters of location-scale models under type II censoring. Statistics: a Journal of Theoretical and Applied Statistics, 41(5):385–394, 2007. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic). See note [Tik09].

Hashorva:2014:TAW

[HW14] Enkelejd Hashorva and Zhichao Weng. Tail asymptotic of Weibull-type risks. Statistics: a Journal of Theoretical and Applied Statistics, 48(5):1155–1165, 2014. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Hempel:1972:RGB

[HWK^{+}72] L. Hempel, H. Weinert, K. Krickeberg, E. Rödel, H. Langer, H. Bandemer, O. Bunke, W. Dück, and H. Toutenburg. Rezensionen. (German) [Book reviews]. Mathematische Operationsforschung und Statistik, 3(1):67–73, 1972. CODEN MTOSBI. ISSN 0047-6277.

Huang:2017:RCM

[HWL17] Jianwen Huang, Jianjun Wang, and Guowang Luo. On the rate of convergence of maxima for the generalized Maxwell distribution. Statistics: a Journal of Theoretical and Applied Statistics, 51(5):1105–1117, 2017. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

He:2020:RSO

[HX20] Xu He and Hongmei Xie. Relative stochastic orders of weighted frailty models. Statistics: a Journal of Theoretical and Applied Statistics, 54(5):989–1004, 2020. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Huang:2010:BDE

[HY10] Mei Ling Huang and Wai Kong Yuen. A bivariate density estimation method based on level crossings. Statistics: a Journal of Theoretical and Applied Statistics, 44(1):31–55, 2010. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Hosseini:2021:BBS

[HZ21] Reyhaneh Hosseini and Mahmoud Zarepour. Bayesian bootstrapping for symmetric distributions. Statistics: a Journal of Theoretical and Applied Statistics, 55(3):711–732, 2021. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Ibazizen:2003:RPC

[ID03] Mohamed Ibazizen and Jacques Dauxois. A robust principal component analysis. Statistics: a Journal of Theoretical and Applied Statistics, 37(1):73–83, 2003. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Iliopoulos:2012:APN

[IDB12] G. Iliopoulos, A. Dembińska, and N. Balakrishnan. Asymptotic properties of numbers of observations near sample quantiles. Statistics: a Journal of Theoretical and Applied Statistics, 46(1):85–97, 2012. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Ibazizen:2003:BEA

[IF03] Mohamed Ibazizen and Hocine Fellag. Bayesian estimation of an AR(1) process with exponential white noise. Statistics: a Journal of Theoretical and Applied Statistics, 37(5):365–372, 2003. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Igarashi:2020:NDD

[Iga20] Gaku Igarashi. Nonparametric direct density ratio estimation using beta kernel. Statistics: a Journal of Theoretical and Applied Statistics, 54(2):257–280, 2020. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Inglot:1990:NTS

[IJL90] Tadeusz Inglot, Tekesa Jurlewicz, and Teresa Ledwina. On Neytman-type smooth tests of fit. Statistics: a Journal of Theoretical and Applied Statistics, 21(4):549–568, 1990. CO-DEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Ishwaran:2008:OPI

[IJZ08] Hemant Ishwaran, Mohammad T. Jahandideh, and Mahmoud Zarepour. Option pricing for infinite variance data. Statistics: a Journal of Theoretical and Applied Statistics, 42(3):245– 260, 2008. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Ittrich:2000:PLQ

[IKR00] C. Ittrich, D. Krause, and W.-D. Richter. Probabilities and large quantiles of noncentral generalized chi-square distributions. Statistics: a Journal of Theoretical and Applied Statistics, 34(1):53–101, 2000. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Ivanov:2015:EHC

[ILRMZ15] A. V. Ivanov, N. N. Leonenko, M. D. Ruiz-Medina, and B. M. Zhurakovsky. Estimation of harmonic component in regression with cyclically dependent errors. Statistics: a Journal of Theoretical and Applied Statistics, 49(1):156–186, 2015. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Iliopoulos:2014:EPI

[IM14] G. Iliopoulos and S. M. T. K. MirMostafaee. Exact prediction intervals for order statistics from the Laplace distribution based on the maximum-likelihood estimators. Statistics: a Journal of Theoretical and Applied Statistics, 48(3):575–592, 2014. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Imhof:2000:EDM

[Imh00] Lorens Imhof. Exact designs minimising the integrated variance in quadratic regression. Statistics: a Journal of Theoretical and Applied Statistics, 34(2):103–115, 2000. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Ishra:1991:BEB

[IR91] M. N. Ishra and B. L. S. Rakasa Rao. Bounds for the equivalence of Bayes and maximum likelihood estimators for a class of diffusion processes. Statistics: a Journal of Theoretical and Applied Statistics, 22(4):613–625, 1991. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Ittrich:2005:ETC

[IR05] C. Ittrich and W.-D. Richter. Exact tests and confidence regions in nonlinear regression. Statistics: a Journal of Theoretical and Applied Statistics, 39(1):13–42, 2005. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Izadkhah:2014:AMR

[IRB14] S. Izadkhah, A. H. Rezaei Roknabadi, and G. R. Mohtashami Borzadaran. Aspects of the mean residual life order for weighted distributions. Statistics: a Journal of Theoretical and Applied Statistics, 48(4):851–861, 2014. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Ibarrola:1998:EMC

[IRV98] P. Ibarrola, R. Rózański, and R. Vélez. On estimation of the mean and covariance parameter for Gaussian random fields. Statistics: a Journal of Theoretical and Applied Statistics, 31 (1):1–20, 1998. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Irle:1984:OSP

[IS84] Albrecht Irle and Norbert Schmitz. On the optimality of the SPRT for processes with continuous time parameter. Statistics (Berlin, DDR), 15(1):91–104, 1984. CODEN MOSSD5. ISSN 0323-3944.

Ibarrola:1989:SDP

[IV89] Pilar Ibarrola and Ricardo Vélez. Sufficiency in decision processes with continuous observations. Statistics: a Journal of Theoretical and Applied Statistics, 20(1):125–134, 1989. CO-DEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Ibarbola:1990:SEM

[IV90a] P. Ibarbola and R. Vélez. On sequential estimation of the mean of a multidimensional Gaussian process. Statistics: a Journal of Theoretical and Applied Statistics, 21(3):471–482, 1990. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Ibarbola:1990:IDP

[IV90b] Pilas Ibarbola and Ricardo Vélez. Invariance in decision processes with continuous observations. Statistics: a Journal of Theoretical and Applied Statistics, 21(1):117–130, 1990. CO-DEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Ibarrola:2002:TCE

[IV02] Pilar Ibarrola and Ricardo Vélez. Testing and confidence estimation of the mean of a multidimensional Gaussian process. Statistics: a Journal of Theoretical and Applied Statistics, 36(4):317–327, 2002. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Ivanon:1983:AED

[IZ83] A. V. Ivanon and S. Zwanzig. An asymptotic expansion of the distribution of least squares estimators in the nonlinear regression model. Statistics (Berlin, DDR), 14(1):7–27, 1983. CODEN MOSSD5. ISSN 0323-3944.

Jilek:1989:BST

[JA89] Milos Jílek and Hanns Ackermann. A bibliography of statistical tolerance regions, II. Statistics: a Journal of Theoretical and Applied Statistics, 20(1):165–172, 1989. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Jacroux:1991:SNM

[Jac91] Mike Jacroux. Some new methods for establishing the optimality of block designs having unequally replicated treatments. Statistics: a Journal of Theoretical and Applied Statistics, 22(1):33–48, 1991. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Jacroux:2013:OBM

[Jac13] Mike Jacroux. Optimal blocked main effects plans. Statistics: a Journal of Theoretical and Applied Statistics, 47(5):1022– 1029, 2013. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Jagers:1980:ILM

[Jag80] P. Jagers. Invariance in the linear model — an argument for χ^{2} and F in nonnormal situations. Statistics (Berlin, DDR), 11(4):455–464, 1980. CODEN MOSSD5. ISSN 0323-3944.

Jahn:1987:BR

[Jah87] W. Jahn. Book review. Statistics: a Journal of Theoretical and Applied Statistics, 18(3):388, 1987. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Jain:1974:BTD

[Jai74] G. C. Jain. A Borel–Tanner distribution and its approximation to the negative binomial distribution. Mathematische Operationsforschung und Statistik, 5(3):207–212, 1974. CO-DEN MTOSBI. ISSN 0047-6277.

Jasiukiewicz:1981:EUE

[Jas81] T. Jasiukiewicz. On existence of UMVUQE for each estimable parametric function of variance covariance components in block designs. Statistics (Berlin, DDR), 12(4):479–485, 1981. CODEN MOSSD5. ISSN 0323-3944.

Jasinski:2018:CGD

[Jas18] Krzysztof Jasiński. Characterizations of geometric distributions based on k-th record values. Statistics: a Journal of Theoretical and Applied Statistics, 52(5):1116–1127, 2018. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Jasinski:2022:SCF

[Jas22] Krzysztof Jasiński. On the status of component failures in a working coherent system when the lifetimes are DNID random variables. Statistics: a Journal of Theoretical and Applied Statistics, 57(1):175–194, 2022. CO-DEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/02331888.2022.2160717.

Jaworski:2013:IDS

[Jaw13] Piotr Jaworski. Invariant dependence structure under univariate truncation: the high-dimensional case. Statistics: a Journal of Theoretical and Applied Statistics, 47(5):1064–1074, 2013. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Joshi:1983:BME

[JB83] P. C. Joshi and N. Balakrishnan. Bounds for the moments of extreme order statistics for large samples. Statistics (Berlin, DDR), 14(3):387–396, 1983. CODEN MOSSD5. ISSN 0323-3944.

Jha:2022:TEC

[JBC22] J. Jha, Atanu Biswas, and Tsung-Chi Cheng. Trimmed estimator for circular-circular regression: breakdown properties and an exact algorithm for computation. Statistics: a Journal of Theoretical and Applied Statistics, 56(2):375–395, 2022. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Jozani:2014:SPC

[JBD14] Mohammad Jafari Jozani, N. Balakrishnan, and Katherine F. Davies. Some Pitman closeness properties pertinent to symmetric populations. Statistics: a Journal of Theoretical and Applied Statistics, 48(6):1380–1393, 2014. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Jacome:2007:ASA

[JC07] M. A. Jácome and R. Cao. Almost sure asymptotic representation for the presmoothed distribution and density estimators for censored data. Statistics: a Journal of Theoretical and Applied Statistics, 41(6):517–534, 2007. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Jana:2022:EOR

[JC22] Nabakumar Jana and Ankur Chakraborty. Estimation of order restricted standard deviations of normal populations with a common mean. Statistics: a Journal of Theoretical and Applied Statistics, 56(4):867–890, 2022. CO-DEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/02331888.2022.2079126.

Jebeli:2020:EMP

[JD20] Maleki Jebeli and E. Deiri. Estimation methods for the probability density function and the cumulative distribution function of the Pareto–Rayleigh distribution. Statistics: a Journal of Theoretical and Applied Statistics, 54(1):135–151, 2020. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Jumad-Eldust:1990:STE

[JE90] A. A. Aly Jumad-Eldust. On some tests for exponentiality against the alternatives. Statistics: a Journal of Theoretical and Applied Statistics, 21(2):217–226, 1990. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Jennergren:1973:NDW

[Jen73] Peter Jennergren. A note on a Dantzig–Wolfe decompositionlike method for solving a particular resource-allocation problem under uncertainty. Mathematische Operationsforschung und Statistik, 4(2):127–132, 1973. CODEN MTOSBI. ISSN 0047-6277.

Jensen:1990:ECC

[Jen90] Uwe Jensen. An example concerning the convergence of conditional expectations. Statistics: a Journal of Theoretical and Applied Statistics, 21(4):609–611, 1990. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Javier:1985:GMV

[JG85] W. R. Javier and A. K. Gupta. On generalized matric variate beta distributions. Statistics: a Journal of Theoretical and Applied Statistics, 16(4):549–558, 1985. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Jimenez-Gamero:2014:TCD

[JGAFBCMG14] M. D. Jiménez-Gamero, V. Alba-Fernández, I. Barranco-Chamorro, and J. Muñoz-García. Two classes of divergence statistics for testing uniform association. Statistics: a Journal of Theoretical and Applied Statistics, 48(2):367–387, 2014. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Jimenez-Gamero:2020:ETB

[JGMO20] M. D. Jiménez-Gamero, Bojana Milosević, and Marko Obradović. Exponentiality tests based on Basu characterization. Statistics: a Journal of Theoretical and Applied Statistics, 54(4):714–736, 2020. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Jiang:2009:TSN

[JH09] Xinxin Jiang and Marjorie G. Hahn. Testing serial nonindependence by self-centring and self-normalizing. Statistics: a Journal of Theoretical and Applied Statistics, 43(4): 315–328, 2009. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Jiang:2018:CQR

[JHYQ18] Rong Jiang, Xueping Hu, Keming Yu, and Weimin Qian. Composite quantile regression for massive datasets. Statistics: a Journal of Theoretical and Applied Statistics, 52(5): 980–1004, 2018. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Jialek:1981:BST

[Jiá81] Milos Jiálek. A bibliography of statistical tolerance regions. Statistics (Berlin, DDR), 12(3):441–456, 1981. CO-DEN MOSSD5. ISSN 0323-3944.

Jiang:2007:TMS

[Jia07] Xinxin Jiang. Testing that marginal sequences of data are not independent via self-normalization. Statistics: a Journal of Theoretical and Applied Statistics, 41(2):119–128, 2007. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Jacome:2010:PED

[JIP10] M. A. Jácome and M. C. Iglesias-Pérez. Presmoothed estimation of the density function with truncated and censored data. Statistics: a Journal of Theoretical and Applied Statistics, 44(3):217–234, 2010. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Jung:2006:TOC

[JJS06] Byoung Cheol Jung, Myoungshic Jhun, and Seuck Heun Song. Testing for overdispersion in a censored Poisson regression model. Statistics: a Journal of Theoretical and Applied Statistics, 40(6):533–543, 2006. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Jansen:1976:ISZ

[JK76] U. Jansen and D. König. Invariante stationäre Zustandswahrscheinlichkeiten für eine Klasse stochastischer Modelle mit funktionellen Abhängigkeiten. (German) [Invariant stationary state probabilities for a class of stochastic models with functional dependencies]. Mathematische Operationsforschung und Statistik, 7(4):497–522, 1976. CODEN MTOSBI. ISSN 0047-6277.

Jain:1979:EF

[JK79] G. C. Jain and M. S. H. Khan. On an exponential family. Statistics (Berlin, DDR), 10(1):153–168, 1979. CODEN MOSSD5. ISSN 0323-3944.

Janicki:1992:CIR

[JK92] Aleksander Janicki and Piotr Kokoszka. Computer investigation of the rate of convergence of Lepage type series to α-stable random variables. Statistics: a Journal of Theoretical and Applied Statistics, 23(4):365–373, 1992. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Johnson:2007:CD

[JK07] Norman L. Johnson and Samuel Kotz. Cloning of distributions. Statistics: a Journal of Theoretical and Applied Statistics, 41(2):145–152, 2007. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Jach:2008:WDT

[JK08] Agnieszka Jach and Piotr Kokoszka. Wavelet-domain test for long-range dependence in the presence of a trend. Statistics: a Journal of Theoretical and Applied Statistics, 42(2):101–113, 2008. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Jamalizadeh:2013:WMO

[JK13] Ahad Jamalizadeh and Debasis Kundu. Weighted Marshall– Olkin bivariate exponential distribution. Statistics: a Journal of Theoretical and Applied Statistics, 47(5):917–928, 2013. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Jin:2023:OSA

[JLM23] Jun Jin, Shuangzhe Liu, and Tiefeng Ma. Optimal subsampling algorithms for composite quantile regression in massive data. Statistics: a Journal of Theoretical and Applied Statistics, 57(4):811–843, 2023. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
02331888.2023.2239507.

Joarder:2012:SCB

[JLO12] Anwar H. Joarder, A. Laradji, and M. Hafidz Omar. On some characteristics of bivariate chi-square distribution. Statistics: a Journal of Theoretical and Applied Statistics, 46(5):577– 586, 2012. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Jha:2021:IRS

[JMK21] Brijesh Kumar Jha, Ajaya Kumar Mahapatra, and Suchandan Kayal. Inadmissibility results for the selected hazard rates. Statistics: a Journal of Theoretical and Applied Statistics, 55(3):580–594, 2021. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Jakob:1975:ONV

[JN75] J. Jakob and P. Neumann. Optimale Nullstellensuche bei vorliegen einer Apriori-Verteilung. (German) [Optimal zeropoint search on an a priori distribution]. Mathematische Operationsforschung und Statistik, 6(3):397–412, 1975. CODEN MTOSBI. ISSN 0047-6277.

Jacob:1995:GAN

[JO95] Pierre Jacob and Paulo Eduardo Oliveira. A general approach to nonparametric histogram estimation. Statistics: a Journal of Theoretical and Applied Statistics, 27(1–2):73–92, 1995. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Jacob:1997:KEG

[JO97] Pierre Jacob and Paulo Eduardo Oliveira. Kernel estimators of general Radon–Nikodym derivatives. Statistics: a Journal of Theoretical and Applied Statistics, 30(1):25–46, 1997. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

John:1980:NDC

[Joh80] J. A. John. New developments in classical design. Statistics (Berlin, DDR), 11(3):389–402, 1980. CODEN MOSSD5. ISSN 0323-3944.

Jones:1976:BSE

[Jon76a] P. W. Jones. Bayes sequential estimation of multinomial parameters. Mathematische Operationsforschung und Statistik, 7(1):123–127, 1976. CODEN MTOSBI. ISSN 0047-6277.

Jones:1976:SRT

[Jon76b] P. W. Jones. Some results for the two armed bandit problem. Mathematische Operationsforschung und Statistik, 7(3):471– 475, 1976. CODEN MTOSBI. ISSN 0047-6277.

Jones:1995:LVB

[Jon95] M. C. Jones. Local and variable bandwidths and local linear regression. Statistics: a Journal of Theoretical and Applied Statistics, 27(1–2):65–71, 1995. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Jorg:1999:CSF

[Jör99] Pawlitschko Jörg. A comparison of survival function estimators in the Koziol–Green model. Statistics: a Journal of Theoretical and Applied Statistics, 32(3):277–291, 1999. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Jozani:2014:NPM

[Joz14] Mohammad Jafari Jozani. A note on Pitman’s measure of closeness with balanced loss function. Statistics: a Journal of Theoretical and Applied Statistics, 48(3):552–557, 2014. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Jahabdah:1986:IGM

[JR86] K. G. Jahabdah and B. Raja Rao. Identifiability of a generalized Markov–Pólya damage model. Statistics: a Journal of Theoretical and Applied Statistics, 17(2):303–310, 1986. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Johnson:2006:IFA

[JR06] Kjell Johnson and William Rayens. Influence function analysis for partial least squares with uncorrelated components. Statistics: a Journal of Theoretical and Applied Statistics, 40 (1):65–93, 2006. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Jokiel-Rokita:2012:BSE

[JR12] Alicja Jokiel-Rokita. Bayes sequential estimation for a timetransformed exponential model. Statistics: a Journal of Theoretical and Applied Statistics, 46(1):123–129, 2012. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Jasinski:2013:MVO

[JR13] Krzysztof Jasiński and Tomasz Rychlik. Maximum variance of order statistics from symmetric populations revisited. Statistics: a Journal of Theoretical and Applied Statistics, 47(2): 422–438, 2013. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Janamiri:2022:REL

[JRCB22] Fariba Janamiri, Abdolrahman Rasekh, Alireza Chaji, and Babak Babadi. Ridge estimation in linear mixed measurement error models using generalized maximum entropy. Statistics: a Journal of Theoretical and Applied Statistics, 56(5):1095– 1112, 2022. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-

4910 (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/02331888.2022.2126474.

Jokiel-Rokita:1999:EDO

[JRM99] Alicja Jokiel-Rokita and Ryszard Magiera. Estimation with delayed observations for the multinomial distribution. Statistics: a Journal of Theoretical and Applied Statistics, 32(4): 353–367, 1999. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Jokiel-Rokita:2004:MED

[JRM04] Alicja Jokiel-Rokita and Ryszard Magiera. Γ-minimax estimation with delayed observations from the multinomial distribution. Statistics: a Journal of Theoretical and Applied Statistics, 38(3):195–206, 2004. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Jaschke:1996:BR

[JRR^{+}96] S. Jaschke, Egmar Rödel, Egmar Rödel, Egmar Rödel, Helmut Lütkepohl, Wolfgang Wagner, Wolfgang Wagner, Wolfgang Wagner, and Ingeborg Küchler. Book reviews. Statistics: a Journal of Theoretical and Applied Statistics, 28(1):83, 1996. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Jokiel-Rokita:2021:MDE

[JRT21] Alicja Jokiel-Rokita and Rafał Topolnicki. Minimum distance estimation of the Lehmann receiver operating characteristic curve. Statistics: a Journal of Theoretical and Applied Statistics, 55(3):618–634, 2021. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Jia:2003:WCL

[JRZ03] Xinwei Jia, M. Bhaskara Rao, and Haimeng Zhang. On weak consistency in linear models with equi-correlated random errors. Statistics: a Journal of Theoretical and Applied Statistics, 37(6):463–473, 2003. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Jockel:1981:CLT

[JS81] K. H. Jöckel and W. Sendler. A Central Limit Theorem for generalized discounting. Statistics (Berlin, DDR), 12(4):605– 608, 1981. CODEN MOSSD5. ISSN 0323-3944.

Jureckova:1982:SEC

[JS82] Jana Jureckova and Paranab Kumar Sen. Simultaneous Mestimator of the common location and the scale-ratio in the two-sample problem. Statistics (Berlin, DDR), 13(2):163– 169, 1982. CODEN MOSSD5. ISSN 0323-3944.

Jahn:1989:BR

[JS89] W. Jahn and W. Schmidt. Book review. Statistics: a Journal of Theoretical and Applied Statistics, 20(1):173–175, 1989. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Joarder:1997:ETS

[JS97] Anwar H. Joarder and Sarjinder Singh. Estimation of the trace of the scale matrix of a multivariate T -model using regression type estimator statistics. Statistics: a Journal of Theoretical and Applied Statistics, 29(2):161–168, 1997. CO-DEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Jang:2014:TRT

[JS14] Myoung Jin Jang and Dong Wan Shin. Tests for random time effects and spatial error correlation in panel regression models. Statistics: a Journal of Theoretical and Applied Statistics, 48(1):101–120, 2014. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Jana:2022:ILQ

[JS22a] Kaushik Jana and Debasis Sengupta. Improving linear quantile regression for replicated data. Statistics: a Journal of Theoretical and Applied Statistics, 56(6):1193–1206, 2022. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/02331888.2022.2152451.

Jose:2022:RER

[JS22b] Jitto Jose and E. I. Abdul Sathar. Rényi entropy on k-records and its applications in characterizing distributions. Statistics: a Journal of Theoretical and Applied Statistics, 56(3):662– 680, 2022. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-

4910 (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/02331888.2022.2070167.

Jureckova:1978:BR

[JSAG78] J. Jurecková, R. Strüby, U. Arndt, and H. J. Girlich. Book reviews. Statistics (Berlin, DDR), 9(2):327–331, 1978. CODEN MOSSD5. ISSN 0323-3944.

Janssen:2007:MKD

[JSV07] Paul Janssen, Jan Swanepoel, and Noël Veraverbeke. Modifying the kernel distribution function estimator towards reduced bias. Statistics: a Journal of Theoretical and Applied Statistics, 41(2):93–103, 2007. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Jiang:2022:ABM

[JSW22] Hui Jiang, Jin Shao, and Shaochen Wang. Asymptotic behaviours for maximum likelihood estimator of drift parameter in α-Wiener bridge process. Statistics: a Journal of Theoretical and Applied Statistics, 56(5):1048–1071, 2022. CO-DEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/02331888.2022.2120878.

Jena:2022:EPP

[JT22] Pravash Jena and Manas Ranjan Tripathy. Estimating positive powers of the scale parameters under order restriction for two normal populations with a common mean. Statistics: a Journal of Theoretical and Applied Statistics, 56(3): 598–630, 2022. CODEN MOSSD5. ISSN 0233-1888 (print),

1029-4910 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/02331888.2022.2079125.

Juhasz:1989:TBC

[Juh89] Ferenc Juhasz. On the theoretical backgrounds of cluster analysis based on the eigenvalue problem of the association matrix. Statistics: a Journal of Theoretical and Applied Statistics, 20(4):573–581, 1989. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Jureckova:1975:NET

[Jur75] Jana Jurecková. Nonparametric estimation and testing linear hypotheses in the linear regression model. Mathematische Operationsforschung und Statistik, 6(2):269–283, 1975. CODEN MTOSBI. ISSN 0047-6277.

Jureckoya:1980:ARE

[Jur80] Jana Jureckoyá. Asymptotic representation of M-estimators of location. Statistics (Berlin, DDR), 11(1):61–73, 1980. CO-DEN MOSSD5. ISSN 0323-3944.

Jurgen:1984:DKD

[Jür84] Kindler Jürgen. Definitheit einer klasse diskreter statistischer Entscheidungsprobleme. (German) [Definiteness of a class of discrete statistical decision problems]. Statistics (Berlin, DDR), 15(1):105–120, 1984. CODEN MOSSD5. ISSN 0323-3944.

Janic-Wroblewska:2004:DDSa

[JW04a] Alicja Janic-Wróblewska. Data-driven smooth test for a location-scale family. Statistics: a Journal of Theoretical and Applied Statistics, 38(4):337–355, 2004. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Janic-Wroblewska:2004:DDSb

[JW04b] Alicja Janic-Wróblewska. Data-driven smooth tests for the extreme value distribution. Statistics: a Journal of Theoretical and Applied Statistics, 38(5):413–426, 2004. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Jeske:2020:SSC

[JY20] Daniel R. Jeske and Weixin Yao. Sample size calculations for mixture alternatives in a control group vs. treatment group design. Statistics: a Journal of Theoretical and Applied Statistics, 54(1):97–113, 2020. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Kalpakam:1988:CRP

[KA88] S. Kalpakam and G. Arivarignan. A continuous review perishable inventory model. Statistics: a Journal of Theoretical and Applied Statistics, 19(3):389–398, 1988. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Kumar:2013:GTP

[KA13] C. Satheesh Kumar and M. R. Anusree. A generalized twopiece skew normal distribution and some of its properties. Statistics: a Journal of Theoretical and Applied Statistics, 47 (6):1370–1380, 2013. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Kalahnea:1976:EUS

[Kal76] Ursula Kalähnea. Existence, uniqueness and some in variance properties of stationary distributions for general single server queues. Mathematische Operationsforschung und Statistik, 7 (4):557–575, 1976. CODEN MTOSBI. ISSN 0047-6277.

Kaluszka:1997:ICM

[Kal97] Marek Kaluszka. Influence of censorship on the minimax risk of decision procedures. Statistics: a Journal of Theoretical and Applied Statistics, 29(2):169–178, 1997. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Kamburowski:1989:BMM

[Kam89] Jerzy Kamburowski. Bounds on the mean of the maximum of a number of normal random variables. Statistics: a Journal of Theoretical and Applied Statistics, 20(3):475–481, 1989. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Kong:2017:CLT

[KAM17] Xiangshun Kong, Mingyao Ai, and Rahul Mukerjee. Central limit theorems for four new types of U-designs. Statistics: a Journal of Theoretical and Applied Statistics, 51(3):655–667, 2017. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Kamila:2022:GHQ

[Kam22] Kare Kamila. General Hannan and Quinn criterion for common time series. Statistics: a Journal of Theoretical and Applied Statistics, 56(1):222–241, 2022. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Khatib:2014:RPL

[KAR14] B. Khatib, Jafar Ahmadi, and M. Razmkhah. Reconstruction of the past lower record values in a proportional reversed hazard rate model. Statistics: a Journal of Theoretical and Applied Statistics, 48(2):421–435, 2014. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Katzenbeisser:1989:EPT

[Kat89] Walter Katzenbeisser. The exact power of two-sample location tests based on exceedance statistics against shift alternatives. Statistics: a Journal of Theoretical and Applied Statistics, 20(1):47–54, 1989. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Kawai:2015:LFS

[Kaw15] Reiichiro Kawai. On the likelihood function of small time variance Gamma lévy processes. Statistics: a Journal of Theoretical and Applied Statistics, 49(1):63–83, 2015. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Koeblok:1985:APS

[KB85] Uwe Koeblok and Tit Braunschweig. Asymptotic properties of a sequential estimator of expectation in the presence of trend. Statistics: a Journal of Theoretical and Applied Statistics, 16(2):203–211, 1985. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Kossler:2000:APR

[KB00] Wolfgang Kössler and Herbert Büning. The asymptotic power and relative efficiency of some c-sample rank tests of homogeneity against umbrella alternatives. Statistics: a Journal of Theoretical and Applied Statistics, 34(1):1–26, 2000. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Kamps:2001:DGO

[KC01] Udo Kamps and Erhard Cramer. On distributions of generalized order statistics. Statistics: a Journal of Theoretical and Applied Statistics, 35(3):269–280, 2001. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Kundu:2017:OPS

[KC17] Amarjit Kundu and Shovan Chowdhury. Ordering properties of sample minimum from Kumaraswamy-G random variables. Statistics: a Journal of Theoretical and Applied Statistics, 52(1):133–146, 2017. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Kim:2016:HBR

[KCL16] Hea-Jung Kim, Taeryon Choi, and Suyeon Lee. A hierarchical Bayesian regression model for the uncertain functional constraint using screened scale mixtures of Gaussian distributions. Statistics: a Journal of Theoretical and Applied Statistics, 50(2):350–376, 2016. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic). URL http://

www.tandfonline.com/doi/abs/10.1080/02331888.2015.
1100616.

Ka:2011:OSS

[KCNB11] C. Y. Ka, P. S. Chan, H. K. T. Ng, and N. Balakrishnan. Optimal sample size allocation for multi-level stress testing with Weibull regression under Type-II censoring. Statistics: a Journal of Theoretical and Applied Statistics, 45(3):257–279, 2011. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Kurotschka:1974:ODT

[KD74] V. G. Kurotschka and P. S. Dwyer. Optimal design of three way layouts without interactions. Mathematische Operationsforschung und Statistik, 5(2):131–145, 1974. CODEN MTOSBI. ISSN 0047-6277.

Kattumannil:2016:GMI

[KD16] Sudheesh K. Kattumannil and Isha Dewan. On generalized moment identity and its applications: a unified approach. Statistics: a Journal of Theoretical and Applied Statistics, 50(5):1149–??, 2016. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Kattumannil:2022:JEL

[KDM22] Sudheesh K. Kattumannil, Isha Dewan, and Litty Mathew. Jackknife empirical likelihood ratio test for testing mean residual life and mean past life ordering. Statistics: a Journal of Theoretical and Applied Statistics, 56(5):1012–1028, 2022. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/02331888.2022.2133120.

Kneip:1996:RKE

[KE96] Alois Kneip and Joachim Engel. A Remedy for kernel estimation under random design. Statistics: a Journal of Theoretical and Applied Statistics, 28(3):201–225, 1996. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Keri:1972:MSS

[Kér72] G. F. Kéri. A modified stepping-stone algorithm for the transportation problem. Mathematische Operationsforschung und Statistik, 3(5):327–331, 1972. CODEN MTOSBI. ISSN 0047-6277.

Kersten:1984:MADa

[Ker84a] Norbert Kersten. On the multidimension asymptotic distribution of covariance estimates of linear stationary processes. Statistics (Berlin, DDR), 15(4):525–532, 1984. CO-DEN MOSSD5. ISSN 0323-3944.

Kersten:1984:MADb

[Ker84b] Norbert Kersten. On the multidimensional asymptotic distribution of covariance estimates of stationary mixing sequences. Statistics (Berlin, DDR), 15(4):533–540, 1984. CO-DEN MOSSD5. ISSN 0323-3944.

Kerns:2016:CSC

[Ker16] Lucy Kerns. Construction of simultaneous confidence bands for multiple logistic regression models over restricted regions. Statistics: a Journal of Theoretical and Applied Statistics, 50 (6):1332–1345, 2016. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Kynclova:2016:CBI

[KFH16] P. Kynclová, P. Filzmoser, and K. Hron. Compositional biplots including external non-compositional variables. Statistics: a Journal of Theoretical and Applied Statistics, 50(5): 1132–??, 2016. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Kotz:1997:MVD

[KFL97] Samuel Kotz, Kai-Tai Fang, and Jia-Juan Liang. On multivariate vertical density representation and its application to random number generation. Statistics: a Journal of Theoretical and Applied Statistics, 30(2):163–180, 1997. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Kohansal:2021:MCS

[KFPG21] Akram Kohansal, Arturo J. Fernández, and Carlos J. Pérez-González. Multi-component stress-strength parameter estimation of a non-identical-component strengths system under the adaptive hybrid progressive censoring samples. Statistics: a Journal of Theoretical and Applied Statistics, 55(4): 925–962, 2021. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Kabe:1972:TSL

[KG72] D. G. Kabe and R. P. Gupta. On testing several linear relations. Mathematische Operationsforschung und Statistik, 3 (3):203–205, 1972. CODEN MTOSBI. ISSN 0047-6277.

Kumar:1981:RETa

[KG81a] Pradesh Kumar and V. K. Gupta. On ratio estimators in two phase sampling under size stratification and estimation over two successive occasions. Statistics (Berlin, DDR), 12 (3):409–417, 1981. CODEN MOSSD5. ISSN 0323-3944.

Kumar:1981:RETb

[KG81b] Pranesh Kumar and V. K. Gupta. On ratio estimators in two phase sampling under size stratification. Statistics (Berlin, DDR), 12(3):419–427, 1981. CODEN MOSSD5. ISSN 0323-3944.

Kundu:2013:PND

[KG13] Debasis Kundu and Rameshwar D. Gupta. Power-normal distribution. Statistics: a Journal of Theoretical and Applied Statistics, 47(1):110–125, 2013. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Kong:2021:HDM

[KH21] Xiaoli Kong and Solomon W. Harrar. High-dimensional MANOVA under weak conditions. Statistics: a Journal of Theoretical and Applied Statistics, 55(2):321–349, 2021. CO-DEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Khattree:1989:RTR

[Kha89] Ravindra Khattree. On robustness of tests for random effects in balanced nested designs. Statistics: a Journal of Theoretical and Applied Statistics, 20(1):41–46, 1989. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Khan:2004:AEF

[Kha04] Rasul A. Khan. Approximation for the expectation of a function of the sample mean. Statistics: a Journal of Theoretical and Applied Statistics, 38(2):117–122, 2004. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Kharroubi:2012:AMD

[Kha12] Samer A. Kharroubi. Approximate marginal densities of independent parameters. Statistics: a Journal of Theoretical and Applied Statistics, 46(4):459–471, 2012. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Khan:2015:REM

[Kha15] Rasul A. Khan. A remark on estimating the mean of a normal distribution with known coefficient of variation. Statistics: a Journal of Theoretical and Applied Statistics, 49(3):705–710, 2015. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Kabera:2015:ACO

[KHN15] Gaëtan M. Kabera, Linda M. Haines, and Principal Ndlovu. The analytic construction of D-optimal designs for the twovariable binary logistic regression model without interaction. Statistics: a Journal of Theoretical and Applied Statistics, 49 (5):1169–1186, 2015. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Khademnoe:2016:ADP

[KHN16] Omid Khademnoe and S. Mohammad E. Hosseini-Nasab. On asymptotic distribution of prediction in functional linear regression. Statistics: a Journal of Theoretical and Applied Statistics, 50(5):974–??, 2016. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Kundu:2020:RSS

[KHN20] Pradip Kundu, Nil Kamal Hazra, and Asok K. Nanda. Reliability study of series and parallel systems of heterogeneous component lifetimes following proportional odds model. Statistics: a Journal of Theoretical and Applied Statistics, 54(2):375–401, 2020. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Kohatsu-Higa:2017:LPE

[KHNT17] Arturo Kohatsu-Higa, Eulalia Nualart, and Ngoc Khue Tran. LAN property for an ergodic diffusion with jumps. Statistics: a Journal of Theoretical and Applied Statistics, 51(2):419– 454, 2017. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Khuri:1995:TDI

[Khu95] André I. Khuri. A test to detect inadequacy of Satterthwaite’s approximation in balanced mixed models. Statistics: a Journal of Theoretical and Applied Statistics, 27(1–2):45–54, 1995.

CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Kateri:2003:CCT

[KI03] Maria Kateri and George Iliopoulos. On collapsing categories in two-way contingency tables. Statistics: a Journal of Theoretical and Applied Statistics, 37(5):443–455, 2003. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Kiesse:2017:FSP

[Kie17] T. Senga Kiessé. On finite sample properties of nonparametric discrete asymmetric kernel estimators. Statistics: a Journal of Theoretical and Applied Statistics, 51(5):1046–1060, 2017. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Kim:2005:CTP

[Kim05] Hea-Jung Kim. On a class of two-piece skew-normal distributions. Statistics: a Journal of Theoretical and Applied Statistics, 39(6):537–553, 2005. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Kim:2007:CWN

[Kim07] Hea-Jung Kim. A class of weighted normal distributions and its variants useful for inequality constrained analysis. Statistics: a Journal of Theoretical and Applied Statistics, 41(5): 421–441, 2007. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Kim:2010:CWM

[Kim10] Hea-Jung Kim. A class of weighted multivariate distributions related to doubly truncated multivariate t-distribution. Statistics: a Journal of Theoretical and Applied Statistics, 44 (1):89–106, 2010. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Kim:2012:CCD

[Kim12] Hea-Jung Kim. On conditional correlations defined by a class of nonlinearly truncated distributions and their applications. Statistics: a Journal of Theoretical and Applied Statistics, 46(3):325–341, 2012. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Kim:2016:CRD

[Kim16] Hea-Jung Kim. A class of ratio distributions of dependent folded normals and its applications. Statistics: a Journal of Theoretical and Applied Statistics, 50(4):791–811, 2016. CO-DEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Kinder:1975:EKF

[Kin75] H. P. Kinder. Zur exponentiellen Konvergenzgeschwindigkeit der Fehlerwahrscheinlichkeiten 2. Art bei einigen Tests im Zweistichprobenfall. (German) [The exponential convergence rate of the error probabilities. 2. Kinds of tests in two random event]. Mathematische Operationsforschung und Statistik, 6 (2):239–252, 1975. CODEN MTOSBI. ISSN 0047-6277.

Kindler:1976:MYG

[Kin76] Jürgen Kindler. Über ein minimaxtheorem von Young. (German) [A minimax theorem of Young]. Mathematische Operationsforschung und Statistik, 7(3):477–480, 1976. CODEN MTOSBI. ISSN 0047-6277.

Kirstein:1976:MCT

[Kir76] B. M. Kirstein. Monotonicity and comparability of timehomogeneous Markov processes with discrete state space. Mathematische Operationsforschung und Statistik, 7(1):151– 168, 1976. CODEN MTOSBI. ISSN 0047-6277.

Kirstein:1983:BR

[Kir83] B. M. Kirstein. Book reviews. Statistics (Berlin, DDR), 14 (3):436, 1983. CODEN MOSSD5. ISSN 0323-3944.

Kozek:2014:DSA

[KJ14] Andrzej S. Kozek and Brian Jersky. Does sequential augmenting of simple linear heteroscedastic regression reduce variances of ordinary least-squares estimators? Statistics: a Journal of Theoretical and Applied Statistics, 48(5):1106–1121, 2014. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Konig:1976:EIA

[KJKR76] D. König, U. Jansen, M. Kotzturek, and H. Rabe. Ergebnisse von Invarianzuntersuchungen für ausgewählte Bedienungsund Zuverlässigkeitssysteme. (German) [Results of invariance analysis for selected operating systems and reliability systems]. Mathematische Operationsforschung und Statistik, 7 (4):523–555, 1976. CODEN MTOSBI. ISSN 0047-6277.

Klemm:1972:AGA

[KJM^{+}72] H. Klemm, Jena, G. Mönch, H. Hollatz, and H. Weinert. Ankündigung. (German) [Announcement]. Mathematische Operationsforschung und Statistik, 3(4):319–324, 1972. CO-DEN MTOSBI. ISSN 0047-6277.

Karpelevich:1980:SBM

[KK80] F. I. Karpelevich and A. Ja Kreinin. Some bounds for the mean waiting time in E_{k} /GX/1 queues. Statistics (Berlin, DDR), 11(1):85–88, 1980. CODEN MOSSD5. ISSN 0323-3944.

Kovalenko:1981:RPR

[KK81] I. N. Kovalenko and N. Ju. Kuznetsov. Renewal process and rare events limit theorems for essentially multidimensional queueing processes. Statistics (Berlin, DDR), 12(2):211–224, 1981. CODEN MOSSD5. ISSN 0323-3944.

Klesov:2011:DFW

[KK11] O. I. Klesov and N. V. Kruglova. The distribution of a functional of the Wiener process and its application to the Brownian sheet. Statistics: a Journal of Theoretical and Applied Statistics, 45(1):19–26, 2011. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Kirch:2014:UCL

[KK14] Claudia Kirch and Joseph Tadjuidje Kamgaing. A uniform central limit theorem for neural network-based autoregressive processes with applications to change-point analysis. Statistics: a Journal of Theoretical and Applied Statistics, 48(6): 1187–1201, 2014. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Kim:2015:CRS

[KK15] Hea-Jung Kim and Hyoung-Moon Kim. A class of rectanglescreened multivariate normal distributions and its applications. Statistics: a Journal of Theoretical and Applied Statistics, 49(4):878–899, 2015. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Kanika:2019:ERE

[KK19a] Kanika and Somesh Kumar. On efficiency and robustness of estimators for a spherical location. Statistics: a Journal of Theoretical and Applied Statistics, 53(3):601–629, 2019. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Kim:2019:HEP

[KK19b] Mihyun Kim and Piotr Kokoszka. Hill estimator of projections of functional data on principal components. Statistics: a Journal of Theoretical and Applied Statistics, 53(4):699–720, 2019. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Kumar:1984:SSI

[KKA84] Pranesh Kumar, O. P. Kathuria, and S. K. Agarwal. On a sampling scheme with inclusion probability proportional to size. Statistics (Berlin, DDR), 15(4):595–600, 1984. CODEN MOSSD5. ISSN 0323-3944.

Koley:2017:ATI

[KKG17] Arnab Koley, Debasis Kundu, and Ayon Ganguly. Analysis of Type-II hybrid censored competing risks data. Statistics: a Journal of Theoretical and Applied Statistics, 51(6):1304– 1325, 2017. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Kowalczyk:1982:SPC

[KL82] T. Kowalczyk and T. Ledwina. Some properties of chosen grade parameters and their rank counterparts. Statistics (Berlin, DDR), 13(4):547–553, 1982. CODEN MOSSD5. ISSN 0323-3944.

Klafszky:1972:DSP

[Kla72] Emil Klafszky. Determination of shortest path in a network with time-dependent edge-lengths. Mathematische Operationsforschung und Statistik, 3(4):255–257, 1972. CODEN MTOSBI. ISSN 0047-6277.

Klar:2003:TEA

[Kla03] B. Klar. On a test for exponentiality against Laplace order dominance. Statistics: a Journal of Theoretical and Applied Statistics, 37(6):505–515, 2003. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Klemm:1972:BFL

[Kle72] H. Klemm. Bericht über die Frühjahrstagung 1971 “Lagerhaltungsmodelle”. (German) [Report on the Spring 1971 meeting on “inventory models”]. Mathematische Operationsforschung und Statistik, 3(1):75, 1972. CODEN MTOSBI. ISSN 0047-6277.

Kleffe:1973:PCR

[Kle73] Jürgen Kleffe. Principal components of random variables with values in a separable Hilbert space. Mathematische Operationsforschung und Statistik, 4(5):391–406, 1973. CODEN MTOSBI. ISSN 0047-6277.

Kleffe:1975:BIQ

[Kle75a] J. Kleffe. Bayes invariant quadratic estimators for variance components in linear models. Mathematische Operationsforschung und Statistik, 6(5):753–767, 1975. CODEN MTOSBI. ISSN 0047-6277.

Kleffe:1975:BQU

[Kle75b] J. Kleffe. Best quadratic unbiased estimators for variance components in the balanced two-way classification model. Mathematische Operationsforschung und Statistik, 6(2):189– 196, 1975. CODEN MTOSBI. ISSN 0047-6277.

Kleffe:1976:NMN

[Kle76] J. Kleffe. A note on MINQUE for normal models. Mathematische Operationsforschung und Statistik, 7(5):707–714, 1976. CODEN MTOSBI. ISSN 0047-6277.

Kleffe:1977:IME

[Kle77a] J. Kleffe. Invariant methods for estimating variance components in mixed linear models. Statistics (Berlin, DDR), 8(2): 233–250, 1977. CODEN MOSSD5. ISSN 0323-3944.

Kleffe:1977:NMV

[Kle77b] J. Kleffe. A note on ∞-MINQUE in variance covariance components models. Statistics (Berlin, DDR), 8(3):337–343, 1977. CODEN MOSSD5. ISSN 0323-3944.

Kleffe:1978:SEE

[Kle78] J. Kleffe. Simultaneous estimation of expectation and covarianee matrix in linear models. Statistics (Berlin, DDR), 9(3): 443–478, 1978. CODEN MOSSD5. ISSN 0323-3944.

Kleffe:1980:RPM

[Kle80] J. Kleffe. On recent progress of MINQUE theory nonnegative estimation, consistency, asymptotic normality and explicite formulae. Statistics (Berlin, DDR), 11(4):563–588, 1980. CO-DEN MOSSD5. ISSN 0323-3944.

Kong:2022:LLQ

[KLF22] Cui-Juan Kong, Han-Ying Liang, and Guo-Liang Fan. Local likelihood of quantile difference under left-truncated, rightcensored and dependent assumptions. Statistics: a Journal of Theoretical and Applied Statistics, 57(1):71–93, 2022. CO-DEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/02331888.2022.2161547.

Klebanov:1978:SNF

[KM78] L. B. Klebanov and I. A. Melamed. Several notes on Fisher information in presence of nuisance parameters. Statistics (Berlin, DDR), 9(1):85–90, 1978. CODEN MOSSD5. ISSN 0323-3944.

Kshirsagar:1982:ERS

[KM82] A. M. Kshirsagar and Bonnie McKee. Estimation of a response surface from — a general incomplete block design. Statistics (Berlin, DDR), 13(4):543–546, 1982. CODEN MOSSD5. ISSN 0323-3944.

Kitsos:1995:RLC

[KM95] Christos P. Kitsos and Christine H. Müller. Robust linear calibration. Statistics: a Journal of Theoretical and Applied Statistics, 27(1–2):93–106, 1995. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Kundu:1996:NCU

[KM96] Debasis Kundu and Amit Mitra. A note on the consistency of the undamped exponential signals model. Statistics: a Journal of Theoretical and Applied Statistics, 28(1):25–33, 1996.

CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Kachiashvili:2013:CBT

[KM13] K. J. Kachiashvili and A. Mueed. Conditional Bayesian task of testing many hypotheses. Statistics: a Journal of Theoretical and Applied Statistics, 47(2):274–293, 2013. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Kubkowski:2017:ASP

[KM17] M. Kubkowski and J. Mielniczuk. Active sets of predictors for misspecified logistic regression. Statistics: a Journal of Theoretical and Applied Statistics, 51(5):1023–1045, 2017. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Kitani:2020:LDC

[KM20] Masato Kitani and Hidetoshi Murakami. The limiting distribution of combining the t and Wilcoxon rank sum tests. Statistics: a Journal of Theoretical and Applied Statistics, 54(4):871–884, 2020. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Kundu:2021:ACR

[KM21] Debasis Kundu and Shuvashree Mondal. Analyzing competing risks data using bivariate Weibull-geometric distribution. Statistics: a Journal of Theoretical and Applied Statistics, 55(2):276–295, 2021. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Kozachenko:2015:DPE

[KMM15] Y. Kozachenko, A. Melnikov, and Y. Mishura. On drift parameter estimation in models with fractional Brownian motion. Statistics: a Journal of Theoretical and Applied Statistics, 49(1):35–62, 2015. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Karlsson:2021:SIT

[KMM21] Sune Karlsson, Stepan Mazur, and Stanislas Muhinyuza. Statistical inference for the tangency portfolio in high dimension. Statistics: a Journal of Theoretical and Applied Statistics, 55(3):532–560, 2021. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Kukush:2010:OQS

[KMSS10] Alexander Kukush, Andrii Malenko, Hans Schneeweiss, and Shalabh. Optimality of quasi-score in the multivariate meanvariance model with an application to the zero-inflated Poisson model with measurement errors. Statistics: a Journal of Theoretical and Applied Statistics, 44(4):381–396, 2010. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Khattree:1990:OTR

[KN90] Ravindra Khattree and D. N. Naik. Optimum tests for random effects in unbalanced nested designs. Statistics: a Journal of Theoretical and Applied Statistics, 21(2):163–168, 1990. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Kleffe:1998:MNE

[KN98] Jürgen Kleffe and Ragnar Norberg. Minimum norm estimation under parameter constraints with an application to insurance. Statistics: a Journal of Theoretical and Applied Statistics, 31(3):215–234, 1998. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Kundu:2006:DDM

[KN06] Debasis Kundu and Swagata Nandi. On discrete-domain multidimensional sinusoidal models. Statistics: a Journal of Theoretical and Applied Statistics, 40(2):129–147, 2006. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Kayser:1972:KWT

[KNS72] Gerhard Kayser, Peter Neumann, and Rainer Salmen. Kritische werte für den X-test von van der Waerden. (German) [Critical values for van der Waerden’s X-test]. Mathematische Operationsforschung und Statistik, 3(5):389–400, 1972. CO-DEN MTOSBI. ISSN 0047-6277.

Kobonacki:1987:KES

[Kob87] Jacek Kobonacki. Kernel estimation of smooth densities using Fabian’s approach. Statistics: a Journal of Theoretical and Applied Statistics, 18(1):37–47, 1987. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Koch:1974:MLS

[Koc74] Rainer Koch. Maximum-Likelihood-Schätzungen von Parametern der Gamma- und Betaverteilung. (German) [Maximum likelihood estimates of the parameters of the gamma and beta distribution]. Mathematische Operationsforschung und Statistik, 5(4–5):333–355, 1974. CODEN MTOSBI. ISSN 0047-6277.

Kochel:1975:SLM

[Köc75] Peter Köchel. Ein stochastisches Lagerhaltungsmodell für mehrere miteinander verbundene Lager. (German) [A stochastic inventory model for multiple interconnected stock]. Mathematische Operationsforschung und Statistik, 6(3):413– 426, 1975. CODEN MTOSBI. ISSN 0047-6277.

Kochel:1976:MLS

[Köc76] Peter Köchel. Ein Minimax-Lagerhaltungsmodell für ein System von mehreren miteinander verbundenen Lagern. (German) [A minimax inventory model for a system of several interconnected stocks]. Mathematische Operationsforschung und Statistik, 7(6):905–925, 1976. CODEN MTOSBI. ISSN 0047-6277.

Kohl:2012:BIE

[Koh12] Matthias Kohl. Bounded influence estimation for regression and scale. Statistics: a Journal of Theoretical and Applied Statistics, 46(4):473–488, 2012. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Konecny:1987:ESI

[Kon87] Franz Konecny. On the estimation of the stochastic intensity and the parameters of Neyman–Scott trigger processes. Statistics: a Journal of Theoretical and Applied Statistics, 18(1): 113–118, 1987. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Konio:1989:BR

[Kon89] D. Konio. Book reviews. Statistics: a Journal of Theoretical and Applied Statistics, 20(4):614, 1989. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Koop:1986:SPS

[Koo86] J. C. Koop. Some problems of statistical inference from sample survey data for analytic studies. Statistics: a Journal of Theoretical and Applied Statistics, 17(2):237–247, 1986. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Korokacki:1984:NRE

[Kor84] J. Korokacki. A note on the révész’ estimator of a regression function. Statistics (Berlin, DDR), 15(2):195–203, 1984. CODEN MOSSD5. ISSN 0323-3944.

Koschitzki:1980:SSP

[Kos80] S. Koschitzki. Some stereological problems for random discs in R^{3}. Statistics (Berlin, DDR), 11(1):75–83, 1980. CODEN MOSSD5. ISSN 0323-3944.

Koukiou:2011:EDP

[Kou11] F. Koukiou. An example of dependence in a physical model. Statistics: a Journal of Theoretical and Applied Statistics, 45 (1):43–47, 2011. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Kovacs:1974:SLI

[Kov74] Lászlo Béla Kovács. Solution of linear integer programming problems by dynamic programming. Mathematische Operationsforschung und Statistik, 5(3):163–176, 1974. CODEN MTOSBI. ISSN 0047-6277.

Kowalczyk:1977:GDS

[Kow77] T. Kowalczyk. General definition and sample counterparts of monotonic dependence functions of bivariate distributions. Statistics (Berlin, DDR), 8(3):351–365, 1977. CODEN MOSSD5. ISSN 0323-3944.

Kowalczyk:1982:SMD

[Kow82] T. Kowalczyk. Shape of the monotone dependence function. Statistics (Berlin, DDR), 13(2):183–192, 1982. CODEN MOSSD5. ISSN 0323-3944.

Koziol:1978:MSR

[Koz78] James A. Koziol. Multivariate signed rank statistics for shift alternatives. Statistics (Berlin, DDR), 9(4):549–562, 1978. CODEN MOSSD5. ISSN 0323-3944.

Kozak:1985:MME

[Koz85] Josef Kozák. Modified minimax estimation of regression coefficients. Statistics: a Journal of Theoretical and Applied Statistics, 16(3):363–371, 1985. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Koziara:1986:ECM

[Koz86] Jan Koziara. Estimation of covariance matrices of vector Wiener process by MINQUE method. Statistics: a Journal of Theoretical and Applied Statistics, 17(4):569–575, 1986. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Koziol:1996:NSR

[Koz96] James A. Koziol. A note on signed rank tests for the changepoint problem. Statistics: a Journal of Theoretical and Applied Statistics, 27(3–4):325–338, 1996. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Kozubowski:2002:VDM

[Koz02] Tomasz J. Kozubowski. On the vertical density of the multivariate exponential power distribution. Statistics: a Journal of Theoretical and Applied Statistics, 36(3):219–221, 2002. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Kleffe:1974:BBQa

[KP74a] J. Kleffe and R. Pincus. Bates and best quadratic unbiased estimators for parameters of the covariance matrix in a normal linear model. Mathematische Operationsforschung und Statistik, 5(1):43–67, 1974. CODEN MTOSBI. ISSN 0047-6277.

Kleffe:1974:BBQb

[KP74b] J. Kleffe and R. Pincus. Bates and best quadratic unbiased estimators for variance components and heteroscedastie variances in linear models. Mathematische Operationsforschung und Statistik, 5(2):147–159, 1974. CODEN MTOSBI. ISSN 0047-6277.

Kyriakoussis:1991:CHT

[KP91] A. Kyriakoussis and H. Papageorgiou. Characterization of hypergeometric type distributions by regression. Statistics: a Journal of Theoretical and Applied Statistics, 22(3):467–477, 1991. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Kutoyants:1994:PEK

[KP94] Yu. Kutoyants and H. Pohlmann. Parameter estimation for Kalman–Bucy filter with small noise. Statistics: a Journal of Theoretical and Applied Statistics, 25(4):307–323, 1994. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Kokonendji:2007:CUM

[KP07] C. C. Kokonendji and D. Pommeret. Comparing UMVU and ML estimators of the generalized variance for natural exponential families. Statistics: a Journal of Theoretical and Applied Statistics, 41(6):547–558, 2007. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Kaminska:2009:RBE

[KP09] Agnieszka Kamińska and Zdzisław Porosiński. On robust Bayesian estimation under some asymmetric and bounded loss function. Statistics: a Journal of Theoretical and Applied Statistics, 43(3):253–265, 2009. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Kim:2015:PLM

[KP15] Seongyoung Kim and Yousung Park. Power-linear models for incomplete contingency tables with nonignorable nonresponses. Statistics: a Journal of Theoretical and Applied Statistics, 49(6):1189–1203, 2015. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Kleffe:1975:RGB

[KPAF75] J. Kleffe, I. Peterson, H. Ahrens, and W. Fischer. Rezensionen. (German) [Book reviews]. Mathematische Operationsforschung und Statistik, 6(5):831–834, 1975. CODEN MTOSBI. ISSN 0047-6277.

Kim:2016:UDW

[KPH^{+}16] Taeyoon Kim, Cheolyong Park, Jeongcheol Ha, Zhi-Ming Luo, and Sun Young Hwang. Using the dependent wild bootstrap for the nonparametric goodness-of-fit test for density functions. Statistics: a Journal of Theoretical and Applied Statistics, 50(4):750–774, 2016. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Koziol:1978:MCA

[KR78] James A. Koziol and Nancy Reid. On multiple comparisons among k populations differing by scale parameters. Statistics (Berlin, DDR), 9(2):179–184, 1978. CODEN MOSSD5. ISSN 0323-3944.

Kundu:2007:DBG

[KR07] Debasis Kundu and Mohammad Z. Raqab. Discriminating between the generalized Rayleigh and log-normal distribution. Statistics: a Journal of Theoretical and Applied Statistics, 41(6):505–515, 2007. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Kozyra:2017:ULB

[KR17] Pawel Marcin Kozyra and Tomasz Rychlik. Upper and lower bounds on the variances of linear combinations of the k-th records. Statistics: a Journal of Theoretical and Applied Statistics, 52(1):177–204, 2017. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Krauth:1976:SLO

[Kra76] Joachim Krauth. some locally optimal two-sample bank tests of scale. Mathematische Operationsforschung und Statistik, 7 (1):113–121, 1976. CODEN MTOSBI. ISSN 0047-6277.

Krauth:1979:GSR

[Kra79] J. Krauth. Generalized sequential ranks and tests of randomness. Statistics (Berlin, DDR), 10(2):291–298, 1979. CODEN MOSSD5. ISSN 0323-3944.

Krafft:1986:MLE

[Kra86] Olaf Krafft. A maximin linear estimator for linear parameters under restrictions in form of inequalities. Statistics: a Journal of Theoretical and Applied Statistics, 17(1):3–8, 1986. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Krafft:1995:D

[Kra95] Olaf Krafft. Discussion. Statistics: a Journal of Theoretical and Applied Statistics, 26(2):162, 1995. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Kim:2019:BAS

[KRC19] Hea-Jung Kim, Taeyoung Roh, and Taeryon Choi. Bayesian analysis of semiparametric Bernstein polynomial regression models for data with sample selection. Statistics: a Journal of Theoretical and Applied Statistics, 53(5):1082–1111, 2019. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Kreinin:1980:BMC

[Kre80] A. Ja Kreinin. Bounds for mean characteristics of E_{k} /GI/1/∞ queues. Statistics (Berlin, DDR), 11(4):515–519, 1980. CODEN MOSSD5. ISSN 0323-3944.

Kremer:1985:BTS

[Kre85] Erhard Kremer. Behavior of two-sample rank tests at infinity. Statistics: a Journal of Theoretical and Applied Statistics, 16(4):527–534, 1985. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Krebs:2018:NPR

[Kre18] Johannes T. N. Krebs. Non-parametric regression for spatially dependent data with wavelets. Statistics: a Journal of Theoretical and Applied Statistics, 52(6):1270–1308, 2018. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Kleffe:1977:BR

[KRF^{+}77] J. Kleffe, W. Randt, K. Fischer, H. Bandemer, O. Bunke, H. Toutenburg, J. Bellach, and P. Küchler. Book reviews. Statistics (Berlin, DDR), 8(2):256–260, 1977. CODEN MOSSD5. ISSN 0323-3944.

Krzatschmer:2005:GFS

[Krz05] Volker Krzätschmer. A generalized framework of sampling inspections by attributes. Statistics: a Journal of Theoretical and Applied Statistics, 39(5):445–455, 2005. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Kaiser:1971:RGB

[KS71] E. Kaiser and H. Sachs. Rezensionen. (German) [Book reviews]. Mathematische Operationsforschung und Statistik, 2 (1):87–88, 1971. CODEN MTOSBI. ISSN 0047-6277.

Konig:1976:BSA

[KS76a] D. König and V. Schmidt. Bedienungsmodelle mit stochastischen Abhängigkeiten. (German) [Operation models with stochastic dependencies]. Mathematische Operationsforschung und Statistik, 7(4):607–639, 1976. CODEN MTOSBI. ISSN 0047-6277.

Kotzturek:1976:QCT

[KS76b] M. Kotzturek and D. Stoyan. A quantitative continuity theorem for the mean stationary waiting time in GI/GI/1. Mathematische Operationsforschung und Statistik, 7(4):595–599, 1976. CODEN MTOSBI. ISSN 0047-6277.

Kuchler:1979:WFS

[KS79] Ingeborg Küchler and Alexander Semjonov. Die Waldsche Fundamentalidentität und ein sequentieller Quotiententest für eine zufällige Irrfahrt über einer homogenen irreduziblen Markovschen Kette mit endlichem Zustandsraum. (German) [The Wald fundamental identity and a sequential quotient test for a random walk on a homogeneous irreducible Markov chain rule with finite state space]. Statistics (Berlin, DDR), 10(2):319–331, 1979. CODEN MOSSD5. ISSN 0323-3944.

Kaul:1983:CHR

[KS83a] C. L. Kaul and Kanwab Sen. Crossings of a height and related bank order statistics. Statistics (Berlin, DDR), 14(4):527– 553, 1983. CODEN MOSSD5. ISSN 0323-3944.

Kaul:1983:DUR

[KS83b] C. L. Kaul and Kanwar Sen. Distribution of upcrossings and related rank order statistics. Statistics (Berlin, DDR), 14(2): 257–261, 1983. CODEN MOSSD5. ISSN 0323-3944.

Kant:1986:SEP

[KS86] Vishnu Kant and Divakar Sharma. Simultaneous estimation of λ^{s} for Poisson distributions. Statistics: a Journal of Theoretical and Applied Statistics, 17(4):539–547, 1986. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Kotlarski:1992:CPS

[KS92] Ignacy Icchak Kotlarski and Zoltán Sasvári. On a characterization problem of statistics. Statistics: a Journal of Theoretical and Applied Statistics, 23(1):85–93, 1992. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Kumar:1993:UIL

[KS93] Somesh Kumar and Divakar Sharma. Unbiased inestimability of the larger of two parameters. Statistics: a Journal of Theoretical and Applied Statistics, 24(2):137–142, 1993. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Kumar:1995:ECL

[KS95] Somesh Kumar and Divakar Sharma. Estimating the common location. Statistics: a Journal of Theoretical and Applied Statistics, 26(3):231–240, 1995. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Korostelev:1996:EAM

[KS96] A. P. Korostelev and V. G. Spokoiny. Exact asymptotics of minimax Bahadur risk in Lipschitz regression. Statistics: a Journal of Theoretical and Applied Statistics, 28(1):13–24, 1996. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Kinkawa:2010:EOP

[KS10] Takuya Kinkawa and Nobuo Shinozaki. Estimation of the optimal portfolio weights by shrinking the mean vector towards a linear subspace. Statistics: a Journal of Theoretical and Applied Statistics, 44(6):629–640, 2010. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Kawasaki:2016:BCT

[KS16] Tamae Kawasaki and Takashi Seo. Bias correction for T^{2} type statistic with two-step monotone missing data. Statistics: a Journal of Theoretical and Applied Statistics, 50(1):76–88, 2016. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910

(electronic). URL http://www.tandfonline.com/doi/abs/
10.1080/02331888.2015.1060238.

Konig:1976:SRB

[KSS76] D. König, V. Schmidt, and D. Stoyan. On some relations between stationary distributions of queue lengths and imbedded queue lengths in G/G/s queueing systems. Mathematische Operationsforschung und Statistik, 7(4):577–586, 1976. CODEN MTOSBI. ISSN 0047-6277.

Kaushal:2011:SPS

[KSS11] R. P. Kaushal, B. V. S. Sisodia, and U. C. Sud. Shrunken predictors in stratified sampling under super-population model. Statistics: a Journal of Theoretical and Applied Statistics, 45(3):281–291, 2011. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Kleffe:1983:IMQ

[KT83] J. Kleffe and R. Thrum. Inequalities for moments of quadratic forms with applications to a.s. convergence. Statistics (Berlin, DDR), 14(2):211–216, 1983. CODEN MOSSD5. ISSN 0323-3944.

Kowalczyk:1989:MGD

[KT89] Teresa Kowalczyk and Joanna Tyrcha. Multivariate gamma distributions — properties and shape estimation. Statistics: a Journal of Theoretical and Applied Statistics, 20(3):465–474, 1989. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Kotz:1996:VDR

[KT96] Samuel Kotz and Marvin D. Troutt. On vertical density representation and ordering of distributions. Statistics: a Journal of Theoretical and Applied Statistics, 28(3):241–247, 1996. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Kattumannil:2012:MID

[KT12] Sudheesh Kumar Kattumannil and Luisa Tibiletti. Moment identity for discrete random variable and its applications. Statistics: a Journal of Theoretical and Applied Statistics, 46(6):767–775, 2012. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Kormann:1987:DMM

[KTW87] U. Kormann, R. Theodorescu, and H. Wolff. A dynamic method of moments. Statistics: a Journal of Theoretical and Applied Statistics, 18(1):131–140, 1987. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Kuchler:1978:SQI

[Küc78] Ingeborg Küchler. Der Sequentielle Quotiententest bei irreduziblen homogenen Markovschen Ketten mit endlichem Zustandsraum. (German) [The sequential ratio test for irreducible homogeneous Markov chains with finite state space]. Statistics (Berlin, DDR), 9(2):227–230, 1978. CODEN MOSSD5. ISSN 0323-3944.

Kuchler:1982:EFMa

[Küc82a] Uwe Küchler. Exponential families of Markov processes — Part I. General results. Statistics (Berlin, DDR), 13(1):57– 69, 1982. CODEN MOSSD5. ISSN 0323-3944.

Kuchler:1982:EFMb

[Küc82b] Uwe Küchler. Exponential families of Markov processes — Part II: Birth- and death processes. Statistics (Berlin, DDR), 13(2):219–230, 1982. CODEN MOSSD5. ISSN 0323-3944.

Kumar:1985:SAC

[Kum85] V. S. Sampath Kumar. Some aspects of continuous sampling plans for Markov-dependent production processes. Statistics: a Journal of Theoretical and Applied Statistics, 16(4):569– 576, 1985. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Kumar:1987:STU

[Kum87] Prenesh Kumar. On sampling of three units using varying probabilities. Statistics: a Journal of Theoretical and Applied Statistics, 18(3):373–377, 1987. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Kumar:1989:SAG

[Kum89] V. S. Sampath Kumar. Some aspects of general m-estimation theory for dependent random variables. Statistics: a Journal of Theoretical and Applied Statistics, 20(1):95–107, 1989. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Kumazawa:1992:TIF

[Kum92] Yoshiki Kumazawa. Tests for increasing failure rate with randomly censored data. Statistics: a Journal of Theoretical and Applied Statistics, 23(1):17–25, 1992. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Kumar:2013:TMU

[Kum13] Nirpeksh Kumar. Test for multiple upper outliers in an exponential sample irrespective of origin. Statistics: a Journal of Theoretical and Applied Statistics, 47(1):184–190, 2013. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Kundu:1997:ATL

[Kun97] Debasis Kundu. Asymptotic theory of the least squares estimators of sinusoidal signal. Statistics: a Journal of Theoretical and Applied Statistics, 30(3):221–238, 1997. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Kundu:2015:BLB

[Kun15] Debasis Kundu. Bivariate log Birnbaum–Saunders distribution. Statistics: a Journal of Theoretical and Applied Statistics, 49(4):900–917, 2015. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Kundu:2017:MGS

[Kun17] Debasis Kundu. Multivariate geometric skew-normal distribution. Statistics: a Journal of Theoretical and Applied Statistics, 51(6):1377–1397, 2017. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Kutoyants:1984:PED

[Kut84] Yu A. Kutoyants. Parameter estimation for diffusion type processes of observation. Statistics (Berlin, DDR), 15(4):541– 551, 1984. CODEN MOSSD5. ISSN 0323-3944.

Kutil:2012:BUE

[Kut12] Rade Kutil. Biased and unbiased estimation of the circular mean resultant length and its variance. Statistics: a Journal of Theoretical and Applied Statistics, 46(4):549–561, 2012. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Kollo:2000:DDA

[KvR00] Tõnu Kollo and Dietrich von Rosen. Distribution and density approximation of the covariance matrix in the growth curve model. Statistics: a Journal of Theoretical and Applied Statistics, 35(1):1–22, 2000. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Kim:2015:GFT

[KW15] Hee-Young Kim and Christian H. Weiß. Goodness-of-fit tests for binomial AR(1) processes. Statistics: a Journal of Theoretical and Applied Statistics, 49(2):291–315, 2015. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Karczewski:2017:LRW

[KW17] Rafal Karczewski and Jacek Wesolowski. Linearity of regression for weak records, revisited. Statistics: a Journal of Theoretical and Applied Statistics, 51(4):878–887, 2017. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Kokoszka:2016:KTF

[KY16] Piotr Kokoszka and Gabriel Young. KPSS test for functional time series. Statistics: a Journal of Theoretical and Applied Statistics, 50(5):957–??, 2016. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Kuran:2022:KLP

[KY22] Özge Kuran and Seçil Yalaz. Kernel Liu prediction approach in partially linear mixed measurement error models. Statistics: a Journal of Theoretical and Applied Statistics, 56(6): 1385–1408, 2022. CODEN MOSSD5. ISSN 0233-1888 (print),

1029-4910 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/02331888.2022.2152816.

Kadyrova:1989:IML

[KYR89a] N. O. Kadyrova, A. Yu Yakovlev, and S. T. Rachev. Isotonic maximum likelihood estimation of bimodal failure rate — a computer-based study. Statistics: a Journal of Theoretical and Applied Statistics, 20(2):271–278, 1989. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Kadyrova:1989:MLE

[KYR89b] N. O. Kadyrova, A. Yu Yakovlev, and S. T. Rachev. Maximum likelihood estimation of the bimodal failure rate for censored and tied observations. Statistics: a Journal of Theoretical and Applied Statistics, 20(1):135–140, 1989. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Kleffe:1978:QEH

[KZ78] J. Kleffe and I. Zöllner. On quadratic estimation of heteroscedastic variances. Statistics (Berlin, DDR), 9(1):27–44, 1978. CODEN MOSSD5. ISSN 0323-3944.

Kleinwachter:1986:SDS

[KZ86] A. Kleinwächter and M. Zähle. Size distribution stereology for quasiellipsoids in R^{n}. Statistics: a Journal of Theoretical and Applied Statistics, 17(3):435–444, 1986. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Klonecki:1987:AIE

[KZ87] W. Klonecki and S. Zontek. On admissible invariant estimators of variance components which dominate unbiased invariant estimators. Statistics: a Journal of Theoretical and Applied Statistics, 18(4):483–498, 1987. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Kuljus:2013:ALL

[KZ13a] Kristi Kuljus and Silvelyn Zwanzig. Asymptotic linearity of a linear rank statistic in the case of symmetric nonidentically distributed variables. Statistics: a Journal of Theoretical and Applied Statistics, 47(1):156–171, 2013. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Kuljus:2013:APR

[KZ13b] Kristi Kuljus and Silvelyn Zwanzig. Asymptotic properties of a rank estimate in linear regression with symmetric nonidentically distributed errors. Statistics: a Journal of Theoretical and Applied Statistics, 47(6):1160–1183, 2013. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Kulawik:2015:REM

[KZ15] Agnieszka Kulawik and Stefan Zontek. Robust estimation in the multivariate normal model with variance components.

Statistics: a Journal of Theoretical and Applied Statistics, 49(4):766–780, 2015. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Khovansky:2020:NCS

[KZ20] Serguey Khovansky and Oleksandr Zhylyevskyy. A note on a cross-sectional GMM estimator in the presence of an observable common shock. Statistics: a Journal of Theoretical and Applied Statistics, 54(2):415–423, 2020. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Lai:2000:ADE

[Lai00] Dejian Lai. Asymptotic distribution of the estimated BDS statistic from the residuals of location-scale type processes. Statistics: a Journal of Theoretical and Applied Statistics, 34(2):117–135, 2000. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

LaMotte:2017:PSN

[LaM17] Lynn R. LaMotte. Proportional subclass numbers in twofactor ANOVA. Statistics: a Journal of Theoretical and Applied Statistics, 52(1):228–238, 2017. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic). See corrigenda [Ano19].

Lando:2021:SOS

[LAO21] Tommaso Lando, Idir Arab, and Paulo Eduardo Oliveira. Second-order stochastic comparisons of order statistics. Statistics: a Journal of Theoretical and Applied Statistics, 55(3):561–579, 2021. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Last:1991:BFD

[Las91] Gunter Last. On Bayes’ formula for doubly stochastic point process on the real half-line. Statistics: a Journal of Theoretical and Applied Statistics, 22(3):431–448, 1991. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Latter:1974:ANU

[Lät74] H. Lätter. On the admissibility and nonadmissibility of the usual estimator for the mean of multivariate normal population and conclusions to optimal design. Mathematische Operationsforschung und Statistik, 5(7–8):591–597, 1974. CODEN MTOSBI. ISSN 0047-6277.

Lauter:1970:OVS

[Läu70] Henning Läuter. Optimale Vorhersage und Schätzung in regulären und singulären Regressionsmodellen. (German) [Optimal prediction and estimation in regular and singular regression models]. Mathematische Operationsforschung und Statistik, 1(3):229–243, 1970. CODEN MTOSBI. ISSN 0047-6277.

Lauter:1971:OSS

[Läu71a] Henning Läuter. Optimale Schätzung bei stochastischen Prozessen mit linearem Regressionsanteil. (German) [Optimal estimation in stochastic processes with linear regression share]. Mathematische Operationsforschung und Statistik, 2 (3):237–246, 1971. CODEN MTOSBI. ISSN 0047-6277.

Lauter:1971:VSPa

[Läu71b] Henning Läuter. Vorhersage bei stochastischen Prozessen mit linearem Regressionsanteil. I. (German) [Prediction in stochastic processes with linear regression share. I]. Mathematische Operationsforschung und Statistik, 2(1):69–85, 1971. CODEN MTOSBI. ISSN 0047-6277.

Lauter:1971:VSPb

[Läu71c] Henning Läuter. Vorhersage bei stochastischen Prozessen mit linearem Regressionsanteil. II. (German) [Prediction in stochastic processes with linear regression share. II]. Mathematische Operationsforschung und Statistik, 2(2):147–166, 1971. CODEN MTOSBI. ISSN 0047-6277.

Lauter:1974:EDC

[Läu74] Elisabeth Läuter. Experimental design in a class of models. Mathematische Operationsforschung und Statistik, 5(4–5): 379–398, 1974. CODEN MTOSBI. ISSN 0047-6277.

Lauter:1975:LME

[Läu75a] H. Läuter. Linear minimax estimation for inadequate models for the I. Petersex approach. Mathematische Operationsforschung und Statistik, 6(5):769–774, 1975. CODEN MTOSBI. ISSN 0047-6277.

Lauter:1975:MLE

[Läu75b] H. Läuter. A minimax linear estimator for linear parameters under restrictions in form of inequalities. Mathematische Operationsforschung und Statistik, 6(5):689–695, 1975. CODEN MTOSBI. ISSN 0047-6277.

Lauter:1976:OMD

[Läu76] E. Läuter. Optimal multipurpose designs for regression models. Mathematische Operationsforschung und Statistik, 7(1): 51–68, 1976. CODEN MTOSBI. ISSN 0047-6277.

Lauter:1985:EEE

[Lau85a] Henning Lauter. An efficient estimator for the error rate in discriminant analysis. Statistics: a Journal of Theoretical and Applied Statistics, 16(1):107–119, 1985. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Lauter:1985:DAU

[Lau85b] Jurcen Lauter. Discriminant analysis under parameter restrictions statistical and computational aspects. Statistics: a Journal of Theoretical and Applied Statistics, 16(1):125–137, 1985. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Lauter:1989:NSC

[Läu89] Henning Läuter. Note on the strong consistency of the least squares estimator in nonlinear regression. Statistics: a Journal of Theoretical and Applied Statistics, 20(2):199–210, 1989. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Lavergne:1995:JME

[Lav95] Christian Lavergne. A jackknife method for estimation of variance components. Statistics: a Journal of Theoretical and Applied Statistics, 27(1–2):1–13, 1995. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Lazar:2016:CMP

[Laz16] Daniel Lazar. Conditional Γ-minimax prediction with a precautionary loss function in a marked point process model. Statistics: a Journal of Theoretical and Applied Statistics, 50 (6):1411–1420, 2016. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Loots:2013:RRQ

[LBAR13] Mattheüs T. Loots, Andriëtte Bekker, Mohammad Arashi, and Jacobus J. J. Roux. On the real representation of quaternion random variables. Statistics: a Journal of Theoretical and Applied Statistics, 47(6):1224–1240, 2013. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Linke:2023:UKT

[LBR23] Y. Y. Linke, I. S. Borisov, and P. S. Ruzankin. Universal kernel-type estimation of random fields. Statistics: a Journal of Theoretical and Applied Statistics, 57(4):785–810, 2023. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/02331888.2023.2231114.

Lopez-Blazquez:2000:EBW

[LBSM00] F. López-Blázquez and B. Salamanca-Miño. Estimation based on the Winsorized mean in the geometric distribution. Statistics: a Journal of Theoretical and Applied Statistics, 35(1): 81–95, 2000. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Lopez-Blazquez:2002:LDU

[LBSM02] F. López-Blázquez and B. Salamanca-Miño. Limit distributions of unbiased estimators in natural exponential families. Statistics: a Journal of Theoretical and Applied Statistics, 36(4):329–338, 2002. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Lopez-Blazquez:2019:OSO

[LBSW19] F. López-Blázquez, Nan-Cheng Su, and Jacek Wesołowski. Order statistics from overlapping samples: bivariate densities and regression properties. Statistics: a Journal of Theoretical and Applied Statistics, 53(5):1052–1081, 2019. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Lombardo:2010:CTW

[LC10] R. Lombardo and I. Camminatiello. CATANOVA for two-way cross classified categorical data. Statistics: a Journal of Theoretical and Applied Statistics, 44(1):57–71, 2010. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Lemonte:2013:ELD

[LC13] Artur J. Lemonte and Gauss M. Cordeiro. An extended Lomax distribution. Statistics: a Journal of Theoretical and Applied Statistics, 47(4):800–816, 2013. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Laha:2015:RTD

[LC15] Arnab Kumar Laha and Mahesh K. C. Robustness of tests for directional mean. Statistics: a Journal of Theoretical and Applied Statistics, 49(3):522–536, 2015. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Lee:2017:RSP

[LC17] Hyunju Lee and Ji Hwan Cha. Reliability sampling plan for repairable items following general failure process and its statistical analysis. Statistics: a Journal of Theoretical and Applied Statistics, 51(5):1159–1178, 2017. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Laumen:2019:PCF

[LC19] Benjamin Laumen and Erhard Cramer. Progressive censoring with fixed censoring times. Statistics: a Journal of Theoretical and Applied Statistics, 53(3):569–600, 2019. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Lee:2021:FSJ

[LCK21] Sang Kyu Lee, Jae Ho Chang, and Hyoung-Moon Kim. Further sharpening of Jensen’s inequality. Statistics: a Journal of Theoretical and Applied Statistics, 55(5):1154–1168, 2021. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Lemonte:2016:NUT

[LCMA16] Artur J. Lemonte, Gauss M. Cordeiro, and Germán Moreno-Arenas. A new useful three-parameter extension of the exponential distribution. Statistics: a Journal of Theoretical and Applied Statistics, 50(2):312–337, 2016. CO-DEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/02331888.2015.1095190.

Louzada:2015:LWN

[LCY15] Francisco Louzada, Vicente G. Cancho, and Bao Yiqi. The log-Weibull-negative-binomial regression model under latent failure causes and presence of randomized activation schemes. Statistics: a Journal of Theoretical and Applied Statistics, 49(4):930–949, 2015. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Li:2017:TIT

[LCY17] Weiming Li, Jiaqi Chen, and Jianfeng Yao. Testing the independence of two random vectors where only one dimension is large. Statistics: a Journal of Theoretical and Applied Statistics, 51(1):141–153, 2017. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Lin:2021:ITI

[LD21] Gwo Dong Lin and Xiaoling Dou. An identity for two integral transforms applied to the uniqueness of a distribution via its Laplace–Stieltjes transform. Statistics: a Journal of Theoretical and Applied Statistics, 55(2):367–385, 2021. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Liang:2018:WEC

[LdCIP18] Han-Ying Liang and María del Carmen Iglesias-Pérez. Weighted estimation of conditional mean function with truncated, censored and dependent data. Statistics: a Journal of Theoretical and Applied Statistics, 52(6):1249–1269, 2018. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Li:2017:SID

[LDF17] Xiaoyin Li, Paul Doukhan, and Jean-Paul Feugeas. Statistical inference for DNA sequences of promoters: a non-stationary qualitative model. Statistics: a Journal of Theoretical and Applied Statistics, 51(1):154–166, 2017. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Ledwina:1978:ATC

[Led78a] Teresa Ledwina. Admissible tests for contingency tables and Poisson’s distributions, II. Statistics (Berlin, DDR), 9(1): 119–125, 1978. CODEN MOSSD5. ISSN 0323-3944.

Ledwina:1978:ATE

[Led78b] Teresa Ledwina. Admissible tests for exponential families with finite support. Statistics (Berlin, DDR), 9(1):105–118, 1978. CODEN MOSSD5. ISSN 0323-3944.

Ledwina:1983:RTI

[Led83] Teresa Ledwina. Robust test of independence of two random variables. Statistics (Berlin, DDR), 14(2):251–256, 1983. CO-DEN MOSSD5. ISSN 0323-3944.

Ledwina:1984:NAS

[Led84] Teresa Ledwina. A note on admissibility of some tests of independence against “positive dependence” in r×c contingeney table. Statistics (Berlin, DDR), 15(4):565–570, 1984. CO-DEN MOSSD5. ISSN 0323-3944.

Lee:1995:D

[Lee95] Carl Lee. Discussion. Statistics: a Journal of Theoretical and Applied Statistics, 26(2):162–166, 1995. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Lee:2008:NPB

[Lee08] Jaeyong Lee. Non-parametric Bayesian analysis of clustered survival data. Statistics: a Journal of Theoretical and Applied Statistics, 42(6):515–526, 2008. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Leifman:1970:MPO

[Lei70] Lev Ja Leifman. The main principles for optimal allocation of capacities in network planning. Mathematische Operationsforschung und Statistik, 1(4):281–288, 1970. CODEN MTOSBI. ISSN 0047-6277.

Lemonte:2013:NAD

[Lem13] Artur J. Lemonte. Nonnull asymptotic distributions of the LR, Wald, score and gradient statistics in generalized linear models with dispersion covariates. Statistics: a Journal of Theoretical and Applied Statistics, 47(6):1249–1265, 2013. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Li:2022:CBN

[LF22] Xiaohu Li and Rui Fang. A count-based nonparametric test on strict bivariate stochastic arrangement increasing property. Statistics: a Journal of Theoretical and Applied Statistics, 56(3):518–536, 2022. CO-DEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/02331888.2022.2074005.

Lopez-Fidalgo:2011:OED

[LFGR11] Jesús López-Fidalgo and Sandra Garcet-Rodríguez. Optimal experimental designs when an independent variable is potentially censored. Statistics: a Journal of Theoretical and Applied Statistics, 45(2):143–154, 2011. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Lauter:2005:TPC

[LG05] Jürgen Läuter and Ekkehard Glimm. A theorem on the principal components inference. Statistics: a Journal of Theoretical and Applied Statistics, 39(3):207–219, 2005. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Lv:2017:ARE

[LGL^{+}17] Jing Lv, Chaohui Guo, Tingting Li, Yuanyuan Hao, and Xiaolin Pan. Adaptive robust estimation in joint meancovariance regression model for bivariate longitudinal data. Statistics: a Journal of Theoretical and Applied Statistics, 52 (1):64–83, 2017. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Liu:2014:NST

[LGW14] Xu-Qing Liu, Feng Gao, and Xun-Qing Wu. New shrinkagetype estimators in a linear regression model when multicollinearity is severe. Statistics: a Journal of Theoretical and Applied Statistics, 48(2):344–358, 2014. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Li:2014:VSG

[LGZL14] Jianbo Li, Minggao Gu, Riquan Zhang, and Heng Lian. Variable selection for general transformation models with ranking data. Statistics: a Journal of Theoretical and Applied Statistics, 48(1):81–100, 2014. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Lorz:1991:NPA

[LH91] U. Lorz and L. Heinrich. Normal and Poisson approximation of infinitely divisible distribution functions. Statistics: a Journal of Theoretical and Applied Statistics, 22(4):627–649, 1991. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Liu:2013:GRP

[LH13] Xu-Qing Liu and Ping Hu. General ridge predictors in a mixed linear model. Statistics: a Journal of Theoretical and Applied Statistics, 47(2):363–378, 2013. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Liu:2015:EPC

[LH15] Xiaoqian Liu and Jianhua Hu. Estimation of patterned covariance in the multivariate linear models: an outer product least-squares approach. Statistics: a Journal of Theoretical and Applied Statistics, 49(2):408–426, 2015. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Lu:2022:EHO

[LH22] Wanbo Lu and Guanglin Huang. Estimating the higher-order co-moment with non-Gaussian components and its application in portfolio selection. Statistics: a Journal of Theoretical and Applied Statistics, 56(3):537–564, 2022. CO-DEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/02331888.2022.2074006.

Lian:2012:SAI

[Lia12] Heng Lian. Stochastic adaptation of importance sampler. Statistics: a Journal of Theoretical and Applied Statistics, 46(6):777–785, 2012. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Liese:1972:AIB

[Lie72] Friedrich Liese. Zur Anwendung der Informationstheorie beim Beweis von Grenzwertsätzen der Wahrscheinlichkeitstheorie. (German) [On the application of information theory in the proof of limit theorems of probability theory]. Mathematische Operationsforschung und Statistik, 3(2):113–128, 1972. CO-DEN MTOSBI. ISSN 0047-6277.

Liero:1982:MDK

[Lie82] H. Liero. On the maximal deviation of the kernel regression function estimate. Statistics (Berlin, DDR), 13(2):171–182, 1982. CODEN MOSSD5. ISSN 0323-3944.

Liese:1986:HID

[Lie86] F. Liese. Hellinger integrals of diffusion processes. Statistics: a Journal of Theoretical and Applied Statistics, 17(1):63–78, 1986. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Liero:1988:BR

[Lie88] H. Liero. Book review. Statistics: a Journal of Theoretical and Applied Statistics, 19(3):464, 1988. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Liebscheb:1990:KEP

[Lie90a] E. Liebscheb. Kernel estimators for probability densities with discontinuities. Statistics: a Journal of Theoretical and Applied Statistics, 21(2):185–196, 1990. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Liese:1990:BR

[Lie90b] F. Liese. Book reviews. Statistics: a Journal of Theoretical and Applied Statistics, 21(4):576, 1990. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Liebscher:1992:DEM

[Lie92] E. Liebscher. Density estimation for Markov chains. Statistics: a Journal of Theoretical and Applied Statistics, 23(1): 27–48, 1992. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Liero:1993:NAB

[Lie93] H. Liero. A note on the asymptotic behaviour of the distance of the k_{n} th nearest neighbour. Statistics: a Journal of Theoretical and Applied Statistics, 24(3):235–243, 1993. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Liebscher:1998:CHS

[Lie98] Eckhard Liebscher. Convergence of Hermite series density estimators under conditions of weak dependence. Statistics: a Journal of Theoretical and Applied Statistics, 31(3):191–214, 1998. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Likes:1985:EQP

[Lik85] Jir Likes. Estimation of the quantiles of Pareto’s distribution. Statistics: a Journal of Theoretical and Applied Statistics, 16(4):541–547, 1985. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Linde:1987:BR

[Lin87] W. Linde. Book review. Statistics: a Journal of Theoretical and Applied Statistics, 18(2):246, 1987. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Lin:2009:APS

[Lin09] Jin-Guan Lin. Asymptotic properties of the score test for varying dispersion in exponential family nonlinear models. Statistics: a Journal of Theoretical and Applied Statistics, 43(2):105–119, 2009. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Lippold:1975:DOT

[Lip75a] Günter Lippold. Dualitätssätze für Optimierungsprobleme in topologischen linearen Räumen. (German) [Duality sets for optimization problems in topological vector spaces]. Mathematische Operationsforschung und Statistik, 6(6):913–928, 1975. CODEN MTOSBI. ISSN 0047-6277.

Lippold:1975:ELA

[Lip75b] Günter Lippold. Über einige lineare Approximationsmethoden in der nichtlinearen Optimierung. Teil I. (German) [Some linear approximation methods in nonlinear optimization. Part I]. Mathematische Operationsforschung und Statistik, 6(1): 71–90, 1975. CODEN MTOSBI. ISSN 0047-6277.

Lippold:1976:ELA

[Lip76] Gühter Lippold. Über einige lineare Approximationsmethoden in der nichtlinearen Optimierung. Teil II. (German) [On some linear approximation methods in nonlinear optimization]. Mathematische Operationsforschung und Statistik, 7 (2):269–279, 1976. CODEN MTOSBI. ISSN 0047-6277.

Lisek:1978:CSD

[Lis78] Bernd Lisek. Comparability of special distributions. Statistics (Berlin, DDR), 9(4):587–598, 1978. CODEN MOSSD5. ISSN 0323-3944.

Lisek:1979:CSS

[Lis79] Bernd Lisek. Construction of stationary state distributions for loss systems. Statistics (Berlin, DDR), 10(4):561–581, 1979. CODEN MOSSD5. ISSN 0323-3944.

Lisek:1983:BR

[Lis83] B. Lisek. Book reviews. Statistics (Berlin, DDR), 14(1):72, 1983. CODEN MOSSD5. ISSN 0323-3944.

Lisek:1988:BR

[Lis88] Bernd Lisek. Book review. Statistics: a Journal of Theoretical and Applied Statistics, 19(2):300, 1988. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Liski:1990:CGM

[Lis90] Erkki P. Liski. Comparing generalized mixed estimators with respect to covariance matrix in a linear regression model. Statistics: a Journal of Theoretical and Applied Statistics, 21 (1):3–8, 1990. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

LitadaSilva:2014:SCS

[Lit14] João Lita da Silva. On sufficient conditions for the strong consistency of least-squares estimates. Statistics: a Journal of Theoretical and Applied Statistics, 48(3):657–667, 2014. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Liu:1992:RLD

[Liu92] Zhijun Liu. Robustness of least distances estimate in ultivariate linear models. Statistics: a Journal of Theoretical and Applied Statistics, 23(2):109–119, 1992. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Liu:2011:BQP

[Liu11] Xu-Qing Liu. Best quadratic predictions in finite populations. Statistics: a Journal of Theoretical and Applied Statistics, 45(6):525–541, 2011. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Liu:2022:EEB

[LJ22] Hanzhong Liu and Jinzhu Jia. On estimation error bounds of the Elastic Net when p ≫ n. Statistics: a Journal of Theoretical and Applied Statistics, 56(3):498–517, 2022. CO-DEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/02331888.2022.2060224.

Lee:2007:DTU

[LKK07] Sangyeol Lee, Eunhee Kim, and Youngjin Kim. Diagnostic test for unstable autoregressive models. Statistics: a Journal of Theoretical and Applied Statistics, 41(3):181–201, 2007. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Lahiri:2014:LAD

[LKM14] Ananya Lahiri, Debasis Kundu, and Amit Mitra. On least absolute deviation estimators for one-dimensional chirp model. Statistics: a Journal of Theoretical and Applied Statistics, 48(2):405–420, 2014. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Lisek:1985:NMT

[LL85] Bernd Lisek and Monika Lisek. A new method for testing whether a point process is Poisson. Statistics: a Journal of Theoretical and Applied Statistics, 16(3):445–450, 1985. CO-DEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Li:2018:SSL

[LL18] Yanfang Li and Jing Lei. Sparse subspace linear discriminant analysis. Statistics: a Journal of Theoretical and Applied Statistics, 52(4):782–800, 2018. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Lachos:2010:MME

[LLBG10] V. H. Lachos, F. V. Labra, H. Bolfarine, and Pulak Ghosh. Multivariate measurement error models based on scale mixtures of the skew-normal distribution. Statistics: a Journal of Theoretical and Applied Statistics, 44(6):541–556, 2010.

CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Levy-Leduc:2011:LSB

[LLBM^{+}11] C. Lévy-Leduc, H. Boistard, E. Moulines, M. S. Taqqu, and V. A. Reisen. Large sample behaviour of some well-known robust estimators under long-range dependence. Statistics: a Journal of Theoretical and Applied Statistics, 45(1):59–71, 2011. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Lee:2016:PCT

[LLC16] Sangyeol Lee, Youngmi Lee, and Cathy W. S. Chen. Parameter change test for zero-inflated generalized Poisson autoregressive models. Statistics: a Journal of Theoretical and Applied Statistics, 50(3):540–557, 2016. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Liero:1998:NVP

[LLK98] H. Liero, H. Läuter, and V. Konakov. Nonparametric versus parametric goodness of fit. Statistics: a Journal of Theoretical and Applied Statistics, 31(2):115–149, 1998. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Li:2017:ELG

[LLL17a] Zhouping Li, Yiming Liu, and Zhi Liu. Empirical likelihood and general relative error criterion with divergent dimension. Statistics: a Journal of Theoretical and Applied Statistics, 51 (5):1006–1022, 2017. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Lin:2017:MMR

[LLL17b] Lu Lin, Yongxin Liu, and Chen Lin. Mini-max-risk and mini-mean-risk inferences for a partially piecewise regression. Statistics: a Journal of Theoretical and Applied Statistics, 51(4):745–765, 2017. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Ling:2016:CME

[LLV16] Nengxiang Ling, Yang Liu, and Philippe Vieu. Conditional mode estimation for functional stationary ergodic data with responses missing at random. Statistics: a Journal of Theoretical and Applied Statistics, 50(5):991–??, 2016. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Larsson:2017:LRT

[LLW17] Rolf Larsson, Johan Lyhagen, and Joakim Westerlund. Likelihood ratio tests for a unit root in panels with random effects. Statistics: a Journal of Theoretical and Applied Statistics, 51(3):627–654, 2017. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Lu:2020:CCM

[LLWW20] Chao Lu, Xiaoqin Li, Rui Wang, and Xuejun Wang. Complete and complete moment convergence for randomly weighted sums of ρ∗-mixing random variables and its applications. Statistics: a Journal of Theoretical and Applied Statistics, 54(1):205–237, 2020. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Liu:2019:TNB

[LLZ19] Mengya Liu, Qi Li, and Fukang Zhu. Threshold negative binomial autoregressive model. Statistics: a Journal of Theoretical and Applied Statistics, 53(1):1–25, 2019. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Lauter:1979:STD

[LM79] Henning Läuter and Norbert Miethe. Some tests in dynamic models with a finite number of parameters. Statistics (Berlin, DDR), 10(3):395–414, 1979. CODEN MOSSD5. ISSN 0323-3944.

Lukic:2023:CBA

[LM23] Zikica Lukić and Bojana Milosević. Characterizationbased approach for construction of goodness-of-fit test for Lévy distribution. Statistics: a Journal of Theoretical and Applied Statistics, 57(5):1087–1116, 2023. CO-DEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/02331888.2023.2238236.

Lee:2010:BSK

[LMP10] Young Kyung Lee, Enno Mammen, and Byeong U. Park. Bandwidth selection for kernel regression with correlated errors. Statistics: a Journal of Theoretical and Applied Statistics, 44(4):327–340, 2010. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Louzada:2014:EEG

[LMR14] Francisco Louzada, Vitor Marchi, and Mari Roman. The exponentiated exponential-geometric distribution: a distribution with decreasing, increasing and unimodal failure rate. Statistics: a Journal of Theoretical and Applied Statistics, 48(1):167–181, 2014. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Lind:1992:AEL

[LMS92] J. C. Lind, K. L. Mehra, and J. N. Sheahan. Asymmetric errors in linear models: Estimation-theory and Monte Carlo. Statistics: a Journal of Theoretical and Applied Statistics, 23(4):305–320, 1992. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Liang:2006:SRM

[LMS06] Han-Ying Liang, Volker Mammitzsch, and Josef Steinebach. On a semiparametric regression model whose errors form a linear process with negatively associated innovations. Statistics: a Journal of Theoretical and Applied Statistics, 40(3): 207–226, 2006. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Louhichi:2017:ANE

[LMV17] Sana Louhichi, Ryozo Miura, and Dalibor Volný. On the asymptotic normality of the R-estimators of the slope parameters of simple linear regression models with associated errors. Statistics: a Journal of Theoretical and Applied Statistics, 51(1):167–187, 2017. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Louzada-Neto:2002:CTW

[LNBR02] Francisco Louzada-Neto, Heleno Bolfarine, and Josemar Rodrigues. Comparing two Weibull models with accelerated data. Statistics: a Journal of Theoretical and Applied Statistics, 36(2):175–184, 2002. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Liski:2014:SIC

[LNO14] Eero Liski, Klaus Nordhausen, and Hannu Oja. Supervised invariant coordinate selection. Statistics: a Journal of Theoretical and Applied Statistics, 48(4):711–731, 2014. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Lee:1986:HON

[LO86] P. A. Lee and S. H. Ong. Higher-order and non-stationary properties of Lampard’s stochastic reversible counter system. Statistics: a Journal of Theoretical and Applied Statistics, 17(2):261–278, 1986. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Loges:1987:NPE

[Log87] Wilfried Loges. Note on parameter estimation for general non-linear time series models. Statistics: a Journal of Theoretical and Applied Statistics, 18(4):587–590, 1987. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Lommatzschi:1970:BDP

[Lom70] Klaus Lommatzschi. Bestimmung der Dimension eines Polyeders mit Hilfe der Simplexmethode. (German) [Determining the dimension of a polyhedron with the help of the simplex method]. Mathematische Operationsforschung und Statistik, 1(3):147–153, 1970. CODEN MTOSBI. ISSN 0047-6277.

Lommatzsch:1974:BKM

[Lom74] K. Lommatzsch. Berichte über die Konferenz “Mathematische Methoden in der ökonomie” Vom 24. 9.–28.9.1973 in Zadov (CSSR). (German) [Reports on the conference “Mathematical Methods in Economics” From 24.9.–28.09.1973 in Zadov (CSSR)]. Mathematische Operationsforschung und Statistik, 5(3):553–554, 1974. CODEN MTOSBI. ISSN 0047-6277.

Lommatzsch:1982:BR

[Lom82] K. Lommatzsch. Book review. Statistics (Berlin, DDR), 13 (4):635–636, 1982. CODEN MOSSD5. ISSN 0323-3944.

Losidis:2023:RTE

[Los23] Sotirios Losidis. Recurrence times and the expected number of renewal epochs over a finite interval. Statistics: a Journal of Theoretical and Applied Statistics, 57(1):195–212, 2023.

CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/02331888.2023.2172172.

Louani:1999:SLD

[Lou99] Djamal Louani. Some large deviations limit theorems in conditional nonparametric statistics. Statistics: a Journal of Theoretical and Applied Statistics, 33(2):171–196, 1999. CO-DEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Lemdani:1997:ETF

[LP97] Mohamed Lemdani and Odile Pons. Estimation and tests in finite mixture models for censored survival data. Statistics: a Journal of Theoretical and Applied Statistics, 29(4):363–388, 1997. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Lin:1997:AIV

[LPC97] Jyh-Jiuan Lin, Nabendu Pal, and Ching-Hui Chang. Applications of improved variance estimators in a multivariate normal mean vector estimation. Statistics: a Journal of Theoretical and Applied Statistics, 30(2):99–125, 1997. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Lapeybe:1990:SLD

[LPS90] Bernard Lapeybe, Gxlles Pages, and Kabam Sab. Sequences with low discrepancy generalisation and application to Robbins–Monro algorithm. Statistics: a Journal of Theoretical and Applied Statistics, 21(2):251–272, 1990. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Lin:2012:CTH

[LQ12] Jin-Guan Lin and Xiao-Yi Qu. A consistent test for heteroscedasticity in semi-parametric regression with nonparametric variance function based on the kernel method. Statistics: a Journal of Theoretical and Applied Statistics, 46(5): 565–576, 2012. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Lin:2017:DRE

[LQZF17] Huiming Lin, Guoyou Qin, Jiajia Zhang, and Wing K. Fung. Doubly robust estimation of partially linear models for longitudinal data with dropouts and measurement error in covariates. Statistics: a Journal of Theoretical and Applied Statistics, 52(1):84–98, 2017. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Lee:2017:NGQ

[LR17] Junbum Lee and Suhasini Subba Rao. A note on general quadratic forms of nonstationary stochastic processes. Statistics: a Journal of Theoretical and Applied Statistics, 51(5): 949–968, 2017. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Lauter:1973:RGB

[LSP^{+}73] H. Läuter, E. Seiffart, András Prékopa, E. Rödel, E. Läuter, and E. E. Borngräber. Rezension. (German) [Book review]. Mathematische Operationsforschung und Statistik, 4(2):173– 176, 1973. CODEN MTOSBI. ISSN 0047-6277.

Land:2012:ERS

[LSS12a] Margaret Land, Sarjinder Singh, and Stephen A. Sedory. Estimation of a rare sensitive attribute using Poisson distribution. Statistics: a Journal of Theoretical and Applied Statistics, 46(3):351–360, 2012. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Li:2012:DBR

[LSS12b] Wenjun Li, Edward J. Stanek III, and Julio M. Singer. Designbased random permutation models with auxiliary information. Statistics: a Journal of Theoretical and Applied Statistics, 46(5):663–671, 2012. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Lu:2018:CBT

[LST18] Changli Lu, Yuqin Sun, and Yongge Tian. A comparison between two competing fixed parameter constrained general linear models with new regressors. Statistics: a Journal of Theoretical and Applied Statistics, 52(4):769–781, 2018. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Lauteb:1985:OCC

[LT85] H. Lauteb and H. Thiele. Optimal combinations of classification procedures. Statistics: a Journal of Theoretical and Applied Statistics, 16(3):389–406, 1985. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

LitadaSilva:2013:SCL

[LT13] João Lita da Silva and João Tiago Mexia. Strong consistency of least squares estimates with i.i.d. errors with mean values not necessarily defined. Statistics: a Journal of Theoretical and Applied Statistics, 47(4):707–714, 2013. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Ling:2016:MCP

[LT16] Chengxiu Ling and Zhongquan Tan. On maxima of chiprocesses over threshold dependent grids. Statistics: a Journal of Theoretical and Applied Statistics, 50(3):579–595, 2016. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Liski:1997:ETD

[LTG97] Erkki P. Liski, Götz Trenkler, and Jürgen Groß. Estimation from transformed data under the linear regression model. Statistics: a Journal of Theoretical and Applied Statistics, 29(3):205–219, 1997. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Larocque:2003:AIR

[LTV03] Denis Larocque, Serge Tardif, and Constance Van Eeden. Affine-invariant rank tests for the bivariate one-way layout and for simple regression models. Statistics: a Journal of Theoretical and Applied Statistics, 37(1):33–50, 2003. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Luceno:1999:ICC

[Luc99] Alberto Luceño. Intra-cluster correlation in the normal model. Statistics: a Journal of Theoretical and Applied Statistics, 33(2):119–128, 1999. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Luceno:2005:RCL

[Luc05] Alberto Luceño. Recursive characterization of a large family of discrete probability distributions showing extra-Poisson variation. Statistics: a Journal of Theoretical and Applied Statistics, 39(3):261–267, 2005. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Luderber:1973:LLO

[Lud73] Bernd Luderber. Lösung einer linearen Optimierungsaufgabe mit Umrandungsstruktur durch doppelte Dekomposition. (German) [Solution of a linear optimization task with fringe structure by double decomposition]. Mathematische Operationsforschung und Statistik, 4(4):251–274, 1973. CO-DEN MTOSBI. ISSN 0047-6277.

Lee:2013:ERS

[LUK13] Gi-Sung Lee, Daiho Uhm, and Jong-Min Kim. Estimation of a rare sensitive attribute in a stratified sample using Poisson distribution. Statistics: a Journal of Theoretical and Applied Statistics, 47(3):575–589, 2013. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Lee:2014:ERS

[LUK14] Gi-Sung Lee, Daiho Uhm, and Jong-Min Kim. Estimation of a rare sensitive attribute in probability proportional to size measures using Poisson distribution. Statistics: a Journal of Theoretical and Applied Statistics, 48(3):685–709, 2014. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Luschgy:1987:ERH

[Lus87] Harald Luschgy. Elimination of randomization and Hunt– Stein type theorems in invariant statistical decision problems. Statistics: a Journal of Theoretical and Applied Statistics, 18 (1):99–111, 1987. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Ling:2018:NMF

[LV18] Nengxiang Ling and Philippe Vieu. Nonparametric modelling for functional data: selected survey and tracks for future. Statistics: a Journal of Theoretical and Applied Statistics, 52(4):934–949, 2018. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Ling:2012:CMK

[LW12] Nengxiang Ling and Yuehua Wu. Consistency of modified kernel regression estimation for functional data. Statistics: a Journal of Theoretical and Applied Statistics, 46(2):149–158, 2012. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Li:2018:SID

[LW18] Zhouping Li and Yang Wei. Statistical inference for the difference of two Lorenz curves. Statistics: a Journal of Theoretical and Applied Statistics, 52(5):1128–1155, 2018. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Liu:2018:TCI

[LWL18] Yi Liu, Qihua Wang, and Xiaohui Liu. Testing conditional independence via integrating-up transform. Statistics: a Journal of Theoretical and Applied Statistics, 52(4):734–749, 2018. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Liang:2021:QRP

[LWS21] Han-Ying Liang, Bao-Hua Wang, and Yu Shen. Quantile regression of partially linear single-index model with missing observations. Statistics: a Journal of Theoretical and Applied Statistics, 55(1):1–17, 2021. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Lin:1993:NSC

[LY93] Fu Yong Lin and Tang Shang Yong. Necessary and sufficient conditions that linear estimators of a mixed effects linear model are admissible under matrix loss function. Statistics: a Journal of Theoretical and Applied Statistics, 24(4):303–309, 1993. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Li:2011:NRT

[LY11] Yalian Li and Hu Yang. A new ridge-type estimator in stochastic restricted linear regression. Statistics: a Journal of Theoretical and Applied Statistics, 45(2):123–130, 2011. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Liu:2016:WQR

[LY16] Tianqing Liu and Xiaohui Yuan. Weighted quantile regression with missing covariates using empirical likelihood. Statistics: a Journal of Theoretical and Applied Statistics, 50(1):89–113, 2016. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910

(electronic). URL http://www.tandfonline.com/doi/abs/
10.1080/02331888.2015.1033164.

Liu:2020:AEL

[LY20] Tianqing Liu and Xiaohui Yuan. Adaptive empirical likelihood estimation with nonignorable nonresponse data. Statistics: a Journal of Theoretical and Applied Statistics, 54(1): 1–22, 2020. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Lu:2019:MTH

[LYB19] Wanbo Lu, Dong Yang, and Kris Boudt. A misspecification test for the higher order co-moments of the factor model. Statistics: a Journal of Theoretical and Applied Statistics, 53(3):471–488, 2019. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Lyska:1976:VEP

[Lys76] Werner Lyska. Ein Verfahren zum Erzeugen von Partialsummen nach Wachsenden Summenwerten. (German) [A method for generating partial sums of growing sum values]. Mathematische Operationsforschung und Statistik, 7(3):371–380, 1976. CODEN MTOSBI. ISSN 0047-6277.

Lyttkens:1977:SHI

[Lyt77] Ejnar Lyttkens. Schätzung mit Hilfe von Instrumentalvariablen. (German) [Estimation using instrumental variables]. Statistics (Berlin, DDR), 8(2):131–172, 1977. CODEN MOSSD5. ISSN 0323-3944.

Liu:2018:EHP

[LZW18] Guangying Liu, Lixin Zhang, and Min Wang. Estimation of the Hurst parameter in the simultaneous presence of jumps and noise. Statistics: a Journal of Theoretical and Applied Statistics, 52(5):1156–1192, 2018. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Lu:2023:BET

[LZW23] Chao Lu, Houlin Zhou, and Xuejun Wang. A Berry– Esseen theorem for sample quantiles under martingale difference sequences. Statistics: a Journal of Theoretical and Applied Statistics, 57(4):844–866, 2023. CO-DEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/02331888.2023.2225668.

Mostafa:1986:EBQ

[MA86] S. M. Mostafa and R. Ahmad. Empirical Bayes quadratic estimators of variance components in normal linear models. Statistics: a Journal of Theoretical and Applied Statistics, 17(3):337–348, 1986. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Mansourvar:2020:ECC

[MA20] Zahra Mansourvar and Majid Asadi. An extension of the Cox– Czanner divergence measure to residual lifetime distributions with applications. Statistics: a Journal of Theoretical and Applied Statistics, 54(6):1311–1328, 2020. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Madan:1976:ISQ

[Mad76] K. C. Madan. Interrupted service queueing with arrivals and departures in batches of variable size and general repair time distribution. Mathematische Operationsforschung und Statistik, 7(1):139–149, 1976. CODEN MTOSBI. ISSN 0047-6277.

Maggu:1972:CCS

[Mag72] Prakash Lal Maggu. On certain channels in series with phase type service. Mathematische Operationsforschung und Statistik, 3(5):355–361, 1972. CODEN MTOSBI. ISSN 0047-6277.

Magiera:1988:SEG

[Mag88] R. Magiera. Sequential estimation for the generalized exponential hyperbolic secant process. Statistics: a Journal of Theoretical and Applied Statistics, 19(2):271–281, 1988. CO-DEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Magiera:1995:BSE

[Mag95] Ryszard Magiera. Bayes sequential estimation procedures in exponential-type processes for a polynomial cost function. Statistics: a Journal of Theoretical and Applied Statistics, 26 (1):61–73, 1995. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Mai:1991:BRP

[Mai91] K. Mai. Book review: P. Sprent, Applied Nonparametric Statistical Methods. Statistics: a Journal of Theoretical and Applied Statistics, 22(1):2, 1991. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Makambi:2006:STP

[Mak06] Kepher H. Makambi. Supplementary test procedures for overall treatment effect. Statistics: a Journal of Theoretical and Applied Statistics, 40(4):289–296, 2006. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Maller:1989:RAE

[Mal89] R. A. Maller. Regression with autoregressive errors-some asymptotic results. Statistics: a Journal of Theoretical and Applied Statistics, 20(1):23–39, 1989. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Malov:2021:UCR

[Mal21] Sergey V. Malov. Uniform convergence rate of the nonparametric maximum likelihood estimator for current status data with competing risks. Statistics: a Journal of Theoretical and Applied Statistics, 55(1):152–172, 2021. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Mammen:1995:QSK

[Mam95] E. Mammen. On qualitative smoothness of kernel density estimates^{1}. Statistics: a Journal of Theoretical and Applied Statistics, 26(3):253–267, 1995. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Mandl:1972:BVP

[Man72] P. Mandl. Bericht über die VI. Prague Conference on Information Theory, Statistical Decision Functions and Random Processes. (German) [Report on the VI. Prague Conference on Information Theory, Statistical Decision Functions and Random Processes]. Mathematische Operationsforschung und Statistik, 3(1):74, 1972. CODEN MTOSBI. ISSN 0047-6277.

Mansourvar:2022:DFM

[Man22a] Zahra Mansourvar. The φ-divergence family of measures based on quantile function. Statistics: a Journal of Theoretical and Applied Statistics, 56(5):1113–1132, 2022. CO-DEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/02331888.2022.2105844.

Mansourvar:2022:DMD

[Man22b] Zahra Mansourvar. A dynamic measure of divergence between two inactivity lifetime distributions. Statistics: a Journal of Theoretical and Applied Statistics, 56(1):147–163, 2022. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Marianna:1997:GAN

[Mar97] Pensky Marianna. A general approach to nonparametric empirical Bayes estimation. Statistics: a Journal of Theoretical and Applied Statistics, 29(1):61–80, 1997. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Marsh:2019:RIN

[Mar19] Patrick Marsh. The role of information in nonstationary regression. Statistics: a Journal of Theoretical and Applied Statistics, 53(3):656–672, 2019. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Mousa:2006:SIR

[MAS06] M. A. M. Ali Mousa and S. A. Al-Sagheer. Statistical inference for the Rayleigh model based on progressively Type-II censored data. Statistics: a Journal of Theoretical and Applied Statistics, 40(2):149–157, 2006. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Massing:2019:LAN

[Mas19] Till Massing. Local asymptotic normality for Student–Lévy processes under high-frequency sampling. Statistics: a Journal of Theoretical and Applied Statistics, 53(4):721–752, 2019. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Matthes:1975:SOS

[Mat75] Bernd Matthes. Stabilitätsanalyse ökonomischer Systeme. (German) [Stability analysis of economic systems]. Mathematische Operationsforschung und Statistik, 6(1):141–153, 1975. CODEN MTOSBI. ISSN 0047-6277.

Mathieu:1981:TSM

[Mat81a] J. R. Mathieu. Tests of χ^{2} in a stationary Markov process and separability of hypotheses. Statistics (Berlin, DDR), 12 (4):529–550, 1981. CODEN MOSSD5. ISSN 0323-3944.

Mathieu:1981:TGL

[Mat81b] J. R. Mathieu. Tests of χ^{2} in the generalized linear model. Statistics (Berlin, DDR), 12(4):509–527, 1981. CODEN MOSSD5. ISSN 0323-3944.

Mathew:1983:NBL

[Mat83] Thomas Mathew. A note on best linear unbiased estimation in the restricted general linear model. Statistics (Berlin, DDR), 14(1):3–6, 1983. CODEN MOSSD5. ISSN 0323-3944.

Mathae:1985:OPO

[Mat85] Rudolf Mathae. Outlier-prone and outlier-resistant multidimensional distributions. Statistics: a Journal of Theoretical and Applied Statistics, 16(3):451–456, 1985. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Mathew:1987:CNE

[Mat87a] Thomas Mathew. On the characterization of nonegatively estimable linear combinations of variance components. Statistics: a Journal of Theoretical and Applied Statistics, 18(2): 227–236, 1987. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Mathieu:1987:SLL

[Mat87b] Jean-René Mathieu. On some linear and log-linear problems about the marginal distributions of two-and threedimensional contingency tables. Statistics: a Journal of Theoretical and Applied Statistics, 18(1):49–63, 1987. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Mathar:1989:OBD

[Mat89] Rudolf Mathar. The outlier behaviour of distributions and the decay of power of optimal tests at decreasing level. Statistics: a Journal of Theoretical and Applied Statistics, 20(2):247– 254, 1989. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Mauermann:1989:KSK

[Mau89] Werner Mauermann. Konstruktion eines sequentiellen Konfidenzschätzverfahrens mit vorgegebener Intervallänge für den Erwartungswert exponentialverteilter Stichprobenwerte. (German) [Construction of a sequential estimation method with a given confidence interval length for the expected value of exponentially distributed sample values]. Statistics: a Journal of Theoretical and Applied Statistics, 20(3):439–448, 1989. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Mayo:2011:EDM

[May11] Nathanaël Mayo. Estimation of dynamic models on the factors of marginal principal component analysis. Statistics: a Journal of Theoretical and Applied Statistics, 45(1):101–120, 2011. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Mukhopadhyay:1991:CBM

[MB91] P. Mukhopadhyay and S. Bhattacharyya. On a comparison between Midzuno strategy and PPSWR strategy. Statistics: a Journal of Theoretical and Applied Statistics, 22(2):291–294, 1991. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic). See addendum [MB92].

Mukhopadhyay:1992:APC

[MB92] P. Mukhopadhyay and S. Bhattacharyya. Addendum to the paper, On a comparison between Midzuno strategy and PP-SWR strategy. Statistics: a Journal of Theoretical and Applied Statistics, 23(2):185, 1992. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic). See [MB91].

Maiti:2018:TSA

[MB18] Raju Maiti and Atanu Biswas. Time series analysis of categorical data using auto-odds ratio function. Statistics: a Journal of Theoretical and Applied Statistics, 52(2):426–444, 2018. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Mukhopadhyay:2021:BMC

[MB21] Sabyasachi Mukhopadhyay and Sourabh Bhattacharya. Bayesian MISE convergence rates of Polya urn based density estimators: asymptotic comparisons and choice of prior parameters. Statistics: a Journal of Theoretical and Applied Statistics, 55(1):120–151, 2021. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Messer:2020:MSA

[MBF^{+}20] Michael Messer, Hendrik Backhaus, Ting Fu, Albrecht Stroh, and Gaby Schneider. A multi-scale approach for testing and detecting peaks in time series. Statistics: a Journal of Theoretical and Applied Statistics, 54(5):1058–1080, 2020. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Mond:1975:CNC

[MC75] B. Mond and B. D. Craven. A class of nondifferentiable complex programming problems. Mathematische Operationsforschung und Statistik, 6(4):581–591, 1975. CODEN MTOSBI. ISSN 0047-6277.

McCartin:2003:GCL

[McC03] Brian J. McCartin. A geometric characterization of linear regression. Statistics: a Journal of Theoretical and Applied Statistics, 37(2):101–117, 2003. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

McCartin:2005:GLR

[McC05] Brian J. McCartin. The geometry of linear regression with correlated errors. Statistics: a Journal of Theoretical and Applied Statistics, 39(1):1–11, 2005. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

McCartin:2006:GCP

[McC06] Brian J. McCartin. Geometric characterization of planar regression. Statistics: a Journal of Theoretical and Applied Statistics, 40(3):187–206, 2006. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Matos:2018:MME

[MCCL18] Larissa A. Matos, Luis M. Castro, Celso R. B. Cabral, and Víctor H. Lachos. Multivariate measurement error models based on Student-t distribution under censored responses. Statistics: a Journal of Theoretical and Applied Statistics, 52(6):1395–1416, 2018. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Massuia:2015:IDS

[MCML15] Monique B. Massuia, Celso Rômulo Barbosa Cabral, Larissa A. Matos, and Víctor H. Lachos. Influence diagnostics for Student-t censored linear regression models. Statistics: a Journal of Theoretical and Applied Statistics, 49(5): 1074–1094, 2015. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Matilainen:2019:SAV

[MCNO19] M. Matilainen, C. Croux, K. Nordhausen, and H. Oja. Sliced average variance estimation for multivariate time series. Statistics: a Journal of Theoretical and Applied Statistics, 53(3):630–655, 2019. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Marin:2003:OQU

[MD03] Jean-Michel Marin and Thierry Dhorne. Optimal quadratic unbiased estimation for models with linear Toeplitz covariance structure. Statistics: a Journal of Theoretical and Applied Statistics, 37(2):85–99, 2003. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Mathew:2022:SSR

[MD22] Angel Mathew and K. R. Deepa. Stress-strength reliability: a quantile approach. Statistics: a Journal of Theoretical and Applied Statistics, 56(1):206–221, 2022. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Mecke:1972:EMA

[Mec72] J. Mecke. Das erlangsche Model mit abhängigen Bedienungszeiten. (German) [The Erlang model with dependent service times]. Mathematische Operationsforschung und Statistik, 3(6):453–464, 1972. CODEN MTOSBI. ISSN 0047-6277.

Mecke:1981:FSP

[Mec81] J. Mecke. Formulas for stationary planar fibre processes III — intersections with fibre systems. Statistics (Berlin, DDR), 12(2):201–210, 1981. CODEN MOSSD5. ISSN 0323-3944.

Mecke:1984:PRM

[Mec84] J. Mecke. Parametric representation of mean values for stationary random mosaics. Statistics (Berlin, DDR), 15(3):437– 442, 1984. CODEN MOSSD5. ISSN 0323-3944.

Mecke:1988:IIR

[Mec88] J. Mecke. An isoperimetrlc inequality for random quadrangle tessellations. Statistics: a Journal of Theoretical and Applied Statistics, 19(1):57–65, 1988. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Meister:2006:SEM

[Mei06] Alexander Meister. Support estimation via moment estimation in presence of noise. Statistics: a Journal of Theoretical and Applied Statistics, 40(3):259–275, 2006. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Mejza:1978:UIB

[Mej78] S. Mejza. Use of inter-block information to obtain uniformly better estimators of treatment contrasts. Statistics (Berlin, DDR), 9(3):335–341, 1978. CODEN MOSSD5. ISSN 0323-3944.

Melzer:1976:DML

[Mel76] Detlef Melzer. Eine duale Methode zur Lösung konvexer Optimierungsprobleme. (German) [A dual method for solving convex optimization problems]. Mathematische Operationsforschung und Statistik, 7(6):871–878, 1976. CODEN MTOSBI. ISSN 0047-6277.

Melas:1978:ODE

[Mel78] V. B. Melas. Optimal designs for exponential regression. Statistics (Berlin, DDR), 9(1):45–59, 1978. CODEN MOSSD5. ISSN 0323-3944.

Melas:1994:BTM

[Mel94] V. B. Melas. Branching technique for Markov chain simulation (finite state case). Statistics: a Journal of Theoretical and Applied Statistics, 25(2):159–171, 1994. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Mensch:1971:IP

[Men71] Gerhard Mensch. The inspection problem. Mathematische Operationsforschung und Statistik, 2(4):261–269, 1971. CO-DEN MTOSBI. ISSN 0047-6277.

Mensch:1973:SSL

[Men73] Gerhard Mensch. Single-stage linear programming zero-one solutions to some job-machine scheduling type problems. Mathematische Operationsforschung und Statistik, 4(1):23– 33, 1973. CODEN MTOSBI. ISSN 0047-6277.

Mihalevic:1970:GSL

[MESS70] V. S. Mihalevič, Ju. M. Ermolev, V. V. Skurba, and N. Z. Sor. Große Systeme und die Lösung von Extremalproblemen. (German) [Large systems and the solution of extremal problems]. Mathematische Operationsforschung und Statistik, 1 (4):309–326, 1970. CODEN MTOSBI. ISSN 0047-6277.

Morais:2017:CRM

[MF17] Alice L. Morais and Silvia L. P. Ferrari. A class of regression models for parallel and series systems with a random number of components. Statistics: a Journal of Theoretical and Applied Statistics, 51(2):294–313, 2017. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Martinez-Florez:2014:APE

[MFBG14a] Guillermo Martínez-Flórez, Heleno Bolfarine, and Héctor W. Gómez. An alpha-power extension for the Birnbaum– Saunders distribution. Statistics: a Journal of Theoretical and Applied Statistics, 48(4):896–912, 2014. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Martinez-Florez:2014:SNA

[MFBG14b] Guillermo Martínez-Flórez, Heleno Bolfarine, and Héctor W. Gómez. Skew-normal alpha-power model. Statistics: a Journal of Theoretical and Applied Statistics, 48(6):1414–1428, 2014. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic). See [SGMFB18].

Martinez-Florez:2022:SNF

[MFTFEO22] Guillermo Martínez-Flórez, Roger Tovar-Falón, and David Elal-Olivero. Some new flexible classes of normal distribution for fitting multimodal data. Statistics: a Journal of Theoretical and Applied Statistics, 56(1):182–205, 2022. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Muller-Gronbach:1996:ODA

[MG96] Thomas Müller-Gronbach. Optimal designs for approximating a stochastic process with respect to a minimax criterion. Statistics: a Journal of Theoretical and Applied Statistics, 27 (3–4):279–296, 1996. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Miethe:1978:BR

[MGLG78] N. Miethe, H. Gillert, H. Läuter, and H.-J. Girlich. Book review. Statistics (Berlin, DDR), 9(4):643–644, 1978. CODEN MOSSD5. ISSN 0323-3944.

Majumder:2020:ORE

[MGM20] Priyanka Majumder, Shyamal Ghosh, and Murari Mitra. Ordering results of extreme order statistics from heterogeneous Gompertz–Makeham random variables. Statistics: a Journal of Theoretical and Applied Statistics, 54(3):595–617, 2020. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Mudholkara:1997:IBP

[MH97] Govind S. Mudholkara and Alan D. Hutson. Improvements in the bias and precision of sample quartiles. Statistics: a Journal of Theoretical and Applied Statistics, 30(3):239–257, 1997. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Mohammadi:2022:IWC

[MH22] Morteza Mohammadi and Majid Hashempour. On interval weighted cumulative residual and past extropies. Statistics: a Journal of Theoretical and Applied Statistics, 56(5):1029– 1047, 2022. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-

4910 (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/02331888.2022.2111429.

Mi:1996:SCE

[Mi96] Jie Mi. Strongly consistent estimation for hazard rate models with a change point. Statistics: a Journal of Theoretical and Applied Statistics, 28(1):35–42, 1996. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Michael:1986:CDT

[Mic86] Sobekseh Michael. Classes of diffusion-type processes with a sufficient reduction. Statistics: a Journal of Theoretical and Applied Statistics, 17(4):585–596, 1986. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Michaletzky:1987:MLE

[Mic87] György Michaletzky. On the maximum likelihood estimation for certain MANOVA models. Statistics: a Journal of Theoretical and Applied Statistics, 18(2):237–245, 1987. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Michel:2009:EAD

[Mic09] René Michel. Estimation of the angular density in bivariate generalized Pareto models. Statistics: a Journal of Theoretical and Applied Statistics, 43(2):187–202, 2009. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Miethe:1979:APT

[Mie79] N. Miethe. Asymptotic properties of tests in autoregressive moving average models. Statistics (Berlin, DDR), 10(2):307– 318, 1979. CODEN MOSSD5. ISSN 0323-3944.

Mielniczuk:1991:SAP

[Mie91] Jan Mielniczuk. Some asymptotic properties of nonparametric regression estimators in case of randomly censored data. Statistics: a Journal of Theoretical and Applied Statistics, 22 (1):85–93, 1991. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Mikosch:1988:ILR

[Mik88] T. Mikosch. Iterated logarithm results for rapidly growing random walk. Statistics: a Journal of Theoretical and Applied Statistics, 19(1):107–115, 1988. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Milhaud:1986:AUM

[Mil86] Xavier Milhaud. Asymptotically uniformly most powerful tests for independent or Markovian observations. Statistics: a Journal of Theoretical and Applied Statistics, 17(1):41–47, 1986. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Mizera:1996:LBM

[Miz96] Beata Mizera. Lower bounds on the minimax risk of sequential estimators. Statistics: a Journal of Theoretical and Applied Statistics, 28(2):123–129, 1996. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Miziula:2017:CDM

[Miz17] P. Miziula. Comparison of dispersions of mixtures of ordered distributions. Statistics: a Journal of Theoretical and Applied Statistics, 51(4):862–877, 2017. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Mousa:2002:BEP

[MJA02] M. A. M. Ali Mousa, Z. F. Jaheen, and A. A. Ahmad. Bayesian estimation, prediction and characterization for the Gumbel model based on records. Statistics: a Journal of Theoretical and Applied Statistics, 36(1):65–74, 2002. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Moyssiadis:1983:EOO

[MK83] Chronis Moyssiadis and Stratis Kounias. Exact d-optimal n observations 2^{k} designs of resolution III, when n[diff between]1or2 mod 4. Statistics (Berlin, DDR), 14(3):367–379, 1983. CODEN MOSSD5. ISSN 0323-3944.

Misra:1995:SBE

[MK95] Neeraj Misra and Narinder Kumar. Selecting the best exponential population when the scale parameters are bounded. Statistics: a Journal of Theoretical and Applied Statistics, 27 (1–2):151–158, 1995. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Mondal:2019:EIM

[MK19] Shuvashree Mondal and Debasis Kundu. Exact inference on multiple exponential populations under a joint type-II progressive censoring scheme. Statistics: a Journal of Theoretical and Applied Statistics, 53(6):1329–1356, 2019. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Mukherjee:2021:TND

[MKM21] Amitava Mukherjee, Wolfgang Kössler, and Hidetoshi Murakami. Two new distribution-free two-sample tests for versatile alternative. Statistics: a Journal of Theoretical and Applied Statistics, 55(5):1123–1153, 2021. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Morvai:1999:REI

[MKN99] Gusztáv Morvai, Sanjeev R. Kulkarni, and Andrew B. Nobel. Regression estimation from an individual stable sequence. Statistics: a Journal of Theoretical and Applied Statistics, 33 (2):99–118, 1999. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Ma:2021:UME

[ML21] Xiaoyu Ma and Lu Lin. Univariate measurement error selection likelihood for variable selection of additive model. Statistics: a Journal of Theoretical and Applied Statistics, 55(4): 875–893, 2021. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Moriyama:2018:SAT

[MM18] Taku Moriyama and Yoshihiko Maesono. Smoothed alternatives of the two-sample median and Wilcoxon’s rank sum tests. Statistics: a Journal of Theoretical and Applied Statistics, 52(5):1096–1115, 2018. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Masihuddin:2021:EEF

[MM21] Masihuddin and Neeraj Misra. Equivariant estimation following selection from two normal populations having common unknown variance. Statistics: a Journal of Theoretical and Applied Statistics, 55(6):1407–1438, 2021. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

MacEachern:2022:RAT

[MM22] Steven N. MacEachern and Koji Miyawaki. A regression approach to the two-dataset problem. Statistics: a Journal of Theoretical and Applied Statistics, 56(6):1225–1241, 2022. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/02331888.2022.2134385.

Martinez:2009:APN

[MMdP09] R. Martínez, M. Mota, and I. del Puerto. On asymptotic posterior normality for controlled branching processes. Statistics: a Journal of Theoretical and Applied Statistics, 43(4):

367–378, 2009. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Marhuenda:2005:CUT

[MMP05] Y. Marhuenda, D. Morales, and M. C. Pardo. A comparison of uniformity tests. Statistics: a Journal of Theoretical and Applied Statistics, 39(4):315–327, 2005. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Marhuenda:2016:TVP

[MMP16] Y. Marhuenda, D. Morales, and M. C. Pardo. Tests for the variance parameter in the Fay–Herriot model. Statistics: a Journal of Theoretical and Applied Statistics, 50(1):27–42, 2016. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910

(electronic). URL http://www.tandfonline.com/doi/abs/
10.1080/02331888.2015.1016026.

Mojirsheibani:2011:CIC

[MMR11] Majid Mojirsheibani, Zahra Montazeri, and Abdolreza Rajaeefard. On classification with incomplete covariates. Statistics: a Journal of Theoretical and Applied Statistics, 45(5): 427–450, 2011. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Martinez-Miranda:2007:LOA

[MMRA07] M. D. Martínez-Miranda, M. Rueda, and A. Arcos. Looking for optimal auxiliary variables in sample survey quantile estimation. Statistics: a Journal of Theoretical and Applied Statistics, 41(3):241–252, 2007. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Mammen:2008:STP

[MN08] Enno Mammen and Swagata Nandi. Some theoretical properties of phase-randomized multivariate surrogates. Statistics: a Journal of Theoretical and Applied Statistics, 42(3):195–205, 2008. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Mathew:2010:NNE

[MNvRS10] Thomas Mathew, Tatjana Nahtman, Dietrich von Rosen, and Bimal Kumar Sinha. Non-negative estimation of variance components in heteroscedastic one-way random-effects ANOVA models. Statistics: a Journal of Theoretical and Applied Statistics, 44(6):557–569, 2010. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Mahmoud:1973:ADS

[MO73] M. W. Mahmoud and A. Oada. The approximate distributions for the sum of equi-correlated lognormal variates. Mathematische Operationsforschung und Statistik, 4(3):229–242, 1973. CODEN MTOSBI. ISSN 0047-6277.

Muller:2003:ANG

[MO03] Ursula U. Müller and Gerhard Osius. Asymptotic normality of goodness-of-fit statistics for sparse Poisson data. Statistics: a Journal of Theoretical and Applied Statistics, 37(2):119–143, 2003. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Milosevic:2019:CES

[MO19] Bojana Milosević and Marko Obradović. Comparison of efficiencies of some symmetry tests around an unknown centre. Statistics: a Journal of Theoretical and Applied Statistics, 53 (1):43–57, 2019. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Moeschlin:1984:NPC

[Moe84] O. Moeschlin. A note on Pratt’s concept of minimal expected length of confidence intervals. Statistics (Berlin, DDR), 15 (4):483–486, 1984. CODEN MOSSD5. ISSN 0323-3944.

Mohner:1986:CSE

[Moh86] M. Mohner. A comparative study of estimators for probabilities in contingency tables. Statistics: a Journal of Theoretical and Applied Statistics, 17(4):557–568, 1986. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Mohner:1987:BR

[Möh87a] M. Möhner. Book review. Statistics: a Journal of Theoretical and Applied Statistics, 18(1):64, 1987. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Mohner:1987:EEA

[Möh87b] M. Möhner. On the estimation of the expected actual error rate in sample-based multinomial classification. Statistics: a Journal of Theoretical and Applied Statistics, 18(4):599–612, 1987. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Mohner:1989:BR

[Moh89] M. Mohner. Book reviews. Statistics: a Journal of Theoretical and Applied Statistics, 20(4):572, 1989. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Mojirsheibani:2007:SAS

[Moj07] Majid Mojirsheibani. Some approximations to L_{p}-statistics of kernel density estimators. Statistics: a Journal of Theoretical and Applied Statistics, 41(3):203–220, 2007. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Molchanov:1996:SVE

[Mol96] I. S. Molchanov. Set-valued estimators for mean bodies related to Boolean models. Statistics: a Journal of Theoretical and Applied Statistics, 28(1):43–56, 1996. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Molchanov:1997:BR

[Mol97] Ilya Molchanov. Book reviews. Statistics: a Journal of Theoretical and Applied Statistics, 29(2):203–204, 1997. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Monch:1972:BTL

[Mön72] G. Mönch. Bericht über die Tagung “Lang-und mittelfristige Planungsmodelle”. (German) [Report of the meeting “Long- and medium-term planning models”]. Mathematische Operationsforschung und Statistik, 3(1):79, 1972. CODEN MTOSBI. ISSN 0047-6277.

Monod:1996:VRM

[Mon96] Hervé Monod. Valid randomization for models with neighbour effects from treatments. Statistics: a Journal of Theoretical and Applied Statistics, 27(3–4):311–324, 1996. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Moszynska:1985:PSS

[Mos85] M. Moszynska. Products of statistical spaces. Statistics: a Journal of Theoretical and Applied Statistics, 16(3):407–418, 1985. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Mukerjee:1986:OCS

[MOS86] Rahul Mukerjee, Kashinath Ohattebjee, and Mausumi Sen. D-optimality of a class of saturated main-effect plans and allied results. Statistics: a Journal of Theoretical and Applied Statistics, 17(3):349–355, 1986. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Mousa:2001:IPB

[Mou01] M. A. M. Ali Mousa. Inference and prediction for the Burr Type X model based on records. Statistics: a Journal of Theoretical and Applied Statistics, 35(4):415–425, 2001. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Mourid:2002:EPF

[Mou02] Tahar Mourid. Estimation and prediction of functional autoregressive processes. Statistics: a Journal of Theoretical and Applied Statistics, 36(2):125–138, 2002. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Mousa:2003:BPB

[Mou03] M. A. M. Ali Mousa. Bayesian prediction based on Pareto doubly censored data. Statistics: a Journal of Theoretical and Applied Statistics, 37(1):65–72, 2003. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Moszynska:1982:ERS

[MP82] M. Moszyńska and E. Pieszozyńska. Equivalence relations for statistical spaces. Statistics (Berlin, DDR), 13(2):193–218, 1982. CODEN MOSSD5. ISSN 0323-3944.

Mammen:1996:BKD

[MP96] Enno Mammen and Byeong U. Park. Behaviour of kernel density estimates and bandwidth selectors for contaminated data sets. Statistics: a Journal of Theoretical and Applied Statistics, 28(2):89–104, 1996. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Macci:2010:LDR

[MP10] Claudio Macci and Lea Petrella. Large deviation results on some estimators for stationary Gaussian processes. Statistics: a Journal of Theoretical and Applied Statistics, 44(2):129– 144, 2010. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Martin:2011:FDS

[MP11] Nirian Martín and Leandro Pardo. Fitting DNA sequences through log-linear modelling with linear constraints. Statistics: a Journal of Theoretical and Applied Statistics, 45(6): 605–621, 2011. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Miguel:2022:MDS

[MP22] Sergio Brenner Miguel and Nathawut Phandoidaen. Multiplicative deconvolution in survival analysis under dependency. Statistics: a Journal of Theoretical and Applied Statistics, 56(2):297–328, 2022. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Melnikov:2023:PEO

[MP23] Alexander Melnikov and Andrey Pak. Parameter estimation in optional semimartingale regression models. Statistics: a Journal of Theoretical and Applied Statistics, 57(5):1165– 1201, 2023. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-

4910 (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/02331888.2023.2242549.

Morales:2004:RST

[MPPV04] D. Morales, L. Pardo, M. C. Pardo, and I. Vajda. Rényi statistics for testing composite hypotheses in general exponential models. Statistics: a Journal of Theoretical and Applied Statistics, 38(2):133–147, 2004. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Menendez:2005:PTC

[MPPZ05] M. L. Menéndez, J. A. Pardo, L. Pardo, and K. Zografos. A preliminary test in classification and probabilities of misclassification. Statistics: a Journal of Theoretical and Applied Statistics, 39(3):183–205, 2005. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Morales:2000:RSD

[MPV00] D. Morales, L. Pardo, and I. Vajda. Rényi statistics in directed families of exponential experiments. Statistics: a Journal of Theoretical and Applied Statistics, 34(2):151–174, 2000. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Mahmoud:1973:OSS

[MR73a] M. W. Mahmoud and Aisha Ragab. On order statistics in samples drawn from the logistic distribution. Mathematische Operationsforschung und Statistik, 4(1):81–88, 1973. CODEN MTOSBI. ISSN 0047-6277.

Mahmoud:1973:PDF

[MR73b] M. W. Mahmoud and Aisha Ragab. The probability density functions of the midrange, the ratio of two independent ranges and other criteria based on order statistics in samples drawn from an exponential distribution. Mathematische Operationsforschung und Statistik, 4(2):165–172, 1973. CODEN MTOSBI. ISSN 0047-6277.

Mahmoud:1975:OSS

[MR75] M. W. Mahmoud and Aisha Ragab. On order statistics in samples drawn from the extreme value distribution. Mathematische Operationsforschung und Statistik, 6(5):809–816, 1975. CODEN MTOSBI. ISSN 0047-6277.

Mnatsakanov:2012:MDE

[MR12] R. M. Mnatsakanov and F. H. Ruymgaart. Moment density estimation for positive random variables. Statistics: a Journal of Theoretical and Applied Statistics, 46(2):215–230, 2012. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Muller:1998:BR

[MRR^{+}98] C. Müller, W.-D. Richter, W.-D. Richter, W.-D. Richter, K. Hoffman, and Ch. Lederer. Book review. Statistics: a Journal of Theoretical and Applied Statistics, 31(4):383–384, 1998. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Mauersberger:1980:BET

[MS80a] R. Mauersberger and W. H. Schmidt. Berry–Esséen type bounds for nonlinear models. Statistics (Berlin, DDR), 11 (4):465–473, 1980. CODEN MOSSD5. ISSN 0323-3944.

Mecke:1980:FSP

[MS80b] J. Mecke and D. Stoyan. Formulas for stationary planar fibre processes I — general theory. Statistics (Berlin, DDR), 11 (2):267–279, 1980. CODEN MOSSD5. ISSN 0323-3944.

Mohner:1982:BR

[MS82] M. Möhner and D. Stoyan. Book review. Statistics (Berlin, DDR), 13(3):454, 1982. CODEN MOSSD5. ISSN 0323-3944.

Misra:1994:EOL

[MS94] Neeraj Misra and Harshinder Singh. Estimation of ordered location parameters: the exponential distribution. Statistics: a Journal of Theoretical and Applied Statistics, 25(3):239– 249, 1994. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Murari:1998:BRF

[MS98] Mitra Murari and K. Basu Sujit. A bound for the renewal function of a process driven by a non-monotonic life distribution function. Statistics: a Journal of Theoretical and Applied Statistics, 31(4):339–345, 1998. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Muller:2016:TLE

[MSCC16] Christine H. Müller, Sebastian Szugat, Nuri Celik, and Brenton R. Clarke. Trimmed likelihood estimators for lifetime experiments and their influence functions. Statistics: a Journal of Theoretical and Applied Statistics, 50(3):505–524, 2016. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Mecke:1976:RGB

[MSK^{+}76] J. Mecke, D. Stoyan, J. Kleffe, H. J. Girlich, and D. König. Rezensionen. (German) [Book reviews]. Mathematische Operationsforschung und Statistik, 7(4):640–644, 1976. CODEN MTOSBI. ISSN 0047-6277.

Muller:1971:RGB

[MSS^{+}71] W. Müller, D. Schreiter, H. J. Sprenger, D. Schreuter, G. Güssefeldt, R. Theodorescu, H. Thiele, H. Toutenburg, P. Neumeann, O. Bunke, and W. Dück. Rezensionen. (German) [Book reviews]. Mathematische Operationsforschung und Statistik, 2(3):247–254, 1971. CODEN MTOSBI. ISSN 0047-6277.

Manteuffel:1972:RGB

[MSW^{+}72] K. Manteuffel, M. Schoch, N. N. Worobjoff, D. Rasch, Vladimir Strnad, H. Toutenburg, W. Dück, E. Griepentrog, K. H. Elster, W. Kempe, and G. Nägler. Rezensionen. (German) [Book reviews]. Mathematische Operationsforschung und Statistik, 3(6):483–491, 1972. CODEN MTOSBI. ISSN 0047-6277.

Muller:2003:EEV

[MSW03] Ursula U. Müller, Anton Schick, and Wolfgang Wefelmeyer. Estimating the error variance in nonparametric regression by a covariate-matched U-statistic. Statistics: a Journal of Theoretical and Applied Statistics, 37(3):179–188, 2003. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Muller:2013:VBE

[MSW13] Ursula U. Müller, Anton Schick, and Wolfgang Wefelmeyer. Variance bounds for estimators in autoregressive models with constraints. Statistics: a Journal of Theoretical and Applied Statistics, 47(3):477–493, 2013. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Morgan:1999:CNB

[MU99] John P. Morgan and Nizam Uddin. A class of neighbor balanced complete block designs and their efficiencies for spatially correlated errors. Statistics: a Journal of Theoretical and Applied Statistics, 32(4):317–330, 1999. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Muhleis:1980:KVP

[Müh80] Wiltrud Mühleis. Konvergenzgeschwindigkeit der Verteilung einer parameterschätzung für stationäre Gaußsche Folgen. (German) [Convergence rate of the distribution of a parameter estimation for stationary Gaussian sequences]. Statistics (Berlin, DDR), 11(3):341–360, 1980. CODEN MOSSD5. ISSN 0323-3944.

Muhleis:1984:APE

[Müh84] Wiltrud Mühleis. On asymptotic properties of estimators stationary Gaussian random sequences. Statistics (Berlin, DDR), 15(2):249–262, 1984. CODEN MOSSD5. ISSN 0323-3944.

Muller:1974:MMD

[Mül74] J. A. Müller. Mathematische Modellierung dynamischer Systeme (Identifikation). (German) [Mathematical modeling of dynamic systems (identification)]. Mathematische Operationsforschung und Statistik, 5(4–5):413–438, 1974. CODEN MTOSBI. ISSN 0047-6277.

Muller:1985:EUC

[Mül85] Jochen Müller. Existence of unbiased covariance components estimators. Statistics: a Journal of Theoretical and Applied Statistics, 16(2):179–183, 1985. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Muller:1995:D

[Mül95] Werner G. Müller. Discussion. Statistics: a Journal of Theoretical and Applied Statistics, 26(2):166–167, 1995. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Munk:1998:TE

[Mun98] Axel Munk. Tchebycheff-experiments. Statistics: a Journal of Theoretical and Applied Statistics, 31(4):289–324, 1998. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Murakami:2015:PMW

[Mur15] Hidetoshi Murakami. The power of the modified Wilcoxon rank-sum test for the one-sided alternative. Statistics: a Journal of Theoretical and Applied Statistics, 49(4):781–794, 2015. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Musiela:1977:SEP

[Mus77] Marek Musiela. Sequential estimation of parameters of a stochastic differential equation. Statistics (Berlin, DDR), 8 (4):483–498, 1977. CODEN MOSSD5. ISSN 0323-3944.

Montanari:2010:IFA

[MV10] Angela Montanari and Cinzia Viroli. The independent factor analysis approach to latent variable modelling. Statistics: a Journal of Theoretical and Applied Statistics, 44(4):397–416, 2010. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Manel:2018:ELV

[MV18] Kacem Manel and Maume-Deschamps Véronique. Estimation of the limit variance for sums under a new weak dependence condition. Statistics: a Journal of Theoretical and Applied Statistics, 52(2):273–287, 2018. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Muhleis:1980:MKS

[MW80] Wiltrud Mühleis and Gisela Wittwer. Minimum-Kontrast-Schätzungen für stochastische Prozesse. (German) [Minimumcontrast estimates for stochastic processes]. Statistics (Berlin, DDR), 11(2):219–227, 1980. CODEN MOSSD5. ISSN 0323-3944.

Miao:2014:MDP

[MW14] Yu Miao and Yanling Wang. Moderate deviation principle for maximum-likelihood estimator. Statistics: a Journal of Theoretical and Applied Statistics, 48(4):766–777, 2014. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Miao:2015:CLE

[MWZ15] Yu Miao, Yanling Wang, and Haojiang Zheng. Consistency of LS estimators in the EV regression model with martingale difference errors. Statistics: a Journal of Theoretical and Applied Statistics, 49(1):104–118, 2015. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Miao:2011:LLL

[MY11] Yu Miao and Guangyu Yang. The loglog law for LS estimator in simple linear EV regression models. Statistics: a Journal of Theoretical and Applied Statistics, 45(2):155–162, 2011. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Matsuura:2018:STH

[MYK18] Shun Matsuura, Haruka Yamashita, and Kimberly K. J. Kinateder. A statistical test for the hypothesis of Gaussian random function. Statistics: a Journal of Theoretical and Applied Statistics, 52(4):801–817, 2018. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Michalski:1996:THV

[MZ96] Andrzej Michalski and Roman Zmyślony. Testing hypotheses for variance components in mixed linear models. Statistics: a Journal of Theoretical and Applied Statistics, 27(3–4):297– 310, 1996. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Mahdizadeh:2018:SER

[MZ18] M. Mahdizadeh and Ehsan Zamanzade. Smooth estimation of a reliability function in ranked set sampling. Statistics: a Journal of Theoretical and Applied Statistics, 52(4):750–768, 2018. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Nekoukhou:2013:DGE

[NAB13] V. Nekoukhou, M. H. Alamatsaz, and H. Bidram. Discrete generalized exponential distribution of a second type. Statistics: a Journal of Theoretical and Applied Statistics, 47(4): 876–887, 2013. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Nadaraya:1980:MDK

[Nad80] E. Nadaraya. On maximal deviation of the kernel type nonparametric density estimates in some degenerate cases. Statistics (Berlin, DDR), 11(4):483–497, 1980. CODEN MOSSD5. ISSN 0323-3944.

Nadarajah:2003:KTD

[Nad03] Saralees Nadarajah. The Kotz-type distribution with applications. Statistics: a Journal of Theoretical and Applied Statistics, 37(4):341–358, 2003. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Nadarajah:2005:EPD

[Nad05] Saralees Nadarajah. Exponentiated Pareto distributions. Statistics: a Journal of Theoretical and Applied Statistics, 39(3):255–260, 2005. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Nadarajah:2006:EDS

[Nad06] Saralees Nadarajah. Exact distributions of XY for some bivariate exponential distributions. Statistics: a Journal of Theoretical and Applied Statistics, 40(4):307–324, 2006. CO-DEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Nadarajah:2007:LCP

[Nad07] Saralees Nadarajah. The linear combination, product and ratio of Laplace random variables. Statistics: a Journal of Theoretical and Applied Statistics, 41(6):535–545, 2007. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Nadarajah:2008:BDA

[Nad08a] Saralees Nadarajah. The bivariate F distribution with application to drought data. Statistics: a Journal of Theoretical and Applied Statistics, 42(6):535–546, 2008. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Nadarajah:2008:GD

[Nad08b] Saralees Nadarajah. On the generalized t(GT) distribution. Statistics: a Journal of Theoretical and Applied Statistics, 42(5):467–473, 2008. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Nadarajah:2008:SGD

[Nad08c] Saralees Nadarajah. Some gamma distributions. Statistics: a Journal of Theoretical and Applied Statistics, 42(1):77–94, 2008. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic). See corrigendum [SN14].

Nadarajah:2010:SEB

[Nad10] Saralees Nadarajah. Simple expressions for a bivariate chisquare distribution. Statistics: a Journal of Theoretical and Applied Statistics, 44(2):189–201, 2010. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Nadarajah:2012:KTR

[Nad12] Saralees Nadarajah. The Kotz type ratio distribution. Statistics: a Journal of Theoretical and Applied Statistics, 46(2): 167–174, 2012. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Nagler:1975:AGA

[Näg75] G. Nägler. Ein allgemeines Grundmodell für Ablaufpläne. (German) [A general basic model for roadmaps]. Mathematische Operationsforschung und Statistik, 6(1):55–70, 1975. CO-DEN MTOSBI. ISSN 0047-6277.

Nagel:1978:GEE

[Nag78] S. Nagel. Goodness of exact experimental designs with respect to a certain family of criteria. Statistics (Berlin, DDR), 9(4): 519–523, 1978. CODEN MOSSD5. ISSN 0323-3944.

Nagel:1983:DSS

[Nag83] W. Nagel. Dünne Schnitte von stationären räumlichen Faserprozessen. (German) [Thin sections of stationary spatial fiber processes]. Statistics (Berlin, DDR), 14(4):569–576, 1983. CODEN MOSSD5. ISSN 0323-3944.

Nagaraja:1992:OSD

[Nag92] H. N. Nagaraja. Order statistics from discrete distributions. Statistics: a Journal of Theoretical and Applied Statistics, 23(3):189–209, 1992. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Nagaev:1997:BR

[Nag97] A. Nagaev. Book review. Statistics: a Journal of Theoretical and Applied Statistics, 29(3):295–297, 1997. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Nigm:2003:BOS

[NAHJ03] A. M. Nigm, Essam K. Al-Hussaini, and Zeinhum F. Jaheen. Bayesian one-sample prediction of future observations under Pareto distribution. Statistics: a Journal of Theoretical and Applied Statistics, 37(6):527–536, 2003. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Natheb:1972:MIL

[Nät72] Wolfgang Nätheb. Minimaxschätzungen im linearen regressionsmodell. (German) [Minimax estimates in the linear regression model]. Mathematische Operationsforschung und Statistik, 3(6):475–482, 1972. CODEN MTOSBI. ISSN 0047-6277.

Nather:1975:CEE

[Nät75a] Wolfgang Näther. The choice of estimators and experimental designs in a linear regression model according to a joint criterion of optimality. Mathematische Operationsforschung und Statistik, 6(5):677–686, 1975. CODEN MTOSBI. ISSN 0047-6277.

Nather:1975:SOB

[Nät75b] Wolfgang Näther. Semi-orderings between distribution functions and their application to robustness of parameter estimators. Mathematische Operationsforschung und Statistik, 6 (2):179–188, 1975. CODEN MTOSBI. ISSN 0047-6277.

Nather:1979:MRB

[Nät79] Wolfgang Näther. Monotony relations between distribution functions and their application to experimental design. Statistics (Berlin, DDR), 10(1):47–53, 1979. CODEN MOSSD5. ISSN 0323-3944.

Nather:1984:OOS

[Nät84a] W. Näther. Optimal observing of stochastic processes. Statistics (Berlin, DDR), 15(2):239–247, 1984. CODEN MOSSD5. ISSN 0323-3944.

Nather:1984:BET

[Nät84b] Wolfgang Näther. Bayes estimation of the trend parameter in random fields. Statistics (Berlin, DDR), 15(4):553–558, 1984. CODEN MOSSD5. ISSN 0323-3944.

Nather:1985:EDR

[Nät85] Wolfgang Näther. Exact designs for regression models with correlated errors. Statistics: a Journal of Theoretical and Applied Statistics, 16(4):479–484, 1985. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Nather:1989:BR

[Nat89] W. Nather. Book reviews. Statistics: a Journal of Theoretical and Applied Statistics, 20(4):558, 1989. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Nather:1991:BR

[Nät91] W. Näther. Book reviews. Statistics: a Journal of Theoretical and Applied Statistics, 22(3):478, 1991. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Nather:1997:LSI

[Nät97] Wolfgang Näther. Linear statistical inference for random fuzzy data. Statistics: a Journal of Theoretical and Applied Statistics, 29(3):221–240, 1997. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Nowzohour:2016:SBC

[NB16] Christopher Nowzohour and Peter Bühlmann. Score-based causal learning in additive noise models. Statistics: a Journal of Theoretical and Applied Statistics, 50(3):471–485, 2016. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Nieto-Barajas:2021:CDD

[NB21] Luis E. Nieto-Barajas. A class of dependent Dirichlet processes via latent multinomial processes. Statistics: a Journal of Theoretical and Applied Statistics, 55(5):1169–1179, 2021. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Nikolakopoulos:2023:EMK

[NCJ23] Stavros Nikolakopoulos, Eric Cator, and Mart P. Janssen. Extending the Mann–Kendall test to allow for measurement uncertainty. Statistics: a Journal of Theoretical and Applied Statistics, 57(3):577–596, 2023. CO-DEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/02331888.2023.2214942.

Nadarajah:2005:MMD

[ND05] Saralees Nadarajah and Dipak K. Dey. Multitude of multivariate t-distributions. Statistics: a Journal of Theoretical and Applied Statistics, 39(2):149–181, 2005. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Ndongo:2016:ESF

[NDDDG16] Mor Ndongo, Abdou Kâ Diongue, Aliou Diop, and Simplice Dossou-Gbété. Estimation for seasonal fractional ARIMA with stable innovations via the empirical characteristic function method. Statistics: a Journal of Theoretical and Applied Statistics, 50(2):298–311, 2016. CO-DEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/02331888.2015.1086769.

Nehse:1975:BEE

[Neh75] Reinhard Nehse. Bemerkungen zu einem ergebnis von J. Zowe. (German) [Remarks on a result of J. Zowe]. Mathematische Operationsforschung und Statistik, 6(6):929–931, 1975. CODEN MTOSBI. ISSN 0047-6277.

Netjdecker:1990:VMM

[Net90] El Netjdecker. The variance matrix of a matrix quadratic form under normality assumptions. Statistics: a Journal of Theoretical and Applied Statistics, 21(3):455–459, 1990. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Neuhaus:1973:VEV

[Neu73] Georg Neuhaus. Zur Verteilungskonvergenz einiger Varianten der Cramér–von Mises-Statistik. (German) [On the convergence in distribution of some variants of the Cramér– von Mises statistic]. Mathematische Operationsforschung und Statistik, 4(6):473–484, 1973. CODEN MTOSBI. ISSN 0047-6277.

Neuhaus:1979:ATG

[Neu79] Georg Neuhaus. Asymptotic theory of goodness of fit tests when parameters are present: a survey. Statistics (Berlin, DDR), 10(3):479–494, 1979. CODEN MOSSD5. ISSN 0323-3944.

Neumann:1992:SOA

[Neu92] Michael H. Neumann. Second order asymptotic risks of smoothed linear estimators in nonparametric regression models. Statistics: a Journal of Theoretical and Applied Statistics, 23(3):217–236, 1992. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Neumann:1994:FDD

[Neu94] Michael H. Neumann. Fully data-driven nonparametric variance estimators. Statistics: a Journal of Theoretical and Applied Statistics, 25(3):189–212, 1994. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Neumann:1997:PCI

[Neu97] Michael H. Neumann. Pointwise confidence intervals in nonparametric regression with heteroscedastic error structure. Statistics: a Journal of Theoretical and Applied Statistics, 29 (1):1–36, 1997. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Neumann:1999:BR

[Neu99] Peter Neumann. Book review. Statistics: a Journal of Theoretical and Applied Statistics, 33(3):287–289, 1999. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Neumann:2002:RMB

[Neu02] Michael H. Neumann. On robustness of model-based bootstrap schemes in nonparametric time series analysis. Statistics: a Journal of Theoretical and Applied Statistics, 36(1): 33–63, 2002. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Nunes:2023:NWM

[NFF^{+}23] Célia Nunes, Dário Ferreira, Sandra S. Ferreira, Miguel Fonseca, Manuela M. Oliveira, and João T. Mexia. Noncentral Wishart matrices, asymptotic normality of vec and smooth statistics. Statistics: a Journal of Theoretical and Applied Statistics, 57(2):261–281, 2023. CO-DEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/02331888.2023.2174117.

Noughabi:2018:EMB

[NFP18] M. Shafaei Noughabi and A. M. Franco-Pereira. Estimation of monotone bivariate quantile residual life. Statistics: a Journal of Theoretical and Applied Statistics, 52(4):919–933, 2018. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Nadarajah:2011:EED

[NH11] Saralees Nadarajah and Firoozeh Haghighi. An extension of the exponential distribution. Statistics: a Journal of Theoretical and Applied Statistics, 45(6):543–558, 2011. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Niemiro:2006:BPA

[Nie06] Wojciech Niemiro. Bayesian prediction with an asymmetric criterion in a nonparametric model of insurance risk. Statistics: a Journal of Theoretical and Applied Statistics, 40(4):

353–363, 2006. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Niebuhr:2017:IOT

[Nie17] Tobias Niebuhr. Irregularly observed time series — some asymptotics and the block bootstrap. Statistics: a Journal of Theoretical and Applied Statistics, 51(5):1118–1131, 2017. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Nigm:1990:PBU

[Nig90] A. M. Nigm. Prediction bounds using Type 1 censoring for the Weibull distribution. Statistics: a Journal of Theoretical and Applied Statistics, 21(2):227–237, 1990. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Nazari:2016:DFE

[NJKK16] Sahar Nazari, Mohammad Jafari Jozani, and Mahmood Kharrati-Kopaei. On distribution function estimation with partially rank-ordered set samples: estimating mercury level in fish using length frequency data. Statistics: a Journal of Theoretical and Applied Statistics, 50(6):1387–1410, 2016. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Naik:1988:QPA

[NK88] V. K. R. Naik and A. Krishnamoorthy. On a queueing process with arrival and service rates depending on the number of units served. Statistics: a Journal of Theoretical and Applied Statistics, 19(1):117–121, 1988. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Nadarajah:2004:MBD

[NK04] Saralees Nadarajah and Samuel Kotz. Multitude of bivariate t distributions. Statistics: a Journal of Theoretical and Applied Statistics, 38(6):527–539, 2004. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Nadarajah:2005:PES

[NK05a] Saralees Nadarajah and Samuel Kotz. On the product XY for the elliptically symmetric Kotz type distribution. Statistics: a Journal of Theoretical and Applied Statistics, 39(3):269–274, 2005. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Nadarajah:2005:SBB

[NK05b] Saralees Nadarajah and Samuel Kotz. Some bivariate beta distributions. Statistics: a Journal of Theoretical and Applied Statistics, 39(5):457–466, 2005. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Nadarajah:2006:MAS

[NK06] Saralees Nadarajah and Samuel Kotz. Models for amplitude statistics. Statistics: a Journal of Theoretical and Applied Statistics, 40(2):175–183, 2006. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Nadarajah:2007:MBD

[NK07] Saralees Nadarajah and Samuel Kotz. Multitude of beta distributions with applications. Statistics: a Journal of Theoretical and Applied Statistics, 41(2):153–179, 2007. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Nandi:2013:EPP

[NK13] Swagata Nandi and Debasis Kundu. Estimation of parameters of partially sinusoidal frequency model. Statistics: a Journal of Theoretical and Applied Statistics, 47(1):45–60, 2013. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Nekoukhou:2017:BDG

[NK17] Vahid Nekoukhou and Debasis Kundu. Bivariate discrete generalized exponential distribution. Statistics: a Journal of Theoretical and Applied Statistics, 51(5):1143–1158, 2017. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Nandi:2023:EPM

[NK23] Swagata Nandi and Debasis Kundu. Estimating parameters in multichannel fundamental frequency with harmonics model. Statistics: a Journal of Theoretical and Applied Statistics, 57(5):1142–1164, 2023. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
02331888.2023.2253992.

Nandi:2004:AMM

[NKI04] Swagata Nandi, Debasis Kundu, and Srikanth K. Iyer. Amplitude modulated model for analyzing non-stationary speech signals. Statistics: a Journal of Theoretical and Applied Statistics, 38(5):439–456, 2004. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Nadarajah:2021:CSD

[NKT21] Saralees Nadarajah, Hok Shing Kwong, and Fatih Tank. Compound sum distributions with dependence. Statistics: a Journal of Theoretical and Applied Statistics, 55(2):409–425, 2021. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Nkurunziza:2008:LRT

[Nku08] Sévérien Nkurunziza. Likelihood ratio test for a special predator-prey system. Statistics: a Journal of Theoretical and Applied Statistics, 42(2):149–166, 2008. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Nkurunziza:2012:RPS

[Nku12] Sévérien Nkurunziza. The risk of pretest and shrinkage estimators. Statistics: a Journal of Theoretical and Applied Statistics, 46(3):305–312, 2012. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Nkurunziza:2023:CNL

[Nku23] Sévérien Nkurunziza. Corrigendum: A note on Liu-type shrinkage estimations in linear models (Statistics 56, 396– 420). Statistics: a Journal of Theoretical and Applied Statistics, 57(4):761–763, 2023. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/02331888.2023.2234061.

See [YAA22].

Nalbach-Leniewska:1979:MDM

[NL79] A. Nalbach-Leniewska. Measures of dependence of the multivariate lognormal distribution. Statistics (Berlin, DDR), 10 (3):381–387, 1979. CODEN MOSSD5. ISSN 0323-3944.

Na:2003:FIT

[NL03] Myung Hwan Na and Sangyeol Lee. A family of IDMRL tests with unknown turning point. Statistics: a Journal of Theoretical and Applied Statistics, 37(5):457–462, 2003. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Nguessan:2005:CCE

[NL05] Assi N’guessan and Claude Langrand. A covariance components estimation procedure when modelling a road safety measure in terms of linear constraints. Statistics: a Journal of Theoretical and Applied Statistics, 39(4):303–314, 2005. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Neudecker:1995:HPS

[NLP95] H. Neudecker, S. Liu, and W. Polasek. The Hadamard product and some of its applications in statistics. Statistics: a Journal of Theoretical and Applied Statistics, 26(4):365–373, 1995. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Nishino:2018:NNN

[NM18] Takuya Nishino and Hidetoshi Murakami. The null and nonnull limiting distributions of the modified multisample Cucconi test. Statistics: a Journal of Theoretical and Applied Statistics, 52(6):1344–1358, 2018. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Nagel:1982:VFG

[NN82] S. Nagel and W. Nagel. Versuchsplanung zu einer Familie von gestauchten Schätzfunktionen im linearen Regressionsmodell. (German) [Experimental design of a family of estimated model functions in the linear regression model]. Statistics (Berlin, DDR), 13(4):521–530, 1982. CODEN MOSSD5. ISSN 0323-3944.

Nair:1991:CGB

[NN91] N. Unnikrishnan Nair and K. R. Muraleedharan Nair. Characterization of the Gumbel’s bivarite exponential distribution. Statistics: a Journal of Theoretical and Applied Statistics, 22 (1):147, 1991. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Nooghabi:2011:SR

[NN11] M. Jabbari Nooghabi and H. Jabbari Nooghabi. Statement of retraction. Statistics: a Journal of Theoretical and Applied Statistics, 45(2):121–122, 2011. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Nawa:2023:CFE

[NN23a] Victor Mooto Nawa and Saralees Nadarajah. Closed form estimators for a multivariate gamma distribution. Statistics: a Journal of Theoretical and Applied Statistics, 57(2):482– 495, 2023. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-

4910 (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/02331888.2023.2183204.

Nawa:2023:NCF

[NN23b] Victor Mooto Nawa and Saralees Nadarajah. New closed form estimators for a bivariate gamma distribution. Statistics: a Journal of Theoretical and Applied Statistics, 57(1): 150–160, 2023. CODEN MOSSD5. ISSN 0233-1888 (print),

1029-4910 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/02331888.2023.2168004.

Nadarajah:2014:TES

[NNM14] Saralees Nadarajah, Vahid Nassiri, and Adel Mohammadpour. Truncated-exponential skew-symmetric distributions. Statistics: a Journal of Theoretical and Applied Statistics, 48(4):872–895, 2014. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Nogales:2020:CEM

[Nog20] A. G. Nogales. Conditional expectation of a Markov kernel given another with some applications in statistical inference and disease diagnosis. Statistics: a Journal of Theoretical and Applied Statistics, 54(2):239–256, 2020. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Nollau:1987:BR

[Nol87] V. Nollau. Book review. Statistics: a Journal of Theoretical and Applied Statistics, 18(1):98, 1987. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Nollau:1988:BR

[Nol88] V. Nollau. Book reviews. Statistics: a Journal of Theoretical and Applied Statistics, 19(1):38, 1988. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Nunes:2013:ADD

[NOM13] Célia Nunes, Manuela M. Oliveira, and João T. Mexia. Application domains for the delta method. Statistics: a Journal of Theoretical and Applied Statistics, 47(2):317–328, 2013. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Norberg:1990:RTS

[Nor90] Ragnar Norberg. Risk theory and its statistics environment. Statistics: a Journal of Theoretical and Applied Statistics, 21(2):273–299, 1990. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Novak:2010:ICE

[Nov10] S. Y. Novak. Impossibility of consistent estimation of the distribution function of a sample maximum. Statistics: a Journal of Theoretical and Applied Statistics, 44(1):25–30, 2010. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Nozicka:1970:ELL

[Nož70] František Nožička. Über die Eindeutigkeit der Lösung von linearen Optimierungsproblemen. (German) [On the uniqueness of the solution of linear programming problems]. Mathematische Operationsforschung und Statistik, 1(1):5–20, 1970. CODEN MTOSBI. ISSN 0047-6277.

Nozicka:1971:GBD

[Nož71] František Nožička. Ein geometrischer Beweis des Dualitätsprinzips in der linearen Optimierung. (German) [A geometric proof of the duality principle in linear optimization]. Mathematische Operationsforschung und Statistik, 2(1):3–17, 1971. CODEN MTOSBI. ISSN 0047-6277.

Nozicka:1972:KLE

[Nož72] František Nožička. Über eine Klasse von linearen einparametrischen Optimierungsproblemen. (German) [On a class of nonlinear one-parameter optimization problems]. Mathematische Operationsforschung und Statistik, 3(3):159– 194, 1972. CODEN MTOSBI. ISSN 0047-6277.

Nadarajah:2012:CFS

[NP12] Saralees Nadarajah and Tibor K. Pogány. Characteristic function of the SGT distribution. Statistics: a Journal of Theoretical and Applied Statistics, 46(4):437–439, 2012. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Nadarajah:2012:CFB

[NPS12] Saralees Nadarajah, Tibor K. Pogány, and Ram K. Saxena. On the characteristic function for Burr distributions. Statistics: a Journal of Theoretical and Applied Statistics, 46(3): 419–428, 2012. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Nather:1981:DOW

[NR81] W. Näther and V. Reinsch. D_{s}-optimality and Whittle’s equivalence theorem. Statistics (Berlin, DDR), 12(3):307– 316, 1981. CODEN MOSSD5. ISSN 0323-3944.

Navarro:2006:MWD

[NRD06] Jorge Navarro, Jose M. Ruiz, and Yolanda Del Aguila. Multivariate weighted distributions: a review and some extensions. Statistics: a Journal of Theoretical and Applied Statistics, 40 (1):51–64, 2006. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Navarro:2008:COP

[NRD08] Jorge Navarro, Jose M. Ruiz, and Yolanda Del Aguila. Characterizations and ordering properties based on log-odds functions. Statistics: a Journal of Theoretical and Applied Statistics, 42(4):313–328, 2008. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Navarro:2007:SCM

[NRS07] Jorge Navarro, Jose M. Ruiz, and Carlos J. Sandoval. Some characterizations of multivariate distributions using products of the hazard gradient and mean residual life components. Statistics: a Journal of Theoretical and Applied Statistics, 41 (1):85–91, 2007. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Nair:2003:FIB

[NS03] N. Unnikrishnan Nair and S. Sunoj. Form-invariant bivariate weighted models. Statistics: a Journal of Theoretical and Applied Statistics, 37(3):259–269, 2003. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Nair:2010:GPC

[NS10a] N. Unnikrishnan Nair and Johny Scaria. General properties and concomitants of Morgenstern bivariate gamma distribution and their applications in estimation. Statistics: a Journal of Theoretical and Applied Statistics, 44(2):155–167, 2010. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Nather:2010:OHW

[NS10b] Wolfgang Näther and Dietrich Stoyan. Obituary: Hans-Walter Bandemer (1 April 1932–15 November 2009). Statistics: a Journal of Theoretical and Applied Statistics, 44(4): 323–326, 2010. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Nosek:2014:CPD

[NS14] Konrad Nosek and Zbigniew Szkutnik. Change-point detection in a shape-restricted regression model. Statistics: a Journal of Theoretical and Applied Statistics, 48(3):641–656, 2014. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Nakajima:2022:PES

[NS22] Shohei Nakajima and Yasutaka Shimizu. Parameter estimation of stochastic differential equation driven by small fractional noise. Statistics: a Journal of Theoretical and Applied Statistics, 56(4):919–934, 2022. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

URL http://www.tandfonline.com/doi/full/10.1080/
02331888.2022.2098960.

Nadarajah:2007:SBF

[NSG07] Saralees Nadarajah, H. M. Srivastava, and Arjun K. Gupta. Skewed Bessel function distributions with application to rainfall data. Statistics: a Journal of Theoretical and Applied Statistics, 41(4):333–344, 2007. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Nather:1980:BR

[NSH^{+}80] W. Näther, W. H. Schmidt, H. P. Höschel, V. Dupac, H. Bandemer, I. Väduva, Volker Nollau, C. Polze, A. Hahnewald-Busch, and M. Goebel. Book reviews. Statistics (Berlin, DDR), 11(1):164–171, 1980. CODEN MOSSD5. ISSN 0323-3944.

Nair:2012:MLQ

[NSK12] N. Unnikrishnan Nair, P. G. Sankaran, and B. Vinesh Kumar. Modelling lifetimes by quantile functions using Parzen’s score function. Statistics: a Journal of Theoretical and Applied Statistics, 46(6):799–811, 2012. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Nadarajah:2017:NBB

[NSN17] Saralees Nadarajah, Shou Hsing Shih, and Daya K. Nagar. A new bivariate beta distribution. Statistics: a Journal of Theoretical and Applied Statistics, 51(2):455–474, 2017. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Nielsen:2009:LLD

[NTJ09] Jens Perch Nielsen, Carsten Tanggaard, and M. C. Jones. Local linear density estimation for filtered survival data, with bias correction. Statistics: a Journal of Theoretical and Applied Statistics, 43(2):167–186, 2009. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Nussbaum:1976:MLL

[Nus76a] M. Nussbaum. Maximum likelihood and least squares estimation of linear functional relationships. Mathematische Operationsforschung und Statistik, 7(1):23–49, 1976. CODEN MTOSBI. ISSN 0047-6277.

Nussbaum:1976:SDM

[Nus76b] Michael Nussbaum. Structural distributions in the multivariate linear model. Mathematische Operationsforschung und Statistik, 7(5):679–683, 1976. CODEN MTOSBI. ISSN 0047-6277.

Nussbaum:1977:AEE

[Nus77a] M. Nussbaum. Asymptotic efficiency of estimators in the multivariate linear model. Statistics (Berlin, DDR), 8(4):439– 445, 1977. CODEN MOSSD5. ISSN 0323-3944.

Nussbaum:1977:AOE

[Nus77b] M. Nussbaum. Asymptotic optimality of estimators of a linear functional relation if the ratio of the error variances is known. Statistics (Berlin, DDR), 8(2):173–198, 1977. CO-DEN MOSSD5. ISSN 0323-3944.

Nussbaum:1979:AEE

[Nus79] Michael Nussbaum. Asymptotic efficiency of estimators of a multivariate linear functional relation. Statistics (Berlin, DDR), 10(4):505–527, 1979. CODEN MOSSD5. ISSN 0323-3944.

Niemiro:2012:LEP

[NW12] Wojciech Niemiro and Jacek Wesolowski. Linear estimation and prediction under model-design approach with small area effects. Statistics: a Journal of Theoretical and Applied Statistics, 46(4):523–547, 2012. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Ozkale:2016:NBE

[ÖA16] M. Revan Özkale and Engin Arican. A new biased estimator in logistic regression model. Statistics: a Journal of Theoretical and Applied Statistics, 50(2):233–253, 2016. CO-DEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/02331888.2015.1123711.

Ozdemir:2021:ELM

[ÖA21] Senay Özdemir and Olcay Arslan. Empirical likelihood-MM (EL-MM) estimation for the parameters of a linear regression model. Statistics: a Journal of Theoretical and Applied Statistics, 55(1):45–67, 2021. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Ozturk:2019:CQC

[OB19] Omer Ozturk and Narayanaswamy Balakrishnan. Constructing quantile confidence intervals using extended simple random sample in finite populations. Statistics: a Journal of Theoretical and Applied Statistics, 53(4):792–806, 2019. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Ollerer:2015:IFP

[ÖCA15] Viktoria Öllerer, Christophe Croux, and Andreas Alfons. The influence function of penalized regression estimators. Statistics: a Journal of Theoretical and Applied Statistics, 49(4): 741–765, 2015. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Ohser:1981:RER

[Ohs81] J. Ohser. A remark on the estimation of the rose of directions of fibre processes. Statistics (Berlin, DDR), 12(4):581–585, 1981. CODEN MOSSD5. ISSN 0323-3944.

Ohser:1983:ERS

[Ohs83] J. Ohser. On estimators for the reduced second moment measure of point processes. Statistics (Berlin, DDR), 14(1):63–71, 1983. CODEN MOSSD5. ISSN 0323-3944.

Ohser:1991:BR

[Ohs91] J. Ohser. Book reviews. Statistics: a Journal of Theoretical and Applied Statistics, 22(1):132, 1991. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Ohyauchi:2013:CRE

[Ohy13] N. Ohyauchi. Comparison of risks of estimators under the LINEX loss for a family of truncated distributions. Statistics: a Journal of Theoretical and Applied Statistics, 47(3):590– 604, 2013. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Obradovic:2015:GFT

[OJM15] Marko Obradović, Milan Jovanović, and Bojana Milosević. Goodness-of-fit tests for Pareto distribution based on a characterization and their asymptotics. Statistics: a Journal of Theoretical and Applied Statistics, 49(5):1026–1041, 2015. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Orvath:2002:CPD

[OK02] Lajos Orváth and Piotr Kokoszka. Change-point detection with non-parametric regression. Statistics: a Journal of Theoretical and Applied Statistics, 36(1):9–31, 2002. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Okolewski:2016:BEE

[Oko16] A. Okolewski. Bounds on expectations of L-estimates for maximally and minimally stable samples. Statistics: a Journal of Theoretical and Applied Statistics, 50(4):903–916, 2016. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Okolewski:2017:DBO

[Oko17] A. Okolewski. Distribution bounds for order statistics when each k-tuple has the same piecewise uniform copula. Statistics: a Journal of Theoretical and Applied Statistics, 51(5): 969–987, 2017. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Oktaba:1986:EMV

[OKW86] Wiktor Oktaba, Andrzej Kornacki, and Jacek Wawrzosek. Estimation of missing values in the general Gauss–Markoff model. Statistics: a Journal of Theoretical and Applied Statistics, 17(2):167–177, 1986. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Olaf:1991:OCS

[OM91] Krafft Olaf and Schaefer Martin. E-optimality of a class of saturated main-effect plans. Statistics: a Journal of Theoretical and Applied Statistics, 22(1):9–15, 1991. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Ong:2010:DDP

[OM10] S. H. Ong and Rahul Mukerjee. Data-dependent probability matching priors of the second order. Statistics: a Journal of Theoretical and Applied Statistics, 44(3):291–302, 2010. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Ong:2011:PMP

[OM11] S. H. Ong and Rahul Mukerjee. Probability matching priors for two-sided tolerance intervals in balanced one-way and two-way nested random effects models. Statistics: a Journal of Theoretical and Applied Statistics, 45(4):403–411, 2011. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Ouimet:2022:RNA

[Oui22] Frédéric Ouimet. Refined normal approximations for the central and noncentral chi-square distributions and some applications. Statistics: a Journal of Theoretical and Applied Statistics, 56(4):935–956, 2022. CO-DEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/02331888.2022.2084544.

Ouyang:1992:INM

[Ouy92] Zhao Ouyang. Invariant nonnegative minimum norm quadratic estimator of variance components and its application. Statistics: a Journal of Theoretical and Applied Statistics, 23(2):145–158, 1992. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Ozkale:2009:PCR

[Özk09] M. Revan Özkale. Principal components regression estimator and a test for the restrictions. Statistics: a Journal of Theoretical and Applied Statistics, 43(6):541–551, 2009. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Pardo:2014:MGM

[PA14] M. C. Pardo and R. Alonso. A more general methodology for fitting global cross-ratio models for discrete longitudinal responses. Statistics: a Journal of Theoretical and Applied Statistics, 48(5):1071–1091, 2014. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Paditz:1981:EFI

[Pad81] Ludwig Paditz. Einseitige Fehlerabschätzungen im zentralen Grenzwertsatz. (German) [One-sided error estimates in the Central Limit Theorem]. Statistics (Berlin, DDR), 12(4):587– 604, 1981. CODEN MOSSD5. ISSN 0323-3944.

Paditz:1984:EEC

[Pad84] Ludwig Paditz. On error-estimates in the Central Limit Theorem for generalized linear discounting. Statistics (Berlin, DDR), 15(4):601–610, 1984. CODEN MOSSD5. ISSN 0323-

Paditz:1987:RCC

[Pad87] Ludwig Paditz. On the rate of convergence in the Central Limit Theorem for discounted sums of martingale differences. Statistics: a Journal of Theoretical and Applied Statistics, 18(1):141–149, 1987. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Paditz:1989:ASC

[Pad89] Ludwig Paditz. On the analytical structure of the constant in the nonuniform version of the Esséen inequality. Statistics: a Journal of Theoretical and Applied Statistics, 20(3):453–464, 1989. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Paditz:1996:EBN

[Pad96] Ludwig Paditz. On the error-bound in the nonuniform version of Esseen’s inequality in the L_{p}-metric. Statistics: a Journal of Theoretical and Applied Statistics, 27(3–4):379–394, 1996. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Panaretos:1984:PIF

[Pan84] J. Panaretos. Partial independence and finite distributions. Statistics (Berlin, DDR), 15(3):397–403, 1984. CODEN MOSSD5. ISSN 0323-3944.

Papadatos:2016:MER

[Pap16] Nickos Papadatos. Maximizing the expected range from dependent observations under mean-variance information. Statistics: a Journal of Theoretical and Applied Statistics, 50(3):596–629, 2016. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Pauly:2012:CSB

[Pau12] Markus Pauly. Consistency of the subsample bootstrap empirical process. Statistics: a Journal of Theoretical and Applied Statistics, 46(5):621–626, 2012. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Pal:2005:ETT

[PAW05] Manisha Pal, M. Masoom Ali, and Jungsoo Woo. Estimation and testing of P (Y > X) in two-parameter exponential distributions. Statistics: a Journal of Theoretical and Applied Statistics, 39(5):415–428, 2005. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Pazman:1980:SED

[Páz80a] A. Pázman. Singular experimental designs (standard and Hilbert-space approaches). Statistics (Berlin, DDR), 11(1): 137–149, 1980. CODEN MOSSD5. ISSN 0323-3944.

Pazman:1980:SFO

[Páz80b] Andrej Pázman. Some features of the optimal design theory — a survey. Statistics (Berlin, DDR), 11(3):415–446, 1980. CODEN MOSSD5. ISSN 0323-3944.

Pazman:1984:NLS

[Páz84] A. Pázman. Nonlinear least squares — uniqueness versus ambiguity. Statistics (Berlin, DDR), 15(3):323–336, 1984. CO-DEN MOSSD5. ISSN 0323-3944.

Pazman:1987:FDN

[Páz87] A. Pázman. On formulas for the distribution of nonlinear L.S. estimates. Statistics: a Journal of Theoretical and Applied Statistics, 18(1):3–15, 1987. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Pazman:1990:AEDa

[Páz90a] Andrej Pázman. Almost exact distributions of estimators I — low dimensional nonlinear regression. Statistics: a Journal of Theoretical and Applied Statistics, 21(1):9–19, 1990. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Pazman:1990:AEDb

[Páz90b] Andrej Pázman. Almost exact distributions of estimators II — flat nonlinear regression models. Statistics: a Journal of Theoretical and Applied Statistics, 21(1):21–33, 1990. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Pazman:1990:IDP

[Páz90c] Andrej Pázman. Invited discussion paper: Small-sample distributional properties of nonlinear regression estimators (a geometric approach). Statistics: a Journal of Theoretical and Applied Statistics, 21(3):323–367, 1990. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Pazman:1991:PVC

[Páz91a] Andrej Pázman. Pivotal variables and confidence regions in flat nonlinear regression models with unknown σ. Statistics: a Journal of Theoretical and Applied Statistics, 22(2):177–189, 1991. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Pazman:1991:S

[Páz91b] Andrej Pázman. Statistics. Statistics: a Journal of Theoretical and Applied Statistics, 22(4):550, 1991. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Pazman:1993:HDN

[Páz93] Andrej Pázman. Higher dimensional nonlinear regression — a statistical use of the Riemannian curvature tensor. Statistics: a Journal of Theoretical and Applied Statistics, 25(1):17–25, 1993. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Pazman:1995:D

[Páz95] Andrej Pázman. Discussion. Statistics: a Journal of Theoretical and Applied Statistics, 26(2):167–168, 1995. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Platen:1982:BR

[PB82] E. Platen and H. Bandemer. Book review. Statistics (Berlin, DDR), 13(3):472, 1982. CODEN MOSSD5. ISSN 0323-3944.

Patriota:2010:MEM

[PB10] Alexandre Galvão Patriota and Heleno Bolfarine. Measurement error models with a general class of error distribution. Statistics: a Journal of Theoretical and Applied Statistics, 44(2):119–127, 2010. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Padellini:2021:SLI

[PB21] Tullia Padellini and Pierpaolo Brutti. Supervised learning with indefinite topological kernels. Statistics: a Journal of Theoretical and Applied Statistics, 55(4):765–786, 2021. CO-DEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Park:2011:FIP

[PBK11] Sangun Park, N. Balakrishnan, and Seong W. Kim. Fisher information in progressive hybrid censoring schemes. Statistics: a Journal of Theoretical and Applied Statistics, 45(6): 623–631, 2011. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Pericleous:2017:ODE

[PCKMK17] K. Pericleous, St. A. Chatzopoulos, F. Kolyva-Machera, and S. Kounias. Optimal designs for estimating linear and quadratic contrasts with three level factors, the case N ≡ 0mod3. Statistics: a Journal of Theoretical and Applied Statistics, 51(5):1061–1081, 2017. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Pleszczynska:1980:POS

[PD80] E. Pleszczyńska and D. Darowska. On partial observability is statistical models. Statistics (Berlin, DDR), 11(1):49–59, 1980. CODEN MOSSD5. ISSN 0323-3944.

Pons:1991:MMB

[PD91] Odile Pons and Elisabeth De Turckheim. Von Mises method, bootstrap and Hadamard differentiability for nonparametric general models. Statistics: a Journal of Theoretical and Applied Statistics, 22(2):205–214, 1991. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Pradhan:2013:NEQ

[PDG13] Biswabrata Pradhan, Anup Dewanji, and Alok Goswami. Nonparametric estimation of quality adjusted lifetime distribution in a simple illness-death model. Statistics: a Journal of Theoretical and Applied Statistics, 47(5):1030–1046, 2013. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Pauly:2015:AHD

[PEB15] Markus Pauly, David Ellenberger, and Edgar Brunner. Analysis of high-dimensional one group repeated measures designs. Statistics: a Journal of Theoretical and Applied Statistics, 49 (6):1243–1261, 2015. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Pederzoli:1986:CRG

[Ped86] Giorgio Pederzoli. Computable representations for the general density of random r-contents of beta distributed random points in an n-ball. Statistics: a Journal of Theoretical and Applied Statistics, 17(3):445–451, 1986. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Pelletier:2011:IFD

[Pel11] Bruno Pelletier. Inference in φ-families of distributions. Statistics: a Journal of Theoretical and Applied Statistics, 45(3):223–236, 2011. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Penev:1985:MEV

[Pen85] S. I. Penev. On a model with errors in variables described by stochastic differential equations. Statistics: a Journal of Theoretical and Applied Statistics, 16(4):485–493, 1985. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Pesta:2013:TLS

[Pes13] Michal Pesta. Total least squares and bootstrapping with applications in calibration. Statistics: a Journal of Theoretical and Applied Statistics, 47(5):966–991, 2013. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Petersen:1973:LME

[Pet73] I. Petersen. Linear minimax estimation for inadequate models in L^{2} metric. Mathematische Operationsforschung und Statistik, 4(6):463–471, 1973. CODEN MTOSBI. ISSN 0047-6277.

Petersen:1986:APE

[Pet86] Ivar Petersen. On the algorithms for parameter estimation in one-dimensional densities and distributions in a general purpose statistical package. Statistics: a Journal of Theoretical and Applied Statistics, 17(1):139–145, 1986. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Pflug:1980:OSZ

[Pfl80] G. Pflug. Optimale sequentielle Zerlegung. (German) [Optimal sequential decomposition]. Statistics (Berlin, DDR), 11 (2):287–295, 1980. CODEN MOSSD5. ISSN 0323-3944.

Pfuff:1978:VBV

[Pfu78] Franz Pfuff. Verallgemeinerte Bayes-Verfahren bei sequentiellen Testproblemen. (German) [Generalized Bayesian method for sequential test problems]. Statistics (Berlin, DDR), 9(2):195–209, 1978. CODEN MOSSD5. ISSN 0323-3944.

Pham-Gia:2009:MSB

[PG09] T. Pham-Gia. The multivariate Selberg beta distribution and applications. Statistics: a Journal of Theoretical and Applied Statistics, 43(1):65–79, 2009. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Park:2019:THR

[PG19] Junyong Park and Iris Ivy Gauran. Testing the homogeneity of risk differences with sparse count data. Statistics: a Journal of Theoretical and Applied Statistics, 53(6):1306–1328, 2019. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Palacios-Gonzalez:2015:FFD

[PGGF15] F. Palacios-González and R. M. García-Fernández. A flexible family of density functions. Statistics: a Journal of Theoretical and Applied Statistics, 49(3):680–704, 2015. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Pham-Gia:2002:OGV

[PGT02] T. Pham-Gia and N. Turkkan. Operations on the generalized-F variables and applications. Statistics: a Journal of Theoretical and Applied Statistics, 36(3):195–209, 2002. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Pham-Gia:2005:DRR

[PGT05] T. Pham-Gia and N. Turkkan. Distributions of ratios of random variables from the power-quadratic exponential family and applications. Statistics: a Journal of Theoretical and Applied Statistics, 39(4):355–372, 2005. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Pham-Gia:2008:BAC

[PGT08] T. Pham-Gia and N. Turkkan. Bayesian analysis of a 2×2 contingency table with dependent proportions and exact sample size. Statistics: a Journal of Theoretical and Applied Statistics, 42(2):127–147, 2008. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Pham-Gia:2010:TSU

[PGT10] T. Pham-Gia and N. Turkkan. Testing sphericity using small samples. Statistics: a Journal of Theoretical and Applied Statistics, 44(6):601–616, 2010. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Peng:2016:AME

[PH16a] Ping Peng and Guikai Hu. The admissible minimax estimator in Gauss–Markov model under a balanced loss function. Statistics: a Journal of Theoretical and Applied Statistics, 50(2):271–277, 2016. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic). URL http://

www.tandfonline.com/doi/abs/10.1080/02331888.2015.
1050023.

Provost:2016:DAM

[PH16b] Serge B. Provost and Hyung-Tae Ha. Distribution approximation and modelling via orthogonal polynomial sequences. Statistics: a Journal of Theoretical and Applied Statistics, 50(2):454–470, 2016. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic). URL http://

www.tandfonline.com/doi/abs/10.1080/02331888.2015.
1053809.

Phelan:1989:LEA

[Phe89] Michael J. Phelan. Lifetesting and estimation with arbitrary distribution function. Statistics: a Journal of Theoretical and Applied Statistics, 20(2):305–318, 1989. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Phelan:1996:EFR

[Phe96] Michael J. Phelan. An estimation function for the rate of drift, dispersion, injection, and decay of transient tracers. Statistics: a Journal of Theoretical and Applied Statistics, 27(3–4):353– 361, 1996. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Platen:1979:BR

[PHG^{+}79] E. Platen, K. Hoffmann, H. D. Gerhardt, R. Strobel, and W. Warmuth. Book reviews. Statistics (Berlin, DDR), 10(2): 347–350, 1979. CODEN MOSSD5. ISSN 0323-3944.

Phoha:1987:ABI

[Pho87] R. C. Phoha. Asymptotic Bayesian inference in some nonstandard cases: Bernstein–von Mises results and regular Bayes’ estimators. Statistics: a Journal of Theoretical and Applied Statistics, 18(2):259–274, 1987. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Provost:2009:ADI

[PHS09] Serge B. Provost, Hyung-Tae Ha, and Deepak Sanjel. On approximating the distribution of indefinite quadratic forms. Statistics: a Journal of Theoretical and Applied Statistics, 43(6):597–609, 2009. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Piehler:1970:VSM

[Pie70] Joachim Piehler. Über die Verschärfung von Schnitten in der Methode von Gomory bei der rein-ganzzahligen linearen Optimierung. (German) [On the escalation of cuts in the method of Gomory in pure-integer linear optimization]. Mathematische Operationsforschung und Statistik, 1(3):207–216, 1970. CODEN MTOSBI. ISSN 0047-6277.

Piehler:1971:SOG

[Pie71] Joachim Piehler. Simulationsmethoden in der Operationsforschung. (German) [Simulation methods in operations research]. Mathematische Operationsforschung und Statistik, 2 (2):117–128, 1971. CODEN MTOSBI. ISSN 0047-6277.

Piehler:1972:MOG

[Pie72a] Joachim Piehler. Modellsysteme der Operationsforschung und Graphentheorie. (German) [Model systems operations research and graph theory]. Mathematische Operationsforschung und Statistik, 3(2):83–89, 1972. CODEN MTOSBI. ISSN 0047-6277.

Piehler:1972:BAK

[Pie72b] Joachim Piehler. Zur Bestimmung aller Kreise in gerichteten Graphen. (German) [On the determination of all circles in the directed graph]. Mathematische Operationsforschung und Statistik, 3(5):339–342, 1972. CODEN MTOSBI. ISSN 0047-6277.

Piehler:1973:BEG

[Pie73] Joachim Piehler. Über eine bemerkenswerte Eigenschaft der Gomoryschen Schnittgruppe in der rein ganzzahligen Optimierung. (German) [A remarkable property of the Gomory cut group in pure integer programming]. Mathematische Operationsforschung und Statistik, 4(5):347–350, 1973. CODEN MTOSBI. ISSN 0047-6277.

Piehler:1974:GSD

[Pie74a] Joachim Piehler. Gomory-Schnitte und Diophantische Gleichungen. (German) [Gomory cuts and Diophantine equations]. Mathematische Operationsforschung und Statistik, 5 (4–5):259–268, 1974. CODEN MTOSBI. ISSN 0047-6277.

Piehler:1974:VGS

[Pie74b] Joachim Piehler. Über eine Verallgemeinerung der Gomoryschen Schnitte. (German) [On a generalization of the Gomory cuts]. Mathematische Operationsforschung und Statistik, 5(7–8):547–554, 1974. CODEN MTOSBI. ISSN 0047-6277.

Piehler:1975:EBS

[Pie75a] Joachim Piehler. Einige Bemerkungen zum Schnittrang in der rein-ganzzahligen linearen Optimierung. (German) [Some remarks on average rank in pure-integer linear optimization]. Mathematische Operationsforschung und Statistik, 6(4):523– 533, 1975. CODEN MTOSBI. ISSN 0047-6277.

Piehler:1975:GBG

[Pie75b] Joachim Piehler. Globale Betrachtungen in der ganzzahligen Optimierung. (German) [Global considerations in integer optimization]. Mathematische Operationsforschung und Statistik, 6(6):839–844, 1975. CODEN MTOSBI. ISSN 0047-6277.

Piehler:1976:SGG

[Pie76] Joachim Piehler. Über Schnittverschärfungen in der gemischtganzzahligen linearen Optimierung. (German) [On average tightening in mixed-integer linear optimization]. Mathematische Operationsforschung und Statistik, 7(3):333–342, 1976. CODEN MTOSBI. ISSN 0047-6277.

Pilz:1981:RBM

[Pil81] J. Pilz. Robust Bayes and minimax-Bayes estimation design in linear regression. Statistics (Berlin, DDR), 12(2):163–177, 1981. CODEN MOSSD5. ISSN 0323-3944.

Pilz:1986:NKM

[Pil86] J. Pilz. A note on Krafft’s maximin linear estimator for linear regression parameters. Statistics: a Journal of Theoretical and Applied Statistics, 17(1):9–14, 1986. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Pilz:1987:BR

[Pil87] J. Pilz. Book review. Statistics: a Journal of Theoretical and Applied Statistics, 18(2):258, 1987. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Pincus:1974:EPC

[Pin74] R. Pincus. Estimability of parameters of the covariance matrix and variance components. Mathematische Operationsforschung und Statistik, 5(3):245–248, 1974. CODEN MTOSBI. ISSN 0047-6277.

Pincus:1975:TLH

[Pin75] R. Pincus. Testing linear hypotheses under restricted alternatives. Mathematische Operationsforschung und Statistik, 6 (5):733–751, 1975. CODEN MTOSBI. ISSN 0047-6277.

Pincus:1977:TVC

[Pin77] R. Pincus. On tests in variance components models. Statistics (Berlin, DDR), 8(2):251–255, 1977. CODEN MOSSD5. ISSN 0323-3944.

Pincus:1989:BR

[Pin89] R. Pincus. Book review. Statistics: a Journal of Theoretical and Applied Statistics, 20(1):160, 1989. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Pewsey:2005:DBM

[PJ05] Arthur Pewsey and M. C. Jones. Discrimination between the von Mises and wrapped normal distributions: just how big does the sample size have to be? Statistics: a Journal of Theoretical and Applied Statistics, 39(2):81–89, 2005. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Park:2017:APV

[PJ17] Taeyoung Park and Seonghyun Jeong. Analysis of Poisson varying-coefficient models with autoregression. Statistics: a Journal of Theoretical and Applied Statistics, 52(1):34–49, 2017. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Patra:2018:ECH

[PK18] Lakshmi Kanta Patra and Somesh Kumar. Estimating the common hazard rate of two exponential distributions with ordered location parameters. Statistics: a Journal of Theoretical and Applied Statistics, 52(5):1040–1059, 2018. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Pleszczynska:1978:DSP

[PKM78] E. Pleszczyńska, A. Kowalski, and A. Matuszewski. Decision schemes of parameter identification in time series analysis. Statistics (Berlin, DDR), 9(1):73–83, 1978. CODEN MOSSD5. ISSN 0323-3944.

Postelnicu:1972:BFC

[PKN72] T. Postelnicu, Jena H. Klemm, and G. Nägler. Bericht über “The Fourth Conference on Probability Theory”, vom 12.– 18.9.1971 in Braşov, SR Rumanian. (German) [Report on the “The Fourth Conference on Probability Theory”, 12.– 18.9.1971 in Braşov, SR Rumania]. Mathematische Operationsforschung und Statistik, 3(2):157–160, 1972. CODEN MTOSBI. ISSN 0047-6277.

Patra:2021:CEO

[PKP21] Lakshmi Kanta Patra, Somesh Kumar, and Constantinos Petropoulos. Componentwise estimation of ordered scale parameters of two exponential distributions under a general class of loss function. Statistics: a Journal of Theoretical and Applied Statistics, 55(3):595–617, 2021. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Platen:1983:BR

[Pla83] E. Platen. Book reviews. Statistics (Berlin, DDR), 14(1):192, 1983. CODEN MOSSD5. ISSN 0323-3944.

Plachky:1994:SPA

[Pla94] D. Plachky. A statistical property of the arrangements of a finite set. Statistics: a Journal of Theoretical and Applied Statistics, 25(3):271–274, 1994. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Plachky:2002:ORS

[Pla02] D. Plachky. From optimality robustness to sufficiency and completeness. Statistics: a Journal of Theoretical and Applied Statistics, 36(2):139–146, 2002. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Perrone:2016:ODC

[PM16] E. Perrone and W. G. Müller. Optimal designs for copula models. Statistics: a Journal of Theoretical and Applied Statistics, 50(4):917–929, 2016. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Palma:2021:DCC

[PM21] Freddy Palma and Ramsés H. Mena. Duality for a class of continuous-time reversible Markov models. Statistics: a Journal of Theoretical and Applied Statistics, 55(1):231–242, 2021. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Pal:2017:ODO

[PMA17] Manisha Pal, N. K. Mandal, and M. L. Aggarwal. A-optimal designs for optimum mixture in an additive quadratic mixture model. Statistics: a Journal of Theoretical and Applied Statistics, 51(2):265–276, 2017. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Pathmanathan:2014:TSB

[PMO14] Dharini Pathmanathan, Rahul Mukerjee, and S. H. Ong. Two-sided Bayesian and frequentist tolerance intervals: general asymptotic results with applications. Statistics: a Journal of Theoretical and Applied Statistics, 48(3):524–538, 2014. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Pohlmann:1981:SFS

[PMS81] S. Pohlmann, J. Mecke, and D. Stoyan. Stereological formulas for stationary surface processes. Statistics (Berlin, DDR), 12 (3):429–440, 1981. CODEN MOSSD5. ISSN 0323-3944.

Pham:1990:SCM

[PN90] Tuain D. Pham and Htjng T. Nguyen. Strong consistency of the maximum likelihood estimators in the change-point hazard rate model. Statistics: a Journal of Theoretical and Applied Statistics, 21(2):203–216, 1990. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Poursaeed:2010:SAM

[PN10] M. H. Poursaeed and A. R. Nematollahi. Some aspects of the mean past lifetime of a parallel system under double regularly checking. Statistics: a Journal of Theoretical and Applied Statistics, 44(5):505–515, 2010. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Pogany:2013:CNR

[PN13] Tibor K. Pogány and Saralees Nadarajah. On the convolution of normal and t random variables. Statistics: a Journal of Theoretical and Applied Statistics, 47(6):1363–1369, 2013. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Pincus:1978:BR

[PNK78] R. Pincus, V. Nollau, and D. König. Book review. Statistics (Berlin, DDR), 9(1):159–164, 1978. CODEN MOSSD5. ISSN 0323-3944.

Pogany:2015:EEP

[Pog15] Tibor K. Pogány. The exponentiated exponential Poisson distribution revisited. Statistics: a Journal of Theoretical and Applied Statistics, 49(4):918–929, 2015. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Polzehl:1985:UJN

[Pol85] Jörg Polzehl. Using jackknife in nonlinear models — confidence regions for a function of the structural parameter. Statistics: a Journal of Theoretical and Applied Statistics, 16(1):139–149, 1985. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Pollett:1986:SPA

[Pol86] P. K. Pollett. Some Poisson approximations for departure processes in general queueing networks. Statistics: a Journal of Theoretical and Applied Statistics, 17(3):393–405, 1986. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Polsehl:1987:BR

[Pol87a] J. Polsehl. Book review. Statistics: a Journal of Theoretical and Applied Statistics, 18(1):16, 1987. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Polzehl:1987:BR

[Pol87b] J. Polzehl. Book review. Statistics: a Journal of Theoretical and Applied Statistics, 18(1):36, 1987. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Polymenis:2008:NVT

[Pol08] Athanase Polymenis. A note on a validity test using the stochastic EM algorithm in order to assess the number of components in a finite mixture model. Statistics: a Journal of Theoretical and Applied Statistics, 42(3):261–274, 2008. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Ponnapalli:1984:ASP

[Pon84] P. Ponnapalli. Asymptotic Studentization of parametric and rank order tests for location. Statistics (Berlin, DDR), 15(2): 277–287, 1984. CODEN MOSSD5. ISSN 0323-3944.

Pons:1986:VCE

[Pon86] O. Pons. Vitesse de convergence des estimateurs à noyau pour l’intensité d’un processus ponctuel. (French) [Speed of convergence of kernel estimators for the intensity of a point process]. Statistics: a Journal of Theoretical and Applied Statistics, 17(4):577–584, 1986. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Pons:1993:HTD

[Pon93] Odile Pons. Homogeneity test in distribution mixtures and application to major genes detection. Statistics: a Journal of Theoretical and Applied Statistics, 24(2):149–160, 1993.

CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Pons:2002:ECR

[Pon02] Odile Pons. Estimation in a Cox regression model with a change-point at an unknown time. Statistics: a Journal of Theoretical and Applied Statistics, 36(2):101–124, 2002. CO-DEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Pons:2004:EMS

[Pon04] Odile Pons. Estimation in a model for a semi-Markov process with covariates under right-censoring. Statistics: a Journal of Theoretical and Applied Statistics, 38(4):273–293, 2004. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Pons:2006:ESM

[Pon06] Odile Pons. Estimation for semi-Markov models with partial observations via self-consistency equations. Statistics: a Journal of Theoretical and Applied Statistics, 40(5):377–388, 2006. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Padmanabhan:1983:TNS

[PP83] A. R. Padmanabhan and Madan L. Puri. Theory of nonparametric statistics for rounded-off data with applications. Statistics (Berlin, DDR), 14(2):301–349, 1983. CODEN MOSSD5. ISSN 0323-3944.

Picard:1992:PEW

[PP92] Dominique Picard and Bernard Prum. Parameter estimation when various models are available. Statistics: a Journal of Theoretical and Applied Statistics, 23(3):237–255, 1992. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Potscher:1994:FUL

[PP94] Benedikt M. Pötscher and Ingmar R. Prucha. On the formulation of uniform laws of large numbers: a truncation approach. Statistics: a Journal of Theoretical and Applied Statistics, 25(4):343–360, 1994. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Pardo:2007:SSN

[PPD07] Julio Angel Pardo, Leandro Pardo, and María Del Carmen Pardo. A simulation study of a nested sequence of binomial regression models. Statistics: a Journal of Theoretical and Applied Statistics, 41(3):253–267, 2007. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Potscher:1989:DCD

[PR89] B. M. Pötscher and E. Reschenhofer. Distribution of the Coates–Diggle test statistic in case of replicated observations. Statistics: a Journal of Theoretical and Applied Statistics, 20(3):417–421, 1989. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Pandhare:2022:RDL

[PR22] S. C. Pandhare and T. V. Ramanathan. The robust desparsified lasso and the focused information criterion for highdimensional generalized linear models. Statistics: a Journal of Theoretical and Applied Statistics, 57(1):1–25, 2022. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/02331888.2022.2154769.

Prakasa:1986:WCL

[Pra86] B. L. S. Prakasa. Weak convergence of lease squares process in the smooth case. Statistics: a Journal of Theoretical and Applied Statistics, 17(4):505–516, 1986. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Prekopa:1972:CSP

[Pré72] András Prékopa. A class of stochastic programming decision problems. Mathematische Operationsforschung und Statistik, 3(5):349–354, 1972. CODEN MTOSBI. ISSN 0047-6277.

Prekopa:1973:GTS

[Pré73] András Prékopa. Generalizations of the theorems of Smirnov with application to a reliability type inventory problem. Mathematische Operationsforschung und Statistik, 4(4):283–297, 1973. CODEN MTOSBI. ISSN 0047-6277.

Prekopa:1974:ESM

[Pré74a] András Prékopa. Eine Erweiterung der sogenannten Methode der zulässigen Richtungen der nichtlinearen Optimierung auf den Fall quasikonkaver Restriktionsfunktionen. (German) [An extension of the so-called method of allowable directions of nonlinear optimization to the case of quasiconcave restriction functions]. Mathematische Operationsforschung und Statistik, 5(4–5):281–293, 1974. CODEN MTOSBI. ISSN 0047-6277.

Prekopa:1974:PUP

[Pré74b] András Prékopa. Programming under probabilistic constraints with a random technology matrix. Mathematische Operationsforschung und Statistik, 5(2):109–116, 1974. CO-DEN MTOSBI. ISSN 0047-6277.

Provost:1989:SIG

[Pro89] Serge B. Provost. On sums of independent gamma random variables. Statistics: a Journal of Theoretical and Applied Statistics, 20(4):583–591, 1989. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Pena:1995:IGT

[PRS95] Edsel A. Peña, Vijay K. Rohatgi, and Gábor J. Székely. Inference for a general Type II censorship model. Statistics: a Journal of Theoretical and Applied Statistics, 26(3):241–252, 1995. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Pruscha:1993:CTS

[Pru93] Helmut Pruscha. Categorical time semes with a recursive scheme and with covariates. Statistics: a Journal of Theoretical and Applied Statistics, 24(1):43–57, 1993. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Plachky:1977:GRC

[PS77] D. Plachky and J. Steinebach. A generalization of a result of Chernoff in large sample theory. Statistics (Berlin, DDR), 8 (3):375–379, 1977. CODEN MOSSD5. ISSN 0323-3944.

Pincus:1984:NBP

[PS84] R. Pincus and W. H. Schmidt. A note on the ban property of the residual sum of squares in regression models. Statistics (Berlin, DDR), 15(1):37–42, 1984. CODEN MOSSD5. ISSN 0323-3944.

Petsa:2010:AQF

[PS10]  Athanasia Petsa and Theofanis Sapatinas.  Adaptive

quadratic functional estimation of a weighted density by model selection. Statistics: a Journal of Theoretical and Applied Statistics, 44(6):571–585, 2010. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Pal:2012:NUQ

[PS12] Sanghamitra Pal and Sarjinder Singh. A new unrelated question randomized response model. Statistics: a Journal of Theoretical and Applied Statistics, 46(1):99–109, 2012. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Park:2014:KLI

[PS14] Sangun Park and Minsuk Shin. Kullback–Leibler information of a censored variable and its applications. Statistics: a Journal of Theoretical and Applied Statistics, 48(4):756–765, 2014. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Peterlin:2023:PAG

[PSB23] Jakob Peterlin, Janez Stare, and Rok Blagus. A permutation approach to goodness-of-fit testing in regression models. Statistics: a Journal of Theoretical and Applied Statistics, 57(1):123–149, 2023. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/02331888.2023.2172173.

Pal:1995:EMN

[PSCC95] Nabendu Pal, Bikas K. Sinha, Gopal Chaudhuri, and Ching-Hui Chang. Estimation of a multivariate normal mean vector and local improvements. Statistics: a Journal of Theoretical and Applied Statistics, 26(1):1–17, 1995. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Polzehl:1994:BR

[PSS^{+}94] J. Polzehl, D. Stoyan, U. Schulze, P. Neumann, G. Siegel, Eckhard Platen, H.-J. Girlich, and H. J. Engelbert. Book review. Statistics: a Journal of Theoretical and Applied Statistics, 25(3):275–283, 1994. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Pieper:1983:ZGS

[PT83] V. Pieper and J. Tiedge. Zuverlässigkeitsmodelle auf der Grundlage stochastischer Modelle von Verschleiß prozessen. (German) [Processes of reliability models based on stochastic models of wear]. Statistics (Berlin, DDR), 14(3):485–502, 1983. CODEN MOSSD5. ISSN 0323-3944.

Pukelsheim:1977:HMR

[Puk77] Friedrich Pukelsheim. On Hsu’s model in regression analysis. Statistics (Berlin, DDR), 8(3):323–331, 1977. CODEN MOSSD5. ISSN 0323-3944.

Pukelsheim:1981:LMC

[Puk81a] Friedrich Pukelsheim. Linear models and convex geometry: aspects of non-negative variance estimation. Statistics (Berlin, DDR), 12(2):271–286, 1981. CODEN MOSSD5. ISSN 0323-3944.

Pukelsheim:1981:ODM

[Puk81b] Friedrich Pukelsheim. On c-optimal design measures. Statistics (Berlin, DDR), 12(1):13–20, 1981. CODEN MOSSD5. ISSN 0323-3944.

Puri:1985:CFR

[PV85] Madan L. Puri and I. Vincze. On the Cramér–Frechet–Rao inequality for translation parameter in the case of finite support. Statistics: a Journal of Theoretical and Applied Statistics, 16(4):495–506, 1985. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Pap:1998:BEB

[PV98] Gyula Pap and Martien C. A. Van Zuijlen. A Berry–Esseen bound for linear combinations of order statistics. Statistics: a Journal of Theoretical and Applied Statistics, 31(2):151–186, 1998. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Pavlenko:2001:EDD

[PvR01] Tatjana Pavlenko and Dietrich von Rosen. Effect of dimensionality on discrimination. Statistics: a Journal of Theoretical and Applied Statistics, 35(3):191–213, 2001. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Petersohn:1974:GAS

[PVW74] U. Petersohn, K. Voss, and K. H. Weber. Genetische Adaptation — ein stochastisches Suchverfahren für diskrete Optimierungsprobleme. (German) [Genetic adaptation — a stochastic search method for discrete optimization problems]. Mathematische Operationsforschung und Statistik, 5 (7–8):555–571, 1974. CODEN MTOSBI. ISSN 0047-6277.

Petrich:1976:ALO

[PW76] F. Petrich and F. Wagner. Ein Algorithmus für das Lineare 0–1-Optimierungsproblem. (German) [An algorithm for the linear 0–1 optimization problem]. Mathematische Operationsforschung und Statistik, 7(3):351–359, 1976. CODEN MTOSBI. ISSN 0047-6277.

Pukelsheim:1995:D

[PW95] F. Pukelsheim and A. Wilhelm. Discussion. Statistics: a Journal of Theoretical and Applied Statistics, 26(2):168–172, 1995. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Pingel:2015:ECC

[PW15] Ronnie Pingel and Ingeborg Waernbaum. Effects of correlated covariates on the asymptotic efficiency of matching and inverse probability weighting estimators for causal inference. Statistics: a Journal of Theoretical and Applied Statistics, 49(4):795–814, 2015. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Pang:2001:FRM

[PYHT01] W. K. Pang, Z. H. Yang, S. H. Hou, and M. D. Troutt. Further results on multivariate vertical density representation and an application to random vector generation. Statistics: a Journal of Theoretical and Applied Statistics, 35(4):463–477, 2001. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Zang:2015:ERC

[pZ15] Qing pei Zang. Exact rates in complete moment convergence of self-normalized sums for multidimensionally indexed random variables. Statistics: a Journal of Theoretical and Applied Statistics, 49(3):537–548, 2015. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Qin:2013:LBC

[QCO13] Hong Qin, Kashinath Chatterjee, and Zujun Ou. A lower bound for the centred L_{2}-discrepancy on combined designs under the asymmetric factorials. Statistics: a Journal of Theoretical and Applied Statistics, 47(5):992–1002, 2013. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Qingguo:2015:ESF

[Qin15] Tang Qingguo. Estimation for semi-functional linear regression. Statistics: a Journal of Theoretical and Applied Statistics, 49(6):1262–1278, 2015. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Qingguo:2008:RCE

[QJ08] Tang Qingguo and Wang Jinde. Reducing component estimation for varying coefficient models. Statistics: a Journal of Theoretical and Applied Statistics, 42(1):25–48, 2008. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Qingguo:2017:QRF

[QK17] Tang Qingguo and Linglong Kong. Quantile regression in functional linear semiparametric model. Statistics: a Journal of Theoretical and Applied Statistics, 51(6):1342–1358, 2017. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Qin:2017:SCR

[QLT17] Ruibing Qin, Weiqi Liu, and Zheng Tian. A strong convergence rate of estimator of variance change in linear processes and its applications. Statistics: a Journal of Theoretical and Applied Statistics, 51(2):314–330, 2017. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Qin:2021:WRT

[QLY21] Ruibing Qin, Xinxin Luo, and Wei Yang. Wilcoxon rank test for change in persistence. Statistics: a Journal of Theoretical and Applied Statistics, 55(3):514–531, 2021. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Quenneville:1986:BPT

[Que86] Benoit Quenneville. Bootstrap procedure for testing linear hypotheses without normality. Statistics: a Journal of Theoretical and Applied Statistics, 17(4):533–538, 1986. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Qi:2023:IHQ

[QXY23] Yongcheng Qi, Mengzi Xie, and Jingping Yang. Inference of high quantiles of a heavy-tailed distribution from block data. Statistics: a Journal of Theoretical and Applied Statistics, 57(4):918–940, 2023. CO-DEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/02331888.2023.2228442.

Raqab:2001:NCB

[RA01] Mohammad Z. Raqab and Adnan M. Awad. A note on characterization based on Shannon entropy of record statistics. Statistics: a Journal of Theoretical and Applied Statistics, 35(4):411–413, 2001. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

RamirezCid:2002:SRC

[RA02] Juan Esteban Ramírez Cid and Jorge Alberto Achcar. Software reliability considering the superposition of nonhomogeneous Poisson processes in the presence of a covariate. Statistics: a Journal of Theoretical and Applied Statistics, 36(3):259–269, 2002. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Raqab:2006:GSH

[RA06] Mohammad Z. Raqab and Mohammad Ahsanullah. Generalized secant hyperbolic order statistics and associated inferences. Statistics: a Journal of Theoretical and Applied Statistics, 40(4):325–338, 2006. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Raqab:2010:SRM

[RA10] Mohammad Z. Raqab and Majid Asadi. Some results on the mean residual waiting time of records. Statistics: a Journal of Theoretical and Applied Statistics, 44(5):493–504, 2010. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Rabier:2015:SPQ

[Rab15] Charles-Elie Rabier. On stochastic processes for quantitative trait locus mapping under selective genotyping. Statistics: a Journal of Theoretical and Applied Statistics, 49(1):19–34, 2015. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Raqab:2007:SIB

[RAD07] Mohammad Z. Raqab, J. Ahmadi, and M. Doostparast. Statistical inference based on record data from Pareto model. Statistics: a Journal of Theoretical and Applied Statistics, 41(2):105–118, 2007. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Rafajlowicz:1988:NLS

[Raf88] Ewaryst Rafajlowicz. Nonparametric least squares estimation of a regression function. Statistics: a Journal of Theoretical and Applied Statistics, 19(3):349–358, 1988. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Rodrigues:1993:BAA

[RALN93] Josemar Rodrigues, Jorge Alberto Achcar, and Francisco Louzada-Neto. A Bayesian analysis of the accelerated life tests via the orthogonal parameters. Statistics: a Journal of Theoretical and Applied Statistics, 24(4):353–357, 1993. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Ramos:2017:BAG

[RAM^{+}17] Pedro L. Ramos, Jorge A. Achcar, Fernando A. Moala, Eduardo Ramos, and Francisco Louzada. Bayesian analysis of the generalized gamma distribution using non-informative priors. Statistics: a Journal of Theoretical and Applied Statistics, 51(4):824–843, 2017. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Rao:1983:ATN

[Rao83] B. L. S. Prakasa Rao. Asymptotic theory for non-linear least squares estimator for diffusion processes. Statistics (Berlin, DDR), 14(2):195–209, 1983. CODEN MOSSD5. ISSN 0323-

Rao:1985:EDD

[Rao85] B. L. S. Prakasa Rao. Estimation of the drift for diffusion process. Statistics: a Journal of Theoretical and Applied Statistics, 16(2):263–275, 1985. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Rao:1998:HIM

[Rao98] B. L. S. Prakasa Rao. Hoeffding identity, multivariance and multicorrelation. Statistics: a Journal of Theoretical and Applied Statistics, 32(1):13–29, 1998. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Raqab:2004:GED

[Raq04] Mohammad Z. Raqab. Generalized exponential distribution: Moments of order statistics. Statistics: a Journal of Theoretical and Applied Statistics, 38(1):29–41, 2004. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Rasch:1985:FSB

[Ras85] Dieter Rasch. Finite sample behaviour of an asymptotic ttest in exponential regression. Statistics: a Journal of Theoretical and Applied Statistics, 16(1):121–124, 1985. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Rasch:1988:RRG

[Ras88] Dieter Rasch. Recent results in growth analysis. Statistics: a Journal of Theoretical and Applied Statistics, 19(4):585–604, 1988. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Rasch:1991:MMS

[Ras91] Dieter Rasch. Mathematical models in selection — an introduction. Statistics: a Journal of Theoretical and Applied Statistics, 22(1):155–160, 1991. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Rasekh:2006:EPB

[Ras06] A. R. Rasekh. Empirical process based on the recursive residuals in functional measurement error models. Statistics: a Journal of Theoretical and Applied Statistics, 40(3):247–258, 2006. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Rataj:1993:RDE

[Rat93] Jan Rataj. Random distances and edge correction. Statistics: a Journal of Theoretical and Applied Statistics, 24(4):377– 385, 1993. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Rossberg:1980:UTC

[RD80] H. J. Rossberg and Chu Duc. Uniqueness theorems for the components of a convolution given on a half line. Statistics (Berlin, DDR), 11(3):361–371, 1980. CODEN MOSSD5. ISSN 0323-3944.

Rabier:2021:SCS

[RD21] Charles-Elie Rabier and Céline Delmas. The SgenoLasso and its cousins for selective genotyping and extreme sampling: application to association studies and genomic selection. Statistics: a Journal of Theoretical and Applied Statistics, 55(1): 18–44, 2021. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Rodriguez-Diaz:2003:BCO

[RDLF03] J. M. Rodríguez-Díaz and J. López-Fidalgo. A bidimensional class of optimality criteria involving ϕ_{p} and characteristic criteria. Statistics: a Journal of Theoretical and Applied Statistics, 37(4):325–334, 2003. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Reimnitz:1982:ANA

[Rei82] P. Reimnitz. Asymptotic near admissibility and asymptotic near optimality by the “two armed bandit” problem. Statistics (Berlin, DDR), 13(2):245–263, 1982. CODEN MOSSD5. ISSN 0323-3944.

Reimnitz:1984:NSA

[Rei84] P. Reimnitz. A new stochastic approximation procedure for obtaining the maximum of an ’unknown regression function. Statistics (Berlin, DDR), 15(2):219–238, 1984. CODEN MOSSD5. ISSN 0323-3944.

Reimnitz:1986:MBS

[Rei86] Peter Reimnitz. Minimax and Bates strategies for the discounted infinite horizon one-armed-bandit. Explicit formulae and structural properties. Statistics: a Journal of Theoretical and Applied Statistics, 17(2):249–260, 1986. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Revesz:1977:HAM

[Rév77] P. Révész. How to apply the method of stochastic approximation in the non-parametric estimation of a regression function. Statistics (Berlin, DDR), 8(1):119–126, 1977. CODEN MOSSD5. ISSN 0323-3944.

Reiser:1994:CB

[RF94] Benjamin Reiser and David Faraggi. Confidence bounds for P r(a′x > b′y). Statistics: a Journal of Theoretical and Applied Statistics, 25(2):107–111, 1994. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Rasekh:2003:IFF

[RF03] A. R. Rasekh and N. R. J. Fieller. Influence functions in functional measurement error models with replicated data. Statistics: a Journal of Theoretical and Applied Statistics, 37(2):169–178, 2003. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Roy:2018:TBE

[RFK18] A. Roy, K. Filipiak, and D. Klein. Testing a block exchangeable covariance matrix. Statistics: a Journal of Theoretical and Applied Statistics, 52(2):393–408, 2018. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Rao:1984:TCP

[RG84] K. S. Madhava Rao and A. P. Gore. Testing concurrence and parallelism of several sample regressions against ordered alternatives. Statistics (Berlin, DDR), 15(1):43–50, 1984. CODEN MOSSD5. ISSN 0323-3944.

Roy:2009:ERP

[RG09] Sugata Sen Roy and Sibnarayan Guria. Estimation of regression parameters in the presence of outliers in the response. Statistics: a Journal of Theoretical and Applied Statistics, 43(6):531–539, 2009. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Reyes:2013:MSD

[RGB13] Jimmy Reyes, Héctor W. Gómez, and Heleno Bolfarine. Modified slash distribution. Statistics: a Journal of Theoretical and Applied Statistics, 47(5):929–941, 2013. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Rasch:1980:EDO

[RH80] D. Rasch and G. Herrendörfer. Experimental design and optimal decision ii estimation of the slope of linear regression. Statistics (Berlin, DDR), 11(1):125–136, 1980. CODEN MOSSD5. ISSN 0323-3944.

Ruckdeschel:2013:ORE

[RH13] Peter Ruckdeschel and Nataliya Horbenko. Optimally robust estimators in generalized Pareto models. Statistics: a Journal of Theoretical and Applied Statistics, 47(4):762–791, 2013. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Roozbeh:2020:USR

[RH20] Mahdi Roozbeh and Nor Aishah Hamzah. Uncertain stochastic ridge estimation in partially linear regression models with elliptically distributed errors. Statistics: a Journal of Theoretical and Applied Statistics, 54(3):494–523, 2020. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Richter:1975:DDM

[Ric75] Knut Richter. Deterministische dynamische Mehrproduktmodelle der Produktionsplanung und Lagerhaltung. (German) [Deterministic dynamic models of multi-product production planning and inventory]. Mathematische Operationsforschung und Statistik, 6(3):385–395, 1975. CODEN MTOSBI. ISSN 0047-6277.

Richter:1976:ALG

[Ric76a] Diete Richter. Ein Algorithmus zur Lösung gemischtganzzahliger, konvexer Optimierungsprobleme. (German) [An algorithm for solving mixed-integer convex optimization problems]. Mathematische Operationsforschung und Statistik, 7 (2):197–208, 1976. CODEN MTOSBI. ISSN 0047-6277.

Richter:1976:BGO

[Ric76b] Dieter Richter. Zur Behandlung ganzzahliger Optimierungsaufgaben ohne Ganzzahligkeitsbedingung an die Zielfunktion. (German) [On the treatment of integer optimization problems without integer conditions on the objective function]. Mathematische Operationsforschung und Statistik, 7(3):343– 349, 1976. CODEN MTOSBI. ISSN 0047-6277.

Richter:1976:RDO

[Ric76c] Knut Richter. Ressourcenaufspaltung für diskrete Optimierungsaufgaben. Teil I. (German) [Resource splitting for discrete optimization tasks. Part I]. Mathematische Operationsforschung und Statistik, 7(6):831–842, 1976. CODEN MTOSBI. ISSN 0047-6277.

Richter:1984:BVQ

[Ric84] M. Richter. Berechnung der Verteilungsfunktion quadratischer formen von normalverteilen Zufallsgroßen and deren Anwendung in der Statistik. (German) [Calculation of the distribution function of quadratic forms of normally distributed random variables and their application in statistics]. Statistics (Berlin, DDR), 15(2):177–194, 1984. CODEN MOSSD5. ISSN 0323-3944.

Richter:1986:DGS

[Ric86] M. Richter. Dominierbarkeit Gauss’scher statistischer Strukturen in Hilberträumen. (German) [Dominatability of Gaussian statistical structures in Hilbert spaces]. Statistics: a Journal of Theoretical and Applied Statistics, 17(3):377–391, 1986. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Richter:1989:BR

[Ric89] M. Richter. Book review. Statistics: a Journal of Theoretical and Applied Statistics, 20(1):124, 1989. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Rieder:1987:CGR

[Rie87] Helmut Rieder. Contamination games in a robust k-sample model. Statistics: a Journal of Theoretical and Applied Statistics, 18(4):527–562, 1987. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Riedel:1989:BRLa

[Rie89a] Manfred Riedel. On the bias-robustness in the location model I. Statistics: a Journal of Theoretical and Applied Statistics, 20(2):223–233, 1989. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Riedel:1989:BRLb

[Rie89b] Manfred Riedel. On the bias-robustness in the location model II. Statistics: a Journal of Theoretical and Applied Statistics, 20(2):235–246, 1989. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Rieder:1989:FSM

[Rie89c] Helmut Rieder. A finite-sample minimax regression estimator. Statistics: a Journal of Theoretical and Applied Statistics, 20(2):211–221, 1989. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Riedel:1990:BR

[Rie90] M. Riedel. Book reviews. Statistics: a Journal of Theoretical and Applied Statistics, 21(4):642, 1990. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Randt:1983:CGG

[RL83] A. Randt and B. Lisek. On the continuity of G/GI/m queues. Statistics (Berlin, DDR), 14(4):577–587, 1983. CO-DEN MOSSD5. ISSN 0323-3944.

Rodrigues:1998:NBA

[RL98] Josemar Rodrigues and José Galvâo Leite. A note on Bayesian analysis in M/M/1 queues derived from confidence intervals. Statistics: a Journal of Theoretical and Applied Statistics, 31 (1):35–42, 1998. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Rocha:2013:IFF

[RLAV13] Gustavo H. M. A. Rocha, Rosangela H. Loschi, and Reinaldo B. Arellano-Valle. Inference in flexible families of distributions with normal kernel. Statistics: a Journal of Theoretical and Applied Statistics, 47(6):1184–1206, 2013. CO-DEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Rocha:2016:BMM

[RLAV16] Gustavo H. M. A. Rocha, Rosangela H. Loschi, and Reinaldo B. Arellano-Valle. Bayesian mismeasurement tmodels for censored responses. Statistics: a Journal of Theoretical and Applied Statistics, 50(4):841–869, 2016. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Ruiz-Medina:2016:FAV

[RM16] M. D. Ruiz-Medina. Functional analysis of variance for Hilbert-valued multivariate fixed effect models. Statistics: a Journal of Theoretical and Applied Statistics, 50(3):689–715, 2016. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Ravi:2017:TBU

[RM17] Sreenivasan Ravi and Mandagere Chandrashekhar Manohar. On tail behaviour of k-th upper order statistics under fixed and random sample sizes via tail equivalence. Statistics: a Journal of Theoretical and Applied Statistics, 52(1):156–176, 2017. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Riedel:1988:BR

[RMGR88] M. Riedel, K. Matthes, H.-J. Girlich, and D. Rasch. Book reviews. Statistics: a Journal of Theoretical and Applied Statistics, 19(3):488–490, 1988. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Rasethuntsa:2018:SSR

[RN18] Tau Raphael Rasethuntsa and Mustafa Nadar. Stressstrength reliability of a non-identical-component-strengths system based on upper record values from the family of Kumaraswamy generalized distributions. Statistics: a Journal of Theoretical and Applied Statistics, 52(3):684–716, 2018. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Reidel:1992:BR

[RND^{+}92] M. Reidel, W. Näther, B. Droge, H. Liero, J. Andël, J. Andël, J. Tiedgel, Marie Husková, J. Polzehl, Werner Plischke, H. Stenger, J. P. Vila, G. Wittwer, W.-D. Richter, Ludwig Fahrmeir, M. Husková, D. Stoyan, F. Liese, M. Riedel, M. H.

Neumannl, M. H. Neumannl, L. Heinrich, and Günter Last. Book review. Statistics: a Journal of Theoretical and Applied Statistics, 23(3):287–304, 1992. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Rezaei:2013:NTP

[RNT13] Sadegh Rezaei, Saralees Nadarajah, and Nahid Tahghighnia. A new three-parameter lifetime distribution. Statistics: a Journal of Theoretical and Applied Statistics, 47(4):835–860, 2013. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Ramesh:1998:SIB

[RO98] C. Gupta Ramesh and Akman Olcay. Statistical inference based on the length-biased data for the inverse Gaussian distribution. Statistics: a Journal of Theoretical and Applied Statistics, 31(4):325–337, 1998. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Rodel:1970:AZW

[Röd70] Egmar Rödel. Abhängigkeitsmaße für Zufallsgrößen mit Werten in einem separablen Hilbertraum. (German) [Measures of dependence for random variables with values in a separable Hilbert space]. Mathematische Operationsforschung und Statistik, 1(2):117–139, 1970. CODEN MTOSBI. ISSN 0047-6277.

Rodel:1971:ZVE

[Röd71] Egmar Rödel. Zweidimensionale Verteilungen mit endlicher Kontingenz und die asymptotische Effizienz von Unabhängigkeitstests. (German) [Two-dimensional distributions with finite contingency and the asymptotic efficiency of tests of independence]. Mathematische Operationsforschung und Statistik, 2(4):291– 309, 1971. CODEN MTOSBI. ISSN 0047-6277.

Rodel:1976:TGG

[Röd76] Egmar Rödel. Tests für den Gestaltsparameter der Gammaverteilung. (German) [Tests for the shape parameter of the gamma distribution]. Mathematische Operationsforschung und Statistik, 7(1):107–111, 1976. CODEN MTOSBI. ISSN 0047-6277.

Rodel:1977:BET

[Röd77] Egmar Rödel. Bivariate exponential-type distributions with linear regression. Statistics (Berlin, DDR), 8(3):387–397, 1977. CODEN MOSSD5. ISSN 0323-3944.

Rodel:1987:NCP

[Röd87a] Egmar Rödel. A necessary condition for positive dependence. Statistics: a Journal of Theoretical and Applied Statistics, 18(3):351–359, 1987. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Rodel:1987:ENB

[Röd87b] Egmar Rödel. R-estimation of normed bivariate density functions. Statistics: a Journal of Theoretical and Applied Statistics, 18(4):573–585, 1987. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Rodriguez:1988:AUR

[Rod88] Alberto Fuentes Rodriguez. Admissibility and unbiasedness of the ridge classification rules for two normal populations with equal covariance matrices. Statistics: a Journal of Theoretical and Applied Statistics, 19(3):383–388, 1988. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Rodel:1989:LBS

[Rod89] Egmar Rodel. Linear bank statistics with estimated scores for testing independence. Statistics: a Journal of Theoretical and Applied Statistics, 20(3):423–438, 1989. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Rodrigues:1993:BAG

[Rod93] Josemar Rodrigues. A Bayesian analysis of the generalized least-squares procedure for functional relationships. Statistics: a Journal of Theoretical and Applied Statistics, 24(4): 347–351, 1993. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Rodel:2004:EOT

[Röd04] Egmar Rödel. R-estimation of optimal transformations via copulas and ACE. Statistics: a Journal of Theoretical and Applied Statistics, 38(2):101–116, 2004. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Rodriguez:2015:LMO

[ROM15] C. Rodríguez, I. Ortiz, and I. Martínez. Locally and maximin optimal designs for multi-factor nonlinear models. Statistics: a Journal of Theoretical and Applied Statistics, 49(5):1157– 1168, 2015. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Ronning:1976:BGS

[Ron76] Gerd Ronning. Bemerkungen zum Gebrauch von Saison-Variablen bei der Schätzung von Distributed-Lag-Modellen unter polynomialen Restriktionen. (German) [Remarks on the use of seasonal variables in the estimation of distributed lag models under polynomial restrictions]. Mathematische Operationsforschung und Statistik, 7(5):755–772, 1976. CODEN MTOSBI. ISSN 0047-6277.

Rosseberg:1971:ABW

[Ros71] Hans-Joachim Rosseberg. Ausbeutung einer bekannten Wiener–Hopf-Faktorisierung beim Wartemodell G/G/1 und einer mit ihm zusammenhängenden Irrfahrt. (German) [Exploitation of a known Wiener–Hopf factorization in the queueing model G/G/1 and a continuous random walk with it]. Mathematische Operationsforschung und Statistik, 2(2):129– 146, 1971. CODEN MTOSBI. ISSN 0047-6277.

Rossberg:1972:CEP

[Ros72] H. J. Rossberg. Characterization of the exponential and the Pareto distributions by means of some properties of the distributions which the differences and quotients of order statistics are subject to. Mathematische Operationsforschung und Statistik, 3(3):207–216, 1972. CODEN MTOSBI. ISSN 0047-6277.

Rosenberg:1975:CCP

[Ros75] I. G. Rosenberg. On Chvátal’s cutting planes in integer linear programming. Mathematische Operationsforschung und Statistik, 6(4):511–522, 1975. CODEN MTOSBI. ISSN 0047-6277.

Ross:1982:NLM

[Ros82] G. J. S. Ross. Non-linear models. Statistics (Berlin, DDR), 13(3):445–453, 1982. CODEN MOSSD5. ISSN 0323-3944.

Rossberg:1989:BEI

[Ros89] Hans-Joachim Rossberg. The Berry–Esséen inequalities derived from restricted convergence. Statistics: a Journal of Theoretical and Applied Statistics, 20(3):449–451, 1989. CO-DEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Rossberg:1991:CSN

[Ros91] Hans-Joachm Rossberg. Characterization of the standard normal probability density within the class of self-adjoint (selfreciprocal) densities. Statistics: a Journal of Theoretical and Applied Statistics, 22(4):651–654, 1991. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Royen:1987:AMN

[Roy87] Thomas Royen. an approximation for multivariate normal probabilities of rectangular regions. Statistics: a Journal of Theoretical and Applied Statistics, 18(3):389–400, 1987. CO-DEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Ristic:2018:DAB

[RPZB18] Miroslav M. Ristić, Bozidar V. Popović, Konstantinos Zografos, and Narayanaswamy Balakrishnan. Discrimination among bivariate beta-generated distributions. Statistics: a Journal of Theoretical and Applied Statistics, 52(2):303–320, 2018. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Ruschendorf:1991:SOLb

[RR91] Ludger Rüschendorf and M. Riedel. Stochastic ordering of likelihood ratios and partial sufficiency. Statistics: a Journal of Theoretical and Applied Statistics, 22(4):551–558, 1991. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Rossberg:2002:LFT

[RR02] Hans-Joachim Rossberg and Manfred Riedel. The Lindeberg– Feller theorem for positive definite densities. Statistics: a Journal of Theoretical and Applied Statistics, 36(4):365–368, 2002. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Raqab:2004:SBE

[RR04] Mohammad Z. Raqab and Tomasz Rychlik. Sharp bounds on the expectations of second record values from symmetric populations. Statistics: a Journal of Theoretical and Applied Statistics, 38(1):49–57, 2004. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Richter:1993:BR

[RRL^{+}93] W.-D. Richter, M. Richter, H. Liero, K. Mai, L. Heinrich, M. Zähle, J. Franz, G. Enderlien, M. Müller, S. v. Weber, Egmar Rödel, and M. Riedel. Book review. Statistics: a Journal of Theoretical and Applied Statistics, 25(1):89–96, 1993. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Rossberg:1974:BKI

[RS74] H. J. Rossberg and G. Siege. Die Bedeutung von Kingmans Integralungleichungen bei der Approximation der stationären Wartezeitverteilung im Modell GI/G/1 mit und ohne Verzögerung beim Beginn einer Beschäftigungsperiode. (German) [The meaning of Kingman’s integral inequalities in the approximation of stationary waiting time distribution in the model GI/G/1 with and without delay when starting an employment period]. Mathematische Operationsforschung und Statistik, 5(9):687–699, 1974. CODEN MTOSBI. ISSN 0047-6277.

Rao:1996:D

[RS96] J. N. K. Rao and R. R. Sitter. Discussion. Statistics: a Journal of Theoretical and Applied Statistics, 27(3–4):246– 247, 1996. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Rafajlowicz:2009:BCC

[RS09] Ewaryst Rafajlowicz and Ansgar Steland. A binary control chart to detect small jumps. Statistics: a Journal of Theoretical and Applied Statistics, 43(3):295–311, 2009. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Rao:1985:ESF

[RT85] Jammalamadaka S. Rao and R. C. Tiwari. Estimation of survival functions and failure rate. Statistics: a Journal of Theoretical and Applied Statistics, 16(4):535–540, 1985. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Redenbach:2013:SOC

[RT13] Claudia Redenbach and Christoph Thäle. Second-order comparison of three fundamental tessellation models. Statistics: a Journal of Theoretical and Applied Statistics, 47(2):237–257, 2013. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Rublik:2021:JST

[Rub21] Frantisek Rublík. On jackknifing the symmetrized Tyler matrix. Statistics: a Journal of Theoretical and Applied Statistics, 55(1):195–230, 2021. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Rudolph:1976:BEL

[Rud76] Peter Rudolph. Bayesian estimation in linear models under different assumptions about the covariance structure: variance component models, equicorrelated models. Mathematische Operationsforschung und Statistik, 7(5):649–663, 1976. CODEN MTOSBI. ISSN 0047-6277.

RuizEspejo:1985:EAO

[Rui85] M. Ruiz Espejo. Equiprecisional allocation and optimum stratification. Statistics: a Journal of Theoretical and Applied Statistics, 16(4):559–562, 1985. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

RuizEspejo:1987:CSS

[Rui87] M. Ruiz Espejo. A control in stratified sampling. Statistics: a Journal of Theoretical and Applied Statistics, 18(2):287–291, 1987. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Rukhin:2007:RNU

[Ruk07] Andrew L. Rukhin. Randomization with non-uniform allocations: fulfillment time distributions and bias properties. Statistics: a Journal of Theoretical and Applied Statistics, 41 (1):11–29, 2007. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Ruschendorf:1981:SOD

[Rüs81] Ludger Rüschendorf. Stochastically ordered distributions and monotonicity of the oc-function of sequential probability ratio tests. Statistics (Berlin, DDR), 12(3):327–338, 1981. CODEN MOSSD5. ISSN 0323-3944.

Ruschendorf:1983:ECM

[Rüs83] Ludger Rüschendorf. Essential completeness of monotone tests and estimators. Statistics (Berlin, DDR), 14(2):243– 250, 1983. CODEN MOSSD5. ISSN 0323-3944.

Ruschendorf:1986:MUT

[Rüs86] Ludger Rüschendorf. Monotonicity and unbiasedness of tests via a.s. constructions. Statistics: a Journal of Theoretical and Applied Statistics, 17(2):221–230, 1986. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Ruschendorf:1987:PPM

[Rüs87] Ludger Rüschendorf. Projections of probability measures. Statistics: a Journal of Theoretical and Applied Statistics, 18(1):123–129, 1987. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Ruschendorf:1990:ERT

[Rüs90] Ludger Rüschendorf. Estimation in random translation models. Statistics: a Journal of Theoretical and Applied Statistics, 21(1):45–55, 1990. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Ruschendorf:1991:SOLa

[Rüs91a] Ludger Rüschendorf. Stochastic ordering of likelihood ratios and partial sufficiency. Statistics: a Journal of Theoretical and Applied Statistics, 22(4):551–558, 1991. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Ruskeepaa:1991:NEE

[Rus91b] Heikki Ruskeepää. A note on estimating equations for linear parameters in discrete-time stochastic processes. Statistics: a Journal of Theoretical and Applied Statistics, 22(3):403–408, 1991. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Ruymgaart:1998:NWC

[Ruy98] Frits H. Ruymgaart. A note on weak convergence of density estimators in Hilbert spaces. Statistics: a Journal of Theoretical and Applied Statistics, 30(4):331–343, 1998. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Rukhin:1997:CPE

[RV97] Andrew L. Rukhin and Ygor Vajda. Change-point estimation as a nonlinear regression problem. Statistics: a Journal of Theoretical and Applied Statistics, 30(3):181–200, 1997. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Rossi:2008:BSM

[RV08] Vivien Rossi and Jean-Pierre Vila. Bayesian selection of multiresponse nonlinear regression model. Statistics: a Journal of Theoretical and Applied Statistics, 42(4):291–311, 2008. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Romeiro:2018:RMB

[RVB18] Renata G. Romeiro, Filidor Vilca, and N. Balakrishnan. A robust multivariate Birnbaum–Saunders distribution: EM estimation. Statistics: a Journal of Theoretical and Applied Statistics, 52(2):321–344, 2018. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Romeiro:2020:RMB

[RVBZ20] Renata G. Romeiro, Filidor Vilca, N. Balakrishnan, and Camila Borelli Zeller. A robust multivariate Birnbaum-Saunders regression model. Statistics: a Journal of Theoretical and Applied Statistics, 54(5):1094–1123, 2020. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Ruli:2022:RCD

[RVM22] Erlis Ruli, Laura Ventura, and Monica Musio. Robust confidence distributions from proper scoring rules. Statistics: a Journal of Theoretical and Applied Statistics, 56(2):455–478, 2022. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Rossmann:1979:UDV

[RW79] H. Rossmann and W. Warmuth. Eine Unschärferelation für diskrete Verteilungen. (German) [An uncertainty principle for discrete distributions]. Statistics (Berlin, DDR), 10(1):173– 178, 1979. CODEN MOSSD5. ISSN 0323-3944.

Rossmann:1984:ECO

[RW84] H. Rossmann and W. Warmuth. On the existence of complete orthonormal systems of pairwise independent random variables over a countable probability space. Statistics (Berlin, DDR), 15(3):405–407, 1984. CODEN MOSSD5. ISSN 0323-3944.

Ronz:1972:RGB

[RWH^{+}72] B. Rönz, H. Weinert, H. Hollatz, O. Bunke, H. Läuter, H. Bunke, H. Bunke, P. H. Müller, O. Bunke, and H. Toutenburg. Rezensionen. (German) [Book reviews]. Mathematische Operationsforschung und Statistik, 3(3):227–236, 1972. CO-DEN MTOSBI. ISSN 0047-6277.

Ronchetti:1986:VSE

[RY86] Elvezio Ronchetti and James H. Yen. Variance stable restimators. Statistics: a Journal of Theoretical and Applied Statistics, 17(2):189–199, 1986. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Ryabko:2014:UHT

[Rya14] Daniil Ryabko. Uniform hypothesis testing for finite-valued stationary processes. Statistics: a Journal of Theoretical and Applied Statistics, 48(1):121–128, 2014. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Rychlik:1987:AMB

[Ryc87] Tomasz Rychlik. An asymptotically most bias-stable estimator of location parameter. Statistics: a Journal of Theoretical and Applied Statistics, 18(4):563–571, 1987. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Rychlik:1993:BRE

[Ryc93a] Tomasz Rychlik. Bias-robustness of l-estimates of location against dependence. Statistics: a Journal of Theoretical and Applied Statistics, 24(1):9–15, 1993. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Rychlik:1993:BEE

[Ryc93b] Tomasz Rychlik. Bounds for expectation of l-estimates for dependent samples. Statistics: a Journal of Theoretical and Applied Statistics, 24(1):1–7, 1993. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Rychlik:2004:OBS

[Ryc04] Tomasz Rychlik. Optimal bounds on l-statistics based on samples drawn with replacement from finite populations. Statistics: a Journal of Theoretical and Applied Statistics, 38(5): 391–412, 2004. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Rychlik:2009:NPU

[Ryc09] Tomasz Rychlik. Non-positive upper bounds on expectations of low rank order statistics from DFR populations. Statistics: a Journal of Theoretical and Applied Statistics, 43(1):53–63, 2009. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Rychlik:2015:MDO

[Ryc15] Tomasz Rychlik. Maximal dispersion of order statistics in dependent samples. Statistics: a Journal of Theoretical and Applied Statistics, 49(2):386–395, 2015. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Rychlik:2021:VEQ

[Ryc21] Tomasz Rychlik. Variation of the expectation and quantiles of the mixture of ordered distributions under imprecise choice of prior. Statistics: a Journal of Theoretical and Applied Statistics, 55(3):682–710, 2021. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Ryden:1995:EOH

[Ryd95] Tobias Rydén. Estimating the order of hidden Markov models. Statistics: a Journal of Theoretical and Applied Statistics, 26(4):345–354, 1995. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Roknossadati:2011:ENU

[RZ11] S. M. Roknossadati and M. Zarepour. M-estimation for near unit roots in spatial autoregression with infinite variance.

Statistics: a Journal of Theoretical and Applied Statistics, 45(4):337–348, 2011. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Solea:2023:HDR

[SA23] Eftychia Solea and Rayan Al Hajj. High-dimensional rankbased graphical models for non-Gaussian functional data. Statistics: a Journal of Theoretical and Applied Statistics, 57(2):388–422, 2023. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/02331888.2023.2201009.

Soliman:2013:RES

[SAEAEA13] Ahmed A. Soliman, A. H. Abd-Ellah, N. A. Abou-Elheggag, and Essam A. Ahmed. Reliability estimation in stressstrength models: an MCMC approach. Statistics: a Journal of Theoretical and Applied Statistics, 47(4):715–728, 2013. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Sahoo:1987:CAU

[Sah87] L. N. Sahoo. On a class of almost unbiased estimators for population ratio. Statistics: a Journal of Theoretical and Applied Statistics, 18(1):119–121, 1987. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Saikkonen:1986:APS

[Sai86] Pentti Saikkonen. Asymptotic properties of some tests for autocorrelation. Statistics: a Journal of Theoretical and Applied Statistics, 17(1):49–61, 1986. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Sakr:1971:AEW

[Sak71] Laila Sakr. Asymptotische Eigenschaften von Wartesystemen mit periodischer Eingangsintensität. (German) [Asymptotic properties of systems with periodic input intensity]. Mathematische Operationsforschung und Statistik, 2(3):227–236, 1971. CODEN MTOSBI. ISSN 0047-6277.

Sarkadi:1982:EQD

[Sar82] K. Sarkadi. The empirical quantile of m-dependent sequences. Statistics (Berlin, DDR), 13(3):405–411, 1982. CODEN MOSSD5. ISSN 0323-3944.

Sarkka:1995:PLA

[Sär95] Aila Särkkä. Pseudo-likelihood approach for Gibbs point processes in connection with field observations. Statistics: a Journal of Theoretical and Applied Statistics, 26(1):89–97, 1995. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Sawitzki:1981:EFE

[Saw81] Günther Sawitzki. Exact filtering in exponential families: Discrete time. Statistics (Berlin, DDR), 12(3):393–401, 1981. CODEN MOSSD5. ISSN 0323-3944.

Sudheesh:2023:USL

[SAX23] K. K. Sudheesh, S. Anjana, and M. Xie. U-statistics for left truncated and right censored data. Statistics: a Journal of Theoretical and Applied Statistics, 57(4):900–917, 2023. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/02331888.2023.2217314.

Schmidt:1982:BRb

[SB82] W. Schmidt and O. Bunké. Book review. Statistics (Berlin, DDR), 13(3):412, 1982. CODEN MOSSD5. ISSN 0323-3944.

Stoimenova:2017:STT

[SB17] Eugenia Stoimenova and N. Balakrishnan. Sidák-type tests for the two-sample problem based on precedence and exceedance statistics. Statistics: a Journal of Theoretical and Applied Statistics, 51(2):247–264, 2017. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Stefanski:1990:DKD

[SC90] Leonard A. Stefanski and Raymond J. Carroll. Deconvolving kernel density estimators. Statistics: a Journal of Theoretical and Applied Statistics, 21(2):169–184, 1990. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Shen:2018:ALM

[SC18] Pao-Sheng Shen and Chyong-Mei Chen. Aalen’s linear model for doubly censored data. Statistics: a Journal of Theoretical and Applied Statistics, 52(6):1328–1343, 2018. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Sperlich:2000:BRa

[SCC^{+}00] S. Sperlich, E. Castell, E. Castell, E. Castell, E. Castell, E. Castell, E. Castell, E. Castell, E. Castell, and E. Castell. Book reviews. Statistics: a Journal of Theoretical and Applied Statistics, 34(2):175–176, 2000. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Schmitz:1970:SMT

[Sch70a] Norbert Schmitz. Sequentielle Minimax-Tests bei Wiener-Prozessen. (German) [Sequential minimax tests for Wiener processes]. Mathematische Operationsforschung und Statistik, 1(4):289–295, 1970. CODEN MTOSBI. ISSN 0047-6277.

Schoch:1970:EKL

[Sch70b] Manfred Schoch. Ein Erweiterungsprinzip als Konzeption zur Lösung kombinatorischer Optimierungsprobleme. (German) [An extension principle as a concept for solving combinatorial optimization problems]. Mathematische Operationsforschung und Statistik, 1(4):265–280, 1970. CODEN MTOSBI. ISSN 0047-6277.

Schulz:1970:LLM

[Sch70c] Günther Schulz. Ein Lösungsalgorithmus für ein lineares mehrstufiges Transportproblem mit proportionalen Mengenänderungen. (German) [A solution algorithm for a linear multi-stage transportation problem with proportional changes in quantity]. Mathematische Operationsforschung und Statistik, 1(3):155–171, 1970. CODEN MTOSBI. ISSN 0047-6277.

Schoch:1973:AEE

[Sch73] Manfred Schoch. Anwendung des Erweiterungsprinzips zur Entwicklung eines Lösungsverfahrens für eine Klasse von Min-Max-Problemen. (German) [Application of the extension principle for the development of a solution procedure for a class of min-max problems]. Mathematische Operationsforschung und Statistik, 4(1):7–22, 1973. CODEN MTOSBI. ISSN 0047-6277.

Schoch:1974:EOG

[Sch74a] Manfred Schoch. Zur Eckpunkt-Optimierung. (German) [On vertex optimization]. Mathematische Operationsforschung und Statistik, 5(2):83–92, 1974. CODEN MTOSBI. ISSN 0047-6277.

Schorr:1974:CCI

[Sch74b] Benno Schorr. On the choice of the class intervals in the application of the chi-square test. Mathematische Operationsforschung und Statistik, 5(4–5):357–377, 1974. CODEN MTOSBI. ISSN 0047-6277.

Schaffer:1975:AVT

[Sch75a] Martin Schaffer. Das asymptotische Verhalten des t-Tests unter allgemeinen Alternativen. (German) [The asymptotic behavior of the t tests under general alternatives]. Mathematische Operationsforschung und Statistik, 6(5):775–785, 1975. CODEN MTOSBI. ISSN 0047-6277.

Schmidt:1975:ANL

[Sch75b] Wolfgang H. Schmidt. Asymptotic normality of least-square estimators in multivariate singular linear models. Mathematische Operationsforschung und Statistik, 6(2):285–299, 1975. CODEN MTOSBI. ISSN 0047-6277.

Schmidt:1975:AOE

[Sch75c] Wolfgang H. Schmidt. Asymptotic optimality of estimators in multivariate linear models. Mathematische Operationsforschung und Statistik, 6(5):713–731, 1975. CODEN MTOSBI. ISSN 0047-6277.

Schulz:1975:AELa

[Sch75d] Günther Schulz. Eine Anwendung des Erweiterungsprinzips auf die Lösung allgemeiner linearer Optimierungsaufgaben. Teil I. (German) [An application of the extension principle to the solution of general nonlinear optimization problems. Part I]. Mathematische Operationsforschung und Statistik, 6 (1):3–14, 1975. CODEN MTOSBI. ISSN 0047-6277.

Schulz:1975:AELb

[Sch75e] Günther Schulz. Eine Anwendung des Erweiterungsprinzips auf die Lösung allgemeiner linearer Optimierungsaufgaben. Teil II. (German) [An application of the expansion principle to the solution of general nonlinear optimization problems. Part II]. Mathematische Operationsforschung und Statistik, 6 (6):855–866, 1975. CODEN MTOSBI. ISSN 0047-6277.

Schwarz:1975:KEM

[Sch75f] Kunibert Schwarz. Die konvexe Erweiterungsmethode — Eine Methode zur Lösung diskreter Optimierungsaufgaben. (German) [The convex extension method — a method for solving discrete optimization problems]. Mathematische Operationsforschung und Statistik, 6(3):331–340, 1975. CODEN MTOSBI. ISSN 0047-6277.

Schafer:1976:AAI

[Sch76a] Thomas Schäfer. Eine Anmerkung zu der Arbeit “Identifiability and estimability” von H. Bunke und O.Bunke. (German) [A note on the work ‘Identifiability and estimability’ of H. Bunke and O. Bunke]. Mathematische Operationsforschung und Statistik, 7(5):697–699, 1976. CODEN MTOSBI. ISSN 0047-6277. See [BB74].

Schafer:1976:EVL

[Sch76b] Wolfgang Schäfer. Die explizite und vollständige Lösung von zwei speziellen Aufgaben der linearen diskreten Approximation. (German) [The explicit and complete solution of two specific tasks of linear discrete approximation]. Mathematische Operationsforschung und Statistik, 7(6):879–888, 1976. CODEN MTOSBI. ISSN 0047-6277.

Schmidt:1976:SCV

[Sch76c] Wolfgange H. Schmidt. Strong consistency of variance estimation and asymptotic theory for tests of the linear hypothesis in multivariate linear models. Mathematische Operationsforschung und Statistik, 7(5):701–705, 1976. CODEN MTOSBI. ISSN 0047-6277.

Schulze:1976:CHW

[Sch76d] U. E. Schulze. Consistency of the Hotelling-, Wilks–Royand Pillai-tests for a general linear hypothesis in multiyariate linear nonnormally distributed models. Mathematische Operationsforschung und Statistik, 7(1):11–21, 1976. CODEN MTOSBI. ISSN 0047-6277.

Schmidt:1977:AML

[Sch77a] Wolfgang H. Schmidt. Asymptotics in multivariate linear models with optimal experimental designs. Statistics (Berlin, DDR), 8(4):447–452, 1977. CODEN MOSSD5. ISSN 0323-3944.

Schutze:1977:AVB

[Sch77b] D. Schütze. Über das asymptotische Verhalten von Bayesschen und Maximum-Likelihood-Schätzungen. (German) [On the asymptotic behavior of Bayesian and maximum likelihood estimates]. Statistics (Berlin, DDR), 8(3):367–373, 1977. CO-DEN MOSSD5. ISSN 0323-3944.

Schmidt:1979:ARE

[Sch79a] Wolfgang H. Schmidt. Asymptotic results for estimation and testing variances in regression models. Statistics (Berlin, DDR), 10(2):209–236, 1979. CODEN MOSSD5. ISSN 0323-3944.

Schmidt:1979:NTA

[Sch79b] Wolfgang H. Schmidt. Normal theory approximations to tests for linear hypotheses. Statistics (Berlin, DDR), 10(3):353– 366, 1979. CODEN MOSSD5. ISSN 0323-3944.

Schonherr:1979:LTS

[Sch79c] Siegfried Schönherr. Limit theorems for single server queues with mixed renewal and Poisson inputs. Statistics (Berlin, DDR), 10(4):601–612, 1979. CODEN MOSSD5. ISSN 0323-3944.

Schonherr:1981:EWT

[Sch81] Siegfried Schönherr. The existence of waiting time moments for single server queues with mixed renewal and Poisson inputs. Statistics (Berlin, DDR), 12(2):225–232, 1981. CODEN MOSSD5. ISSN 0323-3944.

Schafer:1982:RTS

[Sch82a] W. Schäfer. A representation theorem for stochastic kernels and its application in comparison of experiments. Statistics (Berlin, DDR), 13(4):531–541, 1982. CODEN MOSSD5. ISSN 0323-3944.

Schmidt:1982:THN

[Sch82b] Wolfgahg H. Schmidt. Testing hypotheses in nonlinear regressions. Statistics (Berlin, DDR), 13(1):3–19, 1982. CODEN MOSSD5. ISSN 0323-3944.

Schulze:1982:ESR

[Sch82c] U. Schulze. Estimation in segmented regression: known number of regimes. Statistics (Berlin, DDR), 13(2):295–316, 1982. CODEN MOSSD5. ISSN 0323-3944.

Schwarze:1982:BR

[Sch82d] G. Schwarze. Book review. Statistics (Berlin, DDR), 13(2): 351–352, 1982. CODEN MOSSD5. ISSN 0323-3944.

Schmidt:1983:BRa

[Sch83a] W. Schmidt. Book reviews. Statistics (Berlin, DDR), 14(1): 28, 1983. CODEN MOSSD5. ISSN 0323-3944.

Schmidt:1983:BRc

[Sch83b] W. Schmidt. Book reviews. Statistics (Berlin, DDR), 14(3): 380, 1983. CODEN MOSSD5. ISSN 0323-3944.

Schulze:1984:MOU

[Sch84] U. Schulze. Math.operationsforsch.u.statist.,ser.statist. Statistics (Berlin, DDR), 15(2):204, 1984. CODEN MOSSD5. ISSN 0323-3944.

Schmidt:1985:PBM

[Sch85] Volkeb Schmidt. Poisson bounds for moments of shot noise processes. Statistics: a Journal of Theoretical and Applied Statistics, 16(2):253–262, 1985. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Schabe:1987:NAZ

[Sch87a] Hendrik Schäbe. Nichtparametrische Auswertung von zufällig zensierten Stichproben mit unbestimmten Ereignissen. (German) [Nonparametric evaluation of randomly censored sampling with indeterminate events]. Statistics: a Journal of Theoretical and Applied Statistics, 18(3):361–371, 1987. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Schabe:1987:SET

[Sch87b] Hendrik Schäbe. Schiefe und Exzeß als Teststatistiken. (German) [Skewness and excess as test statistics]. Statistics: a Journal of Theoretical and Applied Statistics, 18(4):613–621, 1987. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Schmidt:1987:BRa

[Sch87c] W. Schmidt. Book review. Statistics: a Journal of Theoretical and Applied Statistics, 18(1):150, 1987. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Schmidt:1987:BRb

[Sch87d] W. Schmidt. Book review. Statistics: a Journal of Theoretical and Applied Statistics, 18(2):218, 1987. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Schulze:1987:BR

[Sch87e] U. Schulze. Book review. Statistics: a Journal of Theoretical and Applied Statistics, 18(1):130, 1987. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Schwabe:1987:ADR

[Sch87f] R. Schwabe. On a adaptive design in regression. Statistics: a Journal of Theoretical and Applied Statistics, 18(4):521–525, 1987. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Schabe:1989:NAS

[Sch89a] Hendrik Schabe. Nichtparametrische Auswertung von stochastisch beidseitig zensierten Stichproben. (German) [Non-parametric analysis of stochastic both-sides-censored samples]. Statistics: a Journal of Theoretical and Applied Statistics, 20(3):397–405, 1989. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Schatte:1989:ASC

[Sch89b] Peter Schatte. On almost sure convergence of subsequences in the Central Limit Theorem. Statistics: a Journal of Theoretical and Applied Statistics, 20(4):593–605, 1989. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Schmid:1989:ITO

[Sch89c] W. Schmid. Identification of a Type I outlier in an autoregressive model. Statistics: a Journal of Theoretical and Applied Statistics, 20(4):531–545, 1989. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Schmidt:1989:BRa

[Sch89d] W. Schmidt. Book review. Statistics: a Journal of Theoretical and Applied Statistics, 20(1):12, 1989. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Schmidt:1989:BRb

[Sch89e] W. Schmidt. Book reviews. Statistics: a Journal of Theoretical and Applied Statistics, 20(3):422, 1989. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Schmidt:1989:BRc

[Sch89f] W. Schmidt. Book reviews. Statistics: a Journal of Theoretical and Applied Statistics, 20(3):452, 1989. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Schabe:1990:BRM

[Sch90a] Hendrik Schabe. Bias reduction of maximum likelihood estimators using kernel estimators. Statistics: a Journal of Theoretical and Applied Statistics, 21(4):577–590, 1990. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Schmalfuss:1990:EAL

[Sch90b] Bjorn Schmalfuss. Endlichdimensionale Approximation der Lösung der stochastischen Navier–Stokes-Gleichung. (German) [Finite-dimensional approximation of the solution of the stochastic Navier–Stokes equation]. Statistics: a Journal of Theoretical and Applied Statistics, 21(1):149–157, 1990. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Schabe:1991:GFT

[Sch91a] H. Schabe. Goodness-of-fit-tests for composite hypotheses with censored samples. Statistics: a Journal of Theoretical and Applied Statistics, 22(1):95–109, 1991. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Schaffler:1991:MLE

[Sch91b] S. Schaffler. Maximum likelihood estimation for linear regression models with autoregressive errors. Statistics: a Journal of Theoretical and Applied Statistics, 22(2):191–198, 1991. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Schwabe:1995:D

[Sch95] R. Schwabe. Discussion. Statistics: a Journal of Theoretical and Applied Statistics, 26(2):172–173, 1995. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Schwabe:1996:ODA

[Sch96] Rainer Schwabe. Optimal designs for additive linear models. Statistics: a Journal of Theoretical and Applied Statistics, 27 (3–4):267–278, 1996. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Schwabe:1998:BR

[Sch98] Rainer Schwabe. Book reviews. Statistics: a Journal of Theoretical and Applied Statistics, 32(2):199–201, 1998. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Schwabe:1999:BRb

[Sch99] Rainer Schwabe. Book review. Statistics: a Journal of Theoretical and Applied Statistics, 32(4):395–396, 1999. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Schmisser:2013:PND

[Sch13] Emeline Schmisser. Penalized nonparametric drift estimation for a multidimensional diffusion process. Statistics: a Journal of Theoretical and Applied Statistics, 47(1):61–84, 2013. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Schotz:2022:SLL

[Sch22] Christof Schötz. Strong laws of large numbers for generalizations of fréchet mean sets. Statistics: a Journal of Theoretical and Applied Statistics, 56(1):34–52, 2022. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Sinha:1983:STS

[SD83] K. Sinha and A. Dey. A series of truly self dual PBIB designs. Statistics (Berlin, DDR), 14(1):53–54, 1983. CODEN MOSSD5. ISSN 0323-3944.

Scharm:1986:MED

[SD86] Rolf Scharm and Montka Dewess. Mixtures of exponential distributions in the GI/G/1 queue (with and without warming up). Statistics: a Journal of Theoretical and Applied Statistics, 17(2):291–302, 1986. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Subramanian:2017:TSP

[SD17] Leela Subramanian and Vaijayanti U. Dixit. Tests for the skewness parameter of two-piece double exponential distribution in the presence of nuisance parameters. Statistics: a Journal of Theoretical and Applied Statistics, 52(1):115–132, 2017. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Shen:2019:SCB

[SDC19] Biyi Shen, Fabian Dunker, and Chixiang Chen. Simultaneous confidence bands for growth incidence curves in weighted supnorm metrics. Statistics: a Journal of Theoretical and Applied Statistics, 53(6):1269–1288, 2019. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Shafiei:2016:ABS

[SDJ16] Sobhan Shafiei, Mahdi Doostparast, and Ahad Jamalizadeh. The alpha–beta skew normal distribution: properties and applications. Statistics: a Journal of Theoretical and Applied Statistics, 50(2):338–349, 2016. CO-DEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/02331888.2015.1096938.

Schorning:2018:ODN

[SDKM18] Kirsten Schorning, Holger Dette, Katrin Kettelhake, and Tilman Möller. Optimal designs for non-competitive enzyme inhibition kinetic models. Statistics: a Journal of Theoretical and Applied Statistics, 52(6):1359–1378, 2018. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Sordo:2015:NVO

[SdSSL15] Miguel A. Sordo, Marilia C. de Souza, and Alfonso Suárez-Llorens. A new variability order based on tail-heaviness. Statistics: a Journal of Theoretical and Applied Statistics, 49(5):1042–1061, 2015. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Sebastian:1976:DOE

[Seb76] Hans-Jürgen Sebastian. Dynamische Optimierung unter Einbeziehung zeitabhängiger Einflußparameter. (German) [Dynamic optimization involving time-dependent influence of parameters]. Mathematische Operationsforschung und Statistik, 7(3):427–452, 1976. CODEN MTOSBI. ISSN 0047-6277.

Sulanke:1985:BR

[SEG^{+}85] H. Sulanke, H. J. Engelbert, W. Grecksch, H. J. Girlich, H. Toutenburg, V. Nollau, Hans Peter Höschel, K. Hoffmann, B. Droge, F. Liese, and R. Pincus. Book reviews. Statistics: a Journal of Theoretical and Applied Statistics, 16(1):150–160, 1985. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Seifert:1978:NUC

[Sei78] Burkhardt Seifert. Note on the UMPU-character of a test for the mean in balanced randomized nested classification. Statistics (Berlin, DDR), 9(2):185–189, 1978. CODEN MOSSD5. ISSN 0323-3944.

Seifert:1979:OTF

[Sei79] Burkhardt Seifert. Optimal testing for fixed effects in general balanced mixed classification models. Statistics (Berlin, DDR), 10(2):237–255, 1979. CODEN MOSSD5. ISSN 0323-3944.

Sen:1982:ATS

[Sen82] Pranab Kumar Sen. Asymptotic theory of some tests for constancy of regression relationships over time. Statistics (Berlin, DDR), 13(1):21–31, 1982. CODEN MOSSD5. ISSN 0323-3944.

Sengupta:2011:UET

[Sen11] S. Sengupta. Unbiased estimation of P (X > Y) for twoparameter exponential populations using order statistics. Statistics: a Journal of Theoretical and Applied Statistics, 45(2):179–188, 2011. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Sepanski:2001:RMM

[Sep01] Jungsywan H. Sepanski. On a repeated-measurement model with errors in dependent variable. Statistics: a Journal of Theoretical and Applied Statistics, 35(2):97–112, 2001. CO-DEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Schmidt:1982:BRa

[SFB^{+}82] W. Schmidt, J. Franz, O. Bunke, M. Möhner, W. Näther, H. Bunke, H. Aäuter, H. Bandemer, H. Bandemer, H. J. Girlich, H. J. Rossberg, H. Läuter, R. Mä, K. Henschke, J. Franz, H. H. Fährmann, A. Hahnewald-Busch, E. Platen, Barbara Grabowski, and J. Polzehl. Book reviews. Statistics (Berlin, DDR), 13(1):145–160, 1982. CODEN MOSSD5. ISSN 0323-3944.

Stoyan:1991:SSS

[SG91] Dietrich Stoyan and Pavel Grabarnik. Statistics for the stationary Strauss model by the cusp point method. Statistics: a Journal of Theoretical and Applied Statistics, 22(2):283–289, 1991. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Salvador:2004:BIM

[SGG04] Manuel Salvador, Jose Luis Gallizo, and Pilar Gargallo. Bayesian inference in a matrix normal dynamic linear model with unknown covariance matrices. Statistics: a Journal of Theoretical and Applied Statistics, 38(4):307–335, 2004. CO-DEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Salinas:2018:CSN

[SGMFB18] Hugo S. Salinas, Héctor W. Gómez, Guillermo Martínez-Flórez, and Heleno Bolfarine. Corrigendum: Skew-normal alpha-power model [Statistics 48(2014) 1414–1428]. Statistics: a Journal of Theoretical and Applied Statistics, 52(4): 950–953, 2018. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic). See [MFBG14b].

Schmidt:1983:BRb

[SH83] W. Schmidt and K. Henschke. Book reviews. Statistics (Berlin, DDR), 14(2):242, 1983. CODEN MOSSD5. ISSN 0323-3944.

Stoyan:1986:SMS

[SH86] D. Stoyan and H. Hermann. Some methods for statistical analysis of planar random tessellations. Statistics: a Journal of Theoretical and Applied Statistics, 17(3):407–420, 1986. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Soltani:2009:GTS

[SH09] A. R. Soltani and H. Homei. A generalization for two-sided power distributions and adjusted method of moments. Statistics: a Journal of Theoretical and Applied Statistics, 43(6): 611–620, 2009. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Sepehri:2017:ALM

[SH17] Amir Sepehri and Naftali Harris. The accessible lasso models. Statistics: a Journal of Theoretical and Applied Statistics, 51(4):711–721, 2017. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Sharma:1983:SAC

[Sha83] Divakar Sharma. Sufficient admissibility conditions for scale parameter estimator. Statistics (Berlin, DDR), 14(2):235– 241, 1983. CODEN MOSSD5. ISSN 0323-3944.

Shanmugam:1986:CPB

[Sha86] Ramalingam Shanmugam. A characterization of positive binomial distribution and its goodness-of-fit application. Statistics: a Journal of Theoretical and Applied Statistics, 17(1): 33–39, 1986. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Shao:1991:CJV

[Sha91] Jun Shao. Consistency of jackknife variance estimators. Statistics: a Journal of Theoretical and Applied Statistics, 22 (1):49–57, 1991. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Shao:1996:IDP

[Sha96a] Jun Shao. Invited discussion paper: Resampling methods in sample surveys. Statistics: a Journal of Theoretical and Applied Statistics, 27(3–4):203–237, 1996. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Shao:1996:R

[Sha96b] Jun Shao. Reply. Statistics: a Journal of Theoretical and Applied Statistics, 27(3–4):252–254, 1996. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Shan:2018:ECL

[Sha18] Guogen Shan. Exact confidence limits for the probability of response in two-stage designs. Statistics: a Journal of Theoretical and Applied Statistics, 52(5):1086–1095, 2018. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Shang:2021:REG

[Sha21] Wenpeng Shang. Robust estimation for generalized casecohort design under the proportional hazards model. Statistics: a Journal of Theoretical and Applied Statistics, 55(6): 1231–1270, 2021. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Sheahan:1988:RER

[She88] J. N. Sheahan. Robust estimation of regression and scale parameters in linear models with asymmetric error distributions. Statistics: a Journal of Theoretical and Applied Statistics, 19 (1):27–37, 1988. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Sheena:2003:MNP

[She03] Yo Sheena. On minimaxity of the normal precision matrix estimator of Krishnamoorthy and Gupta. Statistics: a Journal of Theoretical and Applied Statistics, 37(5):387–399, 2003. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Shen:2014:AAI

[She14] Aiting Shen. On asymptotic approximation of inverse moments for a class of nonnegative random variables. Statistics: a Journal of Theoretical and Applied Statistics, 48(6):1371– 1379, 2014. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Shin:2015:SBP

[Shi15] Dong Wan Shin. Stationary bootstrapping for panel cointegration tests under cross-sectional dependence. Statistics: a Journal of Theoretical and Applied Statistics, 49(1):209–223, 2015. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Schoch:1971:RGB

[SHP^{+}71] M. Schoch, H. Hollatz, J. Piehler, P. Franken, W. Richter, W. Richter, H. Läuter, and H. Toutenburg. Rezensionen. (German) [Book reviews]. Mathematische Operationsforschung und Statistik, 2(2):171–178, 1971. CODEN MTOSBI. ISSN 0047-6277.

Siegel:1974:VII

[Sie74] Gerhard Siegel. Verallgemeinerung einer Irrfahrt im R^{1} und deren Bedeutung für Bedienungssysteme mit verzögertem Bedienungsbeginn. (German) [Generalization of a random walk in R^{1} and its relevance for service systems with delayed operation start]. Mathematische Operationsforschung und Statistik, 5(6):465–486, 1974. CODEN MTOSBI. ISSN 0047-6277.

Siegel:1986:CSP

[Sie86] G. Siegel. Convergence of stochastic processes with random parameters. Statistics: a Journal of Theoretical and Applied Statistics, 17(1):121–138, 1986. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Siegel:1987:ACS

[Sie87] G. Siegel. On the Anscombe condition for stochastic processes in a separable Banach space. Statistics: a Journal of Theoretical and Applied Statistics, 18(3):437–451, 1987. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Sielski:2014:NTS

[Sie14] Piotr Sielski. On a new type of sufficiency in undominated models. Statistics: a Journal of Theoretical and Applied Statistics, 48(3):539–551, 2014. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Scalera:2021:RLF

[SIM21] Valentino Scalera, Maria Iannario, and Anna Clara Monti. Robust link functions. Statistics: a Journal of Theoretical and Applied Statistics, 55(4):963–977, 2021. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Sinha:1981:CAP

[Sin81] Kishore Sinha. On the construction of m-associate PBIB designs II. Statistics (Berlin, DDR), 12(4):503–508, 1981. CO-DEN MOSSD5. ISSN 0323-3944.

Singh:1984:SSB

[Sin84] Man Singh. Steady-state behaviour of serial queueing processes with impatient customers. Statistics (Berlin, DDR), 15(2):289–298, 1984. CODEN MOSSD5. ISSN 0323-3944.

Sinha:1991:SES

[Sin91] Kishore Sinha. Some equiblock-sized, unequireplicated cdesigns. Statistics: a Journal of Theoretical and Applied Statistics, 22(4):525–528, 1991. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Siniksaran:2005:GWL

[Sin05] Enis Siniksaran. On the geometry of F, Wald, LR, and LM tests in linear regression models. Statistics: a Journal of Theoretical and Applied Statistics, 39(4):287–299, 2005. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Singh:2009:NMI

[Sin09] Sarjinder Singh. A new method of imputation in survey sampling. Statistics: a Journal of Theoretical and Applied Statistics, 43(5):499–511, 2009. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Singh:2013:DPC

[Sin13] Sarjinder Singh. A dual problem of calibration of design weights. Statistics: a Journal of Theoretical and Applied Statistics, 47(3):566–574, 2013. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Sinha:1997:CPV

[SJK97] Kishore Sinha, Byron Jones, and Sanpei Kageyama. Constructions of pairwise-and variance-balanced designs. Statistics: a Journal of Theoretical and Applied Statistics, 29(3): 241–250, 1997. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Saud:1989:NIT

[SK89] Ibne Saud and M. Z. Khan. A note on information theoretic measures. Statistics: a Journal of Theoretical and Applied Statistics, 20(1):161–164, 1989. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Sahai:1992:CEV

[SK92] Hardeo Sahai and Anwer Khurshid. A comparison of estimators of variance components in a two-way balanced crossed classification random effects model. Statistics: a Journal of Theoretical and Applied Statistics, 23(2):129–143, 1992. CO-DEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Son:2005:BSC

[SK05] Young Sook Son and Seong W. Kim. Bayesian single change point detection in a sequence of multivariate normal observations. Statistics: a Journal of Theoretical and Applied Statistics, 39(5):373–387, 2005. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Sakate:2014:DBC

[SK14a] D. M. Sakate and D. N. Kashid. A deviance-based criterion for model selection in GLM. Statistics: a Journal of Theoretical and Applied Statistics, 48(1):34–48, 2014. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Sankaran:2014:BPM

[SK14b] P. G. Sankaran and Debasis Kundu. A bivariate Pareto model. Statistics: a Journal of Theoretical and Applied Statistics, 48(2):241–255, 2014. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Schmidt:2020:ASL

[SK20] Jacob Peter Schmidt and Udo Kamps. Almost sure limit behaviour of Pfeifer record values. Statistics: a Journal of Theoretical and Applied Statistics, 54(4):830–840, 2020. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Singh:2021:CRR

[SK21] Shivangi Singh and Chanchal Kundu. On cumulative residual Renyi’s entropy for double truncated distribution. Statistics: a Journal of Theoretical and Applied Statistics, 55(6):1310– 1327, 2021. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Shih:2019:ECM

[SKCE19] Jia-Han Shih, Yoshihiko Konno, Yuan-Tsung Chang, and Takeshi Emura. Estimation of a common mean vector in bivariate meta-analysis under the FGM copula. Statistics: a Journal of Theoretical and Applied Statistics, 53(3):673–695, 2019. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Sreedevi:2021:NPT

[SKD21] E. P. Sreedevi, Sudheesh K. Kattumannil, and Isha Dewan. A non-parametric test for independence of time to failure and cause of failure for discrete competing risks data. Statistics: a Journal of Theoretical and Applied Statistics, 55(5):1107– 1122, 2021. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Skinner:1996:D

[Ski96] C. J. Skinner. Discussion. Statistics: a Journal of Theoretical and Applied Statistics, 27(3–4):245–246, 1996. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Sachs:1992:BR

[SKJ92] H. Sachs, W. Krabs, and H. Th. Jongen. Book review. Statistics: a Journal of Theoretical and Applied Statistics, 23(1): 95–96, 1992. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Sultana:2021:TEP

[SKPK21] Farha Sultana, Arnab Koley, Ayan Pal, and Debasis Kundu. On two exponential populations under a joint adaptive type-II progressive censoring. Statistics: a Journal of Theoretical and Applied Statistics, 55(6):1328–1355, 2021. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Sinha:1993:CRD

[SKS93] Kishore Sinha, Sanpei Kageyama, and Mithilesh Kumar Singh. Construction of rectangular designs. Statistics: a Journal of Theoretical and Applied Statistics, 25(1):63–70, 1993. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Safari-Katesari:2020:MCD

[SKSZ20] Hadi Safari-Katesari, S. Yaser Samadi, and Samira Zaroudi. Modelling count data via copulas. Statistics: a Journal of Theoretical and Applied Statistics, 54(6):1329–1355, 2020. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Singh:2018:ERS

[SKV18] G. N. Singh, Amod Kumar, and Gajendra K. Vishwakarma. Estimation of a rare sensitive attribute in a stratified twostage unrelated randomized response model by using Poisson probability distribution. Statistics: a Journal of Theoretical and Applied Statistics, 52(4):901–918, 2018. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Shi:1995:NCR

[SL95] Peide Shi and Guoying Li. A note on the convergent rates of M-estimates for a partly linear model. Statistics: a Journal of Theoretical and Applied Statistics, 26(1):27–47, 1995. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Song:2011:ALS

[SL11] Kai-Sheng Song and Ta-Hsin Li. Asymptotics of least squares for nonlinear harmonic regression. Statistics: a Journal of Theoretical and Applied Statistics, 45(4):309–318, 2011. CO-DEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Sunoj:2012:DCR

[SL12] S. M. Sunoj and M. N. Linu. Dynamic cumulative residual Rényi’s entropy. Statistics: a Journal of Theoretical and Applied Statistics, 46(1):41–56, 2012. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Salhi:2017:BRM

[SL17] Yahia Salhi and Stéphane Loisel. Basis risk modelling: a cointegration-based approach. Statistics: a Journal of Theoretical and Applied Statistics, 51(1):205–221, 2017. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Song:2023:BRE

[SL23] Yuping Song and Hangyan Li. Bias reduction estimation for drift coefficient in diffusion models with jumps. Statistics: a Journal of Theoretical and Applied Statistics, 57(3):597– 616, 2023. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-

4910 (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/02331888.2023.2201504.

Salasar:2015:IML

[SLL15] Luis Ernesto B. Salasar, José Galvão Leite, and Francisco Louzada. On the integrated maximum likelihood estimators for a closed population capture-recapture model with unequal capture probabilities. Statistics: a Journal of Theoretical and Applied Statistics, 49(6):1204–1220, 2015. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Saida:2017:DSC

[SlP17] Abdallah Ben Saïda and Jean luc Prigent. On the diversity score: a copula approach. Statistics: a Journal of Theoretical and Applied Statistics, 51(1):188–204, 2017. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Severance-Lossin:1999:EDA

[SLS99] Eric Severance-Lossin and Stefan Sperlich. Estimation of derivatives for additive separable models. Statistics: a Journal of Theoretical and Applied Statistics, 33(3):241–265, 1999. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Sengupta:2008:UEU

[SM08] Samindranath Sengupta and Sujay Mukhuti. Unbiased estimation of P (X > Y) using ranked set sample data. Statistics: a Journal of Theoretical and Applied Statistics, 42(3): 223–230, 2008. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Song:2019:UMH

[SM19] Yu Song and Reza Modarres. Unified multivariate hypergeometric interpoint distances. Statistics: a Journal of Theoretical and Applied Statistics, 53(4):921–942, 2019. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Samanta:2022:OPO

[SMK22] Debashis Samanta, Shuvashree Mondal, and Debasis Kundu. Optimal plan for ordered step-stress stage life testing. Statistics: a Journal of Theoretical and Applied Statistics, 56(6): 1319–1344, 2022. CODEN MOSSD5. ISSN 0233-1888 (print),

1029-4910 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/02331888.2022.2152453.

Song:2018:MFV

[SMLL18] Anchao Song, Tiefeng Ma, Shaogao Lv, and Changsheng Lin. A model-free variable selection method for reducing the number of redundant variables. Statistics: a Journal of Theoretical and Applied Statistics, 52(6):1212–1248, 2018. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Stoyan:1980:FSP

[SMP80] D. Stoyan, J. Mecke, and S. Pohlmann. Formulas for stationary planar fibre processes II — partially oriented-fibre systems. Statistics (Berlin, DDR), 11(2):281–286, 1980. CO-DEN MOSSD5. ISSN 0323-3944.

Shao:2022:RSI

[SMW22] Yujing Shao, Wei Ma, and Lei Wang. Robust statistical inference for longitudinal data with nonignorable dropouts. Statistics: a Journal of Theoretical and Applied Statistics, 56(5): 1072–1094, 2022. CODEN MOSSD5. ISSN 0233-1888 (print),

1029-4910 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/02331888.2022.2110250.

Sharfuddin:1981:ESR

[SN81] S. M. Sharfuddin and Fadhil Ajab Nahab. Expectation of the square root of an orthosymmetrie determinant. Statistics (Berlin, DDR), 12(4):609–612, 1981. CODEN MOSSD5. ISSN 0323-3944.

Saboor:2014:CSG

[SN14] A. Saboor and S. Nadarajah. Corrigendum to ‘Some gamma distributions’ by Saralees Nadarajah. Statistics: a Journal of Theoretical and Applied Statistics, 48(5):1185, 2014. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic). See [Nad08c].

Schoch:1972:RGB

[SND^{+}72] M. Schoch, J. Nietzsch, W. Dück, H. Damm, H. Hollatz, D. Schreiter, K. Lommatzsch, O. Bunke, N. N. Worobjoff, E. Rödel, H. Toutenburg, J. Läter, D. Pötschke, and O. Bunke. Rezensionen. (German) [Book reviews]. Mathematische Operationsforschung und Statistik, 3(2):145–156, 1972. CODEN MTOSBI. ISSN 0047-6277.

Shin:2009:TSU

[SO09] Dong Wan Shin and Man-Suk Oh. Tests for seasonal unit roots in panels of cross-sectionally correlated time series. Statistics: a Journal of Theoretical and Applied Statistics, 43(2):139–152, 2009. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Soltani:2019:PFL

[Sol19] A. R. Soltani. A pivot function and its limiting distribution: applications in goodness of fit and testing hypothesis. Statistics: a Journal of Theoretical and Applied Statistics, 53(6): 1386–1395, 2019. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Sommermeter:1971:KLO

[Som71] Klaus Sommermeter. Koeffizientenänderung in der linearen Optimierung. (German) [Coefficient change in linear optimization]. Mathematische Operationsforschung und Statistik, 2(4):271–281, 1971. CODEN MTOSBI. ISSN 0047-6277.

Sonnemann:1982:OCL

[Son82] E. Sonnemann. D-optimality of complete latin squares. Statistics (Berlin, DDR), 13(3):387–394, 1982. CODEN MOSSD5. ISSN 0323-3944.

Song:2020:UEE

[Son20] Kyungchul Song. A uniform-in-P Edgeworth expansion under weak Cramér conditions. Statistics: a Journal of Theoretical and Applied Statistics, 54(5):926–950, 2020. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Saran:1998:CRR

[SP98a] Jagdish Saran and Narinder Pushkarna. Correction to a recurrence relation between moments of order statistics. Statistics: a Journal of Theoretical and Applied Statistics, 31(1):79–81, 1998. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Saran:1998:SNI

[SP98b] Jagdish Saran and Narinder Pushkarna. Some new identities for single moments of order statistics. Statistics: a Journal of Theoretical and Applied Statistics, 30(4):345–355, 1998. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Saran:2001:RRM

[SP01] Jadish Saran and Narinder Pushkarna. Recurrence relations for moments of progressive Type-II right censored order statistics from Burr distribution. Statistics: a Journal of Theoretical and Applied Statistics, 35(4):495–507, 2001. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Shah:2008:BEU

[SP08] J. B. Shah and M. N. Patel. Bayes estimation under exponentially weighted minimum expected loss function from multiply Type-II censored Rayleigh data. Statistics: a Journal of Theoretical and Applied Statistics, 42(5):429–434, 2008. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Spjotvoll:1977:RCR

[Spj77] E. Spjotvoll. Random coefficients regression models. a review. Statistics (Berlin, DDR), 8(1):69–93, 1977. CODEN MOSSD5. ISSN 0323-3944.

Sen:2018:FIG

[SPTB18] Tanmay Sen, Biswabrata Pradhan, Yogesh Mani Tripathi, and Ritwik Bhattacharya. Fisher information in generalized progressive hybrid-censored data. Statistics: a Journal of Theoretical and Applied Statistics, 52(5):1025–1039, 2018. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Saran:1991:JDR

[SR91] Jagdish Saran and Sarita Rani. On joint distributions of rank order statistics for equal sample sizes. Statistics: a Journal of Theoretical and Applied Statistics, 22(2):299–312, 1991.

CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Saran:1994:JDR

[SR94] Jagdish Saran and Sarita Rani. On joint distributions of rank order statistics for unequal sample sizes. Statistics: a Journal of Theoretical and Applied Statistics, 25(2):137–157, 1994. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Smith:2002:CGL

[SR02a] Brian Smith and William Rayens. Conditional generalized Liouville distributions on the simplex. Statistics: a Journal of Theoretical and Applied Statistics, 36(2):185–194, 2002. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Smith:2002:MPD

[SR02b] Brian Smith and William Rayens. Multivariate positive dependence on the simplex. Statistics: a Journal of Theoretical and Applied Statistics, 36(1):75–88, 2002. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Sieber:1974:MMO

[SS74] Norbert Sieber and Hans-Jürgen Sebastian. Ein mathematisches Modell zur Optimalregelung von Versorgungssystemen. (German) [A mathematical model for optimal control of power systems]. Mathematische Operationsforschung und Statistik, 5(9):661–685, 1974. CODEN MTOSBI. ISSN 0047-6277.

Saran:1979:FPS

[SS79] Jagdish Saran and Kanwar Sen. On the fluctuations of partial sums. Statistics (Berlin, DDR), 10(3):469–478, 1979. CODEN MOSSD5. ISSN 0323-3944.

Saran:1981:SDR

[SS81] Jagdish Saran and Kanwar Sen. Some distribution results on unconditional ballot problems. Statistics (Berlin, DDR), 12 (4):613–619, 1981. CODEN MOSSD5. ISSN 0323-3944.

Schwandtke:1986:SDS

[SS86] A. Schwandtke and D. Stoyan. Stereological determination of the second moment of mean curvature. Statistics: a Journal of Theoretical and Applied Statistics, 17(3):421–427, 1986. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Sahoo:1994:EFC

[SS94] J. Sahoo and L. N. Sahoo. On the efficiency of four chain-type estimators in two-phase sampling under a model. Statistics: a Journal of Theoretical and Applied Statistics, 25(4):361–366, 1994. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Sato:1998:NCI

[SS98] Ken-Iti Sato and Fred W. Steutel. Note on the continuation of infinitely divisible distributions and canonical measures. Statistics: a Journal of Theoretical and Applied Statistics, 31(4):347–357, 1998. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Schwabe:1999:BRa

[SS99] Rainer Schwabe and Winfried Stute. Book review. Statistics: a Journal of Theoretical and Applied Statistics, 32(3):293– 295, 1999. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Shin:2004:NTU

[SS04] Dong Wan Shin and Beong Soo So. Normal tests for unit roots based on instrumental variable estimators. Statistics: a Journal of Theoretical and Applied Statistics, 38(2):123–132, 2004. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Sliwa:2005:SCM

[ŚS05] Przemysław Śliwa and Wolfgang Schmid. Surveillance of the covariance matrix of multivariate nonlinear time series. Statistics: a Journal of Theoretical and Applied Statistics, 39(3): 221–246, 2005. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Shen:2014:NEL

[sS14] Pao sheng Shen. Nonparametric estimation with lefttruncated and right-censored data when the sample size before truncation is known. Statistics: a Journal of Theoretical and Applied Statistics, 48(2):315–326, 2014. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Sang:2015:SSM

[SS15a] Hailin Sang and Yan Sun. Simultaneous sparse model selection and coefficient estimation for heavy-tailed autoregressive processes. Statistics: a Journal of Theoretical and Applied Statistics, 49(1):187–208, 2015. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Schmidt:2015:OCR

[SS15b] Marius Schmidt and Rainer Schwabe. Optimal cutpoints for random observations. Statistics: a Journal of Theoretical and Applied Statistics, 49(6):1366–1381, 2015. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Shen:2015:FLI

[sS15c] Pao sheng Shen. Full likelihood inference in the Cox model with LTRC data when covariates are discrete. Statistics: a Journal of Theoretical and Applied Statistics, 49(3):602–613, 2015. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Saleh:2020:LTB

[SS20a] A. K. Md. Ehsanes Saleh and Shalabh. On Liu-type biased estimators in measurement error models. Statistics: a Journal of Theoretical and Applied Statistics, 54(6):1171–1213, 2020. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Shen:2020:QRD

[sS20b] Pao sheng Shen. Quantile regression for doubly truncated data. Statistics: a Journal of Theoretical and Applied Statistics, 54(4):649–666, 2020. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Sharifian:2021:JMM

[SS21] Nastaran Sharifian and Ehsan Bahrami Samani. A joint model for mixed longitudinal k-category inflation ordinal and continuous responses. Statistics: a Journal of Theoretical and Applied Statistics, 55(5):1180–1205, 2021. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Sieber:1976:GLO

[SSG76] Norbert Sieber, Hans-Jürgen Sebastian, and Dietrich Gnaudschun. Über geschlossene Lösungen zur Optimalregelung eines speziellen Versorgungssystems mittels dynamischer Optimierung. (German) [On closed-form solutions for optimal control of a special pension scheme using dynamic optimization]. Mathematische Operationsforschung und Statistik, 7(3): 453–469, 1976. CODEN MTOSBI. ISSN 0047-6277.

Shen:2017:ATM

[sSLMJ17] Pao sheng Shen, Yi Liu, Der-Pyng Maa, and Yeunjyr Ju. Analysis of transformation models with right-truncated data. Statistics: a Journal of Theoretical and Applied Statistics, 51(2):404–418, 2017. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Sunoj:2007:CDU

[SSM07] S. M. Sunoj, P. G. Sankaran, and S. S. Maya. Characterizations of distributions using log odds rate. Statistics: a Journal of Theoretical and Applied Statistics, 41(5):443–451, 2007. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Sunoj:2017:QBC

[SSN17] S. M. Sunoj, P. G. Sankaran, and N. Unnikrishnan Nair. Quantile-based cumulative Kullback–Leibler divergence. Statistics: a Journal of Theoretical and Applied Statistics, 52(1):1–17, 2017. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Sieber:1970:IGI

[SSS70] N. Sieber, H.-J Sebastian, and H. Stahn. Informationen. (German) [Information]. Mathematische Operationsforschung und Statistik, 1(3):249–251, 1970. CODEN MTOSBI. ISSN 0047-6277.

Sieber:1974:IGI

[SSS74] N. Sieber, H.-J Sebastian, and H. Stahn. Informationen. (German) [Information]. Mathematische Operationsforschung und Statistik, 5(6):541–543, 1974. CODEN MTOSBI. ISSN 0047-

Sperlich:2000:BRb

[SSS^{+}00] S. Sperlich, S. Sperlich, S. Sperlich, Ulrich Horst, S. Sperlich, S. Sperlich, and S. Sperlich. Book reviews. Statistics: a Journal of Theoretical and Applied Statistics, 34(3):259–260, 2000. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Sommerfield:1996:BR

[SSSS96] Volker Sommerfield, Volker Sommerfield, Volker Sommerfield, and Volker Sommerfield. Book reviews. Statistics: a Journal of Theoretical and Applied Statistics, 28(4):371, 1996. CO-DEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Spokoiny:1999:BR

[SSSS99] V. Spokoiny, S. Sperlich, S. Sperlich, and S. Sperlich. Book reviews. Statistics: a Journal of Theoretical and Applied Statistics, 33(1):93–94, 1999. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Shen:2019:CAM

[sSW19] Pao sheng Shen and Li Ning Weng. The Cox–Aalen model for left-truncated and mixed interval-censored data. Statistics: a Journal of Theoretical and Applied Statistics, 53(5):1152– 1167, 2019. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Schwartz:1975:ZLG

[ST75] B. Schwartz and R. Tichatschke. Über eine Zerlegungsmethode zur Lösung großdimensionierter linearer Optimierungsaufgaben. (German) [A decomposition method for solving largesized linear programs]. Mathematische Operationsforschung und Statistik, 6(1):15–31, 1975. CODEN MTOSBI. ISSN 0047-6277.

Sandvin:1981:NCT

[ST81a] Ottar Sandvin and Dag Tjøstheim. A numerical comparison of two criteria for determining the order of AR processes. Statistics (Berlin, DDR), 12(1):43–51, 1981. CODEN MOSSD5. ISSN 0323-3944.

Schmidt:1981:CAT

[ST81b] Wolfgang H. Schmidt and Rolf Thrum. Contributions toasymptotic theory in regression models with linear covariance structure. Statistics (Berlin, DDR), 12(2):243–269, 1981. CO-DEN MOSSD5. ISSN 0323-3944.

Schiele:1983:MEE

[ST83] K. Schiele and H. Toutenburg. Minimax estimation in econometric models a simulation approach. Statistics (Berlin, DDR), 14(2):217–233, 1983. CODEN MOSSD5. ISSN 0323-3944.

Srivenkataramana:1986:TAS

[ST86] T. Srivenkataramana and D. S. Tracy. Transformations after sampling. Statistics: a Journal of Theoretical and Applied Statistics, 17(4):597–608, 1986. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Szajowski:1987:MCS

[ST87] Krzysztof Szajowski and Stanisław Trubula. Minimax control of a stochastic system with the loss function dependent on parameter of disturbances. Statistics: a Journal of Theoretical and Applied Statistics, 18(1):151–165, 1987. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Stahlecker:1991:LER

[ST91] P. Stahlecker and G. Trenkler. Linear and ellipsoidal restrictions in linear regression. Statistics: a Journal of Theoretical and Applied Statistics, 22(2):163–176, 1991. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Stadtmuller:1995:NRV

[ST95] U. Stadtmüller and A. B. Tsybakov. Nonparametric recursive variance estimation. Statistics: a Journal of Theoretical and Applied Statistics, 27(1–2):55–63, 1995. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Sarkar:2023:ERS

[ST23] Mojammel Haque Sarkar and Manas Ranjan Tripathy. Estimating reciprocals of scale parameters of two exponential populations with common location and ordered scales using censored samples. Statistics: a Journal of Theoretical and Applied Statistics, 57(3):710–739, 2023. CO-DEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/02331888.2023.2209817.

Stadje:1984:NSM

[Sta84] W. Stadje. A note on sample means and variances of dependent normal variables. Statistics (Berlin, DDR), 15(2): 205–218, 1984. CODEN MOSSD5. ISSN 0323-3944.

Stadje:1988:ETD

[Sta88] W. Stadje. Estimation of tie distribution function of a location-scale family. Statistics: a Journal of Theoretical and Applied Statistics, 19(4):537–544, 1988. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Stadje:1990:SEP

[Sta90] W. Stadje. A sequential estimation procedure for the parameter of an exponential distribution. Statistics: a Journal of Theoretical and Applied Statistics, 21(2):239–250, 1990. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Staudte:2017:STC

[Sta17] Robert G. Staudte. The shapes of things to come: probability density quantiles. Statistics: a Journal of Theoretical and Applied Statistics, 51(4):782–800, 2017. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Staudte:2020:EGF

[Sta20] R. G. Staudte. Evidence for goodness of fit in Karl Pearson chi-squared statistics. Statistics: a Journal of Theoretical and Applied Statistics, 54(6):1287–1310, 2020. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Steiner:1976:KSG

[Ste76] Jürgen Steiner. Zur Konstruktion von Schnitten für gemischtganzzahlige LO-Probleme. (German) [On the construction of cuts for mixed integer LO-problems]. Mathematische Operationsforschung und Statistik, 7(6):819–829, 1976. CODEN MTOSBI. ISSN 0047-6277.

Steinebach:1995:VEB

[Ste95] Josef Steinebach. Variance estimation based on invariance principles. Statistics: a Journal of Theoretical and Applied Statistics, 27(1–2):15–25, 1995. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Stepniak:1998:CNL

[Ste98] Czeslaw Stepniak. Comparing normal linear experiments and transformation of observations. Statistics: a Journal of Theoretical and Applied Statistics, 30(4):279–289, 1998. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Stier:1978:VZF

[Sti78] Winfried Stier. Zur Verknüpfung von “Zeitbereichsanalyse” und “Frequenzbereichsanalyse” bei Saisonbereinigungsverfahren — ein Beitrag zur Zeitreihenanalyse. (German) [On linkage of “time-domain analysis” and “frequency-domain analysis” in seasonal adjustment procedures — a contribution to time series analysis]. Statistics (Berlin, DDR), 9(3): 369–381, 1978. CODEN MOSSD5. ISSN 0323-3944.

Stoyan:1971:PBD

[Sto71] Dietrich Stoyan. Probleme der Bedienungs- und Dammtheorie bei “unvollständigen Informationen”. (German) [Problems of operating and dam theory with “incomplete information”]. Mathematische Operationsforschung und Statistik, 2 (1):51–60, 1971. CODEN MTOSBI. ISSN 0047-6277.

Stoyan:1972:SEW

[Sto72a] Dietrich Stoyan. Ein Stetigkeitssatz für einlinige Wartemodelle der Bedienungstheorie. (German) [A discontinuity rate for single-line waiting models of operation theory]. Mathematische Operationsforschung und Statistik, 3(2):103–111, 1972. CODEN MTOSBI. ISSN 0047-6277.

Stoyan:1972:VBD

[Sto72b] Dietrich Stoyan. Verlustsysteme aus der Bedienungs- und Dammtheorie bei “unvollstandigen Informationen” über verteilungsfunktionen. (German) [Loss systems from the operating-systems and dam theory with “incomplete information” about distribution functions]. Mathematische Operationsforschung und Statistik, 3(1):49–59, 1972. CODEN MTOSBI. ISSN 0047-6277.

Stoyan:1973:MSM

[Sto73] Helga Stoyan. Monotonie- und Stetigkeitseigenschaften mehrliniger Wartesysteme der Bedienungstheorie. (German) [Monotonicity and continuity properties of more straightforward response systems of operation theory]. Mathematische Operationsforschung und Statistik, 4(2):155–163, 1973. CO-DEN MTOSBI. ISSN 0047-6277.

Stoyan:1974:SBM

[Sto74] Dietrich Stoyan. Some bounds for many-server systems GI/G/s. Mathematische Operationsforschung und Statistik, 5(2):117–129, 1974. CODEN MTOSBI. ISSN 0047-6277.

Stoyan:1976:CRS

[Sto76a] D. Stoyan. A critical remark on a system approximation in queueing theory. Mathematische Operationsforschung und Statistik, 7(6):953–956, 1976. CODEN MTOSBI. ISSN 0047-6277.

Stoyan:1976:AMG

[Sto76b] Dietrich Stoyan. Approximations for M/G/s queues. Mathematische Operationsforschung und Statistik, 7(4):587–594, 1976. CODEN MTOSBI. ISSN 0047-6277.

Stone:1978:CVR

[Sto78] M. Stone. Cross-validation: a review. Statistics (Berlin, DDR), 9(1):127–139, 1978. CODEN MOSSD5. ISSN 0323-3944.

Stone:1979:PTG

[Sto79] M. Stone. A pictorial treatment of generalized inverses for statistics based on dual vector spaces. Statistics (Berlin, DDR), 10(1):3–17, 1979. CODEN MOSSD5. ISSN 0323-3944.

Stoyan:1983:BR

[Sto83a] D. Stoyan. Book reviews. Statistics (Berlin, DDR), 14(2): 234, 1983. CODEN MOSSD5. ISSN 0323-3944.

Stoyan:1983:IBV

[Sto83b] D. Stoyan. Inequalities and bounds for variances of point processes and fibre processes. Statistics (Berlin, DDR), 14 (3):409–419, 1983. CODEN MOSSD5. ISSN 0323-3944.

Stoyan:1984:FSF

[Sto84] D. Stoyan. Further stereological formulae for spatial fibre processes. Statistics (Berlin, DDR), 15(3):421–428, 1984. CO-DEN MOSSD5. ISSN 0323-3944.

Stoyan:1988:BR

[Sto88a] D. Stoyan. Book review. Statistics: a Journal of Theoretical and Applied Statistics, 19(1):66, 1988. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Stoyan:1988:TPP

[Sto88b] D. Stoyan. Thinnings of point processes and their use in the statistical analysis of a settlement pattern with deserted tillages. Statistics: a Journal of Theoretical and Applied Statistics, 19(1):45–56, 1988. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Stoyan:1989:BR

[Sto89] D. Stoyan. Book reviews. Statistics: a Journal of Theoretical and Applied Statistics, 20(3):416, 1989. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Stoyan:1990:SFR

[Sto90] D. Stoyan. Stereological formulae for a random system of nonintersecting spheres. Statistics: a Journal of Theoretical and Applied Statistics, 21(1):131–136, 1990. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Stoyan:1991:DAM

[Sto91] D. Stoyan. Describing the anisotropy of marked planner point process. Statistics: a Journal of Theoretical and Applied Statistics, 22(3):449–462, 1991. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Stoyan:1993:SSA

[Sto93] D. Stoyan. A spatial statistical analysis of a work of art: did Hans Arp make a “truly random” collage? Statistics: a Journal of Theoretical and Applied Statistics, 24(1):71–80, 1993. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Stoyan:1994:CFP

[Sto94] Dietrich Stoyan. Caution with “fractal” point patterns! Statistics: a Journal of Theoretical and Applied Statistics, 25(3):267–270, 1994. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Stoimenova:1996:TTP

[Sto96] E. Stoimenova. Testing for a trend in partial rankings. Statistics: a Journal of Theoretical and Applied Statistics, 27(3–4): 339–343, 1996. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Stoyan:1998:BR

[Sto98] D. Stoyan. Book reviews. Statistics: a Journal of Theoretical and Applied Statistics, 31(1):89, 1998. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Streller:1982:SPE

[Str82] Arnfried Streller. On stochastic processes with an embedded marked point process. Statistics (Berlin, DDR), 13(4):561– 576, 1982. CODEN MOSSD5. ISSN 0323-3944.

Streit:1995:TWI

[Str95] Franz Streit. Testing whether interpenetration is possible for the grains of a random set. Statistics: a Journal of Theoretical and Applied Statistics, 26(4):355–364, 1995. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Schwetlick:1985:GNL

[STS85] H. Schwetlick, Y. Tilleb, and W. Schellong. Gauss– Newton-like methods for nonlinear least squares with equality constraints-local convergence and applications. Statistics: a Journal of Theoretical and Applied Statistics, 16(2):167–178, 1985. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Stute:1990:BLC

[Stu90] Winfried Stute. Bootstrap of the linear correlation model. Statistics: a Journal of Theoretical and Applied Statistics, 21(3):433–436, 1990. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Stute:1996:CPU

[Stu96] Winfried Stute. Changepoint problems under random censorship. Statistics: a Journal of Theoretical and Applied Statistics, 27(3–4):255–266, 1996. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Styszynski:1975:SVN

[Sty75] Antoni Styszyński. An n-silent-vs.-noisy duel with equal accuracy functions. Mathematische Operationsforschung und Statistik, 6(3):367–383, 1975. CODEN MTOSBI. ISSN 0047-6277.

Sakamoto:2004:ECP

[STY04] Yuji Sakamoto, Yoshikazu Takada, and Nakahiro Yoshida. Expansions of the coverage probabilities of prediction region based on a shrinkage estimator. Statistics: a Journal of Theoretical and Applied Statistics, 38(5):381–390, 2004. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Sankaran:1993:CPR

[SU93] P. G. Sankaran and N. Unnikrishnan. Characterizations by properties of residual life distributions. Statistics: a Journal of Theoretical and Applied Statistics, 24(3):245–251, 1993. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Sarda:1989:EDF

[SV89] P. Sarda and P. Vieu. Empirical distribution function for mixing random variables. application in nonparametric hazard estimation. Statistics: a Journal of Theoretical and Applied Statistics, 20(4):559–571, 1989. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Stummer:2010:DFM

[SV10] Wolfgang Stummer and Igor Vajda. On divergences of finite measures and their applicability in statistics and information theory. Statistics: a Journal of Theoretical and Applied Statistics, 44(2):169–187, 2010. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Sinova:2018:STI

[SV18] Beatriz Sinova and Stefan Van Aelst. A spatial-type intervalvalued median for random intervals. Statistics: a Journal of Theoretical and Applied Statistics, 52(3):479–502, 2018. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Sasvari:1985:WAD

[SW85] Z. Sasvári and W. Wolf. Über die Wiederherstellung der Ausgangsverteilung durch die Statistik X_{1} − X_{3}, X_{2} − X_{3}. (German) [On the restoration of the initial distribution through the statistics X_{1} −X_{3}, X_{2} −X_{3}]. Statistics: a Journal of Theoretical and Applied Statistics, 16(2):233–241, 1985. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Siefert:1989:BR

[SW89a] B. Siefert and W. Wolf. Book reviews. Statistics: a Journal of Theoretical and Applied Statistics, 20(4):530, 1989. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Stone:1989:BSO

[SW89b] M. Stone and J. L. Wang. A Bayes, sweep-operational construction for Rao’s unified theory of linear estimation. Statistics: a Journal of Theoretical and Applied Statistics, 20(1): 3–11, 1989. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Shen:2016:KME

[SW16] Aiting Shen and Xinghui Wang. Kaplan–Meier estimator and hazard estimator for censored negatively superadditive dependent data. Statistics: a Journal of Theoretical and Applied Statistics, 50(2):377–388, 2016. CO-DEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/02331888.2015.1038269.

Sun:2020:ANM

[SX20] Gaoming Sun and Junshan Xie. Asymptotic normality and moderate deviation principle for high-dimensional likelihood ratio statistic on block compound symmetry covariance structure. Statistics: a Journal of Theoretical and Applied Statistics, 54(1):114–134, 2020. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Sacks:1981:VEA

[SY81] Jerome Sacks and Donald Ylvisaker. Variance estimation for approximately linear models. Statistics (Berlin, DDR), 12(2): 147–162, 1981. CODEN MOSSD5. ISSN 0323-3944.

Sevil:2022:DBE

[SY22] Yusuf Can Sevil and Tugba Ozkal Yildiz. Design-based estimators of the distribution function in ranked set sampling with an application. Statistics: a Journal of Theoretical and Applied Statistics, 56(4):891–918, 2022. CO-DEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/02331888.2022.2081690.

Shakhatreh:2016:BGL

[SYM16] Mohammed K. Shakhatreh, Abdullahi Yusuf, and Abdel-Razzaq Mugdadi. The beta generalized linear exponential distribution. Statistics: a Journal of Theoretical and Applied Statistics, 50(6):1346–1362, 2016. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Sun:2022:DRE

[SYW22] Siying Sun, Yaohong Yang, and Lei Wang. Dimensionreduced empirical likelihood estimation and inference for mestimators with nonignorable nonresponse. Statistics: a Journal of Theoretical and Applied Statistics, 56(2):243–270, 2022. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Schmidt:1986:THN

[SZ86] Wolfgrang H. Schmidt and Silvelyn Zwanzig. Testing hypotheses in nonlinear regression for nonnormal distributions. Statistics: a Journal of Theoretical and Applied Statistics, 17(4):483–503, 1986. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Sinha:1997:EEM

[SZ97] Bikas K. Sinha and Ryszard Zieliński. Estimating P {X > Y } in exponential model. Statistics: a Journal of Theoretical and Applied Statistics, 29(4):299–316, 1997. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Sohrabi:2017:BMV

[SZ17] Maryam Sohrabi and Mahmoud Zarepour. Bootstrapping the mean vector for the observations in the domain of attraction of a multivariate stable law. Statistics: a Journal of Theoretical and Applied Statistics, 52(1):50–63, 2017. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Szczesny:1979:ICM

[Szc79] W. Szczesny. Influence curve of the monotonic dependence function for binormal distributions. Statistics (Berlin, DDR), 10(4):541–550, 1979. CODEN MOSSD5. ISSN 0323-3944.

Thomas-Agnan:1990:FSN

[TA90] Christine Thomas-Agnan. A family of splines for nonparametric regression. and their relationships with kriging. Statistics: a Journal of Theoretical and Applied Statistics, 21(4):533–548, 1990. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Tavangar:2011:SRC

[TA11] Mahdi Tavangar and Majid Asadi. Some results on conditional expectations of lower record values. Statistics: a Journal of Theoretical and Applied Statistics, 45(3):237–255, 2011. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Taavoni:2021:HDG

[TA21] M. Taavoni and M. Arashi. High-dimensional generalized semiparametric model for longitudinal data. Statistics: a Journal of Theoretical and Applied Statistics, 55(4):831–850, 2021. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Tabakan:2013:PDB

[Tab13] Gülin Tabakan. Performance of the difference-based estimators in partially linear models. Statistics: a Journal of Theoretical and Applied Statistics, 47(2):329–347, 2013. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Takacs:1986:EPP

[Tak86] R. Takacs. Estimator for the pair-potential of a Gibbsian point process. Statistics: a Journal of Theoretical and Applied Statistics, 17(3):429–433, 1986. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Takagi:2009:SOA

[Tak09] Y. Takagi. Second-order asymptotic comparison of estimators under ϕ-divergence loss. Statistics: a Journal of Theoretical and Applied Statistics, 43(6):571–581, 2009. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Tammer:1974:DHS

[Tam74a] Elisabeth-Charlotte Tammer. Dualitätstheorie für hyperbolische und stückweise lineare konvexe Optimierungsprobleme. (German) [Duality theory for hyperbolic and piecewise linear convex optimization problems]. Mathematische Operationsforschung und Statistik, 5(2):93–108, 1974. CODEN MTOSBI. ISSN 0047-6277.

Tammer:1974:AQO

[Tam74b] Klaus Tammer. Die Abhängigkeit eines quadratischen Optimierungsproblems von einem Parameter in der Zielfunktion. (German) [The dependence of a quadratic optimization problem on a parameter in the objective function]. Mathematische Operationsforschung und Statistik, 5(7–8):573–590, 1974. CO-DEN MTOSBI. ISSN 0047-6277.

Tammer:1976:MAE

[Tam76] Klaus Tammer. Möglichkeiten zur Anwendung der Erkenntnisse der parametrischen Optimierung für die Lösung indefiniter quadratischer Optimierungsprobleme. (German) [Opportunities for applying the findings of the parametric optimization for the solution of indefinite quadratic programming problems]. Mathematische Operationsforschung und Statistik, 7(2):209–222, 1976. CODEN MTOSBI. ISSN 0047-6277.

Tang:1991:EMP

[Tan91] S. Y. Tang. Estimation of mixed parametric functions in the variance components model. Statistics: a Journal of Theoretical and Applied Statistics, 22(4):517–524, 1991. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Tanaka:2011:SCA

[Tan11] H. Tanaka. Sufficient conditions for the admissibility under the LINEX loss function in non-regular case. Statistics: a Journal of Theoretical and Applied Statistics, 45(2):199–208, 2011. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Tang:2014:REF

[Tan14] Qingguo Tang. Robust estimation for functional coefficient regression models with spatial data. Statistics: a Journal of Theoretical and Applied Statistics, 48(2):388–404, 2014. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Targhetta:1988:ALB

[Tar88] Luisa M. Targhetta. On the attainment of a lower bound for the Bates risk in estimating a parametric function. Statistics: a Journal of Theoretical and Applied Statistics, 19(2):233– 239, 1988. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Tasche:2021:CS

[Tas21] Dirk Tasche. Calibrating sufficiently. Statistics: a Journal of Theoretical and Applied Statistics, 55(6):1356–1386, 2021. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Tautenhahn:1984:VKQ

[Tau84] U. Tautenhahn. Verbesserungen der Kleinste-Quadrate-Schätzung durch Schätzungen vom Ridge-typ und vom Steintyp. (German) [Improvements in the least-squares approximation by estimates of Ridge-type and Stein-type]. Statistics (Berlin, DDR), 15(3):337–354, 1984. CODEN MOSSD5. ISSN 0323-3944.

Taylor:1983:CCD

[Tay83] J. M. Taylor. Comparisons of certain distribution functions. Statistics (Berlin, DDR), 14(3):397–408, 1983. CODEN MOSSD5. ISSN 0323-3944.

Tan:1978:EES

[TB78] Peter Tan and Ben Bernholtz. Equivariant estimators for structural models. Statistics (Berlin, DDR), 9(3):351–356, 1978. CODEN MOSSD5. ISSN 0323-3944.

Towhidi:2001:BMN

[TB01] M. Towhidi and J. Behboodian. Bayesian multivariate normal analysis under the extended reflected normal loss function. Statistics: a Journal of Theoretical and Applied Statistics, 35(3):215–223, 2001. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Toutenburg:1983:BR

[TBJ^{+}83] H. Toutenburg, H. Bandemer, J. Jureckova, O. Bunke, and K. Hoffmann. Book reviews. Statistics (Berlin, DDR), 14(4): 632, 1983. CODEN MOSSD5. ISSN 0323-3944.

Tharmaratnam:2013:CRV

[TC13] Kukatharmini Tharmaratnam and Gerda Claeskens. A comparison of robust versions of the AIC based on M-, S- and MM-estimators. Statistics: a Journal of Theoretical and Applied Statistics, 47(1):216–235, 2013. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Teles:2022:UAT

[TC22] Paulo Teles and Wai Sum Chan. The use of aggregate time series for testing conditional heteroscedasticity. Statistics: a Journal of Theoretical and Applied Statistics, 56(6):1242– 1269, 2022. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-

4910 (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/02331888.2022.2134386.

Trommer:1988:BR

[TDB^{+}88] R. Trommer, B. Droge, O. Bunke, O. Bunke, U. Kü, H. Bandemer, B. Droge, Kurt Hoffmann, Kurt Hoffmann, H. Bandemer, J. Franz, M. Richter, D. Stoyan, R. Hansel, D. Stoyan, J. Polzehl, J. Polzehl, J. Polzehl, O. Bunke, and K. H. Fichtner. Book review. Statistics: a Journal of Theoretical and Applied Statistics, 19(1):133–159, 1988. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Tsai:2013:RTA

[TDM13] Yu-Ling Tsai, Debbie J. Dupuis, and Duncan J. Murdoch. A robust test for asymptotic independence of bivariate extremes. Statistics: a Journal of Theoretical and Applied Statistics, 47(1):172–183, 2013. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Tse:2008:OAL

[TDY08] Siu-Keung Tse, Chang Ding, and Chunyan Yang. Optimal accelerated life tests under interval censoring with random removals: the case of Weibull failure distribution. Statistics: a Journal of Theoretical and Applied Statistics, 42(5):435–451, 2008. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Talebi:2015:CSE

[TE15] Hooshang Talebi and Nabaz Esmailzadeh. Comparing search and estimation performances of designs based on a compound criterion. Statistics: a Journal of Theoretical and Applied Statistics, 49(3):629–637, 2015. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Terno:1972:AKM

[Ter72] Johannes Terno. Algorithmen für das klassische Maschinenbelegungsproblem. (German) [Algorithms for the classical machine assignment problem]. Mathematische Operationsforschung und Statistik, 3(3):195–201, 1972. CODEN MTOSBI. ISSN 0047-6277.

Terasvirta:1975:NPS

[Ter75] Timo Teräsvirta. A note on predicting with seemingly unrelated regression equations. Mathematische Operationsforschung und Statistik, 6(5):709–711, 1975. CODEN MTOSBI. ISSN 0047-6277.

Tat:2021:NTM

[TF21] Samaneh Tat and Mohammad Reza Faridrohani. A new type of multivariate records: depth-based records. Statistics: a Journal of Theoretical and Applied Statistics, 55(2):296–320, 2021. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Fang:2001:GMD

[tFYK01] Kai tai Fang, Zhenhai Yang, and Samuel Kotz. Generation of multivariate distributions by vertical density representation. Statistics: a Journal of Theoretical and Applied Statistics, 35(3):281–293, 2001. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Tang:1987:TBI

[TG87] Jen Tang and A. K. Gupta. On the Type-B integral equation and the distribution of Wilks’ statistic for testing independence of several groups of variables. Statistics: a Journal of Theoretical and Applied Statistics, 18(3):379–387, 1987. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Thabane:2004:MVG

[TH04] L. Thabane and M. Safiul Haq. On the matrix-variate generalized hyperbolic distribution and its Bayesian applications. Statistics: a Journal of Theoretical and Applied Statistics, 38(6):511–526, 2004. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Triantafyllopoulos:2008:PMV

[TH08] K. Triantafyllopoulos and P. J. Harrison. Posterior mean and variance approximation for regression and time series problems. Statistics: a Journal of Theoretical and Applied Statistics, 42(4):329–350, 2008. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Thamelt:1975:VLN

[Thä75] Wolfgang Thämelt. Eine Verallgemeinerung des Lemmas von Neyman–Pearson. (German) [A generalization of the lemma of Neyman–Pearson]. Mathematische Operationsforschung und Statistik, 6(6):933–937, 1975. CODEN MTOSBI. ISSN 0047-6277.

Thamelt:1976:LHL

[Thä76] Wolfgang Thämelt. Über Lösbarkeitseigenschaften homogener linearer Ungleichungen in halbgeordneten Räumen. (German) [On solvability properties of homogeneous linear inequalities in partially ordered spaces]. Mathematische Operationsforschung und Statistik, 7(2):281–294, 1976. CODEN MTOSBI. ISSN 0047-6277.

Theodorescu:1972:RP

[The72] Radu Theodorescu. Random programs. Mathematische Operationsforschung und Statistik, 3(1):19–47, 1972. CODEN MTOSBI. ISSN 0047-6277.

Tammer:1976:RGB

[THN^{+}76] K. Tammer, H. Hollatz, D. Niepage, M. B. Malyutov, H. J. Engelbert, W. Dück, H. Körth, N. Sieber, J. Nietzsch, N. Sieber, and W. S. Wittig. Rezensionen. (German) [Book reviews]. Mathematische Operationsforschung und Statistik, 7(3):481–492, 1976. CODEN MTOSBI. ISSN 0047-6277.

Teimouri:2013:CEM

[THN13] Mahdi Teimouri, Seyed M. Hoseini, and Saralees Nadarajah. Comparison of estimation methods for the Weibull distribution. Statistics: a Journal of Theoretical and Applied Statistics, 47(1):93–109, 2013. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Thomsen:1978:FTT

[Tho78] Wolfgang Thomsen. On a Fubini-type theorem and its application in game theory. Statistics (Berlin, DDR), 9(3):419–423, 1978. CODEN MOSSD5. ISSN 0323-3944.

Thompson:1980:MLE

[Tho80] Robin Thompson. Maximum likelihood estimation of variance components. Statistics (Berlin, DDR), 11(4):545–561, 1980. CODEN MOSSD5. ISSN 0323-3944.

Thomas:1984:EPI

[Tho84] C. Thomas. Extremum properties of the intersection densities of stationary Poisson hyperplane processes. Statistics (Berlin, DDR), 15(3):443–449, 1984. CODEN MOSSD5. ISSN 0323-3944.

Thomas:2013:DPI

[Tho13] Naiju M. Thomas. Distribution of products of independently distributed pathway random variables. Statistics: a Journal of Theoretical and Applied Statistics, 47(4):861–875, 2013. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Thrum:1977:OPS

[Thr77] R. Thrum. Optimal prediction of a stochastic process by a finite number of observations. Statistics (Berlin, DDR), 8(4): 457–458, 1977. CODEN MOSSD5. ISSN 0323-3944.

Thrum:1979:LEP

[Thr79] R. Thrum. Linear estimation and prediction in stochastic processes. Statistics (Berlin, DDR), 10(2):333–346, 1979. CO-DEN MOSSD5. ISSN 0323-3944.

TiagoDeOliveira:1991:ISB

[Tia91] J. Tiago De Oliveira. Identifiability of some bivariate extremes models. Statistics: a Journal of Theoretical and Applied Statistics, 22(1):149–154, 1991. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Tijms:1976:SPC

[Tij76] H. C. Tijms. On a switch-over policy for controlling the workload in a queueing system with two constant service rates and fixed switch-over costs. Mathematische Operationsforschung und Statistik, 7(6):933–943, 1976. CODEN MTOSBI. ISSN 0047-6277.

Tiku:2009:NPA

[Tik09] Moti L. Tiku. A note on the paper by Ahmed Hossain and Andrew R. Willan. Statistics: a Journal of Theoretical and Applied Statistics, 43(3):313–314, 2009. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic). See [HW07].

Tiku:2010:MDU

[TIQ10] Moti L. Tiku, M. Qamarul Islam, and Sahar B. Qumsiyeh. Mahalanobis distance under non-normality. Statistics: a Journal of Theoretical and Applied Statistics, 44(3):275–290, 2010. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Tittel:1970:MMP

[Tit70] Bernd Tittel. Zur mathematischen Modellierung der Planung der Arbeitskräfteentwicklung. (German) [On the mathematical modeling of the planning of workforce development]. Mathematische Operationsforschung und Statistik, 1(2):87– 104, 1970. CODEN MTOSBI. ISSN 0047-6277.

Titterington:1980:GAD

[Tit80] D. M. Titterington. Geometric approaches to design of experiment. Statistics (Berlin, DDR), 11(1):151–163, 1980. CO-DEN MOSSD5. ISSN 0323-3944.

Titak:1983:STS

[Tit83] Pham Dinh Titak. A survey of time series analysis through parametric models. Statistics (Berlin, DDR), 14(4):603–631, 1983. CODEN MOSSD5. ISSN 0323-3944.

Titterington:1990:SRR

[Tit90] D. M. Titterington. Some recent research in the analysis of mixture distributions. Statistics: a Journal of Theoretical and Applied Statistics, 21(4):619–641, 1990. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Tahata:2008:MDC

[TIT08] Kouji Tahata, Toshiya Iwashita, and Sadao Tomizawa. Measure of departure from conditional marginal homogeneity for square contingency tables with ordered categories. Statistics: a Journal of Theoretical and Applied Statistics, 42(5): 453–466, 2008. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Tjostheim:1996:MDT

[Tjø96] Dag Tjøstheim. Measures of dependence and tests of independence. Statistics: a Journal of Theoretical and Applied Statistics, 28(3):249–284, 1996. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Tiwari:1989:TIF

[TJZ89a] Ram C. Tiwari, S. Rao Jammalamadaka, and Jyoti N. Zalkikar. Testing an increasing failure rate average distribution with censored data. Statistics: a Journal of Theoretical and Applied Statistics, 20(2):279–286, 1989. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Trwari:1989:CLS

[TJZ89b] Ram C. Trwari, S. Rao Jammalamadaka, and Jyoti N. Zalkikar. A class of large sample tests for bivariate exponentiality versus BIFRA and BNBU(t_{0}, t_{0}) alternatives. Statistics: a Journal of Theoretical and Applied Statistics, 20(2):297–304, 1989. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Tovissode:2022:GSC

[TK22] Chénangnon Frédéric Tovissodé and Romain Glèlè Kakaï. A generalized stochastic condensation mechanism for inducing underdispersion in count models. Statistics: a Journal of Theoretical and Applied Statistics, 56(4):823–843, 2022. CO-DEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/02331888.2022.2084095.

Trong:2015:NRS

[TKTM15] Dang Duc Trong, To Duc Khanh, Nguyen Huy Tuan, and Nguyen Dang Minh. Nonparametric regression in a statistical modified Helmholtz equation using the Fourier spectral regularization. Statistics: a Journal of Theoretical and Applied Statistics, 49(2):267–290, 2015. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Torrado:2015:LRC

[TL15] Nuria Torrado and Rosa E. Lillo. Likelihood ratio comparisons among spacings related to both one or two samples. Statistics: a Journal of Theoretical and Applied Statistics, 49(4):831–841, 2015. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Tracy:1994:UEU

[TM94] D. S. Tracy and B. Mukhopadhyay. On UMVU-estimation under randomized response models. Statistics: a Journal of Theoretical and Applied Statistics, 25(2):173–175, 1994. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Terpstra:2000:HEW

[TMN00] Jeffrey T. Terpstra, Joseph W. McKean, and Joshua D. Naranjo. Highly efficient weighted Wilcoxon estimates for autoregression. Statistics: a Journal of Theoretical and Applied Statistics, 35(1):45–80, 2000. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Tomizawa:2005:PDT

[TMY05] Sadao Tomizawa, Nobuko Miyamoto, and Shunsuke Yamane. Power-divergence-type measure of departure from diagonalsparameter symmetry for square contingency tables with ordered categories. Statistics: a Journal of Theoretical and Applied Statistics, 39(2):107–115, 2005. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Taye:2008:SFT

[TN08] Girma Taye and Peter M. Njuho. Smoothing fertility trends in agricultural field experiments. Statistics: a Journal of Theoretical and Applied Statistics, 42(3):275–289, 2008. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Tomizawa:1990:DAI

[Tom90] Sadao Tomizawa. A decomposition of Akaiee’s information criterion for the conditional symmetry model. Statistics: a Journal of Theoretical and Applied Statistics, 21(4):569–575, 1990. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Topfer:1983:BR

[Töp83] W. Töpfer. Book reviews. Statistics (Berlin, DDR), 14(3): 503–504, 1983. CODEN MOSSD5. ISSN 0323-3944.

Torrado:2017:SCB

[Tor17] Nuria Torrado. Stochastic comparisons between extreme order statistics from scale models. Statistics: a Journal of Theoretical and Applied Statistics, 51(6):1359–1376, 2017. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Toutenburg:1970:OVE

[Tou70a] Helge Toutenburg. Optimale Vorhersage von endogenen Variablen in einem linearen System von strukturellen Gleichungen. (German) [Optimal forecasts of endogenous variables in a linear system of structural equations]. Mathematische Operationsforschung und Statistik, 1(1):69–75, 1970. CODEN MTOSBI. ISSN 0047-6277.

Toutenburg:1970:VIA

[Tou70b] Helge Toutenburg. Vorhersage im allgemeinen linearen Regressionsmodell mit stochastischen Regressoren. (German) [Prediction in the general linear regression model with stochastic regressors]. Mathematische Operationsforschung und Statistik, 1(2):105–116, 1970. CODEN MTOSBI. ISSN 0047-6277.

Toutenburg:1975:MLE

[Tou75] H. Toutenburg. Minimax-linear estimation (MMLE) and 2phase MMLE in a restricted linear regression model. Mathematische Operationsforschung und Statistik, 6(5):703–708, 1975. CODEN MTOSBI. ISSN 0047-6277.

Troutt:1997:FVT

[TP97] Marvin D. Troutt and W. K. Pang. A further VDR-type density representation based on the Box–Muller method. Statistics: a Journal of Theoretical and Applied Statistics, 29(1): 101–108, 1997. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Tanaka:2015:IEG

[TPL15] Hidekazu Tanaka, Nabendu Pal, and Wooi K. Lim. On improved estimation of a gamma shape parameter. Statistics: a Journal of Theoretical and Applied Statistics, 49(1):84–97, 2015. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Tanaka:2019:SOI

[TPL19] Hidekazu Tanaka, Nabendu Pal, and Wooi K. Lim. On second-order improved estimation of a gamma scale parameter. Statistics: a Journal of Theoretical and Applied Statistics, 53(5):968–989, 2019. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Tripathi:2017:ELP

[TPSR17] Yogesh Mani Tripathi, Constantinos Petropoulos, Farha Sultana, and Manoj Kumar Rastogi. Estimating a linear parametric function of a doubly censored exponential distribution. Statistics: a Journal of Theoretical and Applied Statistics, 52 (1):99–114, 2017. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Toutenburg:1978:MLT

[TR78] H. Toutenburg and B. Roedek. Minimax-linear and Theil estimator for restrained regression coefficients. Statistics (Berlin, DDR), 9(4):499–505, 1978. CODEN MOSSD5. ISSN 0323-3944.

Troltzsch:1975:EDL

[Trö75] Fredi Tröltzsch. Existenz- und Dualitätsaussagen für lineare Optimierungsprobleme in reflexiven Banach-Räumen. (German) [Existence and duality statements for linear optimization problems in reflexive Banach spaces]. Mathematische Operationsforschung und Statistik, 6(6):901–912, 1975. CODEN MTOSBI. ISSN 0047-6277.

Troutt:1991:TDD

[Tro91] Marvin D. Troutt. A theorem on the density of the density ordinate and an alternative interpretation of the Box– Muller method. Statistics: a Journal of Theoretical and Applied Statistics, 22(3):463–466, 1991. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Troutt:1993:VDR

[Tro93] Marvin D. Troutt. Vertical density representation and a further remark on the Box–Muller method. Statistics: a Journal of Theoretical and Applied Statistics, 24(1):81–83, 1993. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Trottier:2000:QSM

[Tro00] Catherine Trottier. A quasi-score marginal approach in generalized linear mixed models. Statistics: a Journal of Theoretical and Applied Statistics, 33(4):291–308, 2000. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Truquet:2018:ESE

[Tru18] Lionel Truquet. Efficient semiparametric estimation in timevarying regression models. Statistics: a Journal of Theoretical and Applied Statistics, 52(3):590–618, 2018. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Trybula:1986:SCE

[Try86] S. Trybula. Simultaneous control and estimation of linear stochastic systems with unknown parameters of disturbances. Statistics: a Journal of Theoretical and Applied Statistics, 17(3):365–376, 1986. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Trybula:1988:ESP

[Try88] Stanisław Trybula. Estimation of the same parameters under different data. Statistics: a Journal of Theoretical and Applied Statistics, 19(2):207–213, 1988. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Trybuza:1989:MLC

[Try89] S. Trybuza. Minimax law of control for a multidimensional, time continuous, linear stochastic system. Statistics: a Journal of Theoretical and Applied Statistics, 20(2):319–330, 1989. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Trybula:1991:EDP

[Try91a] Stanisław Trybula. Estimation of the difference of parameters. Statistics: a Journal of Theoretical and Applied Statistics, 22(2):199–204, 1991. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Trybula:1991:SEP

[Try91b] Stanisław Trybula. Systematic estimation of parameters of multinomial and multivariate hypergeometric distribution. Statistics: a Journal of Theoretical and Applied Statistics, 22 (1):59–67, 1991. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Trenkler:1987:QME

[TS87] G. Trenkler and P. Stahlechker. Quasi minimax estimation in the linear regression model. Statistics: a Journal of Theoretical and Applied Statistics, 18(2):219–226, 1987. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Tardivel:2020:SEL

[TSC20] Patrick J. C. Tardivel, Rémi Servien, and Didier Concordet. Simple expressions of the LASSO and SLOPE estimators in low-dimension. Statistics: a Journal of Theoretical and Applied Statistics, 54(2):340–352, 2020. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Tabrizi:2020:IPL

[TSG20] Elham Tabrizi, Ehsan Bahrami Samani, and Mojtaba Ganjali. Identifiability of parameters in longitudinal correlated Poisson and inflated beta regression model with non-ignorable missing mechanism. Statistics: a Journal of Theoretical and Applied Statistics, 54(3):524–543, 2020. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Tichatschke:1976:RGB

[TSL^{+}76] R. Tichatschke, M. Schoch, K. Lommatzsch, D. Oelschlägel, H. Hollatz, and G. Lippold. Rezensionen. (German) [Book reviews]. Mathematische Operationsforschung und Statistik, 7(2):321–327, 1976. CODEN MTOSBI. ISSN 0047-6277.

Taniguchi:2007:IEA

[TSO07] Masanobu Taniguchi, Hiroshi Shiraishi, and Hiroaki Ogata. Improved estimation for the autocovariances of a Gaussian stationary process. Statistics: a Journal of Theoretical and Applied Statistics, 41(4):269–277, 2007. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Tian:2010:ABD

[TT10a] Yongge Tian and Zhe Tian. On additive and block decompositions of WLSEs under a multiple partitioned regression model. Statistics: a Journal of Theoretical and Applied Statistics, 44(4):361–379, 2010. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Tsao:2010:PET

[TT10b] C. Andy Tsao and Yu-Ling Tseng. Power estimation for testing normal means. Statistics: a Journal of Theoretical and Applied Statistics, 44(3):247–259, 2010. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Tuerde:2022:BSA

[TT22] Mulati Tuerde and Niansheng Tang. Bayesian semiparametric approach to quantile nonlinear dynamic factor analysis models with mixed ordered and nonignorable missing data. Statistics: a Journal of Theoretical and Applied Statistics, 56(5): 1166–1192, 2022. CODEN MOSSD5. ISSN 0233-1888 (print),

1029-4910 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/02331888.2022.2121399.

Tuan:1981:NED

[Tua81] Pham Dinh Tuan. Nonparametric estimation of the drift coefficient in the diffusion equation. Statistics (Berlin, DDR), 12(1):61–73, 1981. CODEN MOSSD5. ISSN 0323-3944.

Turkman:1990:BAA

[Tur90] M. A. Amaral Turkman. Bayesian analysis of an autoregressive process with exponential white noise. Statistics: a Journal of Theoretical and Applied Statistics, 21(4):601–608, 1990. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Tuy:1974:CAN

[Tuy74] Hoàangtuy Tuy. On the convex approximation of nonlinear inequalities. Mathematische Operationsforschung und Statistik, 5(6):451–464, 1974. CODEN MTOSBI. ISSN 0047-6277.

Tuy:1976:CAN

[Tuy76] Hoang Tuy. On the convex approximation of nonlinear inequalities. Mathematische Operationsforschung und Statistik, 7(2):319, 1976. CODEN MTOSBI. ISSN 0047-6277.

Truong-Van:1996:ADL

[TVL96] B. Truong-Van and I. Larramendy. Asymptotic distribution of least squares estimators for purely unstable Arma(m, ∞). Statistics: a Journal of Theoretical and Applied Statistics, 28(4):307–346, 1996. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Toutenburg:1978:RSL

[TW78] H. Toutenburg and B. Wargowske. On restricted 2-stage-least-squares (2 SLSE) in a system of structural equations. Statistics (Berlin, DDR), 9(2):167–177, 1978. CODEN MOSSD5. ISSN 0323-3944.

Theodorescu:1983:DSA

[TW83] Radu Theodorescu and Hans Wolff. On dynamic stochastic approximation. Statistics (Berlin, DDR), 14(3):437–448, 1983. CODEN MOSSD5. ISSN 0323-3944.

Tang:2005:EVC

[TW05] Qingguo Tang and Jinde Wang. L_{1}-estimation for varying coefficient models. Statistics: a Journal of Theoretical and Applied Statistics, 39(5):389–404, 2005. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Tang:2006:IDN

[TWZ06] Nian-Sheng Tang, Bo-Cheng Wei, and Wen-Zhuan Zhang. Influence diagnostics in nonlinear reproductive dispersion mixed models. Statistics: a Journal of Theoretical and Applied Statistics, 40(3):227–246, 2006. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Tian:2017:GEL

[TX17] Ruiqin Tian and Liugen Xue. Generalized empirical likelihood inference in partial linear regression model for longitudinal data. Statistics: a Journal of Theoretical and Applied Statistics, 51(5):988–1005, 2017. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Truquet:2012:QLE

[TY12] L. Truquet and J. Yao. On the quasi-likelihood estimation for random coefficient autoregressions. Statistics: a Journal of Theoretical and Applied Statistics, 46(4):505–521, 2012. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Tiwari:1991:NMI

[TZ91] Ram C. Tiwari and Jyoti N. Zalkikar. A new measure of IFRA-ness and its application to two sample test. Statistics: a Journal of Theoretical and Applied Statistics, 22(3):419–430, 1991. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Tang:2013:ELB

[TZ13] Nian-Sheng Tang and Pu-Ying Zhao. Empirical likelihoodbased inference in nonlinear regression models with missing responses at random. Statistics: a Journal of Theoretical and Applied Statistics, 47(6):1141–1159, 2013. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Tziafetas:1987:CBP

[Tzi87] George N. Tziafetas. On the construction of Bayesian prediction limits for the Weibull distribution. Statistics: a Journal of Theoretical and Applied Statistics, 18(4):623–628, 1987. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Uno:2013:HTS

[UI13] Chikara Uno and Eiichi Isogai. Higher than second order approximations for sequential point estimation by a two-stage procedure. Statistics: a Journal of Theoretical and Applied Statistics, 47(2):306–316, 2013. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Ullrich:1976:KOR

[Ull76a] Matthias Ullrich. Konstruktion optimaler Reihenfolgen bei speziellen Reihenfolgeproblemen. (German) [Design of optimal sequences for special sequence problems]. Mathematische Operationsforschung und Statistik, 7(3):361–370, 1976. CO-DEN MTOSBI. ISSN 0047-6277.

Ullrich:1976:ZZR

[Ull76b] Matthias Ullrich. Zusammenhang zwischen dem Rundreiseproblem und einem Reihenfolgeproblem der Maschinenbelegungsplanung. (German) [Connection between the roundtrip problem and a sequence problem of machine assignment planning]. Mathematische Operationsforschung und Statistik, 7(6):843–849, 1976. CODEN MTOSBI. ISSN 0047-6277.

Usha:1981:GAS

[UR81] K. Usha and R. Ramanarayanan. General analysis of systems in which a 2-unit system is a sub-system. Statistics (Berlin, DDR), 12(4):629–637, 1981. CODEN MOSSD5. ISSN 0323-3944.

Vaduva:1977:CGG

[Väd77] I. Väduva. On computer generation of gamma random variables by rejection and composition procedures. Statistics (Berlin, DDR), 8(4):545–576, 1977. CODEN MOSSD5. ISSN 0323-3944.

Valliant:1996:D

[Val96] Richard Valliant. Discussion. Statistics: a Journal of Theoretical and Applied Statistics, 27(3–4):247–251, 1996. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

VanDerLinde:1985:IRF

[Van85] A. Van Der Linde. Interpolation of regression functions in reproducing kernel Hubert spaces. Statistics: a Journal of Theoretical and Applied Statistics, 16(3):351–361, 1985. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

VanderLinde:1988:RFA

[Van88] A. Van der Linde. Rethinking factor analysis as an interpolation problem. Statistics: a Journal of Theoretical and Applied Statistics, 19(3):359–367, 1988. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

VanDerGenugten:1991:IWL

[Van91] B. B. Van Der Genugten. Iterated weighted least squares in heteroscedastic linear models. Statistics: a Journal of Theoretical and Applied Statistics, 22(4):495–516, 1991. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

VanDerLinde:1998:SE

[Van98a] Angelika Van Der Linde. Smoothing errors. Statistics: a Journal of Theoretical and Applied Statistics, 31(2):91–114, 1998. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Vannman:1998:FCI

[Vän98b] Kerstin Vännman. Families of capability indices for onesided specification limits. Statistics: a Journal of Theoretical and Applied Statistics, 31(1):43–66, 1998. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Vineze:1976:RGB

[VBP^{+}76] I. Vineze, O. Bunke, R. Pincus, O. Bunke, E. Läuter, H.- P Höschel, K. Fischer, H. Thiele, H. Bandemer, G. Enderlein, K. Fischer, U. Haverland, K. Fischer, and W. Preuß. Rezensionen. (German) [Book reviews]. Mathematische Operationsforschung und Statistik, 7(5):803–812, 1976. CODEN MTOSBI. ISSN 0047-6277.

Voss:2009:SMF

[VCO09] Florian Voss and Luis M. Cruz-Orive. Second moment formulae for geometric sampling with test probes. Statistics: a Journal of Theoretical and Applied Statistics, 43(4):329–365, 2009. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Volterman:2013:SPC

[VDB13a] William Volterman, Katherine F. Davies, and N. Balakrishnan. Simultaneous Pitman closeness of progressively Type-II right-censored order statistics to population quantiles. Statistics: a Journal of Theoretical and Applied Statistics, 47(2): 439–452, 2013. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Volterman:2013:TSP

[VDB13b] William Volterman, Katherine F. Davies, and N. Balakrishnan. Two-sample Pitman closeness comparison under progressive Type-II censoring. Statistics: a Journal of Theoretical and Applied Statistics, 47(6):1305–1320, 2013. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

vanderVaart:2003:SEM

[vdVvdL03] Aad van der Vaart and Mark van der Laan. Smooth estimation of a monotone density. Statistics: a Journal of Theoretical and Applied Statistics, 37(3):189–203, 2003. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Veraverbeke:1992:ARF

[Ver92] Noël Veraverbeke. Asymptotic results for U-functions of concomitants of order statistics. Statistics: a Journal of Theoretical and Applied Statistics, 23(3):257–264, 1992. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Veraverbeke:2006:RQU

[Ver06] Noël Veraverbeke. Regression quantiles under dependent censoring. Statistics: a Journal of Theoretical and Applied Statistics, 40(2):117–128, 2006. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Vilar-Fernandez:2004:NCC

[VFGM04] J. M. Vilar-Fernández and W. González-Manteiga. Nonparametric comparison of curves with dependent errors. Statistics: a Journal of Theoretical and Applied Statistics, 38(2):81–99, 2004. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Vila:2020:BGD

[VFS^{+}20] Roberto Vila, Letícia Ferreira, Helton Saulo, Fábio Prataviera, and Edwin Ortega. A bimodal gamma distribution: properties, regression model and applications. Statistics: a Journal of Theoretical and Applied Statistics, 54(3):469–493, 2020. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic). See corrigendum [AE21].

Vincze:1981:BR

[VHW^{+}81] I. Vincze, K. Hoffmann, G. Wittwer, G. Linares, W. Warmuth, H. J. Engelbert, P. Lorenz, Jäger, E. Läuter, M. Möhner, and B. M. Kirstein. Book reviews. Statistics (Berlin, DDR), 12(2):287–294, 1981. CODEN MOSSD5. ISSN 0323-3944.

Velez:1990:OIE

[VI90] Ricardo Vélez and Pilar Ibarbola. Optimal invariant estimation of a scale parameter of a continuous time stochastic process. Statistics: a Journal of Theoretical and Applied Statistics, 21(2):197–202, 1990. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Vidal:2023:UEM

[Vid23] Ignacio Vidal. Unforced errors in the matching problem. Statistics: a Journal of Theoretical and Applied Statistics, 57(2):496–514, 2023. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic). URL http://www.

tandfonline.com/doi/full/10.1080/02331888.2023.2201010.

Vieu:1995:OCN

[Vie95] Philippe Vieu. Order choice in nonlinear autoregressive models. Statistics: a Journal of Theoretical and Applied Statistics, 26(4):307–328, 1995. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Vieu:2002:DDM

[Vie02] Philippe Vieu. Data-driven model choice in multivariate nonparametric regression. Statistics: a Journal of Theoretical and Applied Statistics, 36(3):231–246, 2002. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Visek:1986:DTE

[Vís86] Jan Ámos Vísek. On the dependence of the test error probabilities on the contamination level. Statistics: a Journal of Theoretical and Applied Statistics, 17(1):25–32, 1986. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Visek:2011:EDF

[Vís11] Jan Ámos Vísek. Empirical distribution function under heteroscedasticity. Statistics: a Journal of Theoretical and Applied Statistics, 45(5):497–508, 2011. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Volovskiy:2020:OEP

[VK20] Grigoriy Volovskiy and Udo Kamps. Optimal equivariant prediction regions based on multiply type-II censored generalized order statistics from exponential distributions. Statistics: a Journal of Theoretical and Applied Statistics, 54(5):951–968, 2020. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Vandev:1998:ARE

[VN98] D. L. Vandev and N. M. Neykov. About regression estimators with high breakdown point. Statistics: a Journal of Theoretical and Applied Statistics, 32(2):111–129, 1998. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Velilla:2019:NDT

[VN19] Santiago Velilla and Huong Nguyen. A new diagnostic tool for VARMA(p, q) models. Statistics: a Journal of Theoretical and Applied Statistics, 53(4):866–884, 2019. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Viano:1975:RTC

[VO75] M. C. Viano and G. Oppenheim. Robustesse de tests de comparaisons de moyennes vis à vis de la non-indépendance des observations. (French) [Robustness test comparisons of means with respect to the nonindependence of observations]. Mathematische Operationsforschung und Statistik, 6(2):197– 211, 1975. CODEN MTOSBI. ISSN 0047-6277.

VanMontfort:1978:TSP

[VO78] Martin A. J. Van Montfort and Albert Otten. On testing a shape parameter in the presence of a location and a scale parameter. Statistics (Berlin, DDR), 9(1):91–104, 1978. CO-DEN MOSSD5. ISSN 0323-3944.

VonCollani:1987:AOS

[Von87] Elart Von Collani. Approximate A-optimal sampling plans. Statistics: a Journal of Theoretical and Applied Statistics, 18(3):333–344, 1987. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

VonResen:1991:MML

[Von91] Dietrich Von Resen. Moments of maximum likelihood estimators in the growth curve model. Statistics: a Journal of Theoretical and Applied Statistics, 22(1):111–131, 1991. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Viguier-Pla:2004:FBC

[VP04] Sylvie Viguier-Pla. Factor-based comparison of k populations. Statistics: a Journal of Theoretical and Applied Statistics, 38 (1):1–15, 2004. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Vehel:2021:ECE

[VPR21] J. Lévy Véhel, A. Philippe, and C. Robet. Explicit and combined estimators for parameters of stable distributions. Statistics: a Journal of Theoretical and Applied Statistics, 55(4): 736–764, 2021. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

vonRosen:1988:MMN

[vR88] Dietrich von Rosen. Moments for matrix normal variables. Statistics: a Journal of Theoretical and Applied Statistics, 19(4):575–583, 1988. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

vonRosen:1997:DDA

[vR97a] Dietrich von Rosen. Distribution and density approximations of a maximum likelihood estimator in the growth curve model. Statistics: a Journal of Theoretical and Applied Statistics, 29 (1):37–59, 1997. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

vonRosen:1997:MIW

[vR97b] Dietrich von Rosen. On moments of the inverted Wishart distribution. Statistics: a Journal of Theoretical and Applied Statistics, 30(3):259–278, 1997. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

VanHouwelingen:1981:ETB

[VS81] J. C. Van Houwelingen and R. M. Schipper. The efficiency of a test based on the asymptotic distribution of the MLE for a linear functional relationship. Statistics (Berlin, DDR), 12 (1):21–30, 1981. CODEN MOSSD5. ISSN 0323-3944.

VanDerLinde:2008:ARC

[VT08] Angelika Van Der Linde and Gerhard Tutz. On association in regression: the coefficient of determination revisited. Statistics: a Journal of Theoretical and Applied Statistics, 42(1): 1–24, 2008. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

VanDerLaan:1996:ULF

[VV96] Paul Van Der Laan and Constance Van Eeden. On using a loss function in selecting the best of two gamma populations in terms of their scale parameters. Statistics: a Journal of Theoretical and Applied Statistics, 28(4):355–370, 1996. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

VanDerLaan:2000:SSB

[VV00] Mark J. Van Der Laan and Paul Van Der Laan. Subset selection based on order statistics from logistic populations. Statistics: a Journal of Theoretical and Applied Statistics, 34(3): 237–245, 2000. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

vonWolfersdorf:1972:LMAa

[vW72a] Lothar von Wolfersdorf. Zur Lösung einer Minimax-Aufgabe mit Anwendung auf ein Problem der optimalen Steuerung. I. (German) [On solution of a minimax task with application to a problem of optimal control. I]. Mathematische Operationsforschung und Statistik, 3(4):259–271, 1972. CODEN MTOSBI. ISSN 0047-6277.

vonWolfersdorf:1972:LMAb

[vW72b] Lothar von Wolfersdorf. Zur Lösung einer Minimax-Aufgabe mit Anwendung auf ein Problem der optimalen Steuerung II. (German) [On solving a minimax task with application to a problem of optimal control II]. Mathematische Operationsforschung und Statistik, 3(5):343–348, 1972. CODEN MTOSBI. ISSN 0047-6277.

vonWolfersdorf:1974:ZKM

[vW74]  Lothar von Wolfersdorf.  Zurückführung einer Klasse

von Minimum-Problemen der Steuertheorie auf Minimax-Aufgaben. (German) [Previous management of a class of minimum problems of control theory on minimax tasks]. Mathematische Operationsforschung und Statistik, 5(4–5):295–305, 1974. CODEN MTOSBI. ISSN 0047-6277.

VandeVen:1996:PRB

[VW96] R. Van de Ven and N. C. Weber. Percentile related bounds and estimator for the mean of the negative binomial distribution. Statistics: a Journal of Theoretical and Applied Statistics, 27(3–4):345–352, 1996. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Vexler:2009:ECP

[VWL^{+}09] Albert Vexler, Chengqing Wu, Aiyi Liu, Brian W. Whitcomb, and Enrique F. Schisterman. An extension of a change-point problem. Statistics: a Journal of Theoretical and Applied Statistics, 43(3):213–225, 2009. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Vuchkov:1983:COD

[VYD83] I. V. Vuchkov, Ch. A. Yontchev, and D. L. Damgaliev. Continuous D-optimal designs for experiments with mixture and process variables. Statistics (Berlin, DDR), 14(1):33–51, 1983. CODEN MOSSD5. ISSN 0323-3944.

Wu:2022:TPE

[WA22] Jibo Wu and Yasin Asar. A two-parameter estimator in linear measurement error model. Statistics: a Journal of Theoretical and Applied Statistics, 56(4):739–754, 2022. CO-DEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/02331888.2022.2098959.

Wachtler:1986:COP

[Wäc86] M. Wächtler. Cost-optimal performance of special reliability tests using prior knowledge. Statistics: a Journal of Theoretical and Applied Statistics, 17(1):87–103, 1986. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Wafula:1988:EVR

[Waf88] Charles Wafula. On estimating the variance of the ratio estimator. Statistics: a Journal of Theoretical and Applied Statistics, 19(4):565–574, 1988. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Wagner:1983:ALL

[Wag83] Wolfgang Wagner. Zur Abschätzung von Lösungen linearer Integralgleichungen mit Hilfe der Monte-Carlo-Methode. (German) [On the estimation of solutions of linear integral equations by means of the Monte Carlo method]. Statistics (Berlin, DDR), 14(3):449–456, 1983. CODEN MOSSD5. ISSN 0323-3944.

Walter:1988:MFB

[Wal88] Warmuth Walter. Marginal-Fréchet-bounds for multidimensional distribution functions. Statistics: a Journal of Theoretical and Applied Statistics, 19(2):283–294, 1988. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Wang:2007:GCL

[Wan07] Lihong Wang. Gradual changes in long memory processes with applications. Statistics: a Journal of Theoretical and Applied Statistics, 41(3):221–240, 2007. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Wang:2014:CIB

[Wan14] Hsiuying Wang. Confidence intervals for a bounded mean in exponential families. Statistics: a Journal of Theoretical and Applied Statistics, 48(2):327–343, 2014. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Wang:2015:MEE

[Wan15] Shaochen Wang. MDP for estimators in EV regression models with α-mixing errors. Statistics: a Journal of Theoretical and Applied Statistics, 49(1):119–127, 2015. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Warmuth:1985:BR

[War85a] Walter Warmuth. Book review. Statistics: a Journal of Theoretical and Applied Statistics, 16(3):474–475, 1985. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Warmuth:1985:BRK

[War85b] Walter Warmuth. Book review: K. Itô and J. P. Prokhorov (eds.), Probability Theory and Mathematical Statistics: Fourth USSR–Japan Symposium, Proceedings, 1982. Statistics: a Journal of Theoretical and Applied Statistics, 16(2):

184, 1985. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Warmuth:1986:BRb

[War86a] W. Warmuth. Book review. Statistics: a Journal of Theoretical and Applied Statistics, 17(1):62, 1986. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Warmuth:1986:BRc

[War86b] W. Warmuth. Book review. Statistics: a Journal of Theoretical and Applied Statistics, 17(1):104, 1986. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Warmuth:1986:BRa

[War86c] W. Warmuth. Book reviews. Statistics: a Journal of Theoretical and Applied Statistics, 17(1):40, 1986. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Wawrzyniak:1981:CMV

[Waw81] W. Wawrzyniak. A characterization of minimum variance unbiased estimators in the general linear model with restrictions on parameter space. Statistics (Berlin, DDR), 12(4):465–477, 1981. CODEN MOSSD5. ISSN 0323-3944.

Wulff:2005:MVU

[WB05] S. S. Wulff and D. Birkes. Minimum variance unbiased invariant estimation of variance components under normality. Statistics: a Journal of Theoretical and Applied Statistics, 39 (1):53–65, 2005. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Winkler:1981:BR

[WBG^{+}81] W. Winkler, H. Bandemer, J. Gladitz, H. Läuter, J. Gladitz, W. Schmidt, B. Gerlach, E. Platen, K. Arndt, M. Möhner, D. Stoyan, and R. Klette. Book reviews. Statistics (Berlin, DDR), 12(1):135–143, 1981. CODEN MOSSD5. ISSN 0323-3944.

Wiens:2006:EDS

[WBL06] Douglas P. Wiens, Norman C. Beaulieu, and Pavel Loskot. On the exact distribution of the sum of the largest n − k out of n normal random variables with differing mean values. Statistics: a Journal of Theoretical and Applied Statistics, 40(2):165–173, 2006. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Wang:2021:EVE

[WC21] Qing Wang and Xizhen Cai. An efficient variance estimator for cross-validation under partition sampling. Statistics: a Journal of Theoretical and Applied Statistics, 55(3):660–681, 2021. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Weissbach:2022:RSS

[WD22] Rafael Weißbach and Achim Dörre. Retrospective sampling of survival data based on a Poisson birth process: conditional maximum likelihood. Statistics: a Journal of Theoretical and Applied Statistics, 56(4):844–866, 2022. CO-DEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/02331888.2022.2105843.

Wang:2014:CCA

[WDZH14] Xuejun Wang, Xin Deng, Lulu Zheng, and Shuhe Hu. Complete convergence for arrays of rowwise negatively superadditive-dependent random variables and its applications. Statistics: a Journal of Theoretical and Applied Statistics, 48(4):834–850, 2014. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Weba:1979:ISS

[Web79] Michael Weba. Interpolation schwach stationärer Zeitreihen. (German) [Interpolation of weakly stationary time series]. Statistics (Berlin, DDR), 10(3):415–425, 1979. CODEN MOSSD5. ISSN 0323-3944.

Weba:1989:ORM

[Web89] Michael Weba. Optimized regression models for time series. Statistics: a Journal of Theoretical and Applied Statistics, 20(1):109–123, 1989. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Weba:2006:ERC

[Web06] Michael Weba. Estimation of regression coefficients in case of differentiable error processes. Statistics: a Journal of Theoretical and Applied Statistics, 40(2):95–116, 2006. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Weinert:1970:DEL

[Wei70a] Horst Weinert. Doppelt-einparametrische lineare Optimierung. I: Unabhängige Parameter. (German) [Doublesingle-parameter linear optimization. I: Independent parameters]. Mathematische Operationsforschung und Statistik, 1(3): 173–197, 1970. CODEN MTOSBI. ISSN 0047-6277.

Weinert:1970:PLO

[Wei70b] Horst Weinert. Probleme der linearen Optimierung mit nichtlinear-einparametrischen Koeffizienten in der Zielfunktion. (German) [Linear optimization problem with nonlinear one-parameter coefficients in the objective function]. Mathematische Operationsforschung und Statistik, 1(1):21–43, 1970. CODEN MTOSBI. ISSN 0047-6277.

Weinert:1971:DEL

[Wei71] Horst Weinert. Doppelt-einparametrische lineare Optimierung. II: Abhängige Parameter. (German) [Double oneparameter linear optimization. II: Dependent parameters]. Mathematische Operationsforschung und Statistik, 2(1):19– 39, 1971. CODEN MTOSBI. ISSN 0047-6277.

Weinert:1972:PLP

[Wei72] Horst Weinert. On parametric linear programming problems with fixed matrix of constraints. Mathematische Operationsforschung und Statistik, 3(6):407–417, 1972. CODEN MTOSBI. ISSN 0047-6277.

Weinert:1974:UPL

[Wei74] Host Weinert. On uniqueness in parametric linear programming problems with fixed matrix of constraints. Mathematische Operationsforschung und Statistik, 5(3):177–189, 1974. CODEN MTOSBI. ISSN 0047-6277.

Weiss:2009:PCB

[Wei09] Christian H. Weiß. Properties of a class of binary ARMA models. Statistics: a Journal of Theoretical and Applied Statistics, 43(2):131–138, 2009. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Wendler:1976:TPK

[Wen76a] Klaus Wendler. Das Transportproblem mit Parametern als Kapazitätsbeschränkungen. (German) [The transport problem with parameters as capacity restrictions]. Mathematische Operationsforschung und Statistik, 7(3):381–394, 1976. CO-DEN MTOSBI. ISSN 0047-6277.

Wendler:1976:ETP

[Wen76b] Klaus Wendler. Ein erweitertes Transportproblem mit Parametern als Kapazitätsbeschränkungen. (German) [An extended transportation problem with parameters as capacity constraints]. Mathematische Operationsforschung und Statistik, 7(6):859–870, 1976. CODEN MTOSBI. ISSN 0047-6277.

Wenli:1995:UCM

[Wen95] Yang Wenli. MINQE(U, I) and UMVIQUE of the covariance matrix in the growth curve model. Statistics: a Journal of Theoretical and Applied Statistics, 26(1):49–59, 1995. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Wernstedt:1978:EBD

[Wer78] J. Wernstedt. Zur experimentellen Bildung dynamischer Modelle für die automatische Steuerung von Systemen — eine Übersicht. (German) [On the experimental formation of dynamic models for automatic control of systems — an outline]. Statistics (Berlin, DDR), 9(2):293–325, 1978. CODEN MOSSD5. ISSN 0323-3944.

Wertz:1985:SRE

[Wer85] Wolfgang Wertz. Sequential and recursive estimators of the probability density. Statistics: a Journal of Theoretical and Applied Statistics, 16(2):277–295, 1985. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Wesolowski:1993:SRM

[Wes93] Jacek Wesolowski. Some remarks on the multivariate Liouville distribution. Statistics: a Journal of Theoretical and Applied Statistics, 24(2):167–170, 1993. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Westerlund:2014:ADD

[Wes14] Joakim Westerlund. On the asymptotic distribution of the Dickey Fuller–GLS test statistic. Statistics: a Journal of Theoretical and Applied Statistics, 48(6):1233–1253, 2014. CO-DEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Winkler:1982:SSP

[WFK82] W. Winkler, J. Franz, and I. Küchler. Sequential statistical procedures for processes of the exponential class with independent increments. Statistics (Berlin, DDR), 13(1):105–119, 1982. CODEN MOSSD5. ISSN 0323-3944.

Wolfgang:1992:NEE

[WG92] H. Schmidt Wolfgang and Telton Günter. A note on the Edgeworth-expansion for MINQUE. Statistics: a Journal of Theoretical and Applied Statistics, 23(2):121–128, 1992. CO-DEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Wang:2021:SIG

[WG21] Kexin Wang and Wenhao Gui. Statistical inference of generalized progressive hybrid censored step-stress accelerated dependent competing risks model for Marshall–Olkin bivariate Weibull distribution. Statistics: a Journal of Theoretical and Applied Statistics, 55(5):1058–1093, 2021. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Wang:2020:EEN

[WH20] Xiaoguang Wang and Bo Han. Efficient estimation for the non-mixture cure model with current status data. Statistics: a Journal of Theoretical and Applied Statistics, 54(4):756–777, 2020. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Winkler:1990:BRb

[WHD^{+}90] W. Winkler, R. Hanrel, B. Droge, K. Fischer, Walter Warmuth, M. Riedel, B. Tietze, S. Zwanzig, and S. Zwanzig. Book reviews. Statistics: a Journal of Theoretical and Applied Statistics, 21(4):643–648, 1990. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Wenli:1987:BLU

[WHG87] Yang Wenli, Cui Hengjian, and Sun Guowen. On best linear unbiased estimation in the restricted general linear model. Statistics: a Journal of Theoretical and Applied Statistics, 18 (1):17–20, 1987. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Winkler:1980:BR

[WHH^{+}80] W. Winkler, K. Henschke, K. Henschke, O. Bunke, B. Bank, P. Neumann, H. J. Engelbert, D. Stoyan, I. Sonnenburg, H. Bandemer, F. Belchelt, J. Andêl, B. Kummer, K. Hoffmann, and W. Winkler. Book review. Statistics (Berlin, DDR), 11(4):589–600, 1980. CODEN MOSSD5. ISSN 0323-3944.

Weinert:1976:RGB

[WHS^{+}76] H. Weinert, H. Hollatz, N. Sieber, A. Güpfert, W. Thümelt, H. Hollatz, H.-U. Küenle, W. Dück, H.-J. Rossberg, J. Mecke, D. Stoyan, P. Rudolph, K.-H. Bachmann, G. Polze, J. Born, J. Nietzsch, R. Reinhardt, and M. Schäuble. Rezensionen. (German) [Book reviews]. Mathematische Operationsforschung und Statistik, 7(6):957–975, 1976. CODEN MTOSBI. ISSN 0047-6277.

Wu:2003:EPG

[WHT03] Jong-Wuu Wu, Wen-Liang Hung, and Chih-Hui Tsai. Estimation of the parameters of the Gompertz distribution under the first failure-censored sampling plan. Statistics: a Journal of Theoretical and Applied Statistics, 37(6):517–525, 2003. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Wu:2023:SIR

[WHVW23] Yi Wu, Tien-Chung Hu, Andrei Volodin, and Xuejun Wang. Some improved results on Berry–Esséen bounds for strong mixing random variables and applications. Statistics: a Journal of Theoretical and Applied Statistics, 57(3):740–760, 2023. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/02331888.2023.2213460.

Wang:2016:GME

[WHXC16] Siyang Wang, Tao Hu, Liming Xiang, and Hengjian Cui. Generalized m-estimation for the accelerated failure time model. Statistics: a Journal of Theoretical and Applied Statistics, 50(1):114–138, 2016. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic). URL http://

www.tandfonline.com/doi/abs/10.1080/02331888.2015.
1032970.

Wierich:1986:OPD

[Wie86] W. Wierich. Optimality of k-proportional designs for simultaneous inference. Statistics: a Journal of Theoretical and Applied Statistics, 17(2):179–187, 1986. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Wierich:1988:ONO

[Wie88] W. Wierich. On A-optimality and non-optimality of the best proportional design measures in anova-settings. Statistics: a Journal of Theoretical and Applied Statistics, 19(4):503–512, 1988. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Wiens:1992:MQF

[Wie92] Douglas P. Wiens. On moments of quadratic forms in nonspherically distributed variables. Statistics: a Journal of Theoretical and Applied Statistics, 23(3):265–270, 1992. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Wilhelmi:1987:BR

[Wil87] W. Wilhelmi. Book review. Statistics: a Journal of Theoretical and Applied Statistics, 18(1):48, 1987. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Wilczynski:2008:MEE

[Wil08] Maciej Wilczyński. Minimax estimation over ellipsoids in l_{2}. Statistics: a Journal of Theoretical and Applied Statistics, 42 (2):95–100, 2008. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Winter:1981:FSM

[Win81] B. B. Winter. Fourier series methods in nonparametric estimation. Statistics (Berlin, DDR), 12(3):339–355, 1981. CO-DEN MOSSD5. ISSN 0323-3944.

Winkler:1987:BR

[Win87] W. Winkler. Book review. Statistics: a Journal of Theoretical and Applied Statistics, 18(3):372, 1987. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Winkler:1989:BR

[Win89] W. Winkler. Book reviews. Statistics: a Journal of Theoretical and Applied Statistics, 20(4):582, 1989. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Winkler:1990:BRa

[Win90] W. Winkler. Book reviews. Statistics: a Journal of Theoretical and Applied Statistics, 21(1):34, 1990. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Wintenberger:2017:EIU

[Win17] Olivier Wintenberger. Exponential inequalities for unbounded functions of geometrically ergodic Markov chains: applications to quantitative error bounds for regenerative Metropolis algorithms. Statistics: a Journal of Theoretical and Applied Statistics, 51(1):222–234, 2017. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Wisotzki:1977:PAN

[Wis77] C. Wisotzki. Polynomial approximation of nonlinear regression functions. Statistics (Berlin, DDR), 8(3):313–321, 1977. CODEN MOSSD5. ISSN 0323-3944.

Wittwer:1976:ATK

[Wit76a] Gisela Wittwer. Anwendung der Theorie kernreproduzierender Hilbert-Räume in der Statistik zufälliger Vektorprozesse. (German) [Application of the theory of kernel-reproducing Hilbert spaces in the statistics of random vector processes]. Mathematische Operationsforschung und Statistik, 7(5):743– 754, 1976. CODEN MTOSBI. ISSN 0047-6277.

Wittwer:1976:VIS

[Wit76b] Gisela Wittwer. Versuchsplanung im Sinne von Sacks– Ylvisaker for Vektorprozesse. (German) [Experimental design in the sense of Sacks–Ylvisaker for vector processes]. Mathematische Operationsforschung und Statistik, 7(1):95– 105, 1976. CODEN MTOSBI. ISSN 0047-6277.

Wittwer:1978:AOV

[Wit78a] Gisela Wittwer. Über asymptotisch optimale Versuchsplanung im Sinne von Sacks–Ylvisaker. (German) [On asymptotically optimal experimental planning in the sense of Sacks– Ylvisaker]. Statistics (Berlin, DDR), 9(1):61–71, 1978. CO-DEN MOSSD5. ISSN 0323-3944.

Wittwer:1978:AVP

[Wit78b] Gisela Wittwer. Über die asymptotische Verteilung des periodogramms stationärer gaussscher zufälliger Folgen. (German) [On the asymptotic distribution of the periodogram of stationary Gaussian random sequences]. Statistics (Berlin, DDR), 9(3):357–368, 1978. CODEN MOSSD5. ISSN 0323-3944.

Wittwer:1981:MAV

[Wit81] Gisela Wittwer. Über die mehrdimensionalen asymptotischen Verteilungen des Periodogramms Gaussscher stationärer Folgen. (German) [On the multi-dimensional asymptotic distributions of the periodogram of stationary Gaussian sequences]. Statistics (Berlin, DDR), 12(3):361–376, 1981. CO-DEN MOSSD5. ISSN 0323-3944.

Wittwer:1983:BR

[Wit83a] G. Wittwer. Book reviews. Statistics (Berlin, DDR), 14(1): 140, 1983. CODEN MOSSD5. ISSN 0323-3944.

Wittwer:1983:RCD

[Wit83b] Gisela Wittwer. On the rate of convergence of the distribution of the periodogram matrix for Gaussian vector processes. Statistics (Berlin, DDR), 14(1):55–62, 1983. CODEN MOSSD5. ISSN 0323-3944.

Wittwer:1986:DPS

[Wit86] Gisela Wittwer. On the distribution of the periodogram for stationary random sequences. Statistics: a Journal of Theoretical and Applied Statistics, 17(2):201–219, 1986. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Wittwer:1989:BR

[Wit89a] G. Wittwer. Book reviews. Statistics: a Journal of Theoretical and Applied Statistics, 20(4):546, 1989. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Wittwer:1989:SRB

[Wit89b] G. Wittwer. Some remarks on bilinear time series models. Statistics: a Journal of Theoretical and Applied Statistics, 20(4):521–529, 1989. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Wittwer:1990:BR

[Wit90] G. Wittwer. Book reviews. Statistics: a Journal of Theoretical and Applied Statistics, 21(4):612, 1990. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Wittwer:1991:BR

[Wit91] G. Wittwer. Book reviews. Statistics: a Journal of Theoretical and Applied Statistics, 22(1):142, 1991. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Wittwer:1993:CEH

[Wit93] G. Wittwer. Consistent estimates for hidden frequencies in stationary processes for irregularly observed data. Statistics: a Journal of Theoretical and Applied Statistics, 24(2):143– 148, 1993. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Winkleb:1985:NSM

[WK85] W. Winkleb and Y. Kowtsch. A note on sequential maximum likelihood estimates in processes with independent increments. Statistics: a Journal of Theoretical and Applied Statistics, 16(3):419–426, 1985. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Weiss:2013:BAP

[WK13] Christian H. Weiß and Hee-Young Kim. Binomial AR(1) processes: moments, cumulants, and estimation. Statistics: a Journal of Theoretical and Applied Statistics, 47(3):494–510, 2013. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Wu:2015:PHD

[WK15] Jingjing Wu and Rohana J. Karunamuni. Profile Hellinger distance estimation. Statistics: a Journal of Theoretical and Applied Statistics, 49(4):711–740, 2015. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Wu:2001:QSS

[WL01] Jong-Wuu Wu and Wen-Chuan Lee. The quasi-score statistic in quasi-likelihood model. Statistics: a Journal of Theoretical and Applied Statistics, 35(4):523–535, 2001. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Wang:2005:DMU

[WL05] Jiantian Wang and Yingfu Li. Dependency measures under bivariate homogeneous shock models. Statistics: a Journal of Theoretical and Applied Statistics, 39(1):73–80, 2005. CO-DEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Wesolowski:2004:LRP

[WLB04] Jacek Wesolowski and Fernando López-Blázquez. Linearity of regression for the past weak and ordinary records. Statistics: a Journal of Theoretical and Applied Statistics, 38(6):457–464, 2004. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Wan:2018:ELT

[WLD18] Yi Wan, Zhi Liu, and Min Deng. Empirical likelihood test for equality of two distributions using distance of characteristic functions. Statistics: a Journal of Theoretical and Applied Statistics, 52(6):1379–1394, 2018. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Wang:2013:LPQ

[WLF13] Jiang-Feng Wang, Han-Ying Liang, and Guo-Liang Fan. Local polynomial quasi-likelihood regression with truncated and dependent data. Statistics: a Journal of Theoretical and Applied Statistics, 47(4):744–761, 2013. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Wang:2022:IDL

[WLT^{+}22] Liang Wang, Yuhlong Lio, Yogesh Mani Tripathi, Sanku Dey, and Fode Zhang. Inference of dependent left-truncated and right-censored competing risks data from a general bivariate class of inverse exponentiated distributions. Statistics: a Journal of Theoretical and Applied Statistics, 56(2):347–374, 2022. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Wang:2016:NSR

[WLW16] Hongxia Wang, Jinguan Lin, and Jinde Wang. Nonparametric spatial regression with spatial autoregressive error structure.

Statistics: a Journal of Theoretical and Applied Statistics, 50 (1):60–75, 2016. CODEN MOSSD5. ISSN 0233-1888 (print),

1029-4910 (electronic). URL http://www.tandfonline.com/
doi/abs/10.1080/02331888.2015.1094068.

Withers:2012:ACF

[WN12a] Christopher S. Withers and Saralees Nadarajah. Adjusting Cornish–Fisher expansions and confidence intervals for the effect of roundoff. Statistics: a Journal of Theoretical and Applied Statistics, 46(5):627–644, 2012. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Withers:2012:RIU

[WN12b] Christopher S. Withers and Saralees Nadarajah. Repeated integrals of the univariate normal as a finite series with the remainder in terms of Moran’s functions. Statistics: a Journal of Theoretical and Applied Statistics, 46(1):13–22, 2012. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Withers:2013:EDP

[WN13a] Christopher S. Withers and Saralees Nadarajah. Expressions for the distribution and percentiles of the sums and products of chi-squares. Statistics: a Journal of Theoretical and Applied Statistics, 47(6):1343–1362, 2013. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Withers:2013:RFH

[WN13b] Christopher S. Withers and Saralees Nadarajah. A reduction formula for the hypergeometric function using gamma random variables. Statistics: a Journal of Theoretical and Applied Statistics, 47(2):399–404, 2013. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Withers:2013:TIM

[WN13c] Christopher S. Withers and Saralees Nadarajah. Testing if a mixture is from a given location-scale family. Statistics: a Journal of Theoretical and Applied Statistics, 47(5):901–916, 2013. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Withers:2014:AIS

[WN14a] Christopher S. Withers and Saralees Nadarajah. Accurate inference for scale and location families. Statistics: a Journal of Theoretical and Applied Statistics, 48(5):1092–1105, 2014. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Withers:2014:CIO

[WN14b] Christopher S. Withers and Saralees Nadarajah. A class of integral operators generated by random variables. Statistics: a Journal of Theoretical and Applied Statistics, 48(1):159– 166, 2014. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Watson:2017:CVN

[WNBH17] J. Watson, L. Nieto-Barajas, and C. Holmes. Characterizing variation of nonparametric random probability measures using the Kullback–Leibler divergence. Statistics: a Journal of Theoretical and Applied Statistics, 51(3):558–571, 2017. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Wolfersdorf:1975:OSH

[Wol75] L. V. Wolfersdorf. Optimale Steuerung bei Hammersteinschen Integralgleichungen mit schwach singulären Kernen. (German) [Optimal control for Hammerstein integral equations with weakly singular kernels]. Mathematische Operationsforschung und Statistik, 6(4):609–626, 1975. CODEN MTOSBI. ISSN 0047-6277.

Wolf:1984:ZGW

[Wol84] W. Wolf. Der zentrale Grenzwertsatz und Wahrscheinlichkeiten großer Abweichungen für Summen. (German) [The Central Limit Theorem and probabilities of large deviation for sums]. Statistics (Berlin, DDR), 15(3):389–395, 1984. CO-DEN MOSSD5. ISSN 0323-3944.

Wolf:1985:LGG

[Wol85] Werner Wolf. Lokale Grenzwertsätze für gewichtete Summen. (German) [Local limit theorems for weighted sums]. Statistics: a Journal of Theoretical and Applied Statistics, 16(2):243– 247, 1985. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Wong:1995:D

[Won95] Weng Kee Wong. Discussion. Statistics: a Journal of Theoretical and Applied Statistics, 26(2):174, 1995. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Winkler:1979:BR

[WPH^{+}79] W. Winkler, R. Pincus, K. Hoffmann, J. Bellach, R. Strüby, J. Gladitz, W. H. Schmidt, H. J. Girlich, O. Bunke, and R. Siegmund-Schultze. Book reviews. Statistics (Berlin, DDR), 10(1):179–184, 1979. CODEN MOSSD5. ISSN 0323-3944.

Wang:2022:UPM

[WQO22] Kang Wang, Hong Qin, and Zujun Ou. Uniformity pattern of mixed two- and three-level factorials under average projection mixture discrepancy. Statistics: a Journal of Theoretical and Applied Statistics, 56(1):121–133, 2022. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Wang:2023:UPU

[WQO23] Kang Wang, Hong Qin, and Zujun Ou. Uniformity and projection uniformity of combined designs. Statistics: a Journal of Theoretical and Applied Statistics, 57(3):515–533, 2023. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/02331888.2023.2206661.

Wilke:1972:RGB

[WRME72] J. Wilke, K. Reinisch, J. Metz, and G. Enderlein. Rezensionen. (German) [Book reviews]. Mathematische Operationsforschung und Statistik, 3(5):401–403, 1972. CODEN MTOSBI. ISSN 0047-6277.

Walder:1997:MMT

[WS97] O. Wälder and D. Stoyan. Models of markings and thinnings of Poisson processes. Statistics: a Journal of Theoretical and Applied Statistics, 29(2):179–202, 1997. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Wang:2013:BFC

[WS13] Min Wang and Xiaoqian Sun. Bayes factor consistency for unbalanced ANOVA models. Statistics: a Journal of Theoretical and Applied Statistics, 47(5):1104–1115, 2013. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Wang:2022:CPN

[WS22] Rui Wang and Aiting Shen. Consistency properties for the nearest neighbour estimator of the density function and applications based on α-mixing samples. Statistics: a Journal of Theoretical and Applied Statistics, 56(5):991–1011, 2022. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/02331888.2022.2108424.

Wang:2020:SPA

[WT20] Yan Wang and Rui Tuo. Semi-parametric adjustment to computer models. Statistics: a Journal of Theoretical and Applied Statistics, 54(6):1255–1275, 2020. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Wu:2016:JDB

[Wu16] Jibo Wu. A jackknifed difference-based ridge estimator in the partial linear model with correlated errors. Statistics: a Journal of Theoretical and Applied Statistics, 50(6):1363– 1375, 2016. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Warmuth:1983:BR

[WW83] E. Warmuth and W. Warmuth. Book reviews. Statistics (Berlin, DDR), 14(2):210, 1983. CODEN MOSSD5. ISSN 0323-3944.

Warmuth:1991:BR

[WW91] Walter Warmuth and W. Warmut. Book reviews. Statistics: a Journal of Theoretical and Applied Statistics, 22(1):110, 1991. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Worms:2021:EEH

[WW21] Julien Worms and Rym Worms. Estimation of extremes for heavy-tailed and light-tailed distributions in the presence of random censoring. Statistics: a Journal of Theoretical and Applied Statistics, 55(5):979–1017, 2021. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Wang:2022:PEL

[WW22a] Xinyang Wang and Dehui Wang. Penalized empirical likelihood inference for the GINAR(p) model. Statistics: a Journal of Theoretical and Applied Statistics, 56(4):785–804, 2022. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/02331888.2022.2107645.

Wu:2022:UCC

[WW22b] Yi Wu and Xuejun Wang. Uniformly complete consistency of frequency polygon estimation for dependent samples and an application. Statistics: a Journal of Theoretical and Applied Statistics, 56(6):1270–1289, 2022. CO-DEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/02331888.2022.2152452.

Wu:2010:IET

[WWC^{+}10] Shu-Fei Wu, Chin-Chuan Wu, Yung-Lin Chen, Yuh-Ru Yu, and Ying Po Lin. Interval estimation of a two-parameter Burr-XII distribution under progressive censoring. Statistics: a Journal of Theoretical and Applied Statistics, 44(1):77–88, 2010. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Wang:2018:CMC

[WWH18] Xuejun Wang, Yi Wu, and Shuhe Hu. Complete moment convergence for double-indexed randomly weighted sums and its applications. Statistics: a Journal of Theoretical and Applied Statistics, 52(3):503–518, 2018. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Wu:2019:BET

[WWLH19] Yi Wu, Xuejun Wang, Yongming Li, and Shuhe Hu. Berry– Esseen type bounds of the estimators in a semiparametric model under linear process errors with α-mixing dependent innovations. Statistics: a Journal of Theoretical and Applied Statistics, 53(5):943–967, 2019. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Wolff:1990:BR

[WWW^{+}90] Ch. Wolff, W. Winkler, W. Winkler, O. M. Kabahob, V. Schmidt, F. Schmidt, Klaus Fleischmann, Walter Warmuth, A. Kirchheim, Heinz-Uwe Küenle, D. Rasch, W. Klzow, V. Schmidt, W. Winkler, Walter Warmuth, Jürgen Gärtner, Wolfgang Wagner, Heinz-Uwe Küenle, R. Manthey, Walter Warmuth, W. Schmidt, W. Schmidt, W. Schmidt, W. Schmidt, O. Bunke, O. Bunke, O. Bunke, Walter Warmuth, and J. Polzehl. Book reviews. Statistics: a Journal of Theoretical and Applied Statistics, 21(2):300–320, 1990. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Wang:2022:ODN

[WWZ22] Tao Wang, Yunlong Wang, and Qingpei Zang. Outlier detection in non-parametric profile monitoring. Statistics: a Journal of Theoretical and Applied Statistics, 56(4):805–822, 2022. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/02331888.2022.2085707.

Wang:2004:SAP

[WY04] Dabuxilatu Wang and Masami Yasuda. Some asymptotic properties of point estimation with n-dimensional fuzzy data. Statistics: a Journal of Theoretical and Applied Statistics, 38(2):167–181, 2004. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Wu:2013:EAU

[WY13] Jibo Wu and Hu Yang. Efficiency of an almost unbiased twoparameter estimator in linear regression model. Statistics: a Journal of Theoretical and Applied Statistics, 47(3):535–545, 2013. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Wang:2022:BET

[WYLW22] Yan Wang, Wei Yu, Xiaoqin Li, and Xuejun Wang. The Berry–Esseen-type bound for the G-M estimator in a nonparametric regression model with α-mixing errors. Statistics: a Journal of Theoretical and Applied Statistics, 56(1):97–120, 2022. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Wynn:1976:LVF

[Wyn76] H. P. Wynn. The limiting variance function for design sequences. Mathematische Operationsforschung und Statistik, 7 (1):69–74, 1976. CODEN MTOSBI. ISSN 0047-6277.

Wu:2021:BRS

[WYW21] Yi Wu, Wei Yu, and Xuejun Wang. The Bahadur representation of sample quantiles for ϕ-mixing random variables and its application. Statistics: a Journal of Theoretical and Applied Statistics, 55(2):426–444, 2021. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Wei:2021:ESN

[WZW21] Zheng Wei, Xiaonan Zhu, and Tonghui Wang. The extended skew-normal-based stochastic frontier model with a solution to ‘wrong skewness’ problem. Statistics: a Journal of Theoretical and Applied Statistics, 55(6):1387–1406, 2021. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Wang:2015:CCE

[WZXH15] Xuejun Wang, Lulu Zheng, Chen Xu, and Shuhe Hu. Complete consistency for the estimator of nonparametric regression models based on extended negatively dependent errors. Statistics: a Journal of Theoretical and Applied Statistics, 49(2):396–407, 2015. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Xu:2017:SSI

[XB17] Qunfang Xu and Yang Bai. Semiparametric statistical inferences for longitudinal data with nonparametric covariance modelling. Statistics: a Journal of Theoretical and Applied Statistics, 51(6):1280–1303, 2017. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Xu:2019:QRL

[XCJH19] Q. F. Xu, C. Cai, C. X. Jiang, and X. Huang. Quantile regression for large-scale data via sparse exponential transform method. Statistics: a Journal of Theoretical and Applied Statistics, 53(1):26–42, 2019. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Xekalaki:1986:BYD

[Xek86] Evdokia Xekalaki. The bivariate Yule distribution and some of its properties. Statistics: a Journal of Theoretical and Applied Statistics, 17(2):311–317, 1986. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Xu:2007:AED

[XG07] Jin Xu and Arjun K. Gupta. Asymptotic expansions of the distributions of some test statistics in generalized linear models. Statistics: a Journal of Theoretical and Applied Statistics, 41(1):47–64, 2007. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Xie:2009:CML

[Xie09] Yingfu Xie. Consistency of maximum likelihood estimators for the regime-switching GARCH model. Statistics: a Journal of Theoretical and Applied Statistics, 43(2):153–165, 2009. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Xie:2010:IAC

[XW10] Feng-Chang Xie and Bo-Cheng Wei. Influence analysis for count data based on generalized Poisson regression models. Statistics: a Journal of Theoretical and Applied Statistics, 44(4):341–360, 2010. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Xia:2014:APM

[XWJ14] Tian Xia, Xue-Ren Wang, and Xue-Jun Jiang. Asymptotic properties of maximum quasi-likelihood estimator in quasilikelihood nonlinear models with misspecified variance function. Statistics: a Journal of Theoretical and Applied Statistics, 48(4):778–786, 2014. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Xiao:2023:SWU

[XWQN23] Yao Xiao, Shiqi Wang, Hong Qin, and Jianhui Ning. Sequentially weighted uniform designs. Statistics: a Journal of Theoretical and Applied Statistics, 57(3):534–553, 2023. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/02331888.2023.2204438.

Xi:2020:APE

[XWY^{+}20] Mengmei Xi, Rui Wang, Wei Yu, Yan Shen, and Xuejun Wang. Asymptotic properties for the estimators in heteroscedastic semiparametric EV models with α-mixing errors. Statistics: a Journal of Theoretical and Applied Statistics, 54 (6):1232–1254, 2020. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Xia:2016:VSP

[XY16] Xiaochao Xia and Hu Yang. Variable selection for partially time-varying coefficient error-in-variables models. Statistics: a Journal of Theoretical and Applied Statistics, 50(2):278– 297, 2016. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-

4910 (electronic). URL http://www.tandfonline.com/doi/
abs/10.1080/02331888.2015.1074233.

Xiao:2011:MLE

[XZZ11] Wei-Lin Xiao, Wei-Guo Zhang, and Xi-Li Zhang. Maximumlikelihood estimators in the mixed fractional Brownian motion. Statistics: a Journal of Theoretical and Applied Statistics, 45(1):73–85, 2011. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Yanev:2012:CSD

[YA12] George P. Yanev and M. Ahsanullah. Characterizations of Student’s t-distribution via regressions of order statistics. Statistics: a Journal of Theoretical and Applied Statistics, 46(4):429–435, 2012. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Yuzbasi:2022:LTS

[YAA22] Bahadir Yüzbasi, Yasin Asar, and S. Ejaz Ahmed. Liu-type shrinkage estimations in linear models. Statistics: a Journal of Theoretical and Applied Statistics, 56(2):396–420, 2022. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic). See correction [Nku23].

Yao:2000:REI

[Yao00] Jian-Feng Yao. On recursive estimation in incomplete data models. Statistics: a Journal of Theoretical and Applied Statistics, 34(1):27–51, 2000. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Yatracos:2017:DDS

[Yat17] Yannis G. Yatracos. Distributional divergence, statistical experiments and consequences in option pricing. Statistics: a Journal of Theoretical and Applied Statistics, 52(1):18–33, 2017. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Yatracos:2019:PUE

[Yat19] Yannis G. Yatracos. Plug-in L_{2}-upper error bounds in deconvolution, for a mixing density estimate in R^{d} and for its derivatives, via the L_{1}-error for the mixture. Statistics: a Journal of Theoretical and Applied Statistics, 53(6):1251– 1268, 2019. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Yan:2015:ELG

[YC15] Li Yan and Xia Chen. Empirical likelihood for generalized linear models with fixed and adaptive designs. Statistics: a Journal of Theoretical and Applied Statistics, 49(5):978–988, 2015. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Yan:2013:FRP

[YDZ13] Rongfang Yan, Gaofeng Da, and Peng Zhao. Further results for parallel systems with two heterogeneous exponential components. Statistics: a Journal of Theoretical and Applied Statistics, 47(5):1128–1140, 2013. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Yang:2011:AND

[YFZ11] Xiaorong Yang, Ke-Ang Fu, and Li-Xin Zhang. Asymptotic of the L_{r}-norm of density estimators in the autoregressive time series. Statistics: a Journal of Theoretical and Applied Statistics, 45(2):163–178, 2011. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Yang:2015:ONO

[YH15] Chia-Chen Yang and Leng-Cheng Hwang. Optimal, nearly optimal and robust procedures for the exponential distribution with relative linear exponential loss. Statistics: a Journal of Theoretical and Applied Statistics, 49(3):549–563, 2015.

CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Yin:2020:SVD

[Yin20] Yanqing Yin. On the singular value distribution of largedimensional data matrices whose columns have different correlations. Statistics: a Journal of Theoretical and Applied Statistics, 54(2):353–374, 2020. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Yu:2017:RBP

[YL17] Zhuqing Yu and Stephen M. S. Lee. A recentred bootstrap procedure for constructing uniformly correct confidence sets under smooth function models. Statistics: a Journal of Theoretical and Applied Statistics, 51(2):277–293, 2017. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Yang:2016:NPE

[YLL16] Yiping Yang, Gaorong Li, and Heng Lian. Nonconcave penalized estimation for partially linear models with longitudinal data. Statistics: a Journal of Theoretical and Applied Statistics, 50(1):43–59, 2016. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic). URL http://

www.tandfonline.com/doi/abs/10.1080/02331888.2015.
1074232.

Yu:2022:OCR

[YLWL22] Yuncai Yu, Ling Liu, Peng Wu, and Xinsheng Liu. Optimal convergence rates of Bayesian wavelet estimation with a novel empirical prior in nonparametric regression model. Statistics: a Journal of Theoretical and Applied Statistics, 56(3):565– 577, 2022. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-

4910 (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/02331888.2022.2075363.

Yang:2017:QRR

[YLY17] Jing Yang, Fang Lu, and Hu Yang. Quantile regression for robust estimation and variable selection in partially linear varying-coefficient models. Statistics: a Journal of Theoretical and Applied Statistics, 51(6):1179–1199, 2017. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Yang:2018:SIA

[YLY18] Jing Yang, Fang Lu, and Hu Yang. Statistical inference on asymptotic properties of two estimators for the partially linear single-index models. Statistics: a Journal of Theoretical and Applied Statistics, 52(6):1193–1211, 2018. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Yan:2023:SCS

[YN23] Rongfang Yan and Jiale Niu. Stochastic comparisons of second-order statistics from dependent and heterogeneous modified proportional hazard rate observations. Statistics: a Journal of Theoretical and Applied Statistics, 57(2):328– 353, 2023. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-

4910 (electronic). URL http://www.tandfonline.com/doi/
full/10.1080/02331888.2023.2177999.

Yoo:2016:SDR

[Yoo16] Jae Keun Yoo. Sufficient dimension reduction through informative predictor subspace. Statistics: a Journal of Theoretical and Applied Statistics, 50(5):1086–??, 2016. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Yoo:2018:RDR

[Yoo18] Jae Keun Yoo. Response dimension reduction: model-based approach. Statistics: a Journal of Theoretical and Applied Statistics, 52(2):409–425, 2018. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Yang:2023:IPL

[YS23] Lei Yang and Yongzhao Shao. Integrated partially linear model for multi-centre studies with heterogeneity and batch effect in covariates. Statistics: a Journal of Theoretical and Applied Statistics, 57(5):987–1009, 2023. CO-DEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/02331888.2023.2258429.

Young:2001:IDP

[YT01] D. M. Young and D. W. Turner. Independence distribution preserving joint covariance structures for the multivariate two-group case. Statistics: a Journal of Theoretical and Applied Statistics, 35(4):331–347, 2001. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Yu:2004:SCA

[Yu04] Qiqing Yu. Sufficient condition for admissibility of the Wilcoxon test in the classical two-sample problem. Statistics: a Journal of Theoretical and Applied Statistics, 38(4): 295–305, 2004. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Yang:2012:SRC

[YW12] Hu Yang and Jibo Wu. A stochastic restricted k–d class estimator. Statistics: a Journal of Theoretical and Applied Statistics, 46(6):759–766, 2012. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Yanushkevichius:2007:SCI

[YY07] Romanas Yanushkevichius and Olga Yanushkevichiene. Stability of a characterization by the identical distribution of linear forms. Statistics: a Journal of Theoretical and Applied Statistics, 41(4):345–362, 2007. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Yu:2016:GEL

[YYVH16] Jihnhee Yu, Luge Yang, Albert Vexler, and Alan D. Hutson. A generalized empirical likelihood approach for twogroup comparisons given a U-statistic constraint. Statistics: a Journal of Theoretical and Applied Statistics, 50(2):435–453, 2016. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910

(electronic). URL http://www.tandfonline.com/doi/abs/
10.1080/02331888.2015.1050021.

Zahle:1974:AEW

[Zäh74] Matina Zähle. Asymptotische Eigenschaften der Warteschlangenlänge für Bedienungssysteme mit beschränkter Verweilzeit und periodischer Eingangsintensität. (German) [Asymptotic properties of the queue length for operating systems with limited residence time and periodic input intensity]. Mathematische Operationsforschung und Statistik, 5(1):35–42, 1974. CODEN MTOSBI. ISSN 0047-6277.

Zahle:1984:TSS

[Zäh84] M. Zähle. Thick section stereology for random fibres. Statistics (Berlin, DDR), 15(3):429–435, 1984. CODEN MOSSD5. ISSN 0323-3944.

Zahle:2014:MZO

[Zäh14] Henryk Zähle. Marcinkiewicz–Zygmund and ordinary strong laws for empirical distribution functions and plug-in estimators. Statistics: a Journal of Theoretical and Applied Statistics, 48(5):951–964, 2014. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Zang:2014:GRA

[Zan14] Qing-Pei Zang. A general result in almost sure central limit theory for random fields. Statistics: a Journal of Theoretical and Applied Statistics, 48(5):965–970, 2014. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Zang:2015:GRA

[Zan15] Qing-Pei Zang. A general result of the almost sure central limit theorem of uniform empirical process. Statistics: a Journal of Theoretical and Applied Statistics, 49(1):98–103, 2015. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Zaremba:1978:PTT

[Zar78] S. K. Zaremba. Polynomial trends in time series. Statistics (Berlin, DDR), 9(4):625–642, 1978. CODEN MOSSD5. ISSN 0323-3944.

Zhao:2011:SCR

[ZB11] Peng Zhao and N. Balakrishnan. Some characterization results for parallel systems with two heterogeneous exponential components. Statistics: a Journal of Theoretical and Applied Statistics, 45(6):593–604, 2011. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Zhao:2014:RSO

[ZB14] Peng Zhao and N. Balakrishnan. On the right spread ordering of parallel systems with two heterogeneous components. Statistics: a Journal of Theoretical and Applied Statistics, 48(2):447–455, 2014. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Zhu:2020:ELR

[ZBF^{+}20] Xiaojun Zhu, N. Balakrishnan, Chengzhu Feng, Jiacheng Ni, Nanxi Yu, and Weilin Zhou. Exact likelihood-ratio tests for joint type-II censored exponential data. Statistics: a Journal of Theoretical and Applied Statistics, 54(3):636–648, 2020. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Zhang:2010:TIM

[ZC10] Shuhong Zhang and Haitao Cheng. Testing for increasing mean inactivity time. Statistics: a Journal of Theoretical and Applied Statistics, 44(5):467–476, 2010. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Zezula:1993:CCE

[Zez93] Ivan Zezula. Covariance components estimation in the growth curve model. Statistics: a Journal of Theoretical and Applied Statistics, 24(4):321–330, 1993. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Zamini:2020:APE

[ZFA20] Raheleh Zamini, Vahid Fakoor, and Mahboubeh Akbari. Asymptotic properties of the estimator of cumulative residual entropy. Statistics: a Journal of Theoretical and Applied Statistics, 54(2):281–290, 2020. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Zhang:2016:IBA

[ZH16] Shibin Zhang and Xuming He. Inference based on adaptive grid selection of probability transforms. Statistics: a Journal of Theoretical and Applied Statistics, 50(3):667–688, 2016. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Zhang:1997:OBT

[Zha97] Jian Zhang. The optimal breakdown M-test and score test. Statistics: a Journal of Theoretical and Applied Statistics, 30 (1):47–68, 1997. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Zhang:2012:GAE

[Zha12] Shenghai Zhang. A GEE approach for estimating size of hardto-reach population by using capture-recapture data. Statistics: a Journal of Theoretical and Applied Statistics, 46(2): 175–183, 2012. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Zhang:2013:RBM

[Zha13] Jin Zhang. Reducing bias of the maximum-likelihood estimation for the truncated Pareto distribution. Statistics: a Journal of Theoretical and Applied Statistics, 47(4):792–799, 2013. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Zhang:2016:DRE

[Zha16] Biao Zhang. Doubly robust empirical likelihood inference in covariate-missing data problems. Statistics: a Journal of Theoretical and Applied Statistics, 50(5):1173–??, 2016. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Zhang:2018:MVC

[Zha18] Jin Zhang. Minimum-volume confidence sets for normal linear regression models. Statistics: a Journal of Theoretical and Applied Statistics, 52(4):874–884, 2018. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Zhao:2023:RHD

[Zha23] Ping Zhao. Robust high-dimensional alpha test for conditional time-varying factor models. Statistics: a Journal of Theoretical and Applied Statistics, 57(2):444–457, 2023. CO-DEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/02331888.2023.2180003.

Zheng:2010:NDD

[Zhe10] Qi Zheng. A new discrete distribution induced by the Luria– Delbrück mutation model. Statistics: a Journal of Theoretical and Applied Statistics, 44(5):529–540, 2010. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Zheng:2012:IPM

[ZHG12] S. Zheng, J. M. Hardin, and A. K. Gupta. The inverse problem of multivariate and matrix-variate skew normal distributions. Statistics: a Journal of Theoretical and Applied Statistics, 46(3):361–371, 2012. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Zielinski:1985:MBR

[Zie85] Ryszard Zielinski. A most bias-robust estimate of the location parameter: a general existence theorem. Statistics: a Journal of Theoretical and Applied Statistics, 16(4):637–640, 1985. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Zielinski:1988:DFM

[Zie88a] Ryszard Zielinski. A distribution-free median-unbiased quantile estimator. Statistics: a Journal of Theoretical and Applied Statistics, 19(2):223–227, 1988. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Zielinski:1988:SEL

[Zie88b] Ryszard Zieliński. Stable estimation of location parameter — nonasymptotic approach. Statistics: a Journal of Theoretical and Applied Statistics, 19(2):229–231, 1988. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Zielinski:1989:NHP

[Zie89] Wojciech Zielinski. A note on the Hsu problem. Statistics: a Journal of Theoretical and Applied Statistics, 20(2):195–198, 1989. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Zielinski:1995:OSO

[Zie95] Ryszardo Zieliński. One-sample one-sided student T -test under contaminants. Statistics: a Journal of Theoretical and Applied Statistics, 27(1–2):143–150, 1995. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Zielinski:2000:MUE

[Zie00] Ryszard Zieliński. A median-unbiased estimator of the characteristic exponent of a symmetric stable distribution. Statistics: a Journal of Theoretical and Applied Statistics, 34(4): 353–355, 2000. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Zielinski:2001:PON

[Zie01] Ryszard Zieliński. PMC-optimal nonparametric quantile estimator. Statistics: a Journal of Theoretical and Applied Statistics, 35(4):453–462, 2001. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Zilinskasi:1978:SMM

[Zil78] A. Zilinskasi. On statistical models for multimodal optimization. Statistics (Berlin, DDR), 9(2):255–266, 1978. CODEN MOSSD5. ISSN 0323-3944.

Zinzius:1981:MMN

[Zin81] E. Zinzius. Minimaxschätzer für den Mittelwert θ einer normalverteilten Zufallsgröße mit bekannter Varianz bei vorgegebener Oberer und unterer Schranke für θ. (German) [Minimax estimator for θ on the mean of a normally distributed random variable with known variance for a given upper and lower bound for θ]. Statistics (Berlin, DDR), 12 (4):551–557, 1981. CODEN MOSSD5. ISSN 0323-3944.

Zaigraev:2016:NCI

[ZK16] A. Zaigraev and I. Kaniovska. A note on comparison and improvement of estimators based on likelihood. Statistics: a Journal of Theoretical and Applied Statistics, 50(1):219–232, 2016. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910

(electronic). URL http://www.tandfonline.com/doi/abs/
10.1080/02331888.2015.1032971.

Zou:2003:VER

[ZL03] Guohua Zou and Yong Li. On variance estimation of the ratio estimator under two-phase sampling. Statistics: a Journal of Theoretical and Applied Statistics, 37(1):1–16, 2003. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Zhou:2013:SRE

[ZL13] Xing-Cai Zhou and Jin-Guan Lin. Semiparametric regression estimation for longitudinal data in models with martingale difference error’s structure. Statistics: a Journal of Theoretical and Applied Statistics, 47(3):521–534, 2013. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Zhou:2014:ELI

[ZL14] Xing-Cai Zhou and Jin-Guan Lin. Empirical likelihood inference in mixture of semiparametric varying-coefficient models for longitudinal data with non-ignorable dropout. Statistics: a Journal of Theoretical and Applied Statistics, 48(3):668–684, 2014. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Zhao:2016:VSA

[ZL16] Kaifeng Zhao and Heng Lian. Variable selection in additive quantile regression using nonconcave penalty. Statistics: a Journal of Theoretical and Applied Statistics, 50(6):1276– 1289, 2016. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Zou:2017:WED

[ZL17] Yu-Ye Zou and Han-Ying Liang. Wavelet estimation of density for censored data with censoring indicator missing at random. Statistics: a Journal of Theoretical and Applied Statistics, 51(6):1214–1237, 2017. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Zhao:2008:SCS

[ZLLX08] Peng Zhao, Xiaohu Li, Zhouping Li, and Maochao Xu. Stochastic comparisons of spacings of record values from one or two sample sequences. Statistics: a Journal of Theoretical and Applied Statistics, 42(2):167–177, 2008. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Zhou:1995:LRT

[ZM95] S. Zhou and R. A. Maller. The likelihood ratio test for the presence of immunes in a censored sample. Statistics: a Journal of Theoretical and Applied Statistics, 27(1–2):181–201, 1995. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Zhong:2020:SRU

[ZMY20] Weibin Zhong, Kepher H. Makambi, and Ao Yuan. Semiparametric regression using empirical likelihood with shape information. Statistics: a Journal of Theoretical and Applied Statistics, 54(3):441–468, 2020. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Zhou:2014:NEC

[ZN14] Haojin Zhou and Tapan K. Nayak. A note on existence and construction of invariant loss functions. Statistics: a Journal of Theoretical and Applied Statistics, 48(6):1335–1343, 2014. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Zhao:2016:LRO

[ZQB16] Peng Zhao, Jianfei Qiao, and N. Balakrishnan. Likelihood ratio order of the second spacing in multiple-outlier exponential models. Statistics: a Journal of Theoretical and Applied Statistics, 50(1):206–218, 2016. CO-DEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (elec-

tronic). URL http://www.tandfonline.com/doi/abs/10.
1080/02331888.2015.1016028.

Zou:2023:SAL

[ZQC23] Na Zou, Hong Qin, and Kashinath Chatterjee. A study on average Lee discrepancy measure. Statistics: a Journal of Theoretical and Applied Statistics, 57(1):53–70, 2023. CO-DEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/02331888.2023.2169689.

Zhang:2010:NAT

[ZR10] Haimeng Zhang and M. Bhaskara Rao. On Nelson–Aalen type estimation in the partial Koziol–Green model. Statistics: a Journal of Theoretical and Applied Statistics, 44(5):455–465, 2010. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Zinodiny:2017:NCB

[ZRAN17] S. Zinodiny, S. Rezaei, O. Naghshineh Arjmand, and S. Nadarajah. A new class of Bayes minimax estimators of the normal mean matrix for the case of common unknown variances. Statistics: a Journal of Theoretical and Applied Statistics, 51(5):1082–1094, 2017. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Zwanzig:1990:BR

[ZS90] Silvelyn Zwanzig and B. Seifert. Book reviews. Statistics: a Journal of Theoretical and Applied Statistics, 21(1):56, 1990. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Zhao:2015:MLE

[ZT15] Yuan Ying Zhao and Nian Sheng Tang. Maximum-likelihood estimation and influence analysis in multivariate skew-normal reproductive dispersion mixed models for longitudinal data.

Statistics: a Journal of Theoretical and Applied Statistics, 49 (6):1348–1365, 2015. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Zhao:2017:EIP

[ZTJ17] Pu-Ying Zhao, Nian-Sheng Tang, and De-Peng Jiang. Efficient inverse probability weighting method for quantile regression with nonignorable missing data. Statistics: a Journal of Theoretical and Applied Statistics, 51(2):363–386, 2017. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Zuo:2001:FST

[Zuo01] Yijun Zuo. Finite sample tail behaviour of Hodges–Lehmann type estimators. Statistics: a Journal of Theoretical and Applied Statistics, 35(4):557–568, 2001. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Zvara:1979:ECR

[Zvá79] Karel Zvára. On exact confidence regions for linear regression functions. Statistics (Berlin, DDR), 10(1):55–62, 1979. CODEN MOSSD5. ISSN 0323-3944.

Zhao:2010:APE

[ZW10] Jin Zhao and Jinde Wang. Asymptotic properties of an empirical K-function for inhomogeneous spatial point processes. Statistics: a Journal of Theoretical and Applied Statistics, 44(3):261–267, 2010. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Zaigraev:2020:ODH

[ZW20] A. Zaigraev and M. Wilk. Optimal designs for heteroscedastic regression models with two parameters. Statistics: a Journal of Theoretical and Applied Statistics, 54(2):291–309, 2020. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Zwanzig:1980:CAM

[Zwa80] S. Zwanzig. The choice of approximative models in nonlinear regression. Statistics (Berlin, DDR), 11(1):23–47, 1980. CODEN MOSSD5. ISSN 0323-3944.

Zwanzig:1983:NPI

[Zwa83] S. Zwanzig. A note to the paper of Ivanov and Zwanzig on the asymptotic expansion of the least squares estimator. Statistics (Berlin, DDR), 14(1):29–32, 1983. CODEN MOSSD5. ISSN 0323-3944.

Zwanzig:1985:TOA

[Zwa85] S. Zwanzig. A third order asymptotic comparison of least squares, jack-knifing and cross-validation for error variance estimation in nonlinear regression. Statistics: a Journal of Theoretical and Applied Statistics, 16(1):47–54, 1985. CO-DEN MOSSD5. ISSN 0233-1888 1029-4910 (electronic).

Zou:2022:EHM

[ZWFZ22] Yuye Zou, Chengxin Wu, Guoliang Fan, and Riquan Zhang. Estimation for a hybrid model of functional and linear measurement errors regression with missing response. Statistics: a Journal of Theoretical and Applied Statistics, 56(2):271–296, 2022. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Zhang:2014:EPS

[ZWL14] Shangli Zhang, Lichun Wang, and Heng Lian. Estimation by polynomial splines with variable selection in additive Cox models. Statistics: a Journal of Theoretical and Applied Statistics, 48(1):67–80, 2014. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Zhang:2014:EVS

[ZWYG14] Jun Zhang, Xiaoguang Wang, Yao Yu, and Yujie Gai. Estimation and variable selection in partial linear single index models with error-prone linear covariates. Statistics: a Journal of Theoretical and Applied Statistics, 48(5):1048–1070, 2014. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Zeng:2014:GVS

[ZX14] Lingmin Zeng and Jun Xie. Group variable selection via SCAD-L_{2}. Statistics: a Journal of Theoretical and Applied Statistics, 48(1):49–66, 2014. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Zhao:2017:UMP

[ZX17] Guimei Zhao and Xingzhong Xu. Uniformly most powerful unbiased test in univariate linear calibration. Statistics: a Journal of Theoretical and Applied Statistics, 51(3):609–614, 2017. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Zhao:2011:RTD

[ZXT11] Wenzhi Zhao, Zhiming Xia, and Zheng Tian. Ratio test to detect change in the variance of linear process. Statistics: a Journal of Theoretical and Applied Statistics, 45(2):189–198, 2011. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Zhang:2007:CRT

[ZY07] Chunmei Zhang and Hu Yang. The conditional ridge-type estimation in singular linear model with linear equality restrictions. Statistics: a Journal of Theoretical and Applied Statistics, 41(6):485–494, 2007. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Zhou:2021:ODP

[ZYW21] Xiao-Dong Zhou, Rong-Xian Yue, and Yun-Juan Wang. Optimal designs for the prediction of mixed effects in linear mixed models. Statistics: a Journal of Theoretical and Applied Statistics, 55(3):635–659, 2021. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Zhang:2021:UCD

[ZYZ21] Mei Zhang, Feng Yang, and Yong-Dao Zhou. Uniformity criterion for designs with both qualitative and quantitative factors. Statistics: a Journal of Theoretical and Applied Statistics, 55 (1):90–109, 2021. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Zielinski:2005:BEN

[ZZ05] Ryszard Zieliński and Wojciech Zieliński. Best exact nonparametric confidence intervals for quantiles. Statistics: a Journal of Theoretical and Applied Statistics, 39(1):67–71, 2005. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Zhao:2014:GVC

[ZZL^{+}14] Weihua Zhao, Riquan Zhang, Jicai Liu, Zhensheng Huang, and Yazhao Lv. Generalized varying-coefficient single-index model. Statistics: a Journal of Theoretical and Applied Statistics, 48(6):1311–1323, 2014. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Zhang:2020:FCR

[ZZLL20] Fode Zhang, Weiping Zhang, Rui Li, and Heng Lian. Faster convergence rate for functional linear regression in reproducing kernel Hilbert spaces. Statistics: a Journal of Theoretical and Applied Statistics, 54(1):167–181, 2020. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Zhao:2016:EOS

[ZZQ16] Peng Zhao, Yiying Zhang, and Jianfei Qiao. On extreme order statistics from heterogeneous Weibull variables. Statistics: a Journal of Theoretical and Applied Statistics, 50(6):1376– 1386, 2016. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Zhao:2019:PVC

[ZZW^{+}19] Weihua Zhao, Fode Zhang, Xuejun Wang, Rui Li, and Heng Lian. Principal varying coefficient estimator for highdimensional models. Statistics: a Journal of Theoretical and Applied Statistics, 53(6):1234–1250, 2019. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Zhang:2023:TSC

[ZZW23] Xudong Zhang, Ting Zhang, and Lei Wang. Two-stage communication-efficient distributed sparse M-estimation with missing data. Statistics: a Journal of Theoretical and Applied Statistics, 57(3):617–636, 2023. CO-DEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (elec-

tronic). URL http://www.tandfonline.com/doi/full/10.
1080/02331888.2023.2201505.

Zou:2009:STI

[ZZWG09] Guohua Zou, Jie Zeng, Alan T. K. Wan, and Zhong Guan. Stein-type improved estimation of standard error under asymmetric LINEX loss function. Statistics: a Journal of Theoretical and Applied Statistics, 43(2):121–129, 2009. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).

Zheng:2023:WCW

[ZZWW23] Shunping Zheng, Fei Zhang, Chunhua Wang, and Xuejun Wang. Weak convergence for weighted sums of a class of random variables with related statistical applications. Statistics: a Journal of Theoretical and Applied Statistics, 57(4): 867–899, 2023. CODEN MOSSD5. ISSN 0233-1888 (print),

1029-4910 (electronic). URL http://www.tandfonline.com/
doi/full/10.1080/02331888.2023.2227984.

Zhu:2017:MAM

[ZZZ17] Rong Zhu, Guohua Zou, and Xinyu Zhang. Model averaging for multivariate multiple regression models. Statistics: a Journal of Theoretical and Applied Statistics, 52(1):205–227, 2017. CODEN MOSSD5. ISSN 0233-1888 (print), 1029-4910 (electronic).